A Complete Bibliography of Publications in the SIAM Journal on Discrete Mathematics

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

31 May 2024 Version 3.35

Title word cross-reference

(0, 1) [BM15]. (0.5, n) [BE13]. (1, −1) [dLL09]. (1, 2) [DK06, DK10, BKM08]. (1, ≤ l) [Lai05]. (1, m + 1, n + 1) [BD01]. (1 − ε) ex(n, C_{4}) [BS10a]. (2 + ϵ)n [DZ01]. (2, 2) [GZ19]. (2p + 1) [LHHL18]. (2P_{2}, K_{4}) [GH19]. (2s + 1) [LL14b]. (3, 1)^{∗} [Xu09]. (3a: a) [DH20b]. (4, 3) [GZ19]. (4 − ϵ) [KNNW23]. (7, 2) [Mác13, CW09]. (9v, 4, 1) [FHMY01]. (α, β) [HMSZ24]. (d, k) [KS03a]. (Δ + 2) [WHW14]. (k, l) [GSL98]. (k − 2) [dOBMS^{+}17]. (K_{5} \ e) [GL08]. (≤ 4) [DL14]. (Q, x) [Mal89]. (r + 1) [CtJL01]. (S, T) [BBRT23]. (t, m, s) [AS97]. 0 [AZ22, BG91, BCH92, FK21, HL15]. 0, 1/2 [Fio06]. 0.5 [AHS01]. 1 [AZ22, BG91, BCH92, CHZ04, FK21, GL10, HL15, HOD22, LW17a, Rif99, Riz02, Spi95, ZLS08]. 1 − 1/e [ILM20]. 1.1 [NK90]. 1.5 [CSS01]. 1/2 [CD14]. 1/3–2/3 [BW92]. 2 [ADO23, AB94, AK22b, ACF18, BBF99, BM97, BJT92, BH97, BL23, BIT13, BL17b, BCC^{+}22, BJS21, BCKP19, CL15a, CL16, CL07, CSS13, CSS01, CY18, CDK10, DX19, DGM12, FJ09, FMMO93, Fis94, GVW06, GSK91, GY92, GMR^{+}21, HL00, IKM^{+}92, JM97, KSS11a, LRWZ12, LM17, LC12, Nov18, OSW16, OW22a, OZ18, PSW96, RT18, Sav14, Tak22, TSN04, Vaz12, Voi07, Wan02a]. 2/3 [BT14]. 23 [JZ05]. 24 [EK10]. 25 [KKW17]. 2^{m−2} [Car94]. 2^{n} [AFL^{+}20]. 2d + 1 [PVY22]. 2d − 1 [Lei94]. 2 × 4 [MMJF03]. 3 [ABHM00, APS22, AcRS07, BV22, BM22, Bih05, BS16a, BS16b, BK12, BW02, BH22, CM90, CZ97, CH06a, CS18c, CS02, CW09, CL21, CM14, DST01, DPR22, DD13, DJ11, DNR23, DM15, DGM12, DL18b, EHJ01, EGM18, FRMPV15, FRZ16, FZ08, FXYY14, FT12, FZ22, Gab04, GKRS15, GSK91, GLM22, GLP23, GK04, GKL99, HM12a, JKSW17, KRS11, Kan08, KW92, KMMS21, LSS17, LXZZ08, MM12, MRST16, NNO19, OC19, OSW16, PW18, Ran02, TSN04, Vod21, WYZZ14, ZKNS20]. 3/2 [BZ11, RV99]. 3/5 [HK16b]. 31 [KKW17]. 3^{+} [DL12]. 3^{d} [DMN12]. 4 [AHH^{+}10, ALZ96, BCG^{+}10, BH97, CHZ09, Cho94a, CR19, CY03, CLY05a, CLY05b, DX19, DL14, DL17, DL18c, DL18b, Far09, GH19, GZ06, GK04, HWZZ18, HWZZ20, JKSW17, JN16, KiKK17, Kap14a, KP09, LHLZ23, LQ22, LZ18b, LWY22, MR12, OW22b, OZ18, RC98, WX13]. 4, 5 [CCO^{+}13]. 4/3 [CLS15]. 48 [ACM11]. 4n [CF17]. 5 [BW92, CL15a, DX19, DLS11, HLS22, JKSW17, KRS11, iKO16, LHLZ23, Luk20]. 5/6 [CR17b]. 5/8 [LS03a]. 6 [Bal08, GGM^{+}22, iKKKL09, YWWL21]. 6/5 [BL17b]. 7 [Far09, JKSW17, KRS11]. 8 [KSS11b, Ran02]. 84 [FL00]. 9 [Che22, DŠV08, KSS11b]. [32, 17, 8] [CS89]. [4, 4, 4] [FP22]. ^{2} [AZ22]. A [BHJ18, CGG17, Yam16]. A^{ω} [LV89]. A_{n} [FLM12]. AG(3, 2) [KMPR14]. AG(m, q) [BJvV92]. AG(n, q) [TW21]. α [BCCZ11, Lic98, LT01]. B [CGG^{+}16, BBCZ11, GS03, LRT08]. B_{d} [BFK^{+}12]. B_{k} [SS19]. BCH(2^{m}, 8), m [CHZ09]. mod(2p + 1) [Lai08]. modp [BCPP09]. c [KMR^{+}23, KKS22]. C = 6 [BCC^{+}05]. C_{4} [BS10a, CD14, HjK24]. C_{ℓ} [Pic14]. CAT(0) [ABY14]. ·, · [MTV08]. χ [CKPS13, DI22, RW19]. χ = Δ [CR15]. cx^{q+1} + dx^{q} − ax − b [BGM94]. D [KV15, Saf21, YZ17, Bar01, BK21, BK24, Bra05, CC07, KNP05, PVY22, SV08]. D_{2n} × C_{2} [MLMR23]. D_{n} [AR04]. Δ [CKPS13, Cra19, PSWX24]. Δ + 1 [Bon15]. Δ − 1 [CR13, CR17a]. Δ ≥ 9 [WHW14]. Δ ≥ 8 [Bon15]. e [CH19]. E_{7}^{∗} [Wor88]. ℓ [BCPP09, EEFH21, EHV18, GP08b, GKMS04, HW23, TQ09, LM14]. ℓ_{1} [HMM09, CGN^{+}06b]. ℓ_{p} [BGN15]. F [FGPS19, KOT16]. F (3, 3) [GH13]. F_{2^{m}} [CCD00]. F_{3} [BÖ05]. F_{4} [BÖ05]. \frac{10}{7}OP T [HV17]. \frac{3}{4} [BHT16, GW94]. \frac{4}{3} [BCC^{+}22]. G [DHJN02]. G_{k−out} [BFa24]. Γ [Spi95, LN17]. GF (2) [Web08]. GF(p) [Gor93]. GF(q^{2}) [Mou94]. H [ASS17, CKL^{+}21, Cha91, CFG^{+}21, CGSZ20, CKP^{+}21, FS09a, GKY06, GL17, KL19b, KMN23, Let19, MRV17, PPR22, Pfe15, Sig10, Ath14, CH10, Got03, Més16a, Swa05]. H^{∗} [BB23a]. I_{4} [NN22]. k [LWY18, BCdMR08, KS00, AYZ04, AK02, Ana18, AS02, AB07, BNRZ23, BSS14, BKL^{+}15, BBM90, BLR16, BLR17, CDHH14, CL07, CGN^{+}06b, CGN06a, CQX20, CDK^{+}18, Cho94b, COPP12, CKP16, DH90, DHJN02, DŠ09, EIiK08, EG03, EFK05, FKS03, Fed01, FHL^{+}13a, FHL^{+}14, FFV11, FPS13, FHL20, Gab05a, Gab05b, GP18, GH90a, GHHY96, Got03, GKW19, GSS15, Han16a, HVW07, Har18, HI16, HS06, Hur94, HC98, KKKS17, KSY18, KY12, KS19, KPR10, Lev09, Lin97, LWY18, LLZ22, MPŠ08, McL10, MS24, MW22, NY21, Osh21, PSTF00, PP07b, PRS98, RRS07, SS11a, Sav90, SST08, Ste10, Swa05, TRV03, TP97, Tod89, VWY09, YQZ09, Yus03, ZWW23]. K_{r+1}^{r} [LZ09]. K_{(s,t)} [CCH21]. K_{1,2p+1} [Lai08]. K_{2} [Zam21]. K_{2,4} [EMOT16]. K_{3} + e [SWR12]. K_{4} [BFH21, FGRZ21, FS09b, Gyá19, JW18, Sch02b, ZKNS20, GS03]. K_{4} − e [SWR12]. K_{5}^{−} [LZ18b]. K_{5}\e [FT12]. K_{6} [iKMN09, KM14]. K_{Δ} [EK13, HP23]. K_{k+1}^{k} [BBBZ12]. K_{m,n} [dKMP^{+}06]. K_{n} [ÁAFM^{+}18, dKMP^{+}06]. K_{n,n} [LSS17]. K_{n,n} − I [MPS06]. K_{r+1} [GCH22]. K_{t,t} [Mak07]. k-(S, T) [CL13]. L [Ber20, CK11].

L(2, 1) [CK96, FR06, GM05, JSRSW18, Kan08, KST06]. L(p, 1) [HRS12]. L^{p} [Kra18b]. L_{p} [BK17]. λ [LZ05a, RCS88, SS11a, WGM95]. ⌊\frac{n−1}{2}⌋ [FLM12]. ⌊n/k⌋ − k + 1 [NY21]. M [KMT07, Mur06, Kap14b, MRV17, Smi01]. M(K_{5}\e) [MR12]. M_{963}^{9} [BS95b]. E^{3} [Pry23]. F_{!2}^{8} [TV03]. F_{2}^{n} [Mra17, WZZ18]. F_{q^{n}} [Shp10]. R^{n} [Qui10]. Z^{2} [RR18]. Z^{d} [FGLP14, MP13, GKRS15]. Z_{2}^{n} [RRW22]. Z_{3} [DM23]. Z_{6}^{n} [PP22]. Z_{m} [QD14]. Z_{n}^{k} [DV04]. Z_{p} [MR04a, BM15]. Z_{p} × Z_{p^{n}} [Zha23]. Z_{r} ≀ S_{\} [Ath14]. H [KT16]. M [CK08a]. V [KT16]. HJ(4, 2) [Lav16]. PGL_{2}(q) [Shp10]. RLL(d, k) [Lou10]. N [SS91, AS02, DKS16, DH90, DHJN02, EFF91, Hed08, Hur94, JLR20, KP09, LLZ22, OS15, Pin08, Rif99, Smi01, Tör93, dH89]. n + 1 [Tör93]. n + t [DKS16]. n − 1 [dH89]. n − log(n + 1) − 3 [Hed08]. n^{c log n} [BHRZ14]. N^{N} [BJ92]. N_{2} [WLD09]. n ≥ 12 [FLM12]. n ≤ 2k + 1 [DHJN02]. o [ÇMP15]. o(log n) [AKKS89]. O(m log n) [EM99]. O(n) [MM96]. O(n^{α} log n) = o(n^{2.376}) [HTV05]. O(n log n) [KMS98, Spr94, WY10]. O(\sqrt{n} log^{3/4} n) [CS91]. ω [CKPS13]. Ω\sqrt{log n} [RSSW88]. OPT^{O(∞)} [GLSS16]. P [Vav89, CGG^{+}16, BBC11, GP20, Jac92, JO95, LÖ05, MR04a]. p^{2} [LÖ05]. P_{3} [CPRdS13, HR05a]. P_{5} [BM16, CNST24, KPT95, MM12, BF17]. P_{6} [BM16]. P_{8} [CS18c]. P G_{m}(q) [ZG21]. PGL_{3}(q) [MS14]. q [AS97, ACS97, CX08, CEFY24, Etz96a, GSS14, Sca05, WIS12, YZ17]. q, t [LLL18]. Q_{n} [WGM95]. r [BS15b, BCD97, FYK00, GCH22, GRS12, LZ18a, LLZ22, MMSW23, Sid18, Sta11]. R(3, n) [Exo89]. r = 3, 4 [ZGL^{+}09]. R^{d} [LM22b]. R^{n} [BT96]. r × 4 [ZGL^{+}09]. S [Ave13, CD16, MD11, RY91, CGG^{+}16, DH90, Gao15, GKNU10, NS11]. S_{n} [RS93]. S_{Σ} [Abe91]. st [AB00]. T [Bar04, GGW06, GS98, Ray94, WY20, BS10b, BS16c, BF17, CvBdJdVK22, CLWZ22, DŚN21, DM13, Gao15, GNPV20, HS89b, LS03b, Mak07, NS11, RCS88, Spi23, Suk13, Tak22]. {U_{2,5}, U_{3,5}} [CMvZW16]. v [RCS88]. ε [MR15b]. W_{4} [BM94c]. x^{2^{m}} + x [Car94]. Z_{2} [TRV03]. Z_{4} [Ran02]. Z_{m} [LLL17, Tsa96]. Z_{n} [Fil89].

-Analogue [WIS12]. -Approximation [AHS01, BBF99, BZ11, BCC^{+}22, CSS01, RV99, GW94]. -Arboricities [HM12a]. -Arboricity [BCCZ11]. -Arrangements [FHL20]. -Ary [KP09, AS02, Etz96a, Lic98, LÖ05, LT01, PRS98, Sca05]. -Asymptotics [BHRZ14]. -Atom [DGM12]. -Balanced [LLZ22]. -Bernstein [GSS14]. -Biased [MR15b]. -binomial [Mal89]. -Blocks [CDHH14]. -Bounded [RW19]. -Catalan [LLL18]. -Center [KS00]. -Chain [GP08b]. -Choosable [DLS11, Far09, YWWL21]. -Chromatic [DŠ09, GKL99]. -Circuits [CL15a, Luk20]. -Circularity [Saf21]. -Class [AK22b, MS24]. -Claw [MRV17]. -clique [GHHY96]. -Cliques [NY21]. -Closure [KMR^{+}23, KKS22]. -CM [Swa05]. -CNF [HS06, FFV11]. -Colorability [DL18b, AK02]. -Colorable [CS18c, CS02, FT12, GH19, GK04, JKSW17, KRS11, LM17, LS03b, MM12, NNO19, CKP^{+}21, WHW14]. -Colored [DGM12, KW92]. -Coloring [DPR22, GK04, HLS22, KSS11b, KNNW23, Xu09, GKRS15, Sig10, Ray94]. -Colorings [FXYY14, FR06]. -competition [IKM^{+}92]. -complete [Vav89]. -components [JO95]. -Configurations [FYK00]. -Conjecture [BHT16, HK16b]. -Connected [ABHM00, ADO23, BJS21, Che22, CL21, CY03, CLY05a, CLY05b, DL12, DD13, DJ11, EIiK08, EEFH21, EG03, EHJ01, FJ09, GZ06, KKKS17, iKO16, LRWZ12, LC12, LQ22, LZ18b, MR12, OSW16, OW22b, OZ18, ZWW23, BBM90, Voi07, Ste10].

-Connectivity [GGW06, LLL17]. -Constrained [KS03a]. -Convex [Ave13, KMT07, Mur06]. -Convexity [CPRdS13]. -Copies [Let19]. -Core [BKL^{+}15, Wan02a]. -Counting [CEFY24]. -Critical [DL14, DL17, DL18c, DL18b, Har18, iKKKL09]. -Crosses [BE13]. -cube [Rif99]. -Cubes [AS02]. -Cut [CGN06a, CQX20, GH90a]. -Cuts [Fio06, BBRT23]. -Cycle [GGM^{+}22, LHLZ23]. -Cycles [DX19, DL14, DL17, Far09, JKSW17, WX13]. -Decompositions [Lai08]. -Degenerate [KNP05]. -Degree [dOBMS^{+}17, LM14, FGPS19]. -Derangement [CX08]. -Designs [CM90, GL08, Ran02, RCS88]. -Differences [Sav90]. -Dimensional [Bra05, CC07, GMR^{+}21]. -Disjoint [BNRZ23]. -Distance [DX19, NS11]. -Distant [CtJL01]. -Domatic [SV08]. -domination [BCD97]. -Edge [BL17b, BCC^{+}22, CSS01, Gab04, Gab05b, BM97, Jac92]. -Edge-Choosable [CW09, Mác13, Bon15]. -Edge-Colored [KMMS21, LSS17, McL10]. -Edge-Connected [KiKK17, MMSW23, SS11a, Cho94b]. -Element [BM22]. -Equal [Got03]. -error-correcting [CZ97]. -Eulerian [YZ17]. -Factor [CL16, CSS13]. -Factors [BIT13, CL15a, CDK10, Riz02, Tak22, ZLS08, BH97, CY18]. -Fixed-Endpoint [LW17a]. -Flow-Critical [DM23]. -Flows [FZ08, LHLZ23, LXZZ08, WYZZ14, ALZ96]. -Fold [CH10, JLR20, OS15]. -Fragile [CMvZW16]. -Frame [GP20]. -Free [Ave13, BS10a, BFK^{+}12, BM16, BF17, CKL^{+}21, CS18c, CGSZ20, CNST24, CD16, CD14, EK13, GH19, HP23, HjK24, KOT16, MTV08, MM12, Mak07, MD11, NN22, PPR22, Sch02b, ASS17, Pic14, Spi95]. -Games [KMN23]. -Goethals [Ran02]. -Gons [AHH^{+}10, EFF91]. -Graph [AcRS07]. -Graphic [FS09a]. -Graphs [CL07, CM14, FRMPV15, GCH22, LWY18, OC19, Sid18, CFG^{+}21]. -Hamiltonian [Zam21]. -Heavy [GNPV20]. -Hyperbolic [CD14]. -Identifiable [TSN04]. -Identifying [Lai05, Sta11]. -Independence [HC98]. -Infinity [Ber20]. -Intersecting [CLWZ22, CL07, Spi23, KL19b]. -Inversion [CGG^{+}16]. -Joins [Bar04, WY20]. -Kernels [GSL98]. -Kings [SS91]. -Labeling [JSRSW18, Kan08, CK96]. -Labelings [GM05, KST06, HRS12]. -Level [AYZ04, ACF18]. -Line [Gab05a]. -Linear [TRV03]. -Linked [KY12, GKY06, Pfe15]. -List-Colorability [DH20b]. -Matchings [BBCZ11, Har18, LRT08, Mak07, Tak22]. -Matchoid [HW23]. -Matrices [dLL09]. -Matrix [BM15]. -Maximal [Kra18b]. -Median [BBC11]. -Minor [KMPR14, MR12]. -Minor-Free [EMOT16, FGRZ21, JW18]. -Minors [iKMN09, KM14]. -Möbius [JM97]. -Modular [EHV18, OW22a, PSWX24]. -Modules [HMSZ24]. -Monotone [Suk13]. -Monotonicity [GKW19]. -Neighbor [BK21, BK24, DNR23, KS19]. -Neighborhood-Covering [HC98]. -Neighborly [Nov18]. -nets [AS97]. -Noncrossing [MW22]. -NU [FHL^{+}13a, FHL^{+}14]. -Number [LZ05a, WGM95]. -Numberings [AB00]. -of-each [Lin97]. -Orientation [LHHL18]. -Orientations [Lai08, LL14b, MRST16]. -Out [Ana18]. -out-of- [DH90, DHJN02]. -Outerplanar [CGN^{+}06b]. -Packing [Fis94, BCD97]. -Parameter [Vod21]. -Partite [Fed01, GSS15, LWY18, LLZ22, RRS07]. -Partitions [CGG^{+}16]. -Path [BJT92]. -Paths [BHJ18, Yam16]. -Payment [PSTF00]. -Perfect [BS16c, Rif99, GS98]. -Perfection [BS10b, BF17]. -Planar [HOD22]. -Player [Vaz12]. -Plex [GKNU10]. -Point [CF17, Kap14a]. -Polymatroids [BL23, OSW16, Sav14]. -Polynomials [Ath14, ÇMP15, Més16a, BB23a]. -Polytopes [BS16a, BS16b, BW02, DST01, PVY22, KT16]. -Positivity [CH19]. -Preservers [KV15]. -Quasi-planar [FPS13]. -Radius [BLR16, BLR17]. -Ramsey [DI22]. -Rate [CHZ04]. -Regular [CW09, GRS12, MMSW23, McL10, BH97]. -Representation [MRV17]. -Resonance [KSS11a]. -Resonant [YQZ09]. -Robust [KPR10]. -SAT [COPP12]. -Saturated [CCH21]. -Saturating [GL10]. -Schur [BSS14]. -Separators [LQ22]. -Sequences [BCPP09, GKMS04, TQ09, CK11]. -series [ACS97, RY91]. -set [Hur94, Lev09]. -Sets [PP07b, SS19]. -Shape [FL00]. -Space [BK12]. -Sperner [EFK05]. -Sphere [BCKP19]. -Split [HL00]. -Stable [GS03]. -stage [CH06a, DH90]. -State [BS15b]. -Strong [DŚN21]. -Structure [HR05a]. -Subdivision [LZ18b]. -Subdivisions [BFH21]. -Subgraph [SS11a]. -Subgraphs [AB07]. -Subhypergraph [CDK^{+}18]. -Submodular [HI16, Osh21]. -subsets [Hur94]. -Survivable [BKM08]. -Term [SST08]. -Terminal [Bih05]. -Theorem [DMN12]. -thin [BW92]. -Tori [PW18]. -Transversal [GL17]. -Transversal/Packing [GL17]. -tree [Tod89]. -Trees [DM15, GSK91, PSW96, VWY09, TP97]. -tuple [Kap14b]. -Uniform [APS22, BV22, BH22, EGM18, FRZ16, FZ22, GLM22, GLP23, GSS15, Han16a, HWZZ18, HWZZ20, LZ18a, LWY22, LLZ22, RRS07, Yus03, ZKNS20]. -Unimodality [CH19]. -Vectors [LN17, Swa05, Cha91]. -Visibility [HVW07]. -Wheel [Cho94a]. -Wise [GNPV20].

01-Fillings [CWYZ10].

1 [FHL^{+}13a]. 1-Hyperbolic [BC03]. 1-Planar [Suz10]. 1.3-Approximation [DL18a]. 18-Reconstructible [PRS03].

2 [FHL^{+}14, Kra07]. 2-Approximation [BMV22]. 2-Partite [Fue23]. 2-SAT [Kra07].

3 [MRV17]. 3-Arc-Dominated [Tan10]. 3-Choosability [DLŠ10]. 3-Facial [HSŠ08]. 3-Hypergraphs [Raz10]. 3-Uniform [Kha13]. 3/2 [BMV22]. 3/2-Approximation [BMV22]. 3PC [MTV08].

4 [GM93]. 4-Cycles [DLŠ10]. 4-Vertex [Raz10].

5-Coloring [EHLP11]. 5-Cycles [ZLWC12].

91m [GM93].

ABAB [AKP20]. ABAB-Free [AKP20]. Abacus [Loe10]. Abelian [BS95a, BL16, BFGM15, Cai18, CF09, DF94, DR04, DM13, Din06, DF04, FFie98, GF08, GGI07, HZ21, KCL98, Sza06, Zho05]. Absolute [BN01]. Abstract [Koc98, Ngu10]. Acceptance [CZOS98, SW10]. Access [Bon91, DM18a, SWKP10]. Accessible [CGMV19]. Achieves [BN01]. Achlioptas [BBFP09, KS12b]. Achromatic [KK01, KS07]. Acknowledgment [AB05]. Acquaintance [BST14, DP16]. Acquisition [LPW^{+}13]. Acting [MS14]. Actions [QP15]. Acute [Har11, Nov18]. Acyclic [BB16, BCC^{+}11, BMP13, CHM^{+}23, CMM20, Cra19, DS16, FHM03, HH04, Han19a, HN15, iKM10, KPPW15, Sli10, Bal88, BM94a, BFM94]. Acyclicity [BK21, BK24]. Ad [KP04]. Adapted [KZ08]. Adapting [CLMS22]. Adaptive [MMPS10]. Adding [Fri21]. Addition [AB94, Alo13, GS93b]. Additive [BCE05, CY12, EK05, EdJdVLT18, KK10, KLM^{+}03, SS21, Shp13]. Addressing [WIS12]. adjacencies [Sar97]. Adjacency [AKTZ19]. Adjacent [BGLS07, JL20a, WH15]. Admit [AFT12, AHH^{+}10, BJS22]. Admitting [FHL^{+}13a, FHL^{+}14, HSS19]. ADMs [CFG^{+}09, MLS11]. adsorption [Pip89]. Advance [PV10b]. Advancing [CKdAHdF13]. Advantageous [AC90]. Adversarial [AEFT13]. Affine [ACD^{+}13, Ber07, CKPS15, CGHL23, Stu88]. AFPTAS [JK19]. after [ABY11]. against [HMS21, OKS06]. Agents [CDV10, Che16]. Aggregation [BGM22]. Agreement [BHL^{+}15, BLO^{+}22, MS19, vIKLS14]. Ahead [CF10]. Aid [DM11]. Aided [SW99]. Alcuin [CHW10]. Alexander [Rho15]. Algebra [HM12b, JLD^{+}18, KLR22]. Algebraic [BHTK^{+}21, BB03, FG01, GR17, JLD^{+}18, MPSZ19, MW19b, MdCW16, OPVV14, SB10, WWKY11]. Algebras [ADL^{+}09, BHM16, Fer16, TW12]. Algorithm [AHS01, AKP^{+}22, ABGJ15, ÁS09c, BBF99, BHRZ14, BJT92, BHH96, BD20b, BFPP15, BZ11, Bon91, BKS10, Boy96, BCC^{+}22, BCHP08, BCP08, CLMS22, CM05b, Che94, CCDZ23, CMSV17, Dan09, DKM^{+}12, Dji06, DHJ^{+}13, FMV22, FKL^{+}19, Fra95, FL91, FKW10, FKV21, GK07, GW00, GS93a, GH90a, GH90b, GL95, Han98, HM20, Har10, Hli18, HV00, HMM^{+}21, IMS05, JLR20, KPT22, iKK11, KN22, KK90, Köp07, KN95, LS17, LS03a, LW17a, LSL92, Luc98, MPSV21, MC12, MNN18, Moh99, Myu01, New20, OS92, Osh21, PP07a, PV10b, PS22, RC18, RV99, Saw02, Sei01, Śni14, Sot15, Way01, AHS89, LW88a, LM89, Spr94, UWZ97]. Algorithmic [BL04, BK17, BDK^{+}24, CFT93, FKM^{+}16, GKS22, Peg14]. Algorithms [AS09a, AYZ04, AMW00, AD96, ASZ02, AB95, ADKS18, BCSK07, BBC11, BC09a, BDM02, BKS09, BS18, CKNZ05, Che04, CV07, CL13, DDS16, DMNW13, Duc21, Elb09, FJLS03, GDVL17, GS13, GK13b, GMPZ15, GSK91, GS16b, GRS12, GSZ22, HH05, HRS17, HLP^{+}23, HKK^{+}09, HW23, HKST03, JLL16, KMR^{+}23, KN16, KFHR94, KMW24, KPS20, Kim91, KMRR09, KKS22, KY12, KP95, KS06, LS95b, LM16, LPSR12, MT05, MTV08, MV99, MT90, MMPS10, MC93, MSZ10, MT11, NT24, Nao91, PRS18, PRS98, RS10, RS11, SL96, SX13, ST07, SBD^{+}19, Som14, SV11, TP97, TT93, Vaz12, ZM02, dBSS24, vIKLS14, GW94, GHHY96, GMW96, NS89, SS89, Mur96b]. Aligned [Tót08]. All-Pairs [HM94]. Alldifferent [MM11]. Allocating [OPS21]. Allocation [FK98, LNO96, Lyn94, Red17, SX13]. Allocations [ELvS11, LSS24, MS20]. Allowed [AL17]. Almost [BS10a, DR24, DP92, FS09b, GLP^{+}12, GP18, KKSvL20, KK90, LSS17, LWY18, Mon15, PS22, PR20, WZZ09, Zho88]. Almost-Stars [DR24]. Alon [FT05b, Gly10, Sto12]. Alphabetic [KM95]. Alphabets [RSW05]. Alternating [CP10b, DEH20, FLM12, HS22, IKK^{+}22, ZZW13]. Alternative [MS05]. Altshuler [BS95b]. Always [BS10b, KKSvL20]. Amalgam [NS93]. among [CTW93, KS03a]. Amortizing [KOR03]. Ample [CKP22]. Anagram [WW18]. Anagram-Free [WW18]. Analogue [Bra05, KLP12, WIS12]. Analogues [HHLM20]. Analysis [AMW00, AHS89, AT16, BHRZ14, BD20b, BJK13, Bru90, DDS16, FS05, Fis14, FO90, FL91, FKW10, FP13, Gab05b, GF08, HKW15, KRS15, New20, PSTF00, Pip95, PRS98, Sca03, Sot15, Tak08, Yam20, Yam05, Bie88]. Analytic [IK09, MW19a]. Analyzing [KPS20]. Anchored [BKL^{+}15]. Angle [Eti20, LM22b]. Angle-Restricted [Eti20]. Angles [AFT12, AS05, BPV10, DV96, SW04]. Angular [FKK21, MP94]. Annulus [GST23]. Anomaly [Tov90]. Answers [NT05]. Anti [GLS20, GLP23, LW20, LMW23]. Anti-Ramsey [GLS20, GLP23, LW20, LMW23]. Antichain [BPST22]. Antichains [Fra90, DSW90]. Anticoncentration [RY22]. Antiferromagnet [COLMS22]. Antiholes [BBT16]. Antipodality [MAR22]. Antisubstar [CNRS15]. Antiweb [CdV02]. Antiweb-Wheel [CdV02]. Any [AM21, CKP16, DF94]. APN [BBMM09, TQLT13]. Apple [BLM10]. Apple-Free [BLM10]. Application [BBJ^{+}21, Ber20, BP12, CuKŠ07, FL92, GPP04, GL08, KSV05, KŠ08b, KL08, LS17, HS89b]. Applications [AOW15, BYHR10, Bar95, BCdMR08, BK21, BK24, CR16, CF17, CW16, DJKO19, DKLW24, Doh03, ENSZ00, Fra90, Gal18, GKS22, GKLT19, GKS12, GS93b, GMTW15, HY15, Jan00, iKX20, KMRR09, Lu04, MK01, MPP17, MMJF03, ST20b, SS02b, BC88b, SSS95, Stu88]. Applied [HHHH02]. Approach [ACD08, BNMN92, BCE^{+}01, BAM16, BBKS23, CS98, CL91b, CP16b, DMS14, Gab04, IK09, Kim92, LLL18, Roz19, SSS13, SZ94, SV88]. Approximability [Che09, DW24, HKMY22, KvIL^{+}12]. Approximate [dOBMS^{+}17, BLM20, Cal13, CCSW23, CS09, CDN16, DHL^{+}02, FK98, Goe01, HT13, HKP^{+}17a, HKP^{+}17b, HKP^{+}17c, HKP^{+}17d, KV15, LOW10, MR15a, OS13b, Sim90, AA88]. Approximating [ENSZ00, Fei04, FMV22, Fle00, FGT11, Fuk16, Gab04, Gab05b, GGHP22, KK01, KLN10, KS02, Lev09]. Approximation [AS09a, ACLW18, AHS01, AYZ04, ASZ02, AB95, AS14, ÁS09c, BBF99, BYFMR10, BMV22, BZ11, BDGP22, BCC^{+}22, BS18, CFKK17, CCG^{+}11, CH23, CKNZ05, Che04, CSS01, CV07, CL13, CCDZ23, CMSV17, CNP21, Duc21, DL18a, DM15, DFK^{+}21, ELMS11, FLM^{+}16, GLSS16, GPSS01, GHHY96, HKL^{+}14, HLP^{+}23, HMM^{+}21, JSOS03, JP06a, JP18, KP95, KS07, KS06, KN95, LS03a, Pal12, PP07a, RV99, RS10, RZ05, SS11a, SX13, SBD^{+}19, SY11, Svi03, ZM02, dBSS24, vIKLS14, GW94]. Approximations [KS03d]. Arbitrary [BBCZ11, Dan09, DKM^{+}15, DR04, DS05b, Moh99, Sol12, WWKY11]. Arborescence [Kir16]. Arborescences [HPS19, dGNS13]. Arboricities [HM12a]. Arboricity [BCCZ11, ERS19, Wu09, HS89a]. Arc [BHH96, CPS08, FHS14, FIN98, GNS11, LS08, Mac91, NY21, Ram90, Rea15, Saf21, SSR94, Tan10, Yam05, Zho09]. Arc-Disjoint [FIN98]. Arc-Transitive [Mac91, Zho09]. Arc-Weighted [GNS11]. Arcs [BJHY03]. Area [AS16]. Areas [Pin08]. Argument [WL10]. Arguments [FP99, KLMR03, RM97]. Arising [ACF18, GM90, Pip95, ACS97]. Arithmetic [GP14, Het14, JS14, KS23a, PP22]. Arm [ABCG17]. Armstrong [SZ15]. Arrangeable [BKTW15]. Arrangements [ASH18, ARS21, ERSZ22, FHL20, RC18, Sol06]. Array [Lai18]. Arrays [ACLT01, BGS96, CCHZ13, GB12, MMPS10, MHLHL91, SC17, Zhu21, Jed93, JLM93]. Arrovian [Das99, FK97]. Ary [KP09, AS02, Etz96a, Lic98, LÖ05, LT01, PRS98, Sca05]. Ask [ABFRV24]. Aspects [CFT93, MW03, NT05]. assembly [DH90]. Assignment [FYK00, Gol06, JWF05, KS03c, Krá06, dBSS24, Bal89, Ray94]. Assignments [AT90, FFV11, GJ08]. Assisted [Kar20]. Assmus [BR19]. Associahedra [Rho15]. Associahedron [CM18]. Associated [CFM94, RY23, PC97]. Associations [Rab06]. Assortment [Ete22]. Asteroidal [BKKM99, COS97]. Asymmetric [BCGR24, CR96, CDK10, Fis14, LS03a, Bal89]. Asymptotic [AJM08, Bac09, BEL09, Bie88, CH01, DFK^{+}11, GM13, KK09, MS21c, MMS17, McL10, NÖ08, OPR12, Pik03, WS12, Wil05, dJMS16]. Asymptotically [GH90a]. Asymptotics [BHRZ14, KLMR03]. AT-Free [CKOS06]. Atom [DGM12]. Atomicity [MR21]. Attachment [Ja10]. Attractors [JA16]. Augmentation [BJJ98, BJGJS99, CCDZ23, FGT18, HK05, KíP^{+}24, KM22]. Augmenting [ABS10, Bil03, Fra92, Vég11]. Authentication [GMW05, KCTR13, MWW94b, ST14]. Autocorrelation [TQ09]. Autocorrelations [XQ06]. Automata [BBS00, GM91, GJ16, MM93, MRR20, Rom06, GM93]. Automatic [AH21, BDDP24]. Automaton [Jun12]. Automorphism [RD11]. Automorphisms [Kas05, Vse05, Ram97b]. Average [Che94, CCG^{+}00, CF09, DMS08b, DMS08a, Gol06, GRSZ23, GP16, KK89, Wan23, XGG15, AKKS89, Pit89]. Averages [Pol19]. Averse [BS18]. Avis [RC18]. Avoidance [Ehr16]. Avoider [BM09]. Avoider-Enforcer [BM09]. Avoiding [AM06a, CL15a, DGN^{+}20, LR07, MMSJ08, PP22, RR18, Yus09, ZSW11]. Axiomatization [Che24]. Axis [AC14, Kim91, Tót08]. Axis-Aligned [Tót08]. Axis-Parallel [AC14].

Backbone [MŠT09]. Backup [BKM15, Fuk16, HI23]. Bad [GS03, Sch02b]. Bad- [Sch02b]. Balanced [ACLT01, AM11, ABL^{+}20, BLO^{+}22, BG12, EMT15, FL10, LSSZ19, LSS24, LLZ22, Orl93, Wan23, Zhe16, BM94a]. Balancing [AE04, CLGH11, DM18a, SE14, HMP97]. Ball [GJ20, vBJU20]. Ballot [CFJ11]. Balls [BYKKR18, CCM^{+}23, DTW03, LS14, Lut04]. Balog [BC11]. Band [LR04, JM97]. Band- [LR04]. Bandits [KS04]. Bandwidth [AJM08, EG03, JMSW99, JMSW00, KV90, KS02, Mot19, Spr94]. Banner [KN13]. Banzhaf [Bol90]. Bar [BSiT15, CWY21, CHJ^{+}04, HVW07, SiT15]. Bar-Joint [SiT15]. Barbells [LLS^{+}20]. Barely [FGS19]. Bargaining [CGV^{+}14, Vaz12]. Barker [Jed93, JLM93]. Barnette [Kar20]. Barrier [BNRT17, ILM20]. Barriers [DJT15]. Barvinok [Köp07]. Barycentric [Mur10]. Base [AK22a, BDT17, AS97]. Baseball [Way01]. Based [ACKM19, ASZ02, DKSŻ10, EMT15, ES98, EH13, GMPZ15, GKNU10, Hay21, Ja10, Köp07, KCTR13, NLS24, PdCSC21, Sot15, Wag18, WDSH21, dGNS13, vBBC^{+}15]. Bases [AK14, BP09, BP16, Cho09, CP16b, EdJdVLT18, GSS14, GS93b, HKL11, HKW22, HKK24, JS17, NDB07, RY23, Web08, Win16, BGM94, LW88b]. Basic [ADHL13, ADH^{+}14, PRS02]. Basis [BS24b, BS90, Cho09, CS21, GH06, GW07, HM94, KAN90, Kot13, KRZ21, Lub90a]. Basis-Exchange [Kot13]. Batch [DS05b, GLY07]. Baxter [BGRR18]. Bayesian [XY15]. Bazin [KY21]. BCH [CZ97, Li17, NLJM21]. BCH-codes [CZ97]. Be [BS16c, Che16, MPŠ08, PW02, BBS17, Gao18, Kar89]. beat [Tör93]. Beating [GP16]. Beatty [CDR16]. Behavior [BK11, CCT05, CP20a, CCG^{+}00, MW19a]. Belief [BBCZ11, PS17]. Bell [CHZ18, ACFL16, FR94]. Bellman [HS88]. Belong [Més16b]. Belonging [LM12]. Below [COHH17, LPS18]. Bend [DM18b]. Bends [ZN08]. Bent [ÇMP15, CK08a, KCTR13, TKMM19]. Benzenoid [ZLS08]. Berge [GP17, GMOV20, GNPV20, Gyá19, HT19, MTV08, NV22, ZKNS20]. Bergman [Ber20, DD15]. Bermond [LPS09]. Bernoulli [NLS24]. Bernstein [GSS14]. Bertrand [Ete22]. Bessel [Ege10]. Best [GKM12, GMS15, HV00, Jon20]. beta [BDEK06]. beta-Skeletons [BDEK06]. Bethe [CFKK17, CCG^{+}11]. Better [GPSS01]. Between [CCO^{+}15, LN24, SA90, Wil99, ANS16, BYR05, BH20, FHM94, GPW13, HT19, JK24, KL19a, Lab13]. Betweenness [CS98]. Beyond [ABFRV24, ACD08, CLS15, DK22, HW24, KK14a, Neu24, DW24, DJMV21]. Beyond-Planar [HW24]. Bézout [WWKY11]. Bi [BKK16, BGH^{+}17]. Bi-Covering [BGH^{+}17]. Bi-set [BKK16]. Bias [BGY20, FHLT21]. Biased [BGM22, BN96, Cha16, MS24, MR15b, NvdP19, NP20a]. Bichromatic [AHH^{+}10, Pay17]. Bicolorable [BGL07]. Biconnected [ÀB23, PSW96]. Biconnectivity [HK05]. Bicontactual [dR14]. Bicriteria [SSS13]. Bidegree [BR17a]. Bidimensional [DFHT04, DHT06]. Bidirected [DZ09a, EJ01]. Big [CR15]. Bigraphs [Cor24, HH05, Saf21]. Bijection [CMS09, TKA18]. Bijections [PV10a]. Bijective [FW02, HK11]. Bimodality [DH05]. Bin [BDGP22, CCG^{+}00, Eps06, FL91, JK19, Sei01, Woe93]. Bin-Packing [Woe93]. Binarisation [CCJ^{+}17]. Binary [Ald22, ACLT01, BDM02, BZ20, CLVZ96, CM12b, CG02, CSW17, DH05, FMP08, Fis90a, FM24b, GZ06, Gol96, GvZ17, GL10, GPW13, GJ12, HHLÖ95, Han09, Hay21, HWS18, HvIK^{+}07, Kim91, KMP03, KM95, KMPR14, MG19, MR12, Moi08, NN22, NvdP24, OPR12, ÖV04, OS22, PP13, Pon20, RSW12, Saw07, Wag21, Web08, Wil05, Yam20, ZLS08, vIKL^{+}16, HS97]. Binding [HLZ13]. Binomial [BBMM09, KKP15, Mal89]. Binpacking [FK98]. Biodiversity [FS12a]. Biorthogonal [Kim92]. Bipancyclism [HLZ13]. Bipartite [AAHLT10, AKZ17, AS07, AS14, ASSW21, BB16, Bal08, BIKY18, BHT10, BCP08, CHKW24, CWY21, CMPS17, CCH21, CY18, CP16a, CM07, CY20, CDK10, CD11, DZ09a, DEK22, FMMO93, FF06, GL95, GRY08, HvtHLN12, HVLP13, HHLM20, HS10, HK05, Jan20, JQ20, KW08a, KS03b, KS07, Lat23, LLY10, LHHL18, Mak07, MT90, McC21, Mot19, NOO12, Nog24, PT14, Pik03, Pip02, PT94, PSML08, Riz02, Sim13, Tak22, Tan21, Wal19, Wan02b, ZZW13, Zha09, Zha11, ZGH^{+}23, dCLM13, Jac92, dH89]. Bipartition [AAHLT10, LSSZ19]. Bipartition-reversing [AAHLT10]. Bipartitions [KKO16]. bipartization [CNR89]. Birigidity [Ser89]. Birkhoff [DKWL20]. Bisection [AHFM08, KL19a]. Bishellable [ÁAFM^{+}18]. Bisubmodular [BC95, FI05, HK14, Shi12, Fuj97]. Bit [WB90]. Biuniform [BPWX13]. Black [KOS16]. Black-box [KOS16]. Blanks [ASH18]. Block [AFG21, AMPT93, AM95, CM05a, CH01, FSW13, FGKP22, GMZ09, GH22, Han98, LS16, MMJF03, PPR22, Ran02, Wei19, ZGL^{+}09, Lin97]. Block-Type-Decodability [CM05a]. Blocking [BB23b, BD01, FHS14, HKP22]. Blocks [CDHH14, DEE17, DHKM11, FRW12, RR18]. Blow [BKTW15, JMN22]. Blow-Up [JMN22, BKTW15]. Blowups [FJK^{+}23]. Bodies [BH16, BG12, HST22]. Body [BSiT15]. Body-and-Bar [BSiT15]. Boland [LM16]. Bollobás [HWZZ18, HWZZ20]. Bolts [KMS98]. Bond [Ull14]. Bonds [ZWW23]. Bondy [CCO^{+}15]. Bonferroni [Doh03]. Book [BS15a, DGL11, EM99, LP21b, MW90, NOO12, ONN19, dKPS13, BRK89, LB09]. Book-Cycle [LP21b]. book-embeddings [BRK89]. Books [CFS20, LL21]. Boolean [AL95, ARS17, Bac09, BS23, BCE^{+}00, CC03, CK91, DSW90, Eng04, FF06, FLM22, GRR15, Gol06, HSLd88, HKR00, LS09, Lub90a, MNPR17, NW95, PRS02, PZ98, STV21, Tsa96, WS19, dGV05]. Bootstrap [BP15, DNR23, HM20, HL24, SS91]. Boppana [FT05b]. Border [BP16, CH89]. Borel [TA93]. Bottle [iKKKL09, iKMN09]. Bound [AF05, BB13, Bar01, BL19b, BN01, BL17b, Bra05, CTU14, CY08, Chi23, CFM94, CS91, CR17b, Däu22, DSZ05, DH91, DJW12, DvHT20, EP10, FS09a, FLL10, Fra95, Gol06, GC11, GÖ12, HJ18, HTS18, JNiT21, JMU23, KKP11, KK14a, KW13, KL92, KMRR09, KSS09, Kun23, Lat23, Lav16, Lei94, LN24, LM22b, Liu14, LY10, Luc03, Mer99, Nao91, Nie00, OS16, PVY22, Pry23, PP90, RC98, RS15, ST13, SS19, SS11b, SS18b, SL95, Ste07, Tod14, o09, Sar97, SG16, Tar88]. Boundaries [AF05, KiT13]. Boundary [AE03, BB23a, Lai18, vD11, LS89]. Bounded [AKP^{+}22, AD96, ABC^{+}15, ASSW21, Bab92, BLP21, BST20, BFGR17, BHLP92, CCVZ10, CCSW23, CKK^{+}04, Che94, CCL^{+}06, DHT06, DP92, DEJ^{+}20, FJK^{+}19, GGHP22, GVW06, GnN06, GWZ18, HKM20, HKK^{+}09, JK99, JB22, Jon05, Joo16, KiK12, KLWY21, KPP13, KR23, KKS17, LR04, LM11, OPR12, Oze13, RW19, SW14, Sei01, Woe93, Yos19, Zak14, ZM02, BCLR89, LP88]. Bounded-Degree [HKK^{+}09, KR23, KKS17, BCLR89]. Bounded-Genus [DHT06]. Bounded-Skew [ZM02]. Bounded-Space [Woe93]. Bounding [AH96, BCG^{+}24, BDF^{+}18, BM11, BELP21, BC17, CP10a, EK13, GI97, HW16, JLL16, Lab13, LS22, Zuc92]. Bounds [AHKiO22, AM21, AS10, BST20, BEL09, BHL^{+}15, BGJ^{+}12, BGS96, BJK14, Buk16, CKL^{+}21, CHZ18, CHLZ00, CL05, CNP21, DMR11, DW10, DW11, DHUZ01, DNR23, DSST13, DMS12, DS16, EMRPS14, FFie98, FP01, FP04, FHNN18, FL22, FXZ23, GG15, GK07, GP91, GRS12, GY23, HT13, HHLÖ95, HBDS23, HK02, HP21, JWF05, KKW17, KMÖ18, KT19, KM95, KW14, KT17, KŠ08b, KO06a, KOP^{+}21, LN17, LL99, Lam20, LRTW11, LZ09, LNO96, LS16, Mac18, MPSS20, MP13, MSD19, MS19, MNP08, MTGK05, NW95, NÖ08, NS11, PdCSC21, RTS00, RN23, Raz20, RS22, RZ05, RR03, SC17, Sta11, SS02b, Swa05, TW21, VVY15, XSR11, YAT16, Yu17, Zha94, Zha99, dKMP^{+}06, dKPS13, CET97, Car88, SSS95, Tam88, Tsa96]. Box [FHK96, KOS16]. Box-Greedy [FHK96]. Boxes [AC14, CC07, CFP11]. Boxicity [ABC11]. Bracelets [ACM11]. Braess [LRTW11]. Brambles [MPRS22]. Branch [DMR11, DL19, JKiO21, KML05]. Branch-Decompositions [JKiO21]. Branched [KML05]. Branches [Yam20]. Breaker [CFG^{+}15a, CHMP21, EFKP15, Han19b]. Breaking [ILM20, LB09, GPS88]. Brenti [YZ17]. Bricks [CFLZ19]. Bridge [BJS22]. Bridge-Depth [BJS22]. Bridgeless [Mác13]. Bridges [Wol10]. Bridging [GHW05]. Bringing [SS10]. Broadcast [AR17, BMRT20, EK05, FY04, GV92, GPP01, GP91, KP95, LP88, dBSS24]. Broadcasting [BHLP92, BFPP08, FP01, FP04, Gab05a, KP04]. Broken [ACM11, Swa05]. Brook [AKW05, KLP12]. Brooks [FHH08, FKS05a, HS90]. Brooks-Type [FHH08, FKS05a]. Brouwer [BV10, Gu21]. Brown [Sha23]. Brownian [Pal12]. Bruhat [Abe91]. Bruijn [ES98, KRR16, Mit97, Obr93, RSW12]. Brunn [AF10, CIN18, CL22, OV12]. Brushes [AaW09]. Buchberger [UWZ97]. Budgeted [BDGP22, CMPS17, Jüt06, KPS20]. Buffer [ELvS11]. Building [CFG^{+}15a]. Bull [CS18b, LM08, Pen12, dFM04]. Bull-Free [CS18b, LM08, Pen12, dFM04]. Bundles [Zhe16]. Buneman [DHKM11]. Buser [LMP19]. Buy [LPSR12].

C. [Naa01]. Caccetta [ABC^{+}23]. CAD [SW04]. Calcul [Zen90]. calculation [Zen90]. calculus [PU95]. calls [Ser88]. Can [ASS09, BBS17, BS16c, CCNV11, Che16, Gao18, KSW17, MPŠ08, PP12, Kar89].

Canceling [IMS05]. Cancellative [BS95a]. Canonical [AM95, AL24, Fra90, HKL99, HKW15, Rea15]. Capacitated [ACLW18, BYFMR10, EEL09, HI23, KS00, NI92]. Capacities [BCF^{+}10, KS03a, Kor18, Yam05]. Capacity [ACG^{+}20, BOS01, DKMS17, GR99, GZ98, HTS18, Kas03, KSF19, Pip91, KM06]. Caps [GL10, TW21]. Carathéodory [BCD^{+}12]. card [KLW89]. Cardinality [Eps06, HKMY22, LM22b, Luc03, MNN18]. Carlo [SV11]. Carries [Alo13]. Cartesian [BK07, CHM^{+}07, CHK10, PZ05, Špa07]. Case [ASZ02, BCE^{+}01, BCC^{+}22, CCG^{+}00, DF04, FHL^{+}13a, FHL^{+}14, HK16a, KM05a, PV10b, SS95, SZ15, Tak08, HS89b, Tas97]. Catalan [LLL18, Abe91, SZ15]. Categorical [BNR96, Tót10]. Category [ST23]. Caterpillar [ETT13]. Catlin [LHLZ23]. Causal [LRS23]. Cause [BHH94]. Cause-Effect [BHH94]. Cayley [ABM14, ALZ96, BK19, DGS96, DF04, ES98, FFie98, GGI07, HXZ18, Jia95, Lev15, MWWZ20, MM93, Mra17, PRS^{+}23, RS93, Shp15, TA93, Tsu23, Zho05, Zho09]. ce [KvIL^{+}12]. Cell [Smi01]. Cells [RC18]. Cellular [MM93]. Census [FT17, MW08]. Center [BDLR01, KS00, Tam88]. Centered [BH16, Len98]. Centers [Bev09, CR10]. central [WW91]. Centrally [Nov18]. Certain [BC11, Bro11, GM90, LW88a, PR91, Ros09, Zho93]. Certificates [KT16]. Certifying [HH05]. c’est [KvIL^{+}12]. Chain [AK22b, BKS10, CLY05a, DL12, DM19, Enc05, EMT15, GVKSS06, GP08b, HN23, LSL92, LCV03, MMR06, MW20a, Spi19, SV11]. Chains [BK90, CDMT21, CEFY24, CLY05b, DJW12, DL18b, HH92, KLR22, Rea08, RS22, SYKY18, DSW90, Pat88]. Chairs [ABF^{+}14]. Chang [IMR14]. Change [PSTF00, Zha94]. Changes [AM06a, FMP08]. Changing [BBG08]. Channel [DS05b, GS95, GJ08, JWF05, KSV05, KS03c, Krá06, Lei94]. Channels [GF08]. Characteristic [BPEY22, DQW^{+}15, Gly12, LÖ05, RT18, Wu09, XY15]. Characterization [Bol90, BP16, CW16, CLiO18, CH11, EMOT16, FHK96, FG01, GSZ22, HH04, HSS19, HMS93, HHH^{+}02, LW17b, LS05, Mar09, NOP14, Tuz08, Wag18, WH15, WZZ09, Zho93, vWW94]. Characterizations [BL04]. Characterizes [BJS22, Tod89]. Characterizing [BTW08, GW19]. Characters [BS15b, HMS05, KMW24, LGS11]. Charges [DW10, DW11]. Charts [GW00]. Cheaters [BS91]. Cheating [OKS06]. Chebyshev [Mun06]. Check [BM15, CF09, CHW88]. Chernoff [SSS95]. Chessboard [ABFV23]. Chinese [GJW16, SL96]. Chip [Eri96, GK16, JSZ15, Tar88]. Chip-Firing [Eri96, JSZ15]. Chirality [DH91]. Choice [FLM22, GMS15, JL23, Jon20, KM13c, XZ22, Yok19]. Choices [Red17]. Choosability [CDWZ17, DLŠ10, WL02, WX13, WL10]. Choosable [CW09, DLS11, Far09, Mác13, YWWL21, Bon15]. Choose [GKM12]. Choosing [CL16]. Chord [Saw02]. Chordal [AAS22, AS06, BGL07, BDPR18, BDCV98, CP16b, FGP10, HC98, JP18, KN16, Krá04, LS15, LC04, RLWY21, Sak94, SB94, BCD97, LM89]. Chordality [GHvtHP15, HHCHL21]. Chorded [CGH^{+}15, GHM14]. Chromatic [APdMOZ23, AM06b, BHLY08, BF24, BCHW17, CLS09, CP20a, CH19, CHK10, CS21, DK06, DK10, DJ11, DŠ09, EK13, EHJ01, FKS05b, FHL13b, GHM21, GKL99, Har19, HZ10, HP23, KLMR13, KZ04, KR15, Liu14, LM10, Mra17, Ngu24, Per16, PS21, PdCSC21, RS15, Sch92, Smo07, dS91, dMP93]. Chvatal [BCH92, ACD^{+}13, CSS13, CDD^{+}15]. Circle [Fra90, HIMW24, Sch91a, VZ93, Zhe16]. Circles [ERSZ22, SW04]. Circuit [BCSK07, BFH15, FM24b, KPS19, MRS19, PL94, Swa05, Wag18, WY20]. Circuit-Switched [BCSK07]. Circuit-Switching [PL94]. Circuits [AB94, AS02, CL15a, CH13, CH17, CL91a, CG07, FG14, GV21, Juk21, KiK12, KW90, Kot13, Luk20, SB91, SR94]. Circulant [Cre04, FKK05, FKK07, GW19, LLZ99]. Circulants [Col98, LSX14]. Circular [BHH96, BH13a, BFN20, CL20b, DEG^{+}07, FHM03, FHS14, HIMW24, KNNW23, KW17, LS08, LZ05a, LM10, MS21b, MR04b, Saf21, SSR94]. Circular-Arc [FHS14, LS08, Saf21, SSR94]. Circularity [Saf21]. Circularly [Saf21]. Circulations [LS08]. CIS [WZZ09]. city [RX88]. Class [AK22b, BILL23, Che22, Dan09, DS05a, KCTR13, LLW22, Lou10, MS24, Sca03, Zho09, van94, BS88]. Classes [AS06, AFK^{+}22, AS10, BG11, BDJ^{+}20, BL23, BPT91, BLL^{+}15, BC23a, BH15, CKP^{+}21, dOCHLO19, CW09, DMS21, DFK^{+}11, DEJ^{+}20, DKT22, FG22, FKKL98, FH21, FKV21, GPR^{+}21, HH04, dAHFdFK10, HN20, KM23, KNNW23, Kie97, jKiO20, OYY13, OS22, PRS98, WZZ18]. Classical [CHZ04, XSR11, Exo89]. Classification [BÖ05, EK10, Fed01, FF06, Sim13]. Classifying [AA96]. Clatworthy [GG11]. Claw [BS10b, BS16c, CR13, CR17a, GPvL15, MRV17, PRR02]. Claw-Free [BS10b, BS16c, CR13, CR17a, GPvL15, PRR02]. Clean [ACGT22]. Cleaning [AaW09]. Clique [BMRT20, BDJ^{+}20, BCD97, BC23a, CCL^{+}06, CD18, DMS21, Fei04, FRRS09, GnN06, HJ18, Lam20, LR04, Mor21, Mra17, SB94, Wan23, XGG15, GHHY96, Jac92]. Clique-Free [HJ18]. Clique-Width [BDJ^{+}20, BC23a, FRRS09, GnN06, Lam20, LR04]. Cliques [AHDHL22, BGG^{+}04, BCKN21, BHH96, BBC^{+}23, CR15, Dan01, DŚN21, FW20, FPS18, JY23, KR23, KO06a, LiO15, McK19, MNS14, NY21, Yus14]. Clock [CKK^{+}04]. Cloning [CS12]. Clos [CH06a]. Closed [BDJ^{+}20, BJS22, CJL^{+}23, DR04, DEJ^{+}20, GPR^{+}21, HN20, OS22, Per16]. Closed-Convexity [CJL^{+}23]. Closedness [MD16]. Closely [SV11]. Closer [BIT13]. Closest [BGW20, DSL19, MPSV21, MGC14, SFS09]. Closing [BLS19, BCD21, FMP17, MS21c]. Closure [BGJ^{+}12, HQ03, KMR^{+}23, KKS22]. Closures [ACD^{+}13]. Cluster [BGG^{+}20, GKNU10, Tza08]. Clustering [ADPR03, BV21, Flu24, GKNU10, MRT11, Ram04]. Clusterings [BG12]. Clusters [CLW09]. Clutter [NS93]. Clutters [AP18, ACGT22, CG02]. CM [Swa05]. CNF [FFV11, HS06]. Co [BLP21, GM04, HR05a, MH09, NS16]. Co- [HR05a]. Co-2-plex [MH09]. Co-Cyclic [NS16]. Co-degree [BLP21]. Co-Site [GM04]. Coalescing [CEOR13]. Cocircuits [OW22b, Wag21]. Cocircumference [sJS24]. Cocomparability [CDHK16, HHLM20, HHL24, KS93, MC12, MNN18]. Code [BB13, BN01, Kas08, LT01, NLJM21, Saw07, CS89, Lic98, WS96]. Codegree [APS22, CM14, FRMPV15, FRZ16, LZ18a, Sid18, ZG21]. Codegrees [Spi23]. Codensity [CCD^{+}23]. Codes [Adr23, AM07, ABNR22, ACLT01, AGK^{+}23, AMPT93, BOP94, BvH03, BM00, Bar02, BZ04, BGY20, BGÖ17, BHL05, BGS96, Bla03, BEN08, BK91, BHTK^{+}21, BLL^{+}15, BÖ05, Buk16, BJ23, BE13, BR17b, BR19, BC23b, CD93, CCD00, CJL^{+}23, CGL10, CKPS15, CGK^{+}19, CH01, CS89, CGHL23, CR23, CHLZ00, CF09, DR04, Din13, Dow91b, EK10, Enc05, Etz96b, EV98, FG01, FL22, FT05b, FHMY01, FKMS10, GDCM20, GF08, GMW05, GOR20, GKS12, GRSZ23, GÖ12, GV21, HHLÖ95, Han98, Hed08, HKW15, HWS18, HXZ18, Jan00, KMP03, Kit02, KCTR13, KM11, Lai05, Li17, LÖ05, LS06, LWW10, MG19, MWWZ20, MMR06, MP22, MP98, MP21, Moi08, MMSJ08, NÖ08, NLJM21, PG06, Pon20, Ran02, RR03, ST14, SS09, SS10, Sta11, SW98, TRV03, TSN04, UV15, VV94, Wil05]. Codes [WC12, ZSW11, Dow91a, Etz96a, Mou94, PL99, Sav97, TZ97, CZ97]. Codewords [Adr23]. Codimension [CMSM^{+}18, PPTY22]. Coding [GRR15, Gad18, JLR20, KSV05, ST20b, SZ94, ZLS08]. Coefficients [BI13, Lla06, TKA18, Mal89]. coexistence [GKRS15]. Cofactors [IT09]. Coffman [CM05b]. Cograph [BCHP08]. CoGraphic [GS16b]. Cographs [BM93]. Cohen [Bro11, BCW96, CN12, DDL20, PPTY22]. Coin [AD11, BN96, Cha16, CGG88, Sak89]. Collapses [MST21]. Collection [DO08]. Collective [DYL06, Sak89]. Collector [MW03]. colliding [KM06]. Collinearities [BDJ^{+}15]. Collisions [CTW93]. Color [AS10, DFJS15, FPS20, FR06, FHLT21, FKV21, GJXZ20, JLR20, Mot19, Yus09]. Color-Avoiding [Yus09]. Color-Bias [FHLT21]. Colorability [DL18b, DH20b, HKSS08, AK02]. Colorable [CS18c, CKP^{+}21, CS02, Dan01, FT12, GH19, GK04, JKSW17, KRS11, LM17, LS03b, MM12, NNO19, WHW14]. coloration [HS97]. Colorations [Bah23]. Colored [AHH21, AF19, AL24, BFa24, BS09, BCF^{+}20, CH17, CGK^{+}19, CFP11, DGM12, EFK18, KW92, KY18, KMMS21, LSS17, Lo14, MP08, McL10, MWW94a, CH13, IS93, Zen90]. colorés [Zen90]. Colorful [ARS23, BM07, CKMU14, CFH23]. Coloring [AAD^{+}18, AKP20, AH03, ANP14, ADKS18, BGG^{+}04, BCC^{+}11, BTU09, BLM20, BHE05, BSY21, BKP21, BKM13, BDDP24, BCP08, CKP13b, CLiO18, CGSZ20, CM16, CR13, CR17a, Cra19, DEG^{+}07, DPR22, DX19, DŠT08, DH20c, EEFH21, EST14, EEL09, EHLP11, FKS05a, GKLT19, GN08, GKR^{+}18, GK04, GS20, GMOW21, HSZ13, HLS22, HLS24, HS90, HM11, HSŠ08, HK16a, HK96, JKSY22, JL20a, JSRSW18, KSS11b, KP16, KNNW23, iKM10, KPT94, KMŠ^{+}09, KNP05, KZ08, KSŠ08, KLS18, Krá04, KS08c, Lam20, LM08, Lu04, LSSY10, MN18, Nao91, NK90, PZ10, Pen12, Riz02, Rya07, Sig10, SV08, Wan08, WW22, Xu09, Yus03, vBEM17, GKRS15, Kie88, KPT95, Ray94]. Colorings [AS10, AG19, BGLS07, BL19a, BVVW11, BW02, BD24, CtJL01, CT21, DJKP09, FXYY14, FHM03, FHH08, FKLW98, FW02, FR06, Fue23, GL15, GvdHM^{+}08, HJ18, HGY20, JK99, JT11, JS14, KMŠ^{+}09, KS23a, LM21, LM11, MŠT09, Nor11, Off08, Prz13, WW18, KKV11]. Colors [CR13, CR17a, Cra19, DF10, HJ18, KK14a]. Colourful [DMS14]. Colouring [Mac91, BJHM88, BC88b]. Column [OPR12, PSWX24]. Combinations [AL24, GMTW15]. Combinatorial [AYZ04, ABNR22, ACF18, BHM16, BSS14, BDT17, BG11, BJK13, BGL24, Boy96, CHM^{+}23, CC07, DMS14, DO08, EGR08, FS05, FL92, Fra95, FP13, GMW05, GRS12, HRS17, Hir11, JJiT14, KPS19, Kim92, LLL18, Mar09, Mat19, MW19a, Mun06, Ngu13, Prz13, Sim90, SW98, Wil16, Sav97, SV88, Yen97, MPP17]. Combinatorially [SGM20]. Combinatorics [ABY14, BCE10, DJS12, GR17, Mar20, MWW94b, SW99, SS21, ST10, Tha08, WDSH21]. Combined [Pfe15]. Combing [GST23]. Committees [WS12]. Commodities [Fle00]. Commodity [GS16b]. Commodity-Flow [GS16b]. Common [Ald22, BCdMR08, BM14, BH20, DSST13, HKL11, NvZ15, Web08]. Communication [AP92, BYKKR18, BOT92, BH93, CT93, KS92, KKN95, Kus92, Orl93]. Communication-Space [AP92]. Communications [Bon91]. Community [LN21]. Commuter [AL17]. Compact [ACL^{+}06, DKWL20, Har01, HKW15, KM01, LLS04, Lu10, dS91]. Compactions [FGZ19]. Compactness [GP99]. Comparability [HHLM20]. Comparative [KvIL^{+}12, MK09]. Comparing [FKS03, FKM^{+}06, HLR92]. Comparison [AB94, RC18, ST14]. Comparisons [DZ01]. Compatibility [ASY14]. Compatible [KR92, Saf21]. Compensation [AG06]. Competition [Ete22, FLM^{+}95, Gui98, MSS14, IKM^{+}92]. Competitive [GL95]. Complement [Ram98]. Complementarity [AGM23]. Complementary [Fra10, LSX14, LHHL18, LW88b]. Complementation [BDJ^{+}20, BH13b]. Complements [CY12, KK14b, OS22]. Completability [JJiT14]. Complete [ACGH20, AS07, ARS21, BGM08, BCCZ11, BCLR95, BK91, BL01, CWY21, CDM00, CDM04, CHLT24, CP16a, Con05, CS21, DHJ^{+}13, FKK20, FRRS09, FH10, Fox10, GG11, GMA15, GMS00, GS00, GKP19, GSZ22, Gut93, Han98, Jan20, iKSZ04, LLY10, LZ18a, LLZ22, Pip02, PSML08, Sid18, SZ13, Tót10, Wal19, dKPS13, Jac92, Vav89, dH89]. Completely [FG01, GS00]. Completing [CDGO22]. Completion [BFPP15, ELSS17, FG00, Som14]. Completions [CKP22]. Complex [BTU09, BCKP19, FP99, GGHP22, KLMR03, SS08, RM97]. Complex-Valued [GGHP22]. Complexes [ABY14, BDM^{+}24, BHM16, BW99, BV20, Bro11, Cha91, CKP22, CDGO22, CN12, DDL20, DD15, EGS13, Hir11, Jon05, KN13, LS03b, Més16b, PPTY22, Shu23, Swa05, Tza08]. Complexities [Do19]. Complexity [AR08, AM11, BZ04, BBRT23, BGM22, BN01, BT96, BCF12, BKP21, Boy96, BCKP19, CEHS08, CC07, CT93, CI07, CLST12, DSL19, DL89a, DM18b, FKL93, FMP17, FJLS03, FGZ19, FLM^{+}18, GGHP22, GKMT17, GS93b, GNS11, GvdHM^{+}08, HLS22, HKP01, HM20, Hoà10, HKL^{+}21, JLR^{+}17, Juk16, KS92, KKN95, Kas08, KBE^{+}05, KLS18, Kus92, KOP^{+}21, LZ03, Mac91, May08, MC93, MC06, Onn91, Pad16, PW02, Saa93, ST13, SW23a, Sno13, Zer11, BJHM88, KW96, PU95, RSSW88, Tam88, Win88]. Component [Blo10, CL06, CF16, CY21, De 24, KM19b]. Components [ÀB23, BAH10, CFDK20, GW99, Luo23, MP08, SW01, JO95]. Componentwise [HK05]. Composition [BJK14, HS04, WC12]. Compositions [ADM^{+}15, BM94a, BM94b, BM94c, BFM94, KR93]. Compound [FL91]. Comprehensive [AT16]. Compression [CCM^{+}23, MPSZ19]. Computable [JMS90, KR04]. Computation [DM13, GDVL17, KR98, KM05b]. Computational [BM11, BGM22, GS93b, Onn91]. Computations [BCE^{+}00, KOR03, KOP^{+}21, Car88, KM97]. Compute [BKS09]. Computer [CDP94, Kar20, SW99]. Computer-Aided [SW99]. Computer-Assisted [Kar20]. Computing [ACD08, BCG^{+}24, BBDK00, BCdMR08, CCH14, Dan09, DDS16, Duc21, DJT15, FvIKS15, Gal18, GnN06, GS16a, HKP01, HTV05, HvtHLN12, IIL14, IT09, JP06b, KN16, KL19a, KV90, LZZ22, MP13, NI92, NNI97, Nag10, Owe11, ST23, WY10, LW88a]. Concatenation [BB23b]. Concave [AAMM24, BJHY00, DSS12, FOST10, Vaz13]. Concave-Round [BJHY00]. Concavity [Brä10, GMTW15]. Concentration [CV09, FTS24, GLS15, Mra17]. Concentrators [Pip91]. Concept [COCF10]. Concepts [KT24, KS23b]. Concerning [DKS16, Kla06, LdCKM18]. Concise [LL14a]. Concrete [FKKL98, Hof95]. condensed [CH89]. Condition [CHKW24, CSS13, CS19, CKKO10, CM14, Enc05, FZ08, FZ22, GMS00, GM03, HLZ13, KM14, LGS11, OC19, PPTY22, Sak94, Sch10, ZLWC12, GY92]. Conditional [Che24, HL10]. Conditionally [MB18]. Conditions [APS22, BPSS19, BPSS20, BR17a, CL20a, GJ06, GMR^{+}21, JJiT14, Pfe15, Ste00, TZ15, WL03, ZZBL17]. Conductor [ACM11]. Cone [AADM18, PZ10]. Cones [BBK^{+}16, JL20b]. conference [Ser88]. Configuration [ABCG17]. Configurations [FPS20, FYK00, GH13, MM93, Raz10, SW04, SW99]. Conflict [AAD^{+}18, ADKS18, BKP21, CKP13b, HGY20, MMSJ08, ZSW11]. Conflict-Avoiding [MMSJ08, ZSW11]. Conflict-Free [AAD^{+}18, BKP21, CKP13b, HGY20]. Conflicts [CM18]. Conformal [Che17]. Conformance [SL96]. Congested [BMRT20]. Congestion [CEP18, MPRS22, Ram97b, Ram98]. Congestion-free [Ram97b, Ram98]. congruences [Jae89]. congruent [Kle89]. Conic [MD16]. Conics [BVdZ16]. Conjecture [AK14, ABC^{+}23, AAMM24, ACM^{+}18, ÀB23, BT14, BP13, BLS19, BBJ^{+}21, BCD21, BLM^{+}22, BBT16, BBM09, BLO^{+}22, BHT16, CCO^{+}13, CCD^{+}23, CEOT15, CKPS13, CG02, CR19, DS21a, DKS16, EKK^{+}15, EEFH21, FGS19, GM20, GH06, GW07, Gu21, GJWZ23, Han16b, HK16a, HK16b, HKP^{+}17a, HKP^{+}17b, HKP^{+}17c, HKP^{+}17d, HWZZ18, HWZZ20, HKS07, JL20a, KKS10, Kar20, KSY18, Kun23, LPS09, Lic14, MB18, NvdP24, Roz19, SS21, SGM20, ST17a, Sto12, Sza08, TW19, WS06, Zha23, BW92, Dow88, HL92, CCO^{+}15, Cho09, CK11, HRS12, HY13, dH04, KO12, Rab08, Shp10, SZ15]. Conjectures [Gly10, KKL19, LW03]. Connected [ABHM00, ADO23, BC02, BJHY03, BAH10, BL17b, BCC^{+}22, BJS21, CWY00, Che22, CSS01, CL13, CDP94, CDMO16, CGOvZ21, CEOR13, CL21, CY03, CLY05a, CLY05b, Der12, DL12, DD13, DJ11, EIiK08, EKM^{+}19, EEFH21, EG03, EHJ01, FJ09, FGPS19, FLM^{+}95, FM13, Gab04, Gab05b, GZ06, GvZ19, HW16, HV17, Hof98, JW18, KiKK17, KKKS17, KPT22, iKO16, KPS20, KKO16, KRS15, LRWZ12, LC12, LQ22, LPS18, LZ18b, Luo23, MMSW23, MS16, MW14, MR12, MZ22, Ngu24, OSW16, OW22b, OZ18, PTT16, SS11a, SW01, Wol10, ZWW23, BM97, BBM90, Cho94b, RX88, Voi07, Ste10]. Connectedness [HT90]. Connecting [MW19a]. Connection [DFT15, ERS19, LMP19]. Connections [BGL24, Car09, CHW10, GV21]. Connective [GL14]. Connectivities [HvZ14, JA16, LiO23]. Connectivity [AKM22, Ana18, BJFJ95, BJJ98, BJGJS99, BILS22, BK21, BK24, Bev10, BMNS22, Cal13, CDHH14, CK14, CW14a, Cho92b, CG02, DMS08b, DMS08a, DGS96, DP16, Fin09, Fle05, FGT18, Fra92, FL10, GGW06, GM90, HLST00, HI23, JS03, Kao96, KW90, LLL17, LRS^{+}21, NI92, OC10, Pfe15, Ram04, Ste22, Vég11]. Connectors [Kar92]. Conolly [EIJ^{+}12]. Conolly-like [EIJ^{+}12]. Conquer [ARS95, ÁS09c]. Consecutive [DHJN02, Ehr16, GMZ09, MR21, RR18]. Consecutive- [DHJN02]. Consensus [BJ91, BFN20, MP95, MP04]. Consequences [BGW20, HK96, HKK^{+}09]. Consistency [SY11, Tod89]. Consistent [BK11, CK18, Abe91]. Consisting [EIiK08]. Constant [BDGP22, CGK^{+}19, CEP18, FR94, Het14, Jan10, Kun23, MPRS22, MMSJ08, SS11a, WC12, Car94]. Constant-Ratio [BDGP22]. Constant-Weight [CGK^{+}19]. Constants [DDS16, GL14, MLMR23, OS13a]. Constrained [BGS17, BHM^{+}22, BDK^{+}24, CM05a, CMM^{+}24, Gol96, HMP04, JNiT21, JP06a, Jor03, KS03a, Kas03, LS95a, Mar20, SL95, Tov90, Hef97, RTW97]. Constraining [SW04, SW99]. Constraint [BK11, CCJ^{+}17, CM12a, DKLW24, EM20, FF06, FK17, GM04, KJJ04, KL08, MMS15, MT11, MRT11, Yos19, ZK11]. Constraints [ALM^{+}18, BJGJS99, BKK16, BK21, BK24, BMN13, DH91, DLŠ10, Eps06, FKT06, FGP12, FGP10, GS13, GM90, HKMY22, JPZ21, KT14, Kam17, Kam19, Kas03, KiT13, KNS05, KM94, LMNS10, Lou10, MW90, OR04, Ste22, dMP93, PS97]. constructed [TZ97]. Constructible [TZ19, TZ20]. Constructing [ASZ02, BT18, BHTK^{+}21, BB03, BLR16, BV21, Che07, GS89, ILM^{+}16, KM95, KM94, WC12, BLR17]. Construction [Ald90, Bon10, Boy01, Cap03, Cha03, CKPS15, CGK^{+}19, CCG05, FS91, HJ94, KST06, Lu08, Pip95, ST14, SWKP10, AS97, CGG17]. Construction- [CGG17]. Constructions [AM07, AB94, ACG^{+}20, BER11, BV20, CTU14, CS02, DA10, GG15, GMW05, SG16, SW98]. Constructive [CL05, CPR99, Nag17, XSR11]. Contain [ARTV12, MM15]. Contained [KLMR18]. Containing [CM24, CY18, FJK^{+}19, Het14, ZL11]. Containment [IN22, KT16, SSSU89]. Content [BYKKR18]. Contiguity [AG19]. Contingency [Sul05]. Continued [Het14]. Continuous [FFLP20, HK99, SA90]. Contour [Git99]. Contours [FP99]. Contractible [DS06, DL14, EIiK08]. Contractile [LM08, MT05]. Contracting [DMR11, HVLP13]. Contraction [AFL^{+}20]. Contractions [BDD^{+}19]. Contrast [BDDS03]. Control [FPR18]. Controllable [SSS13]. Converge [PP12]. Convergence [ACD^{+}13, CS14, PS17, SV11]. Convex [AS16, Ave13, BJHY00, BH16, BT93, BJ23, CD16, CG17, DST01, DMN21, GVW06, HW96, HST22, HQ03, HR05b, KMW24, KT24, KMT07, LS05, MD11, MPS^{+}09, MvLvL13, Mur06, NO09, Onn03, Sch09, SA90, ST07, dCST20, Vaz12, EFF91, MRS89]. Convex-Ear [Sch09]. Convex-Round [BJHY00]. Convexity [BCD^{+}12, CPRdS13, CJL^{+}23, DNB99, HI23, HRS93, LSTY17]. Convexly [HMS05]. Convolutional [Kit02, RF12]. Convolutions [Cha16]. Cooperative [Bil03]. Coordinated [SST22]. Cop [BELP21, CCNV11, CCPP14]. Copies [CL07, CM24, Let19, Pit23]. Coprime [KSY18]. Cops [CNST24, FGP12, HKM20, SS11b]. Core [BKL^{+}15, FK98, GMRT11, HW18, SZ15, VVY15, Wan02a]. Cores [CS02]. Corners [HW18]. Corollary [GLW11]. Corona [BPS07]. Correcting [BGS96, CD93, CGL10, CGHL23, DA10, FT05b, GMZ09, GOR20, KM11, CZ97]. Correction [ADH^{+}14, AG06, JMSW00]. Correctness [PS17]. Correlated [Orl93]. Correlation [AL95, Bac09, Wag07]. Correlations [DJKO19, KM11]. Corrigendum [DJM^{+}18a]. coset [CZ97]. Cosets [CHZ09]. Cost [BCG^{+}10, BL17b, CL13, CGP98, FJ17, GRY08, HSS19, HR12a, HQ03, IMS05, Mak07, MTGK05, KP06]. Costs [BYKKR18, CV07, DSS12, FKK21, FOST10, FT05a, FJ17, SS11a, vWW94]. Coteries [MK01]. Count [Har19]. Counterexample [CCO^{+}13]. Counterexamples [BBS00, BLM^{+}22, CEOT15, Tha08]. Counting [ASH18, AFS12, ACD08, Bac09, BBRT23, Cai93, CCG^{+}11, CEFY24, DM19, DJMV21, ERSZ22, FKLL15, FGZ19, FGRZ21, GGMM21, GRS11, Hof98, KLL13, MMS15, MO22, MPP20, MR04a, MR04b, MRST16, NT12, NS16, Pal12, PRR02, PRS^{+}23, PC97, vzGVZ13, BJ92]. Counts [dJMS16]. Coupon [MW03]. Covectors [CLGH11]. Cover [ÁS09c, BYFMR10, BYHR10, BFRS16, BPRS13, BJS22, CMPS17, CF05, CFP19, CHW10, FKK21, FPT22, GSS15, HMM09, HKP22, KG98, Kra18a, Lev09, LW17a, Lub90a, MPSZ19, SL95, Yam07, Zeh17, Zha93, Zuc92, Jac92]. Cover-Decomposition [BPRS13]. Coverable [CCO^{+}13]. Coverage [AS14, BNRT17, CMPS17, CH15, HW23]. Covered [BN05, JY23, LdCKM18, ZGH^{+}23]. Covering [BKK16, BEGK22, BGS17, BGH^{+}17, BR17b, CKMU14, CL20a, CCHZ13, CH01, DFG^{+}21, EM20, FRZ16, Fan92, FKK05, FKK07, FMV22, FGLS18, FXZ23, GKR^{+}18, Gu18, Hof95, Hof98, HK02, HC98, Jan00, Jon05, KMR11, KMMS21, KCL98, jKR21, LL99, Lee17, MMPS10, Pry23, SC17, Wan02b, WY20, Yus24, Zha94, BS88, CFGG88, Fra89, GHHY96]. Coverings [FKKL98, FL00, HHLÖ95, KCL98, KML05, Bou97]. Covers [CHKW24, CL16, DZ09a, DHS14, Dre12, JWF05, KKL^{+}10, KKN95, KS03b, LHLZ23, Luk20, MRS19, MS21a, MHLHL91, Muñ05, MRS89]. Coxeter [Lub90b, Mar09, Pet13, Sim13]. Creating [KS12b]. Creation [ADHL13, ADH^{+}14, BFLM24]. Crick [BDT17]. Crime [RT09]. Criteria [KM21a, Mur96a]. Criterion [DP15, SP88]. Critical [Age94, De 24, DL14, DL17, DL18c, DL18b, DM23, GRSZ23, GL10, Har18, JLD^{+}18, iKKKL09, LM12, Zha90]. Cross [BJK14, Bor10, KS05, KG93]. Cross-Composition [BJK14]. Cross-Intersecting [Bor10]. Cross-Intersections [KS05]. Crosscorrelation [CCD00]. Crosses [BE13]. Crossing [AFT12, AADM18, BGS17, BFM06, BK07, Chi11, HT18a, KM02, OZ18, TT07, dKMP^{+}06, dKPS13]. Crossings [CSW17, DEW17, DLS11, EM99, EHLP11, Kla06, PSŠ10, PY09]. Crux [HHK^{+}22]. Cryptographic [TKMM19]. Cryptography [BDDS03]. CSPs [BM22, BZ20, TZ15]. Cube [FFHJ94, RKDD13, Rif99]. Cubes [AS02, Bah23, KM02, Shp22]. Cubic [AC14, ABHM00, BB16, Bev10, BZ11, BIT13, CCOY17, CL15a, CL16, Cap03, CK08a, CFLZ19, CLS15, DL18a, KSS09, LM10, Luk20, Mác13, MS21a, MS21b, NT24, PRR02, PV22, KKV11]. Cubical [ABY14, BGJW21, Cha91]. Cuboids [BCM^{+}12]. Curvature [LMP19, OV12, Yos19]. Curve [ACM11, GS94a, GC11]. Curved [OV12]. Curves [BK12, BHTK^{+}21, CM90, CC03, RT18, Suk13, WWKY11, vBEM17, RTW97]. Cusick [DKS16, Ros09]. Cut [AHFM08, BDM^{+}24, Bih05, CKL^{+}21, CMM20, CGN06a, CQX20, Cho92a, COHK21, COLMS22, CFP16, Cor24, DZ09a, DHKM11, GKS22, GH90a, GvZ19, Gün07, GY23, HM94, IMS05, JPvL21, iKX20, LM22a, LRT08, Mic21, Sot15]. Cut-Rank [DZ09a]. Cutoff [HY89]. Cuts [AEPT24, BBRT23, BK90, BCH92, CR04, Fio06, GVW06, Har01, KŠ08a, LL17, NNI97]. Cutting [Boy96, ERSZ22]. Cutwidth [HLMP11, HvtHLN12]. Cyclability [GKMT17]. Cycle [ARTV12, AS06, ABS13, ALM^{+}18, BGL07, BCH92, CL16, Cap99, Che93, CFGJ06, CLiO18, CM03, CL21, DSZ23, DPSW08, DHJN02, FKS12, FS91, GKPP18, GGM^{+}22, GvZ19, HM94, JPvL21, KKL^{+}10, KRS11, KW13, LS15, LP21b, LHLZ23, Luk20, LLZ22, MPS06, MN18, MS21a, Mic21, NT12, NPSW23, OC19, PPR22, PY90, PSML08, RLWY21, ST17b, Zha93, BP89, BC88a, FH94, TZ97, KvIL^{+}12]. Cycle-Bicolorable [BGL07]. Cycles [AS06, ABHJ21, AHH21, AF10, AK22a, AB18, AF19, ARS17, APS22, BM19, BF12, BCKN21, dOBMS^{+}17, BSKS11, BBC^{+}23, BG11, BY08, BCF^{+}20, CdVL11, CFH23, CF08, CGH^{+}15, CFLZ19, CKKO10, CPPT20, CGK94, Con05, CFK10, CY03, CHHM09, CM14, DKM^{+}15, DX19, DX24, Dre12, Duc21, Dvo05, DLŠ10, DL14, DL17, EIiK08, EFMN08, EGM18, EFK18, Fan92, FJ09, Far09, Fed01, FKMS20, FKP15, FHLT21, FKT99, GM20, GHM21, GJXZ20, GKPP18, GK13a, GKMS22, GHM14, GkoKK23, GLS20, HW15, HHK^{+}22, HZ95, HY12, Hng22, HH13, HM19, IKK^{+}22, Jan22, JY23, JKSW17, KŠ08a, KLN10, KSS12, KS12a, KO06b, LSS17, LLY10, LL17, LPS09, Lic14, LZ05b, LY21, LQ22, LMRS17, LMSZ19, MT90, NV22, OS17, PZ05, PS20, PP90, RS14, Raz20, Špa07, Ste22, Tan21, Wan02b, Wan08]. Cycles [WX13, WWY23, Zam22, ZZW13, Zho92, Zho93, ZLWC12, Hur94, Hut88, KP06, RS93, YZ97]. Cyclic [ANP91, BGÖ17, BER11, BW02, BST23, CS14, CCD00, Din13, EHJ01, Fér15, HK16a, JS17, KML05, MWW21, MP98, NS16, QP15, Stu21, CET97, GÖ12, Jia95]. Cyclically [Ehr16, GM91, GM93]. Cyclotomic [CGSM16, Mom13, SS18a]. Cylinder [TT91]. Cylindrical [Ful14, dS91].

D [Naa01]. DAGs [Chi23, DGL11]. Data [CKN^{+}15, GJ19, GKNU10, Kao96, SV11, Tam88]. Databases [AA11, BHM00]. Dates [GQS^{+}02]. Davenport [Pet11, Pet15]. Deadlock [Lyn94]. Decay [KM11]. Decidable [BJ23]. Deciding [ACFL16, BI13, CGG17, HT18a]. Decimation [COPP12]. Decimations [BCPP09, CK11, GKMS04, TQ09]. Decision [LNNW95]. Decisions [WS12]. Decodability [CM05a]. Decoders [AM95]. Decoding [BZ04, CR23, EH13, Han98, KRR16, RR03]. Decomposable [CDGO22]. Decompose [KN22]. Decomposing [BH97, GJK^{+}23, dH89]. Decomposition [BCdMR08, BP12, BPRS13, BPT91, CKMU14, CdMR12, Che17, CP16a, CD18, DH20a, Erd17, EMT15, FL10, Gab04, Gij05, HSV23, HKP^{+}17a, IM96, KT19, LLZ22, MWvZ11, PP13, Sch09, SWKP10, JO95, Spi89]. Decompositions [BDF^{+}18, BAM16, Cap99, CGK^{+}19, CS09, CK99, CLY05a, Dro16, DJM^{+}18a, DJM^{+}18b, Erd20, FM24b, JKiO21, KiT13, Köp07, Lai08, MPS06, MS17b, OS13b, Shp13, Spi19]. Decycling [PZ05]. Deep [DMR11, Pip06]. Defective [BLM20]. defectives [CHW88]. Define [Fio06, HMS05]. Defined [AKP20, ADL^{+}09, Bón09, Gij05, GM91, GM93]. Defining [BS15b]. Definite [Tro15]. Degeneracy [BJSS22, FGL^{+}20]. Degenerate [GKLT19, JL23, KMŠ^{+}09, KMS23, KNP05, Sha23]. Degree [AKP^{+}22, ABC^{+}15, AK22b, AM96, ASSW21, dOBMS^{+}17, BC94, BHLP92, BPSS19, BPSS20, BBC^{+}23, Bon08, BH22, BCG22, BMN13, BGS23, CCJS22, CKKO10, COL10, CFK10, CR19, CNG19, DLMT21, DŚN21, DLMO18, DFJS15, DP92, DF04, Dro16, DH20a, DK14, ERS19, EMT15, FGPS19, FJP22, FT20, FTS24, FZ22, GGHP22, GP18, GKY06, HPS09, HRS18, HN15, HZ10, HK16b, HKK^{+}09, HOD22, HLT19, JN16, Joo16, KKL^{+}10, Kan08, KSS11b, KPP13, Kha13, KL14, KW08b, KR23, KM13b, KSŠ08, KKS17, LS95a, LM14, LR04, LM11, LXZZ08, OPVV14, OC19, Pfe15, PVUY18, PS21, Shp15, SE14, ST17b, Ste10, Ste22, UV15, WH15, Wan23, WWY23, Yus14, Zak14, ZZBL17, BLP21, BCLR89, Car94, CHO^{+}89, LP88]. Degree-Constrained [LS95a]. Degree-Diameter [DŚN21, DF04]. Degree-Doubling [KM13b]. Degree-Restricted [BC94]. Degrees [ABC^{+}23, BBLM13, CL06, CDK10, FP22, GM13, Hng22, IT09, Oze13, STV21, Wil99]. Delannoy [AS03]. Delaunay [Str20]. Delay [LS95a]. Delaying [CDP94]. Delays [AB05]. Deletion [CGL10, FSV13, GLSS16, GPR^{+}21, JP18, PRS18, Sno13]. Deletion-Correcting [CGL10]. Deletions [ABY11, CGHL23]. Delineation [DO08]. Delivery [BYKKR18, HV00]. Delsarte [AL07]. Delta [AP18, BC95, BH13b, FF06, KT14, Tak14]. Delta-Matroid [FF06, KT14].

Delta-Matroids [BC95, BH13b, Tak14]. Demand [Che04, Myu01, NS89]. Demands [CEP18, Che04]. Democratic [May96]. Dense [AHP19, AHH21, AHDHL22, AZ22, AK22a, BD20a, CFP19, DHJ^{+}13, DKT22, FHLT21, GKR^{+}18, HW15, JN17, MP14, McC21, McK19, PS22, Sol06]. Denser [Nel15]. Densest [BT96, CDK^{+}18]. Densities [BBBZ12, DS21a, FXZ23, Sha23]. Density [BPST22, BHL05, CF09, CFS20, DM23, Hua14, LM14, LZ18a, LZ09, Meh12, MMP13, Ric14, Sid18, Suk13, YP22, ZG21, WS96]. Depend [Kra07]. Dependence [BPEY22, DM03]. Dependencies [BBG08, SB10]. Dependent [Ete22, FS12a, JL23]. Depending [AB95]. Depth [AB94, BJS22, DMS14, HBDS23, KW90, RTS00, RJS93, SR94]. Depth- [AB94]. Depth-Two [RTS00]. Derangement [CX08, MS14]. derangements [FZ88, Zen90, Zen90]. Derivation [EW19]. derivatives [Car94]. Derived [DF02, HPS96]. Deriving [AMB11]. Descents [HMZ21]. Description [ART14, Wal19]. Descriptions [May08]. Design [ACLW18, BLMA^{+}08, BKM08, BH93, CMV10, CS09, CKN^{+}15, CV07, FL92, GM90, HKS07, IM96, KNS05, SW99, DM88]. Design-Variable [IM96]. Designing [BAH10]. Designs [BS24a, BCS04, BV10, BRS09, CM90, GG11, GL08, GHM10, GH22, Jan00, Kas05, LS16, MMJF03, Ran02, Zha94, ZGL^{+}09, RCS88]. Desire [dAHFdFK10]. Detachment [Fle05]. Detachments [JS03]. Detecting [CK08b, FKLL15, KLL13]. Detection [BS91, CHW88, KG93]. Determinant [CHX15, LN24, MV99]. Determinants [CV09, EFMN08, GGMM21, KY21]. Determination [AJM08, LCV03]. Determine [CM18, PS22]. Determined [Pin08]. Determining [BN01, LS89]. Deterministic [BMS12, BBS00, BPT91, DKM^{+}12, FKK18, GMPZ15, HV00, JPvL21, KP04, Lu04, SYKY18, Som14]. Detours [BCDF19]. Deviation [DK22, Yam20, WW91]. device [CHW88]. DFS [Cai93]. Diagnosis [SW01]. Diagonal [OS17]. Diagonals [PR98, EFF91]. Diagram [FG22, dAHFdFK10]. Diagrams [ANP91, BG12, DFG^{+}21, DEH20, KvRRS23, MRAS19, Mor94, Rea15, Saw02, Stu88]. Dial [BD20b]. Dial-a-Ride [BD20b]. Diameter [BCS04, BDF^{+}18, BC09a, BFH15, CvBdJdVK22, Cap03, CFM94, CKdAHdF13, CLST12, DPR22, DŚN21, DS21b, DF04, FFV11, FR99, FM11, FFie98, FK18a, HKM20, HP21, KPS19, LLM19, LY10, Mer99, RC98, RKDD13, SE14, Tod14, BC88a, Zho88]. Diameter-Reducing [HP21]. Diameters [BGL24, NŠ07, Ram04]. Diametral [BM13]. Diamond [AG15, MTR14]. Diamond-Free [AG15, MTR14]. Dice [ABFRV24]. Dichotomies [WS12]. Dichotomy [ASS17, Fed06, HLS22, HR12a, MW20a, Sig10]. Dickson [QD14]. DICUT [AHS01]. Difference [CCG05, Fan20, FK18b, KT99a, Mom13, NYKY20, FMRR88, Ste88]. Differences [Sav90]. Different [DF10, HY12, KMS08, LN21, Tan21, KS88, KMS12]. Differential [LZ06]. Difficult [BFR12]. Difficulty [HM20]. Diffie [CY08, FS01]. Diffusion [BGM22]. Diffusion-Limited [BGM22]. Diffusions [FPR18]. Digirth [LM17]. Digit [MS05]. Digital [AF05, CK18, DM13, FL15]. Digits [DKS16]. Digraph [GSPRM91, Gut93, KLMR13, Ram04, WWY23]. Digraphs [ACZ23, APS^{+}24, BM19, BJT92, BJHY03, BJSS22, BPS18, CH13, CH17, CGK^{+}19, CL13, CY18, CKKO10, CM03, DKM^{+}15, DS16, EFK14, ES98, FMMO93, FHM03, FIN98, FLM^{+}95, GSL98, GGI07, GŻ10, Gui98, GRY08, HM11, HR12a, HR12b, HHMR20, HY12, HPS19, LS18, LM17, LPS09, Mac91, OS92, OS13b, PMM98, SS94, Ste22, Tan10, ZZW13, ZZBL17, BJHM88, IP91, Jia95]. Dijoins [ACZ23]. Dilate [Sha20]. Dilates [RN21]. Dilworth [iT12]. Dimension [ABC11, ABC^{+}15, AE03, BGN15, BDF^{+}18, BFGR17, BLL^{+}15, BT93, CHM^{+}07, Cai18, CH23, DFG^{+}21, FHM94, Fel14, FMP17, GKI^{+}23, Gly10, Hed08, HWS18, JMW17, KT24, Kra18b, LCV03, PZ98, SGM20, ST23, SS10, BT97, Sen97]. Dimension-Free [Kra18b]. Dimension-Normalized [AE03]. Dimensional [Ale10, ABZ15, BP15, Bra05, CK18, CC07, DDS16, FKT06, FM11, GPP04, GW00, GMR^{+}21, MW19a, MR15a, MW20b, NH91, SBD^{+}19, Zun11, HM20]. dimensioned [Tod89]. Dimensions [AC14, AK14, ANP91, AS05, Bar01, CP22, DNR23, KKW17, KM21a, LSW18, RS16, SS95, VVY15]. Diophantine [Sis21]. Dirac [CS19, CP20b, Lo14]. Direct [AAHLT10, BIKZ05, OS22, WZ18, WY23]. Directed [AFG^{+}16, AFG^{+}09, AB18, ABHW13, BCS04, BW99, BMP13, CLMS22, CMM20, CF08, CEP18, CY18, Cho92a, CGK94, Erd20, FFie98, GHM10, GMS15, GNS11, Han19a, HN15, Hua14, JLR^{+}17, KiK12, KvIL^{+}12, KNS05, KT99b, KPPW15, KM05a, LL17, Lic14, MPRS22, Sli10, Yus24, ZZW13, Bal89]. Directional [ATPRU91, AB00]. Directions [DMN12, SW99]. Dirichlet [CW92]. Discharging [BDDP24]. Disconnecting [GS00]. Discordant [CDFR18]. Discovery [FKW10, LRS23]. Discrepancies [DF10, GDCM20]. Discrepancy [FKK18, Kun23, LN24, MMPPG23]. Discrete [AF10, AG06, AG18, AAMM24, BGJW21, BL09, BILL23, BL90, BHH94, Brä10, CHX15, CIN18, CL22, CCG^{+}00, DFJS15, DGP06, DGM12, GM93, Gor93, GLW11, HT13, HI23, IM96, KLR22, KVZ23, OV12, Rio98, Sca05, SS02b, SB10, Vin07, vD11, Con89, HR88]. Discrimination [Vaz13]. Discriminator [BM22]. Discursive [Fer24]. Disentangling [Sul12]. Disjoint [AK22a, ARS17, BM19, BNRZ23, CGH^{+}15, CFLZ19, Chi23, CFH16, Con05, DSS92, EJ01, FIN98, Ful14, GM20, GPvL15, HKL11, HY12, HMM^{+}21, HMP04, KiKK17, iKK11, LLL17, LPS09, Lic14, LW20, LMW23, MG19, MMSW23, MHLHL91, NY21, Sch91b, Sli10, Ste22, Tan21, WWY23, Yus09, Yus14, HR88]. Disjointness [ALM^{+}18, SW14]. Disk [GC11, JSRSW18, Sol12]. Disks [AKP20, FS05, Ric14]. Dispersers [RTS00]. Dispersion [JPT12]. Dissections [CMR18, DST01, Alt89]. Dissolution [vBBC^{+}15]. Distance [AKP^{+}22, AGQS23, ANS16, BS24b, BDD^{+}19, BCE05, Che98, CE06, CKdAHdF13, DMS08b, DMS08a, DOS94, DX19, FG01, GP08a, GMS00, GM03, GJ06, GÖ12, HKL^{+}21, HHH^{+}02, IK09, JK24, KN16, KL19a, LZZ22, LL99, LLM19, Len98, Li17, LZ05a, LZ05b, NLJM21, NS11, Pon20, Sak94, Vin11, WL03, Wil99, WC12, XGG15, XG20, Yu17, ZLWC12, FGK89, GY92, PS97]. Distance-Hereditary [DOS94, HHH^{+}02, KN16]. Distance-Increasing [WC12]. Distance-Preserving [BDD^{+}19, Che98]. Distance-Regular [FG01]. Distance-Uniform [LLM19]. Distances [ABR05, CR16, CP22, CF09, GJ12, HKP22, KNZ14, Lab13, Owe11]. Distant [CtJL01]. Distinct [ASS09, CFG^{+}15b, DFX21, GWZ18, Jev95, Pin08, Tak90, KP06]. Distinctness [TQ09, RSSW88]. Distinguished [HK15, NÖ08]. Distinguishes [BKL^{+}22, Erd17]. Distinguishing [ACD08, BGLS07, CHK10, JL20a, Prz13, WH15]. Distorted [GC11]. Distortion [ABFRV24, SBD^{+}19]. Distributed [DM18a, HKP01, HJ94]. Distribution [Bón09, CHP^{+}90, CFKK17, CP22, ERS19, FS01, FPS18, LS09, RJS93, RF12, Shp10, Ste88, WS17b, HM88].

Distributions [AAMM24, CY08, COL10, COHK21, Gao13, Gar92, GMTW15, KR04, Orl93, CZ97]. Distributive [ADL^{+}09, Nao91]. Divergence [FTS24]. Divide [ARS95, ÁS09c]. Divide-and-Conquer [ARS95, ÁS09c]. Divisibility [BMVW17, CCD00]. Divisible [GG11]. Division [AFP^{+}18, BILS22, BCC^{+}05, Bon91, MS21c]. Divisors [Kle89, NÖ08, Car94]. DLP [BN01]. DM [BIKY18]. DM-Irreducible [BIKY18]. DNF [SST08]. Do [ARTV12, MS20, NS22, TV03, CM18, KPT95]. Does [GC11, Kra07, MM15]. Doignon [CHZ18]. Domains [Das99, FK97]. Domatic [LHC90, SV08, MM96]. Domes [Gra04]. Dominants [CFP16]. Dominated [Tan10]. Dominating [CCOY17, DK02, EKM^{+}19, HK93, KLMN14, KPS20, LPS18, PP10]. Domination [AS09a, BM16, BKR10, BHT16, CWY00, CKPW09, Che98, DMKŠ08, DHL^{+}13, Fis94, GM16, GLS15, GPRT11, HHHH02, HY10, HK16b, KMR^{+}23, KWZ13, KS03b, KS93, LC12, Tuz08, BG88, BCD97]. Don’t [ABFRV24]. Double [AFK12, AG18, CY08, CFS17, HY10, KS08c, LHLZ23]. Double-Resolution [AG18]. Double-split [AFK12]. Doubles [NT21]. Doubling [BGN15, GK13b, GL10, KK10, KM13b, Lev22]. Doubling-Critical [GL10]. Doubly [NvdP18]. Dowling [Got03]. Down [DZ09b]. Downset [FM11]. Drawing [KPP13, Žit94]. Drawings [AFT12, ARS21, Ful14, HH17, HIMW24, ZN08, ÁAFM^{+}18, Tod89]. Drawn [BBS17]. Dream [Més16b]. Drop [BCG^{+}10]. Dual [AC90, BOP94, BYR05, BBKS23, BM22, BFH15, CCM95, CCZ12, CMSV17, DZ09a, FMV22, HK11, Jan00, LL99, MV18, MAR22, MAR23, PG06, PdCSC21, dCST20, UV15, VV94, BGM94]. Dualities [ETT13, Hir11]. Duality [CGM^{+}15, CDHZ12, DEE17, LT11, NT05, Rho15]. Dualization [Elb09, YS95]. Dually [BDCV98, BCD97]. Duals [Moi08, NŠ07]. Duplication [FTS24, LZZ22]. Dyadic [ACGT22, ACP22, CMR18, CGOvZ21]. Dyck [AG15, Fer16]. Dyck-Paths [AG15]. Dynamic [AB05, AH21, ABL^{+}20, BBKS23, GPST15, HKL^{+}14, IW91, Juk16, LNO96, Ram04, SS00, dBSS24, Tam88]. Dynamical [Gad18]. Dynamics [FGKP22, LM11, OS11, RSV^{+}14, Ull14].

Each [CY21, Lin97]. Ear [CK99, Gab04, Sch09]. Early [KM23, WDSH21]. Easier [AH21]. Easy [BAH10, CMV10, DRW98, Vaz13]. Economic [vWW94]. Eden [MM93]. Edge [AJM08, AK22a, AS10, ASS17, BJKV07, BGLS07, BJFJ95, BJJ98, BJGJS99, BS09, BCC^{+}11, BG20, BL17b, BCC^{+}22, BMR^{+}10, BCF^{+}20, CH13, CH17, CLS09, CGK^{+}19, CSS01, Chi23, Cho92b, CM18, CMM^{+}10, Con05, CF22, CM07, CW09, Cra19, DJKKV23, DMS08b, DMS08a, DJKO19, DS06, EJ01, EFK18, Fle05, Fra92, FPS18, Gab04, Gab05b, GS94a, GK07, GPR^{+}21, Goe01, GPW09, HMS21, HK93, HT18b, HMM^{+}21, HMP04, JMSW99, JMSW00, Jor03, JS03, JL20a, KŠ08a, KiKK17, KSS11b, KMMS21, Lai05, LSS17, Lev15, LLL17, Lo14, LW20, LMW23, MMSW23, Mác13, MW14, McL10, Meh12, MLS11, NI92, NK90, OC10, Prz13, Riz02, SS11a, Sch91b, SST22, Sim13, Sli10, Ste10, WL02, Yus14, dMP93, BM97, Bon15, Cho94b, Jac92, BFH21]. Edge-Bandwidth [AJM08, JMSW99, JMSW00]. Edge-Bipartite [Sim13]. Edge-Chromatic [dMP93]. Edge-Colored [BS09, BCF^{+}20, CH17, CGK^{+}19, Lo14, CH13]. Edge-Coloring [BCC^{+}11, Cra19, NK90, Riz02]. Edge-Colorings [BGLS07]. Edge-Connectivity [BJFJ95, BJJ98, BJGJS99, DMS08b, DMS08a, Fle05, Fra92, JS03, NI92, OC10]. Edge-Cuts [KŠ08a]. Edge-Disjoint [Chi23, Con05, EJ01, HMM^{+}21, HMP04, KiKK17, LW20, LMW23, Sch91b, Sli10, Yus14]. Edge-Face [CLS09, KSS11b]. Edge-Homogeneous [GPW09]. Edge-Independent [HT18b]. Edge-Injective [BMR^{+}10]. Edge-Isoperimetric [Lev15]. Edge-Labeling [Meh12]. Edge-Markovian [CMM^{+}10]. Edge-Partitioning [MLS11]. Edge-Robust [Lai05]. Edge-Splitting [Jor03]. Edge-Surjective [BMR^{+}10]. Edge-Weights [BJKV07]. Edges [AD11, BS10a, BGH^{+}17, CCH21, CL07, Dvo05, EIiK08, EM99, FPS13, FT17, Fri21, Ful14, HV17, HVLP13, Kla06, LS22, Luo23, NT21, PY09, ZGH^{+}23]. Edgewise [Ath14]. Edit [Lab13]. Editing [FGPS19, GKNU10]. Effect [BHH94, Fri21]. Effective [CPP23]. Efficiency [CDHZ12, GK02]. Efficient [BM16, CGHL23, COCF10, DP96, DMNW13, GOR20, GS93a, HKK^{+}09, KG93, KRR16, Myu01, PW02, TKMM19, Vaz12]. Efficiently [ABY14, BNCPR20, HHH^{+}02]. Ehrhart [ST10]. Eigenpolytope [BS24a]. Eigenspace [Iri16]. Eigenvalue [CKNV16, Iri16]. Eigenvalues [BG20, CFM94, HRS17, JiT22, OC10, Ost23, Ste07, Kah97]. Eigenvectors [LSO03]. Eight [KS23b]. Eisenberg [CDV10]. ELAs [MMPS10]. Electric [HHHH02]. Electrical [BK90, KW17]. Elegantly [EFK18]. Element [BM22, CK14, RSSW88]. Element-Connectivity [CK14]. Elementary [MR04a, MR04b, SS08]. Elements [Che07, KS23b, Sav14, Jed93, JLM93]. Elimination [AKP^{+}22, BKS09, Che98, HKP22, KSW17, Way01, Vav89]. Elliptic [ACM11, CM90, CF17]. Embeddability [DM15, HMM09]. Embedded [AD96, CdVL11, CCH14, DH20b, PP90, Car88, Hut88]. Embedding [BPSS19, BPSS20, BHT10, CGN^{+}06b, CNG19, EP10, EM99, Gol96, HKP^{+}17d, Hor14, Kri10, Moh99, MW90, NOO12, Obr93, ONN19, JM97]. Embeddings [AS02, BS15a, BGM08, Ber07, BCLR95, Cai93, CK99, DGL11, EN22, KFHR94, MR15a, PSW96, SBD^{+}19, BRK89, SP88, Suz10]. Emerging [FPT22]. Empty [AHH^{+}10, BDJ^{+}15]. EMSO [AZ22]. Emulators [HP21]. Encoders [AM95, RF12]. Encoding [Gra04]. Encodings [HKL99, STT92]. Encryption [KOS16]. Endowed [BR17b]. Endpoint [LW17a]. Ends [Ste10]. Endvertices [DG08]. Energies [GP20]. Energy [BDvL13, FPS20, KK10, Mus21]. Enforcer [BM09]. engineering [Tod89]. Entanglement [AP18]. Entirely [WHW14]. Entries [LN17, Vin12]. Entropic [IMR14]. entropies [KM88]. Entropy [FGP10, MWW21, MP13, NW95]. Enumerating [EM20, FKK05, FKK07, JB22, NP18]. Enumeration [GM13, Hof95, KLMN14, KBE^{+}05, KCL98, KML05, MRR20, McL10, OPR12, Pip01, Pip02, RC18, Sav14, VZ93]. Enumerative [MPP17]. Enumerator [BK91, DM13]. Enumerators [Bar02, Kap14b]. Envy [MS20, PR20]. Envy-Free [MS20]. Envy-Freeness [PR20]. EPT [vIKL^{+}16]. EPTAS [Jan10]. Equal [CER98, Got03]. Equalities [FJZ15, KS03a]. Equals [Kao96]. Equation [Hor19, Sis21]. Equations [BILL23, KLL13]. Equiangular [Buk16, KT19, Yu17]. Equicolorable [Pip01]. Equidistribution [SS95]. Equidistributions [HMZ21]. Equilibria [AGQS23, AGM23]. Equireplicate [ACLT01]. Equitable [KNP05, LSSY10, Yus03]. Equivalence [BYR05, BK21, BK24, BCGR24, CHZ04, CW16, LGS11, PRS98]. Equivalent [Cho92a, Cho92b]. Equivariant [GXY22].

Erdos [BFH21, Sha23, DK16, AFH^{+}18, BBT16, BMPS21, BHJ18, BJS21, CSS13, CDT23, dOCHLO21, CP96, DT16, Dow88, DFX21, Han16b, HLO17, HjK24, IK09, JO18, KSY18, LMSZ19, MS14, Roz19, Vin11, XG20, vBJU20]. Erlang [FG89]. Errata [GM93]. Erratum [BK24, BLR17, DW11, Dow91a, FKK07, HWZZ20, TZ20]. Error [AG06, BZ04, BGS96, CD93, DA10, FT05b, GMZ09, GOR20, KM11, CZ97]. Error-Correcting [BGS96, CD93, FT05b, GOR20]. Errors [MMPS10]. Escape [FGLP14]. Estimate [Gol06, Hor19, WWKY11, LRN11]. Estimates [KR13]. Estimating [Ful14, HKL^{+}21, Luc98, PV10b]. Estimation [ERS19, SS02b]. Estimators [CV09]. Euclidean [BT22, DGN^{+}20, GL15, HM88, Har11, Kar89, SE14, Tas97]. Euler [FG14, IKZ08, Wu09]. Eulerian [BOP94, BPS18, Cap99, CCM95, CH13, CH17, DMNW13, FIN98, IP91, KiKK17, MS17a, MS17b, MRS19, YZ17, ZL11]. Evaluation [HKR00]. Evaluations [MR04a, MR04b]. Evans [GH22]. Evans-Style [GH22]. Evasion [DKSŻ10, IKK06, SS89]. Even [Adr23, BL09, BCPP09, CCOY17, CHZ09, CGK94, DZ09b, DQW^{+}15, DX24, GJXZ20, GB12, GKMS22, GvZ19, IT08, NNO15, PPR22, PSŠ10, Rio98, YZ97, Jed93]. Even-Cut [GvZ19]. Even-Cycle [GvZ19]. Eventown [SV18]. Events [YAT16]. Every [AcRS07, KSS11b, OZ18, Sha23, HS89b, Zho88]. Evolution [Blo10, DKS18, KM23]. Evolutionary [CHP^{+}90]. Evolving [CMM^{+}10, STT92]. Exact [ADKS18, BHL05, BDM02, Cre04, GH13, HMS93, KYDN09, KMRR09, LZ09, MSK93, PSTF00, RY22, Sch04, dJMS16]. Examples [BL16, Lut04]. Excedances [Ath14, HMZ21]. Excess [PVUY18]. Exchange [BS24b, FPP22, Kot13, KRZ21, Obr93].

Excluded [BL23, Che22, CNP21, EW19, FHJV17, FGT11, Tak22]. Excluded-Minor [CNP21]. Excluding [CNRS15, NW22, SSS22]. Exclusion [Doh03, Fér15]. Exhaustive [BHRZ14, Kas05]. Exist [CHHM09, MS20, NS22, TV03]. Existence [CL91a, CSS13, CM14, GGLS21, GL08, HZ95, LWY13, MMJF03, ÖV04, OC19, WLD09, ZGL^{+}09, RCS88]. Expanded [GCH22]. Expander [BZ04]. Expanders [HN23, Kri18, PRS^{+}23, Vin13]. Expanding [AMNV18, NSiTW18, OS13b]. Expansion [ACKM19, DN16]. Expansions [BSS14, KMV15]. Expected [BHL^{+}15, CL06, FS12a, MS19, Pip91, Pit23, PRS88]. Experimental [BLMA^{+}08]. Explicit [AB94, Cap03, CY08, DA10, DO08, GPST15, Lu08, SS02b]. Exploiting [CP16b, KMR^{+}23, KKS22]. Exponential [CY08, CHZ09, Ege10, FKL^{+}19, GRS12, IK09, Mac18, MPSS20, NPSW23, SW21, Zhu21]. Exponentially [HM12a, NvdP18, NT05, PP22]. Exponents [BZ04, FS01, ST20a]. Expressing [EGMR20]. Ext [JKSW17]. Ext-Triangular [JKSW17]. extend [RS93]. Extendability [AS06, ANP91, CFGJ06, RLWY21, VW09, ZZBL17]. Extendable [ABS13, LS15]. Extended [KPT12, MP22, dCST20]. Extendibility [LZ03]. Extending [ARS21, CDGO22, GR17, KKK^{+}12]. Extension [BYHR10, Bol90, BKTW15, DMNW13, FR99, FM11, GMOW21, NNO15, NNO19, Oxl19, Pad16, Peg14, Voi07, JM97]. Extensions [AKW05, AMPT93, BM13, Che07, GGMM21, Naa00, PR91, Sta92, Ten09, FH89]. Extinction [FS12a]. Extractors [RTS00]. Extremal [AE03, BGV07, BL19a, CDM04, CL20a, CKPW09, CL07, CFS20, DS21a, DST01, DFJS15, DNR23, DJS12, FJK^{+}23, GP17, GH13, HJ94, Jan20, Jia95, JS17, KLM14, KWZ13, LLZ99, Liu24, MT20, NW22, NT05, Ost23, Pon20, ST20a, Sid18, Siv23, WY23, ZZBL17, dLL09, BG91]. Extremality [Fue23]. Extremum [HM88].

Face [BW02, Bro11, CLS09, CN12, KSS11b, KM14, Mur10]. Face-Width [KM14]. Faces [Set21, Tza08]. Facet [CLM03, CdV02]. Facet-Inducing [CdV02]. Facets [BKG99, Chi11, DF02, Fio06, FLMY09, Pad16, Bal89]. Facial [HSŠ08, MM11]. Facility [AYZ04, BB09, CCZ12, KVZ23, RS10, SX13, Tam91]. Factor [CL16, CSS13, CDMT21, IT08, KS03d, Mur06, PSML08, Rah16, SS11a, SV08]. factorable [FH89]. Factories [NLS24]. Factoring [Rón92, Sza06]. Factorization [AAHLT10, Tro15, BGM94]. Factorizations [BD24, CL05]. Factors [AS09b, BIT13, BMN13, CL15a, CDK10, LWY13, LW17b, MP98, Riz02, Tak22, ZLS08, BH97, CY18, HKKK88, Mit97]. Fail [MMO20]. Fails [AZ22]. Failures [HMS21]. Fair [AFP^{+}18, BILS22, MS21c]. Fairly [OPS21]. Falconer [IK09, Vin11, XG20]. Families [BLP21, BKK16, BSS14, Bor10, BCF^{+}10, CLWZ22, CGK^{+}19, CLW09, DK10, EFK05, GZ19, GP08a, GLP^{+}12, Ger23, ILM^{+}16, Jev95, KL19b, KM01, KMS12, KM13b, LÖ05, MZ22, OS16, PMM98, Per16, Ran02, SG16, WZ18]. Family [Ave13, CCT05, FFLP20, LL15, MP22, MTR14, Pin14, dM07]. Fan [CH06a]. Fan-Out [CH06a]. Fano [dOCHLO21]. Fans [Ber20, CR17b, Sha13]. Far [FGL^{+}20]. Farthest [HHLÖ95]. Farthest-Off [HHLÖ95]. Fast [CR23, Dah93, DM13, EMT15, FH10, GH90b, GRS12, MP17a, Saw02, Som14, SV11, Car88]. Faster [AFL^{+}20, DPR22, Duc21, KP04, LMSZ19, Ull14, Way01, YZ97]. Fat [GZ19]. Faudree [JO18]. Fault [AS02, AU91, BCSK07, CFH16, HL10, PMM98, PL94, SX13, UBHS93].

Fault-Tolerant [AS02, BCSK07, PMM98, PL94, UBHS93]. Faults [DP96, GV92]. Faulty [CL91a, DG08]. FCC [EIH18]. FCSR [XQ06]. Feasible [GVW06, GS95, LS08]. Feedback [BBF99, BNN90, CPPW13, ENSZ00, GKI^{+}23, KvIL^{+}12, NY21, PPR22, Ko88]. Feige [AAMM24]. Fermat [CW14b, EA11, LY18a, OS11]. Ferrero [Boy01]. Few [AFT12, Bal08, BHNP16, BKKZ17, BS16a, BCF^{+}20, CH10, DL19, HT19, HVLP13, HS06, Hor14, KPP13, LQ22, NT21, OPS21, Pad16, SV20, Zam22, Stu88]. Fewer [AEPT24, BS10a]. Fibonacci [CCT05]. Field [AM21, Che07, HLP^{+}23, LS24, LSW18, MP22, Sch02a, Gor93]. Fields [BGY20, BGL03, BS90, CKP16, DQW^{+}15, IK09, KMP03, MWW20, MP13, Mat19, NvZ15, NW22, NÖ08, QP15, Rón92, Ros09, Shp13, Shp15, SW21, Vin11, Vin12, Vin13, WB90, XG20, vzGVZ13, LW88b, LRN11]. Files [Orl93]. Filling [FMV22, LSS24]. Fillings [CWYZ10]. Final [Pic14]. Find [BNRZ23]. Finding [Age94, BCDF19, BZ11, BIT13, BCKP19, CdVL11, CKP^{+}21, COL10, Dji06, DHJ^{+}13, FO08, FL96, Gut93, HIKT99, Hoà10, JKiO21, KY12, Kri18, LZ06, LM16, MM96, MT90, MGC14, NYKY20, PPR22, Riz02, SS04, SW23a, SW01, SFS09, Wan02a, YZ97, Zha90]. Finds [GS93a]. Fine [DL17, DL18c, DL18b, KLS18]. Fine-Grained [KLS18]. Finer [HKP^{+}17c, Lam20]. Finite [AM07, AMB11, Bab92, BCE10, BBS00, BFGM15, BGL03, BY08, BS90, CW92, Che07, CW16, CDFR18, CKP16, DF94, DM13, DQW^{+}15, Din06, FPP22, Gad18, GKM^{+}18, GGI07, HN23, HZ21, IK09, JS17, KM01, KCL98, LZ03, LS24, LT11, LSW18, MWW20, Mal15, MK09, MRR20, MMR06, Mat19, MW20b, PRS03, Pin14, QP15, RSW05, RS22, Rom06, Rón92, Ros09, Sca03, Sca05, Sch10, Shp13, Shp15, SW21, Sza06, Vin11, Vin12, Vin13, WB90, XG20, YAT16, vzGVZ13, FMRR88, Jia95, LS89, LW88b, LRN11, Lin89]. Firefighter [CCVZ10, LW10]. Firing [Eri96, GK16, JSZ15, Tar88]. First [BHLY08, BKS10, BKM13, CH06a, DJW12, LS17, MGC14, MW20a, Kie88, KPT95]. First-Fit [BHLY08, BKS10, BKM13, DJW12, MW20a, Kie88, KPT95]. First-Stage [CH06a]. Fit [BHLY08, BKS10, BKM13, DJW12, MW20a, Kie88, KPT95]. Fitch [SHS23]. Five [BBF^{+}23, CPPT20, CP22, CNG19, DH20c, DH20b, LSS17, Obr93]. Fixed [AKP^{+}22, ARS17, ADKS18, BJKV07, BCDF19, BV10, CI07, CFK10, CPPW13, DW10, DW11, Fed06, FJP22, FT20, GRR15, HN15, HT19, KPT22, KPPW15, LW17a, LSSZ19, PMM98, PSV08, PPTY22, RSW12, WS96]. Fixed-Parameter [AKP^{+}22, BCDF19, CPPW13, HN15, KPPW15, LSSZ19]. Fixed-Weight [RSW12]. Flag [Bro11, CN12, KN13, LN17, dM07]. Flammable [Pra13]. Flex [Sch10]. Flip [FKMS20]. Flipping [GKW19, CGG88, Sak89]. Flips [BDFP10, GKW19]. Flooding [ADL13, CMM^{+}10]. Floor [Fra10]. Floorplanning [YS95]. Flora [Fra10]. Flow [DM23, Fle00, GS16b, Gün07, IMS05, KNK93, KíP^{+}24, LLW22, LMS19, MS21b, NLS24, Ram90, Svi03, TH11]. Flow-Based [NLS24]. Flows [CLLZ18, CDW07, CL20b, DKLW24, EdJdVLT18, FZ08, FO00, GR99, Gün07, HLL^{+}21, KR16, Koc98, KK09, LHLZ23, LLS^{+}20, LXZZ08, MS17a, MMO20, WCLZ15, WYZZ14, ALZ96, YCH97]. FO [AZ22]. Fold [CH10, JLR20, OS15]. Folder [Hir11]. Folkman [HRS18, LL15, Lu08]. Forbidden [AFK12, CFP16, DMS21, FXYY14, FH21, FPS20, FLM^{+}16, FW20, FM13, FK21, HH04, HLS22, HLS24, Hav19, LiO15, Let19, PSWX24, PP07b, Raz10, ST17a, Tuz08].

Force [FM13]. Forced [dOBMS^{+}17].

Forcing [Dan01, GkoKK23, KKS19, KM14, ST17b, SZ13]. Ford [HS88]. Ford-orderable [HS88]. Forest [Cha19, CY21, CKN13, HSV23]. Forests [AT90, BK14, CK14, KMR11, Tak90, Tak14, vIKLS14]. Form [Jev95, WS17b, Exo89]. Formal [ASMF10, BJ91, MP95]. Forms [BCE^{+}00, CD93, CS14]. Formula [CF17, Han09, NW95]. Formulae [Cre04, Lla06, PRS02, Sto12]. Formulas [Bac09, FFV11, HS06, MPP17, MSK93, MNPR17]. Formulation [CKNZ05]. Formulations [KPT12, dCST20]. Forward [OS92]. Forwarding [Saa93]. Foundations [BL16]. Four [AS05, HT18b, HMS05, LGS11, LM17, San96, Vin13]. Four-Variable [Vin13]. Fourier [BBMM09, Car88, DDS16, IK09, Mal15, Sca03, TQLT13]. FPT [CLMS22, HW23, KMR^{+}23]. FPT-Algorithms [HW23]. FPTASes [AH21]. Fractals [FHNN18]. Fraction [KKS10]. Fractional [ACGT22, ACP22, Bar04, Dro16, DH20c, EK13, Fis94, Fle00, FO00, GKLT19, HZ10, HP23, KKN95, KKS10, KKV11, KLP12, KKK^{+}12, Liu14, PS21, XZ22, CFGG88]. Fractionally [DKLW24, KiK12]. Fractions [Het14, HKW15]. Fragile [CMvZW16]. Frame [FMOS20, Fra95, GP20]. Frameproof [Bla03, SW98]. Frames [BHJ18, CW16]. Framework [BC23a, BDK^{+}24, Mur06, SB10, Tak22, WW22]. Frameworks [BSiT15, CW96, CMM^{+}24, FSW13, FRW12, JNiT21, KiT13, NOP12, NOP14, NSiTW18, SiT15, SST22, Whi88]. Free [AP18, AAD^{+}18, AKP20, AG15, ADKS18, Ave13, BS10a, BCL^{+}23, BFK^{+}12, BKP21, BLM10, BM16, BKKM99, BS10b, BS16c, BF17, CKL^{+}21, CKP13b, Cho94a, CS18c, CS18b, CGSZ20, CNST24, Cib13, CD16, CF10, CKOS06, CD14, CR13, CR17a, DI22, DK10, DJ11, DPRS10, DNB99, DLŠ10, DL17, DM17, DL18c, DL18b, EK13, EMOT16, Eti20, Fer24, FGRZ21, FL96, FKMS10, GH19, GPvL15, GKL99, HJ18, Hav19, HHL24, HKW22, HKW15, HP23, HGY20, HT93, HjK24, IKM99, JW18, Juk21, KM19a, Kas05, KS08c, Kra18b, KOT16, LM08, Liu14, MTV08, MM12, Mak07, MS20, MTR14, MD11, NN22, OS16, PPR22, PRR02, Pen12, PT94, RY91, SHS23, Sch02b, WW18, dFM04, ASS17, BH97, COS97, Pic14, Ram97b, Ram98, Spi95]. Freeness [AKKR08, BBJ^{+}21, PR20]. French [Zen90]. Frequencies [Nag10]. frequency [Ray94]. Friendly [Mon15]. Frieze [DKM^{+}12]. Frobenius [AOW15]. Front [Kim11]. Fugitive [RT11]. Fuglede [Zha23]. Fukuda [RC18]. Full [Din06, FR06, Hyu10, ÖV04, TV03]. Full-Rank [ÖV04, TV03]. Fullerene [FKS12, Kar20, YQZ09]. Fullerenes [Gra04, KSS11a]. Fully [HKL^{+}14, HMM^{+}21]. Function [AFH^{+}18, APdMOZ23, BD20a, BGS17, BS23, Cre04, CS21, EFMW18, FI05, GSWW92, GS13, GJ16, Har19, HKMY22, Jun12, KY21, KG98, LL99, LS09, Loe10, Lou10, MR04a, MR04b, NW22, NÖ08, Osh21, Sch02a, Yos19, FV97]. Functionals [GM91, GM93]. Functions [AH96, Ale10, ABZ15, AL95, AL17, Bac09, BBK^{+}16, BBDK00, BBM24, BSS14, BGS96, BT18, BB03, BCL^{+}18, BBMM09, BFH^{+}08, Bru90, BG19, CCD00, CW92, ÇMP15, CH15, CK08a, CS18a, CEFY24, CC03, CH19, DvW18, Dei15, EMRPS14, FO90, FLM22, FiT14, GSPRM91, GVKSS06, GSS14, GM91, Gru17, HKR00, HK14, HR05b, JMS90, KMÖ18, KMT07, KL08, KCTR13, LMNS10, Lev15, Lla06, Mur06, NW95, PZ98, STV21, SS02b, TQLT13, TKMM19, WS19, Yok19, Zun11, dGV05, GM93, HSLd88, Lin97, Tsa96]. fundamental [YH88]. Further [HVW07, KMR^{+}23, Ray94].

Gabriel [BDEK06]. Gain [EHV18]. Gale [CDV10, Stu88]. Gallai [BL19a, HM11]. Galois [KCTR13, LÖ05, Rón92]. Game [BKR10, BHT16, CCPP14, DEF19, DS05b, FS05, Fei10, Fra10, FHL13b, GG15, Han19b, HK16b, JSZ15, KWZ13, KLW89, Tar88]. Games [AGQS23, ABS10, AEFT13, ADHL13, ADH^{+}14, BM09, BMS12, Bil03, BFLM24, Bol90, BHKL08, CDR16, CGV^{+}14, CCNV11, CFG^{+}15a, CHMP21, Eri96, EFKP15, FK98, Gad18, HKSS08, Jon20, KMN23, Vaz12, HR88]. Gamete [LGS11]. Gamma [DJKKV23]. Gamma-Vector [DJKKV23]. Gantt [GW00]. Gap [BCD21, BDG^{+}17, FMP17, GW19, LN24, RY22, Sha20]. Gap-Hamming [RY22]. Gaps [CL21, HMM09, LR07, Lon21, MS21c, Sul05]. Garden [MM93]. Gates [Hås94]. Gaussian [Vav89]. Gaussians [RSD17]. Gelfand [LMS19]. Gene [SHS23]. Genera [iKSZ04]. General [ART14, AKKR08, AH11, Blo10, Bon91, COL10, CPS08, DS05a, DSS92, FS09a, FP01, GS95, GK16, LN21, LWY13, Mar20, MP22, Mur06, PW13, PR20, WW22, LRN11, BST20, DL18c]. Generalization [ACM^{+}18, BC11, BCC^{+}19, CuKŠ07, GG11, Han09, Kam22, KM05a, SS94, Tod89]. Generalizations [AMW00, AFP^{+}18, BS90, GS13, Pet15]. Generalized [AB00, AS03, BMNP22, BYHR10, BAM16, BKM15, DHL^{+}13, Fan20, FKS05a, FT05b, Fuk16, Jan22, JLD^{+}18, Kar92, KY21, KK90, Lee17, Len98, Lev15, LWW10, LY23, MSD19, MŠT09, MUWY18, PRS18, Sch92, SL95, TW21, TW12, Tza08, VV94, FG89, LB09]. Generalizing [LGS11]. Generated [BFGM15, Shp15, Web08]. Generating [AMNV18, BBK^{+}16, FO90, FLMY09, GSWW92, GSS14, PR91, Sav90, Saw02]. Generation [AH21, BBJ^{+}21, CHM^{+}23, GDVL17, GVW06, Kas05]. Generator [EA11]. Generators [GMA15]. Generic [SvM08, iT12]. Generically [GGLS21, NOP14]. Genetic [GVKSS06]. Genomic [BH13a]. Genus [AF05, AD96, BELP21, Che94, DHT06, DEK22, GMTW15, Neu24, RT18, ST10]. Geodesic [Gra04, KvRRS23, Owe11]. Geodesics [CM18]. Geodetic [BDPR18, CPRdS13]. Geometric [ACLW18, Adr23, AFK^{+}22, BNRZ23, BFLM24, CTU14, CFM^{+}09, CS98, DMP07, DSST13, EMRPS14, GM16, JDT13, KM21a, LRS23, MP17b, RS16, SSSU89, ST23, Smo07]. Geometries [KT24, NO09, NvdP24]. Geometry [AM07, BCE10, BC09b, DDS16, FFie98, KLR22, Nel15, NW22, Onn91, Sis21]. Gessel [Bón09]. Giant [CL06, FPT22, Kra07]. Girth [AA10, BGV07, Bal08, BBF^{+}23, Cha03, CH06b, DŠV08, DH20c, DH20b, KP16, KLMR13, KO06a, Lub90b, PS21, WL03, WY10, KKV11]. Given [AHS01, CGHW22, CL06, CKNV16, GHM21, GM13, HW15, KW08b, KM94, LM10, MR04a, MR04b, MS19, RMS01, Yus24, ZL11, DK89, FGK89]. Glauber [LM11, RSV^{+}14]. Global [CL11, CPS08, CMM^{+}24, GJ20, GK97, GMR^{+}21, NÖ08, Tre04]. Globally [KM22, MP21]. Go [Che16]. Goethals [Ran02]. Going [FGL^{+}20]. Golf [DTW03]. Gomory [ACD^{+}13, CDD^{+}15, NS22]. Gonality [ADM^{+}21, Hen18]. Gons [AHH^{+}10, EFF91]. Good [LCV03, Meh12, TKMM19]. Goodey [Kar20]. Goodness [Mor21, PS20]. Goods [OPS21]. Goresky [ACM^{+}18, CK11]. Gossiping [CGP98, DP92, DP96, FP01, FP04, Ser88]. Gowers [BC11]. Gradient [Hor19]. Graham [CM05b, WIS12]. Grained [KLS18]. Gram [Sim13]. Graph [AS06, AC14, ALQ^{+}24, AADM18, AM21, AL07, ASS09, AHFM08, ACFL16, AcRS07, ABL^{+}20, BD20a, BDF^{+}18, BDK^{+}21, BST14, BCD21, BP10, BGS17, BDD^{+}19, BDJ^{+}20, BKP21, BPT91, BELP21, BNR96, CCH14, CW98, CDM00, Cha03, CR04, CHJ^{+}04, CFG^{+}21, CK14, CK99, CSS13, CNR89, Cho94a, CKP^{+}21, CFM94, CL06, CKNV16, CTW93, CK08c, CDHK16, CHW10, CD11, DR24, DSS92, DE93, DW22, DM11, DHKM11, DFK^{+}11, DHJ^{+}13, DEJ^{+}20, DM18b, DŠV08, DKT22, EGR08, EEFH21, EJK^{+}09, FO08, FKKL98, FKK05, FKK07, FH21, FJLS03, FKPR05, Fin09, Fis94, FG00, FGPS19, FH10, FJ17, GMP22, GP08a, GKPP18, GM04, GT13, GPR^{+}21, GS00, GLS15, GN08, GH90a, GKMT17, GKY06, GJ06, GKNU10, GK04, HK99, HW15, HVLP13, HW24, HS10, HT18a, Hoà10, HK96, Hof98]. Graph [HKL^{+}21, HLWY23, HOD22, Jan20, JLD^{+}18, JN17, KMR11, KSS11b, KM23, Kao96, KW13, KPT94, KíP^{+}24, KKS22, KY12, Kor18, KMS08, KMS12, KM13b, KŠ08b, KM05a, KM11, KCL98, LW10, LM12, LLS04, LS08, LMP19, Liu24, LW17b, ŁR91, MS21b, MS14, Mer99, MC06, Mra17, Mur06, Naa00, New20, Ost23, OZ18, PRS^{+}23, PY90, PSML08, Rab06, RD11, RT09, RZ05, Rom06, SS05, ST13, SL96, Ste07, SWKP10, Wag07, Wan02b, WS17a, WY10, WW18, YS95, Yus24, Zha09, Zha11, ZGH^{+}23, dCLM13, Bal89, BVW88, BB97, GM94, HS89a, KW96, MM96, RW89, SP88, Spi89, Zho88, dH89, DMS21, LB09]. Graph-Based [GKNU10]. Graph-Coloring [HK96]. Graph-Different [KMS08, KMS12]. Graph-Theoretic [KM05a]. Graph-Theoretical [Wag07]. Graph-TSP [New20]. Graphic [FS09a, GS16b, KMPR14, PP13, Wag18, Che22]. Graphical [BS24a, CR96, FGP10, MUWY18]. Graphicality [BR17a]. Graphs [AAHLT10, AAD^{+}18, ARTV12, AAS22, ABS13, AFT12, AA10, Adl08, AH03, AD11, AKP^{+}22, AKZ17, ABM14, AHKiO22, ADM^{+}21, AHH21, AHDHL22, AZ22, AJM08, AH16, AH96, AGH14, ABHM00, AFK^{+}22, AD96, AFK12, APdMOZ23, ADO23, ACG94, AKKR08, AaW09, ABC^{+}15, AK22a, AGK^{+}23, AFKK23, AKTZ19, AHP09, AGH11, ADL13, AP14, AF19, AFG21, ABY11, AS07, AS09b, AS14, AG15, ARS21, ACD08, AM06b, AB07, ABHW13, ASSW21, AE03, BB16, BJKV07, BGG^{+}04, BGV07, Bal08, BCG^{+}24, BHLY08, BS10a, BS15a, BL19a, BCL^{+}23, BC02, BC03, BHRZ14, BJFJ95, BJHY00, BCS04, BGJW21, BPS07, BF96, BS09, BBF^{+}23, BCC^{+}11, BST20, BFP12, BBCZ11, BDM^{+}24, BGM08, BTU09, BC09a, BFGR17, BSKS11, BEL09, BIKY18, BCdMR08, BHLP92, BGL07, BDPR18, Bev10, BHH96, BG20]. Graphs [BBC^{+}23, BW99, BILL23, BGG^{+}20, Blo10, BKK95, BDJ^{+}98, BDFP10, BGP22, BMP13, Bon08, BZ11, BDEK06, BKM13, BK23, BHT10, BKTW15, BCL^{+}18, BLL^{+}15, BDDP24, BCG22, BCGR24, BIT13, BGS24, BMNS22, BD24, BDCV98, BL04, BLM10, BM16, BC09b, BK07, BFH^{+}08, BMN13, BMPS21, BGS23, BS93, BC23a, BKKM99, BCP08, BN05, BY08, BS10b, BS16c, BF17, CR10, CdVL11, CHM^{+}07, CCOY17, Cai93, CCVZ10, CDP08, CEHS08, CL15a, CL16, CTU14, CWY21, CCJS22, Cap03, CKL^{+}21, CDHH14, CW92, CMPS17, CDM04, CL20a, CCH21, CFH23, CHLT24, CCM^{+}23, CLS09, Cha19, CMM20, CFG^{+}21, CDM^{+}14, CF08, CL07, CKP13b, CGN^{+}06b, CEP18, Che94, CFGJ06, CEOT15, Che17, CM24, CL11, CLLZ18, CY18, CHK10, CLiO18, Cho92a, Cho94a, CKPS13, CS18c, CS18b, CPPT20, CGSZ20]. Graphs [CNST24, CGK94, CP16a, CH11, CHMP21, CMM^{+}10, CP10a, COL10, COLMS22, CCL^{+}06, Con05, CN12, CDKS21, CF05, CFK10, CFR10, CF10, CEOR13, CDFR18, CFP19, CFS96, CKOS06, COS10, CDHK16, CLS15, CD14, CW09, CR13, CR15, CR17a, CR17b, Cra19, CR19, CL20b, CY20, Cre04, CNG19, CL21, CY03, CLY05a, CLY05b, CLST12, CDK10, CM14, Dan01, DNS94, DI22, DOS94, De 24, DLMT21, DKWL20, DPR22, Dei15, DZ09a, DKRR12, DHT06, DSZ23, DK02, DV96, DMP07, Die10, DEE17, DL12, DD13, Do19, DEK22, DK06, DK10, DJ11, DX19, DMKŠ08, DHL^{+}13, DP15, DGS96, DF04, DPRS10, DKS18, DNB99, Dre12, Dro16, Duc21, DFT15, DP16, DP17, DEW17, DJM^{+}18b, DSST13, DMS12, DS23, DŠT08, DŠV08, DŠ09, DLŠ10, DLS11, DL14, DL17, DM17]. Graphs [DL18c, DL18b, DH20c, DH20b, DMN21, DM23, EK13, EIiK08, EKM^{+}19, EFMN08, EMOT16, EP10, EEFH21, EG03, EM99, EHJ01, EST14, EW19, Eri96, EJK^{+}09, EHLP11, EO16, EdJdVLT18, EGG21, EFK18, ELR98, ENSZ00, FRMPV15, Fan92, FJ09, FKS12, Far09, FMMO93, FHL^{+}13a, FHL^{+}14, Fei10, Fel90, FKMS20, FFie98, FG01, FPS20, FKLW98, FW02, Fle05, FKLL15, FPR18, FGRZ21, FJP22, FPS13, FW20, FGP10, FHS14, FJJ18, FT17, Fra92, FLM^{+}95, FL96, FHLT21, FK18a, FKS05b, FHL13b, FPS18, FT20, Fue23, FKV21, FP13, FYK00, Ful14, FK18b, GRR15, GGHP22, GGLS21, Gao13, GHM21, GCH22, GJ20, GH19, GP08a, GMS00, GM03, GM05, GS98, GM16, GS03, GKLT19, GnN06, GKP19, GR17, GK13a, GB12, GKMS22, GH90b, GHvtHP15, GPvL15, GHP20, GHV06, Gra07, GL14]. Graphs [GKR^{+}18, GK92, Gui98, GKL99, HMSZ24, HH04, HLS24, HPW09, HLST00, HRS18, Han19a, HLL^{+}21, HST19, HSS19, HVW07, HN15, HM94, Har18, HM12a, HHK^{+}22, HZ10, HSŠ08, HHHH02, HKL99, HZ08, HLR92, HLMP11, HvtHLN12, HH05, HHLM20, HHCHL21, HHL24, HRS17, Hen18, HY10, HK16b, HW24, HIMW24, HKP^{+}17b, HKP^{+}17c, HL00, HR05a, Hor19, HKM20, HHH^{+}02, HP23, HXZ18, HGY20, HT93, HM12b, HjK24, HC98, IKM99, JPvL21, Ja10, JT11, JY23, JMSW99, JMSW00, JS12, JL23, JKSW17, JLR^{+}17, Jon05, JRS14, Joo15, JN16, Joo16, JS03, JDT13, JW18, JMU23, JSRSW18, KKL^{+}10, KRS11, KR16, KiKK17, KKS19, KKP11, KLWY21, KMM12, KKP15, KM23, KNR92, KW92, KN16, KFHR94, Kar20, KNNW23, KM19a, iKSZ04, iKKKL09, iKM10, iKO16, KPP13, KNK93, KN22, KZ04]. Graphs [KMŠ^{+}09, KL14, KR04, KM22, KM19b, KM02, KV90, KMRR09, KMS23, KK14b, KMMS21, KW14, KM88, KS07, KNP05, KZ08, KMV15, KSŠ08, KM14, Krá04, KST06, KS08c, KSS09, KSS12, KT99b, KS93, KS02, KPPW15, KR20, Kri10, KLS10, KS12a, KRS15, KKS17, Kri18, KMN23, KO06a, KO06b, KOT09, KO12, KZ18, jKiO20, LS15, LTBL20, Lat23, LC04, LS22, LLM19, LPW^{+}13, LLS13, LiO15, Len98, Let19, Lev15, LM08, LLY10, LRWZ12, LL14b, LW17a, LHHL18, LLW22, LL17, LLZ99, LS95b, LS03b, LZ05a, LC12, Liu14, LRS^{+}21, LY23, Liu24, LR04, LHC90, Lu08, LWY13, LWY18, LZ18b, LLS^{+}20, LMW23, Lub90b, LdCKM18, LM10, Luk20, LXZZ08, LSSY10, Lyn94, MN18, MWWZ20, MMSW23, Mác13, MS17a, MS17b, MRS19, MS21a, MM93, MT05, MTV08, MM12, MP14, Mak07, MP94, MS16]. Graphs [MW14, MPRS22, McC21, MN15, MSS14, McK19, McL10, MWW94a, Meh12, MMP13, MP17a, MSZ10, MC12, Mer15, MNN18, MT03, MŠT09, MP17b, MTR14, Moh99, Mom13, Mor21, Mot19, MNS14, MvLvL13, MLS11, NI92, NOO12, NY23, NvdP19, NP20a, New20, Nie00, Nor11, OC10, Obr93, OYY13, OC19, Oze13, ONN19, PPR22, Pal12, PRR02, PZ10, PS22, Pen12, PT14, Per16, Pfe15, Pik03, Pip02, PT94, Pra13, PP90, Ram90, Riz02, RN21, RS08, RW19, RLWY21, RS15, Sak94, SMR08, SHS23, Sch91a, Sch92, Sch91b, Sch02b, SS94, SS04, Shp15, Shu23, Sid18, Sim13, Sli10, SZ94, Sol12, ST17a, Sta92, Ste10, SZ13, Sud08, SV08, Suk13, SSR94, Suz10, SB94, TT16, Tak22, TT91, Tam91, TW19, TT07, TW12, Tót10, Tuz08, Voi07, Wal19]. Graphs [WL02, WL03, Wan08, WX13, WHW14, WH15, Wan23, WIS12, Wei19, Wil99, Wol10, WZZ09, Wu09, WYZZ14, Xu09, XZ22, Yam16, YWWL21, YQZ09, Yus14, Zak14, ZZ92, Zam21, Zam22, ZLS08, ZZW13, ZWW23, Zho05, ZN08, Zho09, ZLWC12, Žit94, dFM04, dKPS13, o09, sG24, ALZ96, ACS97, BM97, BM94a, BM94b, BM94c, BFM94, BG88, BH97, BCLR89, Bon15, BCD97, CET97, CH89, CK96, COS97, DYL06, FH89, GPS88, GY92, HS88, Hut88, IS93, IKM^{+}92, Jac92, JO95, Kan08, KKV11, Kie88, KPT95, Kie97, KM06, LM89, MRS89, PS97, Pip89, PC97, RX88, RS93, Spr94, TA93, Tar88, TZ97, WGM95, dH89, BM94c, DJM^{+}18a]. Gravity [BG12]. Gray [Lic98, KCTR13, LT01, Sav97, Saw07, WS96]. Greater [FS12a]. Greedy [DW24, FHK96, Har10, HMS93, KOT16, Lev09, LRS23, Luc98, WS06]. Grid [AFK^{+}22, ABY11, BHL05, BDvL13, CW98, CLMS22, CHLZ00, CH06b, DPSW08, DKSŻ10, HPL13, JPT12, KFHR94, MMJF03, Sta11, ZZ92, ZGL^{+}09]. Grid-Block [MMJF03, ZGL^{+}09]. Grids [AJM08, CDP08, CFP11, DGM12, FKK18, GPRT11, Lag00, MP08, RX88, SS89]. Griggs [HRS12]. Grigoriev [KM24]. Gröbner [BP09, CP16b, GS93b, RY23]. Grooming [BCC^{+}05, BCG^{+}10, CFG^{+}09, MLS11]. Grothendieck [FR94]. Ground [DO08]. Group [ACKM19, AR04, Che93, DF94, DR04, GG11, GLY07, HH01, Hyu10, LM21, LLL17, Pet13, QP15, RD11, RKDD13, Yam16, HY89, YH88]. Group-Based [ACKM19]. Group-Labelled [Yam16]. Groups [ABM14, AMNV18, Bab92, BFGM15, CS14, CF09, DM13, Din06, FLM12, FFie98, GF08, GGI07, HZ21, JLD^{+}18, JS17, Kas05, KCL98, Lev22, Lub90b, MG19, MM93, MWW21, Mar09, MR15a, MW20b, Rón92, Sza06, Zho05, ALZ96, CS89, Jia95]. Growing [BYKKR18, CS12, KS12b]. Growth [ABH^{+}11, BG19, GPW09, GPS19, MW20b, RF12]. Grundy [BKL^{+}22, GSPRM91]. Guaranteed [GS93a]. Guarantees [AB95, ELMS11, EH13, GPSS01, GHP20, Fra89]. Guarding [HK96]. Guess [GSWW92]. Guessing [BHKL08, GRR15, KMS23]. Gupta [CCD^{+}23].

Hadamard [Mom13, Orr08]. Hadwiger [EKK^{+}15, GHvtHP15, GJWZ23, Pry23, ST17a]. Häggkvist [ABC^{+}23]. Hahn [BBM09]. Hajnal [BBT16, BMPS21, NP20b]. Hajós [BV20, PU95]. Hajós-Type [BV20]. Hales [Lav16]. Half [BYKKR18, BCC^{+}22, Brä10, DZ09a, Fuk16, iKK11, KS23b]. Half-Balls [BYKKR18]. Half-Disjoint [iKK11]. Half-Integral [BCC^{+}22, DZ09a, Fuk16]. Half-Plane [Brä10, KS23b]. Halfspaces [LS14]. Halin [CLS09, CEOT15, CEOT17, CS19]. Hall [BBK^{+}16, BLM^{+}22]. Halting [Cai18]. Halved [Lai18]. Hamilton [AHH21, AK22a, AB18, AF19, APS22, BF12, BNN90, BSKS11, BY08, CFH23, CF08, CKKO10, CFK10, CM14, DKM^{+}15, FHLT21, FKT99, GK13a, KMS12, KS12a, KO06b, LLZ22, OS13b, RS93, Sta92, ST17b, ZZW13]. Hamiltonian [Zam21, AS06, ABS13, AS02, dOBMS^{+}17, CL91a, CFGJ06, Dvo05, FH94, FMMO93, FLM^{+}95, Gui98, KP06, Kan08, Kar20, iKO16, LS15, LY21, LQ22, NPSW23, OZ18, RLWY21, ZZ92]. Hamiltonian-Connected [iKO16]. Hamiltonicity [AFG21, BSKS11, BP10, CFG^{+}21, DMP07, LRWZ12, PS22]. Hamming [AJM08, AE03, GPS19, LTBL20, LWW10, LS14, ÖV04, Pol19, RY22, VV94, XG20]. Hanani [PSS09, Sch22]. Hankel [CHX15, FP99]. Hanoi [Rom06]. Hard [BDPR18, DGM12, FGKP22, GMRT11, VVY15, LZZ22]. Hard-Core [VVY15]. Hardness [AS09a, ACG^{+}20, ASS17, BFN20, CH23, GKW19, GSZ22, GK04, GS20, HN15, ST23, HM20]. Harmonic [Bru90, CW92, FP13, FV97]. Harris [BLM^{+}22]. Harvesting [JPS^{+}14]. Hash [SG16]. Hashing [DHL^{+}02, MNP08]. Hastens [Ram04]. Having [AB07, BL23, CNG19, DK06, KCL98, LXZZ08]. Hazard [Juk21]. Hazard-Free [Juk21]. Heat [BILL23, Hor19]. Heavy [GNPV20, LRWZ12]. Hedetniemi [TW19]. Height [BMNP22, HKW22, JMW17, Luc98]. Heilbronn [Bar01, Bra05]. Hellman [CY08, FS01]. Helly [Ave13, CD16, DS21b, IN22, MT03]. Helly-Type [DS21b, IN22]. help [GK97]. Helps [CGV^{+}14]. Henning [Lic14]. Hereditary [ACFL16, BLL^{+}15, BC23a, CKP^{+}21, DOS94, DF10, GS13, HKL^{+}21, HHH^{+}02, KN16, LN24, Tuz08]. Hermitian [BM00, BN01]. Heroes [AAS22]. Heterogeneous [KVZ23, MHLHL91]. Heuristic [MTGK05, BVW88, Fra89]. Heuristics [GS94a, HS89b, KK89]. Hex [Sta11]. Hexagon [JMN22, Lai18]. Hexagonal [CHLZ00, HM88]. HHD [DNB99]. HHD-Free [DNB99]. Hidden [AA05, NYKY20]. Hide [CCNV11]. Hierarchical [Flu24, Sch22]. Hierarchies [LCV03]. Hierarchy [BM00, SA90, HSLd88]. High [BB90, BCHW17, Cha03, CK18, CFDK20, DDS16, Hen18, HOD22, LCV03, Wan23, Yus14]. High-Dimensional [DDS16]. High-Order [CFDK20]. Higher [FKT06, KM21a, LMP19, MW19a]. Higher-Dimensional [FKT06]. Highly [CCD00, CGOvZ21, FM13, GvZ19, KKO16, MW14, Ngu24, PTT16, Wol10]. Highways [DHX99]. Hilbert [HKW22, Shp22]. Hill [BLS19]. Hilton [CR19]. HITS [PP12]. Hitting [AK22a, BST20, CP10a, FLM^{+}16, GDVL17, GKMS22, JP12, JPS^{+}14, LMRS17, Tak08].

Hive [TKA18]. Hoc [KP04]. Hoeffding [SSS95]. Hoffman [AL07, PdCSC21]. Holds [Zha23]. Hole [CtJL01, CCL^{+}06, FSW13, RWW88]. Holes [AS16, BBT16, CS18b, DS23, FRW12, HjK24, MSS14, ST17a]. Homogeneous [GPW09, Kim17, KPT20, SV11, JO95]. Homological [JLD^{+}18, KM14]. Homology [BGJW21, Got03, LN17]. Homomorphic [DF94, DŠV08, KOS16, Zho93]. Homomorphism [BP17, EST14, HR12a, NT05, Wro20]. Homomorphisms [AFS12, BDK^{+}24, CL15c, Fed01, FHM03, FHH08, FGZ19, FGRZ21, GRY08, HZ95, MR15a, Zha11]. Homothet [IN22]. Homotopic [Sch91b]. Homotopy [Dan09, Wro20]. Homotopy-Like [Dan09]. Honeycomb [CHY13]. Hook [GGMM21, Han09, MPP17, RY91]. Hop [BHM^{+}22, HMP04]. Hop-Constrained [BHM^{+}22, HMP04]. Hopf [BHM16, FLS10, HM12b]. Horizontal [DGM12, IM96, SHS23]. Horn [BBM24]. Horton [WW91, Yam20]. Hu [NS22]. Hub [Yam05]. Huffman [FT05a]. Huge [OS15]. Hull [ACD^{+}13, BDPR18, CPRdS13, DPRS10, GVW06, KN16, SA90, Sen97]. Hulls [MD16]. Hurwitzian [Het14]. Hybridization [KvIL^{+}12, vIKL^{+}16]. Hyper [Moi08]. Hyper-Kloosterman [Moi08]. Hyperbolic [BC03, CD14, FK18a, KM19b]. Hyperbolicity [CCPP14]. Hypercontractivity [Pol19]. Hypercube [AKS07, Ana18, BGM08, Fin09, GT12, HPS96, HKK^{+}09, HL10, Kra18b, Off08, OV12, Shu23, Sid18, Bal88, Ram97b]. Hypercube-Derived [HPS96]. Hypercube-Like [HL10]. Hypercubes [AA96, BFPP08, BIKZ05, CL91a, Dvo05, DG08, GSV12, LLY10, MW08, Mol11, Spi19, VW09, YCH97, Chu89]. Hyperelliptic [CC03]. Hypergraph [BBLM16, BCE^{+}01, BBM24, CP20a, FJK^{+}23, HL24, KOT16, Lu04, NW95, ST20a, Tak08, WW22, YP22]. Hypergraphic [CKP13a]. Hypergraphs [AKP20, Adl08, AS19, APS22, ADKS18, AG19, BF12, BBLM13, BV22, BBBZ12, BF24, BCCZ11, BP20, BH22, BL01, BHT16, BK14, BCF^{+}20, CGHW22, CKP13b, dOCHLO19, CFDK20, CGH^{+}21, CM16, CF22, DJKO19, DMS12, EGM18, FRZ16, FGT11, FY04, FJK^{+}19, FZ22, GP17, GMOV20, GNPV20, GLS20, GLM22, GLP23, GSS15, HPS09, Han16a, Han16b, Han18, HST19, HT19, HY10, HY15, HWZZ18, HM19, JKSY22, JKiO21, Jon05, KLM14, Kha13, KR23, KZ08, KR15, LZ06, LZ18a, LZ09, LY18b, LYY21, LWY22, LP19, LLZ22, MMPPG23, MT03, OS17, OR04, Raz10, RRS07, SW14, SS04, ST20b, Sha23, Smo07, SV21, Spi23, Tim08, Yus03, ZKNS20, CFGG88, HWZZ20, Rya96]. Hypermaps [dR14]. Hyperoctahedral [Che93]. Hyperplane [RC18]. Hyperplanes [AA96, AS07]. Hypersurfaces [Gly12]. Hypertrees [CFDK20].

IC [YWWL21]. IC-Planar [YWWL21]. ICCSI [BC17]. ICSI [BC17]. Ideal [BTW08, BM22, CG02]. Ideals [BPEY22, CP16b, GMA15, GGI07, OPVV14, RY23]. Identical [AE04, PP07a]. Identically [PG06]. Identifiability [FKV21]. Identifiable [BEN08, TSN04]. Identify [FGP10]. Identifying [BCE^{+}01, BHL05, BLL^{+}15, CHLZ00, FL22, Lai05, Muñ05, Sta11]. Identities [CS94, Doh03, Kap14b, KLMR03, Kla06, Loe10, Mal89]. Identity [HK11, KY21]. Idleness [BCL^{+}18]. Ihara [Dei15]. II [BM94b, CEOT17, DMS08b, DF10, DL17, EK10, Gao18, GM94, HN23, Han16a, HKP^{+}17b, JLM93, MR04b, MAR23, MPP17, Mur96b, MRNS17, Sha23, SS02b]. II. [KW96]. IID [Rah16]. III [BM94c, DL18c, HKP^{+}17c, KLW89, KMV15].

IIS [Rya96]. IIS-hypergraphs [Rya96]. ILPs [JLR20]. Image [AF05]. Images [DS97, Kim91]. Imagination [BKR10]. Imaging [AG18]. Immersion [BS09, FW20, GPR^{+}21]. Immersion-Closed [GPR^{+}21]. Immersions [DK14, MW14]. Immune [Bac09]. Impact [BMRT20]. Imperfection [GM04]. Implementation [TKMM19]. Implicants [SST08]. Implications [FiT14, Hay21]. Implicit [KNR92]. Implying [GKY06]. Impossibility [BGN15, dH89]. Impossible [JSZ15]. Improper [CLiO18, WX13]. Improved [AYZ04, AMW00, ACG^{+}20, AS14, Bac09, BL19b, BDM02, CH23, Che04, CCDZ23, CNP21, Däu22, Doh03, DJW12, DS16, DvHT20, ELMS11, Gab05b, GHP20, HJ18, JNiT21, KKW17, iKK11, KW13, KS07, KT17, KO06a, LLS04, LÖ05, Lu10, MSZ10, New20, NÖ08, Osh21, Sot15, Sta11, TW21, Tam88, Tod14, VVY15, Woe93, dKMP^{+}06, dKPS13, LRN11]. Improving [CSS01, FL22]. Imset [XY15]. In-Community [LN21]. Inapproximability [EK05, GSS15, GL17]. Incidence [AS07, Bal08, CF22, HM12b, IK09, LS16, Mac18, Pip02, RS16, SS08]. Include [Pin14]. Inclusion [Doh03, Fér15, KS05]. Inclusion-Exclusion [Doh03, Fér15]. Incomparability [BKM13]. Incomparable [DJW12]. Incompatibility [HLWY23]. Incomplete [BHH94]. Increasing [BJFJ95, HSS24, WC12]. Incremental [DKLW24]. Independence [ACL^{+}06, ASSW21, BNR96, Che24, DW22, DMS12, FMMO93, FLMY09, Fox10, FGKP22, HC98, KM13a, LOW10, RN23, Swa05, Tót10, SSS95]. Independent [Age94, AFKK23, BLM10, BKKM99, CW98, CHKW24, CER98, CCG^{+}11, CPPT20, CY21, DM17, DJMV21, FRMPV15, FKS12, FO08, Fle00, FL96, Für91, GS93a, HST19, HT18b, HT93, IT08, Kir16, LZ06, OR04, PP07a, Pit23, RSV^{+}14, SS04, TT16, Vse05, Zha90, Bou97, GS89, Sag88]. Independently [PSŠ10]. Index [BCHW17, FKS05b, FL15, KR15, LW03, LLW22, LM10, MS21b, Saa93, WH15]. Indicators [DSS12]. Indices [Wag07]. Indifference [SS94]. Individual [HH01]. induce [KPT95]. Induced [AHJ^{+}24, AFK12, ADO23, AKS08, AS19, AKTZ19, AB07, BB16, BP89, Che93, CPPT20, CDKS21, CF16, Cor24, CM07, DF02, DS16, DS23, EIiK08, Fis90a, GP18, GPvL15, GRS12, HLS22, HLS24, Hua14, Jan05, JRS14, JN16, Joo16, JW18, KMM12, KM19a, jKR21, MST21, McC21, Nag10, PTT16, Pit23, Tuz08, BM94a]. Induced-Sidorenko [Cor24]. Induced-Universal [AKTZ19]. Inducibility [CSW17]. Inducing [CdV02]. Inequalities [AB95, BCCZ10, BDP19, BH16, BCH92, CPP23, CdV02, CR96, CL22, Doh03, DJS12, FR94, FP13, GL14, Kra18b, LRR21, LMP19, MWW21, SS05, SL95, TZ19, TZ20, Kah97, LR94]. Inequality [AF10, Bar95, BL90, CIN18, DDS16, IMR14, OV12, RSD17]. Inference [DM03, SHS23]. Infinite [BPST22, BCE10, BCCZ10, Bon10, BCF^{+}10, CW92, Dei15, JK24, KLMR18, KK09, LN21, PT14, Ran02, Śni14, Ste10, SZ13, AHS89, KM06, Wen97]. Infinitesimal [Dew20, KiT13, SiT15]. Infinity [BL17a, Ber20]. Influence [BS23, Che09, CGG88]. Information [WS17a, KG93]. Ingleton [NvdP18]. Inhomogeneous [McK19]. Injective [BMR^{+}10]. Input [RF12]. Input-Output [RF12]. Inputs [SY11]. Insertions [CGHL23]. Insertions/Deletions [CGHL23]. Insights [MV99]. Instance [PSV08]. Instances [ASZ02]. Instantaneous [FO00]. Integer [AOW15, BCCZ10, BDP19, BBCZ11, BP09, BP12, BCF23, BH93, CS09, CLLZ18, Dan09, Dow91b, GVW06, Gij05, GM90, HLL^{+}21, HS04, JB22, KYDN09, KM21b, MD16, MvLvL13, Myu01, NP18, Onn91, PPU92, She18, Tro15, WCLZ15, WS17b, Yam07, Dow91a, FKS97, TT89, UWZ97]. Integer-Valued [CLLZ18]. Integers [AHP19, AL24, HST19, Han19b, Hav19, KSY18, KLMR18, MW19b, Rya07]. Integrable [CHX15]. Integral [ABM14, BCC^{+}22, Che17, DZ09a, Fuk16, Jan10, Jev95, MH09]. Integrality [BB09, CCZ12, DZ09a, GW19, HMM09, Lon21, dCST20]. Integrity [BEL09]. Interactive [Bab92, Orl93]. Interchange [Die10, SS00]. Interconnecting [DHX99]. Interconnection [CKN^{+}15, Zho09, DM88]. Interference [FKLW98, FW02]. Interference-Minimizing [FKLW98]. Interior [Ave12, DJT15]. Interlacing [CDH^{+}04, FJZ15]. Interlacings [Die10]. Interleaving [JCB06, MP98]. Interlocked [CFK19]. Internal [DM03]. Internally [GZ06, MR12]. Internet [BAG03]. Interplay [FHM94]. Interpolation [CW14b, CF17]. Interpretation [CS94, Mun06]. Intersecting [BLP21, BKK16, Bor10, CLWZ22, CL07, CLW09, GLP^{+}12, Ger23, KŠ08a, KL19b, Luk20, MB18, Spi23, WL02, WZ18, ZWW23]. Intersection [ABS13, AC14, BBC11, BTU09, BKYY23, Blo10, CH89, DMN21, FF06, GMA15, JSRSW18, KS92, KMW06, Koc98, KM01, MO22, MSZ10, MvLvL13, Pet13, PR98, Sha13, Suk13, Mur96a, Mur96b]. Intersections [FKM13, GDCM20, KS05, ST10]. Intertwines [Bon10, GI97]. Intertwining [CW14a, HvZ14, LiO23]. Interval [BNCPR20, BFPP15, CFGJ06, COS10, DJ11, EST14, EEL09, FHM94, Fer16, FG00, FSV13, FJJ18, GP99, GP08a, HH05, HR12b, KV90, KS06, KS02, LC04, LW17a, LHC90, MN15, Mer15, BG88, Kie88, KW96, MM96, PS97, Spr94]. Intervals [BCdMR08, BEGK22, DK10, Fer16, KJJ04]. Intractable [CM12a]. Invariant [CD93, CDHH14, CDR16, Gly12, KLR22, KG93]. Invariants [BC23b, DEH20]. Inverse [BILL23, GXY22, LW03, Mal15]. Inversion [CGG^{+}16, Mal15]. Invertibility [APS^{+}24, Con89]. Invertible [GK16]. Inverting [SS02b]. Involutions [AAHLT10]. Involutory [TKA18]. Involving [HZ21, LL21, SWR12]. Irrational [Köp07, YP22]. Irreducible [RMS01, vzGVZ13, BIKY18]. irredundance [BG88]. Irreflexive [FHL^{+}14]. Irregular [AT90, CFS96]. Irregularity [BBLM16, KKP11, MP14, Nie00, o09]. Ising [COLMS22, GGHP22, HLP^{+}23]. Islands [EO16]. Isogenies [BT18]. Isolating [PTT16]. Isometric [BGM08, EGG21, FHJV17, Mol11]. Isometry [Hyu10]. Isomorph [Kas05]. Isomorph-Free [Kas05]. Isomorphic [AAFL06, BR20, Con05, GJXZ20, GJK^{+}23, dH89]. Isomorphism [Che94, FKKL98, Neu24]. Isoperimetric [BL90, HLST00, Kah97, KW14, Lev15]. Isostatic [FSW13]. Issues [BAG03, Nas14]. Item [CCG^{+}00]. Iterated [Fra10, Zhu18, Mal89]. Iterating [Lic98, LT01]. Iterative [SFS09]. IV [BFM94, HKP^{+}17d]. IX [SSS24].

J [GM93]. Jackson [BBM09]. Jeu [Śni14]. Jewett [Lav16]. Job [GPSS01, JSOS03]. Job-Shop [GPSS01]. Jobs [Jan10, PP07a]. Jogs [TH90]. Johnson [Etz96b, RD11, SG16]. Johnson-type [SG16]. Join [Rea15]. Joins [ACP22, Bar04, BK07, WY20]. Joint [AK22b, EMT15, SiT15]. Joints [Qui10]. Jordan [vBEM17]. Judicious [LSSZ19]. Jump [BC95, KMT07, Mur06, ST07, Shi12, Sza08]. Jumps [HLR13]. Just [MR15b].

Kac [TW12]. Kaiser [CCO^{+}13]. Kakeya [LSW18, MWW20]. Kalai [Tod14]. Kannan [DKM^{+}12]. Karakhanyan [BL09]. Katona [Buk12]. Kazhdan [GXY22]. Kernel [BJS22, FSV13, GLSS16, JPvL21, KL19a, PRS18]. Kernelization [ALM^{+}18, BJK13, BJK14, CFG^{+}21, FHNN18, FLM^{+}16, FG21, JP18, KKS22, Kra18a]. Kernels [BFRS16, EKM^{+}19, GSPRM91, GSL98, GPST15, GPR^{+}21, Hed08, HKP22, KMR^{+}23]. Kerov [Śni14]. Key [FGPS20, GLY07]. Khintchine [DDS16]. KillerQu’est [KvIL^{+}12]. KillerQu’est-ce [KvIL^{+}12]. Killing [KM21b]. Kimura [MRV17, Vod21]. Kimura-3 [MRV17]. Kind [MGC14, QD14]. Kinds [HMZ21]. Kinetic [KvRRS23]. Kinetically [Mar20]. Kings [SS91]. Kissing [KKW17, MSD19]. Kite [LRS^{+}21]. Kite-Linked [LRS^{+}21]. KKM [AFP^{+}18]. KKM-type [AFP^{+}18]. Klapper [CK11, ACM^{+}18, Shp10]. Klein [iKKKL09, iKMN09, RS16]. Kleitman [Tod14]. Kloosterman [Moi08]. Knapsack [DKMS17, DKLW24, KSF19, LMNS10, SL95, Yam07, Yos19]. Knee [Lei94]. Kneser [CMSM^{+}18, EJK^{+}09, FK18b, HPW09, Hui23]. Kneser-Type [Hui23]. Knock [Lei94]. Knock-Knee [Lei94]. Knowledge [DF94]. Known [CPS08]. Knuth [Śni14]. Ko [DT16, DK16, MS14]. Komlós [HLT19, Kun23]. Kotzig [CuKŠ07]. Kruskal [Buk12]. Krylov [BGL03]. Kühn [MMS17].

L [BL17a]. L-Infinity [BL17a]. Label [BBC^{+}19a, BKS09]. Labeled [Ald22, BJS21, CX08, CFJ11, GM03, MS16]. Labeling [ABR05, AKTZ19, BBC^{+}19a, BGP22, CKNZ05, DSZ05, GP08a, GMS00, JPT12, JSRSW18, Kan08, Kim91, LZ05a, Meh12, WL03, ZLWC12, CK96]. Labelings [BJKV07, GM05, Gra07, GK92, KST06, Lag00, LZ05b, MMP13, HRS12]. Labelled [Ald90, Yam16]. Labelling [GY92, Sak94, Zho05]. Labellings [BMR^{+}10, EJK^{+}09, GJ06, KŠ08b, FGK89]. Laced [Sim13]. Lagrangian [YP22]. Laguerre [FZ88]. Languages [CHM^{+}23, ETT13, FKL93]. Laplacian [BMV92, CFM94, CL15b, GM94, HRS17, Iri16, LMP19, Ost23]. Laplacians [CDH^{+}04, JLD^{+}18, Ste07, TH11]. Large [AA10, AKZ17, AHH^{+}10, AHDHL22, ADO23, AKS08, AS07, AS16, BBC^{+}19a, BFK^{+}12, dOBMS^{+}17, BW02, CCJS22, Che07, CKP^{+}21, CP10b, CDKS21, Cra19, DLMT21, Dro16, DM17, FG14, FKP15, FM13, GP18, GJ12, HPS09, HS04, KL19b, Kha13, KZ04, Kim17, KPT20, KST06, KO06a, LLM19, Lee17, LP21b, LL21, MP08, MNS14, Ngu24, PPR22, SS04, Sha20, SZ13, Ste22, Sul05, WH15, Wan23, Yam20, EH91, KKV11, RCS88, WW91, RX88]. Large-Girth [AA10]. Larger [KMP03]. Largest [Ald22, GW99, Iri16, KW08b, KM19b, SST08]. Lariat [DvW18]. Lassos [HK15]. Last [JZ05, KKS10]. Latency [LRTW11]. Latin [BCM^{+}12, MW08, WLD09]. Lattice [Ave12, BHE05, BH16, BT96, BS16a, BS16b, Can93, CL90, DD15, DMN12, EIH18, FGPS20, FL00, FXZ23, Got03, HST22, HKW22, KNK93, KT17, KŠ08b, KS03d, MGC14, MW19a, NT12, NDB07, OPVV14, Onn91, Rea08, SFS09, Tsu23]. Lattice-Free [HKW22]. Lattice-Simplex [FL00]. Lattice-Width [DMN12]. Lattices [ADL^{+}09, ABH^{+}11, BFGM15, CGG17, CCGG18, EHV18, FM11, GJ08, GV21, KL08, Lec90, MPSV21, NDB07, NS16, RSD17, Sch09, WZ08, DSW90]. Laurent [CHX15, KM24]. Law [AZ22, CK91, Ja10]. Lawler [GW00]. Layer [Bah23, GS95, Lei94]. Layer-Rainbow [Bah23]. Layered [DEJ^{+}20, IM96]. Layers [DM18b]. Laying [HLR92]. Layouts [GSV12, KLN91, KS02, HP97, Pat88]. Lazy [JJL23, RT11]. LBFS [COS10]. Leading [CDR16]. Leads [FS12a, Ja10]. Leaf [Ald22, Alt89, vIM18, HS97]. Leaf-Labeled [Ald22]. Leaf-Reconstructibility [vIM18]. Leakage [AEFT13]. Learning [AA05, BBS00]. Lease [LPSR12]. Least [ARTV12, BBF^{+}23, CPPT20, DŠV08, DH20b, HK16b]. Leaves [AFG^{+}09, BKKZ17, Bon08, BZ11, CWY00, KW91]. Lecture [BBK^{+}16]. Lee [Špa07]. Lee-Type [Špa07]. Left [Cor24]. Left-Cut-Percolation [Cor24]. Lehman [AFG^{+}16]. Lekkerkerker [LM16]. Leman [FKV21, KN22]. Lemma [BKTW15, BK17, CS18a, DKM^{+}12, DEMT22, FPST06, NT12, Peg14]. Lempel [CI07]. Length [BDL21, BZ23, BH20, CCD00, CGH^{+}15, CPPT20, EK10, FJ09, Han09, HT19, Hed08, HY12, KW13, KS03d, NLJM21, SS09, Ste22, TSN04, sG24, vD11, BFGR17, DL89b, HS97, Tas97]. Lengths [CHP^{+}90, GS94a, GGMM21, GKPP18, JA16, KRS11, KJJ04, MN18, PY90, SW99, Tan21]. Lens [BNRZ23]. Lenses [ERSZ22]. Letter [AFK^{+}22]. Level [AYZ04, ACF18]. Levelness [HKT20]. Levine [AFKK23]. LexBFS [BCHP08, HH05]. Lexicographic [BBM09]. Lexicographically [KRR16]. Li [RY23]. Liar [DS05b]. Lie [ADL^{+}09]. Lifetime [BNRT17]. Lift [AT16]. Lift-and-Project [AT16]. Lifting [BDP19, BK12]. Liftings [SL95]. Lifts [ACKM19, CF08, FT12, GS16b]. Light [BT22, DS06, ES11, ENS15, Sol12]. Like [BHM^{+}22, Dan09, HL10, EIJ^{+}12]. Likelihood [CFKK17, SS02b]. Limit [BK11, CP10b, WW91]. Limited [BGM22, OR04, SSS95]. Limiting [Gar92, RJS93]. Limits [Liu24]. Line [AE04, BD20b, BKS10, CvBdJdVK22, CH11, CR17b, CMM^{+}24, Gab05a, GSPRM91, GKM^{+}18, GKL99, HT90, HV00, KFHR94, KT99a, Kap14a, KPT94, McD15, MZ22, MT03, Sch91b, Woe93, BCP08, CKPS13, Con10, Fra89, KPT95, MSS14].

Line-Polar [CH11]. Linear

[AGQS23, AD96, AL24, BB13, BNMN92, Bar02, BZ04, BBCZ11, BJL^{+}24, BL17a, BHH96, BGS96, BP09, BKG99, Boy96, BGS24, BCHP08, BM13, CD93, CS14, CdMR12, CKNZ05, Che94, CGHL23, CKOS06, DDJ^{+}21, Dji06, DHJ^{+}13, EGM18, FR99, FM11, Fio06, GCH22, GPST15, GGMM21, GPR^{+}21, GH90a, GMS15, Han98, HKL99, HvtHLN12, HK05, HTS18, JSOS03, JLR20, JMW17, KPT22, KMP03, KK90, LW17a, LL15, LM16, LWW10, MG19, MO22, MN15, Mer15, MNN18, Moh99, Naa00, PS17, Pon20, PR91, Ram98, RS15, San96, She18, ST07, Sis21, SY11, Spe08, Sta92, TRV03, Ten09, Wan02a, WWKY11, Wen97, Wu09, Yam16, o09, HSLd88, IS93, LM89, Mou94, KG93]. Linear-Complexity [BZ04]. Linear-Interval [Mer15]. Linear-Time [Che94, Dji06, DHJ^{+}13, HKL99, LM16, MN15, MNN18]. linearity [Kie88]. Linearized [Wal19]. Linearly [JNiT21]. Lines [Buk16, KT19, PV22, Pay17, SW04, Vin07, Yu17]. Linkage [GST23]. Linkages [BP10]. Linked [GKY06, KY12, LRS^{+}21, Pfe15]. Linking [Pip06, IP91]. Links [CDP94, DEH20, MMO20, MAR23]. List [BSY21, CEHS08, CGSZ20, CR17a, DSZ05, DŠT08, DH20b, EST14, FHKM03, FHH08, FKS05a, GJWZ23, GMOW21, HLS22, HLS24, HM11, KZ08, RR03, Saw07, Tak08]. List- [HLS22]. List-3-Coloring [HLS24]. List-Coloring [CR17a, DŠT08]. List-Decoding [RR03]. Listen [MP04]. Listing [CGMV19, PRS98]. Lists [FKS03]. Little [MR15b]. Littlewood [BI13, LR94, TKA18]. Liu [HL92]. LKS [HKP^{+}17b, HKP^{+}17c]. LLL [BGW20]. Load [AE04, HMP97]. Loading [Däu22, Myu01, SSW98, Sku16]. Local [Ath14, BJFJ95, BYR05, BGJ^{+}12, CER98, CL20a, CCGG18, CSS13, CL11, CKPS13, DFG^{+}21, DSS12, DFHT04, EH13, FPS20, GK02, IKK06, JS03, KW08a, Lev09, Peg14, Roz19, Špa07, Tre04]. Local-Optimality [EH13]. Local/Global [CL11]. Locality [BMRT20, HSV23, MNP08]. Localization [ASMF10]. Localized [APS22, LSO03]. Locally [BK11, BY08, BDK^{+}24, DEW17, HW17a, iKM10, Kri18, Lub90b, MG19, MT20, Sol06, KS88]. Locating [Ber07, MMPS10]. Location [AYZ04, BB09, CCZ12, CFH16, KVZ23, RS10, Tam91, Yam05, HM88, Tam88]. Locked [ZK11]. Loebl [HKP^{+}17a, HKP^{+}17b, HKP^{+}17c, HKP^{+}17d]. Log [AAMM24, GMTW15, CL22]. LOG-Brunn [CL22]. Log-Concave [AAMM24]. Log-Concavity [GMTW15]. Logarithmic [GV92, KW90]. logarithms [Gor93]. Logic [Fer16]. Lollipop [DvW18]. Lonely [CKdAHdF13]. Long [AB05, BBC^{+}19b, BBT16, DJW12, EMM14, FG14, HHK^{+}22, HjK24, JY23, KW08a, KS23a, LLS^{+}20, NT05, Ste00, ZWW23, Kar89]. Longest [BM14, BH20, CGH^{+}21, EFMN08, FJ09, Gut93, KW13, LMM21, MC12, RS14, Zam22]. Look [CF10]. Look-Ahead [CF10]. Loop [BH13b, HJ94]. Loops [HHCHL21]. Loose [dOBMS^{+}17, CM14, OS17, OC19]. Loss [FS12a]. Lossless [EN22]. Lossy [EKM^{+}19]. Lot [DW10, DW11, vWW94]. Lot-Sizing [DW10, DW11, vWW94]. Lotteries [KMR95]. Lovász [KG98, CEOT15, Cre04, LS05]. Low [CF09, DK89, ES11, GS20, JJiT14, MR15a, SBD^{+}19, Tót08]. Low-Density [CF09]. Low-Dimensional [MR15a, SBD^{+}19]. Low-Distortion [SBD^{+}19]. Low-Rank [JJiT14]. Lower [Bar01, BGJ^{+}12, BJK14, Cap03, CKL^{+}21, Chi23, DSZ05, DHUZ01, DMS12, EMRPS14, FS09a, FP01, FP04, FHNN18, FLL10, Fra95, FXZ23, GY23, HBDS23, HP21, JWF05, KKW17, KL92, KM95, KW14, KSS09, Kun23, KOP^{+}21, Lab13, Lei94, LN24, LY10, Luc03, MNP08, Nao91, NW95, OS92, PVY22, PP90, ST13, SS18b, Swa05, Tsa96, XSR11, YAT16, Zuc92, dKPS13, Car88, Sar97]. Lower-Diameter [Cap03]. Lozenge [Lai18]. LP [CKP13a, CQX20, GW19]. Lusztig [GXY22]. Lyapunov [GM93, GM91]. Lyndon [KRR16, RMS01].

Macaulay

[Bro11, BCW96, CN12, DDL20, PPTY22]. Machine [CMSV17, GQS^{+}02, HV00, MV18]. Machines [AE04, ELSS17, JMS90, SS02a, SSS13]. MacWilliams [Kap14b]. Made [AH21, PR98]. Magic [Gra07, OS13a]. Mahonian [IKZ08]. Main [HKP^{+}17d, MLMR23]. Maiorana [TKMM19]. Majorities [EGMR20, Lec90]. Majority [AB94, EGMR20, Lon21, WS12]. Majorization [Mus21, Ste07]. Maker [CFG^{+}15a, CHMP21, Han19b]. Makespan [JSOS03, Svi03]. Making [BIKY18, CKMU14]. Management [Car09, GLY07]. Manhattan [CDF08]. Manifolds [Mur10]. Many [AH16, AFG^{+}09, BM07, BMV22, Bon08, BZ11, BCG22, CWY00, CPS08, GKW19, HM12a, KKSvL20, KR23, KW91, Let19, Luo23, MPŠ08, Mon15, NvdP18, OPS21]. Many-Visits [BMV22]. Map [KCTR13, QP15]. Mappings [Fed06, WC12, DS97]. Maps [AL17, BC94, CMS09, Fel14, GW99, iKX20, MAR22, MAR23, BT97]. Marching [Che16]. Marginals [FGP10]. Marked [APdMOZ23, FFLP20, JS14]. Market [BKK21, Ete22, Vaz13]. Markets [BBKS23, CDV10]. Markov [ART14, AK22b, BK90, EMT15, HH92, MP13, RS22, SV11, Win16]. Markovian [CMM^{+}10]. Marriage [EMM14, GSZ22, HW17a, PSV08]. Mars [MTGK05]. Mass [DK22]. Master [DRW98]. Matched [SV08].

Matched-Factor [SV08]. Matching [ABFRV24, BH22, CGHW22, CDHZ12, CCDZ23, CS91, DSS12, ELMS11, Fin09, FJ17, GPP04, GH90b, GL95, GMPZ15, Han16b, KMR11, KT14, Kam22, KMS98, LdCKM18, LLZ22, MS21b, MNN18, NH91, Tak14, VW09, Wal19, WZ08, Zha99, ZZW13, BC88a, Hef97]. Matchings [AKZ17, ACG94, ABY11, ARS23, BBCZ11, Can93, CL11, Che16, CY18, CDKS21, CHK17, CFKP22, DJKP09, DM19, FJJ18, FKT99, FKM13, GMP22, GLM22, GLP23, HPS09, Han16a, Han16b, Han18, HKP01, Har18, HR02, HMS21, HK13, HOD22, IKK^{+}22, JRS14, JN16, Joo16, JP06b, Kam17, Kam19, KMM12, Kav24, KY18, KMW24, Kha13, Kla06, KSS09, KO06b, LRT08, LWY18, LY18b, LYY21, LWY22, ŁR91, MMSW23, Mak07, Nas14, OC10, Pip02, Tak22, WY23, dCLM13, Pit89, RS93]. Matchoid [HW23]. Matchstick [LS22]. Math [GM93]. mathematical [SU89]. Matrices [AK22b, AM06a, Bal08, BS95a, BK17, CP10b, EM20, FHK96, Gij05, GK16, HN23, HHL24, IM96, IT09, JJiT14, JiT22, KS05, KM21b, Kun23, LL14a, MP22, MW19b, Mon15, Ngu13, OPR12, Orr08, OW22a, PSWX24, Vin12, dLL09, BG91, DK89, KP06, Spi95]. Matrix [BM15, BR17b, CF22, EMT15, FLM^{+}18, GWZ18, HS22, Iri16, KK90, LS17, LSO03, MW20b, OPVV14, Som14, Tro15, WS17b]. Matroid [BPWX13, BKK16, BKK21, BKYY23, BLMA^{+}08, BCF23, FF06, FKM13, GMA15, GGW06, Hli18, HKMY22, HN20, KT14, Kam17, Kam19, Kam22, Kas08, KiT13, KS03b, LMNS10, MO22, May08, Onn03, Oxl19, PVS22, Sza08, Yam16, dGNS13, Mur96a, Mur96b]. Matroid-Based [dGNS13]. Matroidal [Sha13, Yok19]. Matroids [BBJ^{+}21, BSY21, BS24b, Ber20, Bon10, BL23, BC95, BST23, BH13b, BRS09, CW14a, Che22, Che24, CKP22, CDMO16, CMvZW16, CGOvZ21, DD15, DM88, FMOS20, FGLS18, FM24b, GXY22, GHW05, GH06, GZ06, GvZ17, GvZ19, Gra91, GPW13, GS16b, GV21, HKL11, HKK24, HvZ14, HN20, Jan05, JKiO21, Kap14a, KBE^{+}05, Kot13, KRZ21, KS23b, KMPR14, MWvZ11, MR12, Nel15, NvZ15, NvdP18, NN22, Ngu10, OS22, OW22b, PG06, PP13, Sno13, Tak14, Tak22, iT12, Wag18, Wag21, Web08, Wil05, dM07, CH89, Whi88]. matter [CHO^{+}89]. Matters [KPT12]. Mattson [BR19]. Max [ADKS18, CKL^{+}21, COLMS22, ELvS11, FiT14, Gün07, HSZ13, HR12b, KSS12, PS17, Sot15, SS00, Fuj97, Hef97, AHS01]. Max-Conflict-Free [ADKS18]. Max-Cut [CKL^{+}21, Sot15]. Max-Flow [Gün07]. Max-min [ELvS11]. Max-Product [PS17]. Maxima [BNMN92]. Maximal [Ave13, BGG^{+}04, BKS09, CER98, CP20a, CD16, CGMV19, Dji06, DSW90, Für91, GPS19, HKP01, HL24, HT93, KW08a, Kir16, Kra18b, MD11, Muñ05, RRW22, Sza06, TSN04, GS89]. Maximin [ARS95]. Maximization [DKLW24, DW24, HKMY22, ILM20, KT99a, KSF19, Osh21, FKS97]. Maximize [Car09]. Maximizing [AAH14, BNRT17, CvBdJdVK22, CKNV16, HIKT99, Kas03, KL08, LMNS10, Yos19]. Maximum [AD11, ANS16, AS14, ASSW21, BB16, BB13, BLP21, BP15, BHL^{+}15, BPSS20, BHH96, BLO^{+}22, BW02, BLM10, CCD00, CCJS22, CMPS17, CFKK17, CKP13b, CDHZ12, CPPT20, Cib13, CM07, CR19, DSST13, DK14, EJ01, Fei04, FO08, FvIKS15, FGP10, FT20, FTS24, GP18, GL95, GRS12, GY23, HJ18, HZ10, HRS17, HV00, Hua14, JT11, Joo16, Kan08, KSS11b, Kav24, KL14, KW08b, KY12, KLN10, KSŠ08, KL08, KMPR14, LS03a, LRTW11, LSL92, Luc03, Mak07, MMP13, MNN18, MS19, Oze13, PS21, Pon20, SY11, SS02b, Suz10, WH15, XQ06, Zak14, Zeh17, GW94]. Maximum-Cardinality [MNN18].

Maximum-Length [CCD00]. Maximum-Weight [CDHZ12]. Maxmaxflow [RS15]. MaxMinSum [DM18a]. McFarland [TKMM19]. MDS [BHTK^{+}21]. Mean [AGQS23, HH92]. Means [Kra18b, LRR14]. Measurable [GL15]. Measure [AGQS23]. measurement [FMRR88]. measurements [LS89]. Measures [JK24, Śni14]. Mechanisms [HLR92]. Medial [Kim91]. Median [BBC11, BH13a, BK07, CL11, DHUZ01, DZ01, IKM99, MP95]. Medians [BC02, Lec90]. Meet [Fra92]. Meets [Lev09]. Mélange [DS21b]. Melas [VV94]. Membership [BM22]. Memoryless [BBFP09]. Menger [AHJ^{+}24, BK19]. Menger-Type [AHJ^{+}24]. Mergings [Han19a]. Meromorphic [Lla06]. Mersenne [EA11]. Mesh [CDP94]. Mesh-Connected [CDP94]. Message [FKPR05, KPR10, MT11]. Meta [CCT05]. Meta-Fibonacci [CCT05]. Metabolism [WDSH21]. Method [BM11, CP10a, MRR20, MdCW16, Ros09, Tak08]. Methods [AT16, DS05a, DA10, GKLT19, LZ06]. Metric [BDF^{+}18, BFGR17, BMV22, BFKT24, BDG^{+}17, BR17b, BR19, CHM^{+}07, CKNZ05, CV07, COHK21, DM11, GKI^{+}23, Gao15, GRSZ23, SS11a, Sol12, XG20, Ste88, Win88]. Metrics [Ban90, BS09, BHM^{+}22, Dah93, HMM09, SP88]. Meyniel [DFK^{+}23]. Midpoints [EFF91]. Migration [JK19]. MILP [Jan10]. Min [Cal13, CMSV17, FiT14, Gün07, HM94, HR12b, HHMR20, KSS12, SS00, Vse05, ELvS11, Fuj97, Hef97]. Min-Cut [Gün07]. Min-Max [FiT14, HR12b, KSS12, SS00, Fuj97, Hef97]. Min-Orderable [HHMR20]. Min-Power [Cal13]. Min-Sum [CMSV17]. Min-Wise [Vse05]. Minconvex [AS09b]. Minima [LL14a]. Minimal [ABZ15, ABNR22, BNN90, BCCZ10, BDP19, BKS09, BBC^{+}23, BB03, BCG22, BC23a, CY12, DDL20, Gab05a, GDVL17, GMA15, GH90a, HRS18, HTV05, HK15, KLMN14, KRR16, LPS18, Mus21, RS08, Saw07, SL95, Sis21, Tak08, Zeh17, Zhe16]. Minimal-Time [Gab05a]. Minimally [AG15, OW22b, Ste10]. Minimax [HKST03, Ram90]. Minimization [BDvL13, FI05, GS13, JSOS03, Svi03, CNR89]. minimize [DL89b]. Minimized [FT05a]. Minimizers [GP20]. Minimizing [Alo13, AE03, BYKKR18, CCH21, CKK^{+}04, CFG^{+}09, ELSS17, FKLW98, HRS17, HQ03, HV00]. Minimum [BGV07, BLP21, BJHY03, BHL05, BBRT23, BPSS20, BBC^{+}23, Bon08, BS15b, BL17b, BCC^{+}22, CCJS22, CR04, CtJL01, CQX20, CH01, CL13, CF09, CDK10, Das99, DLMT21, Dro16, DH20a, ENSZ00, FKK21, FJ17, FPS18, FK18b, FZ22, GV92, GKY06, GP91, GÖ12, GSS15, GRY08, HIKT99, HPS09, HRS18, HM94, HR12a, HK16b, Hng22, HK93, HOD22, IMS05, KKL^{+}10, KN16, iKX20, KS05, KKM94, KP95, LRT08, Li17, LL17, Lon21, NLJM21, OC19, Pin08, Ram04, SS11a, SMNF09, Spi23, Ste22, Str20, UBHS93, WY20, XGG15, Yus14, ZN08, BBM90]. Minimum-Cost [CL13]. Minimum-Time [KP95]. Minimum-Weight [FPS18, BBM90]. Minkowski [AF10, CIN18, CL22, DMN12, GS93b, LS06, OV12, SS09, Tsu23]. Minor [BFM06, BJS22, CDMO16, CNP21, DHJ^{+}13, DEJ^{+}20, EMOT16, FGRZ21, HKSS08, HN20, JW18, KMPR14, MR12, Per16, BM94c]. Minor-Closed [BJS22, DEJ^{+}20, HN20, Per16]. Minors [AP18, AFH^{+}18, BST20, BL23, BJS21, Che22, CFP16, DDJ^{+}21, EW19, FKPR05, FHJV17, FGT11, FLM^{+}16, Fox10, GMA15, GT13, HW24, iKMN09, KM14, KNZ14, LiO23, NW22, SZ13]. Minrank [CH23]. Minsquare [Mur06]. Minus [Gly10]. Mirror [CL11]. Misère [KLW89]. Mixed [BJFJ95, BH13a, CWYZ10, IM96, IT09, RV99, SW23a, GJW16]. Mixing [AK22b, BPS18, BP17, CDW07, EMT15, KS02, MS24, RS22, SYKY18, Win16, GKRS15]. Mixture [MM15]. Mixtures [MM15, Sul12]. MMSNP [BCF12]. Mobile [BNRT17]. Möbius [CFM^{+}09, JM97, Lin97]. Mod [GVW06, LL14b, LHHL18]. Mod- [GVW06]. Model [ART14, AFG21, BMRT20, CHZ04, ELMS11, FGP10, GGHP22, GMRT11, GKNU10, HLP^{+}23, HL10, HKMY22, LN21, LNNW95, MSZ10, SW10, TKA18, Vaz13, VVY15, Vod21, GKRS15, SU89]. Models [BDT17, BGG^{+}20, Che24, Fer24, FTS24, Mar20, MUWY18, WDSH21, Zho09, Bie88]. Moderate [Lev22]. Modes [Bon91]. Modification [ASS17]. Modifications [LSO03]. Modifying [TKMM19]. Modular [BCdMR08, EHV18, Kap14a, OW22a, PSWX24, Yam05]. Modules [CCNP24, HMSZ24]. Modulo [CLLZ18, FGRZ21, HLL^{+}21, MS16]. Molien [ACM11]. Moment [BMM20]. Moments [Bar95, ERS19]. Mömke [New20]. Monadic [KZ18]. Monochromatic [AHH^{+}10, AHP19, BMM20, BL01, BCF^{+}20, CH13, CH17, CM24, CFP11, ELR98, FS09b, FM13, KS23a, KMMS21, LR07, LSS17, MP08, ZGH^{+}23]. Monoid [BS95a]. Monoids [GR17]. Monomial [BPEY22, CK08a, GGI07, JL20b]. Monomials [Din13]. Monotone [AH21, ARS17, AcRS07, BDL21, GM91, HH17, HR12b, HKMY22, HRS93, JS14, KW90, ST13, Suk13, Yos19, GM93]. Monotonic [DSZ05, GHM10]. Monotonicity [BGJ^{+}12, GKW19, HY89, YH88]. Monte [SV11]. Moody [TW12]. Moon [CWYZ10]. Moore [Lat23, MP22]. Morphisms [CI07]. Morse [JP06b, SvM08]. Moser [Peg14]. Most [CvBdJdVK22, CNG19, DD13, HZ10, HV17, JN16, JMW17, KK14a, San96, Zho88]. Motif [FKW10]. Motion [ATPRU91].

Motions [GGLS21, SST22]. Motors [JSZ15]. Moves [BHNP16]. Moving [ABY14]. MPS [PV10b]. MR [GM93]. MSRD [MP22]. MST [Sol12]. Mubayi [MB18]. Muller [BGY20]. Multicandidate [Bol90]. Multicast [CH06a]. Multichromosomal [BH13a]. Multicolor [BZ23, DP17, JKSY22, SWR12]. Multicolored [AAFL06, Con05, KO06b]. Multicommodity [Fle00, Gün07, RS10, YCH97]. Multicovering [HK02]. Multicriteria [JL20b]. Multicut [KPPW15]. Multidemand [BBKS23]. Multidimensional [AJM08, ARS95, Buk12, CDF08, GL15, Kit02, OR04]. Multiflow [Fuk16, Hir11, HI23, FKS97]. Multigraph [Gab04, Gab05b]. Multigraphs [GM13, HLR13, KY18, MT20, NI92, NK90, WH04, CCM95]. Multilevel [LZ05b]. Multimatroids [Bou97]. multinomial [HY89]. Multiobjective [BP09]. Multipartite [CS21, Gut93, MMS17, Tót10]. Multipartition [Wag96]. Multipartitions [HRS93]. Multiparty [KOR03]. Multiple [BCE^{+}01, Bon91, BMN13, CS18a, CFR10, DGN^{+}20, FKW10, GK02, GZ98, HHLÖ95, PZ10]. Multiplexed [BCC^{+}05]. Multiplication [QP15, SR94, WB90]. Multiplicative [Shp22, TW19]. Multiplicities [FKK20]. Multiplicity [DSST13, HS04, RY91]. Multiplicity-Free [RY91]. Multiplihedra [FLS10]. Multiply [GHM14]. Multiprocessor [BLMS^{+}00, Tov90]. Multiset [OS16, Yen94]. Multiset-Union-Free [OS16]. Multisets [Siv23]. Multislope [LPSR12]. Multistate [GMRT11]. Multisubgraph [BCC^{+}22]. Multitasking [ACG^{+}20]. Multitolerance [GM16]. Multitype [KKP15, MSK93]. Multivalued [FS12b]. Multivariable [MMR06]. Multivariate [DEMT22, vzGVZ13]. Multiway [CMM20, JPvL21]. Murnaghan [Wil16]. Musical [ABF^{+}14]. Must [Pin14]. Mutating [Eri96]. Mutations [LN21]. Mycielski [MŠT09].

Nagy [DS21a]. Nakayama [Wil16]. Name [ACL^{+}06]. Narrow [ES11, Li17]. Narrow-Sense [Li17]. Narrow-Shallow-Low-Light [ES11]. Nash [AGM23, CGV^{+}14, HSV23, Vaz12]. Nathanson [CY12]. Natural [BL23, BCF23, MPSZ19, dM07]. Navigability [FLL10]. Navigating [DM11]. NC [GSK91]. Near [BFH^{+}08, CGOvZ21, CY20, CHHM09, DP16, Han16a, HK13, JLR20, LT11, MWvZ11, Pon20, dLL09, Zho88]. Near-Bipartite [CY20]. Near-Extremal [Pon20, dLL09]. Near-Linear [JLR20]. Near-Popular [HK13]. Near-Regular [CGOvZ21, MWvZ11]. near-traceable [Zho88]. Near-Unanimity [BFH^{+}08, LT11]. Near-Universal [CHHM09]. Nearest [BS09, Tas97]. Nearest-Neighbor [BS09]. Nearly [AKS08, DJ11, Gij05, GK13b, HH92, HZ08, LYY21, Meh12, NDB07]. Necessary [TZ15]. Necklaces [AEPT24, GL15, JPZ21, Stu21, WS96]. Negative [HMM09, Wu09]. Neighbor [BS09, BK21, BK24, DGS96, DNR23, FFHJ94, Har10, KS19, Prz13, Shi12, Tas97]. Neighborhood [BFPP08, BV20, CFT93, DHL^{+}02, GPS19, HGY20, HC98, KPS20, MW90]. Neighborhoods [FRMPV15, dJMS16]. Neighborly [Nov18]. Neighbors [ARTV12, HW17b, MP04]. Nemhauser [BYHR10]. Nested [DS05a, EIJ^{+}12, ILM^{+}16, JL23]. Nestings [Kla06, PY09]. Nestohedra [Gru17]. Nets [CF16, DM13, AS97]. Network [ACLW18, ADHL13, ADH^{+}14, ASMF10, BKM08, BFLM24, BOS01, CS09, CV07, DM03, Fer24, FvIKS15, FLL10, GM90, HKS07, JM17, KYDN09, KNS05, KK09, LS95a, vBBC^{+}15, Bie88, BCW96, NNI97]. Network-Based [vBBC^{+}15]. Networks [ARS17, AU91, BCSK07, BYKKR18, BTU09, BAH10, BK90, BKL^{+}15, BL16, BH93, CHZ04, CH06a, Che09, CHY13, CDF08, DP92, DP96, ES98, GRR15, GV92, GVKSS06, GL08, GR99, GM91, GP91, Hay21, HHHH02, HBDS23, HPS96, HKK^{+}09, HL10, HJ94, HKS07, JLD^{+}18, JA16, KW17, KL92, Koc98, KP04, KK09, KM05b, Lu10, May96, PL94, Pip95, Pip06, SW01, UBHS93, XY15, Zho09, vIM18, Bal88, BBM90, FFP88, GM93, LS89, LP88, PSW97, Tam88, TH11]. Neumann [DKWL20]. Neural [Bal88, CJL^{+}23, HBDS23]. Neuronal [JA16]. Newton [FGPS20]. Next [EMM14, SS10]. Niho [dH04]. Nikodym [LSW18]. Nilsequences [CS14]. Nine [KSŠ08, SvM08]. No [BPV10, CH13, CH17, CtJL01, CL15b, DK06, GLSS16, KM19a, KMW06, Kim17, KST06, MS21a, MR12, Svi03, BM94c, GM93, KMPR14]. No-Hole [CtJL01]. No-Wait [Svi03]. Node [ARTV12, HI23, TT89, Vég11]. Node-Connectivity [HI23, Vég11]. Node-packing [TT89]. Nodes [XY15, BVW88]. Noise [AG06]. Noisy [ASMF10, GJ19, KM05b]. Non [Eng04, GM05, KY12]. Non- [KY12]. Non-Boolean [Eng04]. Non-Surjective [GM05]. Nonadaptive [HH01]. Nonadjacent [MS05]. Nonapproximability [Eng04]. Nonbinary [vIKLS14]. Nonbipartite [Tak22]. Nonblocking [CHZ04, Pip95, FFP88]. Noncommutative [BSS14]. Nonconstant [CW92, LS95a]. Nonconstructible [Lut04]. Noncontractible [PP90, Hut88]. Nonconvex [BFP12]. Noncrossing [AR04, CEFY24, KLN91, LL17, MW22, Rea08, Rea15, Tza08]. noncryptographic [Sak89]. Nondegenerate [CJL^{+}23]. Nondominated [MK01]. Nonequivalent [Etz96a]. Nonexistence [Etz96b]. Nonexpansive [Fed06]. Nonextendible [BCM^{+}12]. Nonflowerable [AD11]. Nonhamiltonian [ABHM00]. Nonhomogeneous [CDR16]. Noninclusion [RT98]. Noninteractive [KOS16]. Nonintersecting [BP20]. Nonisomorphic [Saw02]. Nonjumps [HLR13]. Nonlinear [BI05, BLMA^{+}08, CCD00, CCM^{+}15, LOW10, FMV22]. Nonlinearities [KMÖ18]. Nonlinearity [CC03, LS09, Pet11]. Nonmonotone [LMNS10]. Nonmultiplicativity [Zho92]. Nonnegative [AAH14, FHK96, KM21b]. Nonorientable [iKSZ04, KML05]. Nonplanar [CGH^{+}10]. Nonpreemptive [ILM20, PP07a]. Nonrainbow [DJKP09]. Nonredundant [Spi95]. Nonrepetitive [KM13c]. Nonseparable [Vaz13]. Nonseparating [CY03, CLY05b, EIiK08, Wag21]. Nonsuccinct [May08]. Nonsymmetric [CGV^{+}14, Vaz12]. Nonsystematic [PL99]. Nontiles [CM12b]. Nontrivial [AF10, CLWZ22, Ger23, NV22]. Nonuniform [ABC^{+}23]. Nonunique [BDP19]. Normal [BGM94, Bón09, BCE^{+}00, BS90, Gar92, HL15, HW17b, WS17b, LW88b]. Normality [Vod21]. Normalized [AE03, CDH^{+}04, WZ08]. Normed [Dew20, KS23a]. Note [Adr23, AD11, BPST22, CD11, DFX21, FHLT21, GHV06, GJ12, HS10, Hof98, HS06, dH04, HGY20, KS04, KL19b, LM21, LM22a, LT01, Mer99, Naa01, PV10b, Rab08, Sch91a, Sku16, Yen94, Zun11, HS89a, Sag88]. Notes [Juk21]. Notion [dCST20]. Notions [KLMN14]. Nowhere [ALZ96, FZ08, GKR^{+}18, KR16, LXZZ08, MS17a, WYZZ14]. Nowhere-Zero [FZ08, KR16, LXZZ08, MS17a, WYZZ14, ALZ96]. NP [CH23, DGM12, FRRS09, HM20, HN15, LZZ22, ST23]. NP-Complete [FRRS09].

NP-Hard [LZZ22, DGM12]. NP-Hardness [HM20, CH23, HN15, ST23]. NU [FHL^{+}13a, FHL^{+}14, LS18]. Nullity [BH13b]. Nullstellensatz [Prz13]. Number [AD11, Ale10, AADM18, Alo13, AAH14, ARS17, ACFL16, Ave13, AM06a, AM06b, AB07, ABHW13, ASSW21, BM19, BHLY08, BJHY03, BCS04, BCD^{+}12, BJL^{+}24, BC94, BF24, BDPR18, BFM06, BS15b, BELP21, BVdZ16, CW98, CWY21, Car09, CPRdS13, CCH21, CLS09, CR04, CHJ^{+}04, CP20a, Chi11, CHK10, CD16, CHW10, DMS21, DW22, DPRS10, DMS12, DŠ09, DFK^{+}21, EK13, Eli09, EHJ01, EJK^{+}09, FJZ15, Fle00, Fox10, FPS13, FW20, FHL13b, Fue23, FL10, Ful14, GHM21, GV92, GKPP18, GGM^{+}22, GR17, GWZ18, GK13a, GLS15, GH13, GHvtHP15, GPRT11, GJ06, GJ08, GLP23, GV21, Gyá19, HJ18, Har19, HZ10, HT18a, HLZ13, HP23, HW17b, HT93, Jan10, Jan20, JMN22, JY23, Jon05, JMU23, KKS19, KLWY21, KLMR13, KvIL^{+}12, KZ04, KMŠ^{+}09, KWZ13, KW08b, KK01, KS07, KM13c]. Number [KŠ08b, KSS09, LS22, Lav16, LiO15, LZ05a, Liu14, LY21, LY23, LHC90, LW18, LW20, LMW23, LS05, MS21b, MS16, MRAS19, Mot19, MNS14, NvdP19, Ngu24, Nor11, OZ18, PSWX24, Pin08, PS21, PSV08, Pit23, PR98, RMS01, SST08, Smo07, Sul12, Tak90, TH90, WWKY11, Wei19, Wil05, XZ22, ZKNS20, Zun11, dCLM13, vIKL^{+}16, CFGG88, Gor93, Jac92, Jed93, JLM93, KW96, Pit89, Ram97a, WGM95]. Numberings [AB00]. Numbers [ACD08, BV22, BCG^{+}24, BZ23, BPRS13, Bro11, BK14, CFT93, CX08, CFJ11, Con10, CPR99, DFG^{+}21, DF10, EA11, FP99, GLS20, HRS18, JKSY22, JO18, JS12, JQ20, KKW17, KPPR15, KLMR03, Kla06, KMS23, KW14, LLL18, LLY10, LP21a, LP21b, LL21, MSS14, MSD19, Mra17, NV22, OS17, PdCSC21, Sch92, SS91, SWR12, SS08, Sud08, Tót08, XSR11, dKMP^{+}06, dKPS13, Abe91, Exo89, LW88a, RM97]. Numerical [BP12, CGSM16, EA11, SS18a]. Nuts [KMS98].

Obeying [LM22b]. Objective [HW23, HR05b]. objects [Ko88]. Obnoxious [BDLR01, CR10, Tam91]. Observations [BHH94]. Obstacle [BCG^{+}24, MWvZ11]. Obstruction [CLiO18, FHS14]. Obstructions [BK12, CJL^{+}23, CGSZ20, JM97, Mon15, Ram97a]. Obtained [EGS13, Mou94]. Obtaining [CLM03, HVLP13]. Obtuse [LM22b]. occupancy [LW88a]. Occurrence [MRR20]. Occurrences [HS06]. Octants [CKMU14]. Odd [AK14, ABHJ21, BGY20, BCH92, CMM20, CFLZ19, CS18b, CCL^{+}06, CG07, DZ09b, DŠV08, FG22, FKS12, GHM21, JPvL21, KRS11, KiKK17, KSS12, LRT08, LMRS17, MN18, MS17b, RT18, Wan08, JLM93]. Odd-Girth [DŠV08]. Oddtown [SV18]. Off [AP92, HHLÖ95]. Offord [LR94]. Offs [CGP98]. Offset [GS95]. Old [DZ09b, MV99]. Oligo [ACLT01]. Olives [DEF19]. Ollivier [BCL^{+}18]. On-Line [AE04, BKS10, GKL99, HV00, KT99a, KPT94, McD15, Woe93, BCP08, Con10, KPT95]. One [BMS12, Che16, CK91, CHK17, DGS96, DK22, EMM14, Gly10, Kam19, MW90, Mra17, PSML08, SA90, UV15, Vég11, Wal19, NS89]. One-Factor [PSML08]. One-Page [MW90]. One-Player [BMS12]. One-Point [Mra17]. One-Sided [CHK17, Kam19]. Ones [PPTY22, Saf21]. Online [AB18, BBC^{+}19a, BCD21, BFKT24, BD20b, DSZ05, DE11, Eps06, EEL09, ELvS11, Ete22, HKST03, JK19, JSRSW18, KP00, Sei01]. Only [Kar20, Wan08]. Onto [FKK05, FKK07, KMP03]. Opaque [DJT15]. Open [HGY20, YZ17]. Open-Neighborhood [HGY20]. Operations [FHL^{+}13a, FHL^{+}14, KMT07, Orr08]. Optical [AM07, FHMY01, GL08, KSV05]. Optima [GK02]. Optimal [BSKS11, BCG^{+}10, BDDS03, BLR16, BLR17, Bon91, BG12, BC17, CHW88, CZOS98, CCG^{+}00, DDS16, DH90, DHJN02, DJS12, EJK^{+}09, EH13, Fei10, FM24a, FO00, FPR18, FHMY01, FS12b, GP08a, GMW05, GH90a, GLY07, GS95, HSS19, HZ08, HH17, HK05, HLZ13, HH01, HKST03, HRS93, JCB06, JP06b, KFHR94, KPS20, Kim91, KST06, KS12a, Lai05, LSL92, MG19, MR15b, Sei01, SR94, Špa07, SS89, Tak14, TZ97, Zha99, Ram97b, RX88]. Optimality [BHE05, EIH18, EH13, FP01, SP88, Mur96a]. Optimally [CM05b, GM03]. Optimization [AOW15, BL17a, Ber20, BLMA^{+}08, BP09, BCH92, BHM00, BS18, CMV10, CC07, DLMO18, DO08, GJ19, HR05b, JL20b, Jüt06, LOW10, Onn03, RT98, ST07, SSS13, Sim90]. Optimizing [GT13, MP21, Ten09, dFM04]. Optimum [OKS06]. Oracle [HK14]. Oracles [BKYY23]. Orbits [Shp10]. Order [BGV07, BCD^{+}12, BDF^{+}18, BT93, BR20, Che07, CKNV16, CW96, CG17, CFDK20, DF02, FKT06, FL92, Fel14, FMP17, GB12, GJK^{+}23, GMS15, HKT20, JS14, KP09, KZ18, LSX14, LMP19, LMS19, MR21, Pet13, Abe91, BT97, CET97]. Order-Isomorphic [BR20, GJK^{+}23]. Orderable [HHMR20, HS88]. Ordered [ATPRU91, BV22, BK23, CL05, Elb09, Fis90b, GT98, GvdHM^{+}08, HLS24, IKZ08, Rif99, SSS22, XY15]. Ordering [BKG99, Fio06, MSK93, OS92, iO08, Rio98]. Orderings [BKS09, Che98, CKOS06, HKL99, HR12b, HHL24, MR21, Sch04, Wag18, RW89]. Orderly [Lu10]. Orders [ANP91, Boy01, CM05b, EGS13, HH13, HRS93, JZ05, Mer15, Sch91a, BB97, SSSU89]. Ordinary [BVdZ16]. Ore [CNG19, FZ08, KOT09]. Ore-Degree [CNG19]. Ore-type [KOT09]. Orientable [CMS09, iKSZ04]. Orientation [BCKN21, CKN13, KNS05, LHHL18, SHS23]. Orientations [AAS22, AH16, CHM^{+}23, DKM^{+}15, DE93, GkoKK23, HLL^{+}21, Iri16, Lai08, LL14b, MRST16, PY90, Bal88, RX88]. Oriented [AGH11, BCF^{+}10, Che24, CKP22, DLMT21, Fed01, HZ95, Zho92, Zho93]. Origins [Pet11]. Orthogonal [AM07, BGS96, BÖ05, DJM^{+}18b, FHMY01, GB12, NDB07, Spi19, ZN08, Bal88, DJM^{+}18a]. Orthogonality [CGG17, CH23, GN08]. orthogonally [MRS89]. Osthus [MMS17]. Other [FPS18, GKI^{+}23, GRS11]. Out-Community [LN21]. Outdegree [CH17]. Outerplanar [CGN^{+}06b, LZ05a]. Outerstring [RW19]. Output [RF12]. Overfullness [CCJS22]. Overinterval [CM05b]. Overlap [Jan00]. Overlapping [BH20, WS12].

Packable [GŻ10]. Packets [DP96]. Packing [ACZ23, ALM^{+}18, AB18, BF12, BP13, Bar04, BSKS11, BCD21, BDGP22, CLM03, CDM^{+}14, CK14, CQX20, CCG^{+}00, CG07, DR19, DFJS15, DKMS17, EIH18, Eps06, FKT06, Fis94, FL91, GMPZ15, GMW96, Gu18, GL17, Har18, HPS19, Jan05, JSO12, JLL16, JK19, KiK12, KiKK17, Kir16, jKR21, LMSZ19, NY23, Sei01, Woe93, Yam16, Yus24, Zak14, dGNS13, vBJU20, BCD97, HS89b, TT89]. Packings [ACGT22, ACP22, AMB11, BT96, CCG^{+}00, KT17, KS12a, KOT09, Nag17, Ric14]. Page [EM99, JMU23, MW90]. Pagenumber [VWY09]. Pages [Obr93]. Paint [XZ22]. Pair [BJvV92, DSL19, DK02, FHH08, FGK89, GK92, HK11]. Pair-List [FHH08]. Pair-splitting [BJvV92]. Pairs [BS24b, BDT17, Bon10, BCGR24, Boy01, BM13, GPW13, HM94, LRWZ12, MO22, PSV08, SSS22, SSS24, SV20, Alt89]. Pairwise [CE06, KV15, KSY18, KM06, MMSW23, RN23]. Pancyclicity [HL10, KLS10]. Paradox [LRTW11].

Paradoxical [GGLS21]. Paraffin [ACM11]. Parallel [AC14, ALSY11, AM96, BB90, BGL24, BC09b, BH15, Cho92a, Cho94a, Dah93, EP10, ELSS17, GPS88, GS93a, HS90, HHH^{+}02, JSZ15, KR16, KFHR94, Kim91, KN95, NS89, PP07a, RS15, SSS13, ZN08, BM97, DL89a, GS89, LM89, RSSW88]. Parallel-Task [KN95]. Parallelisms [AAFL06]. Parallelohedra [SGM20]. Parameter [AKP^{+}22, ADKS18, BCDF19, CPPW13, GJ19, HN15, KPT22, KPPW15, Kra18a, KS12b, LSSZ19, Vod21]. Parameterizations [BJS22]. Parameterized [BLM20, BBRT23, BFPP15, BKP21, BCKP19, DFK^{+}21, FLM^{+}18, FKL^{+}19, GKI^{+}23, GKMT17, GNS11, GSZ22, GJW16, GP16, GMOW21, JLR^{+}17, NPSW23, Neu24, PRS^{+}23, SW23a, Sli10, Zeh17]. Parameters [DFHT04, FGT11, GKI^{+}23, MP17b]. Parametric [BGW20]. Parametrized [AB95]. Parametrizing [She18]. Parent [BCE^{+}01, BEN08, TSN04]. Parents [BCE^{+}01]. Pareto [Kam19]. Parities [MO22]. Parity [BM15, CF09, HT18a, MO22, SMR08, Stu21, Sza08, Yam16, Zha93, CHW88]. Parity-Check [CF09]. Parity-Unimodality [Stu21]. Parking [CEFY24]. Parsimonious [WH04]. Parsimony [FvIKS15]. Part [EFK05, ACM^{+}18, CCG^{+}00, FHL^{+}13a, FHL^{+}14, Gao18, Lla06]. Partial [BYFMR10, BP09, BHTK^{+}21, Bor10, CMPS17, Ege10, FKM^{+}06, FL92, FMP17, FY04, GSK91, HS22, Hor14, HRS93, JJL23, JT11, JS14, KPS20, KM02, Koc98, PSW96, Sch22, WWY23, dGV05, SSSU89, TP97]. Partially [BFRS16, Elb09]. participant [CGG88]. Particles [BDvL13]. Partite [Fed01, Fue23, GKP19, GSS15, LWY18, LLZ22, RRS07]. Partition [BJGJS99, BD20a, BSY21, BGS17, CEHS08, Can93, CS94, DF02, HR05b, KKSvL20, Rea08, Smi01, KS88, MM96]. Partition-Optimization [HR05b]. Partitioning [AD96, ABL^{+}20, BKS10, BCF^{+}20, Cho94a, Gal18, HK99, LSS17, MLS11, Sot15, BVW88, HM88, TP97]. Partitionings [BDM02]. Partitions [ASS09, AM11, AR04, BL01, CX08, CGG^{+}16, COL10, CY21, DG08, FHKM03, Fel90, GS94b, HW18, HMSW14, IKZ08, Kim11, KR93, LV89, Len98, Lin97, MPP20, MW20a, Mol11, Onn91, PY09, RY23, Rif99, Smi01, SZ15, Tza08, Zen90]. Parts [AHS01, HS04]. Party [KOS16]. Passage [HH92]. Passing [MT11]. Path [AFL^{+}20, AHP09, BJT92, BMV22, CL15c, DNS94, De 24, DD15, DD13, Erd20, Gao15, GMS15, Gut93, Hli18, KM19a, KM05a, LW17a, LMM21, MMP13, MC12, SB94, ZZBL17, sG24]. Path-Decompositions [Erd20]. Path-Tough [DNS94]. Path-Width [AHP09, DD13, Hli18]. Pathology [AL95]. Paths [AK10, AHJ^{+}24, AP14, AG15, AS03, BNN90, BPV10, BCD^{+}12, BNRZ23, BK90, BDL21, BZ23, BCF^{+}10, BHJ18, BK14, CFJ11, Chi23, CGH^{+}21, DR19, DSS92, DS23, DG08, EJ01, EFK18, ELR98, EH13, FJ09, FIN98, GPvL15, GLS20, GZ98, HW10, HT19, Hng22, HMM^{+}21, HMP04, iKK11, KMS12, LL14a, LL17, LZ05b, LW18, LP19, MT90, Mor21, MNPR17, NT12, RS14, Rom06, Sch91b, Sli10, Špa07, Sta92, Vin07, Yam16, ZWW23, KP06]. Pathwidth [ALQ^{+}24, BKL^{+}22, BM93, BKK95, Der12, DEJ^{+}20, GJW16, KPT22, Kas08, KMRR09, sJS24]. Patricia [RJS93]. Pattern [AF19, CFKK17, FKLL15, GPP04, MRR20, DK89]. Patterns [AL17, AM06a, BK05, CT21, DZ09b, FH21, FW02, FHL20, JSZ15]. Patterson [KMÖ18]. Paving [GH06]. Payment [PSTF00]. Payoff [AGQS23]. Pebbling [AGH14, BC09a, Chu89, CLST12, CH06b, MC06]. Pedigrees [Tha08]. Peeling [HPL13]. Penalizing [AB05]. Pentagons [BDJ^{+}15]. Peptoids [LW03]. Perceptron [Gol06]. Percolation [BP15, BW09, Cor24, DNR23, FJP22, GKM^{+}18, HM20, HL24, MdCW16, Rah16, SS91]. Perfect [ABY11, BKS09, BH22, BL04, BS16c, BE13, CGL10, CY18, CCG^{+}00, CL15b, DM19, Etz96b, EV98, Fan20, FKT99, GMP22, GS98, GH90b, HPS09, Han16a, Han16b, Han18, Hed08, HL00, HR05a, HXZ18, IKK^{+}22, Kha13, KSS09, KO06b, KOT09, LGS11, LWY18, LYY21, LWY22, LLZ22, MWWZ20, MMSW23, MS21b, MWW94b, MP98, MRS89, Pen12, Rif99, SW23a, Vaz13, dCLM13, dFM04, Etz96a, Mit97, PL99]. Perfect-Matching [GH90b]. Perfection [BS10b, BF17, Sch02b]. Perfectly [LM08, MT05]. Performance [AB95, Fra89, PV10b, KK89]. Period [BCG^{+}10, XQ06]. Periodic [BSiT15, CS14, MK09, SU89]. Periodic-Finite-Type [MK09]. Periodicity [FPST06, GPP04, KK09]. Periods [MK09]. Permanent [CV09, DJMV21, Eti20]. Permanents [BMV92, Vin12]. Permutation [AMNV18, BT18, BKK95, BCF^{+}10, CHM^{+}23, DQW^{+}15, Ehr16, EST14, HvtHLN12, HS22, HKK^{+}09, KL92, WZZ18]. Permutations [ALSY11, AFK^{+}22, ACG94, BHNP16, BCdMR08, Bón09, Bor10, BCPP09, BR20, Cap99, CR16, Cib13, CFK19, CKdAHdF13, DR04, Eli09, EFK05, FKMS10, GJK^{+}23, KMS08, KS19, Lab13, MRR20, MR21, MS24, Ram98, Sav90, Vse05, Yen94, KM06]. Permute [QD14]. Permutohedra [MUWY18]. person [KLW89]. Perspectives [ABNR22]. Perturbations [KK09, Lin89]. Perturbed [AHP19, AHH21, AHDHL22, CHMP21, KKS17]. Petersen [LB09, DŠV08]. Pfaffian [MO22]. Phase [CMR18, CFKK17, DH05, FMP08, GKRS15, KKP15, LN21, MMPPG23, RSV^{+}14, VVY15]. Phenomena [BER11].

Phenomenon [BMM20, CL11, LP19, Stu21]. Phylogenetic [ART14, ANS16, BL17a, BS15b, FvIKS15, Hay21, HMS05, KL19a, SV11, Sul12, dJMS16, vIM18, Ste88]. Phylogenetics [Ber20]. Phylogenies [DMR11]. Phylogeny [LGS11, SW23a]. Piecewise [WWKY11]. Piercing [ABFV23]. Pillar [KT19]. Pipe [Més16b]. Pivot [DDJ^{+}21, LiO23]. Pivot-Minors [DDJ^{+}21, LiO23]. pivoting [LM89, Vav89]. PL [ST23]. Placement [BH05, MLS11]. Placements [SV20]. Planar [AH03, ABHM00, AST94, AHP09, ACD08, BB16, BBF^{+}23, BCC^{+}11, Bon15, BGP22, BDDP24, BS93, BY08, BJ23, Cai93, CTU14, CEP18, Chi23, CLiO18, Cra19, CL20b, CL21, CLY05b, DV96, DSS92, Dji06, DX19, DM18b, DŠV08, DLŠ10, DM17, DH20c, Far09, Fel14, Für91, GW99, GGM^{+}22, GKMS22, GHP20, HM94, HM12a, HKL99, HW24, HMM^{+}21, HOD22, JPvL21, JMW17, JMU23, KMM12, KM23, KFHR94, iKM10, KW17, KPP13, KNK93, KMŠ^{+}09, KL92, KSS12, KR20, LM17, LL17, LLS04, LY21, LQ22, Lu10, LSSY10, MP94, MRST16, SMR08, Sch91a, Sch91b, Suz10, WL02, WL03, WX13, WHW14, WH15, WY10, YWWL21, ZLWC12, BCLR89, BT97, FPS13, GI97, DL17, DL18b]. Planarity [DGL10, DEK22, DP15, HKSS08, Sch22]. Plancherel [FP13]. Plane [BBS17, BFP12, BDT17, Boy96, Brä10, Cha91, CGM^{+}15, CEOT15, CR16, CEOT17, dOCHLO21, DEG^{+}07, DJ11, DM15, EHJ01, FHJV17, FT17, GZ19, GPS19, HST22, HSŠ08, HZ08, JKSW17, KSS11b, KSŠ08, KS23b, MZ22, MS14, Nog24, ONN19, Pay17, PRS03, PSS09, PP07b, Pin08, PSW96, SW04, SW99, Xu09, CCM95, RWW88, Ser89]. Planes [Bal08, Dew20, GKM^{+}18, GB12]. Planning [ATPRU91]. Planted [COL10]. Planting [ZK11]. Plates [DEF19]. play [KLW89]. Played [KMN23]. Player [BMS12, Vaz12]. Playing [DTW03]. Plethystic [Wil16]. Plex [GKNU10, MH09]. Plummer [CEOT15]. Plus [PSML08, BC88a]. PM [DKWL20]. PM-Compact [DKWL20]. PMDS [MP22]. PMU [BH05]. PoA [ÀB23]. Poincaré [FG22, FP13]. Point [AHH^{+}10, ATPRU91, AS16, Ave12, BDJ^{+}15, BH16, BDG^{+}17, BV10, CF17, Dan09, DSL19, DGP06, Kap14a, Mat19, MGC14, Mra17, SS95, SFS09, HY89]. Point-Sets [DSL19]. Points [AS07, ARS17, BS16a, BS16b, CI07, CKP16, ES11, FGPS20, Fed06, GRR15, HHLÖ95, JB22, KML05, Lag00, LY18a, MPŠ08, NP18, Pin08, PR98, VZ93, dH89]. Polar [CH11, DSL19]. Polar-Pair [DSL19]. Polarity [TT16]. Policy [CZOS98]. Pollak [WIS12]. Polya [FP13]. Polychromatic [BPRS13, Off08]. Polygon [HK96, KvRRS23, MvLvL13, PR98]. Polygon-Guarding [HK96]. Polygonal [ASH18, CDMT21]. Polygons [GGS17, Lub90a, SS09, Fra89]. Polyhedra [AC90, BBC11, BDL21, BFH15, BC95, Cho92a, Cho92b, CPS08, CL90, EM20, Gij05, GM90, HKW22, HMP04, NP18, Shi12, BM94a, BM94b, BM94c, BFM94, Fuj97]. Polyhedral [AOW15, AT16, BP16, DHS14, NS93, Yam05, vWW94]. Polyhedron [Bih05, CR96, CPS08, Tod14, Yam07, Cho94b]. Polyline [AFT12, FT17]. Polylogarithmic [EK05]. Polymatroid [BCF23]. Polymatroids [BL23, OSW16, Sav14]. Polymer [WDSH21]. Polymer-Based [WDSH21]. Polymorphism [BM22]. Polymorphisms [GS20, LS18, LT11]. Polynomial [AP92, BJT92, BFRS16, Bon91, BJS22, BC92, Bru90, BS16c, BG19, CDV10, CKN^{+}15, CP16b, DM13, DN16, EGR08, FSV13, FKL^{+}19, GLSS16, GnN06, HKL^{+}14, IMS05, IT09, JMS90, JP06a, JK19, JPvL21, KN16, Kra18a, KS03d, MT90, MPSV21, MC12, Mer15, NPSW23, PS22, ST07, SWKP10, Svi03, Tar88, Zha90, dGV05, BC88b, PRS88]. Polynomial-Factor [KS03d]. Polynomial-Time [CKN^{+}15, JP06a, PS22, ST07]. polynomially [BS88]. Polynomials [Ath14, BB23a, Bar02, ÇMP15, DK10, DJ11, Ege10, FGPS20, FG22, GXY22, Jan00, Kim92, Més16a, Mun06, QD14, RT18, Rón92, Ros09, RMS01, Shp15, SW21, YZ17, Zhu18, vzGVZ13, FZ88, Mal89, Tsa96]. Polyominoes [CWYZ10]. Polytope [AHFM08, BBC11, BKM08, BKG99, BH93, Chi11, Cho94a, DF02, Fis14, FLMY09, Ger23, GS03, GMTW19, MRV17, MH09, Mic21, PVUY18, RC98, XY15, Bal89, Sar97]. Polytopes [ACF18, ABH^{+}11, BB09, BB23a, BS16a, BS16b, BW02, BT93, CFS17, CdV02, CG17, DJKKV23, Dan09, DST01, FGPS20, FFLP20, FLM12, FP22, GS93b, HL15, HKT20, HS22, JK99, JB22, JS14, KPS19, KT16, LM22a, LMS19, Més16b, NLS24, Nov18, Pad16, PVY22, PVS22, Set21, SL95, Tsu23, UV15, Wal19, BM97, Stu88]. POMSET [FKL93]. Popular [CHK17, CFKP22, HK13, Kam17, Nas14]. Popularity [Kav24]. Posa [LMSZ19, AFH^{+}18, BFH21, BHJ18, BJS21, HjK24, KO12, vBJU20]. Poset [ABC11, BMVW17, CPP23, CFS17, FFLP20, FR99, Hyu10]. Poset-Isometry [Hyu10]. Posets [BMNP22, CL11, DFG^{+}21, DJW12, FM11, Fer16, GGMM21, JMW17, LZ03, MW22, PR91, Sta92, BW92, HP97]. Posimodular [SMNF09]. Position [BOT92, PW13]. Positive [BLP21, Sim13, Spi23, Tro15]. Positive-Definite [Tro15]. Positivity [BI13, CH19, Zhu18, Zhu21, FZ88]. Possible [Boy01, HHCHL21, HV00, KMÖ18]. Postman [DMNW13, GJW16, RV99, SL96, Zha93]. Postnikov [Han09]. Potential [EFMW18]. Potentially [FS09a]. Potentials [Che17].

Potts [BGG^{+}20]. Power

[BI05, BG12, Cal13, CDHK16, DMKŠ08, DHL^{+}13, ELSS17, FLM22, HHHH02, Ja10, KN22, LSX14, LSS24, May96, NT21, Red17, SB91, BP89]. Power-of-Two-Choices [Red17]. Powerful [vzGVZ13]. Powers [AH03, AL07, BMPS21, GMS15, Hng22, KP16, Krá04, LC04, MS16, PPU92]. Practical [ZM02]. Pre [GMOW21]. Pre-Coloring [GMOW21]. Precedence [Tov90]. Precise [CHKW24, KPPR15]. Precision [Gol06, HSS24]. Precolored [DL17]. Precoloring [AKW05, Voi07]. Precolorings [KKK^{+}12]. Predicates [Eng04]. Predicting [BHH94, WS19]. Preemption [ELSS17]. Preemptive [DE11, JP06a, KN95]. Preferences [CHK17, HW17a, FK97]. Prefix [CM03, HKW15, HvIK^{+}07]. Prefix-Free [HKW15]. Prefixing [DA10]. Preimages [Zho93, DS97]. Preprocessing [BJK13]. Prescribed [AS09b, BGV07, Dvo05, DG08, KŠ08a, KiK12, KiKK17, Kas05, Lub90b]. Preservers [BCE05, CE06, KV15]. Preserving [BJFJ95, BJJ98, BDD^{+}19, CK14, Che98, DHL^{+}02, Fle05, JS03, KNZ14]. Prestress [CW96, CG17]. Preventing [BKL^{+}15]. Price [BILS22, Vaz13, GK97]. Pricing [BKK21, BBKS23]. Primal [BYR05, CMSV17, FMV22, Köp07]. Primal-Dual [BYR05, CMSV17, FMV22]. Prime [BGY20, Gly10, Gly12, LSX14, MWW21, MS16, MR04a, NW22, SST08, Spi89, CET97, FH89]. Prime-Power [LSX14]. Primitive [Li17]. Principal [IM96]. Priori [MLS11]. Prioritized [EN22]. Prioritizing [AR17]. Prism [CK08b]. Privacy [CK91, FY04, Kus92]. Private [FFHJ94, HKS07, KR98, KOR03, KOP^{+}21, KM97]. Probabilistic [BMS12, BAM16, EP10, FKW10, HKW15, KS92, MRR20, MUWY18, Nao91, STV21]. Probabilities [Fis90a]. Probability [AFKK23, Bar95, BM11, BB90, COHK21, DK22, DJS12, EGS13, Gol96, GJ16, JK24, Jun12, Lyn94, Pip06, RN23, YAT16]. Probe [BGL07, MN15, PW02]. Probing [HSS24]. Problem [AFG^{+}16, AYZ04, AR08, AFKK23, AS14, BBF99, BJT92, BNMN92, Bar01, BDK^{+}21, BAH10, BKM15, BM22, BKL^{+}15, BKM08, BN96, BKS10, BH13a, BCC^{+}22, BHM00, Bra05, BH05, BDLR01, CCVZ10, CEHS08, CFKK17, CKNZ05, CGN06a, CDHZ12, CY12, CKN^{+}15, Che04, CCDZ23, CHY13, CDK^{+}18, Cho92b, Cho09, CL91b, CPPT20, CM12a, dOCHLO21, CM07, Däu22, DI22, DŚN21, DFJS15, DF04, DFX21, DGM12, EK05, EJ01, FMP17, Fio06, FOST10, FS09b, FYK00, Fuk16, GDVL17, GS94a, Gao15, GPP01, GJXZ20, GL95, GRSZ23, GMS15, GRS12, GW19, GP16, HM94, HW10, HLO17, HK13, HW23, HN20, IK09, IT08, JWF05, Jan22, JPT12, KYDN09, KP00, Kam22, iKK11, KS00, KR92, KKM94, KS03c, Krá06, KM05a, LZ03, LW10, Lev15, Lev09, LW03, LW17a, Lon21, LSL92, LHC90, Lub90a, MWW20]. Problem [MC93, MPSV21, MPSZ19, MC12, Mur06, MRNS15, MRNS17, MW03, Myu01, PSTF00, PW13, PSV08, Pry23, Qui10, RV99, RY22, Rya07, Saa93, SSW98, Sha23, SS91, Sku16, SS95, Str20, Vin11, WY23, WIS12, XG20, Yam05, YZ17, dBSS24, vWW94, Bal89, BG91, BC17, CK96, GW94, GHHY96, Hef97, IW91, JM97, Lin97, SS89, Tas97, Yen97, DRW98, DvHT20, GJW16, SS11b]. Problems [AMW00, AGM23, AM11, ABFRV24, ASS17, BLS17, BILL23, BK11, BGW20, BFN20, Cai18, CLM03, CKMU14, CKPW09, CS09, CCZ12, CMSV17, CC07, Cho92b, CKPV91, CP22, CCJ^{+}17, CKN13, DSS12, DJS12, EV98, FL92, FMV22, FK17, FK21, FJK^{+}23, GMOV20, GM16, GPR^{+}21, GT12, GMPZ15, GM90, GNS11, GMOW21, HR12a, HK96, HHH^{+}02, HJ94, Hua14, HC98, JLL16, Jor03, Jüt06, KT14, KMR^{+}23, KKSvL20, iKX20, KLM14, KBE^{+}05, KPS20, KíP^{+}24, KWZ13, KKS22, KMV15, KS06, KJJ04, Liu24, LNNW95, LdCKM18, MO22, MT11, MRT11, NH91, PT14, RS10, RT98, RS16, ST20a, SA90, SSS13, Sid18, Sim90, SR94, Sno13, Som14, SV21, SS00, SV18, ZK11, Zha90, van94, BS88, BCD97, HS89b, NS89, PC97, RWW88, Ray94, SU89, Tam88, TP97, TT89]. Procedure [MP95]. Procedures [CLM03]. Process [BCD21, COPP12, CS94, Die10, Gao18, GMP22, KM23, KPS16, KOT16, Pic14, Red17]. Processes [AK22a, BBFP09, CDFR18, KS12b, ŁR91]. Processing [SSS13]. Processors [Jan10, PP07a, DL89b]. Product [AAHLT10, ALSY11, Ath14, BL19b, CCM^{+}15, CKPS15, DMKŠ08, KR13, KL08, MM15, MRNS15, MRNS17, PS17, Sha13, AHS89, LRN11]. Product-Type [KR13]. Products [BBM09, BIKZ05, BK07, CHM^{+}07, CHK10, CH10, Elb09, Fis94, GMS00, HW24, Lee17, Mat19, MM15, PZ05, RN21, Smi01, Špa07, WZ18, WY23]. Profile [DR24, DH05, FKK20, GP08b, Ger23]. profit [PRS88]. Profits [JSO12]. Program [Dow91b, She18, Dow91a]. Programming [ASZ02, BGS96, BP12, CKL^{+}21, CKNZ05, GPST15, HK99, HTS18, Juk16, KG98, LP21a, PS17, PdCSC21, SA90, UWZ97]. Programs [AH21, BCCZ10, BBCZ11, Boy96, HKL^{+}14, HMM09, OS15, PPU92, Vaz12]. Progress [JL20a]. Progression [Het14]. Progressions [GP14, KS23a, PP22]. Project [AT16]. Projected [LRR21, TZ19, TZ20]. Projection [DF02, Gar92]. Projections [DGM12, FKK05, FKK07, Hof95, Hof98, KMP03, MPŠ08, NP18, PZ98, PP10]. Projective [Adr23, Bal08, CW16, Enc05, GKM^{+}18, GB12, LQ22, MS14, MP21, Nel15, Nog24, ONN19, PSS09, Vse05]. Prolific [BEN08]. Proof [BL09, CK11, CS91, EMT15, FGS19, Gu21, HK11, HKP^{+}17d, dH04, HKS07, IMR14, Kar20, Lic14, Sig10, Wil16, HK16b, Yen97]. Proofs [Bab92, EFK05, Loe10, MPP17]. Propagation [AS09a, BBCZ11, PS17]. Proper [BHH96, BFPP15, BMVW17, CFG^{+}21, CT21, FSV13, HH05, HR12b, LC04, Saf21]. Properly [KY18]. Properties [AC90, ACFL16, BMPS21, CdVL11, CM05a, CGMV19, DST01, DP17, EGR08, FMMO93, FPS20, FPP22, GZ19, HST19, HKL^{+}21, JLD^{+}18, NS93, OPVV14, RRS07, ST13, Shp22, SW98, TKMM19, ZZ92, dLL09, CCM95, HM88, HY89, TT89]. Property [AcRS07, BBC11, BCE^{+}01, BEN08, Brä10, BFH21, BHJ18, BJS21, CHX15, CM24, HjK24, KS23b, TSN04, WZ08, Way01]. Prophet [EHLM17]. Proportional [NLJM21, PT90]. protection [KG93]. Protocol [ACMW17, Sak89]. Protocols [FP04, KOS16]. Prototype [MTGK05]. Proximity [GK13b]. Pruning [DMR11]. Pseudo [AKP20, HSLd88]. pseudo-Boolean [HSLd88]. Pseudo-Disks [AKP20]. Pseudocircles [ARS21]. Pseudodisc [Pin14]. Pseudodiscs [Pin14]. Pseudoforest [PRS18]. Pseudoforests [FHJ^{+}10]. Pseudomoments [KM24]. Pseudorandom [KLS10, KO06b]. Pseudorandomness [OS11]. Pumpkins [JPS^{+}14]. Pure [SSS22, SSS24]. Pursuit [BMP13, DKSŻ10, IKK06, SS89]. Pursuit-Evasion [DKSŻ10, IKK06, SS89]. Push [MP17a]. Push&Pull [ACMW17]. Pushing [KKSvL20]. Pyramidally [van94]. Pyramids [PW18].

Quadrangulations [Nog24]. Quadratic [BCH92, Fis14, HK99, LS05, Wal19]. Quadratically [BM07]. Quadric [RS16]. Quadrilinear [Mac18]. Quadruple [FHS14, GMZ09, JZ05, EH91]. Qualitative [EGS13]. Quantifiers [MMS15]. Quantitative [CHZ18, IN22]. Quartet [ASY14, ANS16, SY11]. Quasi [AGH11, BJHY03, BGÖ17, CM90, Che22, CH19, CKPS13, CT93, CPR99, FPS13, GSPRM91, GHV06, GÖ12, KPPR15, MR21, MSS14, iO08, SMR08]. Quasi-Cyclic [BGÖ17, GÖ12]. Quasi-graphic [Che22]. Quasi-line [CKPS13, MSS14]. Quasi-Order [MR21]. Quasi-Parity [SMR08]. Quasi-Ramsey [CPR99, KPPR15]. Quasi-Random [AGH11]. Quasi-Symmetric [CM90, CH19]. Quasi-transitive [BJHY03]. Quasi-triangulated [GHV06]. Quasigroups [KP09]. Quasirandom [FH10, GkoKK23]. Quasirandom-Forcing [GkoKK23]. Quasisymmetric [Gru17]. Quaternary [TRV03]. Quaternions [BL19b, SS08]. que [KvIL^{+}12]. Queries [DHL^{+}02, OPS21]. Query [ABFRV24, HKL^{+}21]. Question [JO18, Naa01, Tha08, CP96]. Questions [BGS24, Dow88, FM24a, NT05]. Queue [GSV12, HP97]. queueing [FG89]. Queues [HLR92]. Quick [Ser88]. Quickly [CFG^{+}15a]. Quiet [ZK11]. Quorum [PW02, HMP97]. Quotients [CW14b, OS11].

r [CF22]. r-Out [CF22]. Rademacher [DK22]. Radical [RY23]. Radii [HK02]. Radio [EK05, KP04, KM05b]. Radius [BLR16, BLR17, BR17b, CH01, GCH22, KZ04, LL99, RR03]. Rado [DT16, DK16, Han19b, MS14]. Radon [Con89, DV04, Fil89, Onn91, Sca03]. Rainbow [ABHJ21, ABC^{+}23, AHH21, AHDHL22, Bah23, BFa24, BJL^{+}24, CHLT24, dOCHLO21, CP20b, DX24, DFT15, FKMS20, FM13, GS20, HLO17, HKK24, Jan22, KY18, Kor18, KLS18, LY18b, LW20, LWY22, LMW23, OYY13, WY23, sG24]. Ramanujan [Mor94, TZ97]. Ramsey [AL07, BV22, BLS17, BCKN21, BMS12, BBC^{+}23, BZ23, BCG22, BCGR24, BGS24, CET97, Con10, CPR99, DI22, DP17, EMRPS14, Exo89, FS09b, GJXZ20, GMOV20, GT12, GLS20, GLP23, HRS18, HST19, JKSY22, JO18, KLWY21, KPPR15, KSV05, Kie97, KKL19, LP21a, LP21b, LL21, LW18, LW20, LMW23, Mor21, Mot19, NP20a, NP20b, NV22, OS17, Pik03, PS20, RW89, RRS07, RS08, SWR12, Sud08, XSR11]. Ramsey-Type [FS09b, RW89]. Random [AZ22, Ald90, Ald22, AK22a, AEPT24, AGH11, ADL13, AP14, AF19, AFG21, AS16, BBLM13, BBLM16, BHRZ14, BF96, BTU09, BSKS11, BCD21, BF24, BK90, Bev09, Bev10, BG20, BILL23, BGG^{+}20, Blo10, BMP13, BPEY22, BGL03, BGS23, CMR18, CFH23, CF08, CZLfW05, CL06, CP16a, CP10a, COL10, COPP12, COLMS22, CM12a, CFDK20, CDKS21, CGH^{+}21, CF05, CFK10, CFR10, CF10, CEOR13, CF16, CTW93, CFS96, CV09, CSW17, De 24, DKRR12, DSZ23, DMP07, DJKO19, Die10, DP96, DEK22, DKS18, DH05, DFT15, DP16, DP17, DEF19, DS16, DS23, DJS12, EFK18, FT12, FO08, FFV11, FMP08, FJP22, FK18a, FKK18, FKT99, FKS05b, FHL13b, FJ17, FPS18, FT20, FL15, FK18b, GW99, Gao13, GK07, Git99, GK13a, GLS15, GS94b, GH90a, GH90b]. Random [GJK^{+}23, Han19b, HLP^{+}23, HS04, HPS19, HN20, JP12, JT11, JL23, Joo15, JiT22, JA16, KKP15, KM23, KW13, KL14, KR04, KM19b, KLMR18, KMMS21, KW14, KPS16, Kri10, KLS10, KS12a, KMN23, KO12, KOT16, KZ18, KR15, Luc98, Lyn94, MMPPG23, MP13, McK19, MP17a, MP17b, MRT11, Mor21, MSK93, Mra17, NY21, NDB07, NT21, Ngu13, Pip89, PSV08, Pit23, Ric14, RRS07, Sch04, Sch92, SYKY18, SS02b, WS17b, Yam20, ACS97, BC88a, BB97, DS97, HS97, LR94, Kra07]. Random-Cluster [BGG^{+}20]. Random-Edge [GK07]. Randomized [FH10, IKK06, Kra18a, LPSR12, Osh21, SS02a]. Randomly [AHP19, AHH21, AHDHL22, BFa24, CHMP21, Fri21, KKS17, Zho88]. Randomness [CT93, KM97, KR98, KOR03, KOP^{+}21, MR15b]. Randomness-Rounds [KR98]. Range [dBSS24]. Range-Assignment [dBSS24]. Rank [ADL^{+}09, BBJ^{+}21, BKK21, BR17b, BR19, CDD^{+}15, CF22, DDJ^{+}21, DZ09a, Din06, FLM12, GRSZ23, JJiT14, Ja10, KS05, MPSZ19, MPS^{+}09, ÖV04, iO08, PP12, Sch09, TV03, DK89]. Rank-Based [Ja10]. Rank-Metric [GRSZ23]. Rank-Selected [Sch09]. Rank-Width [DDJ^{+}21, iO08]. Ranking [Alo06, CPR99, KRR16, McD15, CH89, PRS88]. Rankings [BDJ^{+}98, FKM^{+}06, Fis90a, KT99b]. Ranks [FK18b, GGS17]. Rapid [AK22b, MS24, Win16]. Rapidly [SYKY18]. Rate [BC17, CHZ04, LW10, RF12, WW91]. Rates [BG19, CCVZ10, FGLP14]. Ratio [BYR05, BCC^{+}05, BCG^{+}10, BDG^{+}17, BLM^{+}22, BDEK06, BDGP22, BNR96, CFG^{+}09, Gün07, PV10b, Tót10, WS17a, HS89b]. Rational [BB23a, BT18, DEH20, Rho15, Vaz12, Jae89]. Rationality [CDV10]. ratios [Kle89, LW88a, WW91]. rays [DGP06]. Re [Suz10]. Re-embeddings [Suz10]. Reach [MP04]. Reached [Che16]. Real [Bón09, EJK^{+}09, GJ06, GJ08, KŠ08b, MP21, YZ17]. Real-Rootedness [YZ17]. Reality [dAHFdFK10]. Realizability [FHJV17, KM21a]. Realization [CFM^{+}09, GMTW19, Win88]. Realizations [EMT15, Ost23]. Realized [Eli09]. Realizer [LLS04]. Realizing [AM96, BNCPR20, HN15, LXZZ08, PS97]. Rearrangeable [CH06a]. Reassembling [Vyg16]. Receivers [AR17]. Reciprocal [KR93]. Reciprocity [HZ21, JS14]. Recognition [Ban90, BDK^{+}21, BCHP08, CH11, CCL^{+}06, COS10, CD14, Dah93, HH05, LS17, MN15, Mer15, San96]. Recognized [BS16c]. Recognizing [BGL07, CH15, GSK91, HS88, HL00, LC04]. Recoloring [BBF^{+}23]. Reconciliation [SHS23]. Reconciliation-Free [SHS23]. Reconfiguration [IKK^{+}22, MNPR17, Wro20]. Reconfigurations [CDP08]. Reconnection [KL19a]. Reconstructability [LN21]. Reconstructed [MPŠ08]. Reconstructibility [PVS22, vIM18]. Reconstructible [PRS03]. Reconstructing [DGM12, GVKSS06, KK14b]. Reconstruction [ADM^{+}15, BM11, BVVW11, BGJ^{+}12, CCM^{+}15, MRT11, Ste00, SV11, Tha08]. Reconstructions [vD11]. Recovery [CR23]. Recski [Sza08]. Rectangle [BDM02, KS06, MPP20, SV20]. Rectangles [Lub90a, Fra89, RWW88]. Rectangular [CMR18, Lag00, MW19b]. rectilinear [Fra89]. Recurrence [EIJ^{+}12, Pip95]. Recurrences [ARS95]. Recurrent [KK09]. Recursion [AL95]. Recursions [ILM^{+}16]. Recursive [BPT91, CCG05, FMP08, FG89, KPT94]. Rédei [QP15]. Reduced [DEH20]. Reducible [BHTK^{+}21, Ram90, vzGVZ13]. Reducing [HP21]. Reduction [BGN15, BSY21, CK14, CKN^{+}15, Cho09, LHLZ23, Més16a, Tov90, Hli18]. Reductions [PV10a]. Reed [BGY20, BK91, CKPS13, Rab08, RR03]. Refined [GJWZ23, KR20]. Refinement [LP23]. Refining [Erd17]. Reflexive [BFH^{+}08, FHL^{+}13a, LS18]. Region [Kim91, TZ19, TZ20, Wor88]. Regions [Eti20, vBEM17]. Regression [AGQS23]. Regular [AKS08, ASS09, AaW09, AP14, Bal08, BS15a, BCS04, BF24, BGG^{+}20, BK17, CP20a, CCG05, CKNV16, CGOvZ21, COLMS22, CF05, CFR10, CFS96, CW09, DJKO19, DHL^{+}13, DFT15, ETT13, FP22, FG01, FKLW98, FW02, FGLS18, GMRT11, GM03, GGS17, Gra07, GRS12, GV21, Hof95, JK24, KW14, KCL98, LLW22, LWY13, MMSW23, MK01, MWvZ11, McL10, Meh12, MTR14, Mom13, MLS11, OC10, OR04, OS13b, PS22, PW18, PR98, RSV^{+}14, Riz02, Zam22, dR14, BH97]. Regularity [BK17, COCF10, DKM^{+}12, NT24]. Reiner [FGS19]. Rejection [BLMS^{+}00]. Related [Adl08, BBBZ12, Bar02, BAG03, BH93, BGS24, Buk16, CKMU14, CY08, CGK^{+}19, CY12, CP22, DJS12, GP08a, GJXZ20, Hua14, KM23, KLMN14, KS23b, MP17b, SR94, SV11, Dow88, WGM95]. Relating [DM18b]. Relation [BBCZ11]. Relations [EIJ^{+}12, HMM09, KSS12]. Relationship [CCO^{+}15]. Relationships [BHH94, GPW13]. Relative [EKK^{+}15, FJ09, HK02, LWW10, SV21]. Relatively [vzGVZ13]. Relatives [BGRR18]. Relaxation [BCCZ10, CQX20, Fuk16, HY13, HI16, Jan10, KG98]. Relaxations [CKP13a, Mic21, SA90]. Relaxed [ALM^{+}18, GKNU10]. Release [GQS^{+}02]. Relevance [HKR00]. Reliability [Ber07, BCE^{+}00, BC92, DM03, Doh03, Bie88, BC88b, BCW96]. Reliable [FKPR05]. ReLU [HBDS23]. Remark [Cha16]. Remarks [GLW11, SV18]. Removal [DT16, FJZ15]. Removed [Lai18]. Removing [Fei04, PSŠ10]. Rendezvous [GJ16]. Rent [LPSR12]. Rental [LPSR12]. Reordering [DE11, ELvS11, LM89]. Repair [MG19]. Repairable [MG19]. Repeated [AS05, BBLM13, CL07, CT21, GP20, Het14]. Repelling [BDvL13]. Repetition [DA10]. Repetitions [DGN^{+}20]. Replacement [LL14a]. Replicated [MHLHL91]. Replicating [Vin93]. Representability [BPWX13]. Representable [BS95a, HvZ14, NvZ15]. Representation [AKZ17, Do19, GKS12, KNR92, Kim11, LLS04, MRV17, PT90]. Representations [ATPRU91, AM21, ADL^{+}09, BFP12, BS93, FK18b, GM16, GPW13, HPW09, Har01, HZ08, KM01, LS08, MR15a, MS05, MvLvL13, Rea15, Saw07, SA90, TT91, TA93, dGV05]. Representative [GMPZ15]. Representatives [Jev95, KR92]. Represented [DEH20]. Representing [AC14, BHNP16, HMSW14, PG06, SV20, Tsa96]. Require [AEPT24, KW90]. Required [DH20a, Gol06]. Requirements [Fra92]. Requires [DZ01, KKSvL20]. Rerouting [MMO20]. Rescaling [LP19]. Reserving [BOS01]. Residues [BCPP09]. Resilience [BSKS11]. Resilient [BGS96, BOS01, KLS10]. resistances [LS89]. Resistor [JLD^{+}18]. Resolution [AG18, MP94]. Resolvability [LTBL20]. Resolvable [JZ05, Siv23, WLD09]. Resolving [ÖV04]. Resonance [KSS11a, ZLS08]. Resonant [YQZ09]. Resource [FYK00, HKST03, JP06a, Lyn94]. Respect [AL24, Fue23, HRS93]. Restarts [JP12]. Restricted [ABS10, BC94, BKYY23, DEW17, Eti20, FP04, FJJ18, GKPP18, HL10, Hui23, KS05, OS92, SS00, Tak22, Vin12, WZ08, HKKK88, Lin97]. Restrictions [AS10]. Result [CDH^{+}04, CR16, GH22, GLM22, HKP^{+}17d, HK96, LZ09, Oxl19, Ric14]. Resulting [FJZ15]. Results [ACG^{+}20, ASS17, AFP^{+}18, BMNP22, BFN20, CS14, CKPV91, dOCHLO19, FKL93, GP17, GSZ22, Har19, HVW07, Mac18, Pik03, SW10, HL92, Ray94]. Retaining [CDMO16]. Retractions [FHJ^{+}10]. Revealing [SB10]. Revenues [Ete22]. Reversals [Cap99, CI01, HvIK^{+}07, RSW05]. Reverse [Cib13, CDD^{+}15, FKMS10, RC18]. Reverse-Free [Cib13, FKMS10]. Reversible [BK90, JMS90]. Reversibly [JMS90]. reversing [AAHLT10]. Revisited [CdMR12, Fel14, LRT08, Naa00, WL10, BS95b]. Revisiting [CD18, FL22]. Revolution [NOP14]. Ricci [BCL^{+}18, OV12]. Rich [ERSZ22]. Richardson [BI13, TKA18]. Ride [BD20b, CCNV11]. Riesz [Lee17]. Rigid [AH96, AG19, GGLS21, Gu18, JDT13, KM22, NOP14]. Rigidity [CW96, CG17, CMM^{+}24, Dew20, FRW12, FLM^{+}18, Fra95, GJ20, Gra91, GMR^{+}21, JNiT21, JiT22, KiT13, Ngu10, NOP12, NSiTW18, SiT15, iT12, Whi88]. Rim [RY91]. Ring [Däu22, HN23, HKS07, KCTR13, MLS11, Myu01, Nao91, SSW98, Sku16]. Rings [BvH03, BCC^{+}05, Car09, Che04, CFG^{+}09, LÖ05, MMR06, BCW96]. Riordan [BL09, Zhu21]. Rips [Shu23]. Risk [BS18]. Risk-Averse [BS18]. River [TH90]. RNA [BDT17, MW22]. Robber [CCNV11, CCPP14, FGP12]. Robbers [CNST24, HKM20, SS11b]. Robinson [CCNP24, Śni14]. Robinsonian [LS17]. Robotic [ABCG17]. Robots [ABY14]. Robust [BDGP22, BMNS22, BD24, DDS16, FKM13, HR02, JK19, KM13a, KPR10, Lai05, Sak89]. Robustly [OS13b]. Roll [ABFRV24]. Rollercoasters [BBC^{+}19b]. Roman [CKPW09, LC12]. Roommates [CL11, HK13]. Root [ABH^{+}11, CGOvZ21, FR94, Més16b, Set21]. Rooted [BC94, CMS09, EGR08, Hay21, KiT13, vIKL^{+}16]. Rooted-Tree [KiT13]. Rootedness [YZ17]. Roots [AA10, BC92, LS95b, Per16, RS15]. Rota [AK14, Cho09, GH06, GW07, Gly10]. Rothschild [dOCHLO21, HLO17]. Rotor [AH11, Cha19, FGLP14]. Rough [HKP^{+}17b]. Round [Bab92, BJHY00]. Rounding [FO00, Knu95, SS02a, TT89]. Rounds [KR98, AA88]. Routing [ACG94, AB00, ACL^{+}06, AP92, BCSK07, BAG03, CEP18, Che04, CPS08, GP99, GS95, GZ98, HSZ13, HKS07, Lei94, LLY10, LRTW11, Lu10, TH90, Zha99]. Routings [PMM98, Ram98, Ram97b]. Rover [MTGK05]. Row [LL14a, OPR12]. Rows [GWZ18]. Rubik [RKDD13]. Rule [Wil16].

Rules [BBFP09]. Rumor [ACMW17, FH10]. Rumors [JM17]. Run [MSK93, Sch04]. Runlength [OR04]. Runlength-Limited [OR04]. Running [HL24, JLL16]. Runs [BBC^{+}19b]. Rural [DMNW13, SL96]. Ryser [BP20].

Saalschutz [Zen90]. saalschützien [Zen90]. Safe [DP92]. Salesman [CR96, Fis14, GS94a, Gao15, GW19, GP16, RC98, SS95, Vyg16, van94, Bal89, Kar89, Sar97, Tas97]. Salvage [BCLR95]. Salvage-Embeddings [BCLR95]. Same [CGH^{+}15, RR18, Wil99]. Sample [CCM^{+}23, KM94]. Sampled [SY11]. Samples [BDG^{+}17, SW23a]. Sampling [BGG^{+}20, MRST16]. Sandpiles [Cai18, CGM^{+}15]. Sárközy [LS24]. SAT [COPP12, Kra07]. Satellite [Bon91]. Satellite-Switched [Bon91]. Satisfaction [BK11, CCJ^{+}17, CM12a, FF06, FK17, KJJ04, KL08, MMS15, MT11, MRT11, ZK11]. Satisfiability [COHH17, CM12a, Fed01, KKM94, Tre04, GW94, KK89]. Satisfiable [FFV11]. satisfied [Mal89]. Satisfying [CSS13, CGMV19, Enc05, FFV11, Gol06, KM94, PPTY22]. Saturated [CCH21]. Saturating [GL10]. Saturation [BJL^{+}24, BK23, CHLT24, DSZ23, FK21, GKP19, Kor18]. Savage [Naa01]. SC-Hamiltonicity [BP10]. Scale [CF10, Fer24, Shu23]. Scale-Free [CF10, Fer24]. Scan [FKK21]. Scarf [CHZ18]. Scattered [jKiO20]. schedule [DL89b]. Scheduling [BLMS^{+}00, CMSV17, DE11, DL89b, GW00, GQS^{+}02, GPSS01, HKST03, JP06a, Jan10, JLL16, KN95, MV18, PP07a, SS02a, SSS13, SS00, Tov90, dS91, dMP93, DL89a, PSW97, SU89]. Schema [BYR05]. Scheme [CZOS98, CZLfW05, Etz96b, JSOS03, OKS06, SY11, Svi03]. Schemes [ACLW18, ABR05, AKTZ19, BDDS03, BGP22, BS91, CCM^{+}23, DF94, DFK^{+}21, HKL^{+}14, JP06a, MWW94b, SW98, SWKP10, CGG88, SV88].

Schensted [Śni14]. Schinzel [Pet11, Pet15]. Schnyder [LLS04]. Schroder [SS91]. Schur [AHP19, BSS14, HRS93, HR05b, KY21, Loe10]. Schwartz [DEMT22]. Score [FvIKS15, PP12]. Scott [HWZZ18, HWZZ20]. SDR [HS89b]. Search [BHRZ14, BDK^{+}21, BKS09, DH05, FMP08, GK02, LS17, Lev09, LNNW95, Luc03, MTGK05, RC18]. Searches [KW08a]. Searching [CK08c, CDHK16, FG00, KK90, Ko88, RT09, RT11, AA88]. Second [BMM20, CKNV16, CW96, CG17, KM19b, KZ18, QD14]. Second-Moment [BMM20]. Second-Order [CW96, CG17]. Secondary [BDT17]. Secrecy [EHV18]. Secret [BBDK00, BI05, BTW08, CCM^{+}15, OKS06, SWKP10]. Secret-Sharing [BI05]. Secretary [AMW00, EHLM17, HN20, KM05a]. Section [MT11]. Secure [FY04, HMS21, OKS06]. Security [AA11, Kao96]. Seemingly [CM12a]. Segmentation [Ete22]. Segments [BEGK22, CK18, Tör93]. Selected [LMRS17, Sch09]. Selecting [DHUZ01]. Selection [BB90, CHY13, DZ01, DKT22, PW13]. Selections [IM96]. Self [BOP94, BÖ05, Hli18, Jan00, LLS13, LW88b, LSX14, MAR22, MAR23, PG06, WS19, BGM94]. Self-Complementary [LSX14, LW88b]. Self-Dual [BOP94, Jan00, MAR22, MAR23, PG06, BGM94]. Self-Orthogonal [BÖ05]. Self-Predicting [WS19]. Self-reduction [Hli18]. Self-Similarity [LLS13]. Selfish [LRTW11]. Semi [BGRR18, ELMS11]. Semi-Baxter [BGRR18]. Semi-streaming [ELMS11]. Semialgebraic [ART14, Do19]. Semicomplete [BJT92, GRY08, BJHM88]. Semidefinite [ASZ02, CKL^{+}21, HMM09, KG98, LP21a, PdCSC21, dCST20]. Semiexact [CKKO10]. Semifields [GRSZ23]. Semigraphoids [MPS^{+}09]. Semigroups [BP12, CGSM16, EA11, Mal15, SS18a].

Semijoin [Gar92]. Semikernels [GSPRM91, GSL98]. Semimodular [Lec90]. Seminormality [LM22a]. Semiorders [SS94]. Semiovals [BD01]. Semirandom [GMP22]. Semiregular [Lat23]. Semisimple [ADL^{+}09]. Sense [Li17, FFP88]. Sensitive [MNP08]. Sensitivity [BPS18, GS20, JDT13]. Sensor [ASMF10]. Sensors [BNRT17]. Sentences [KZ18]. separability [HSLd88]. Separated [KS19]. Separately [HI23]. Separating [KS88, SG16]. Separation [ABC^{+}15, KíP^{+}24, GMW96]. Separations [GHW05]. Separator [BEL09]. Separators [AST94, CD18, DN16, DMN21, LQ22]. Sequence [BGRR18, BH22, CCHZ13, CFK10, DFJS15, FJP22, HLT19, KRR16, LÖ05, SV20]. Sequences [ADM^{+}21, AM96, BBG08, BNCPR20, BBC^{+}19b, BLR16, BLR17, BCPP09, BGS23, BR17a, CCT05, CCD00, CDR16, CGG^{+}16, Col98, DLMO18, FS09a, FKW10, GMZ09, GKMS04, GMTW15, HZ21, HN15, LR07, LXZZ08, Pet11, Pet15, RSW12, ST17b, Ste00, TQ09, XQ06, Zhu18, CK11, KS88, Mit97]. Sequential [KRZ21, SL95, Tak08, Pip89]. Serial [GF08, Kot13, MMR06, WB90]. Seriation [GJ19]. Series [ACM11, ABH^{+}11, BGL24, BC09b, BFN20, Cho92a, Cho94a, Ege10, EP10, KR16, RY91, RS15, Wag96, ZN08, ACS97, BM97]. Series-Parallel [BC09b, Cho92a, Cho94a, EP10, KR16, RS15, ZN08, BM97]. Serre [PPTY22]. Server [CL91b, CKPV91]. Servers [Ber07]. Set [BBC^{+}19a, BBF99, BNMN92, BFRS16, BS23, CLM03, CdV02, CHY13, CPPT20, CDW07, CY21, CH10, CPPW13, DO08, DKT22, EKM^{+}19, Ete22, FGS19, FO08, FFV11, FPP22, GDCM20, GDVL17, GKI^{+}23, GLSS16, GS03, GT98, GS94b, Han19b, HQ03, IKZ08, JPY10, KS92, KN16, KMÖ18, KS05, KMW06, KvIL^{+}12, KPS20, Kim11, Lev09, LPS18, PPR22, Pin08, PY09, RRW22, Rya07, Sha20, WZ18, BKK16, BS88, BC88b, EH91, GS89, Hur94]. Set-Codes [GDCM20]. set-covering [BS88]. Set-Dependent [Ete22]. Set-Systems [FPP22]. Set-Valued [FGS19]. Sets [Adr23, Age94, AHH^{+}10, AHP19, ATPRU91, ABZ15, AFKK23, AS16, AMNV18, Ave13, AE03, BNN90, BDJ^{+}15, BB23b, Bon10, BLM10, BKKM99, BM14, BK05, CW98, CCSW23, CER98, CCG^{+}11, CLW09, CLiO18, CCG05, Cib13, CD16, CF16, DSL19, DA10, DM17, DJMV21, Elb09, EFK14, Eti20, ENSZ00, FKS12, FM24a, Fis90b, FL96, Für91, GZ19, GVW06, GLP^{+}12, GS93a, GMPZ15, GL10, HST19, Har11, Hav19, HKP22, HK93, HT93, JK99, Jev95, Juk16, KLMN14, Kim17, KSY18, KLMR18, KM01, KT17, KMN23, LRT08, LM12, LZ06, LM22b, LSW18, MB18, MZ22, Mom13, MR15b, MD16, MD11, MS05, NY21, Nov18, NS11, OR04, PP22, PP07b, RSV^{+}14, Ros09, SS19, SS04, SS95, TT16, Tak08, Vse05, Yu17, Zha90, Abe91, Bal88, BM94b, BJvV92, Bou97, HR88]. sets [HS89b, Sag88]. Settings [ACF18, LS24]. Seven [CFG^{+}09, KK14a, Sav14]. Several [CHX15, DF04, Fed06]. Seymour [CG02]. Shadows [FZ22, KMV15]. Shallow [ES11, HW24]. Shannon [AL07, HTS18, Kas03]. Shape [CP10b, FL00, MS19]. Shapes [DMN21, MPP17]. Shard [Pet13]. Shared [BBDK00, Pit23]. Shares [HSS19]. Sharing [BI05, BTW08, CCM^{+}15, OKS06, SWKP10]. Sharp [DMP07]. Shatter [BG19]. Shattering [FS12b]. Sheddings [MST21]. Shellability [BCE^{+}00]. Shellable [Cha91]. Shellings [CDGO22, MST21]. Shepp [CZLfW05]. Shermer [DJT15]. Shift [Eli09, MK09]. shifts [Lin89]. Shop [GPSS01]. Shops [JSOS03, Svi03]. Short [BF96, BBC^{+}19b, CL16, CR17b, DMR11, EW19, FKP15, HGY20, KKL^{+}10, KM21b, KLN10, KZ18, Luk20, MS21a, NT05, PV10b, PP90, RSM^{+}96, STT92, Hut88]. Shortcuts [HP21]. Shorter [BGP22]. Shortest [AP14, BNRZ23, BGW20, Duc21, IKK^{+}22, KS03d, LL17, MNPR17, Rom06, Zha93]. Shortest-Weight [AP14]. Shrinking [CL90]. Shuffle [Obr93]. Shuffle-Exchange [Obr93]. Shuffles [JJL23]. SIAM [GM93]. Sided [CHK17, Kam19]. Sidon [Juk16, KLMR18, KT17, RRW22]. Sidorenko [Cor24]. Sieve [Sch02a, Gor93]. Sieving [BER11, Cha16, Stu21]. Sign [CP10b, HS22, DK89]. Signatures [NY23]. Signed [Che17, CLLZ18, GPW13, HLL^{+}21, KR16, KNNW23, LLS^{+}20, MS17a, MRS19, NY23, PP13, Saw07, SZ94, WYZZ14, WY20]. Signed-Circuit [WY20]. Signed-Graphic [PP13]. Signless [HRS17]. Silverman [HR88]. Similar [GJ12, LRR14]. Similarity [LS17, LLS13]. Similarity-First [LS17]. Simple [BBJ^{+}21, BL09, BCHP08, CR23, CS91, De 24, Ful14, KvRRS23, MC12, Mer15, MD16, Saw07]. Simple-Triangle [Mer15]. Simpler [Hli18]. Simplex [ABGJ15, FL00, GK07, JPY10]. Simplex-Algorithm [GK07]. Simplices [Ave12, BM07, FXZ23, KT17]. Simplicial [BHM16, BCKP19, Dan09, DDL20, DMS14, EGS13, Jon05, Lut04, PPTY22, TT93]. Simplicity [BJJ98]. Simplified [Tov90]. Simplotopes [SS18b]. Simply [Sim13]. Simulation [HKK^{+}09]. Simultaneity [dMP93]. Simultaneous [KS03b, SL95, SZ15]. Sine [Sca05]. Single [Ave12, CMSV17, FKPR05, FKL^{+}19, GQS^{+}02, GS95, GS16b, HV00, KLW89, KSW17, KPR10, NPSW23, Ull14, YP22, Zha94, CGG88]. Single-Bond [Ull14]. Single-Change [Zha94]. Single-Elimination [KSW17]. Single-Exponential [FKL^{+}19]. Single-Layer [GS95]. Single-Machine [CMSV17]. Single-Message [KPR10]. Single-suit [KLW89]. Singleton [KMW06]. Singular [BGJW21]. Singularity [FO90, Ngu13]. Sink [EFK14]. Sink-Stable [EFK14]. Site [GM04]. Situations [BOT92]. Six [BS16b, KM19a, PP22, WZZ18]. Six-Term [PP22]. Six-Vertex [KM19a]. Sixth [CGOvZ21]. Sixth-Root-of-Unity [CGOvZ21]. Size [AS09b, Ave12, BLP21, BDM02, BHL^{+}15, BW02, BGS24, CDM00, Cib13, CCL^{+}06, Das99, GW99, GS93a, Hås94, HST22, HW15, HKK^{+}09, HOD22, JO18, KLWY21, KL92, KMPR14, LW18, MS19, NW95, Pic14, Pik03, Ran02, RR18, SC17, Sud08, WS12, GI97, HS89b, PS97]. Size-Linear [BGS24]. Size-Ramsey [LW18]. Sized [BBRT23, Sei01]. Sizes [AHS01, AB07, CGHW22, CGL10, CCG^{+}00, De 24, Gar92, KPT12, KŠ08a, MP22, OS16, TW21, Zak14, Lin97]. Sizing [DW10, DW11, vWW94]. Skeletons [BDEK06]. Skew [CKK^{+}04, HK14, Mom13, MPP17, ZM02]. Ski [LPSR12]. Skrekovski [CCO^{+}13]. Slack [GMTW19]. Slash [HHL24]. Slash-Free [HHL24]. Sliceable [YS95]. Slicing [SFS09]. Sliding [AMPT93, AM95, GOR20]. Sliding-Block [AMPT93]. Slightly [LM22b]. Slopes [KPP13]. Slotting [Che04]. Slow [Gao18, ILM^{+}16]. Slowly [JM17, PP12]. Small [Adr23, ABR05, AM06a, AB07, BCS04, BPV10, BB23b, BBRT23, BGH^{+}17, BCG22, CCJS22, CY21, DPR22, DFK^{+}21, FKLL15, FKV21, GDCM20, GHvtHP15, GRS11, JM17, JN17, KV15, KM94, KLL13, KS12b, Lev22, Lu08, Lut04, MW08, MP21, Mot19, NŠ07, OW22b, PP22, PG06, Pin14, RSM^{+}96, RRW22, Shp15, SB91, Ste10, Sul05, TH90, Wan08, NNI97]. Small-Diameter [DPR22]. Small-Sized [BBRT23]. Small-World [JM17]. Smaller [Kra18a, MP22]. Smallest [Adr23, BBS00, CSS01, Gab04, Gab05b, OS22]. Smartstart [BD20b]. Smith [WS17b]. Smoothed [KRS15]. Smoothing [DH91]. Snoop [PW02]. Social [BKL^{+}15, Che09]. Soft [EIH18]. Solid [CFLZ19]. Solomon [BK91, RR03]. Solution [BT14, DGM12, MNPR17, Sha23, Sza08]. Solutions [BBCZ11, CHX15, EIJ^{+}12, EV98, FGP10, HT13, Hor19, ILM^{+}16, LW03, Sim90, Sis21, Str20, FG89]. Solvability [CDV10, SS05]. Solvable [AMNV18, HHH^{+}02, van94, ALZ96, BS88]. Solves [CM05b]. Solving [Boy96, KKSvL20, MPSV21, Jae89]. Some [AL24, BB09, Bal08, BBBZ12, Bar02, BGW20, Cai18, CCZ12, Cho92b, DR24, Din13, DK10, DSST13, Har19, HL92, KPS19, KNNW23, iKSZ04, KBE^{+}05, Li17, Lub90a, MW20b, MW03, NT24, PMM98, Rém02, RY91, SWR12, Ste07, Stu88, SW10, Zha94, Bal89, Bie88, DP17]. Sometimes [DRW98]. Somewhat [KOS16]. SONET [CFG^{+}09]. Sorting [AKKS89, AA88, AU91, BP98, BNN90, BFR12, Cap99, CDP94, CI01, dAHFdFK10]. Sós [CDT23, HKP^{+}17a, HKP^{+}17b, HKP^{+}17c, HKP^{+}17d, Roz19, Sha23]. Sourcewise [CE06]. Space [ABCG17, AP92, BDM02, BK12, DHS14, FKL^{+}19, GMTW19, KW17, MNPR17, NPSW23, Owe11, Pol19, Sei01, Web08, Woe93, Car88, IS93]. Spaced [Lag00, LSO03]. Spacefilling [GS94a]. Spaces [BS09, BL17a, BFKT24, BDG^{+}17, BGL03, BP17, BN05, CLWZ22, CCNP24, CLVZ96, DGN^{+}20, FGLS18, GK13b, GL10, GL15, Har11, IK09, KS23a, KM94, LSTY17, MP21, MW22, ÖV04, Rém02, SBD^{+}19, Sol12, Spe08, Vin11, Vse05, XG20]. Spacial [Hor19]. Span [CtJL01, Sza06]. Spanners [BGJ^{+}12, BT22, BFKT24, BCE05, CC95, ENS15, HPS96, HP21, KV15, KLM^{+}03, LS95a, SE14, DYL06]. Spanning [AAFL06, Ald90, AFG^{+}09, AKM22, BFa24, BJHY03, Bon08, BZ11, BDEK06, BL17b, CWY00, CGM^{+}15, Cha19, CEOT15, CEOT17, CS19, CSS01, CFG^{+}15a, CP10a, DL19, DM11, Gab04, Gab05b, Gao13, GL95, GPS19, Gu18, HV17, HT18b, KKKS17, KW91, Kri10, KKS17, LLL17, Lu10, LW20, LMW23, Nog24, Oze13, RSM^{+}96, SW23b, BFM94, BBM90, Cho94b]. Sparse [AM11, ADL13, BT22, BCE05, BK17, BMPS21, CR23, COCF10, CM16, CE06, CY20, Cre04, DEK22, DMS12, DŠT08, EKM^{+}19, FO08, FS01, FHL13b, Gao13, GK13a, GM13, Gu18, Hen18, HKP^{+}17a, HP21, JT11, JLR^{+}17, Joo15, JW18, KKKS17, KMRR09, KLMR18, KS12a, Kri18, KM11, KZ18, LSSY10, MP17a, MT20, Pra13, RN21, ST20b, Sha23, SW21, SY11, BG91, GPS88]. Sparsification [BZ20, FK17, GHP20]. Sparsifiers [KR20]. Sparsity [MC93]. Spatial [BP17, DM03, JM17]. Special [BL23, HMZ21, LR07, MT11, RY91, Sch02a]. Species [SV11]. Specified [BGS23, CDMO16]. Spectra [BBMM09, CM03, CF09, TQLT13]. Spectral [CDT23, FGKP22, GCH22, KLM14, Lat23, LP23, Liu24, Sot15]. Spectrum [BCL^{+}23, BPS07, CGL10, CL21, KM24, Lee17, Sca05, GM94]. Speed [MV18]. Sperner [AFP^{+}18, EFK05]. Sperner- [AFP^{+}18]. Sphere [AMB11, BC94, BCKP19, EIH18, GGLS21, Vin11, Zhe16, BS95b]. Spheres [FGKP22, LN17, Lut04, Mus21, NSiTW18]. Spherical [BV10, Buk16, FRW12, Kra18b, MWW20, Pol19, Yu17]. Spider [ABS13, Pip06]. Spider-Web [Pip06]. Spikes [BCC^{+}19]. Spin [Mar20]. Spine [EM99]. Spirality [DGL10]. spirals [Wen97]. Split [AGH14, BS24b, BL04, CdMR12, HLMP11, HL00, LC04, Tim08, AFK12, KM88, Spi89]. Split-Perfect [BL04]. Splits [BBS17, LGS11]. Splitter [GZ06]. Splitting [GMW05, Goe01, GL15, JPZ21, Jor03, Myu01, BJvV92]. Spotting [BKKZ17]. Spread [CFH16, JM17]. Spreading [ACMW17, FH10]. Spreads [ÇMP15]. Springer [CFJ11]. Sprinkling [NT21].

Square [BHL05, BDDP24, KM19a, LS95b, MTV08, ST17b, FH94]. Square-3PC [MTV08]. Square-Free [KM19a]. Squaregraphs [BCE10]. Squares [DŠT08, GS94a, JSO12, KT16, WLD09]. Stabbed [AKP20]. Stabbing [KS06, Tót08]. Stability [AG06, CW96, CG17, dOCHLO19, DT16, DW22, EFMW18, GLM22, PdCSC21, DK16]. Stabilization [CM05a]. Stable [CDHZ12, CdV02, CL11, Che16, CFKP22, EMM14, EFK14, GS03, GSZ22, HSZ13, HW10, HW17a, JK99, Kam19, Kam22, NH91, PSV08, dBSS24, Pit89, BM94b]. Stack [HP97]. Stacks [HLR92]. Stage [Kar92, CH06a, DH90]. Stalls [COHH17]. Standard [DK22]. Stanley [Bón09, CR16, CS18a]. Star [HSV23, iKM10, LLZ99]. Star-Forest [HSV23]. Stars [DR24, GRS11, Hua14, TT07]. Start [vWW94]. Start-Up [vWW94]. Starting [Dan09]. State [BN01, BS15b, CL15b, LGS11]. Stationary [MP13]. Statistical [AA11, BHM00, DM03]. Statistics [Bón09, CWYZ10, FHL20, IKZ08, MSK93, Sch04]. Steinberg [HY13]. Steiner [AFG^{+}16, BM97, BAH10, BT22, BHM^{+}22, BH15, BCHW17, CKP13a, CGN06a, CK14, CKN13, DM18a, DOS94, DR19, DFK^{+}21, ES11, EH91, FKL^{+}19, Für91, GMZ09, GMW96, GNS11, Hor14, IW91, JZ05, JLR^{+}17, NO09, PC97, Rif99, RZ05, RTW97, VZ93, Wen97]. Steiner-Tree [VZ93]. Step [CK14, LPS09]. Stiffness [JiT22]. Stirling [Bón09, FP99, KLMR03, RM97]. Stochastic [AH21, AFG21, BS18, CS94, HKL^{+}14, HSS24]. Stock [DW10, DW11]. Stockhausen [Yen97]. Storage [BYKKR18, DM18a, LNO96]. Story [LMS19]. Strahler [Yam20]. Straight [KFHR94, Sch91b]. Straight-Line [KFHR94, Sch91b]. Strategic [Nas14]. Strategies [Fei10]. Strategy [BKR10].

Streaming [HKMY22, ELMS11]. Street [CDF08, RX88]. Strength [Gal18, HKR00, KKP11, MP14, Nie00, NS11, o09]. Strengthening [BH22, CKPS13, DFK^{+}23, GHM21]. Stretch [CCH14, CDMT21, KV15]. Strict [GL14, HW17a, FK97]. Strictly [CHZ04]. String [ADM^{+}15, RT98]. Strings [BFN20, CI01, Gol96, HvIK^{+}07, MR04a, MR04b, PRS98, RSW12]. Stripping [Gao18]. Strong [AH16, AM06b, BB23b, BP10, BGRR18, BP17, BS10b, Cal13, DŚN21, DE93, DK14, FKS05b, GL95, HHCHL21, HHL24, HY10, JPY10, PSS09, Sch02b, XZ22]. Strong-Baxter [BGRR18]. Stronger [LRTW11]. Strongly [BJHY03, CDV10, CGMV19, DN16, FLM^{+}95, IMS05, KMS23, LLL17, MTR14, Mom13, RLWY21, TW19, ZSW11, RX88]. Structural [BJS22, GKI^{+}23, SW01]. Structure [BL19a, CCSW23, CER98, CP16b, CMvZW16, COS10, DMS21, DE93, DEW17, DL17, DL18c, DL18b, EW19, GM05, GT13, Hay21, HKP^{+}17b, HKP^{+}17c, HR05a, IM96, KNK93, KM01, LM21, MM11, Nas14, NN22, Rém02, SB10, Wan08, BBM90, BB97, Ram97a]. Structured [AGK^{+}23, FGT18, DL89b]. Structures [BDT17, BG11, Che93, CF16, Dan01, FLS10, GLY07, HIKT99, HMS93, LT11, MW22, SW23b, STT92, Špa07, SWKP10, Tam88]. Study [BDF^{+}18, DFK^{+}11, FYK00, MK09, ACS97]. Style [GH22]. Subadditive [DKLW24]. Subclass [CK08a]. Subclasses [CS18c]. Subcodes [BGÖ17, LWW10]. Subcolorings [FJLS03]. Subcritical [CFDK20, DFK^{+}11]. Subcubes [Off08]. Subcubic [DŠT08, JRS14, KMM12, Liu14]. Subdeterminants [GWZ18, JB22]. Subdigraphs [BJHY03]. Subdivided [BGM08, JS12]. Subdivision [FW20, LiO15, LZ18b]. Subdivisions [Ath14, BFH21, Jan20, JQ20, jKR21, Mur10, Tót08, Wan23, WW18]. Subexponential [BFPP15]. Subfamilies [BFK^{+}12, CKOS06]. Subfamily [EJK^{+}09]. Subfield [NvZ15]. Subgraph [AA05, AM11, AK22a, BL17b, CEOT15, CEOT17, CSS01, Cho92a, Cho92b, CM07, DJKO19, Dji06, Gab04, Gab05b, Gu18, GRS12, HLS22, HLS24, SS11a, Zha93, BM97, Cho94b]. Subgraphs [AFK12, ADO23, AKS08, AGH11, AFS12, AS07, AB07, BB16, BMM20, CCOY17, CvBdJdVK22, CDM00, CDM04, CS19, CKP^{+}21, CGMV19, DS16, FXYY14, Fei04, FKK20, FG14, FKP15, FPS18, FM13, Gao13, GS98, GP18, GRS11, HH04, HV17, Jan05, JN17, Joo15, JiT22, JW18, KiKK17, KKKS17, KW08a, KLN10, KLL13, KLN91, KS12b, LRWZ12, LM16, McC21, Ngu24, PTT16, PT94, SW01, Tuz08, ZL11, BM94a, BFM94, BP89, FH89]. Subgroup [Hyu10, MWWZ20]. Subgroups [Shp13]. Subhypergraph [CDK^{+}18]. Subhypergraphs [Nag10]. Sublinear [CHZ18, DN16, DMN21, ERS19, GRS11, JMU23, LMM21, RS11]. Sublinear-Time [GRS11]. Submatrices [FK21, KPT20, LM16, PSWX24]. Submodular [BW09, DW24, FiT14, GS13, HI16, HKMY22, IMS05, KSF19, LMNS10, Osh21, SSS13, Yos19]. Submodularity [CGV^{+}14, NYKY20]. Subpattern [AR08]. Subpermutations [GJK^{+}23]. Subposets [Sch09]. Subquartic [New20]. Subsequence [BH20, CP96]. Subsequences [BM14]. Subset [CKN^{+}15, CPPW13, DFX21, ENSZ00, JKSY22, PW13, SS00, DM88]. Subset-Restricted [SS00]. Subsets [AAH14, BGN15, BD20a, BGH^{+}17, CFG^{+}15b, CHHM09, DO08, Für91, GJ12, HW96, HIKT99, LPS18, PRS03, PP10, Sza06, Zha90, Hur94]. Subspace [WZ08]. Substar [CNRS15]. Substring [ADM^{+}15]. Subtour [GW19]. Subtournaments [Kim17]. Subtree [Ald22, BHL^{+}15, BLO^{+}22, FMP08, MS19]. Subtrees [KW08b, NS22, Pit23]. Succeed [Ste00]. Success [AFKK23]. Succinct [GKS12, HKL99]. Suffice [Cra19, HMS05, SV20]. Sufficient [BR17a, GMR^{+}21, Sch10, Ste00]. Sufficiently [Ste00]. suit [KLW89]. Sum [BL19b, Bev10, CMSV17, DKS16, HZ21, Hav19, HR05b, Jev95, KT16, KR13, MLMR23, MRNS15, MRNS17, Ram04, RR18, Zak14, LRN11]. Sum-Distinct [Jev95]. Sum-Free [Hav19]. Sum-of-Diameters [Ram04]. Sum-Product [BL19b, MRNS15, LRN11]. Sums [CY08, CHZ09, CKP16, CH10, DFX21, DK22, Ege10, FKT99, GS94a, HKW15, IK09, KR93, LS06, Mac18, MPSS20, MWW21, OPR12, OS22, RN21, SW21, Smi01, Tsu23, CP96]. Sumset [CCSW23]. Sumsets [Fan20, Hui23]. Sun [Hoà10]. Super [Das99, FK97, HT90, Kam22, KW90, KMÖ18]. Super-Logarithmic [KW90]. Super-Set [KMÖ18]. Super-Stable [Kam22]. Superconcentrators [RTS00]. Supercritical [Fer24, KW13]. Superlinear [Lon21]. Superlogarithmic [McK19]. Supermodular [BGS17, KL08]. Supersolvable [Sch09]. Superstring [WS06]. Supertree [DS05a]. supply [NS89]. supply-demand [NS89]. Supported [NOP12]. Surface [LS06, Moh99, SS09]. Surfaces [BK12, BMNS22, CMS09, CGH^{+}10, DKS18, DŠ09, DL14, DL18c, EO16, Fio06, GKLT19, KM14, KML05, NOP12, NOP14, Oze13, PV22, Wu09, Žit94]. Surjective [AMPT93, BMR^{+}10, FGZ19, GM05, KST06]. survey [Sav97]. Survivable [BKM08, BH93]. Surviving [CCVZ10, LW10]. Svensson [New20]. Swapping [FRW12, Zha11]. Swaps [JJL23]. sweep [Fra89]. sweep-line [Fra89].

Swendsen [Ull14]. Switch [AK22b, DM19]. Switched [BCSK07, Bon91]. Switches [FG21]. Switching [Bon91, CHZ04, HH04, OYY13, Orr08, PL94, Wil99]. Symbolic [RT18]. Symmetric [APdMOZ23, AL17, Bah23, BS95a, BvH03, BGS17, CM90, CEP18, CR96, CS21, DJKKV23, DvW18, GF08, GS13, Kot13, Lub90b, MR04a, MR04b, NYKY20, Nov18, RC98, SS18a, SiT15, Sol12, Spi19, STV21, WZ18, dGV05, van94, CH19, Ste88]. Symmetries [LB09]. Symmetry [BGS23, KPT12, LLL18, MAR23, Sch10, TKA18, CS89, GPS88]. symmetry-breaking [GPS88]. Symplectic [HK11]. Synchronizing [GJ16, Jun12]. Synchronous [GPP01]. Synthesis [KYDN09]. System [BZ23, FLMY09, MM11, BC88b, DH90]. Systematic [KCTR13, Mou94]. Systems [ABS10, Bil03, BC95, BH15, BCHW17, CS14, CM05a, CHX15, CM05b, CFH16, DM18a, DR19, FPP22, Für91, Gad18, GMA15, GPP01, GMZ09, Gyá19, Har10, Hor14, HLWY23, IM96, JZ05, JPY10, KM13a, KS03a, Kas03, KS05, KMW06, KMT07, KN95, LOW10, MH09, Mur06, Nag17, NT24, PW02, PSML08, Rif99, SMNF09, ST07, Shi12, Sza08, ZLS08, DL89a, EH91, FG89, HMP97, HS89b, KS88]. Syzygies [BPEY22]. Szemerédi [BC11, NP20b].

T [ACP22, FKS05a, Sch02b]. T-Coloring [FKS05a]. T-Joins [ACP22]. T-Perfection [Sch02b]. Tableau [PV10a]. Tableaux [FGS19, Śni14]. Tables [Lu10, Sul05, KG93]. Tabloids [RY91]. tabulated [KG93]. Tail [MSD19, Raz20]. Tails [BG20]. Takagi [Lev15]. Tandem [LZZ22]. Tanglegrams [CSW17]. Tangles [DEE17, Erd17, Flu24, GS16a]. Taquin [Śni14]. Tardos [Peg14]. Target [CHY13, DKT22]. Tarsi [Gly10, Sto12].

Task [CM05b, KN95, PSW97, DL89a]. Tasks [AE04, DL89b]. Taxa [DS05a]. TCP [AB05]. Teaching [ABZ15]. Technique [BYR05, CGK^{+}19, SS18b, Zuc92]. Techniques [BB03, HKP^{+}17d, MV18, Ste00]. Templates [GvZ17]. Temporal [Fer16]. Temporary [AE04]. Tenner [FGS19]. Tensegrities [KM21a]. Tensegrity [CW96]. Tensions [Che17]. Tensor [BD24, BC23b, PPU92]. Tent [AL17]. Terao [BBJ^{+}21]. Term [PP22, SST08, Wal19]. Terminal [BKM15, Bih05, Fuk16, HI23, KNZ14]. terminals [RTW97]. Terms [RS15]. Ternary [HvIK^{+}07, OS22]. Tessellations [GPW09, Vin93]. Test [BFRS16, KG93]. Testable [AcRS07]. Testing [AK02, ADPR03, AKKR08, BD20a, CCGG18, DJKO19, DGL10, EG03, GKS12, GKW19, HH01, LRR14, MMPS10, Neu24, PRS02, SL96, HY89, Spi89, YH88]. Tests [MPS^{+}09]. Tetrahedra [SvM08]. Their [BL19a, BC23b, HY15, JMW17, iKX20, KK14b, MK01, MMJF03, LW88a, LWW10]. Theorem [AFG^{+}16, AOW15, AHJ^{+}24, BL09, BS90, Buk12, BR19, CLMS22, CHZ18, CDT23, CuKŠ07, CP20b, DFK^{+}23, DK16, DSS92, DL12, DMN12, DS06, EW19, Fed06, FKS05a, FiT14, GZ06, GT13, HM11, Hay21, HLY23, Hui23, HLT19, IN22, Koc98, KS03c, KOT09, LGS11, LP23, LS24, Lo14, MMS17, MWvZ11, MS14, MW20a, NP20a, NP20b, OSW16, PVY22, Ram90, Yam20, Fuj97, Hef97, RW89, WW91, AKW05, AKS07, BYHR10, BV10, BC11, KLP12]. Theorems [AL24, BEL09, CCG^{+}00, DFJS15, DEE17, DS21b, FHH08, FR06, FT05b, IK09, MSD19, SS08, Suk13]. Theoretic [BCS04, KM05a]. Theoretical [Wag07]. Theory [ACGH20, ASMF10, BLS17, BJ91, DHT06, FL92, GRR15, Gad18, GR17, KSV05, KPT94, KKL19, MP95, MT20, PT90, ST20b, ST10]. Theta [CS18a, CK08b, Cre04, KG98, LY23, LS05]. Thickness [DM18b]. Thicknesses [Fis90b]. thin [BW92]. Third [HK16a]. Thomassen [CCO^{+}15, LPS09, WL10]. Three [AC14, ABNR22, AS05, BCD^{+}12, BL23, Bon08, Cai18, CH17, CMSM^{+}18, Cho09, CGSZ20, DD13, DNR23, EM99, FH21, GPP04, GLW11, HJ18, HZ10, KKL^{+}10, Kar92, KZ04, LGS11, LSW18, Mot19, MW20b, NH91, Pet15, vIKL^{+}16, IP91, IS93, KP06]. Three-Color [Mot19]. three-colored [IS93]. Three-Coloring [CGSZ20]. Three-Dimensional [GPP04, MW20b, NH91]. three-linking [IP91]. Three-Stage [Kar92]. Three-State [LGS11]. Threshold [Ale10, ABZ15, AKM22, BTW08, BDDS03, BB03, BH22, BGS23, BS91, Bru90, CM12a, DF94, DMP07, DP16, FRMPV15, Hås94, HLMP11, KPPR15, SB91, SR94, Zun11, FV97, SV88]. Thresholds [BCKN21, CH06b, FRZ16]. Throughput [ILM20]. Ties [CHK17, Kam17]. Tight [AFH^{+}18, APS22, BF12, BD20b, BFN20, BCF^{+}20, Chi23, CL21, CM14, DSZ05, EP10, FJK^{+}19, GP91, HM19, KL19a, KL92, Lam20, LW18, LNO96, LLZ22, Mic21, NY21, Nie00, OC19, RN23]. Tighter [RZ05]. Tile [JWF05]. Tiling [BDM02, CD11, HW17b, Zha09]. Tilings [BE13, CMR18, CLVZ96, DLMT21, Din06, EV98, Han18, HS10, HLWY23, Lai18, ÖV04, OC19, Rém02, TV03]. Time [AK22a, BNMN92, BP15, BST14, Bon91, BCHP08, BS16c, BFN20, CdMR12, Che94, CKN^{+}15, CMM^{+}10, CF05, CFP19, CGP98, Dji06, DP16, DHJ^{+}13, ERS19, ELSS17, FP04, FO00, FKL^{+}19, FPT22, Gab05a, GV92, GRS11, GJ16, GPS19, HKL^{+}14, HL24, HKL99, HTV05, HvtHLN12, HV00, HK05, JMS90, JSOS03, JP06a, JLL16, JLR20, KN16, KK90, KMS98, KP95, LW17a, LM16, MN15, MPSV21, MNN18, Moh99, NT24, NPSW23, PS22, RS22, RS11, ST07, SY11, SWKP10, Svi03, Wan02a, WY10, Zuc92, AKKS89, Car88, IS93, MM96, PRS88]. Time-Division [Bon91]. time-space [Car88]. Times [BPS18, CP10a, HH92, JP12, SSS13]. Together [Més16b]. Tolerance [GM16, MSZ10]. Tolerant [AS02, AU91, BCSK07, PMM98, PL94, SX13, UBHS93]. Tolerating [GV92]. Tomaszewski [DvHT20]. Tomescu [EEFH21]. Tomography [AG06, AG18, DFJS15, DGM12, GLW11, HT13, SB10, vD11]. Too [PP12]. Top [FKS03]. Topological [AM21, CdVL11, CDM00, CDM04, DGL11, EW19, Ful14, HK15, KLN91, KO06a, SZ13, TT07]. Topologically [ACGH20]. Topologies [VZ93]. Topology [BDM^{+}24, BAG03]. Tori [JCB06, PW18]. Toric [BDvL13, BC09b, GMA15, dAHFdFK10, LS06, SS09, SS10, UV15]. Toroidal [iKO16, NOO12]. torpid [GKRS15]. Torsion [Lev22]. Torsor [SW23b]. Torus [AF10, BL09, BL90, KS08c, Rio98]. Total [BG88, BHT16, CLS09, CDWZ17, CCZ12, DLMT21, DZ09a, ELSS17, Gra07, HY10, HY15, KSŠ08, LPW^{+}13, dCST20, Zhu21, o09, KW96]. Total-Coloring [KSŠ08]. Totally [EM20, Gij05]. Tough [DNS94]. Toughness [Gu21]. Tour [DRW98, SL96, Zha93, Kar89]. Tournament [CPR99, DLMT21, KSW17, NY21, PRS88]. Tournaments [Alo06, APS^{+}24, BNN90, HW96, KKKS17, KKO16, Kim17, MT90, Tan21]. Tours [BIT13, CLS15, Tas97]. Tower [Rom06]. Trace [GR17, Mat19, RMS01]. Traceability [SW98]. traceable [Zho88]. Tractability [HN15, HK14, KPPW15, Sli10, TZ15]. Tractable [AKP^{+}22, ADKS18, BCDF19, CPPW13, KPT22, LSSZ19]. Trade [AP92, CGP98]. Trade-Off [AP92]. Trade-Offs [CGP98]. Tradeoff [KR98, Lou10]. tradeoffs [Car88]. Trades [KAN90]. Traffic [BCC^{+}05, MLS11]. Trait [FS12a]. Trait-Dependent [FS12a]. Transfer [CL15b, SHS23]. Transference [MSD19]. Transform [DV04, Kim91, Sca05, Con89, Fil89]. Transformation [KCL98]. Transformations [BV21, MK01]. Transforming [CGH^{+}10]. Transforms [Sca03, Car88]. Transhipment [AC90]. Transient [Cha19]. Transients [JA16]. Transition [FKK21, KKP15, LN21, MMPPG23, RSV^{+}14, VVY15]. Transitions [CMR18, CH13, CH17, CFKK17, DH05]. Transitive [BGJ^{+}12, DLMT21, GL14, HLST00, Iri16, Mac91, Zho09, BJHY03]. Transitive-Closure [BGJ^{+}12]. Transitivity [HT90]. Transmission [AR17, FKPR05, KPR10]. Transport [DSS12]. Transportation [BFH15, JK24, KP00]. Transposition [CKdAHdF13, LY10, RS93]. Transpositions [ALSY11, BP98, BFR12, CI01, dAHFdFK10, PR91, RSW05]. Transversal [BHT16, FiT14, JPvL21, PPR22, SSS24]. Transversal/Packing [GL17]. Transversals [CHKW24, CMSM^{+}18, FKS12, Fel90, HY10, HY15, LMM21, MZ22, RS14, SMNF09]. Travel [MTGK05]. Traveling [CR96, Fis14, GS94a, Gao15, GW19, GP16, RC98, SS95, Vyg16, van94, Bal89, Kar89, Sar97, Tas97]. Travelling [DRW98]. Tree [AFG^{+}16, Adl08, BP13, Ban90, BDF^{+}18, BFGR17, BHM^{+}22, BS15b, BDLR01, CC95, CKP13b, CQX20, CP10a, DDJ^{+}21, Dah93, DHS14, DM11, DJM^{+}18a, DJM^{+}18b, Erd17, FJZ15, FS05, FKL^{+}19, FJK^{+}19, GMRT11, GKS22, GLSS16, GL95, GLY07, HW16, Hay21, HMS05, HKS07, JL23, JK24, JLR^{+}17, KiT13, KL19a, KLM^{+}03, Let19, LSTY17, LN21, LNNW95, LR04, Luc03, ŁR91, MRV17, MS19, MdCW16, Owe11, Pit23, RZ05, San96, Sca05, SSS22, Ste00, VZ93, Wan02a, Yam20, Alt89, DYL06, DL89b, HS97, IW91, Tod89].

Tree- [LR04]. Tree-Based [Hay21].

Tree-Cut [GKS22]. Tree-Decomposition [Erd17]. Tree-Degenerate [JL23]. Tree-length [BFGR17]. Tree-Like [BHM^{+}22]. Tree-Matchings [ŁR91]. Tree-Related [Adl08]. tree-structured [DL89b]. Tree-Width [HW16, Luc03, San96]. Treedepth [CNP21, GJW16, NPSW23]. Treelength [CDN16]. Trees [AT90, AAFL06, Ald90, Ald22, ART14, AFG^{+}09, ASY14, ANS16, AKM22, ABR05, AL95, ÀB23, AH11, ANP14, BFa24, BS95a, BDK^{+}21, BL17a, BPSS19, BPSS20, Bev09, BM11, BVVW11, BCLR95, BKKZ17, BDT17, BHM^{+}22, Bon08, BZ11, BLO^{+}22, BMV92, Cai93, CWY00, CHP^{+}90, CKP13a, Cha91, CGM^{+}15, CDWZ17, CKK^{+}04, CK14, CFG^{+}15a, Con05, CL15b, CL15c, CSW17, DL19, DZ09b, DH05, DM15, DS23, DFK^{+}21, ES11, EJ01, FMP08, FT05a, FKT99, FL15, FJK^{+}23, Git99, GSK91, HSZ13, Han09, HH17, HKW15, HT18b, HMSW14, IIL14, KL19a, KZ04, KW08b, KM95, KW91, KMMS21, KM13c, KW96, Kri10, KKS17, LLL17, Lu10, LW20, LMW23, LM11, Luc98, McD15, Més16a, Mol11, NO09, Oze13, Pip01, PSW96, Rah16, RSM^{+}96, RSV^{+}14, SHS23, SSS24, SW23b, VWY09, Vyg16]. Trees [XGG15, ZM02, Zha99, dJMS16, vIKL^{+}16, BCLR89, GMW96, PC97, RTW97, Sag88, Ste88, TP97, Wen97]. Treewidth [ALQ^{+}24, BST20, BKL^{+}22, BM93, BKK95, BJK13, CCVZ10, CDN16, DMS21, DFHT04, GSZ22, HIMW24, KLWY21, MP17b, SSR94, Ram97a]. Trellis [BT96, Kas08]. Triadic [KPS16]. Triangle [AKKR08, AB95, BCL^{+}23, Bar01, BCD21, Bra05, CM24, CCG05, DI22, DPRS10, Dro16, DH20a, DLŠ10, DL17, DM17, DL18c, DL18b, FL96, GKL99, Har19, HT93, IKM99, JS14, KS08c, Liu14, Mer15, NN22, PT94, BH97]. Triangle-Free [BCL^{+}23, DI22, DPRS10, DLŠ10, DL17, DM17, DL18c, DL18b, FL96, GKL99, HT93, IKM99, KS08c, Liu14, NN22, PT94, BH97]. Triangle-Freeness [AKKR08]. Triangles [CDM^{+}14, CFS20, DL18b, Lai18, Pin08, WL02, Yus09]. Triangular [DV96, JKSW17, KŠ08b]. Triangularization [TT93]. Triangulated [CG17, ELR98, GHV06]. Triangulating [IS93, KW92, MWW94a]. Triangulation [Str20]. Triangulations [BBK^{+}16, CFM^{+}09, CEOT15, CEOT17, CGH^{+}10, HTV05, iKMN09, LY21, LQ22, Més16b, MRST16, NNO15, NNO19, Nog24, SS18b, WWKY11, Zhe16]. Trick [Zha11]. Tries [RJS93]. Trilinear [Mac18, MPSS20]. Trinomial [MPSS20]. Trinomials [Din13, DQW^{+}15, WZZ18]. Tripartite [BV22, iKSZ04]. Triple [BZ23, BKKM99, BH15, BCHW17, DR19, GJ16, Gyá19, Hor14, Nag17, NT24, Rif99, COS97]. Triple-Free [BKKM99, COS97]. Triples [AHP19, FS01, Hor14]. Trisection [RT18]. Tropical [AGQS23, AGM23, BK19, BK12, JL20b, Juk16, LY18a, PV22, Sha13, Spe08, ST10]. Tropicalizing [ABGJ15, DR24]. Trotter [BYHR10]. True [CR19]. Truncations [iT12]. Trunk [FJK^{+}19]. Truthful [KVZ23]. Tsetlin [LMS19]. TSP [BMV22, BIT13, CLS15, DL18a, LS03a, New20]. Tucker [LM16]. Tumor [SW23a]. Tunnel [ABCG17]. tuple [Kap14b]. Turán [AKS07, BBBZ12, BLS17, BMNP22, BK14, Dow88, Fue23, GGM^{+}22, GH13, Gyá19, Jan22, JMN22, JS12, JQ20, KKL19, KMV15, LP23, LY23, LM14, LZ18a, LZ09, NY21, NP20b, Nor11, PT14, ST20a, Sha23, Sid18, SV21, ZKNS20]. Turbo [GF08]. Turing [JMS90]. Tutte [GnN06, LW17b, PSS09, Sch22]. Tutte-Type [LW17b]. Tuza [BCD21, HLY23, NvdP24]. Twenty [FM24a, RKDD13]. Twice [ABFRV24]. Twin [AHKiO22]. Twin-Width [AHKiO22]. Twins [BR20, GKM12]. Twist [KS04]. Two [ARTV12, ATPRU91, AHJ^{+}24, Ale10, ABZ15, ADL^{+}09, ABFRV24, BLS17, BMNP22, BPS07, BC09a, BMS12, BP15, BSY21, BCG^{+}10, BGH^{+}17, BGRR18, CDV10, CFS17, CMSM^{+}18, dOCHLO19, CHK17, CLST12, DFJS15, DJW12, DNR23, DTW03, DS05b, DLS11, DL17, DL18b, EA11, EFK05, FM11, FK17, FR94, FIN98, GK07, GMS00, GM03, GW00, Gol96, HJ18, HM20, HKL11, HV17, HK16b, HMP04, KVZ23, Kas03, KL19a, KOS16, KMŠ^{+}09, KL14, KKL19, KS23a, Kla06, Knu95, Lei94, LW03, LZ05a, Lla06, LSL92, LdCKM18, MN18, MP08, PZ05, Pit23, PY09, RTS00, RR18, Red17, RS16, Rya07, Sak94, ST20a, Smi01, Ste00, SV18, WL03, Web08, Yu17, Zak14, ZLWC12, Zun11, CH89, CHW88, DL89b, Exo89, HKKK88, KP06, KLW89, Ko88, Zho88]. Two-Batch [DS05b]. Two-Chain [LSL92]. Two-Color [DFJS15]. Two-Colored [MP08]. Two-Coloring [KMŠ^{+}09]. Two-Colorings [KS23a]. Two-Connected [HV17]. Two-Dimensional [Ale10, ABZ15, BP15, FM11, GW00, Zun11, HM20]. Two-Directional [ATPRU91]. Two-Distance [Yu17]. Two-Facility [KVZ23]. two-factors [HKKK88]. Two-Generator [EA11]. Two-Layer [Lei94]. Two-Part [EFK05]. Two-Party [KOS16]. Two-Period [BCG^{+}10]. two-person [KLW89]. Two-Player [BMS12]. Two-Query [ABFRV24]. Two-Set [Rya07]. Two-Sided [CHK17]. Two-Variable [FK17]. Two-Way [Knu95]. Type [Adr23, AHJ^{+}24, BD01, BV20, CM05a, CGG^{+}16, CIN18, CL22, DS21b, EK10, FHH08, FP22, FKS05a, FS09b, HMM09, Hui23, IN22, KR13, LW17b, MK09, RY23, Špa07, TKMM19, YZ17, AFP^{+}18, KOT09, Lin89, RW89, SG16]. Types [BL23, BRK89]. Typical [BL19a, CCSW23, FKK05, FKK07].

UET [CM05b]. Ultimate [BNR96, Tót10]. Ultrametrics [Dah93]. Unanimity [BFH^{+}08, LT11]. Unavoidable [ADO23, BK05, CDMO16, Kim17]. Unbalanced [AG15, Red17]. Unbounded [BC23a, FKK18, NP18]. Unbreakable [FMOS20]. Uncapacitated [Yam05]. Uncertain [KSF19]. Uncertainty [BDGP22]. Uncoupled [HH92]. Undecidable [Cai18]. Understanding [CFKP22]. underweight [Ko88]. Undirected [BBF99, CS09, ENSZ00, GR99, HSS19, Vég11, NNI97]. Unicyclic [BAH10]. Unidirectional [BCC^{+}05, CFG^{+}09]. Unified [CK08c, DW24, Tak22]. Uniform [Ald90, APS22, BF12, BBLM13, BV22, BCCZ11, BH22, BL01, CFKK17, CKP22, ELSS17, EGM18, FRZ16, FZ22, GXY22, GLM22, GLP23, GSS15, HPS09, Han16a, Han18, HWZZ18, HWZZ20, Jan10, Kha13, KR15, LLM19, Lla06, LZ18a, LYY21, LWY22, LLZ22, OS17, PRS18, RRS07, SV21, Yus03, ZKNS20, HM88]. Uniformity [BDG^{+}17, Sha23, SS21]. Unimodality [CH19, Stu21, HSLd88]. Unimodular [EM20, Gij05, MW19b, OS15]. Union [AS10, BFK^{+}12, GP14, OS16, YAT16, Whi88]. Union-free [BFK^{+}12]. Unions [ALQ^{+}24, GMZ09, Sta92, WZ18]. Unique [CKP13b, FMRR88, HSZ13, JJiT14, KMR95, MMP13]. Unique-Max [HSZ13]. Unique-Maximum [CKP13b]. Unique-Path [MMP13]. Uniquely [Dan01, FJJ18, Siv23]. Uniqueness [BGG^{+}20, GLW11, OS13a]. Unit [BEGK22, Che04, GJ20, GR99, HKW15, JSRSW18, KS23a, LS08, Ric14, Vin11, Kle89]. unit-congruent [Kle89]. Unitary [CW16, dAHFdFK10]. Unity [CGOvZ21]. Universal [AS19, AKTZ19, BS09, BCLR89, BG11, CG17, CHHM09, EGG21, HH13, Pat88, Hur94]. Universality [BS24a, DKRR12, KL14]. Unknot [MRAS19]. Unknown [DKMS17, SW01].

Unraveling [BKL^{+}15]. Unrelated [SS02a]. Unsatisfiable [HS06, Kun23]. Unsolved [LdCKM18]. Untangling [CTU14]. Unweighted [Lev09]. update [FV97]. Updates [GLY07, GK97]. Upper [BST20, BG20, Bra05, CTU14, Car88, CFM94, Däu22, DW10, DW11, EP10, FL22, KKP11, KMÖ18, KW13, KT19, KM95, KOP^{+}21, Lav16, Liu14, LZ09, Mer99, OS16, Raz20, RS22, SC17, SS19, SL95]. Upright [CS91]. Upward [DGL11, DGL10, JMU23]. Urn [CZOS98, CZLfW05, SW10]. Use [BL04, CDN16, KOS16, RS16]. Using [ABY14, BDF^{+}18, BNRZ23, BDG^{+}17, Cai93, CCG^{+}11, EFK05, HH04, HRS93, Jan10, KY21, KNZ14, Kri18, Luc03, MMPS10, NW95, NÖ08, SL96, Str20, Gor93, Tam88]. Utilities [Vaz13].

V [CHM^{+}23]. Vacant [CF16]. Valency [CKNV16]. Valid [BDT17]. Valuated [Mur96a, Mur96b, Tak14]. Valuations [BKK21, PR20]. Value [ABS10, Bol90, BOT92, DM13, Ros09]. Valued [CLLZ18, CCJ^{+}17, FGS19, FK17, GGHP22, TZ15]. values [KP06]. Vanishing [BGJW21]. Variable [BJKV07, BJSS22, FK17, HS06, IM96, Krá06, Sei01, Vin13]. Variable-Sized [Sei01]. Variables [BDP19, BS23, HKR00, Jan10, Lla06, LR94]. Variance [DH05, Str20]. Variant [DEF19, NP20b, UWZ97]. Variation [HTS18]. Variations [MRNS15, MRNS17]. Varieties [BFH^{+}08, DF04, FMP08]. Variety [BK19]. Various [LS24]. Varying [MV18]. VC [BDF^{+}18, BLL^{+}15, HWS18]. VC-Dimension [BLL^{+}15, HWS18]. Vector [BGW20, BN05, CLWZ22, DJKKV23, IK09, KS03d, MPSV21, Vin11, WCLZ15, Web08, XG20]. Vectors [Cha91, CN12, FGLS18, GP08b, GJ12, Jev95, LN17, Mur10, Swa05, Bal88]. Vehicle [GZ98]. Vershik [Śni14]. Version [AK14, dOBMS^{+}17, BGW20, CZLfW05, GJWZ23, Lo14]. Versions [ABC^{+}23]. Verstraëte [GM20]. Versus [BBC^{+}23, Mor21, Tre04, ÀB23, CPRdS13, CFS20, DMS21, MS21b, PPTY22]. Vertex [Age94, ABY11, AS14, ÁS09c, BBF99, BM19, BGLS07, BYFMR10, BYHR10, BMP13, BJS22, BH22, BMR^{+}10, BRK89, CMPS17, CDGO22, CF22, CY21, CHW10, CPPW13, Dan01, DZ09a, FJZ15, Fel90, FSV13, FL10, GKI^{+}23, GM20, GT12, GHP20, Gra07, GSS15, HT90, HLST00, HPS09, HRS18, HMM09, HY12, HKP22, JP18, JPvL21, JL20a, KM19a, iKX20, KvIL^{+}12, Kha13, KG98, Kra18a, KR20, KMN23, Lai05, LiO23, Lic14, LR04, Mac91, McD15, MWW94a, MPSZ19, Mol11, MW90, OS13a, PPR22, Raz10, Tan21, WH15, Zeh17, Zer11, Zha90, o09, vBBC^{+}15, HS89a, Kie97, TP97]. Vertex-Colored [MWW94a]. Vertex-Disjoint [BM19, GM20, Lic14]. Vertex-Edge [CF22]. Vertex-Magic [Gra07]. Vertex-Minors [LiO23]. Vertex-Neighborhood [MW90]. Vertex-Pursuit [BMP13]. Vertex-Robust [Lai05]. Vertex-Transitive [HLST00, Mac91]. Vertical [DGM12]. Vertically [OW22b]. Vertices [Bal08, BHL05, BGH^{+}17, BCG22, CH17, CHLZ00, DL19, DHKM11, DFK^{+}21, EM20, FH21, Fle05, GP18, HT19, KiK12, KiKK17, KKSvL20, LM12, Pad16, PVY22, Ste10, ZL11, Stu88]. Very [Sol06]. Vestergaard [HLY23]. VI [SSS22]. Via [ACG94, AM06a, AFS12, AB00, AL17, BD20a, BMS12, BV10, CHM^{+}23, CKL^{+}21, CNR89, CFKP22, DSL19, FPS20, FGKP22, Fuk16, Ful14, GPST15, GS20, HKL99, HTS18, IKK^{+}22, JLR20, KLMR13, KT17, KLL13, LW88a, LL14a, LL17, LLS04, Lu10, NT12, Prz13, Wro20, WC12, Yam16]. Vietoris [Shu23]. View [CPS08, CK08c, FKM^{+}16, FGKP22, Mat19]. Viewpoint [FLM^{+}18]. Violation [iKX20]. Virtual [HKS07]. Visibility [ABHW13, CWY21, CHJ^{+}04, HVW07, HZ08, IKK06, LLS04]. Visible [RT11]. Vision [DKSŻ10]. Vision-Based [DKSŻ10]. Visits [BMV22]. Visual [BDDS03]. Vizing [KK14a]. VLSI [FL92]. Volume [CL06, CFG^{+}15b]. Volumes [LRR21, TZ19, TZ20]. Voronoi [KvRRS23, MGC14, SGM20, Wor88]. Voter [Fer24]. Voting [AK10, Bol90, CDFR18, FKM^{+}16, IIL14, CEOR13]. VPN [FOST10]. vs [sJS24].

Wait [Svi03]. Wakeup [GPP01]. Walk [Ald90, CF16]. Walker [EFKP15]. Walkers [DPSW08]. Walks [AH11, BF96, Bev09, BILL23, Cha19, CP10a, CFR10, CF10, CEOR13, CTW93, CFS96, Die10, DJS12, FGLP14, FKK18, GR17, JP12, KR04, MW19a, SYKY18]. Walksat [COHH17]. Wandering [Vin07]. Wang [Ull14]. Water [FMV22]. Water-Filling [FMV22]. Watson [BDT17]. Wave [RT09]. Wavelength [BCC^{+}05, BCG^{+}10, Car09]. Wavelength-Division [BCC^{+}05]. Wax [ACM11]. Way [EMM14, Knu95, BJ92]. WDM [Car09, CFG^{+}09]. Weak [AK14, DF02, GT13, HH13, IP91, SS19, Abe91]. Weakly [AGH11]. Weakness [GT98]. Weaving [ACGH20]. Web [Pip06]. Weber [LY18a]. Weight [Adr23, AP14, BB13, BM00, Bar02, BHH96, BK91, BB03, BLM10, CCD00, CER98, CDWZ17, CHZ09, CGK^{+}19, CDHZ12, CPPT20, DM13, FPS18, JT11, Kap14b, LS14, LCV03, MMSJ08, RF12, RSW12, SE14, BBM90, CZ97]. Weighted [ACZ23, ADL13, ANP14, BB16, BS95a, BBCZ11, BTW08, BM16, BG12, CDM^{+}14, Dei15, Die10, Duc21, DJMV21, ELMS11, Fri21, FP13, Gal18, GKMS22, GNS11, GY23, HRS17, Jon20, KíP^{+}24, LOW10, MC93, BG88, FZ88]. Weights [BJKV07, CM07, FPS18, Hås94, Krá06, LWW10, SB91, VV94]. Weisfeiler [FKV21, KN22]. Well [BN05, COHH17, FL92, Lag00, MR21, iO08, Rif99]. Well-Covered [BN05]. Well-Ordered [Rif99]. Well-Partial-Order [FL92]. Well-Quasi-Ordering [iO08]. Well-Spaced [Lag00]. Wheel [AFH^{+}18, CdV02, Cho94a]. Wheels [OSW16]. Wheels-and-Whirls [OSW16]. Which [PPTY22, BJS22, Erd17, HS89b]. Whirls [OSW16]. Whitney [FG21]. Who [KSW17]. Whose [Jev95]. Wide [FFP88]. Wide-sense [FFP88]. Widely [LSO03]. Width [AHKiO22, AHP09, BDJ^{+}20, BC23a, DDJ^{+}21, DD13, DMN12, FRRS09, FGT11, GKS22, GnN06, HW16, HKW22, Hli18, KM14, Lam20, LR04, Luc03, iO08, San96]. Widths [Adl08]. Wiedemann [KMÖ18]. Wiener [FL15, LW03]. Williams [HSV23]. Win [KSW17]. Windows [GOR20]. Windy [CPS08]. Winkler [GG15]. Winners [KMR95]. Wired [Cha19]. Wirelength [CKK^{+}04]. Wiring [SW01, Pat88]. Wise [GNPV20, Vse05]. Without [CDM00, CL07, CH06a, CEOT15, CEOT17, CS12, DJW12, DL14, MNS14, BBC^{+}19b, BH15, CDM04, CPPT20, DMR11, DX19, ES11, Far09, FKP15, GCH22, GS03, JPY10, JKSW17, KS23a, KSY18, Kle89, KLN10, LLS^{+}20, Tan21, WL02, WX13, ZLWC12, sG24]. Word [BK05, Sch04]. Word-Run [Sch04]. Words [BM14, BH20, Fra10, KM21b, KRR16, MRR20, MR21, RMS01]. World [JM17]. Worst [ASZ02, PV10b, SS95, Tak08, Tas97, HS89b]. Worst-Case [PV10b, SS95, Tak08, HS89b]. Wreath [Ath14].

X [DGP06]. X-rays [DGP06].

Yeh [HRS12]. Yeo [Lic14]. Yields [GVW06]. Yong [FGS19]. Young [DZ09b, DFG^{+}21, PV10a].

Zagier [CR16]. Zeckendorf [BAM16]. Zero [CKK^{+}04, CK91, DF94, DK10, DJ11, Eti20, FZ08, HZ21, KR16, KKS19, LXZZ08, MS17a, MLMR23, SA90, WYZZ14, ZM02, ALZ96, NS89]. Zero- [ZM02]. Zero-Free [DK10, DJ11, Eti20]. Zero-Knowledge [DF94]. Zero-One [CK91, SA90, NS89]. Zero-Sum [HZ21, MLMR23]. Zeros [Bón09, DK06]. Zeta [Dei15]. Zhao [CR19]. Zippel [DEMT22]. Ziv [CI07]. Zone [Zer11]. Zonotopal [MPSV21].

References

Alon:1988:SAS

[AA88] Noga Alon and Yossi Azar. Sorting, approximate sorting, and searching in rounds. SIAM Journal on Discrete Mathematics, 1(3):269–280, August 1988. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Aichholzer:1996:CHH

[AA96] Oswin Aichholzer and Franz Aurenhammer. Classifying hyperplanes in hypercubes. SIAM Journal on Discrete Mathematics, 9(2): 225–232, May 1996. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Alon:2005:LHS

[AA05] Noga Alon and Vera Asodi. Learning a hidden subgraph. SIAM Journal on Discrete Mathematics, 18 (4):697–712, 2005. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/43107.

Adamaszek:2010:LGR

[AA10] Anna Adamaszek and Micha Adamaszek. Large-girth roots of graphs. SIAM Journal on Discrete Mathematics, 24(4): 1501–1514, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ahlswede:2011:SSD

[AA11] R. Ahlswede and H. Aydinian. On security of statistical databases. SIAM Journal on Discrete Mathematics, 25(4):1778–1791, ???? 2011. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v25/i4/p1778_s1.

Abel:2018:CFC

[AAD^{+}18] Zachary Abel, Victor Alvarez, Erik D. Demaine, Sándor P. Fekete, Aman Gour, Adam Hesterberg, Phillip Keldenich, and Christian Scheffer. Conflict-free coloring of graphs. SIAM Journal on Discrete Mathematics, 32(4):2675–2702, ???? 2018. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Alfaro:2018:CNC

[AADM18] Carlos A. Alfaro, Alan Arroyo, Marek Dernár, and Bojan Mohar. The crossing number of the cone of a graph. SIAM Journal on Discrete Mathematics, 32(3): 2080–2093, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Akbari:2006:MPI

[AAFL06] S. Akbari, A. Alipour, H. L. Fu, and Y. H. Lo. Multicolored parallelisms of isomorphic spanning trees. SIAM Journal on Discrete Mathematics, 20(3):564–567, January 2006. CODEN SJD-MEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Abrego:2018:BD

[ÁAFM^{+}18] Bernardo M. Ábrego, Oswin Aichholzer, Silvia Fernández-Merchant, Dan McQuillan, Bojan Mohar, Petra Mutzel, Pedro Ramos, R. Bruce Richter, and Birgit Vogtenhuber. Bishellable drawings of K_{n}. SIAM Journal on Discrete Mathematics, 32(4): 2482–2492, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Alon:2014:MNN

[AAH14] Noga Alon, Harout Aydinian, and Hao Huang. Maximizing the number of nonnegative subsets. SIAM Journal on Discrete Mathematics, 28 (2):811–816, ???? 2014. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Abay-Asmerom:2010:DPF

[AAHLT10] Ghidewon Abay-Asmerom, Richard H. Hammack, Craig E. Larson, and Dewey T. Taylor. Direct product factorization of bipartite graphs with bipartition-reversing involutions. SIAM Journal on Discrete Mathematics, 23(4): 2042–2052, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Alqasem:2024:CFD

[AAMM24] Abdulmajeed Alqasem, Heshan Aravinda, Arnaud Marsiglietti, and James Melbourne. On a conjecture of feige for discrete log-concave distributions. SIAM Journal on Discrete Mathematics, 38(1):93–102, January 2024. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Aboulker:2022:HOC

[AAS22] Pierre Aboulker, Guillaume Aubian, and Raphael Steiner. Heroes in orientations of chordal graphs. SIAM Journal on Discrete Mathematics, 36(4):2497–2505, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1481427.

Alon:2009:CRG

[AaW09] Noga Alon, Pawe Pra at, and Nicholas Wormald. Cleaning regular graphs with brushes. SIAM Journal on Discrete Mathematics, 23(1): 233–250, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Alon:1994:ECD

[AB94] Noga Alon and Jehoshua Bruck. Explicit constructions of depth-2 majority circuits for comparison and addition. SIAM Journal on Discrete Mathematics, 7(1): 1–8, February 1994. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Andreae:1995:PGA

[AB95] Thomas Andreae and Hans-Jürgen Bandelt. Performance guarantees for approximation algorithms depending on parametrized triangle inequalities. SIAM Journal on Discrete Mathematics, 8 (1):1–16, February 1995. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Annexstein:2000:DRG

[AB00] Fred S. Annexstein and Kenneth A. Berman. Directional routing via generalized st-numberings. SIAM Journal on Discrete Mathematics, 13(2):268–279, 2000. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/33329.

Albers:2005:DTA

[AB05] Susanne Albers and Helge Bals. Dynamic TCP acknowledgment: Penalizing long delays. SIAM Journal on Discrete Mathematics, 19(4):938–951, 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/44165.

Axenovich:2007:GHS

[AB07] Maria Axenovich and József Balogh. Graphs having small number of sizes on induced k-subgraphs. SIAM Journal on Discrete Mathematics, 21 (1):264–272, ???? 2007. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Anastos:2018:PDH

[AB18] Michael Anastos and Joseph Briggs. Packing directed Hamilton cycles online. SIAM Journal on Discrete Mathematics, 32(2):1505–1539, ????

2018. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Alvarez:2023:PCT

[ÀB23] Carme Àlvarez and Arnau Messegué Buisan. On the PoA conjecture: Trees versus biconnected components. SIAM Journal on Discrete Mathematics, 37 (2):1030–1052, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1466426.

Adiga:2011:BPD

[ABC11] Abhijin Adiga, Diptendu Bhowmick, and L. Sunil Chandran. Boxicity and poset dimension. SIAM Journal on Discrete Mathematics, 25(4):1687–1698, ???? 2011. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v25/i4/p1687_s1.

Alon:2015:SDB

[ABC^{+}15] Noga Alon, Manu Basavaraju, L. Sunil Chandran, Rogers Mathew, and Deepak Rajendraprasad. Separation dimension of bounded degree graphs. SIAM Journal on Discrete Mathematics, 29(1): 59–64, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Aharoni:2023:NDR

[ABC^{+}23] Ron Aharoni, Eli Berger, Maria Chudnovsky, He Guo, and Shira Zerbib. Nonuniform degrees and rainbow versions of the Caccetta– Häggkvist conjecture. SIAM Journal on Discrete Mathematics, 37(3):1704–1714, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL

https://epubs.siam.org/
doi/10.1137/22M1529658.

Ardila:2017:CSR

[ABCG17] Federico Ardila, Hanner Bastidas, Cesar Ceballos, and John Guo. The configuration space of a robotic arm in a tunnel. SIAM Journal on Discrete Mathematics, 31(4): 2675–2702, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Abello:1991:WBO

[Abe91] James Abello. The weak Bruhat order of S_{Σ} consistent sets, and Catalan numbers. SIAM Journal on Discrete Mathematics, 4(1): 1–16, February 1991. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Afek:2014:MC

[ABF^{+}14] Yehuda Afek, Yakov Babichenko, Uriel Feige, Eli Gafni, Nati Linial, and Benny Sudakov. Musical chairs. SIAM Journal on Discrete Mathematics, 28(3):1578–1600, ???? 2014. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Amanatidis:2024:DRD

[ABFRV24] Georgios Amanatidis, Georgios Birmpas, Aris Filos-Ratsikas, and Alexandros A. Voudouris. Don’t roll the dice, ask twice: The twoquery distortion of matching problems and beyond. SIAM Journal on Discrete Mathematics, 38(1):1007– 1029, March 2024. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ambrus:2023:PC

[ABFV23] Gergely Ambrus, Imre Bárány, Péter Frankl, and Dániel Varga. Piercing the chessboard. SIAM Journal on Discrete Mathematics, 37 (3):1457–1471, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M146048X.

Allamigeon:2015:TSA

[ABGJ15] Xavier Allamigeon, Pascal Benchimol, Stéphane Gaubert, and Michael Joswig. Tropicalizing the simplex algorithm. SIAM Journal on Discrete Mathematics, 29(2): 751–795, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ardila:2011:RPG

[ABH^{+}11] Federico Ardila, Matthias Beck, Serkan Hosten, Julian Pfeifle, and Kim Seashore. Root polytopes and growth series of root lattices. SIAM Journal on Discrete Mathematics, 25(1):360–378, ???? 2011. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:

//epubs.siam.org/sidma/
resource/1/sjdmec/v25/i1/
p360_s1.

Aharoni:2021:ROC

[ABHJ21] Ron Aharoni, Joseph Briggs, Ron Holzman, and Zilin Jiang. Rainbow odd cycles. SIAM Journal on Discrete Mathematics, 35(4): 2293–2303, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Aldred:2000:NCC

[ABHM00] R. E. L. Aldred, S. Bau, D. A. Holton, and Brendan D. McKay. Nonhamiltonian 3-connected cubic planar graphs. SIAM Journal on Discrete Mathematics, 13(1):25–32, 2000.

CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/34866.

Axenovich:2013:VND

[ABHW13] Maria Axenovich, Andrew Beveridge, Joan P. Hutchinson, and Douglas B. West. Visibility number of directed graphs. SIAM Journal on Discrete Mathematics, 27(3): 1429–1449, ???? 2013. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Avin:2020:DBG

[ABL^{+}20] Chen Avin, Marcin Bienkowski, Andreas Loukas, Maciej Pacut, and Stefan Schmid. Dynamic balanced graph partitioning. SIAM Journal on Discrete Mathematics, 34(3):1791–1812, ???? 2020. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ahmady:2014:ICG

[ABM14] Azhvan Ahmady, Jason P. Bell, and Bojan Mohar. Integral Cayley graphs and groups. SIAM Journal on Discrete Mathematics, 28(2): 685–701, ???? 2014. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Alfarano:2022:TCP

[ABNR22] Gianira N. Alfarano, Martino Borello, Alessandro Neri, and Alberto Ravagnani. Three combinatorial perspectives on minimal codes. SIAM Journal on Discrete Mathematics, 36(1):461–489, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1391493.

Alstrup:2005:LSS

[ABR05] Stephen Alstrup, Philip Bille, and Theis Rauhe. Labeling schemes for small distances in trees. SIAM Journal on Discrete Mathematics, 19(2):448–462, 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/43340.

Algaba:2010:VGR

[ABS10] E. Algaba, J. M. Bilbao, and M. Slikker. A value for games restricted by augmenting systems. SIAM Journal on Discrete Mathematics, 24 (3):992–1010, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Abueida:2013:HSI

[ABS13] Atif Abueida, Arthur Busch, and R. Sritharan. Hamiltonian spider intersection graphs are cycle extendable. SIAM Journal on Discrete Mathematics, 27(4): 1913–1923, ???? 2013. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Anstee:2011:PMG

[ABY11] R. P. Anstee, J. Blackman, and Hangjun Yang. Perfect matchings in grid graphs after vertex deletions. SIAM Journal on Discrete Mathematics, 25(4): 1754–1767, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v25/i4/p1754_s1.

Ardila:2014:MRE

[ABY14] Federico Ardila, Tia Baker, and Rika Yatchak. Moving robots efficiently using the combinatorics of CAT(0) cubical complexes. SIAM Journal on Discrete Mathematics, 28(2):986–1007, ???? 2014. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Alekseyev:2015:MTS

[ABZ15] Max A. Alekseyev, Marina G. Basova, and Nikolai Yu. Zolotykh. On the minimal teaching sets of two-dimensional threshold functions. SIAM Journal on Discrete Mathematics, 29(1):157–165, ???? 2015. CODEN SJDMEC.

ISSN 0895-4801 (print), 1095-7146 (electronic).

Adler:1990:APD

[AC90] I. Adler and S. Cosares.

Advantageous properties of dual transhipment polyhedra. SIAM Journal on Discrete Mathematics, 3(2): 157–167, May 1990. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Adiga:2014:RCG

[AC14] Abhijin Adiga and L. Sunil Chandran. Representing a cubic graph as the intersection graph of axis-parallel boxes in three dimensions. SIAM Journal on Discrete Mathematics, 28(3):1515–1539, ???? 2014. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Arvind:2008:CDN

[ACD08] V. Arvind, Christine T.

Cheng, and Nikhil R. Devanur. On computing the distinguishing numbers of planar graphs and beyond: a counting approach. SIAM Journal on Discrete Mathematics, 22(4):1297–1324, ???? 2008. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Averkov:2013:CAH

[ACD^{+}13] Gennadiy Averkov, Michele Conforti, Alberto Del Pia, Marco Di Summa, and Yuri Faenza. On the convergence of the affine hull of the Chvátal–Gomory closures. SIAM Journal on Discrete Mathematics, 27(3): 1492–1502, ???? 2013. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Aprile:2018:LPA

[ACF18] Manuel Aprile, Alfonso Cevallos, and Yuri Faenza. On 2-level polytopes arising in combinatorial settings. SIAM Journal on Discrete Mathematics, 32(3):1857–1886, ???? 2018. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Atminas:2016:DBN

[ACFL16] Aistis Atminas, Andrew Collins, Jan Foniok, and Vadim V. Lozin. Deciding the Bell number for hereditary graph properties. SIAM Journal on Discrete Mathematics, 30(2):1015–1031, ???? 2016. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Alon:1994:RPG

[ACG94] Noga Alon, F. R. K. Chung, and R. L. Graham. Routing permutations on graphs via matchings. SIAM Journal on Discrete Mathematics, 7(3): 513–530, August 1994. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Alon:2020:MCH

[ACG^{+}20] Noga Alon, Jonathan D. Cohen, Thomas L. Griffiths, Pasin Manurangsi, Daniel Reichman, Igor Shinkar, Tal Wagner, and Alexander Yu. Multitasking capacity: Hardness results and improved constructions. SIAM Journal on Discrete Mathematics, 34 (1):885–903, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Akleman:2020:TCT

[ACGH20] Ergun Akleman, Jianer Chen, Jonathan L. Gross, and Shiyu Hu. A topologically complete theory of weaving. SIAM Journal on Discrete Mathematics, 34(4):2457–2480, ???? 2020. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Abdi:2022:CCD

[ACGT22] Ahmad Abdi, Gérard Cornuéjols, Bertrand Guenin, and Levent Tunçel. Clean clutters and dyadic fractional packings. SIAM Journal on Discrete Mathematics, 36 (2):1012–1037, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1397325.

Agarwal:2019:EGB

[ACKM19] Naman Agarwal, Karthekeyan Chandrasekaran, Alexandra Kolla, and Vivek Madan. On the expansion of groupbased lifts. SIAM Journal on Discrete Mathematics, 33(3): 1338–1373, ???? 2019. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Arias:2006:CRN

[ACL^{+}06] Marta Arias, Lenore J. Cowen, Kofi A. Laing, Rajmohan Rajaraman, and Orjeta Taka. Compact routing with name independence. SIAM Journal on Discrete Mathematics, 20(3): 705–726, January 2006. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Alon:2001:EBB

[ACLT01] Noga Alon, Charles J. Colbourn, Alan C. H. Ling, and Martin Tompa. Equireplicate balanced binary codes for oligo arrays. SIAM Journal on Discrete Mathematics, 14(4):481–497, 2001. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/38389.

Adamaszek:2018:ASC

[ACLW18] Anna Adamaszek, Artur Czumaj, Andrzej Lingas, and Jakub Onufry Wojtaszczyk. Approximation schemes for capacitated geometric network design. SIAM Journal on Discrete Mathematics, 32(4):2720–2746, ???? 2018. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Amdeberhan:2011:BBM

[ACM11] Tewodros Amdeberhan, Mahír Bílen Can, and Victor H. Moll. Broken bracelets, Molien series, paraffin wax, and an elliptic curve of conductor 48. SIAM Journal on Discrete Mathematics, 25(4): 1843–1859, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v25/i4/p1843_s1.

Alsulmi:2018:GGK

[ACM^{+}18] Badria Alsulmi, Todd Cochrane, Michael J. Mossinghoff, Vincent Pigno, Chris Pinner, C. J. Richardson, and Ian Thompson. A generalization of the Goresky–Klapper conjecture, Part I. SIAM Journal on Discrete Mathematics, 32(4):2973–3002, ???? 2018. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Acan:2017:PPR

[ACMW17] Hüseyin Acan, Andrea Collevecchio, Abbas Mehrabian, and Nick Wormald. On the Push&Pull protocol for rumor spreading. SIAM Journal on Discrete Mathematics, 31 (2):647–668, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Abdi:2022:DFP

[ACP22] Ahmad Abdi, Gérard Cornuéjols, and Zuzanna Palion. On dyadic fractional packings of T-joins. SIAM Journal on Discrete Mathematics, 36 (3):2445–2451, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1445260.

Avart:2007:EMG

[AcRS07] Christian Avart, Vojt ch Rödl, and Mathias Schacht. Every monotone 3-graph property is testable. SIAM Journal on Discrete Mathematics, 21(1):73–92, ???? 2007. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Andrews:1997:SAS

[ACS97] George E. Andrews, Davide Crippa, and Klaus Simon. qseries arising from the study of random graphs. SIAM Journal on Discrete Mathematics, 10(1):41–56, February 1997. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Abdi:2023:PDD

[ACZ23] Ahmad Abdi, Gérard Cornuéjols, and Michael Zlatin. On packing dijoins in digraphs and weighted digraphs. SIAM Journal on Discrete Mathematics, 37(4):2417–2461, October 2023. CODEN SJD-MEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Aleksandrov:1996:LAP

[AD96] L. Aleksandrov and H. Djidjev. Linear algorithms for partitioning embedded graphs of bounded genus. SIAM Journal on Discrete Mathematics, 9(1):129–150, February 1996. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Agnarsson:2011:NMN

[AD11] Geir Agnarsson and Jill Bigley Dunham. A note on the maximum number of edges of nonflowerable coin graphs. SIAM Journal on Discrete Mathematics, 25(4):1714– 1721, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v25/i4/p1714_s1.

Alon:2014:CBN

[ADH^{+}14] Noga Alon, Erik D. Demaine, MohammadTaghi Hajiaghayi, Panagiotis Kanellopoulos, and Tom Leighton. Correction: Basic network creation games. SIAM Journal on Discrete Mathematics, 28(3):1638–1640, ???? 2014.

CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). See [ADHL13].

Alon:2013:BNC

[ADHL13] Noga Alon, Erik D. Demaine, Mohammad T. Hajiaghayi, and Tom Leighton. Basic network creation games. SIAM Journal on Discrete Mathematics, 27(2):656–668, ???? 2013. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). See correction [ADH^{+}14].

Ashok:2018:EFP

[ADKS18] Pradeesha Ashok, Aditi Dudeja, Sudeshna Kolay, and Saket Saurabh. Exact and fixed parameter tractable algorithms for maxconflict-free coloring in hypergraphs. SIAM Journal on Discrete Mathematics, 32(2): 1189–1208, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Adler:2008:TRW

[Adl08] Isolde Adler. Tree-related widths of graphs and hypergraphs. SIAM Journal on Discrete Mathematics, 22(1): 102–123, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Alverson:2009:DLD

[ADL^{+}09] L. Wyatt Alverson II, Robert G. Donnelly, Scott J. Lewis, Marti McClard, Robert Pervine, Robert A. Proctor, and N. J. Wildberger. Distributive lattices defined for representations of rank two semisimple Lie algebras. SIAM Journal on Discrete Mathematics, 23(1): 527–559, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Amini:2013:FWS

[ADL13] Hamed Amini, Moez Draief, and Marc Lelarge. Flooding in weighted sparse random graphs. SIAM Journal on Discrete Mathematics, 27 (1):1–26, ???? 2013. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Acharya:2015:SRS

[ADM^{+}15] Jayadev Acharya, Hirakendu Das, Olgica Milenkovic, Alon Orlitsky, and Shengjun Pan. String reconstruction from substring compositions. SIAM Journal on Discrete Mathematics, 29(3): 1340–1371, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Aidun:2021:GSG

[ADM^{+}21] Ivan Aidun, Frances Dean, Ralph Morrison, Teresa Yu, and Julie Yuan. Gonality sequences of graphs. SIAM Journal on Discrete Mathematics, 35(2):814–839, ???? 2021. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Allred:2023:UIS

[ADO23] Sarah Allred, Guoli Ding, and Bogdan Oporowski. Unavoidable induced subgraphs of large 2-connected graphs. SIAM Journal on Discrete Mathematics, 37 (2):684–698, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1447593.

Alon:2003:TC

[ADPR03] Noga Alon, Seannie Dar, Michal Parnas, and Dana Ron. Testing of clustering. SIAM Journal on Discrete Mathematics, 16 (3):393–417, 2003. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/41097.

Adriaensen:2023:NSW

[Adr23] Sam Adriaensen. A note on small weight codewords of projective geometric codes and on the smallest sets of even type. SIAM Journal on Discrete Mathematics, 37 (3):2072–2087, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1556277.

Azizouglu:2003:ESM

[AE03] M. Cemil Azizouglu and Ömer Eugeciouglu. Extremal sets minimizing dimensionnormalized boundary in Hamming graphs. SIAM Journal on Discrete Mathematics, 17(2):219–236, 2003. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/37505.

Azar:2004:LLB

[AE04] Yossi Azar and Leah Epstein. On-line load balancing of temporary tasks on identical machines. SIAM Journal on Discrete Mathematics, 18(2):347–352, 2004. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/32929.

Alon:2013:ALG

[AEFT13] Noga Alon, Yuval Emek, Michal Feldman, and Moshe Tennenholtz. Adversarial leakage in games. SIAM Journal on Discrete Mathematics, 27(1):363–385, ???? 2013. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Alon:2024:RNR

[AEPT24] Noga Alon, Dor Elboim, János Pach, and Gábor Tardos. Random necklaces require fewer cuts. SIAM Journal on Discrete Mathematics, 38(2):1381–1408, April 2024. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Abrams:2005:GBD

[AF05] Lowell Abrams and Donniell E. Fishkind. A genus bound for digital image boundaries. SIAM Journal on Discrete Mathematics, 19(3):807–813, 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/44569.

Alon:2010:BMI

[AF10] Noga Alon and Ohad N. Feldheim. The Brunn– Minkowski inequality and nontrivial cycles in the discrete torus. SIAM Journal on Discrete Mathematics, 24 (3):892–894, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Anastos:2019:PCH

[AF19] Michael Anastos and Alan Frieze. Pattern colored Hamilton cycles in random graphs. SIAM Journal on Discrete Mathematics, 33(1): 528–545, ???? 2019. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Alon:2009:SDT

[AFG^{+}09] Noga Alon, Fedor V. Fomin, Gregory Gutin, Michael Krivelevich, and Saket Saurabh. Spanning directed trees with many leaves. SIAM Journal on Discrete Mathematics, 23 (1):466–476, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Abdi:2016:LTD

[AFG^{+}16] Ahmad Abdi, Andreas Emil Feldmann, Bertrand Guenin, Jochen Könemann, and Laura Sanità. Lehman’s theorem and the directed Steiner tree problem. SIAM Journal on Discrete Mathematics, 30 (1):141–153, ???? 2016. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Anastos:2021:HRG

[AFG21] Michael Anastos, Alan Frieze, and Pu Gao. Hamiltonicity of random graphs in the stochastic block model. SIAM Journal on Discrete Mathematics, 35(3):1854–1880, ???? 2021. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Aboulker:2018:TEP

[AFH^{+}18] Pierre Aboulker, Samuel Fiorini, Tony Huynh, Gwenaël Joret, Jean-Florent Raymond, and Ignasi Sau. A tight Erdős–Pósa function for wheel minors. SIAM Journal on Discrete Mathematics, 32(3):2302–2312, ???? 2018. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Alexeev:2012:FIS

[AFK12] Boris Alexeev, Alexandra Fradkin, and Ilhee Kim. Forbidden induced subgraphs of double-split graphs. SIAM Journal on Discrete Mathematics, 26(1):1–14, ???? 2012. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:

//epubs.siam.org/sidma/
resource/1/sjdmec/v26/i1/
p1_s1.

Alecu:2022:LGG

[AFK^{+}22] Bogdan Alecu, Robert Ferguson, Mamadou Moustapha Kanté, Vadim V. Lozin, Vincent Vatter, and Victor Zamaraev. Letter graphs and geometric grid classes of permutations. SIAM Journal on Discrete Mathematics, 36 (4):2774–2797, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1449646.

Alon:2023:SPL

[AFKK23] Noga Alon, Ehud Friedgut, Gil Kalai, and Guy Kindler.

The success probability in Levine’s hat problem, and independent sets in graphs. SIAM Journal on Discrete Mathematics, 37(4):2717– 2729, November 2023. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Agrawal:2020:PCF

[AFL^{+}20] Akanksha Agrawal, Fedor V.

Fomin, Daniel Lokshtanov, Saket Saurabh, and Prafullkumar Tale. Path contraction faster than 2^{n}. SIAM Journal on Discrete Mathematics, 34(2):1302–1325, ???? 2020. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Asada:2018:FDG

[AFP^{+}18] Megumi Asada, Florian Frick, Vivek Pisharody, Maxwell Polevy, David Stoner, Ling Hei Tsang, and Zoe Wellner. Fair division and generalizations of Sperner- and KKM-type results. SIAM Journal on Discrete Mathematics, 32(1): 591–610, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Amini:2012:CSH

[AFS12] Omid Amini, Fedor V. Fomin, and Saket Saurabh. Counting subgraphs via homomorphisms. SIAM Journal on Discrete Mathematics, 26(2): 695–717, ???? 2012. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ackerman:2012:GAP

[AFT12] Eyal Ackerman, Radoslav Fulek, and Csaba D. Tóth. Graphs that admit polyline drawings with few crossing angles. SIAM Journal on Discrete Mathematics, 26(1): 305–320, ???? 2012. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Alpers:2006:SEC

[AG06] Andreas Alpers and Peter Gritzmann. On stability, error correction, and noise compensation in discrete tomography. SIAM Journal on Discrete Mathematics, 20(1): 227–239, January 2006. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Apollonio:2015:MUD

[AG15] Nicola Apollonio and Anna Galluccio. Minimally unbalanced diamond-free graphs and Dyck-paths. SIAM Journal on Discrete Mathematics, 29(4):1837–1863, ???? 2015. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Alpers:2018:DRI

[AG18] Andreas Alpers and Peter Gritzmann. On doubleresolution imaging and discrete tomography. SIAM Journal on Discrete Mathematics, 32(2):1369–1399, ???? 2018. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ayre:2019:RCH

[AG19] Peter Ayre and Catherine Greenhill. Rigid colorings of hypergraphs and contiguity. SIAM Journal on Discrete Mathematics, 33(3): 1575–1606, ???? 2019. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ageev:1994:FCI

[Age94] Alexander A. Ageev. On finding critical independent and vertex sets. SIAM Journal on Discrete Mathematics, 7 (2):293–295, May 1994. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Amini:2011:SWQ

[AGH11] Omid Amini, Simon Griffiths, and Florian Huc. Subgraphs of weakly quasi-random oriented graphs. SIAM Journal on Discrete Mathematics, 25 (1):234–259, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v25/i1/p234_s1.

Alcon:2014:PSG

[AGH14] Liliana Alcón, Marisa Gutierrez, and Glenn Hurlbert. Pebbling in split graphs. SIAM Journal on Discrete Mathematics, 28(3):1449–1466, ???? 2014. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Alon:2023:SCG

[AGK^{+}23] Noga Alon, Anna Gujgiczer, János Körner, Aleksa Milojević, and Gábor Simonyi. Structured codes of graphs. SIAM Journal on Discrete Mathematics, 37(1):379–403, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1487989.

Allamigeon:2023:TCP

[AGM23] Xavier Allamigeon, Stéphane Gaubert, and Frédéric Meunier. Tropical complementarity problems and Nash equilibria. SIAM Journal on Discrete Mathematics, 37 (3):1645–1665, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1446861.

Akian:2023:TLR

[AGQS23] Marianne Akian, Stéphane Gaubert, Yang Qi, and Omar Saadi. Tropical linear regression and mean payoff games:

Or, how to measure the distance to equilibria. SIAM Journal on Discrete Mathematics, 37(2):632–674, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL

https://epubs.siam.org/
doi/10.1137/21M1428297.

Albertson:1996:BFR

[AH96] Michael O. Albertson and Ruth Haas. Bounding functions and rigid graphs. SIAM Journal on Discrete Mathematics, 9(2):269–273, May 1996. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Agnarsson:2003:CPP

[AH03] Geir Agnarsson and Magnús M.

Halldórsson. Coloring powers of planar graphs. SIAM Journal on Discrete Mathematics, 16(4):651–662, 2003. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/36795.

Angel:2011:RWG

[AH11] Omer Angel and Alexander E. Holroyd. Rotor walks on general trees. SIAM Journal on Discrete Mathematics, 25(1):423–446, ???? 2011. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.

                                        siam.org/sidma/resource/
                                        1/sjdmec/v25/i1/p423_s1.

Aksoy:2016:GMS

[AH16] Sinan Aksoy and Paul Horn. Graphs with many strong orientations. SIAM Journal on Discrete Mathematics, 30(2): 1269–1282, ???? 2016. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Alon:2021:AGF

[AH21] Tzvi Alon and Nir Halman. Automatic generation of FP-TASes for stochastic monotone dynamic programs made easier. SIAM Journal on Discrete Mathematics, 35(4): 2679–2722, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Aigner-Horev:2022:LRC

[AHDHL22] Elad Aigner-Horev, Oran Danon, Dan Hefetz, and Shoham Letzter. Large rainbow cliques in randomly perturbed dense graphs. SIAM Journal on Discrete Mathematics, 36(4):2975–2994, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL

                                        https://epubs.siam.org/
                                        doi/10.1137/21M1423117.

Armbruster:2008:GBC

[AHFM08] Michael Armbruster, Christoph Helmberg, Marzena Fügenschuh, and Alexander Martin. On the graph bisection cut polytope. SIAM Journal on Discrete Mathematics, 22(3): 1073–1098, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Aichholzer:2010:LBP

[AHH^{+}10] Oswin Aichholzer, Thomas Hackl, Clemens Huemer, Ferran Hurtado, and Birgit Vogtenhuber. Large bichromatic point sets admit empty monochromatic 4-gons. SIAM Journal on Discrete Mathematics, 23(4): 2147–2155, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Aigner-Horev:2021:RHC

[AHH21] Elad Aigner-Horev and Dan Hefetz. Rainbow Hamilton cycles in randomly colored randomly perturbed dense graphs. SIAM Journal on Discrete Mathematics, 35(3): 1569–1577, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Albrechtsen:2024:MTT

[AHJ^{+}24] Sandra Albrechtsen, Tony Huynh, Raphael W. Jacobs, Paul Knappe, and Paul Wollan. A Mengertype theorem for two induced paths. SIAM Journal on Discrete Mathematics, 38(2): 1438–1450, May 2024. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ahn:2022:BTW

[AHKiO22] Jungho Ahn, Kevin Hendrey, Donggyu Kim, and Sang il Oum. Bounds for the twinwidth of graphs. SIAM Journal on Discrete Mathematics, 36(3):2352–2366, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1452834.

Amini:2009:PWP

[AHP09] Omid Amini, Florian Huc, and Stéphane Pérennes. On the path-width of planar graphs. SIAM Journal on Discrete Mathematics, 23(3): 1311–1316, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Aigner-Horev:2019:MST

[AHP19] Elad Aigner-Horev and Yury Person. Monochromatic Schur triples in randomly perturbed dense sets of integers. SIAM Journal on Discrete Mathematics, 33(4): 2175–2180, ???? 2019. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Allouche:1989:AIP

[AHS89] J.-P. Allouche, P. Hajnal, and J. O. Shallit. Analysis of an infinite product algorithm. SIAM Journal on Discrete Mathematics, 2(1): 1–15, February 1989. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ageev:2001:AAM

[AHS01] Alexander Ageev, Refael Hassin, and Maxim Sviridenko. A 0.5-approximation algorithm for MAX DICUT with given sizes of parts. SIAM Journal on Discrete Mathematics, 14(2):246–255, 2001. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/36813.

Akhtar:2008:ADE

[AJM08] Reza Akhtar, Tao Jiang, and Zevi Miller. Asymptotic determination of edgebandwidth of multidimensional grids and Hamming graphs. SIAM Journal on Discrete Mathematics, 22(2): 425–449, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Alon:2002:TC

[AK02] Noga Alon and Michael Krivelevich. Testing kcolorability. SIAM Journal on Discrete Mathematics, 15(2):211–227, 2002. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/35865.

Abraham:2010:VP

[AK10] David J. Abraham and Telikepalli Kavitha. Voting paths. SIAM Journal on Discrete Mathematics, 24(2): 520–537, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Aharoni:2014:WVR

[AK14] Ron Aharoni and Daniel Kotlar. A weak version of Rota’s bases conjecture for odd dimensions. SIAM Journal on Discrete Mathematics, 28(1): 385–393, ???? 2014. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Alon:2022:HTE

[AK22a] Yahav Alon and Michael Krivelevich. Hitting time of edge disjoint Hamilton cycles in random subgraph processes on dense base graphs. SIAM Journal on Discrete Mathematics, 36 (1):728–754, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/20M1375838.

Amanatidis:2022:RMS

[AK22b] Georgios Amanatidis and Pieter Kleer. Rapid mixing of the switch Markov chain for 2-class joint degree matrices. SIAM Journal on Discrete Mathematics, 36(1):118–146, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1352697.

Alon:2008:TTF

[AKKR08] Noga Alon, Tali Kaufman, Michael Krivelevich, and Dana Ron. Testing triangle-freeness in general graphs. SIAM Journal on Discrete Mathematics, 22(2): 786–819, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ajtai:1989:SAT

[AKKS89] M. Ajtai, D. Karabeg, J. Komlós, and E. Szemerédi. Sorting in average time o(log n). SIAM Journal on Discrete Mathematics, 2(3):285–292, August 1989. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Alon:2022:STC

[AKM22] Yahav Alon, Michael Krivelevich, and Peleg Michaeli. Spanning trees at the connectivity threshold. SIAM Journal on Discrete Mathematics, 36(3):1612–1626, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/20M1380521.

Ackerman:2020:CHD

[AKP20] Eyal Ackerman, Balázs Keszegh, and Dömötör Pálvölgyi. Coloring hypergraphs defined by stabbed pseudo-disks and ABAB-free hypergraphs. SIAM Journal on Discrete Mathematics, 34(4):2250– 2269, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Agrawal:2022:FPT

[AKP^{+}22] Akanksha Agrawal, Lawqueen Kanesh, Fahad Panolan, M. S. Ramanujan, and Saket Saurabh. A fixed-parameter tractable algorithm for elimination distance to bounded degree graphs. SIAM Journal on Discrete Mathematics, 36(2):911–921, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1396824.

Alon:2007:TTH

[AKS07] Noga Alon, Anja Krech, and Tibor Szabó. Turán’s Theorem in the hypercube. SIAM Journal on Discrete Mathematics, 21(1):66–72, ???? 2007. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Alon:2008:LNR

[AKS08] Noga Alon, Michael Krivelevich, and Benny Sudakov. Large nearly regular induced subgraphs. SIAM Journal on Discrete Mathematics, 22(4): 1325–1337, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Alstrup:2019:ALS

[AKTZ19] Stephen Alstrup, Haim Kaplan, Mikkel Thorup, and Uri Zwick. Adjacency labeling schemes and induceduniversal graphs. SIAM Journal on Discrete Mathematics, 33(1):116–137, ???? 2019. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Albertson:2005:PEB

[AKW05] Michael O. Albertson, Alexandr V.

Kostochka, and Douglas B. West. Precoloring extensions of Brooks’ Theorem. SIAM Journal on Discrete Mathematics, 18(3):542–553, 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/42594.

Aharoni:2017:RLM

[AKZ17] Ron Aharoni, Dani Kotlar, and Ran Ziv. Representation of large matchings in bipartite graphs. SIAM Journal on Discrete Mathematics, 31(3):

1726–1731, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Althofer:1995:CBF

[AL95] Ingo Althöfer and Imre Leader. Correlation of Boolean functions and pathology in recursion trees. SIAM Journal on Discrete Mathematics, 8(4):526–535, November 1995. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Alon:2007:GPD

[AL07] Noga Alon and Eyal Lubetzky. Graph powers, Delsarte, Hoffman, Ramsey, and Shannon. SIAM Journal on Discrete Mathematics, 21(2): 329–348, ???? 2007. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Archer:2017:APS

[AL17] Kassie Archer and Scott M. LaLonde. Allowed patterns of symmetric tent maps via commuter functions. SIAM Journal on Discrete Mathematics, 31(1):317–334, ???? 2017. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Axenovich:2024:CTC

[AL24] Maria Axenovich and Hanno Lefmann. Canonical theorems for colored integers with respect to some linear combinations. SIAM Journal on Discrete Mathematics, 38(1): 609–628, February 2024. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Aldous:1990:RWC

[Ald90] David J. Aldous. The random walk construction of uniform spanning trees and uniform labelled trees. SIAM Journal on Discrete Mathematics, 3(4):450–465, November 1990. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Aldous:2022:LCS

[Ald22] David J. Aldous. On the largest common subtree of random leaf-labeled binary trees. SIAM Journal on Discrete Mathematics, 36 (1):299–314, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1347504.

Alekseyev:2010:NTD

[Ale10] Max A. Alekseyev. On the number of two-dimensional threshold functions. SIAM Journal on Discrete Mathematics, 24(4):1617–1631, ???? 2010. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.

                                        siam.org/sidma/resource/
                                        1/sjdmec/v24/i4/p1617_s1.

Agrawal:2018:KCP

[ALM^{+}18] Akanksha Agrawal, Daniel Lokshtanov, Diptapriyo Majumdar, Amer E. Mouawad, and Saket Saurabh. Kernelization of cycle packing with relaxed disjointness constraints. SIAM Journal on Discrete Mathematics, 32(3): 1619–1643, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Alon:2006:RT

[Alo06] Noga Alon. Ranking tournaments. SIAM Journal on Discrete Mathematics, 20(1): 137–142, January 2006. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Alon:2013:MNC

[Alo13] Noga Alon. Minimizing the number of carries in addition. SIAM Journal on Discrete Mathematics, 27(1): 562–566, ???? 2013. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Alecu:2024:TPG

[ALQ^{+}24] Bogdan Alecu, Vadim V. Lozin, Daniel A. Quiroz, Roman Rabinovich, Igor Razgon, and Viktor Zamaraev. The treewidth and pathwidth of graph unions. SIAM Journal on Discrete Mathematics, 38(1):261–276, January 2024. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Albert:2011:PPP

[ALSY11] Chase Albert, Chi-Kwong Li, Gilbert Strang, and Gexin Yu. Permutations as product of parallel transpositions. SIAM Journal on Discrete Mathematics, 25(3): 1412–1417, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v25/i3/p1412_s1.

Altschul:1989:LPT

[Alt89] Stephen F. Altschul. Leaf pairs and tree dissections. SIAM Journal on Discrete Mathematics, 2(3):293–299, August 1989. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Alspach:1996:NZF

[ALZ96] Brian Alspach, Yi-Ping Liu, and Cun-Quan Zhang. Nowhere-zero 4-flows and Cayley graphs on solvable groups. SIAM Journal on Discrete Mathematics, 9(1): 151–154, February 1996. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ashley:1995:CES

[AM95] Jonathan Ashley and Brian Marcus. Canonical encoders for sliding block decoders. SIAM Journal on Discrete Mathematics, 8(4): 555–605, November 1995. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Arikati:1996:RDS

[AM96] Srinivasa R. Arikati and Anil Maheshwari. Realizing degree sequences in parallel. SIAM Journal on Discrete Mathematics, 9(2):317–338, May 1996. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Axenovich:2006:APM

[AM06a] Maria Axenovich and Ryan Martin. Avoiding patterns in matrices via a small number of changes. SIAM Journal on Discrete Mathematics, 20(1): 49–54, January 2006. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Axenovich:2006:SCN

[AM06b] Maria Axenovich and Ryan Martin. On the strong chromatic number of graphs. SIAM Journal on Discrete Mathematics, 20(3):741–747, January 2006. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Alderson:2007:COO

[AM07] T. L. Alderson and Keith E.

Mellinger. Constructions of optical orthogonal codes from finite geometry. SIAM Journal on Discrete Mathematics, 21(3):785–793, ???? 2007. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Alon:2011:SBP

[AM11] Noga Alon and Dániel Marx. Sparse balanced partitions and the complexity of subgraph problems. SIAM Journal on Discrete Mathematics, 25(2): 631–644, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v25/i2/p631_s1.

Alishahi:2021:TBG

[AM21] Meysam Alishahi and Frédéric Meunier. Topological bounds for graph representations over any field. SIAM Journal on Discrete Mathematics, 35(1): 91–104, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Arkus:2011:DFS

[AMB11] Natalie Arkus, Vinothan N.

Manoharan, and Michael P. Brenner. Deriving finite sphere packings. SIAM Journal on Discrete Mathematics, 25(4):1860–1901,

????  2011.  CODEN

SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v25/i4/p1860_s1.

Arvind:2018:EGS

[AMNV18] V. Arvind, Partha Mukhopadhyay, Prajakta Nimbhorkar, and Yadu Vasudev. Expanding generating sets for solvable permutation groups. SIAM Journal on Discrete Mathematics, 32(3):1721– 1740, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ashley:1993:SES

[AMPT93] Jonathan Ashley, Brian Marcus, Dominique Perrin, and Selim Tuncel. Surjective extensions of sliding-block codes. SIAM Journal on Discrete Mathematics, 6(4):582– 611, November 1993. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ajtai:2000:IAA

[AMW00] Miklos Ajtai, Nimrod Megiddo, and Orli Waarts. Improved algorithms and analysis for secretary problems and generalizations. SIAM Journal on Discrete Mathematics, 14(1):1–27, 2000. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29001.

Anastos:2018:CH

[Ana18] Michael Anastos. Connectivity of the k-out hypercube. SIAM Journal on Discrete Mathematics, 32(3): 2194–2216, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Alles:1991:EDD

[ANP91] Peter Alles, Jaroslav Nešetřil, and Svatopluk Poljak. Extendability, dimensions, and diagrams of cyclic orders. SIAM Journal on Discrete Mathematics, 4(4):453–471, November 1991. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Araujo:2014:WCT

[ANP14] Julio Araujo, Nicolas Nisse, and Stéphane Pérennes. Weighted coloring in trees. SIAM Journal on Discrete Mathematics, 28(4):2029– 2041, ???? 2014. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Alon:2016:MQD

[ANS16] Noga Alon, Humberto Naves, and Benny Sudakov. On the maximum quartet distance between phylogenetic trees. SIAM Journal on Discrete Mathematics, 30(2):

718–735, ???? 2016. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Adjiashvili:2015:PFT

[AOW15] David Adjiashvili, Timm Oertel, and Robert Weismantel. A polyhedral Frobenius theorem with applications to integer optimization. SIAM Journal on Discrete Mathematics, 29(3):1287–1302, ???? 2015. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Awerbuch:1992:RPC

[AP92] Baruch Awerbuch and David Peleg. Routing with polynomial communication-space trade-off. SIAM Journal on Discrete Mathematics, 5(2): 151–162, May 1992. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Amini:2014:SWP

[AP14] Hamed Amini and Yuval Peres. Shortest-weight paths in random regular graphs. SIAM Journal on Discrete Mathematics, 28(2):656–672, ???? 2014. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Abdi:2018:DMD

[AP18] Ahmad Abdi and Kanstantsin Pashkovich. Delta minors, delta free clutters, and entanglement. SIAM Journal on Discrete Mathematics, 32(3): 1750–1774, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Aliste-Prieto:2023:MGC

[APdMOZ23] José Aliste-Prieto, Anna de Mier, Rosa Orellana, and José Zamora. Marked graphs and the chromatic symmetric function. SIAM Journal on Discrete Mathematics, 37 (3):1881–1919, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M148046X.

Araujo:2022:LCC

[APS22] Pedro Araújo, Simón Piga, and Mathias Schacht. Localized codegree conditions for tight Hamilton cycles in 3uniform hypergraphs. SIAM Journal on Discrete Mathematics, 36(1):147–169, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL

https://epubs.siam.org/
doi/10.1137/21M1408531.

Alon:2024:IDT

[APS^{+}24] Noga Alon, Emil Powierski, Michael Savery, Alex Scott, and Elizabeth Wilmer. Invertibility of digraphs and tournaments. SIAM Journal on Discrete Mathematics, 38(1):327–347, January 2024. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Athanasiadis:2004:NPG

[AR04] Christos A. Athanasiadis and Victor Reiner. Noncrossing partitions for the group D_{n}. SIAM Journal on Discrete Mathematics, 18(2):397–417, 2004. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/43219.

Ahal:2008:CSP

[AR08] Shlomo Ahal and Yuri Rabinovich. On complexity of the subpattern problem. SIAM Journal on Discrete Mathematics, 22(2):629–649, ???? 2008. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Alon:2017:BTP

[AR17] Noga Alon and Guy Rutenberg. Broadcast transmission to prioritizing receivers. SIAM Journal on Discrete Mathematics, 31(4): 2517–2529, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Alonso:1995:MDC

[ARS95] Laurent Alonso, Edward M. Reingold, and René Schott.

Multidimensional divide-and-conquer maximin recurrences. SIAM Journal on Discrete Mathematics, 8(3): 428–447, August 1995. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Aracena:2017:NFP

[ARS17] Julio Aracena, Adrien Richard, and Lilian Salinas. Number of fixed points and disjoint cycles in monotone Boolean networks. SIAM Journal on Discrete Mathematics, 31(3): 1702–1725, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Arroyo:2021:EDC

[ARS21] Alan Arroyo, R. Bruce Richter, and Matthew Sunohara. Extending drawings of complete graphs into arrangements of pseudocircles. SIAM Journal on Discrete Mathematics, 35(2):1050–1076, ???? 2021. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Arman:2023:CM

[ARS23] Andrii Arman, Vojtech Rödl, and Marcelo Tadeu Sales. Colorful matchings. SIAM Journal on Discrete Mathematics, 37(2):925–950, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL

                                        https://epubs.siam.org/
                                        doi/10.1137/21M139997X.

Allman:2014:SDG

[ART14] Elizabeth S. Allman, John A. Rhodes, and Amelia Taylor. A semialgebraic description of the general Markov model on phylogenetic trees. SIAM Journal on Discrete Mathematics, 28(2):736–755, ???? 2014. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Aboulker:2012:GDC

[ARTV12] Pierre Aboulker, Marko Radovanović, Nicolas Trotignon, and Kristina Vusković. Graphs that do not contain a cycle with a node that has at least two neighbors on it. SIAM Journal on Discrete Mathematics, 26(4):1510–1531, ???? 2012. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Adams:1997:CNB

[AS97] Michael J. Adams and Bryan L. Shader. A construction for (t, m, s)-nets in base q. SIAM Journal on Discrete Mathematics, 10(3): 460–468, August 1997. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ashir:2002:FTE

[AS02] Yaagoub A. Ashir and Iain A. Stewart. Fault-tolerant embeddings of Hamiltonian circuits in k-ary n-cubes. SIAM Journal on Discrete Mathematics, 15(3):317–328, 2002. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31118.

Autebert:2003:GDP

[AS03] Jean-Michel Autebert and Sylviane R. Schwer. On generalized Delannoy paths. SIAM Journal on Discrete Mathematics, 16(2): 208–223, 2003. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/38740.

Apfelbaum:2005:RAT

[AS05] Roel Apfelbaum and Micha Sharir. Repeated angles in three and four dimensions. SIAM Journal on Discrete Mathematics, 19 (2):294–300, 2005. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/44394.

Abueida:2006:CEH

[AS06] Atif Abueida and R. Sritharan. Cycle extendability and Hamiltonian cycles in chordal graph classes. SIAM Journal on Discrete Mathematics, 20(3):669–681, January 2006. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Apfelbaum:2007:LCB

[AS07] Roel Apfelbaum and Micha Sharir. Large complete bipartite subgraphs in incidence graphs of points and hyperplanes. SIAM Journal on Discrete Mathematics, 21(3): 707–725, ???? 2007. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Aazami:2009:AAH

[AS09a] Ashkan Aazami and Kael Stilp. Approximation algorithms and hardness for domination with propagation. SIAM Journal on Discrete Mathematics, 23(3): 1382–1399, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Apollonio:2009:MFP

[AS09b] Nicola Apollonio and András Sebö. Minconvex factors of prescribed size in graphs. SIAM Journal on Discrete Mathematics, 23(3):1297– 1310, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Asgeirsson:2009:DCA

[ÁS09c] Eyjólfur Ingi Ásgeirsson and Cliff Stein. Divide-and-conquer approximation algorithm for vertex cover. SIAM Journal on Discrete Mathematics, 23(3):1261–1280, ???? 2009. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Aravind:2010:BEC

[AS10] N. R. Aravind and C. R.

Subramanian. Bounds on edge colorings with restrictions on the union of color classes. SIAM Journal on Discrete Mathematics, 24(3): 841–852, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Apollonio:2014:IAM

[AS14] Nicola Apollonio and Bruno Simeone. Improved approximation of maximum vertex coverage problem on bipartite graphs. SIAM Journal on Discrete Mathematics, 28(3): 1137–1151, ???? 2014. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Arizmendi:2016:LAC

[AS16] Octavio Arizmendi and Gelasio Salazar. Large area convex holes in random point sets. SIAM Journal on Discrete Mathematics, 30(3): 1866–1875, ???? 2016. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Alon:2019:IUH

[AS19] Noga Alon and Nadav Sherman. Induced universal hypergraphs. SIAM Journal on Discrete Mathematics, 33(2): 629–642, ???? 2019. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Akopyan:2018:CBP

[ASH18] Arseniy Akopyan and Erel Segal-Halevi. Counting blanks in polygonal arrangements. SIAM Journal on Discrete Mathematics, 32(3): 2242–2257, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Anderson:2010:FTN

[ASMF10] Brian D. O. Anderson, Iman Shames, Guoqiang Mao, and Bariş Fidan. Formal theory of noisy sensor network localization. SIAM Journal on Discrete Mathematics, 24(2): 684–698, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Alon:2009:CGD

[ASS09] Noga Alon, Asaf Shapira, and Uri Stav. Can a graph have distinct regular partitions? SIAM Journal on Discrete Mathematics, 23(1): 278–287, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Aravind:2017:DRH

[ASS17] N. R. Aravind, R. B.

Sandeep, and Naveen Sivadasan. Dichotomy results on the hardness of H-free edge modification problems. SIAM Journal on Discrete Mathematics, 31(1):542–561, ???? 2017. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Axenovich:2021:BIN

[ASSW21] Maria Axenovich, Jean-Sébastien Sereni, Richard Snyder, and Lea Weber. Bipartite independence number in graphs with bounded maximum degree. SIAM Journal on Discrete Mathematics, 35(2):1136–1148, ???? 2021. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Alon:1994:PS

[AST94] Noga Alon, Paul Seymour, and Robin Thomas. Planar separators. SIAM Journal on Discrete Mathematics, 7 (2):184–193, May 1994. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Alon:2014:CQT

[ASY14] Noga Alon, Sagi Snir, and Raphael Yuster. On the compatibility of quartet trees. SIAM Journal on Discrete Mathematics, 28(3):1493– 1507, ???? 2014. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Alon:2002:CWC

[ASZ02] Noga Alon, Benny Sudakov, and Uri Zwick. Constructing worst case instances for semidefinite programming based approximation algorithms. SIAM Journal on Discrete Mathematics, 15(1):58–72, 2002. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/37971.

Aigner:1990:IAT

[AT90] M. Aigner and E. Triesch. Irregular assignments of trees and forests. SIAM Journal on Discrete Mathematics, 3 (4):439–449, November 1990. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Au:2016:CAP

[AT16] Yu Hin Au and Levent Tunçel. A comprehensive analysis of polyhedral liftand-project methods. SIAM Journal on Discrete Mathematics, 30(1):411–451, ???? 2016. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Athanasiadis:2014:ESL

[Ath14] Christos A. Athanasiadis. Edgewise subdivisions, local h-polynomials, and excedances in the wreath product Z_{r} ≀ S_{\}. SIAM Journal on Discrete Mathematics, 28(3): 1479–1492, ???? 2014. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Al-Thukair:1991:MPT

[ATPRU91] Fawzi Al-Thukair, Andrzej Pelc, Ivan Rival, and Jorge Urrutia. Motion planning, two-directional point representations, and ordered sets. SIAM Journal on Discrete Mathematics, 4(2): 151–163, May 1991. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Assaf:1991:FTS

[AU91] Shay Assaf and Eli Upfal.

Fault tolerant sorting networks. SIAM Journal on Discrete Mathematics, 4(4):472– 480, November 1991. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Averkov:2012:SLS

[Ave12]  Gennadiy Averkov.  On

the size of lattice simplices with a single interior lattice point. SIAM Journal on Discrete Mathematics, 26(2): 515–526, ???? 2012. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Averkov:2013:MFS

[Ave13] Gennadiy Averkov. On maximal S-free sets and the Helly number for the family of Sconvex sets. SIAM Journal on Discrete Mathematics, 27(3): 1610–1624, ???? 2013. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ageev:2004:ICA

[AYZ04] Alexander Ageev, Yinyu Ye, and Jiawei Zhang. Improved combinatorial approximation algorithms for the k-level facility location problem. SIAM Journal on Discrete Mathematics, 18(1):207–217, 2004. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/41721.

Akhmejanova:2022:EFL

[AZ22] Margarita Akhmejanova and Maksim Zhukovskii. EMSO(FO^{2}) 0–1 law fails for all dense random graphs. SIAM Journal on Discrete Mathematics, 36 (3):1788–1799, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1429655.

Babai:1992:BRI

[Bab92] László Babai. Bounded round interactive proofs in finite groups. SIAM Journal on Discrete Mathematics, 5(1):

88–111, February 1992. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bach:2009:IAF

[Bac09] Eric Bach. Improved asymptotic formulas for counting correlation immune Boolean functions. SIAM Journal on Discrete Mathematics, 23(3): 1525–1538, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ben-Ameur:2003:IRR

[BAG03] Walid Ben-Ameur and Eric Gourdin. Internet routing and related topology issues. SIAM Journal on Discrete Mathematics, 17 (1):18–49, 2003. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/37742.

Ben-Ameur:2010:DSN

[BAH10] Walid Ben-Ameur and Makhlouf Hadji. Designing Steiner networks with unicyclic connected components: An easy problem. SIAM Journal on Discrete Mathematics, 24(4): 1541–1557, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v24/i4/p1541_s1.

Bahmanian:2023:SLR

[Bah23] Amin Bahmanian. Symmetric layer-rainbow colorations of cubes. SIAM Journal on Discrete Mathematics, 37(4): 2617–2625, November 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Baldi:1988:NNA

[Bal88] Pierre Baldi. Neural networks, acyclic orientations of the hypercube, and sets of orthogonal vectors. SIAM Journal on Discrete Mathematics, 1(1):1–13, February 1988. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Balas:1989:AAP

[Bal89] Egon Balas. The asymmetric assignment problem and some new facets of the traveling salesman polytope on a directed graph. SIAM Journal on Discrete Mathematics, 2 (4):425–451, November 1989. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Balbuena:2008:IMP

[Bal08] C. Balbuena. Incidence matrices of projective planes and of some regular bipartite graphs of girth 6 with few vertices. SIAM Journal on Discrete Mathematics, 22(4): 1351–1363, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ben-Ari:2016:PAG

[BAM16] Iddo Ben-Ari and Steven J.

Miller. A probabilistic approach to generalized Zeckendorf decompositions. SIAM Journal on Discrete Mathematics, 30(2):1302–1332, ???? 2016. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bandelt:1990:RTM

[Ban90] Hans-Jürgen Bandelt. Recognition of tree metrics. SIAM Journal on Discrete Mathematics, 3(1):1–6, February 1990. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Barnes:1995:IPM

[Bar95] Earl R. Barnes. An inequality for probability moments with applications. SIAM Journal on Discrete Mathematics, 8(3):347–358, August 1995. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Barequet:2001:LBH

[Bar01] Gill Barequet. A lower bound for Heilbronn’s triangle problem in d dimensions. SIAM Journal on Discrete Mathematics, 14(2):230–236, 2001. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/36585.

Barg:2002:SPR

[Bar02] Alexander Barg. On some polynomials related to weight enumerators of linear codes. SIAM Journal on Discrete Mathematics, 15(2): 155–164, 2002. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/36414.

Barahona:2004:FPJ

[Bar04] Francisco Barahona. Fractional packing of T -joins. SIAM Journal on Discrete Mathematics, 17(4): 661–669, 2004. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/40961.

Bollobas:1990:PSH

[BB90] Béla Bollobás and Graham Brightwell. Parallel selection with high probability. SIAM Journal on Discrete Mathematics, 3(1):21–31, February 1990. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bollobas:1997:SRG

[BB97] Béla Bollobás and Graham Brightwell. The structure of random graph orders. SIAM Journal on Discrete Mathematics, 10(2):318–335, May 1997. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bohossian:2003:ATC

[BB03] Vasken Bohossian and Jehoshua Bruck. Algebraic techniques for constructing minimal weight threshold functions. SIAM Journal on Discrete Mathematics, 16(1):114–126, 2003. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/32604.

Baiou:2009:ISF

[BB09] Mourad Baïou and Francisco Barahona. On the integrality of some facility location polytopes. SIAM Journal on Discrete Mathematics, 23(2): 665–679, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ball:2013:BMW

[BB13] Simeon Ball and Aart Blokhuis.

A bound for the maximum weight of a linear code. SIAM Journal on Discrete Mathematics, 27(1):575–583, ???? 2013. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

BaIou:2016:MWI

[BB16] Mourad BaÏou and Francisco Barahona. Maximum weighted induced bipartite subgraphs and acyclic subgraphs of planar cubic graphs. SIAM Journal on Discrete Mathematics, 30(2): 1290–1301, ???? 2016. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bajo:2023:BPR

[BB23a] Esme Bajo and Matthias Beck. Boundary H^{∗}- polynomials of rational polytopes. SIAM Journal on Discrete Mathematics, 37 (3):1952–1969, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1508911.

Bartoli:2023:SSB

[BB23b] Daniele Bartoli and Martino Borello. Small strong blocking sets by concatenation. SIAM Journal on Discrete Mathematics, 37(1): 65–82, ???? 2023. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:/

                                        /epubs.siam.org/doi/10.
                                        1137/21M145032X.

Balogh:2012:TDS

[BBBZ12] József Balogh, Tom Bohman, Béla Bollobás, and Yi Zhao.

Turán densities of some hypergraphs related to K_{k+1}^{k}. SIAM Journal on Discrete Mathematics, 26(4):1609– 1617, ???? 2012. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Baiou:2011:MPI

[BBC11] Mourad Baïou, Francisco Barahona, and Jose Correa. On the p-median polytope and the intersection property: Polyhedra and algorithms. SIAM Journal on Discrete Mathematics, 25(1): 1–20, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v25/i1/p1_s1.

Babka:2019:OLL

[BBC^{+}19a] Martin Babka, Jan Bulánek, Vladim’r Cunát, Michal Koucký, and Michael E. Saks. On online labeling with large label set. SIAM Journal on Discrete Mathematics, 33(3): 1175–1193, ???? 2019. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Biedl:2019:RLS

[BBC^{+}19b] Therese Biedl, Ahmad Biniaz, Robert Cummings, Anna Lubiw, Florin Manea, Dirk Nowotka, and Jeffrey Shallit. Rollercoasters: Long sequences without short runs.

SIAM Journal on Discrete Mathematics, 33(2):845–861, ???? 2019. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bishnoi:2023:MDM

[BBC^{+}23] Anurag Bishnoi, Simona Boyadzhiyska, Dennis Clemens, Pranshu Gupta, Thomas Lesgourgues, and Anita Liebenau. On the minimum degree of minimal Ramsey graphs for cliques versus cycles. SIAM Journal on Discrete Mathematics, 37(1):25–50, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL

                                        https://epubs.siam.org/
                                        doi/10.1137/21M1444953.

Bayati:2011:BPW

[BBCZ11] Mohsen Bayati, Christian Borgs, Jennifer Chayes, and Riccardo Zecchina. Belief propagation for weighted b-matchings on arbitrary graphs and its relation to linear programs with integer solutions. SIAM Journal on Discrete Mathematics, 25(2): 989–1011, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sidma/resource/
                                        1/sjdmec/v25/i2/p989_s1.

Beimel:2000:CFS

[BBDK00] Amos Beimel, Mike Burmester, Yvo Desmedt, and Eyal Kushilevitz. Computing functions of a shared secret. SIAM Journal on Discrete Mathematics, 13(3):324–345, 2000. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28881.

Bafna:1999:AAU

[BBF99] Vineet Bafna, Piotr Berman, and Toshihiro Fujito. A 2approximation algorithm for the undirected feedback vertex set problem. SIAM Journal on Discrete Mathematics, 12(3):289–297, August 1999. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/30512.

Bartier:2023:RPG

[BBF^{+}23] Valentin Bartier, Nicolas Bousquet, Carl Feghali, Marc Heinrich, Benjamin Moore, and Théo Pierron. Recoloring planar graphs of girth at least five. SIAM Journal on Discrete Mathematics, 37(1):332–350, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1463598.

Beveridge:2009:MRA

[BBFP09] Andrew Beveridge, Tom Bohman, Alan Frieze, and Oleg Pikhurko. Memoryless rules for Achlioptas processes. SIAM Journal on Discrete Mathematics, 23(2): 993–1008, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Balister:2008:SCD

[BBG08] Paul Balister, Béla Bollobás, and Stefanie Gerke. Sequences with changing dependencies. SIAM Journal on Discrete Mathematics, 22(3): 1149–1154, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Barakat:2021:GRS

[BBJ^{+}21] Mohamed Barakat, Reimer Behrends, Christopher Jefferson, Lukas Kühne, and Martin Leuner. On the generation of rank 3 simple matroids with an application to Terao’s freeness conjecture. SIAM Journal on Discrete Mathematics, 35(2):1201–1223, ???? 2021. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Beck:2016:GFT

[BBK^{+}16] Matthias Beck, Benjamin Braun, Matthias Köppe, Carla D. Savage, and Zafeirakis Zafeirakopoulos. Generating functions and triangulations for lecture hall cones. SIAM Journal on Discrete Mathematics, 30(3):1470–1479, ???? 2016. CODEN SJDMEC.

ISSN 0895-4801 (print), 1095-7146 (electronic).

Berczi:2023:DAD

[BBKS23] Kristóf Bérczi, Erika R.

Bérczi-Kovács, and Evelin Szögi. A dual approach for dynamic pricing in multidemand markets. SIAM Journal on Discrete Mathematics, 37(3):1771–1787, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1535474.

Balister:2013:RDR

[BBLM13] Paul Balister, Béla Bollobás, Jenö Lehel, and Michal Morayne. Repeated degrees in random uniform hypergraphs. SIAM Journal on Discrete Mathematics, 27(1): 145–154, ???? 2013. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Balister:2016:RHI

[BBLM16] Paul Balister, Béla Bollobás, Jenö Lehel, and Michal Morayne. Random hypergraph irregularity. SIAM Journal on Discrete Mathematics, 30(1):465–473, ???? 2016. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bienstock:1990:SMW

[BBM90] Daniel Bienstock, Ernest F.

Brickell, and Clyde L.

Monma. On the structure of minimum-weight k-connected spanning networks. SIAM Journal on Discrete Mathematics, 3(3):320–329, August 1990. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bondy:2009:LPC

[BBM09] J. A. Bondy, X. Buchwalder, and F. Mercier. Lexicographic products and a conjecture of Hahn and Jackson. SIAM Journal on Discrete Mathematics, 23(2): 882–887, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Berczi:2024:HHF

[BBM24] Kristóf Bérczi, Endre Boros, and Kazuhisa Makino. Hypergraph horn functions. SIAM Journal on Discrete Mathematics, 38(2):1417– 1437, May 2024. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bracken:2009:FSB

[BBMM09] Carl Bracken, Eimear Byrne, Nadya Markin, and Gary McGuire. Fourier spectra of binomial APN functions. SIAM Journal on Discrete Mathematics, 23(2): 596–608, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Berge:2023:PCC

[BBRT23] Pierre Bergé, Wassim Bouaziz, Arpad Rimmel, and Joanna Tomasik. On the parameterized complexity of counting small-sized minimum (S, T)- cuts. SIAM Journal on Discrete Mathematics, 37 (2):964–996, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1398203.

Birkendorf:2000:LDF

[BBS00] Andreas Birkendorf, Andreas Böker, and Hans Ulrich Simon. Learning deterministic finite automata from smallest counterexamples. SIAM Journal on Discrete Mathematics, 13(4):465–491, 2000. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/34094.

Balvociute:2017:WCS

[BBS17] Monika Balvociute, David Bryant, and Andreas Spillner. When can splits be drawn in the plane? SIAM Journal on Discrete Mathematics, 31 (2):839–856, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bonamy:2016:EHC

[BBT16] Marthe Bonamy, Nicolas Bousquet, and Stéphan Thomassé.

The Erdős–Hajnal conjecture for long holes and antiholes. SIAM Journal on Discrete Mathematics, 30(2): 1159–1164, ???? 2016. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bollobas:1988:DCP

[BC88a] B. Bollobás and F. R. K. Chung. The diameter of a cycle plus a random matching. SIAM Journal on Discrete Mathematics, 1(3):328– 333, August 1988. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Brown:1988:SSP

[BC88b] Jason I. Brown and Charles J. Colbourn. A set system polynomial with colouring and reliability applications. SIAM Journal on Discrete Mathematics, 1(2):151–157, May 1988. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Brown:1992:RRP

[BC92] Jason I. Brown and Charles J. Colbourn. Roots of the reliability polynomial. SIAM Journal on Discrete Mathematics, 5(4):571–585, November 1992. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bender:1994:NDR

[BC94] Edward A. Bender and E. Rodney Canfield. The number of degree-restricted rooted maps on the sphere. SIAM Journal on Discrete Mathematics, 7(1):9– 15, February 1994. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bouchet:1995:DMJ

[BC95] André Bouchet and William H.

Cunningham. Delta-matroids, jump systems, and bisubmodular polyhedra. SIAM Journal on Discrete Mathematics, 8(1):17–32, February 1995. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bandelt:2002:GCM

[BC02] Hans-Jürgen Bandelt and Victor Chepoi. Graphs with connected medians. SIAM Journal on Discrete Mathematics, 15(2):268–282, 2002. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/36360.

Bandelt:2003:HG

[BC03] Hans-Jürgen Bandelt and Victor Chepoi. 1-hyperbolic graphs. SIAM Journal on Discrete Mathematics, 16 (2):323–334, 2003. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/38090.

Bekmetjev:2009:PAD

[BC09a] Airat Bekmetjev and Charles A. Cusack. Pebbling algorithms in diameter two graphs. SIAM Journal on Discrete Mathematics, 23(2): 634–646, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Brennan:2009:TGS

[BC09b] Joseph Brennan and Guantao Chen. Toric geometry of series-parallel graphs. SIAM Journal on Discrete Mathematics, 23(2):754–764, ???? 2009. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Borenstein:2011:CGB

[BC11] Evan Borenstein and Ernie Croot. On a certain generalization of the Balog– Szemerédi–Gowers Theorem. SIAM Journal on Discrete Mathematics, 25(2):685–694, ???? 2011. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sidma/resource/
                                        1/sjdmec/v25/i2/p685_s1.

Byrne:2017:BOR

[BC17] Eimear Byrne and Marco Calderini. Bounding the optimal rate of the ICSI and ICCSI problem. SIAM Journal on Discrete Mathematics, 31(2):1403–1427, ???? 2017. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Brignall:2023:FMH

[BC23a] R. Brignall and D. Cocks.

A framework for minimal hereditary classes of graphs of unbounded clique-width. SIAM Journal on Discrete Mathematics, 37(4):2558– 2584, October 2023. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Byrne:2023:TCT

[BC23b] Eimear Byrne and Giuseppe Cotardo. Tensor codes and their invariants. SIAM Journal on Discrete Mathematics, 37(3):1988–2015, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1487734.

Bermond:2005:TGU

[BCC^{+}05] Jean-Claude Bermond, Charles J.

Colbourn, David Coudert, Gennian Ge, Alan C. H. Ling, and Xavier Muñoz. Traffic grooming in unidirectional wavelength-division multiplexed rings with grooming ratio C = 6. SIAM Journal on Discrete Mathematics, 19(2):523–542, 2005.

CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/44431.

Basavaraju:2011:AEC

[BCC^{+}11] Manu Basavaraju, L. Sunil Chandran, Nathann Cohen, Frédéric Havet, and Tobias Müller. Acyclic edgecoloring of planar graphs. SIAM Journal on Discrete Mathematics, 25(2):463–478, ???? 2011. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sidma/resource/
                                        1/sjdmec/v25/i2/p463_s1.

Brettell:2019:GS

[BCC^{+}19] Nick Brettell, Rutger Campbell, Deborah Chun, Kevin Grace, and Geoff Whittle. On a generalization of spikes. SIAM Journal on Discrete Mathematics, 33(1): 358–372, ???? 2019. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Boyd:2022:AAM

[BCC^{+}22] Sylvia Boyd, Joseph Cheriyan, Robert Cummings, Logan Grout, Sharat Ibrahimpur, Zoltán Szigeti, and Lu Wang. A \frac{4}{3}-approximation algorithm for the minimum 2edge connected multisubgraph problem in the halfintegral case. SIAM Journal on Discrete Mathematics, 36(3):1730–1747, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1372822.

Basu:2010:MII

[BCCZ10] Amitabh Basu, Michele Conforti, Gérard Cornuéjols, and Giacomo Zambelli. Minimal inequalities for an infinite relaxation of integer programs. SIAM Journal on Discrete Mathematics, 24(1): 158–168, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bermond:2011:ACU

[BCCZ11] J.-C. Bermond, Y. M. Chee, N. Cohen, and X. Zhang. The α-arboricity of complete uniform hypergraphs. SIAM Journal on Discrete Mathematics, 25(2):600–610, ???? 2011. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v25/i2/p600_s1.

Brandstadt:1997:CDC

[BCD97] Andreas Brandstädt, Victor D. Chepoi, and Feodor F. Dragan. Clique r-domination and clique r-packing problems on dually chordal graphs. SIAM Journal on Discrete Mathematics, 10(1):

109–127, February 1997. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Barbosa:2012:CNC

[BCD^{+}12] Rommel M. Barbosa, Erika M. M. Coelho, Mitre C. Dourado, Dieter Rautenbach, and Jayme L. Szwarcfiter. On the Carathéodory number for the convexity of paths of order three. SIAM Journal on Discrete Mathematics, 26(3): 929–939, ???? 2012. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bennett:2021:CRG

[BCD21] Patrick Bennett, Ryan Cushman, and Andrzej Dudek. Closing the random graph gap in Tuza’s conjecture through the online triangle packing process. SIAM Journal on Discrete Mathematics, 35(3):2145–2169, ???? 2021. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bezakova:2019:FDF

[BCDF19] Ivona Bezáková, Radu Curticapean, Holger Dell, and Fedor V. Fomin. Finding detours is fixed-parameter tractable. SIAM Journal on Discrete Mathematics, 33(4): 2326–2345, ???? 2019. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bergeron:2008:CCI

[BCdMR08] Anne Bergeron, Cedric Chauve, Fabien de Montgolfier, and Mathieu Raffinot. Computing common intervals of K permutations, with applications to modular decomposition of graphs. SIAM Journal on Discrete Mathematics, 22(3): 1022–1039, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Boros:2000:BNF

[BCE^{+}00] Endre Boros, Yves Crama, Oya Ekin, Peter L. Hammer, Toshihide Ibaraki, and Alexander Kogan. Boolean normal forms, shellability, and reliability computations. SIAM Journal on Discrete Mathematics, 13 (2):212–226, 2000. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/32180.

Barg:2001:HAI

[BCE^{+}01] Alexander Barg, Gérard Cohen, Sylvia Encheva, Gregory Kabatiansky, and Gilles Zémor. A hypergraph approach to the identifying parent property: The case of multiple parents. SIAM Journal on Discrete Mathematics, 14(3):423–431, 2001. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/37684.

Bollobas:2005:SDP

[BCE05] Béla Bollobás, Don Coppersmith, and Michael Elkin. Sparse distance preservers and additive spanners. SIAM Journal on Discrete Mathematics, 19(4):1029–1055, January 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bandelt:2010:CGF

[BCE10] Hans-Jürgen Bandelt, Victor Chepoi, and David Eppstein. Combinatorics and geometry of finite and infinite squaregraphs. SIAM Journal on Discrete Mathematics, 24(4): 1399–1440, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Brightwell:2010:PCF

[BCF^{+}10] Graham Brightwell, Gérard Cohen, Emanuela Fachini, Marianne Fairthorne, János Körner, Gábor Simonyi, and Ágnes Tóth. Permutation capacities of families of oriented infinite paths. SIAM Journal on Discrete Mathematics, 24 (2):441–456, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bodirsky:2012:CM

[BCF12] Manuel Bodirsky, Hubie Chen, and Tomás Feder. On the complexity of MM-SNP. SIAM Journal on Discrete Mathematics, 26(1): 404–414, ???? 2012. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bustamante:2020:PEC

[BCF^{+}20] Sebastián Bustamante, Jan Corsten, Nóra Frankl, Alexey Pokrovskiy, and Jozef Skokan. Partitioning edge-colored hypergraphs into few monochromatic tight cycles. SIAM Journal on Discrete Mathematics, 34(2):1460–1471, ???? 2020. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bonin:2023:NMI

[BCF23] Joseph E. Bonin, Carolyn Chun, and Tara Fife. The natural matroid of an integer polymatroid. SIAM Journal on Discrete Mathematics, 37(3):1751–1770, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1521122.

Bermond:2010:DCW

[BCG^{+}10] Jean-Claude Bermond, Charles J.

Colbourn, Lucia Gionfriddo, Gaetano Quattrocchi, and Ignasi Sau. Drop cost and wavelength optimal two-period grooming with ratio 4. SIAM Journal on Discrete Mathematics, 24(2):400–419, ???? 2010. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Boyadzhiyska:2022:MRG

[BCG22] Simona Boyadzhiyska, Dennis Clemens, and Pranshu Gupta. Minimal Ramsey graphs with many vertices of small degree. SIAM Journal on Discrete Mathematics, 36(3):1503–1528, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1393273.

Balko:2024:BCO

[BCG^{+}24] Martin Balko, Steven Chaplick, Robert Ganian, Siddharth Gupta, Michael Hoffmann, Pavel Valtr, and Alexander Wolff. Bounding and computing obstacle numbers of graphs. SIAM Journal on Discrete Mathematics, 38(2):1537–1565, May 2024. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Boyadzhiyska:2024:REA

[BCGR24] Simona Boyadzhiyska, Dennis Clemens, Pranshu Gupta, and Jonathan Rollin. Ramsey equivalence for asymmetric pairs of graphs. SIAM Journal on Discrete Mathematics, 38(1):55–74, January 2024. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Boros:1992:CCO

[BCH92] E. Boros, Y. Crama, and P. L. Hammer. Chvatal cuts and odd cycle inequalities in quadratic 0-1 optimization. SIAM Journal on Discrete Mathematics, 5(2): 163–177, May 1992. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bretscher:2008:SLT

[BCHP08] Anna Bretscher, Derek Corneil, Michel Habib, and Christophe Paul. A simple linear time LexBFS cograph recognition algorithm. SIAM Journal on Discrete Mathematics, 22(4): 1277–1296, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bryant:2017:STS

[BCHW17] Darryn Bryant, Charles J.

Colbourn, Daniel Horsley, and Ian M. Wanless. Steiner triple systems with high chromatic index. SIAM Journal on Discrete Mathematics, 31(4):2603–2611, ???? 2017. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Barros:2021:ORT

[BCKN21] Gabriel Ferreira Barros, Bruno Pasqualotto Cavalar, Yoshiharu Kohayakawa, and Tássio Naia. Orientation Ramsey thresholds for cycles and cliques. SIAM Journal on Discrete Mathematics, 35(4): 2844–2857, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Burton:2019:PCF

[BCKP19] Benjamin Burton, Sergio Cabello, Stefan Kratsch, and William Pettersson. The parameterized complexity of finding a 2-sphere in a simplicial complex. SIAM Journal on Discrete Mathematics, 33(4):2092–2110, ???? 2019. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bourne:2018:ORI

[BCL^{+}18] D. P. Bourne, D. Cushing, S. Liu, F. Münch, and N. Peyerimhoff. Ollivier– Ricci idleness functions of graphs. SIAM Journal on Discrete Mathematics, 32(2): 1408–1424, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Balogh:2023:STF

[BCL^{+}23] József Balogh, Felix Christian Clemen, Bernard Lidický, Sergey Norin, and Jan Volec. The spectrum of trianglefree graphs. SIAM Journal on Discrete Mathematics, 37 (2):1173–1179, ???? 2023.

CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M150767X.

Bhatt:1989:UGB

[BCLR89] Sandeep N. Bhatt, F. R. K.

Chung, F. T. Leighton, and Arnold L. Rosenberg. Universal graphs for boundeddegree trees and planar graphs. SIAM Journal on Discrete Mathematics, 2(2): 145–155, May 1989. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bhatt:1995:SEC

[BCLR95] Sandeep N. Bhatt, Fan R. K. Chung, Frank Thomson Leighton, and Arnold L. Rosenberg. Salvage-embeddings of complete trees. SIAM Journal on Discrete Mathematics, 8(4):617–637, November 1995. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bryant:2012:NLC

[BCM^{+}12] Darryn Bryant, Nicholas J.

Cavenagh, Barbara Maenhaut, Kyle Pula, and Ian M. Wanless. Nonextendible Latin cuboids. SIAM Journal on Discrete Mathematics, 26 (1):239–249, ???? 2012. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Broersma:2008:NAL

[BCP08] Hajo J. Broersma, Agostino Capponi, and Daniël Paulusma. A new algorithm for online coloring bipartite graphs. SIAM Journal on Discrete Mathematics, 22(1):72–91, ???? 2008. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bourgain:2009:DSP

[BCPP09] Jean Bourgain, Todd Cochrane, Jennifer Paulhus, and Christopher Pinner. Decimations of ℓ-sequences and permutations of even residues mod p. SIAM Journal on Discrete Mathematics, 23(2): 842–857, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Banks:2004:NTD

[BCS04] William D. Banks, Alessandro Conflitti, and Igor E. Shparlinski. Number theoretic designs for directed regular graphs of small diameter. SIAM Journal on Discrete Mathematics, 17 (3):377–383, 2004. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/39667.

Bagchi:2007:AFT

[BCSK07] Amitabha Bagchi, Amitabh Chaudhary, Christian Scheideler, and Petr Kolman. Algorithms for fault-tolerant routing in circuit-switched networks. SIAM Journal on Discrete Mathematics, 21(1): 141–157, ???? 2007. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Brown:1996:CMR

[BCW96] Jason I. Brown, Charles J.

Colbourn, and David G. Wagner. Cohen–Macaulay rings in network reliability. SIAM Journal on Discrete Mathematics, 9(3):377–392, August 1996. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Batten:2001:BST

[BD01] Lynn M. Batten and Jeremy M.

Dover. Blocking semiovals of type (1, m + 1, n + 1). SIAM Journal on Discrete Mathematics, 14(4):446–457, 2001. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/33800.

Barvinok:2020:TDS

[BD20a] Alexander Barvinok and Anthony Della Pella. Testing for dense subsets in a graph via the partition function. SIAM Journal on Discrete Mathematics, 34(1):308–327, ???? 2020. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Birx:2020:TAS

[BD20b] Alexander Birx and Yann Disser. Tight analysis of the smartstart algorithm for online dial-a-ride on the line. SIAM Journal on Discrete Mathematics, 34(2): 1409–1443, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Brakensiek:2024:RFC

[BD24] Joshua Brakensiek and Sami Davies. Robust factorizations and colorings of tensor graphs. SIAM Journal on Discrete Mathematics, 38(1): 883–916, February 2024. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Brandstadt:1998:DCG

[BDCV98] Andreas Brandstädt, Feodor Dragan, Victor Chepoi, and Vitaly Voloshin. Dually chordal graphs. SIAM Journal on Discrete Mathematics, 11(3):437–455, August 1998. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bernstein:2019:DPG

[BDD^{+}19] Aaron Bernstein, Karl Däubel, Yann Disser, Max Klimm, Torsten Mütze, and Frieder Smolny. Distance-preserving graph contractions. SIAM Journal on Discrete Mathematics, 33(3):1607–1636, ????

2019. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bousquet:2024:SCP

[BDDP24] Nicolas Bousquet, Quentin Deschamps, Lucas De Meyer, and Théo Pierron. Square coloring planar graphs with automatic discharging. SIAM Journal on Discrete Mathematics, 38(1):504–528, January 2024. CODEN SJD-MEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Blundo:2003:COT

[BDDS03] C. Blundo, P. D’Arco, A. De Santis, and D. R. Stinson. Contrast optimal threshold visual cryptography schemes. SIAM Journal on Discrete Mathematics, 16(2):224–261, 2003. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/33668.

Bose:2006:SRG

[BDEK06] Prosenjit Bose, Luc Devroye, William Evans, and David Kirkpatrick. On the spanning ratio of Gabriel graphs and beta-skeletons. SIAM Journal on Discrete Mathematics, 20(2):412–427, January 2006. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Beaudou:2018:BOG

[BDF^{+}18] Laurent Beaudou, Peter Dankelmann, Florent Foucaud, Michael A. Henning, Arnaud Mary, and Aline Parreau. Bounding the order of a graph using its diameter and metric dimension: a study through tree decompositions and VC dimension. SIAM Journal on Discrete Mathematics, 32(2):902–918, ???? 2018. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bohman:2010:FG

[BDFP10] Tom Bohman, Andrzej Dudek, Alan Frieze, and Oleg Pikhurko. Flips in graphs. SIAM Journal on Discrete Mathematics, 24(3):1046– 1055, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bishnu:2017:UPS

[BDG^{+}17] Arijit Bishnu, Sameer Desai, Arijit Ghosh, Mayank Goswami, and Subhabrata Paul. Uniformity of point samples in metric spaces using gap ratio. SIAM Journal on Discrete Mathematics, 31(3):2138–2171, ???? 2017. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bougeret:2022:CRA

[BDGP22] Marin Bougeret, György Dósa, Noam Goldberg, and Michael Poss. Constant-ratio approximation for robust bin packing with budgeted uncertainty. SIAM Journal on Discrete Mathematics, 36 (4):2534–2552, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1457199.

Bodlaender:1998:RG

[BDJ^{+}98] Hans L. Bodlaender, Jitender S. Deogun, Klaus Jansen, Ton Kloks, Dieter Kratsch, Haiko Müller, and Zsolt Tuza. Rankings of graphs. SIAM Journal on Discrete Mathematics, 11(1):168–181, February 1998. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Barat:2015:EPP

[BDJ^{+}15] János Barát, Vida Dujmović, Gwenaël Joret, Michael S. Payne, Ludmila Scharf, Daria Schymura, Pavel Valtr, and David R. Wood. Empty pentagons in point sets with collinearities. SIAM Journal on Discrete Mathematics, 29 (1):198–209, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Blanche:2020:CWG

[BDJ^{+}20] Alexandre Blanché, Konrad K. Dabrowski, Matthew Johnson, Vadim V. Lozin, Daniël Paulusma, and Viktor Zamaraev. Clique-width for graph classes closed under complementation. SIAM Journal on Discrete Mathematics, 34(2):1107–1147, ???? 2020. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Beisegel:2021:RPG

[BDK^{+}21] Jesse Beisegel, Carolin Denkert, Ekkehard Köhler, Matjaz Krnc, Nevena Pivac, Robert Scheffler, and Martin Strehler. The recognition problem of graph search trees. SIAM Journal on Discrete Mathematics, 35(2):1418–1446, ???? 2021. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bulteau:2024:AFL

[BDK^{+}24] Laurent Bulteau, Konrad K. Dabrowski, Noleen Köhler, Sebastian Ordyniak, and Daniël Paulusma. An algorithmic framework for locally constrained homomorphisms. SIAM Journal on Discrete Mathematics, 38(2): 1315–1350, April 2024. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Blanchard:2021:LMP

[BDL21] Moïse Blanchard, Jesús A. De Loera, and Quentin Louveaux. On the length of monotone paths in polyhedra. SIAM Journal on Discrete Mathematics, 35(3):

1746–1768, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Burkard:2001:OCP

[BDLR01] Rainer E. Burkard, Helidon Dollani, Yixun Lin, and Günter Rote. The obnoxious center problem on a tree. SIAM Journal on Discrete Mathematics, 14(4):498–509, 2001. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/34096.

Berman:2002:ESB

[BDM02] Piotr Berman, Bhaskar Das-Gupta, and S. Muthukrishnan. Exact size of binary space partitionings and improved rectangle tiling algorithms. SIAM Journal on Discrete Mathematics, 15(2):252–267, 2002. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/38434.

Bayer:2024:TCC

[BDM^{+}24] Margaret Bayer, Mark Denker, Marija Jeli Milutinovi, Rowan Rowlands, Sheila Sundaram, and Lei Xue. Topology of cut complexes of graphs. SIAM Journal on Discrete Mathematics, 38(2):1630– 1675, May 2024. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Basu:2019:NLI

[BDP19] Amitabh Basu, Santanu S. Dey, and Joseph Paat. Nonunique lifting of integer variables in minimal inequalities. SIAM Journal on Discrete Mathematics, 33(2): 755–783, ???? 2019. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bessy:2018:GHN

[BDPR18] Stéphane Bessy, Mitre C. Dourado, Lucia D. Penso, and Dieter Rautenbach. The geodetic hull number is hard for chordal graphs. SIAM Journal on Discrete Mathematics, 32(1):543–547, ???? 2018. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Black:2017:VPT

[BDT17] Francis Black, Elizabeth Drellich, and Julianna Tymoczko. Valid plane trees: Combinatorial models for RNA secondary structures with Watson–Crick base pairs. SIAM Journal on Discrete Mathematics, 31(4): 2586–2602, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bouman:2013:EMR

[BDvL13] Niek Bouman, Jan Draisma, and Johan S. H. van Leeuwaarden. Energy minimization of repelling particles on a toric grid. SIAM Journal on Discrete Mathematics, 27(3): 1295–1312, ???? 2013. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Buzaglo:2013:TCP

[BE13] Sarit Buzaglo and Tuvi Etzion. Tilings by (0.5, n)- crosses and perfect codes. SIAM Journal on Discrete Mathematics, 27(2):1067– 1081, ???? 2013. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bergren:2022:CSU

[BEGK22] Dan Bergren, Eduard Eiben, Robert Ganian, and Iyad Kanj. On covering segments with unit intervals. SIAM Journal on Discrete Mathematics, 36(2):1200–1230, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL

https://epubs.siam.org/
doi/10.1137/20M1336412.

Benko:2009:ABI

[BEL09] D. Benko, C. Ernst, and D. Lanphier. Asymptotic bounds on the integrity of graphs and separator theorems for graphs. SIAM Journal on Discrete Mathematics, 23(1):265–277, ???? 2009. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bowler:2021:BCN

[BELP21] Nathan Bowler, Joshua Erde, Florian Lehner, and Max Pitz. Bounding the cop number of a graph by its genus. SIAM Journal on Discrete Mathematics, 35(4): 2459–2489, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Blackburn:2008:PCI

[BEN08] Simon R. Blackburn, Tuvi Etzion, and Siaw-Lynn Ng. Prolific codes with the identifiable parent property. SIAM Journal on Discrete Mathematics, 22(4):1393–1410, ???? 2008. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Berman:2007:LSR

[Ber07] Kenneth A. Berman. Locating servers for reliability and affine embeddings. SIAM Journal on Discrete Mathematics, 21(3):637–646, ???? 2007. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Berget:2011:CCS

[BER11] Andrew Berget, Sen-Peng Eu, and Victor Reiner. Constructions for cyclic sieving phenomena. SIAM Journal on Discrete Mathematics, 25(3): 1297–1314, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v25/i3/p1297_s1.

Bernstein:2020:IOB

[Ber20] Daniel Irving Bernstein.

L-infinity optimization to Bergman fans of matroids with an application to phylogenetics. SIAM Journal on Discrete Mathematics, 34(1): 701–720, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Beveridge:2009:CRW

[Bev09] Andrew Beveridge. Centers for random walks on trees. SIAM Journal on Discrete Mathematics, 23(1): 300–318, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Beveridge:2010:CRC

[Bev10] Andrew Beveridge. Connectivity of random cubic sum graphs. SIAM Journal on Discrete Mathematics, 24(3): 895–909, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Barnes:1996:SRW

[BF96] Greg Barnes and Uriel Feige.

Short random walks on graphs. SIAM Journal on Discrete Mathematics, 9(1): 19–28, February 1996. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bal:2012:PTH

[BF12] Deepak Bal and Alan Frieze. Packing tight Hamilton cycles in uniform hypergraphs. SIAM Journal on Discrete Mathematics, 26(2):435–451, ???? 2012. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bruhn:2017:PFG

[BF17] Henning Bruhn and Elke Fuchs. t-perfection in P_{5}-free graphs. SIAM Journal on Discrete Mathematics, 31(3): 1616–1633, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bennett:2024:CNR

[BF24] Patrick Bennett and Alan Frieze. On the chromatic number of random regular hypergraphs. SIAM Journal on Discrete Mathematics, 38(2): 1369–1380, April 2024. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bal:2024:RST

[BFa24] Deepak Bal, Alan Frieze, and Pawe Pra at. Rainbow spanning trees in randomly colored G_{k−out}. SIAM Journal on Discrete Mathematics, 38 (1):867–882, February 2024. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bottcher:2015:LGF

[BFGM15] Albrecht Böttcher, Lenny Fukshansky, Stephan Ramon Garcia, and Hiren Maharaj. On lattices generated by finite Abelian groups. SIAM Journal on Discrete Mathematics, 29(1):382–404, ???? 2015. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Belmonte:2017:MDB

[BFGR17] Rémy Belmonte, Fedor V.

Fomin, Petr A. Golovach, and M. S. Ramanujan. Metric dimension of bounded treelength graphs. SIAM Journal on Discrete Mathematics, 31(2):1217–1243, ???? 2017. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Brewster:2008:NUF

[BFH^{+}08] Richard C. Brewster, Tomas Feder, Pavol Hell, Jing Huang, and Gary MacGillivray. Near-unanimity functions and varieties of reflexive graphs. SIAM Journal on Discrete Mathematics, 22(3): 938–960, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Borgwardt:2015:CDD

[BFH15] Steffen Borgwardt, Elisabeth Finhold, and Raymond Hemmecke. On the circuit diameter of dual transportation polyhedra. SIAM Journal on Discrete Mathematics, 29(1): 113–121, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bruhn-Fujimoto:2021:SEE

[BFH21] Henning Bruhn-Fujimoto and Matthias Heinlein. K_{4}- subdivisions have the Edge– Erdős–Pósa property. SIAM Journal on Discrete Mathematics, 35(1):392–430, ???? 2021. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Barat:2012:LFU

[BFK^{+}12] János Barát, Zoltán Füredi, Ida Kantor, Younjin Kim, and Balázs Patkós. Large B_{d}-free and union-free subfamilies. SIAM Journal on Discrete Mathematics, 26(1): 71–76, ???? 2012. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bhore:2024:OSM

[BFKT24] Sujoy Bhore, Arnold Filtser, Hadi Khodabandeh, and Csaba D. Tóth. Online spanners in metric spaces. SIAM Journal on Discrete Mathematics, 38(1):1030– 1056, March 2024. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bilo:2024:GNC

[BFLM24] Davide Bilò, Tobias Friedrich, Pascal Lenzner, and Anna Melnichenko. Geometric network creation games. SIAM Journal on Discrete Mathematics, 38(1):277–315, January 2024. CODEN SJD-MEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Barahona:1994:CGPd

[BFM94] Francisco Barahona, Jean Fonlupt, and Ali Ridha Mahjoub. Compositions of graphs and polyhedra. IV. Acyclic spanning subgraphs. SIAM Journal on Discrete Mathematics, 7(3): 390–402, August 1994. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bokal:2006:MCN

[BFM06] Drago Bokal, Gasper Fijavz, and Bojan Mohar. The minor crossing number. SIAM Journal on Discrete Mathematics, 20(2):344–356, January 2006. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bulteau:2020:THR

[BFN20] Laurent Bulteau, Vincent Froese, and Rolf Niedermeier. Tight hardness results for consensus problems on circular strings and time series. SIAM Journal on Discrete Mathematics, 34(3): 1854–1883, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Battista:2012:NRP

[BFP12] Giuseppe Di Battista, Fabrizio Frati, and Maurizio Patrignani. Nonconvex representations of plane graphs. SIAM Journal on Discrete Mathematics, 26(4):1670– 1681, ???? 2012. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bermond:2008:NBH

[BFPP08] Jean-Claude Bermond, Afonso Ferreira, Stéphane Pérennes, and Joseph G. Peters. Neighborhood broadcasting in hypercubes. SIAM Journal on Discrete Mathematics, 21(4): 823–843, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bliznets:2015:SPA

[BFPP15] Ivan Bliznets, Fedor V. Fomin, Marcin Pilipczuk, and Michal Pilipczuk. A subexponential parameterized algorithm for proper interval completion. SIAM Journal on Discrete Mathematics, 29(4): 1961–1987, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bulteau:2012:STD

[BFR12] Laurent Bulteau, Guillaume Fertin, and Irena Rusu. Sorting by transpositions is difficult. SIAM Journal on Discrete Mathematics, 26(3): 1148–1180, ???? 2012. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Basavaraju:2016:PPK

[BFRS16] Manu Basavaraju, Mathew C.

Francis, M. S. Ramanujan, and Saket Saurabh. Partially polynomial kernels for set cover and test cover. SIAM Journal on Discrete Mathematics, 30(3):1401–1423, ???? 2016. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bertossi:1988:TDI

[BG88] Alan A. Bertossi and Alessandro Gori. Total domination and irredundance in weighted interval graphs. SIAM Journal on Discrete Mathematics, 1(3):317–327, August 1988. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bienstock:1991:EPS

[BG91] Dan Bienstock and Ervin Györi. An extremal problem on sparse 0-1 matrices. SIAM Journal on Discrete Mathematics, 4(1):17–27, February 1991. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Blanca:2011:UCN

[BG11] Antonio Blanca and Anant P. Godbole. On universal cycles for new classes of combinatorial structures. SIAM Journal on Discrete Mathematics, 25(4):1832–1842, ???? 2011. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v25/i4/p1832_s1.

Brieden:2012:OWB

[BG12] Andreas Brieden and Peter Gritzmann. On optimal weighted balanced clusterings: Gravity bodies and power diagrams. SIAM Journal on Discrete Mathematics, 26(2):415–434, ???? 2012. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bukh:2019:SFP

[BG19] Boris Bukh and Xavier Goaoc. Shatter functions with polynomial growth rates. SIAM Journal on Discrete Mathematics, 33(2): 784–794, ???? 2019. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bhattacharya:2020:UTE

[BG20] Bhaswar B. Bhattacharya and Shirshendu Ganguly. Upper tails for edge eigenvalues of random graphs. SIAM Journal on Discrete Mathematics, 34(2):1069–1083, ???? 2020. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bacso:2004:CMC

[BGG^{+}04] Gábor Bacsó, Sylvain Gravier, András Gyárfás, linebreak Myriam Preissmann, and András Sebo. Coloring the maximal cliques of graphs. SIAM Journal on Discrete Mathematics, 17 (3):361–376, 2004. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/35999.

Blanca:2020:SUP

[BGG^{+}20] Antonio Blanca, Andreas Galanis, Leslie Ann Goldberg, Daniel Stefankovic, Eric Vigoda, and Kuan Yang. Sampling in uniqueness from the Potts and random-cluster models on random regular graphs. SIAM Journal on Discrete Mathematics, 34(1): 742–793, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bhangale:2017:BCC

[BGH^{+}17] Amey Bhangale, Rajiv Gandhi, Mohammad T. Hajiaghayi, Rohit Khandekar, and Guy Kortsarz. Bi-covering: Covering edges with two small subsets of vertices. SIAM Journal on Discrete Mathematics, 31(4):2626–2646, ???? 2017. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bhattacharyya:2012:LBL

[BGJ^{+}12] Arnab Bhattacharyya, Elena Grigorescu, Madhav Jha, Kyomin Jung, Sofya Raskhodnikova, and David P. Woodruff. Lower bounds for local monotonicity reconstruction from transitive-closure spanners. SIAM Journal on Discrete Mathematics, 26(2):618–646, ???? 2012. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Barcelo:2021:VDS

[BGJW21] Hélène Barcelo, Curtis Greene, Abdul Salam Jarrah, and Volkmar Welker. On the vanishing of discrete singular cubical homology for graphs. SIAM Journal on Discrete Mathematics, 35(1): 35–54, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Brent:2003:RKS

[BGL03] Richard P. Brent, Shuhong Gao, and Alan G. B. Lauder. Random Krylov spaces over finite fields. SIAM Journal on Discrete Mathematics, 16(2):276–287, 2003. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/39388.

Berry:2007:RCP

[BGL07] Anne Berry, Martin Charles Golumbic, and Marina Lipshteyn. Recognizing chordal probe graphs and cyclebicolorable graphs. SIAM Journal on Discrete Mathematics, 21(3):573–591, ???? 2007. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Borgwardt:2024:CDP

[BGL24] Steffen Borgwardt, Weston Grewe, and Jon Lee. On the combinatorial diameters of parallel and series connections. SIAM Journal on Discrete Mathematics, 38(1): 485–503, January 2024. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Balister:2007:AVD

[BGLS07] P. N. Balister, E. Györi, J. Lehel, and R. H. Schelp. Adjacent vertex distinguishing edge-colorings. SIAM Journal on Discrete Mathematics, 21(1):237–250, ???? 2007. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Blake:1994:NSD

[BGM94] Ian F. Blake, Shuhong Gao, and Ronald C. Mullin. Normal and self-dual normal bases from factorization of cx^{q+1} + dx^{q} − ax − b. SIAM Journal on Discrete Mathematics, 7(3):499–512, August 1994. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Beaudou:2008:IES

[BGM08] Laurent Beaudou, Sylvain Gravier, and Kahina Meslem. Isometric embeddings of subdivided complete graphs in the hypercube. SIAM Journal on Discrete Mathematics, 22(3):1226–1238, ???? 2008. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bitar:2022:CCB

[BGM22] Nicolas Bitar, Eric Goles, and Pedro Montealegre. Computational complexity of biased diffusion-limited aggregation. SIAM Journal on Discrete Mathematics, 36 (1):823–866, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/18M1215815.

Bartal:2015:IDR

[BGN15] Yair Bartal, Lee-Ad Gottlieb, and Ofer Neiman. On the impossibility of dimension reduction for doubling subsets of ℓ_{p}. SIAM Journal on Discrete Mathematics, 29(3): 1207–1222, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Belfiore:2017:QCS

[BGÖ17] Jean-Claude Belfiore, Cem Güneri, and Buket Özkaya. Quasi-cyclic subcodes of cyclic codes. SIAM Journal on Discrete Mathematics, 31(2):1454–1475, ???? 2017. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bonamy:2022:SLS

[BGP22] Marthe Bonamy, Cyril Gavoille, and Michał Pilipczuk. Shorter labeling schemes for planar graphs. SIAM Journal on Discrete Mathematics, 36 (3):2082–2099, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1330464.

Bouvel:2018:SBS

[BGRR18] Mathilde Bouvel, Veronica Guerrini, Andrew Rechnitzer, and Simone Rinaldi. Semi-Baxter and strong-Baxter: Two relatives of the Baxter sequence. SIAM Journal on Discrete Mathematics, 32(4): 2795–2819, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bierbrauer:1996:OAR

[BGS96] Jürgen Bierbrauer, K. Gopalakrishnan, and D. R. Stinson. Orthogonal arrays, resilient functions, error-correcting codes, and linear programming bounds. SIAM Journal on Discrete Mathematics, 9(3):424–452, August 1996. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bernath:2017:PCC

[BGS17] Attila Bernáth, Roland Grappe, and Zoltán Szigeti. Partition constrained covering of a symmetric crossing supermodular function by a graph. SIAM Journal on Discrete Mathematics, 31(1): 335–382, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Brick:2023:TSR

[BGS23] Lochlan Brick, Pu Gao, and Angus Southwell. The threshold of symmetry in random graphs with specified degree sequences. SIAM Journal on Discrete Mathematics, 37(1):94–113, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1395296.

Brada:2024:RSL

[BGS24] Domagoj Brada, Lior Gishboliner, and Benny Sudakov. On Ramsey sizelinear graphs and related questions. SIAM Journal on Discrete Mathematics, 38(1): 225–242, January 2024. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Balbuena:2007:MOE

[BGV07] C. Balbuena and P. García-Vázquez. On the minimum order of extremal graphs to have a prescribed girth. SIAM Journal on Discrete Mathematics, 21(1): 251–257, ???? 2007. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bogart:2020:PVL

[BGW20] Tristram Bogart, John Goodrick, and Kevin Woods. A parametric version of LLL and some consequences: Parametric shortest and closest vector problems. SIAM Journal on Discrete Mathematics, 34(4): 2363–2387, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Beame:2020:BRM

[BGY20] Paul Beame, Shayan Oveis Gharan, and Xin Yang. On the bias of Reed–Muller codes over odd prime fields. SIAM Journal on Discrete Mathematics, 34(2):1232–1247, ???? 2020. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Boyd:1993:IPR

[BH93] Sylvia C. Boyd and Tianbao Hao. An integer polytope related to the design of survivable communication networks. SIAM Journal on Discrete Mathematics, 6(4):612– 630, November 1993. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bertram:1997:DRG

[BH97] Edward Bertram and Peter Horák. Decomposing 4regular graphs into trianglefree 2-factors. SIAM Journal on Discrete Mathematics, 10 (2):309–317, May 1997. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Brueni:2005:PPP

[BH05] Dennis J. Brueni and Lenwood S. Heath. The PMU placement problem. SIAM Journal on Discrete Mathematics, 19(3):744–761, 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/43255.

Boyd:2013:MCM

[BH13a] Sylvia Boyd and Maryam Haghighi. Mixed and circular multichromosomal genomic median problem. SIAM Journal on Discrete Mathematics, 27(1):63–74, ???? 2013. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Brijder:2013:NLC

[BH13b] Robert Brijder and Hendrik Jan Hoogeboom. Nullity and loop complementation for delta-matroids. SIAM Journal on Discrete Mathematics, 27(1):492–506, ???? 2013. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bryant:2015:STS

[BH15] Darryn Bryant and Daniel Horsley. Steiner triple systems without parallel classes. SIAM Journal on Discrete Mathematics, 29(1):693–696, ???? 2015. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Berg:2016:LPI

[BH16] Sören Lennart Berg and Martin Henk. Lattice point inequalities for centered convex bodies. SIAM Journal on Discrete Mathematics, 30(2): 1148–1158, ???? 2016. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bukh:2020:LLC

[BH20] Boris Bukh and Raymond Hogenson. Length of the longest common subsequence between overlapping words. SIAM Journal on Discrete Mathematics, 34(1):721–729, ???? 2020. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bowtell:2022:DSS

[BH22] Candida Bowtell and Joseph Hyde. A degree sequence strengthening of the vertex degree threshold for a perfect matching in 3-uniform hypergraphs. SIAM Journal on Discrete Mathematics, 36(2):1038–1063, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/20M1364825.

Ben-Haim:2005:OCL

[BHE05] Yael Ben-Haim and Tuvi Etzion. On the optimality of coloring with a lattice. SIAM Journal on Discrete Mathematics, 18(4):844–878, 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/43958.

Boros:1994:PCE

[BHH94] E. Boros, P. L. Hammer, and J. N. Hooker. Predicting cause-effect relationships from incomplete discrete observations. SIAM Journal on Discrete Mathematics, 7 (4):531–543, November 1994. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bhattacharya:1996:LAM

[BHH96] Binay Bhattacharya, Pavol Hell, and Jing Huang. A linear algorithm for maximum weight cliques in proper circular arc graphs. SIAM Journal on Discrete Mathematics, 9(2):274–289, May 1996. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bruhn:2018:FPE

[BHJ18] Henning Bruhn, Matthias Heinlein, and Felix Joos. Frames, A-paths, and the Erdős–Pósa property. SIAM Journal on Discrete Mathematics, 32(2):1246–1260, ???? 2018. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Butler:2008:GG

[BHKL08] Steve Butler, Mohammad T.

Hajiaghayi, Robert D. Kleinberg, and Tom Leighton. Hat guessing games. SIAM Journal on Discrete Mathematics, 22(2):592–605, ???? 2008. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ben-Haim:2005:EMD

[BHL05] Yael Ben-Haim and Simon Litsyn. Exact minimum density of codes identifying vertices in the square grid. SIAM Journal on Discrete Mathematics, 19(1):69–82, 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/44408.

Bernstein:2015:BES

[BHL^{+}15] Daniel Irving Bernstein, Lam Si Tung Ho, Colby Long, Mike Steel, Katherine St. John, and Seth Sullivant. Bounds on the expected size of the maximum agreement subtree. SIAM Journal on Discrete Mathematics, 29(4): 2065–2074, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bermond:1992:BBD

[BHLP92] Jean-Claude Bermond, Pavol Hell, Arthur L. Liestman, and Joseph G. Peters. Broadcasting in bounded degree graphs. SIAM Journal on Discrete Mathematics, 5(1): 10–24, February 1992. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Balogh:2008:FFC

[BHLY08] József Balogh, Stephen G. Hartke, Qi Liu, and Gexin Yu. On the first-fit chromatic number of graphs. SIAM Journal on Discrete Mathematics, 22(3):887–900, ???? 2008. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Brankovic:2000:OPS

[BHM00] Ljiljana Brankovic, Peter Horak, and Mirka Miller. An optimization problem in statistical databases. SIAM Journal on Discrete Mathematics, 13(3):346–353, 2000. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/34500.

Benedetti:2016:CHA

[BHM16] Carolina Benedetti, Joshua Hallam, and John Machacek. Combinatorial Hopf algebras of simplicial complexes. SIAM Journal on Discrete Mathematics, 30(3):1737– 1757, ???? 2016. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bohm:2022:HCS

[BHM^{+}22] Martin Böhm, Ruben Hoeksma, Nicole Megow, Lukas Nölke, and Bertrand Simon. On hopconstrained Steiner trees in tree-like metrics. SIAM Journal on Discrete Mathematics, 36(2):1249–1273, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1425487.

Bereg:2016:RPF

[BHNP16] Sergey Bereg, Alexander E. Holroyd, Lev Nachmanson, and Sergey Pupyrev. Representing permutations with few moves. SIAM Journal on Discrete Mathematics, 30(4): 1950–1977, ???? 2016. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Banderier:2014:AES

[BHRZ14] Cyril Banderier, Hsien-Kuei Hwang, Vlady Ravelomanana, and Vytas Zacharovas. Analysis of an exhaustive search algorithm in random graphs and the n^{c log n}-asymptotics. SIAM Journal on Discrete Mathematics, 28(1):342–371, ???? 2014. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bottcher:2010:EBG

[BHT10] Julia Böttcher, Peter Heinig, and Anusch Taraz. Embedding into bipartite graphs. SIAM Journal on Discrete Mathematics, 24(4):1215– 1233, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bujtas:2016:TGH

[BHT16] Csilla Bujtás, Michael A.

Henning, and Zsolt Tuza. Transversal game on hypergraphs and the \frac{3}{4}-conjecture on the total domination game. SIAM Journal on Discrete Mathematics, 30(3): 1830–1847, ???? 2016. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bogart:2021:CPM

[BHTK^{+}21] Tristram Bogart, Anna-Lena Horlemann-Trautmann, David Karpuk, Alessandro Neri, and Mauricio Velasco. Constructing partial MDS codes from reducible algebraic curves. SIAM Journal on Discrete Mathematics, 35(4):2946–2970, ???? 2021. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Beimel:2005:PNS

[BI05] Amos Beimel and Yuval Ishai.

On the power of nonlinear secret-sharing. SIAM Journal on Discrete Mathematics, 19(1):258–280, 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/41286.

Burgisser:2013:DPL

[BI13] Peter Bürgisser and Christian Ikenmeyer. Deciding positivity of Littlewood–Richardson coefficients. SIAM Journal on Discrete Mathematics, 27(4):

1639–1681, ???? 2013. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bienstock:1988:AAS

[Bie88] Daniel Bienstock. Asymptotic analysis of some network reliability models. SIAM Journal on Discrete Mathematics, 1(1):14–21, February 1988. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Biha:2005:TCP

[Bih05] Mohamed Didi Biha. On the 3-terminal cut polyhedron. SIAM Journal on Discrete Mathematics, 19 (3):575–587, 2005. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/44514.

Berczi:2018:MBG

[BIKY18] Kristóf Bérczi, Satoru Iwata, Jun Kato, and Yutaro Yamaguchi. Making bipartite graphs DM-Irreducible. SIAM Journal on Discrete Mathematics, 32(1):560–590, ???? 2018. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bresar:2005:HDP

[BIKZ05] Bostjan Bresar, Wilfried Imrich, Sandi Klavzar, and Blaz Zmazek. Hypercubes as direct products. SIAM Journal on Discrete Mathematics, 18(4):778–786, 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/43835.

Bilbao:2003:CGU

[Bil03] Jesús Mario Bilbao. Cooperative games under augmenting systems. SIAM Journal on Discrete Mathematics, 17(1): 122–133, January 2003. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Blaasten:2023:IPD

[BILL23] Emilia Blåsten, Hiroshi Isozaki, Matti Lassas, and Jinpeng Lu. Inverse problems for discrete heat equations and random walks for a class of graphs. SIAM Journal on Discrete Mathematics, 37(2):831–863, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1439936.

Bei:2022:PCF

[BILS22] Xiaohui Bei, Ayumi Igarashi, Xinhang Lu, and Warut Suksompong. The price of connectivity in fair division. SIAM Journal on Discrete Mathematics, 36 (2):1156–1186, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/20M1388310.

Boyd:2013:FFC

[BIT13] Sylvia Boyd, Satoru Iwata, and Kenjiro Takazawa. Finding 2-factors closer to TSP tours in cubic graphs. SIAM Journal on Discrete Mathematics, 27(2):918–939, ???? 2013. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Barthelemy:1991:FTC

[BJ91] J.-P. Barthélémy and M. F. Janowitz. A formal theory of consensus. SIAM Journal on Discrete Mathematics, 4(3): 305–322, August 1991. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Benjamin:1992:AWC

[BJ92] Arthur T. Benjamin and Fritz Juhnke. Another way of counting N^{N}. SIAM Journal on Discrete Mathematics, 5(3):377–379, August 1992. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bukh:2023:PCC

[BJ23] Boris Bukh and R. Amzi Jeffs. Planar convex codes are decidable. SIAM Journal on Discrete Mathematics, 37(2):951–963, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1511187.

Bang-Jensen:1995:PIL

[BJFJ95] Jørgen Bang-Jensen, András Frank, and Bill Jackson. Preserving and increasing local edge-connectivity in mixed graphs. SIAM Journal on Discrete Mathematics, 8(2): 155–178, May 1995. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bang-Jensen:1999:ECA

[BJGJS99] Jørgen Bang-Jensen, Harold N.

Gabow, Tibor Jordán, and Zoltán Szigeti. Edgeconnectivity augmentation with partition constraints. SIAM Journal on Discrete Mathematics, 12(2):160–207, May 1999. CODEN SJD-MEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/32470.

Bang-Jensen:1988:CCS

[BJHM88] Jørgen Bang-Jensen, Pavol Hell, and Gary MacGillivray. The complexity of colouring by semicomplete digraphs. SIAM Journal on Discrete Mathematics, 1(3): 281–298, August 1988. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bang-Jensen:2000:CRC

[BJHY00] Jørgen Bang-Jensen, Jing Huang, and Anders Yeo. Convex-round and concaveround graphs. SIAM Journal on Discrete Mathematics, 13(2):179–193, 2000. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/32215.

Bang-Jensen:2003:SCS

[BJHY03] Jorgen Bang-Jensen, Jing Huang, and Anders Yeo. Strongly connected spanning subdigraphs with the minimum number of arcs in quasitransitive digraphs. SIAM Journal on Discrete Mathematics, 16(2):335–343, 2003. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/35422.

Bang-Jensen:1998:ECA

[BJJ98] Jørgen Bang-Jensen and Tibor Jordán. Edge-connectivity augmentation preserving simplicity. SIAM Journal on Discrete Mathematics, 11(4): 603–623, November 1998. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bodlaender:2013:PTC

[BJK13] Hans L. Bodlaender, Bart M. P. Jansen, and Stefan Kratsch. Preprocessing for treewidth: a combinatorial analysis through kernelization. SIAM Journal on Discrete Mathematics, 27(4): 2108–2142, ???? 2013. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bodlaender:2014:KLB

[BJK14] Hans L. Bodlaender, Bart M. P. Jansen, and Stefan Kratsch. Kernelization lower bounds by crosscomposition. SIAM Journal on Discrete Mathematics, 28 (1):277–305, ???? 2014. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Babilon:2007:LGF

[BJKV07] Robert Babilon, Vít Jelínek, Daniel Král’, and Pavel Valtr. Labelings of graphs with fixed and variable edgeweights. SIAM Journal on Discrete Mathematics, 21(3): 688–706, ???? 2007. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Behague:2024:RSN

[BJL^{+}24] Natalie Behague, Tom Johnston, Shoham Letzter, Natasha Morrison, and Shannon Ogden. The rainbow saturation number is linear. SIAM Journal on Discrete Mathematics, 38(2):1239–1249, April 2024. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bruhn:2021:EPP

[BJS21] Henning Bruhn, Felix Joos, and Oliver Schaudt. Erdős– Pósa property for labeled minors: 2-connected minors. SIAM Journal on Discrete Mathematics, 35(2): 893–914, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bougeret:2022:BDC

[BJS22] Marin Bougeret, Bart M. P. Jansen, and Ignasi Sau. Bridge-depth characterizes which minor-closed structural parameterizations of vertex cover admit a polynomial kernel. SIAM Journal on Discrete Mathematics, 36(4):2737–2773, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1400766.

Bang-Jensen:2022:DVD

[BJSS22] Jørgen Bang-Jensen, Thomas Schweser, and Michael Stiebitz. Digraphs and variable degeneracy. SIAM Journal on Discrete Mathematics, 36(1):578–595, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/20M1386827.

Bang-Jensen:1992:PAP

[BJT92] Jørgen Bang-Jensen and Carsten Thomassen. A polynomial algorithm for the 2path problem for semicomplete digraphs. SIAM Journal on Discrete Mathematics, 5(3):366–376, August 1992. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Beutelspacher:1992:PSS

[BJvV92] A. Beutelspacher, D. Jungnickel, P. C. van Oorschot, and S. A. Vanstone. Pairsplitting sets in AG(m, q). SIAM Journal on Discrete Mathematics, 5(4):451–459, November 1992. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Berman:1990:RPC

[BK90] Kenneth A. Berman and Mokhtar H. Konsowa. Random paths and cuts, electrical networks, and reversible Markov chains. SIAM Journal on Discrete Mathematics, 3(3):311–319, August 1990. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Blake:1991:CWE

[BK91] Ian F. Blake and Khun Kith.

On the complete weight enumerator of Reed–Solomon codes. SIAM Journal on Discrete Mathematics, 4(2):

164–171, May 1991. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Burstein:2005:USW

[BK05] Alexander Burstein and Sergey Kitaev. On unavoidable sets of word patterns. SIAM Journal on Discrete Mathematics, 19 (2):371–381, 2005. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/44567.

Bresar:2007:CGJ

[BK07] Boštjan Brešar and Sandi Klavžar. Crossing graphs as joins of graphs and Cartesian products of median graphs. SIAM Journal on Discrete Mathematics, 21(1): 26–32, ???? 2007. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bodirsky:2011:LBL

[BK11] Manuel Bodirsky and Daniel Král’. Limit behavior of locally consistent constraint satisfaction problems. SIAM Journal on Discrete Mathematics, 25(2): 916–933, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v25/i2/p916_s1.

Bogart:2012:OLT

[BK12] Tristram Bogart and Eric Katz. Obstructions to lifting tropical curves in surfaces in 3-space. SIAM Journal on Discrete Mathematics, 26(3): 1050–1067, ???? 2012. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bushaw:2014:TNF

[BK14] Neal Bushaw and Nathan Kettle. Turán numbers for forests of paths in hypergraphs. SIAM Journal on Discrete Mathematics, 28(2): 711–721, ???? 2014. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Brazitikos:2017:ARL

[BK17] Silouanos Brazitikos and Thodoris Karageorgos. An algorithmic regularity lemma for L_{p} regular sparse matrices. SIAM Journal on Discrete Mathematics, 31(4): 2301–2313, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bernstein:2019:TCM

[BK19] Daniel I. Bernstein and Robert Krone. The tropical Cayley–Menger variety. SIAM Journal on Discrete Mathematics, 33(3): 1725–1742, ???? 2019. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bergougnoux:2021:MAN

[BK21] Benjamin Bergougnoux and Mamadou Moustapha Kanté. More applications of the dneighbor equivalence: Acyclicity and connectivity constraints. SIAM Journal on Discrete Mathematics, 35(3): 1881–1926, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). See erratum [BK24].

Boskovic:2023:SOG

[BK23] Vladimir Bosković and Balázs Keszegh. Saturation of ordered graphs. SIAM Journal on Discrete Mathematics, 37(2):1118–1141, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1485735.

Bergougnoux:2024:EMA

[BK24] Benjamin Bergougnoux and Mamadou M. Kanté. Erratum: More applications of the d-neighbor equivalence: Acyclicity and connectivity constraints. SIAM Journal on Discrete Mathematics, 38(1): 1191–1192, March 2024. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). See [BK21].

Bolotashvili:1999:NFL

[BKG99] G. Bolotashvili, M. Kovalev, and E. Girlich. New facets of the linear ordering polytope. SIAM Journal on Discrete Mathematics, 12 (3):326–336, August 1999. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/30014.

Bodlaender:1995:TPP

[BKK95] Hans L. Bodlaender, Ton Kloks, and Dieter Kratsch. Treewidth and pathwidth of permutation graphs. SIAM Journal on Discrete Mathematics, 8(4):606–616, November 1995. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Berczi:2016:CIB

[BKK16] Kristóf Bérczi, Tamás Király, and Yusuke Kobayashi. Covering intersecting bi-set families under matroid constraints. SIAM Journal on Discrete Mathematics, 30(3): 1758–1774, ???? 2016. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Berczi:2021:MPM

[BKK21] Kristóf Bérczi, Naonori Kakimura, and Yusuke Kobayashi. Market pricing for matroid rank valuations. SIAM Journal on Discrete Mathematics, 35(4): 2662–2678, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Broersma:1999:ISA

[BKKM99] Hajo Broersma, Ton Kloks, Dieter Kratsch, and Haiko Müller. Independent sets in asteroidal triple-free graphs. SIAM Journal on Discrete Mathematics, 12(2):276–287, May 1999. CODEN SJD-MEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/32634.

Bjorklund:2017:STF

[BKKZ17] Andreas Björklund, Vikram Kamat, Lukasz Kowalik, and Meirav Zehavi. Spotting trees with few leaves. SIAM Journal on Discrete Mathematics, 31(2):687–713, ???? 2017. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bhawalkar:2015:PUS

[BKL^{+}15] Kshipra Bhawalkar, Jon Kleinberg, Kevin Lewi, Tim Roughgarden, and Aneesh Sharma. Preventing unraveling in social networks: The anchored k-core problem. SIAM Journal on Discrete Mathematics, 29(3): 1452–1475, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Belmonte:2022:GDT

[BKL^{+}22] Rémy Belmonte, Eun Jung Kim, Michael Lampis, Valia Mitsou, and Yota Otachi. Grundy distinguishes treewidth from pathwidth. SIAM Journal on Discrete Mathematics, 36(3):1761–1787, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1385779.

Biha:2008:PSN

[BKM08] Mohamed Didi Biha, Hervé Kerivin, and A. Ridha Mahjoub. On the polytope of the (1, 2)-survivable network design problem. SIAM Journal on Discrete Mathematics, 22(4):1640–1666, ???? 2008. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bosek:2013:FFC

[BKM13] Bartlomiej Bosek, Tomasz Krawczyk, and Grzegorz Matecki. First-fit coloring of incomparability graphs. SIAM Journal on Discrete Mathematics, 27(1):126–140, ???? 2013. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bernath:2015:GTB

[BKM15] Attila Bernáth, Yusuke Kobayashi, and Tatsuya Matsuoka. The generalized terminal backup problem. SIAM Journal on Discrete Mathematics, 29(3): 1764–1782, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bodlaender:2021:PCC

[BKP21] Hans L. Bodlaender, Sudeshna Kolay, and Astrid Pieterse. Parameterized complexity of conflict-free graph coloring. SIAM Journal on Discrete Mathematics, 35(3):2003– 2038, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bresar:2010:DGI

[BKR10] Boštjan Brešar, Sandi Klavžar, and Douglas F. Rall. Domination game and an imagination strategy. SIAM Journal on Discrete Mathematics, 24 (3):979–991, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Berry:2009:MLS

[BKS09] A. Berry, R. Krueger, and G. Simonet. Maximal label search algorithms to compute perfect and minimal elimination orderings. SIAM Journal on Discrete Mathematics, 23 (1):428–446, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bosek:2010:FFA

[BKS10] Bartłomiej Bosek, Tomasz Krawczyk, and Edward Szczypka. First-fit algorithm for the on-line chain partitioning problem. SIAM Journal on Discrete Mathematics, 23(4):1992–1999, ???? 2010. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bottcher:2015:EBL

[BKTW15] Julia Böttcher, Yoshiharu Kohayakawa, Anusch Taraz, and Andreas Würfl. An extension of the blow-up lemma to arrangeable graphs. SIAM Journal on Discrete Mathematics, 29(2):962–1001, ???? 2015. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Berczi:2023:MIU

[BKYY23] Kristóf Bérczi, Tamás Király, Yutaro Yamaguchi, and Yu Yokoi. Matroid intersection under restricted oracles. SIAM Journal on Discrete Mathematics, 37 (2):1311–1330, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M152579X.

Bollobas:1990:IID

[BL90] Béla Bollobás and Imre Leader. An isoperimetric inequality on the discrete torus. SIAM Journal on Discrete Mathematics, 3(1): 32–37, February 1990. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Brys:2001:MPC

[BL01] Krzysztof Bryś and Zbigniew Lonc. Monochromatic partitions of complete uniform hypergraphs. SIAM Journal on Discrete Mathematics, 14(3):286–290, 2001. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/35771.

Brandstadt:2004:SPG

[BL04] Andreas Brandstädt and Van Bang Le. Split-perfect graphs: Characterizations and algorithmic use. SIAM Journal on Discrete Mathematics, 17(3):341–360, 2004. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/36767.

Bezrukov:2009:SPK

[BL09] Sergei L. Bezrukov and Uwe Leck. A simple proof of the Karakhanyan–Riordan theorem on the even discrete torus. SIAM Journal on Discrete Mathematics, 23(3): 1416–1421, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bond:2016:ANF

[BL16] Benjamin Bond and Lionel Levine. Abelian networks I. Foundations and examples. SIAM Journal on Discrete Mathematics, 30(2): 856–874, ???? 2016. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bernstein:2017:IOL

[BL17a] Daniel Irving Bernstein and Colby Long. L-infinity optimization to linear spaces and phylogenetic trees. SIAM Journal on Discrete Mathematics, 31(2):875–889, ???? 2017. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Boyd:2017:TBM

[BL17b] Sylvia Boyd and Philippe Legault. Toward a 6/5 bound for the minimum cost 2edge connected spanning subgraph. SIAM Journal on Discrete Mathematics, 31(1): 632–644, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Balogh:2019:TSG

[BL19a] József Balogh and Lina Li.

The typical structure of Gallai colorings and their extremal graphs. SIAM Journal on Discrete Mathematics, 33(4):2416–2443, ???? 2019. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Basit:2019:ISP

[BL19b] Abdul Basit and Ben Lund. An improved sum-product bound for quaternions. SIAM Journal on Discrete Mathematics, 33(2):1044–1060, ???? 2019. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bonin:2023:EMT

[BL23] Joseph E. Bonin and Kevin Long. The excluded minors for three classes of 2-polymatroids having special types of natural matroids. SIAM Journal on Discrete Mathematics, 37 (3):1715–1737, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1521134.

Blackburn:2003:FC

[Bla03] Simon R. Blackburn. Frameproof codes. SIAM Journal on Discrete Mathematics, 16(3):499–510, 2003. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/38463.

Bousquet:2015:ICH

[BLL^{+}15] Nicolas Bousquet, Aurélie Lagoutte, Zhentao Li, Aline Parreau, and Stéphan Thomassé. Identifying codes in hereditary classes of graphs and VC-dimension. SIAM Journal on Discrete Mathematics, 29(4):2047–2064, ???? 2015. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Brandstadt:2010:ISM

[BLM10] Andreas Brandstädt, Vadim V.

Lozin, and Raffaele Mosca. Independent sets of maximum weight in apple-free graphs. SIAM Journal on Discrete Mathematics, 24(1): 239–254, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Belmonte:2020:PAD

[BLM20] Rémy Belmonte, Michael Lampis, and Valia Mitsou. Parameterized (approximate) defective coloring. SIAM Journal on Discrete Mathematics, 34(2):1084–1106, ???? 2020. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Blumenthal:2022:CCH

[BLM^{+}22] Adam Blumenthal, Bernard Lidický, Ryan R. Martin, Sergey Norin, Florian Pfender, and Jan Volec. Counterexamples to a conjecture of Harris on Hall ratio. SIAM Journal on Discrete Mathematics, 36 (3):1678–1686, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/18M1229420.

Berstein:2008:NMO

[BLMA^{+}08] Yael Berstein, Jon Lee, Hugo Maruri-Aguilar, Shmuel Onn, Eva Riccomagno, Robert Weismantel, and Henry Wynn. Nonlinear matroid optimization and experimental design. SIAM Journal on Discrete Mathematics, 22(3): 901–919, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bartal:2000:MSR

[BLMS^{+}00] Yair Bartal, Stefano Leonardi, Alberto Marchetti-Spaccamela, Jirí Sgall, and Leen Stougie. Multiprocessor scheduling with rejection. SIAM Journal on Discrete Mathematics, 13(1):64–78, 2000. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/30052.

Bloznelis:2010:CEG

[Blo10] Mindaugas Bloznelis. Component evolution in general random intersection graphs. SIAM Journal on Discrete Mathematics, 24(2):639–654, ???? 2010. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bordewich:2022:MAS

[BLO^{+}22] Magnus Bordewich, Simone Linz, Megan Owen, Katherine St.John, Charles Semple, and Kristina Wicke. On the maximum agreement subtree conjecture for balanced trees. SIAM Journal on Discrete Mathematics, 36(1):336–354, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1379678.

Balogh:2021:MSI

[BLP21] Jozsef Balogh, Nathan Lemons, and Cory Palmer. Maximum size intersecting families of bounded minimum positive co-degree. SIAM Journal on Discrete Mathematics, 35(3): 1525–1535, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bondy:2016:COR

[BLR16] Adrian Bondy, Zbigniew Lonc, and Pawel Rzazewski. Constructing optimal kradius sequences. SIAM Journal on Discrete Mathematics, 30(1):452–464, ???? 2016. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). See erratum [BLR17].

Bondy:2017:ECO

[BLR17] Adrian Bondy, Zbigniew Lonc, and Pawel Rzazewski. Erratum: Constructing optimal k-radius sequences. SIAM Journal on Discrete Mathematics, 31(1):645–646, ???? 2017. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). See [BLR16].

Balogh:2017:TPR

[BLS17] József Balogh, Hong Liu, and Maryam Sharifzadeh. On two problems in Ramsey–Turán theory. SIAM Journal on Discrete Mathematics, 31(3): 1848–1866, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Balogh:2019:CHC

[BLS19] József Balogh, Bernard Lidický, and Gelasio Salazar. Closing in on Hill’s conjecture. SIAM Journal on Discrete Mathematics, 33(3):1261–1276, ???? 2019. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bodlaender:1993:PTC

[BM93] Hans L. Bodlaender and Rolf H. Möhring. The pathwidth and treewidth of cographs. SIAM Journal on Discrete Mathematics, 6(2): 181–188, May 1993. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Barahona:1994:CGPa

[BM94a] Francisco Barahona and Ali Ridha Mahjoub. Compositions of graphs and polyhedra. I. Balanced induced subgraphs and acyclic subgraphs. SIAM Journal on Discrete Mathematics, 7(3): 344–358, August 1994. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Barahona:1994:CGPb

[BM94b] Francisco Barahona and Ali Ridha Mahjoub. Compositions of graphs and polyhedra. II. Stable sets. SIAM Journal on Discrete Mathematics, 7(3):359–371, August 1994. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Barahona:1994:CGPc

[BM94c] Francisco Barahona and Ali Ridha Mahjoub. Compositions of graphs and polyhedra. III. Graphs with no W_{4} minor. SIAM Journal on Discrete Mathematics, 7(3): 372–389, August 1994. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Baiou:1997:SEC

[BM97] Mourad Baïou and Ali Ridha Mahjoub. Steiner 2-edge connected subgraph polytopes on series-parallel graphs. SIAM Journal on Discrete Mathematics, 10(3):505–514, August 1997. CODEN SJD-MEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Barbero:2000:WHH

[BM00] Angela I. Barbero and Carlos Munuera. The weight hierarchy of Hermitian codes. SIAM Journal on Discrete Mathematics, 13(1):79–104, 2000. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/34342.

Barany:2007:QMC

[BM07] Imre Bárány and Jiří Matoušek. Quadratically many colorful simplices. SIAM Journal on Discrete Mathematics, 21(1):191–198, ???? 2007. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Balogh:2009:AEG

[BM09] József Balogh and Ryan Martin. On avoider-enforcer games. SIAM Journal on Discrete Mathematics, 23(2): 901–908, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bhatnagar:2011:CMB

[BM11] Nayantara Bhatnagar and Elitza Maneva. A computational method for bounding the probability of reconstruction on trees. SIAM Journal on Discrete Mathematics, 25 (2):854–871, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v25/i2/p854_s1.

Brightwell:2013:DPL

[BM13] Graham Brightwell and Mareike Massow. Diametral pairs of linear extensions. SIAM Journal on Discrete Mathematics, 27(2):634–649, ???? 2013. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bukh:2014:LCS

[BM14] Boris Bukh and Jie Ma.

Longest common subsequences in sets of words. SIAM Journal on Discrete Mathematics, 28(4):2042– 2049, ???? 2014. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bshouty:2015:PCM

[BM15] Nader H. Bshouty and Hanna Mazzawi. On parity check (0, 1)-matrix over Z_{p}. SIAM Journal on Discrete Mathematics, 29(1):631–657, ???? 2015. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Brandstadt:2016:WED

[BM16] Andreas Brandstädt and Raffaele Mosca. Weighted efficient domination for P_{5}-free and P_{6}-free graphs. SIAM Journal on Discrete Mathematics, 30(4):2288–2303, ???? 2016. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bai:2019:NVD

[BM19] Yandong Bai and Yannis Manoussakis. On the number of vertex-disjoint cycles in digraphs. SIAM Journal on Discrete Mathematics, 33(4): 2444–2451, ???? 2019. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bharathi:2022:IMP

[BM22] Arpitha P. Bharathi and Monaldo Mastrolilli. Ideal membership problem over 3-element CSPs with dual discriminator polymorphism. SIAM Journal on Discrete Mathematics, 36(3):1800– 1822, ???? 2022. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:/

                                        /epubs.siam.org/doi/10.
                                        1137/21M1397131.

Bhattacharya:2020:SMP

[BMM20] Bhaswar B. Bhattacharya, Somabha Mukherjee, and Sumit Mukherjee. The second-moment phenomenon for monochromatic subgraphs. SIAM Journal on Discrete Mathematics, 34(1): 794–824, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Brewster:2013:FMD

[BMN13] Richard C. Brewster, Sean McGuinness, and Morten Hegner Nielsen. Factors with multiple degree constraints in graphs. SIAM Journal on Discrete Mathematics, 27(4): 1734–1747, ???? 2013. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Balogh:2022:GTR

[BMNP22] József Balogh, Ryan R. Martin, Dániel T. Nagy, and Balázs Patkós. On generalized Turán results in height two posets. SIAM Journal on Discrete Mathematics, 36 (2):1483–1495, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1457254.

Bradshaw:2022:RCG

[BMNS22] Peter Bradshaw, Tomás Masarík, Jana Novotná, and Ladislav Stacho. Robust connectivity of graphs on surfaces. SIAM Journal on Discrete Mathematics, 36 (2):1416–1435, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1417077.

Bonato:2013:VPR

[BMP13] Anthony Bonato, Dieter Mitsche, and Pawel Pralat. Vertex-pursuit in random directed acyclic graphs. SIAM Journal on Discrete Mathematics, 27(2):732–756, ???? 2013. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Brianski:2021:EHP

[BMPS21] Marcin Briański, Piotr Micek, Michał Pilipczuk, and Michał T. Seweryn. Erdős–Hajnal properties for powers of sparse graphs. SIAM Journal on Discrete Mathematics, 35(1): 447–464, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Brandt:2010:EIE

[BMR^{+}10] Stephan Brandt, Jozef Miškuf, Dieter Rautenbach, Friedrich Regen, and Imre Z. Ruzsa. Edge-injective and edgesurjective vertex labellings. SIAM Journal on Discrete Mathematics, 24(2):666–683, ???? 2010. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Becker:2020:ILB

[BMRT20] F. Becker, P. Montealegre, I. Rapaport, and I. Todinca. The impact of locality in the broadcast congested clique model. SIAM Journal on Discrete Mathematics, 34 (1):682–700, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Belfrage:2012:POP

[BMS12] Michael Belfrage, Torsten Mütze, and Reto Spöhel. Probabilistic one-player Ramsey games via deterministic two-player games. SIAM Journal on Discrete Mathematics, 26(3):1031–1049, ???? 2012. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Botti:1992:LPT

[BMV92] Phillip Botti, Russell Merris, and Cheryl Vega. Laplacian permanents of trees. SIAM Journal on Discrete Mathematics, 5(4):460–466, November 1992. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Berczi:2022:AMM

[BMV22] Kristóf Bérczi, Matthias Mnich, and Roland Vincze. A 3/2-approximation for the metric many-visits path TSP. SIAM Journal on Discrete Mathematics, 36 (4):2995–3030, ???? 2022.

CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1483414.

Bolognini:2017:PPD

[BMVW17] Davide Bolognini, Antonio Macchia, Emanuele Ventura, and Volkmar Welker. The poset of proper divisibility. SIAM Journal on Discrete Mathematics, 31(3): 2093–2109, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Boppana:1996:BCP

[BN96] Ravi B. Boppana and Babu O. Narayanan. The biased coin problem. SIAM Journal on Discrete Mathematics, 9(1): 29–36, February 1996. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Blackmore:2001:DWA

[BN01] T. Blackmore and G. H. Norton. Determining when the absolute state complexity of a Hermitian code achieves its DLP bound. SIAM Journal on Discrete Mathematics, 15(1):14–40, 2001. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/37643.

Brown:2005:WCV

[BN05] J. I. Brown and R. J.

Nowakowski. Well-covered vector spaces of graphs. SIAM Journal on Discrete Mathematics, 19(4):952–965, January 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bar-Noy:2020:ERI

[BNCPR20] Amotz Bar-Noy, Keerti Choudhary, David Peleg, and Dror Rawitz. Efficiently realizing interval sequences. SIAM Journal on Discrete Mathematics, 34(4):2318–2337, ???? 2020. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bar-Noy:1992:LTA

[BNMN92] Amotz Bar-Noy, Rajeev Motwani, and Joseph Naor. A linear time approach to the set maxima problem. SIAM Journal on Discrete Mathematics, 5(1):1–9, February 1992. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bar-Noy:1990:SMF

[BNN90] Amotz Bar-Noy and Joseph Naor. Sorting, minimal feedback sets, and Hamilton paths in tournaments. SIAM Journal on Discrete Mathematics, 3(1):7–20, February 1990. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Brown:1996:UCI

[BNR96] Jason I. Brown, Richard J. Nowakowski, and Douglas Rall. The ultimate categorical independence ratio of a graph. SIAM Journal on Discrete Mathematics, 9(2): 290–300, May 1996. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bar-Noy:2017:MBC

[BNRT17] Amotz Bar-Noy, Dror Rawitz, and Peter Terlecky. Maximizing barrier coverage lifetime with mobile sensors. SIAM Journal on Discrete Mathematics, 31(1):573–596, ???? 2017. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bentert:2023:UGL

[BNRZ23] Matthias Bentert, André Nichterlein, Malte Renken, and Philipp Zschoche. Using a geometric lens to find kdisjoint shortest paths. SIAM Journal on Discrete Mathematics, 37(3):1674–1703, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL

                                        https://epubs.siam.org/
                                        doi/10.1137/22M1527398.

Bouyukliev:2005:CSO

[BÖ05] Iliya Bouyukliev and Patric R. J. Östergard. Classification of self-orthogonal codes over F_{3} and F_{4}. SIAM Journal on Discrete Mathematics, 19(2):363–370, 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/44108.

Bolger:1990:CEB

[Bol90] E. M. Bolger. A characterization of an extension of the Banzhaf value to multicandidate voting games. SIAM Journal on Discrete Mathematics, 3(4):466–477, November 1990. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bonuccelli:1991:PTO

[Bon91] M. A. Bonuccelli. A polynomial time optimal algorithm for satellite-switched time-division multiple access satellite communications with general switching modes. SIAM Journal on Discrete Mathematics, 4(1): 28–35, February 1991. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bonsma:2008:STM

[Bon08] Paul S. Bonsma. Spanning trees with many leaves in graphs with minimum degree three. SIAM Journal on Discrete Mathematics, 22(3): 920–937, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bona:2009:RZN

[Bón09] Miklós Bóna. Real zeros and normal distribution for statistics on Stirling permutations defined by Gessel and Stanley. SIAM Journal on Discrete Mathematics, 23(1): 401–406, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bonin:2010:CIS

[Bon10] Joseph E. Bonin. A construction of infinite sets of intertwines for pairs of matroids. SIAM Journal on Discrete Mathematics, 24(4): 1742–1752, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sidma/resource/
                                        1/sjdmec/v24/i4/p1742_s1.

Bonamy:2015:PGE

[Bon15]  Marthe Bonamy.  Planar

graphs with Δ ≥ 8 are (Δ+1)-edge-choosable. SIAM Journal on Discrete Mathematics, 29(3):1735–1763, ???? 2015. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Babai:1994:ESD

[BOP94] L. Babai, H. Oral, and K. T. Phelps. Eulerian self-dual codes. SIAM Journal on Discrete Mathematics, 7(2): 325–330, May 1994. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Borg:2010:CIF

[Bor10] Peter Borg. Cross-intersecting families of partial permutations. SIAM Journal on Discrete Mathematics, 24(2): 600–608, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Brightwell:2001:RRC

[BOS01] G. Brightwell, G. Oriolo, and F. B. Shepherd. Reserving resilient capacity in a network. SIAM Journal on Discrete Mathematics, 14(4):524–539, 2001. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/36818.

Borm:1992:PVC

[BOT92] Peter Borm, Guillermo Owen, and Stef Tijs. On the position value for communication situations. SIAM Journal on Discrete Mathematics, 5(3): 305–320, August 1992. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bouchet:1997:MCI

[Bou97] André Bouchet. Multimatroids. I. Coverings by independent sets. SIAM Journal on Discrete Mathematics, 10 (4):626–646, November 1997.

CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Boyd:1996:CCP

[Boy96]  E. Andrew Boyd.  On

the complexity of a cutting plane algorithm for solving combinatorial linear programs. SIAM Journal on Discrete Mathematics, 9(3): 365–376, August 1996. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Boykett:2001:CFP

[Boy01] Tim Boykett. Construction of Ferrero pairs of all possible orders. SIAM Journal on Discrete Mathematics, 14(3):283–285, 2001. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/33873.

Bermond:1989:ISP

[BP89] Jean-Claude Bermond and Claudine Peyrat. Induced subgraphs of the power of a cycle. SIAM Journal on Discrete Mathematics, 2(4):452– 455, November 1989. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bafna:1998:ST

[BP98] Vineet Bafna and Pavel A. Pevzner. Sorting by transpositions. SIAM Journal on Discrete Mathematics, 11(2): 224–240, May 1998. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Blanco:2009:PGB

[BP09] Víctor Blanco and Justo Puerto. Partial gröbner bases for multiobjective integer linear optimization. SIAM Journal on Discrete Mathematics, 23(2):571–595, ???? 2009. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Benvenuti:2010:SHL

[BP10] Daniel K. Benvenuti and Abraham P. Punnen. SC-Hamiltonicity and its linkages with strong Hamiltonicity of a graph. SIAM Journal on Discrete Mathematics, 23(4): 2035–2041, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Blanco:2012:AIP

[BP12] Víctor Blanco and Justo Puerto. An application of integer programming to the decomposition of numerical semigroups. SIAM Journal on Discrete Mathematics, 26 (3):1210–1237, ???? 2012. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Balogh:2013:TPC

[BP13] József Balogh and Cory Palmer. On the tree packing conjecture. SIAM Journal on Discrete Mathematics, 27(4): 1995–2006, ???? 2013. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Benevides:2015:MPT

[BP15] Fabricio Benevides and Michal Przykucki. Maximum percolation time in two-dimensional bootstrap percolation. SIAM Journal on Discrete Mathematics, 29(1):224–251, ???? 2015. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Braun:2016:PCB

[BP16] Gábor Braun and Sebastian Pokutta. A polyhedral characterization of border bases. SIAM Journal on Discrete Mathematics, 30(1): 239–265, ???? 2016. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Briceno:2017:SSM

[BP17] Raimundo Briceño and Ronnie Pavlov. Strong spatial mixing in homomorphism spaces. SIAM Journal on Discrete Mathematics, 31(3): 2110–2137, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bishnoi:2020:NRH

[BP20] Anurag Bishnoi and Valentina Pepe. Nonintersecting Ryser hypergraphs. SIAM Journal on Discrete Mathematics, 34 (1):230–240, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Booms-Peot:2022:CDS

[BPEY22] Caitlyn Booms-Peot, Daniel Erman, and Jay Yang. Characteristic dependence of syzygies of random monomial ideals. SIAM Journal on Discrete Mathematics, 36 (1):682–701, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1392474.

Bollobas:2013:CDP

[BPRS13] Béla Bollobás, David Pritchard, Thomas Rothvoss, and Alex Scott. Cover-decomposition and polychromatic numbers. SIAM Journal on Discrete Mathematics, 27(1): 240–256, ???? 2013. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Barik:2007:SCT

[BPS07] S. Barik, S. Pati, and B. K.

Sarma. The spectrum of the corona of two graphs. SIAM Journal on Discrete Mathematics, 21(1):47–56, ????

2007. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Boczkowski:2018:SMT

[BPS18] Lucas Boczkowski, Yuval Peres, and Perla Sousi. Sensitivity of mixing times in Eulerian digraphs. SIAM Journal on Discrete Mathematics, 32 (1):624–655, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Besomi:2019:DCE

[BPSS19] Guido Besomi, Matías Pavez-Signé, and Maya Stein. Degree conditions for embedding trees. SIAM Journal on Discrete Mathematics, 33(3): 1521–1555, ???? 2019. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Besomi:2020:MMD

[BPSS20] Guido Besomi, Matías Pavez-Signé, and Maya Stein. Maximum and minimum degree conditions for embedding trees. SIAM Journal on Discrete Mathematics, 34(4): 2108–2123, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Balister:2022:NIA

[BPST22] Paul Balister, Emil Powierski, Alex Scott, and Jane Tan. A note on infinite antichain density. SIAM Journal on Discrete Mathematics, 36(1):573–577, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M143025X.

Borie:1991:DDR

[BPT91] Richard B. Borie, R. Gary Parker, and Craig A. Tovey. Deterministic decomposition of recursive graph classes. SIAM Journal on Discrete Mathematics, 4(4):481–501, November 1991. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Barany:2010:PNS

[BPV10] Imre Bárány, Attila Pór, and Pavel Valtr. Paths with no small angles. SIAM Journal on Discrete Mathematics, 23(4):1655–1666, ???? 2010. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ball:2013:RBM

[BPWX13] Simeon Ball, Carles Padró, Zsuzsa Weiner, and Chaoping Xing. On the representability of the biuniform matroid. SIAM Journal on Discrete Mathematics, 27(3): 1482–1491, ???? 2013. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Burstein:2017:SCG

[BR17a] David Burstein and Jonathan Rubin. Sufficient conditions for graphicality of bidegree sequences. SIAM Journal on Discrete Mathematics, 31(1): 50–62, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Byrne:2017:CRM

[BR17b] Eimear Byrne and Alberto Ravagnani. Covering radius of matrix codes endowed with the rank metric. SIAM Journal on Discrete Mathematics, 31(2):927–944, ???? 2017. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Byrne:2019:AMT

[BR19] Eimear Byrne and Alberto Ravagnani. An Assmus– Mattson theorem for rank metric codes. SIAM Journal on Discrete Mathematics, 33(3):1242–1260, ???? 2019. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bukh:2020:OIT

[BR20] Boris Bukh and Oleksandr Rudenko. Order-isomorphic twins in permutations. SIAM Journal on Discrete Mathematics, 34(3):1620–1622, ???? 2020. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Brass:2005:UBD

[Bra05] Peter Brass. An upper bound for the d-dimensional analogue of Heilbronn’s triangle problem. SIAM Journal on Discrete Mathematics, 19(1):192–195, 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/43581.

Branden:2010:DCH

[Brä10] Petter Brändén. Discrete concavity and the half-plane property. SIAM Journal on Discrete Mathematics, 24(3): 921–933, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Buss:1989:VTB

[BRK89] Jonathan F. Buss, Arnold L.

Rosenberg, and Judson D. Knott. Vertex types in bookembeddings. SIAM Journal on Discrete Mathematics, 2 (2):156–175, May 1989. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Browder:2011:FNC

[Bro11] Jonathan Browder. Face numbers of certain Cohen– Macaulay flag complexes. SIAM Journal on Discrete Mathematics, 25(4):1768– 1777, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sidma/resource/
                                        1/sjdmec/v25/i4/p1768_s1.

Britz:2009:DM

[BRS09] T. Britz, G. Royle, and K. Shiromoto. Designs from matroids. SIAM Journal on Discrete Mathematics, 23(2): 1082–1099, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bruck:1990:HAP

[Bru90] Jehoshua Bruck. Harmonic analysis of polynomial threshold functions. SIAM Journal on Discrete Mathematics, 3(2):168–177, May 1990. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bertolazzi:1988:CPS

[BS88] P. Bertolazzi and A. Sassano. A class of polynomially solvable set-covering problems. SIAM Journal on Discrete Mathematics, 1(3): 306–316, August 1988. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bshouty:1990:GNB

[BS90] Nader H. Bshouty and Gadiel Seroussi. Generalizations of the normal basis theorem of finite fields. SIAM Journal on Discrete Mathematics, 3(3):

330–337, August 1990. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Brickell:1991:DCT

[BS91] E. F. Brickell and D. R. Stinson. The detection of cheaters in threshold schemes. SIAM Journal on Discrete Mathematics, 4(4):502–510, November 1991. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Brightwell:1993:RPG

[BS93] Graham R. Brightwell and Edward R. Scheinerman. Representations of planar graphs. SIAM Journal on Discrete Mathematics, 6(2): 214–229, May 1993. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bandelt:1995:SMR

[BS95a] Hans-Jürgen Bandelt and Michael Anthony Steel. Symmetric matrices representable by weighted trees over a cancellative Abelian monoid. SIAM Journal on Discrete Mathematics, 8(4):517–525, November 1995. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bokowski:1995:ASR

[BS95b] Jürgen Bokowski and Peter Schuchert. Altshuler’s sphere M_{963}^{9} revisited. SIAM Journal on Discrete Mathematics, 8(4):670–677, November 1995. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bartal:2009:UIS

[BS09] Yair Bartal and Leonard J. Schulman. Universal immersion spaces for edge-colored graphs and nearest-neighbor metrics. SIAM Journal on Discrete Mathematics, 23(2): 1110–1115, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Balogh:2010:AAF

[BS10a] József Balogh and Wojciech Samotij. Almost all C_{4}- free graphs have fewer than (1 − ε) ex(n, C_{4}) edges. SIAM Journal on Discrete Mathematics, 24(3):1011–1018, ???? 2010. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bruhn:2010:PAS

[BS10b] Henning Bruhn and Maya Stein. t-perfection is always strong for claw-free graphs. SIAM Journal on Discrete Mathematics, 24(3): 770–781, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Balogh:2015:BER

[BS15a] József Balogh and Gelasio Salazar. Book embeddings of regular graphs. SIAM Journal on Discrete Mathematics, 29(2):811–822, ???? 2015. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bordewich:2015:DPT

[BS15b] Magnus Bordewich and Charles Semple. Defining a phylogenetic tree with the minimum number of r-state characters. SIAM Journal on Discrete Mathematics, 29(2): 835–853, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Blanco:2016:LPF

[BS16a] M. Blanco and F. Santos.

Lattice 3-polytopes with few lattice points. SIAM Journal on Discrete Mathematics, 30 (2):669–686, ???? 2016. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Blanco:2016:LPS

[BS16b] M. Blanco and F. Santos.

Lattice 3-polytopes with six lattice points. SIAM Journal on Discrete Mathematics, 30 (2):687–717, ???? 2016. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bruhn:2016:CFP

[BS16c] Henning Bruhn and Oliver Schaudt. Claw-free t-perfect graphs can be recognized in polynomial time. SIAM Journal on Discrete Mathematics, 30(2):832–855, ???? 2016. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Byrka:2018:AAS

[BS18] Jaroslaw Byrka and Aravind Srinivasan. Approximation algorithms for stochastic and risk-averse optimization. SIAM Journal on Discrete Mathematics, 32(1): 44–63, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Biswas:2023:ISV

[BS23] Aniruddha Biswas and Palash Sarkar. Influence of a set of variables on a Boolean function. SIAM Journal on Discrete Mathematics, 37 (3):2148–2171, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1503531.

Babecki:2024:EUG

[BS24a] Catherine Babecki and David Shiroma. Eigenpolytope universality and graphical designs. SIAM Journal on Discrete Mathematics, 38(1):

947–964, March 2024. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Berczi:2024:EDB

[BS24b] Kristóf Bérczi and Tamás Schwarcz. Exchange distance of basis pairs in split matroids. SIAM Journal on Discrete Mathematics, 38(1): 132–147, January 2024. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Borcea:2015:PBB

[BSiT15] Ciprian Borcea, Ileana Streinu, and Shin ichi Tanigawa. Periodic body-and-bar frameworks. SIAM Journal on Discrete Mathematics, 29(1): 93–112, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ben-Shimon:2011:RHO

[BSKS11] Sonny Ben-Shimon, Michael Krivelevich, and Benny Sudakov. On the resilience of Hamiltonicity and optimal packing of Hamilton cycles in random graphs. SIAM Journal on Discrete Mathematics, 25(3):1176– 1193, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v25/i3/p1176_s1.

Berg:2014:CEF

[BSS14] Chris Berg, Franco Saliola, and Luis Serrano. Combinatorial expansions for families of noncommutative k-Schur functions. SIAM Journal on Discrete Mathematics, 28(3): 1074–1092, ???? 2014. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Benjamini:2014:ATG

[BST14] Itai Benjamini, Igor Shinkar, and Gilad Tsur. Acquaintance time of a graph. SIAM Journal on Discrete Mathematics, 28(2):767–785, ???? 2014. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Baste:2020:HMB

[BST20] Julien Baste, Ignasi Sau, and Dimitrios M. Thilikos. Hitting minors on bounded treewidth graphs. I. General upper bounds. SIAM Journal on Discrete Mathematics, 34(3):1623–1648, ???? 2020. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Brettell:2023:CM

[BST23] Nick Brettell, Charles Semple, and Gerry Toft. Cyclic matroids. SIAM Journal on Discrete Mathematics, 37 (2):1212–1236, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1468747.

Berczi:2021:LCT

[BSY21] Kristóf Bérczi, Tamás Schwarcz, and Yutaro Yamaguchi. List coloring of two matroids through reduction to partition matroids. SIAM Journal on Discrete Mathematics, 35(3):2192–2209, ???? 2021. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Brightwell:1993:ODC

[BT93] Graham Brightwell and William T.

Trotter. The order dimension of convex polytopes. SIAM Journal on Discrete Mathematics, 6(2):230–245, May 1993. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Blake:1996:TCD

[BT96] Ian F. Blake and Vahid Tarokh. On the trellis complexity of the densest lattice packings in R^{n}. SIAM Journal on Discrete Mathematics, 9(4):597–601, November 1996. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Brightwell:1997:ODP

[BT97] Graham R. Brightwell and William T. Trotter. The order dimension of planar maps. SIAM Journal on Discrete Mathematics, 10(4):

515–528, November 1997. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Baber:2014:SC

[BT14] Rahil Baber and John Talbot. A solution to the 2/3 conjecture. SIAM Journal on Discrete Mathematics, 28(2): 756–766, ???? 2014. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bisson:2018:CPR

[BT18] Gaetan Bisson and Mehdi Tibouchi. Constructing permutation rational functions from isogenies. SIAM Journal on Discrete Mathematics, 32(3): 1741–1749, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bhore:2022:ESS

[BT22] Sujoy Bhore and Csaba D. Tóth. Euclidean Steiner spanners: Light and sparse. SIAM Journal on Discrete Mathematics, 36(3):2411– 2444, ???? 2022. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:/

                                        /epubs.siam.org/doi/10.
                                        1137/22M1502707.

Behrisch:2009:CRI

[BTU09] Michael Behrisch, Anusch Taraz, and Michael Ueckerdt. Coloring random intersection graphs and complex networks. SIAM Journal on Discrete Mathematics, 23(1): 288–299, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Beimel:2008:CIW

[BTW08] Amos Beimel, Tamir Tassa, and Enav Weinreb. Characterizing ideal weighted threshold secret sharing. SIAM Journal on Discrete Mathematics, 22(1):360–397, ???? 2008. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bukh:2012:MKK

[Buk12] Boris Bukh. Multidimensional Kruskal–Katona theorem. SIAM Journal on Discrete Mathematics, 26(2): 548–554, ???? 2012. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bukh:2016:BEL

[Buk16] Boris Bukh. Bounds on equiangular lines and on related spherical codes. SIAM Journal on Discrete Mathematics, 30(1):549–554, ???? 2016. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bondarenko:2010:SDB

[BV10] Andriy V. Bondarenko and Maryna S. Viazovska. Spherical designs via Brouwer Fixed Point Theorem. SIAM Journal on Discrete Mathematics, 24(1):207–217, ???? 2010. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Braun:2020:HTC

[BV20] Benjamin Braun and Julianne Vega. Hajós-type constructions and neighborhood complexes. SIAM Journal on Discrete Mathematics, 34(4): 2424–2447, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Borgwardt:2021:CCT

[BV21] Steffen Borgwardt and Charles Viss. Constructing clustering transformations. SIAM Journal on Discrete Mathematics, 35(1):152–178, ???? 2021. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Balko:2022:ORN

[BV22] Martin Balko and Máté Vizer. On ordered Ramsey numbers of tripartite 3uniform hypergraphs. SIAM Journal on Discrete Mathematics, 36(1):214–228, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL

https://epubs.siam.org/
doi/10.1137/21M1404958.

Boys:2016:NOC

[BVdZ16] Thomas Boys, Claudiu Valculescu, and Frank de Zeeuw. On the number of ordinary conics. SIAM Journal on Discrete Mathematics, 30(3): 1644–1659, ???? 2016. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Barbe:2003:SCR

[BvH03] André Barbé and Fritz von Haeseler. Symmetric codes over rings. SIAM Journal on Discrete Mathematics, 17(2):254–263, 2003. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/41377.

Bhatnagar:2011:RCT

[BVVW11] Nayantara Bhatnagar, Juan Vera, Eric Vigoda, and Dror Weitz. Reconstruction for colorings on trees. SIAM Journal on Discrete Mathematics, 25(2):809–826, ???? 2011. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v25/i2/p809_s1.

Barnes:1988:NHP

[BVW88] Earl R. Barnes, Anthony Vannelli, and James Q. Walker. A new heuristic for partitioning the nodes of a graph. SIAM Journal on Discrete Mathematics, 1(3): 299–305, August 1988. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Brightwell:1992:CP

[BW92] Graham Brightwell and Colin Wright. The 1/3–2/3 conjecture for 5-thin posets. SIAM Journal on Discrete Mathematics, 5(4):467–474, November 1992. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bjorner:1999:CDG

[BW99] Anders Björner and Volkmar Welker. Complexes of directed graphs. SIAM Journal on Discrete Mathematics, 12 (4):413–424, November 1999. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/33872.

Borodin:2002:CCP

[BW02] Oleg V. Borodin and Douglas R. Woodall. Cyclic colorings of 3-polytopes with large maximum face size. SIAM Journal on Discrete Mathematics, 15(2):143–154, 2002. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/30248.

Brightwell:2009:SP

[BW09] Graham R. Brightwell and Peter Winkler. Submodular percolation. SIAM Journal on Discrete Mathematics, 23(3):1149–1178, ???? 2009. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bruhn:2008:HCP

[BY08] Henning Bruhn and Xingxing Yu. Hamilton cycles in planar locally finite graphs. SIAM Journal on Discrete Mathematics, 22(4):1381–1392, ???? 2008. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bar-Yehuda:2010:APC

[BYFMR10] Reuven Bar-Yehuda, Guy Flysher, Julián Mestre, and Dror Rawitz. Approximation of partial capacitated vertex cover. SIAM Journal on Discrete Mathematics, 24(4): 1441–1469, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bar-Yehuda:2010:ENT

[BYHR10] Reuven Bar-Yehuda, Danny Hermelin, and Dror Rawitz. An extension of the Nemhauser– Trotter Theorem to generalized vertex cover with applications. SIAM Journal on Discrete Mathematics, 24(1): 287–300, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bar-Yehuda:2018:GHB

[BYKKR18] Reuven Bar-Yehuda, Erez Kantor, Shay Kutten, and Dror Rawitz. Growing halfballs: Minimizing storage and communication costs in content delivery networks. SIAM Journal on Discrete Mathematics, 32(3):1903–1921, ???? 2018. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bar-Yehuda:2005:EBP

[BYR05] Reuven Bar-Yehuda and Dror Rawitz. On the equivalence between the primaldual schema and the local ratio technique. SIAM Journal on Discrete Mathematics, 19(3):762–797, 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/62538.

Barg:2004:EEE

[BZ04] Alexander Barg and Gilles Zémor. Error exponents of expander codes under linearcomplexity decoding. SIAM Journal on Discrete Mathematics, 17(3):426–445, 2004. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/40379.

Bonsma:2011:AAF

[BZ11] Paul Bonsma and Florian Zickfeld. A 3/2approximation algorithm for finding spanning trees with many leaves in cubic graphs. SIAM Journal on Discrete Mathematics, 25(4):1652– 1666, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v25/i4/p1652_s1.

Butti:2020:SBC

[BZ20] Silvia Butti and Stanislav Zivný. Sparsification of binary CSPs. SIAM Journal on Discrete Mathematics, 34 (1):825–842, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bohman:2023:MRN

[BZ23] Tom Bohman and Emily Zhu.

On multicolor Ramsey numbers of triple system paths of length 3. SIAM Journal on Discrete Mathematics, 37 (3):1419–1435, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1518402.

Cai:1993:CEP

[Cai93] Jiazhen Cai. Counting embeddings of planar graphs using DFS trees. SIAM Journal on Discrete Mathematics, 6(3):335–352, August 1993. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cairns:2018:SHP

[Cai18] Hannah Cairns. Some halting problems for Abelian sandpiles are undecidable in dimension three. SIAM Journal on Discrete Mathematics, 32(4):2636–2666, ???? 2018. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Calinescu:2013:AMP

[Cal13] G. Calinescu. Approximate min-power strong connectivity. SIAM Journal on Discrete Mathematics, 27(3): 1527–1543, ???? 2013. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Canfield:1993:MPL

[Can93] E. Rodney Canfield. Matchings in the partition lattice. SIAM Journal on Discrete Mathematics, 6(1):100– 109, February 1993. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Caprara:1999:SPR

[Cap99] Alberto Caprara. Sorting permutations by reversals and Eulerian cycle decompositions. SIAM Journal on Discrete Mathematics, 12 (1):91–110, February 1999.

CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31994.

Capalbo:2003:ECL

[Cap03] M. Capalbo. An explicit construction of lower-diameter cubic graphs. SIAM Journal on Discrete Mathematics, 16(4):630–634, 2003. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/33026.

Carlson:1988:ULB

[Car88] David A. Carlson. Upper and lower bounds on timespace tradeoffs for computations with embedded fast Fourier transforms. SIAM Journal on Discrete Mathematics, 1(1):22–37, February 1988. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Carlet:1994:DCD

[Car94] Claude Carlet. The divisors of x^{2^{m}} + x of constant derivatives and degree 2^{m−2}. SIAM Journal on Discrete Mathematics, 7(2):238–244, May 1994. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Caragiannis:2009:WMW

[Car09] Ioannis Caragiannis. Wavelength management in WDM rings to maximize the number of connections. SIAM Journal on Discrete Mathematics, 23 (2):959–978, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cai:1995:TS

[CC95] Leizhen Cai and Derek G. Corneil. Tree spanners. SIAM Journal on Discrete Mathematics, 8(3):359–387, August 1995. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cheon:2003:NBF

[CC03] Jung Hee Cheon and Seongtaek Chee. Nonlinearity of Boolean functions and hyperelliptic curves. SIAM Journal on Discrete Mathematics, 16(3):354–365, 2003. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/36625.

Chlebik:2007:CCO

[CC07] Miroslav Chlebík and Janka Chlebíková. The complexity of combinatorial optimization problems on d-dimensional boxes. SIAM Journal on Discrete Mathematics, 21(1): 158–169, ???? 2007. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Canteaut:2000:WDC

[CCD00] Anne Canteaut, Pascale Charpin, and Hans Dobbertin. Weight divisibility of cyclic codes, highly nonlinear functions on F_{2^{m}}, and crosscorrelation of maximumlength sequences. SIAM Journal on Discrete Mathematics, 13(1):105–138, 2000. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/35005.

Cao:2023:GCC

[CCD^{+}23] Yan Cao, Guantao Chen, Guoli Ding, Guangming Jing, and Wenan Zang. On Gupta’s codensity conjecture. SIAM Journal on Discrete Mathematics, 37(3):1666–1673, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL

https://epubs.siam.org/
doi/10.1137/22M1514003.

Cheriyan:2023:IAA

[CCDZ23] J. Cheriyan, R. Cummings, J. Dippel, and J. Zhu. An improved approximation algorithm for the matching augmentation problem. SIAM Journal on Discrete Mathematics, 37(1):163–190, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL

https://epubs.siam.org/
doi/10.1137/21M1453505.

Coffman:2000:PDI

[CCG^{+}00] E. G. Coffman, Jr., C. Courcoubetis, M. R. Garey, D. S. Johnson, P. W. Shor, R. R. Weber, and M. Yannakakis. Bin packing with discrete item sizes, Part I: Perfect packing theorems and the average case behavior of optimal packings. SIAM Journal on Discrete Mathematics, 13(3):384–402, 2000. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/32593.

Chu:2005:RCR

[CCG05] Wensong Chu, Charles J. Colbourn, and Solomon W. Golomb. A recursive construction for regular difference triangle sets. SIAM Journal on Discrete Mathematics, 18(4):741–748, 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/43676.

Chandrasekaran:2011:CIS

[CCG^{+}11] Venkat Chandrasekaran, Misha Chertkov, David Gamarnik, Devavrat Shah, and Jinwoo Shin. Counting independent sets using the Bethe approximation. SIAM Journal on Discrete Mathematics, 25(2): 1012–1034, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v25/i2/p1012_s1.

Chandrasekaran:2018:LTL

[CCGG18] Karthekeyan Chandrasekaran, Mahdi Cheraghchi, Venkata Gandikota, and Elena Grigorescu. Local testing of lattices. SIAM Journal on Discrete Mathematics, 32(2): 1265–1295, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cabello:2014:CSE

[CCH14] Sergio Cabello, Markus Chimani, and Petr Hlinený. Computing the stretch of an embedded graph. SIAM Journal on Discrete Mathematics, 28(3):1391–1401, ???? 2014. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chakraborti:2021:MNE

[CCH21] Debsoumya Chakraborti, Da Qi Chen, and Mihir Hasabnis. Minimizing the number of edges in K_{(s,t)}-saturated bipartite graphs. SIAM Journal on Discrete Mathematics, 35(2):1165–1181, ???? 2021. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chee:2013:SCA

[CCHZ13] Yeow Meng Chee, Charles J.

Colbourn, Daniel Horsley, and Junling Zhou. Sequence covering arrays. SIAM Journal on Discrete Mathematics, 27(4):1844–1861, ???? 2013. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cohen:2017:BVC

[CCJ^{+}17] David A. Cohen, Martin C. Cooper, Peter G. Jeavons, Andrei Krokhin, Robert Powell, and Stanislav Zivný. Binarisation for valued constraint satisfaction problems. SIAM Journal on Discrete Mathematics, 31(4):2279– 2300, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cao:2022:OGS

[CCJS22] Yan Cao, Guantao Chen, Guangming Jing, and Songling Shan. The overfullness of graphs with small minimum degree and large maximum degree. SIAM Journal on Discrete Mathematics, 36 (3):2258–2270, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1432776.

Conforti:2006:OHR

[CCL^{+}06] Michele Conforti, Gérard Cornuéjols, Xinming Liu, Kristina Vuskovic, and Giacomo Zambelli. Odd hole recognition in graphs of bounded clique size. SIAM Journal on Discrete Mathematics, 20(1):42–48, January 2006. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Carlson:1995:DEP

[CCM95] Bradley S. Carlson, C. Y. Roger Chen, and Dikran S. Meliksetian. Dual Eulerian properties of plane multigraphs. SIAM Journal on Discrete Mathematics, 8(1): 33–50, February 1995. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cascudo:2015:SSN

[CCM^{+}15] Ignacio Cascudo, Ronald Cramer, Diego Mirandola, Carles Padró, and Chaoping Xing. On secret sharing with nonlinear product reconstruction. SIAM Journal on Discrete Mathematics, 29(2): 1114–1131, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chalopin:2023:SCS

[CCM^{+}23] Jérémie Chalopin, Victor Chepoi, Fionn McInerney, Sébastien Ratel, and Yann Vaxès. Sample compression schemes for balls in graphs. SIAM Journal on Discrete Mathematics, 37(4): 2585–2616, November 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Carmona:2024:MRS

[CCNP24] Mikhael Carmona, Victor Chepoi, Guyslain Naves, and Pascal Préa. Modules in Robinson spaces. SIAM Journal on Discrete Mathematics, 38(1):190–224, January 2024. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chalopin:2011:CRG

[CCNV11] Jérémie Chalopin, Victor Chepoi, Nicolas Nisse, and Yann Vaxès. Cop and robber games when the robber can hide and ride. SIAM Journal on Discrete Mathematics, 25 (1):333–359, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sidma/resource/
                                        1/sjdmec/v25/i1/p333_s1.

Cada:2013:CCC

[CCO^{+}13] Roman Cada, Shuya Chiba, Kenta Ozeki, Petr Vrána, and Kiyoshi Yoshimoto. 4, 5 is not coverable: a counterexample to a conjecture of Kaiser and Skrekovski. SIAM Journal on Discrete Mathematics, 27(1): 141–144, ???? 2013. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cada:2015:RBT

[CCO^{+}15] Roman Cada, Shuya Chiba, Kenta Ozeki, Petr Vrána, and Kiyoshi Yoshimoto. A relationship between Thomassen’s Conjecture and Bondy’s Conjecture. SIAM Journal on Discrete Mathematics, 29(1): 26–35, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cada:2017:DES

[CCOY17] Roman Cada, Shuya Chiba, Kenta Ozeki, and Kiyoshi Yoshimoto. On dominating even subgraphs in cubic graphs. SIAM Journal on Discrete Mathematics, 31(2): 890–907, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chalopin:2014:CRG

[CCPP14] Jérémie Chalopin, Victor Chepoi, Panos Papasoglu, and Timothée Pecatte. Cop and robber game and hyperbolicity. SIAM Journal on Discrete Mathematics, 28(4): 1987–2007, ???? 2014. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Campos:2023:TAS

[CCSW23] Marcelo Campos, Matthew Coulson, Oriol Serra, and Maximilian Wötzel. The typical approximate structure of sets with bounded sumset. SIAM Journal on Discrete Mathematics, 37 (3):1386–1418, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1511801.

Callaghan:2005:BFM

[CCT05] Joseph Callaghan, John J. Chew III, and Stephen M. Tanny. On the behavior of a family of meta-Fibonacci sequences. SIAM Journal on Discrete Mathematics, 18(4):794–824, 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/42139.

Cai:2010:SRG

[CCVZ10] Leizhen Cai, Yongxi Cheng, Elad Verbin, and Yuan Zhou. Surviving rates of graphs with bounded treewidth for the firefighter problem. SIAM Journal on Discrete Mathematics, 24(4):1322–1335, ???? 2010. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chen:2012:TDI

[CCZ12] Xujin Chen, Zhibin Chen, and Wenan Zang. Total dual integrality in some facility location problems. SIAM Journal on Discrete Mathematics, 26(3):1022–1030, ???? 2012. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Calderbank:1993:ECC

[CD93] A. R. Calderbank and P. Delsarte. On error-correcting codes and invariant linear forms. SIAM Journal on Discrete Mathematics, 6(1): 1–23, February 1993. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Czygrinow:2011:NBG

[CD11] Andrzej Czygrinow and Louis DeBiasio. A note on bipartite graph tiling. SIAM Journal on Discrete Mathematics, 25(4):1477–1489, ???? 2011. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v25/i4/p1477_s1.

Coudert:2014:RFH

[CD14] David Coudert and Guillaume Ducoffe. Recognition of C_{4}-free and 1/2-hyperbolic graphs. SIAM Journal on Discrete Mathematics, 28(3): 1601–1617, ???? 2014. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Conforti:2016:MFC

[CD16] Michele Conforti and Marco Di Summa. Maximal S-free convex sets and the Helly number. SIAM Journal on Discrete Mathematics, 30(4): 2206–2216, ???? 2016. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Coudert:2018:RDC

[CD18] David Coudert and Guillaume Ducoffe. Revisiting decomposition by clique separators. SIAM Journal on Discrete Mathematics, 32(1): 682–694, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Conforti:2015:RCG

[CDD^{+}15] Michele Conforti, Alberto Del Pia, Marco Di Summa, Yuri Faenza, and Roland Grappe. Reverse Chvátal– Gomory rank. SIAM Journal on Discrete Mathematics, 29 (1):166–181, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Comellas:2008:MMS

[CDF08] F. Comellas, C. Dalfó, and M. A. Fiol. Multidimensional Manhattan street networks. SIAM Journal on Discrete Mathematics, 22(4): 1428–1447, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cooper:2018:DVP

[CDFR18] Colin Cooper, Martin Dyer, Alan Frieze, and Nicolás Rivera. Discordant voting processes on finite graphs.

SIAM Journal on Discrete Mathematics, 32(4):2398– 2420, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Coleman:2022:CES

[CDGO22] Michaela Coleman, Anton Dochtermann, Nathan Geist, and Suho Oh. Completing and extending shellings of vertex decomposable complexes. SIAM Journal on Discrete Mathematics, 36 (2):1291–1305, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1445119.

Chen:2004:IRN

[CDH^{+}04] Guantao Chen, George Davis, Frank Hall, Zhongshan Li, Kinnari Patel, and Michael Stewart. An interlacing result on normalized Laplacians. SIAM Journal on Discrete Mathematics, 18 (2):353–361, 2004. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/43858.

Carmesin:2014:BCI

[CDHH14] J. Carmesin, R. Diestel, M. Hamann, and F. Hundertmark. k-blocks: a connectivity invariant for graphs. SIAM Journal on Discrete Mathematics, 28(4): 1876–1891, ???? 2014. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Corneil:2016:PGS

[CDHK16] Derek G. Corneil, Jérémie Dusart, Michel Habib, and Ekkehard Köhler. On the power of graph searching for cocomparability graphs. SIAM Journal on Discrete Mathematics, 30(1):569–591, ???? 2016. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chen:2012:MWS

[CDHZ12] Xujin Chen, Guoli Ding, Xiaodong Hu, and Wenan Zang. The maximum-weight stable matching problem: Duality and efficiency. SIAM Journal on Discrete Mathematics, 26(3):1346–1360, ???? 2012. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Czygrinow:2010:FBG

[CDK10] Andrzej Czygrinow, Louis DeBiasio, and H. A. Kierstead. 2-factors of bipartite graphs with asymmetric minimum degrees. SIAM Journal on Discrete Mathematics, 24 (2):486–504, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chlamtac:2018:DSP

[CDK^{+}18] Eden Chlamtác, Michael Dinitz, Christian Konrad, Guy Kortsarz, and George Rabanca. The densest k-subhypergraph problem. SIAM Journal on Discrete Mathematics, 32(2):1458– 1477, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cooley:2021:LIM

[CDKS21] Oliver Cooley, Nemanja Draganić, Mihyun Kang, and Benny Sudakov. Large induced matchings in random graphs. SIAM Journal on Discrete Mathematics, 35(1): 267–280, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cera:2000:SGT

[CDM00] M. Cera, A. Diánez, and A. Márquez. The size of a graph without topological complete subgraphs. SIAM Journal on Discrete Mathematics, 13(3):295–301, 2000. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31594.

Cera:2004:EGT

[CDM04] M. Cera, A. Diánez, and A. Márquez. Extremal graphs without topological complete subgraphs. SIAM Journal on Discrete Mathematics, 18(2):388–396, 2004. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/37867.

Chapuy:2014:PTW

[CDM^{+}14] Guillaume Chapuy, Matt DeVos, Jessica McDonald, Bojan Mohar, and Diego Scheide. Packing triangles in weighted graphs. SIAM Journal on Discrete Mathematics, 28(1):226–239, ???? 2014. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chun:2016:UCM

[CDMO16] Carolyn Chun, Guoli Ding, Dillon Mayhew, and James Oxley. Unavoidable connected matroids retaining a specified minor. SIAM Journal on Discrete Mathematics, 30(3):1590–1606, ???? 2016. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Charbit:2012:LTS

[CdMR12] Pierre Charbit, Fabien de Montgolfier, and Mathieu Raffinot. Linear time split decomposition revisited. SIAM Journal on Discrete Mathematics, 26 (2):499–514, ???? 2012. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chen:2021:SFP

[CDMT21] Ke Chen, Adrian Dumitrescu, Wolfgang Mulzer, and Csaba D. Tóth. On the stretch factor of polygonal chains. SIAM Journal on Discrete Mathematics, 35(3):1592–1614, ???? 2021. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Coudert:2016:ATU

[CDN16] David Coudert, Guillaume Ducoffe, and Nicolas Nisse. To approximate treewidth, use treelength! SIAM Journal on Discrete Mathematics, 30(3):1424–1436, ???? 2016. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chlebus:1994:SMC

[CDP94] Bogdan S. Chlebus, Krzysztof Diks, and Andrzej Pelc. Sorting on a mesh-connected computer with delaying links. SIAM Journal on Discrete Mathematics, 7(1):119–132, February 1994. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Calinescu:2008:RGG

[CDP08] Gruia Călinescu, Adrian Dumitrescu, and János Pach. Reconfigurations in graphs and grids. SIAM Journal on Discrete Mathematics, 22(1): 124–138, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cassaigne:2016:NBS

[CDR16] Julien Cassaigne, Eric Duchêne, and Michel Rigo. Nonhomogeneous Beatty sequences leading to invariant games. SIAM Journal on Discrete Mathematics, 30(3):1798– 1829, ???? 2016. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cioaba:2023:SES

[CDT23] Sebastian Cioaba, Dheer Noal Desai, and Michael Tait. A spectral Erdős–Sós theorem. SIAM Journal on Discrete Mathematics, 37 (3):2228–2239, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M150650X.

Cheng:2002:FIA

[CdV02] Eddie Cheng and Sven de Vries. On the facetinducing antiweb-wheel inequalities for stable set polytopes. SIAM Journal on Discrete Mathematics, 15 (4):470–487, 2002. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/39105.

Chakrabarty:2010:RSP

[CDV10] Deeparnab Chakrabarty, Nikhil R.

Devanur, and Vijay V. Vazirani. Rationality and strongly polynomial solvability of Eisenberg–Gale markets with two agents. SIAM Journal on Discrete Mathematics, 24(3): 1117–1136, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cabello:2011:FCT

[CdVL11] Sergio Cabello, Éric Colin de Verdière, and Francis Lazarus. Finding cycles with topological properties in embedded graphs. SIAM Journal on Discrete Mathematics, 25(4):1600– 1614, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v25/i4/p1600_s1.

Conforti:2007:MSF

[CDW07] Michele Conforti, Marco Di Summa, and Laurence A. Wolsey. The mixing set with flows. SIAM Journal on Discrete Mathematics, 21(2): 396–407, ???? 2007. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chang:2017:TWC

[CDWZ17] Gerard Jennhwa Chang, Guan-Huei Duh, Tsai-Lien Wong, and Xuding Zhu. Total weight choosability of trees. SIAM Journal on Discrete Mathematics, 31(2): 669–686, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Coppersmith:2006:SSP

[CE06] Don Coppersmith and Michael Elkin. Sparse sourcewise and pairwise distance preservers. SIAM Journal on Discrete Mathematics, 20(2): 463–501, January 2006. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cheng:2024:BQC

[CEFY24] Yen-Jen Cheng, Sen-Peng Eu, Tung-Shan Fu, and Jyun-Cheng Yao. On q-counting of noncrossing chains and parking functions. SIAM Journal on Discrete Mathematics, 38(1):917–946, March 2024. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cameron:2008:CLP

[CEHS08] Kathie Cameron, Elaine M. Eschen, Chính T. Hoàng, and R. Sritharan. The complexity of the list partition problem for graphs. SIAM Journal on Discrete Mathematics, 21 (4):900–929, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cooper:2013:CRW

[CEOR13] Colin Cooper, Robert Elsässer, Hirotaka Ono, and Tomasz Radzik. Coalescing random walks and Voting on connected graphs. SIAM Journal on Discrete Mathematics, 27(4):1748–1758, ???? 2013. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chen:2015:PTS

[CEOT15] Guantao Chen, Hikoe Enomoto, Kenta Ozeki, and Shoichi Tsuchiya. Plane triangulations without a spanning Halin subgraph: Counterexamples to the Lovász– Plummer conjecture on Halin graphs. SIAM Journal on Discrete Mathematics, 29(3): 1423–1426, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chen:2017:PTS

[CEOT17] Guantao Chen, Hikoe Enomoto, Kenta Ozeki, and Shoichi Tsuchiya. Plane triangulations without a spanning Halin subgraph II. SIAM Journal on Discrete Mathematics, 31(4):2429–2439, ???? 2017. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chekuri:2018:CCR

[CEP18] Chandra Chekuri, Alina Ene, and Marcin Pilipczuk. Constant congestion routing of symmetric demands in planar directed graphs. SIAM Journal on Discrete Mathematics, 32(3):2134–2160, ???? 2018. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Caro:1998:LSW

[CER98] Yair Caro, M. N. Ellingham, and J. E. Ramey. Local structure when all maximal independent sets have equal weight. SIAM Journal on Discrete Mathematics, 11 (4):644–654, November 1998. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Calkin:1997:NRB

[CET97] Neil J. Calkin, Paul Erdős, and Craig A. Tovey. New Ramsey bounds from cyclic graphs of prime order. SIAM Journal on Discrete Mathematics, 10(3):381–387, August 1997. CODEN SJD-MEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cooper:2005:CTR

[CF05] Colin Cooper and Alan Frieze. The cover time of random regular graphs. SIAM Journal on Discrete Mathematics, 18(4):728–740, 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/42847.

Chebolu:2008:HCR

[CF08] Prasad Chebolu and Alan Frieze. Hamilton cycles in random lifts of directed graphs. SIAM Journal on Discrete Mathematics, 22(2): 520–540, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Como:2009:ASM

[CF09] Giacomo Como and Fabio Fagnani. Average spectra and minimum distances of lowdensity parity-check codes over Abelian groups. SIAM Journal on Discrete Mathematics, 23(1):19–53, ???? 2009. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cooper:2010:RWL

[CF10] Colin Cooper and Alan Frieze. Random walks with look-ahead in scale-free random graphs. SIAM Journal on Discrete Mathematics, 24(3):1162–1176, ???? 2010. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cooper:2016:VSV

[CF16] Colin Cooper and Alan Frieze. Vacant sets and vacant nets: Component structures induced by a random walk. SIAM Journal on Discrete Mathematics, 30(1): 166–205, ???? 2016. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chen:2017:PEI

[CF17] Sandy H. L. Chen and Amy M. Fu. A 4n-point elliptic interpolation formula and its applications. SIAM Journal on Discrete Mathematics, 31(2):758–765, ???? 2017. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cooper:2022:RVE

[CF22] Colin Cooper and Alan Frieze. Rank of the vertexedge incidence matrix of rout hypergraphs. SIAM Journal on Discrete Mathematics, 36(3):2238–2257, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1467572.

Cooley:2020:SRH

[CFDK20] Oliver Cooley, Wenjie Fang, Nicola Del Giudice, and Mihyun Kang. Subcritical random hypergraphs, highorder components, and hypertrees. SIAM Journal on Discrete Mathematics, 34(4): 2033–2062, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Colbourn:2009:MSA

[CFG^{+}09] Charles J. Colbourn, Hung-Lin Fu, Gennian Ge, Alan C. H. Ling, and Hui-Chuan Lu. Minimizing SONET ADMs in unidirectional WDM rings with grooming ratio seven. SIAM Journal on Discrete Mathematics, 23(1):109–122, ???? 2009. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Clemens:2015:BST

[CFG^{+}15a] Dennis Clemens, Asaf Ferber, Roman Glebov, Dan Hefetz, and Anita Liebenau. Building spanning trees quickly in maker–breaker games. SIAM Journal on Discrete Mathematics, 29(3):1683–1705, ???? 2015. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Conlon:2015:DVS

[CFG^{+}15b] David Conlon, Jacob Fox, William Gasarch, David G. Harris, Douglas Ulrich, and Samuel Zbarsky. Distinct volume subsets. SIAM Journal on Discrete Mathematics, 29 (1):472–480, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chaplick:2021:KGH

[CFG^{+}21] Steven Chaplick, Fedor V.

Fomin, Petr A. Golovach, Dusan Knop, and Peter Zeman. Kernelization of graph Hamiltonicity: Proper Hgraphs. SIAM Journal on Discrete Mathematics, 35(2): 840–892, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chung:1988:FCN

[CFGG88] F. R. K. Chung, Z. Füredi, M. R. Garey, and R. L. Graham. On the fractional covering number of hypergraphs. SIAM Journal on Discrete Mathematics, 1(1): 45–49, February 1988. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chen:2006:CEH

[CFGJ06] Guantao Chen, Ralph J. Faudree, Ronald J. Gould, and Michael S. Jacobson. Cycle extendability of Hamiltonian interval graphs. SIAM Journal on Discrete Mathematics, 20(3):682–689, January 2006. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Colbourn:2016:DSS

[CFH16] Charles J. Colbourn, Bingli Fan, and Daniel Horsley. Disjoint spread systems and fault location. SIAM Journal on Discrete Mathematics, 30(4): 2011–2026, ???? 2016. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chakraborti:2023:CHC

[CFH23] Debsoumya Chakraborti, Alan M.

Frieze, and Mihir Hasabnis. Colorful Hamilton cycles in random graphs. SIAM Journal on Discrete Mathematics, 37(1):51–64, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1403291.

Chen:2011:LBP

[CFJ11] William Y. C. Chen, Neil J. Y. Fan, and Jeffrey Y. T. Jia. Labeled ballot paths and the Springer numbers. SIAM Journal on Discrete Mathematics, 25(4):1530– 1546, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v25/i4/p1530_s1.

Cooper:2010:HCR

[CFK10] Colin Cooper, Alan Frieze, and Michael Krivelevich. Hamilton cycles in random graphs with a fixed degree sequence. SIAM Journal on Discrete Mathematics, 24(2): 558–569, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cohen:2019:IP

[CFK19] Gérard Cohen, Emanuela Fachini, and János Körner. Interlocked permutations.

SIAM Journal on Discrete Mathematics, 33(3):1662– 1668, ???? 2019. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chan:2017:PTU

[CFKK17] Chun Lam Chan, Winston Fernandes, Navin Kashyap, and Manjunath Krishnapur. Phase transitions for the uniform distribution in the pattern maximum likelihood problem and its Bethe approximation. SIAM Journal on Discrete Mathematics, 31 (1):597–631, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cseh:2022:UPM

[CFKP22] Ágnes Cseh, Yuri Faenza, Telikepalli Kavitha, and Vladlena Powers. Understanding popular matchings via stable matchings. SIAM Journal on Discrete Mathematics, 36(1):188–213, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL

https://epubs.siam.org/
doi/10.1137/19M124770X.

Chen:2019:DOC

[CFLZ19] Guantao Chen, Xing Feng, Fuliang Lu, and Lianzhu Zhang. Disjoint odd cycles in cubic solid bricks. SIAM Journal on Discrete Mathematics, 33(1):393–397, ????

2019. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chung:1994:UBD

[CFM94] F. R. K. Chung, V. Faber, and Thomas A. Manteuffel. An upper bound on the diameter of a graph from eigenvalues associated with its Laplacian. SIAM Journal on Discrete Mathematics, 7(3):443– 457, August 1994. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chavez:2009:GRM

[CFM^{+}09] María Jose Chávez, Gašper Fijavž, Alberto Márquez, Atsuhiro Nakamoto, and Esperanza Suárez. Geometric realization of Möbius triangulations. SIAM Journal on Discrete Mathematics, 23(1): 221–232, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cooper:2011:MBC

[CFP11] Joshua Cooper, Stephen Fenner, and Semmy Purewal. Monochromatic boxes in colored grids. SIAM Journal on Discrete Mathematics, 25(3): 1054–1068, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v25/i3/p1054_s1.

Conforti:2016:CDF

[CFP16] Michele Conforti, Samuel Fiorini, and Kanstantsin Pashkovich. Cut dominants and forbidden minors. SIAM Journal on Discrete Mathematics, 30(3):1571–1589, ???? 2016. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cooper:2019:CTD

[CFP19] Colin Cooper, Alan M. Frieze, and Wesley Pegden. On the cover time of dense graphs. SIAM Journal on Discrete Mathematics, 33(3): 1374–1389, ???? 2019. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cooper:2010:MRW

[CFR10] Colin Cooper, Alan Frieze, and Tomasz Radzik. Multiple random walks in random regular graphs. SIAM Journal on Discrete Mathematics, 23(4):1738–1761, ???? 2010. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Coppersmith:1996:RWR

[CFS96] Don Coppersmith, Uriel Feige, and James Shearer. Random walks on regular and irregular graphs. SIAM Journal on Discrete Mathematics, 9(2):301–308, May 1996. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chappell:2017:TDP

[CFS17] Thomas Chappell, Tobias Friedl, and Raman Sanyal. Two double poset polytopes. SIAM Journal on Discrete Mathematics, 31(4): 2378–2413, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Conlon:2020:BVT

[CFS20] David Conlon, Jacob Fox, and Benny Sudakov. Books versus triangles at the extremal density. SIAM Journal on Discrete Mathematics, 34 (1):385–398, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chang:1993:AAN

[CFT93] Gerard J. Chang, Martin Farber, and Zsolt Tuza. Algorithmic aspects of neighborhood numbers. SIAM Journal on Discrete Mathematics, 6(1): 24–29, February 1993. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cornuejols:2002:IBC

[CG02] Gérard Cornuéjols and Bertrand Guenin. Ideal binary clutters, connectivity, and a conjecture of Seymour. SIAM Journal on Discrete Mathematics, 15(3):329–352, 2002. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/37138.

Conforti:2007:POC

[CG07] Michele Conforti and Bert Gerards. Packing odd circuits. SIAM Journal on Discrete Mathematics, 21(2): 273–302, ???? 2007. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Connelly:2017:PST

[CG17] Robert Connelly and Steven J. Gortler. Prestress stability of triangulated convex polytopes and universal secondorder rigidity. SIAM Journal on Discrete Mathematics, 31(4):2735–2753, ???? 2017. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chor:1988:ISP

[CGG88] Benny Chor and Mihály Geréb-Graus. On the influence of single participant in coin flipping schemes. SIAM Journal on Discrete Mathematics, 1(4):411–415, November 1988. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chen:2016:ISP

[CGG^{+}16] William Y. C. Chen, Alan J. X. Guo, Peter L. Guo, Harry H. Y. Huang, and Thomas Y. H. Liu. sinversion sequences and P - partitions of type B. SIAM Journal on Discrete Mathematics, 30(3):1632–1643, ???? 2016. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chandrasekaran:2017:DOC

[CGG17] Karthekeyan Chandrasekaran, Venkata Gandikota, and Elena Grigorescu. Deciding orthogonality in Construction-A lattices. SIAM Journal on Discrete Mathematics, 31(2): 1244–1262, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cortes:2010:TTN

[CGH^{+}10] C. Cortés, C. I. Grima, F. Hurtado, A. Márquez, F. Santos, and J. Valenzuela. Transforming triangulations on nonplanar surfaces. SIAM Journal on Discrete Mathematics, 24(3): 821–840, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chen:2015:DCC

[CGH^{+}15] Guantao Chen, Ronald J.

Gould, Kazuhide Hirohata, Katsuhiro Ota, and Songling Shan. Disjoint chorded cycles of the same length. SIAM Journal on Discrete Mathematics, 29(2):1030–1041, ???? 2015. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cooley:2021:LPR

[CGH^{+}21] Oliver Cooley, Frederik Garbe, Eng Keat Hng, Mihyun Kang, Nicolás Sanhueza-Matamala, and Julian Zalla. Longest paths in random hypergraphs. SIAM Journal on Discrete Mathematics, 35(4): 2430–2458, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cheng:2023:ELA

[CGHL23] Kuan Cheng, Venkatesan Guruswami, Bernhard Haeupler, and Xin Li. Efficient linear and affine codes for correcting Insertions/Deletions. SIAM Journal on Discrete Mathematics, 37(2):748–778, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL

                                        https://epubs.siam.org/
                                        doi/10.1137/21M142798X.

Chang:2022:MGS

[CGHW22] Yulin Chang, Huifen Ge, Jie Han, and Guanghui Wang. Matching of given sizes in hypergraphs. SIAM Journal on Discrete Mathematics, 36(3):2323–2338, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1430327.

Chung:1994:ECD

[CGK94] F. R. K. Chung, Wayne Goddard, and Daniel J. Kleitman. Even cycles in directed graphs. SIAM Journal on Discrete Mathematics, 7(3): 474–483, August 1994. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chee:2019:DEC

[CGK^{+}19] Yeow Meng Chee, Fei Gao, Han Mao Kiah, Alan Chi Hung Ling, Hui Zhang, and Xiande Zhang. Decompositions of edge-colored digraphs: a new technique in the construction of constant-weight codes and related families. SIAM Journal on Discrete Mathematics, 33(1):209–229, ???? 2019. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chee:2010:SSP

[CGL10] Yeow Meng Chee, Gennian Ge, and Alan C. H. Ling. Spectrum of sizes for perfect deletion-correcting codes. SIAM Journal on Discrete Mathematics, 24(1): 33–55, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chan:2015:SST

[CGM^{+}15] Melody Chan, Darren Glass, Matthew Macauley, David Perkinson, Caryn Werner, and Qiaoyu Yang. Sandpiles, spanning trees, and plane duality. SIAM Journal on Discrete Mathematics, 29(1):

461–471, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Conte:2019:LMS

[CGMV19] Alessio Conte, Roberto Grossi, Andrea Marino, and Luca Versari. Listing maximal subgraphs satisfying strongly accessible properties. SIAM Journal on Discrete Mathematics, 33(2):587–613, ???? 2019. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chekuri:2006:SCP

[CGN06a] Chandra Chekuri, Sudipto Guha, and Joseph (Seffi) Naor. The Steiner k-cut problem. SIAM Journal on Discrete Mathematics, 20(1): 261–271, January 2006. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chekuri:2006:EOG

[CGN^{+}06b] Chandra Chekuri, Anupam Gupta, Ilan Newman, Yuri Rabinovich, and Alistair Sinclair. Embedding kouterplanar graphs into ℓ_{1}. SIAM Journal on Discrete Mathematics, 20(1):119–136, January 2006. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Clark:2021:HCD

[CGOvZ21] Ben Clark, Kevin Grace, James Oxley, and Stefan H. M. van Zwam. On the highly connected dyadic, near-regular, and sixth-root-of-unity matroids. SIAM Journal on Discrete Mathematics, 35(2):1356–1380, ???? 2021. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Czumaj:1998:TCT

[CGP98] Artur Czumaj, Leszek Gasieniec, and Andrzej Pelc. Time and cost trade-offs in gossiping. SIAM Journal on Discrete Mathematics, 11(3): 400–413, August 1998. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ciolan:2016:CNS

[CGSM16] Emil-Alexandru Ciolan, Pedro A. García-Sánchez, and Pieter Moree. Cyclotomic numerical semigroups. SIAM Journal on Discrete Mathematics, 30(2):650–668, ???? 2016. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chudnovsky:2020:OTC

[CGSZ20] Maria Chudnovsky, Jan Goedgebeur, Oliver Schaudt, and Mingxian Zhong. Obstructions for three-coloring and list three-coloring H-free graphs. SIAM Journal on Discrete Mathematics, 34(1): 431–469, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chakrabarty:2014:SHN

[CGV^{+}14] Deeparnab Chakrabarty, Gagan Goel, Vijay V. Vazirani, Lei Wang, and Changyuan Yu. Submodularity helps in Nash and nonsymmetric bargaining games. SIAM Journal on Discrete Mathematics, 28 (1):99–115, ???? 2014. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Camerini:1989:ITM

[CH89] Paolo M. Camerini and Horst W. Hamacher. Intersection of two matroids: (condensed) border graphs and ranking. SIAM Journal on Discrete Mathematics, 2(1): 16–27, February 1989. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chen:2001:AMC

[CH01] Po-Ning Chen and Yunghsiang S. Han. Asymptotic minimum covering radius of block codes. SIAM Journal on Discrete Mathematics, 14(4):549–564, 2001. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/37999.

Chen:2006:MRS

[CH06a] Hong-Bin Chen and Frank K.

Hwang. On multicast rearrangeable 3-stage Clos networks without first-stage fanout. SIAM Journal on Discrete Mathematics, 20(2): 287–290, January 2006. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Czygrinow:2006:GPG

[CH06b] Andrzej Czygrinow and Glenn Hurlbert. Girth, pebbling, and grid thresholds. SIAM Journal on Discrete Mathematics, 20(1):1–10, January 2006. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Croot:2010:FSS

[CH10] Ernie Croot and Derrick Hart. h-fold sums from a set with few products. SIAM Journal on Discrete Mathematics, 24(2):505–519, ???? 2010. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Churchley:2011:LPG

[CH11] Ross Churchley and Jing Huang. Line-polar graphs: Characterization and recognition. SIAM Journal on Discrete Mathematics, 25(3): 1269–1284, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.

                                        siam.org/sidma/resource/
                                        1/sjdmec/v25/i3/p1269_s1.

Carraher:2013:ECN

[CH13] James M. Carraher and Stephen G. Hartke. Eulerian circuits with no monochromatic transitions in edgecolored digraphs. SIAM Journal on Discrete Mathematics, 27(4):1924–1939, ???? 2013. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chakrabarty:2015:RCF

[CH15] Deeparnab Chakrabarty and Zhiyi Huang. Recognizing coverage functions. SIAM Journal on Discrete Mathematics, 29(3):1585–1599, ???? 2015. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Carraher:2017:ECN

[CH17] James M. Carraher and Stephen G. Hartke. Eulerian circuits with no monochromatic transitions in edgecolored digraphs with all vertices of outdegree three. SIAM Journal on Discrete Mathematics, 31(1):190–209, ???? 2017. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cho:2019:PUC

[CH19] Soojin Cho and JiSun Huh. On e-positivity and e-unimodality of chromatic quasi-symmetric functions. SIAM Journal on Discrete Mathematics, 33(4): 2286–2315, ???? 2019. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chawin:2023:INH

[CH23] Dror Chawin and Ishay Haviv. Improved NP-hardness of approximation for orthogonality dimension and minrank. SIAM Journal on Discrete Mathematics, 37(4): 2670–2688, November 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chan:1991:PTV

[Cha91] Clara Chan. Plane trees and H-vectors of shellable cubical complexes. SIAM Journal on Discrete Mathematics, 4 (4):568–574, November 1991. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chandran:2003:HGG

[Cha03]  L. Sunil Chandran.  A

high girth graph construction. SIAM Journal on Discrete Mathematics, 16 (3):366–370, 2003. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/38789.

Chang:2016:RSB

[Cha16] Mei-Chu Chang. A remark on sieving in biased coin convolutions. SIAM Journal on Discrete Mathematics, 30(2): 1181–1192, ???? 2016. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chan:2019:RWT

[Cha19] Swee Hong Chan. Rotor walks on transient graphs and the wired spanning forest. SIAM Journal on Discrete Mathematics, 33(4): 2369–2393, ???? 2019. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chen:1993:ICS

[Che93] William Y. C. Chen. Induced cycle structures of the hyperoctahedral group. SIAM Journal on Discrete Mathematics, 6(3):353–362, August 1993. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chen:1994:LTA

[Che94] Jianer Chen. A linear-time algorithm for isomorphism of graphs of bounded average genus. SIAM Journal on Discrete Mathematics, 7(4):614– 631, November 1994. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chepoi:1998:DPD

[Che98] Victor Chepoi. On distancepreserving and domination elimination orderings. SIAM Journal on Discrete Mathematics, 11(3):414–436, August 1998. CODEN SJD-MEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cheng:2004:IAA

[Che04] Christine T. Cheng. Improved approximation algorithms for the demand routing and slotting problem with unit demands on rings. SIAM Journal on Discrete Mathematics, 17(3):384–402, 2004. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/38672.

Cheng:2007:CFF

[Che07] Qi Cheng. Constructing finite field extensions with large order elements. SIAM Journal on Discrete Mathematics, 21 (3):726–730, ???? 2007. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chen:2009:AIS

[Che09] Ning Chen. On the approximability of influence in social networks. SIAM Journal on Discrete Mathematics, 23(3): 1400–1415, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cheng:2016:SMC

[Che16] Christine T. Cheng. On the stable matchings that can be reached when the agents go marching in one by one. SIAM Journal on Discrete Mathematics, 30(4): 2047–2063, ???? 2016. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chen:2017:CDI

[Che17] Beifang Chen. Conformal decomposition of integral tensions and potentials of signed graphs. SIAM Journal on Discrete Mathematics, 31(4): 2457–2478, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chen:2022:CEM

[Che22] Rong Chen. The 9-connected excluded minors for the class of quasi-graphic matroids. SIAM Journal on Discrete Mathematics, 36 (3):1627–1644, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M142784X.

Chen:2024:AMO

[Che24] Xiangying Chen. An axiomatization of matroids and oriented matroids as conditional independence models. SIAM Journal on Discrete Mathematics, 38(2): 1526–1536, May 2024. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Curtis:2009:NUC

[CHHM09] Dawn Curtis, Taylor Hines, Glenn Hurlbert, and Tatiana Moyer. Near-universal cycles for subsets exist. SIAM Journal on Discrete Mathematics, 23(3):1441–1449, ???? 2009. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chimani:2011:FCN

[Chi11] Markus Chimani. Facets in the crossing number polytope. SIAM Journal on Discrete Mathematics, 25(1): 95–111, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v25/i1/p95_s1.

Chitnis:2023:TLB

[Chi23] Rajesh Chitnis. A tight lower bound for edge-disjoint paths on planar DAGs. SIAM Journal on Discrete Mathematics, 37(2):556–572, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1395089.

Chang:2004:BVN

[CHJ^{+}04] Yi-Wu Chang, Joan P. Hutchinson, Michael S. Jacobson, Jeno Lehel, and Douglas B. West. The bar visibility number of a graph. SIAM Journal on Discrete Mathematics, 18(3):462–471, 2004. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/34345.

Choi:2010:DCN

[CHK10] Jeong Ok Choi, Stephen G. Hartke, and Hemanshu Kaul. Distinguishing chromatic number of Cartesian products of graphs. SIAM Journal on Discrete Mathematics, 24(1): 82–100, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cseh:2017:PMT

[CHK17] Ágnes Cseh, Chien-Chung Huang, and Telikepalli Kavitha. Popular matchings with twosided preferences and onesided ties. SIAM Journal on Discrete Mathematics, 31(4): 2348–2377, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cambie:2024:PCI

[CHKW24] Stijn Cambie, Penny Haxell, Ross J. Kang, and Ronen Wdowinski. A precise condition for independent transversals in bipartite covers. SIAM Journal on Discrete Mathematics, 38(2): 1451–1461, May 2024. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chakraborti:2024:RSC

[CHLT24] Debsoumya Chakraborti, Kevin Hendrey, Ben Lund, and Casey Tompkins. Rainbow saturation for complete graphs. SIAM Journal on Discrete Mathematics, 38(1): 1090–1112, March 2024. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cohen:2000:BCI

[CHLZ00] Gérard D. Cohen, Iiro Honkala, Antoine Lobstein, and Gilles Zémor. Bounds for codes identifying vertices in the hexagonal grid. SIAM Journal on Discrete Mathematics, 13(4):492–504, 2000. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/36099.

Caceres:2007:MDC

[CHM^{+}07] José Cáceres, Carmen Hernando, Mercè Mora, Ignacio M. Pelayo, María L. Puertas, Carlos Seara, and David R. Wood. On the metric dimension of Cartesian products of graphs. SIAM Journal on Discrete Mathematics, 21(2):423–441, ???? 2007. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cardinal:2023:CGP

[CHM^{+}23] Jean Cardinal, Hung P. Hoang, Arturo Merino, Ondrej Micka, and Torsten Mütze. Combinatorial generation via permutation languages. v. acyclic orientations. SIAM Journal on Discrete Mathematics, 37 (3):1509–1547, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/23M1546567.

Clemens:2021:MBG

[CHMP21] Dennis Clemens, Fabian Hamann, Yannick Mogge, and Olaf Parczyk. Maker– breaker games on randomly perturbed graphs. SIAM Journal on Discrete Mathematics, 35(4):2723–2748, ???? 2021. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chartrand:1989:MD

[CHO^{+}89] G. Chartrand, H. Hevia, O. Oellermann, F. Saba, and A. Schwenk. A matter of degree. SIAM Journal on Discrete Mathematics, 2(4):456– 466, November 1989. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chopra:1992:ESD

[Cho92a] Sunil Chopra. The equivalent subgraph and directed cut polyhedra on series-parallel graphs. SIAM Journal on Discrete Mathematics, 5(4): 475–490, November 1992. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chopra:1992:PES

[Cho92b] Sunil Chopra. Polyhedra of the equivalent subgraph problem and some edge connectivity problems. SIAM Journal on Discrete Mathematics, 5(3):321–337, August 1992. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chopra:1994:GPP

[Cho94a] Sunil Chopra. The graph partitioning polytope on seriesparallel and 4-Wheel free graphs. SIAM Journal on Discrete Mathematics, 7(1): 16–31, February 1994. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chopra:1994:ECS

[Cho94b] Sunil Chopra. The k-edge-connected spanning subgraph polyhedron. SIAM Journal on Discrete Mathematics, 7 (2):245–259, May 1994. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chow:2009:RRB

[Cho09] Timothy Y. Chow. Reduction of Rota’s Basis Conjecture to a problem on three bases. SIAM Journal on Discrete Mathematics, 23(1): 369–371, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Carter:1990:DLE

[CHP^{+}90] M. Carter, M. Hendy, D. Penny, L. A. Szekely, and N. C. Wormald. On the distribution of lengths of evolutionary trees. SIAM Journal on Discrete Mathematics, 3(1): 38–47, February 1990. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chung:1989:PH

[Chu89] Fan R. K. Chung. Pebbling in hypercubes. SIAM Journal on Discrete Mathematics, 2 (4):467–472, November 1989. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chang:1988:ODT

[CHW88] X. M. Chang, F. K. Hwang, and J. F. Weng. Optimal detection of two defectives with a parity check device. SIAM Journal on Discrete Mathematics, 1(1):38–44, February 1988. CODEN SJDMEC.

ISSN 0895-4801 (print), 1095-7146 (electronic).

Csorba:2010:ANG

[CHW10] Péter Csorba, Cor A. J.

Hurkens, and Gerhard J. Woeginger. The Alcuin number of a graph and its connections to the vertex cover number. SIAM Journal on Discrete Mathematics, 24(3): 757–769, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chang:2015:HDS

[CHX15] Xiang-Ke Chang, Xing-Biao Hu, and Guoce Xin. Hankel determinant solutions to several discrete integrable systems and the Laurent property. SIAM Journal on Discrete Mathematics, 29(1): 667–682, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chiang:2013:TSS

[CHY13] Chun-Ying Chiang, Liang-Hao Huang, and Hong-Gwa Yeh. Target set selection problem for honeycomb networks. SIAM Journal on Discrete Mathematics, 27(1): 310–328, ???? 2013. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chen:2004:ERM

[CHZ04] W. R. Chen, F. K. Hwang, and Xuding Zhu. Equivalence of the 1-rate model to the classical model on strictly nonblocking switching networks. SIAM Journal on Discrete Mathematics, 17(3):446–452, 2004. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/41480.

Charpin:2009:CWE

[CHZ09] Pascale Charpin, Tor Helleseth, and Victor Zinoviev. On cosets of weight 4 of BCH(2^{m}, 8), m even, and exponential sums. SIAM Journal on Discrete Mathematics, 23(1):59–78, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chestnut:2018:SBQ

[CHZ18] Stephen R. Chestnut, Robert Hildebrand, and Rico Zenklusen. Sublinear bounds for a quantitative Doignon–Bell– Scarf theorem. SIAM Journal on Discrete Mathematics, 32 (1):352–371, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Christie:2001:SSR

[CI01] David A. Christie and Robert W. Irving. Sorting strings by reversals and by transpositions. SIAM Journal on Discrete Mathematics, 14 (2):193–206, 2001. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/33199.

Constantinescu:2007:LZC

[CI07] Sorin Constantinescu and Lucian Ilie. The Lempel–Ziv complexity of fixed points of morphisms. SIAM Journal on Discrete Mathematics, 21(2): 466–481, ???? 2007. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cibulka:2013:MSR

[Cib13] Josef Cibulka. Maximum size of reverse-free sets of permutations. SIAM Journal on Discrete Mathematics, 27(1): 232–239, ???? 2013. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cifre:2018:DBM

[CIN18] María A. Hernández Cifre, David Iglesias, and Jesús Yepes Nicolás. On a discrete Brunn–Minkowski type inequality. SIAM Journal on Discrete Mathematics, 32(3): 1840–1856, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chan:2023:NNC

[CJL^{+}23] Patrick Chan, Katherine Johnston, Joseph Lent, Alexander Ruys de Perez, and Anne Shiu. Nondegenerate neural codes and obstructions to closed-convexity. SIAM Journal on Discrete Mathematics, 37(1):114–145, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1452147.

Chor:1991:ZOL

[CK91] Benny Chor and Eyal Kushilevitz. A zero-one law for Boolean privacy. SIAM Journal on Discrete Mathematics, 4(1):36–47, February 1991. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chang:1996:LPG

[CK96] Gerard J. Chang and David Kuo. The L(2, 1)-labeling problem on graphs. SIAM Journal on Discrete Mathematics, 9(2):309–316, May 1996. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chen:1999:GED

[CK99] Jianer Chen and Saroja P. Kanchi. Graph ear decompositions and graph embeddings. SIAM Journal on Discrete Mathematics, 12(2):229–242, May 1999.

CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/30423.

Charpin:2008:CMB

[CK08a] Pascale Charpin and Gohar M. Kyureghyan. Cubic monomial bent functions: a subclass of M. SIAM Journal on Discrete Mathematics, 22 (2):650–665, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chudnovsky:2008:DTP

[CK08b] Maria Chudnovsky and Rohan Kapadia. Detecting a theta or a prism. SIAM Journal on Discrete Mathematics, 22(3):1164–1186, ???? 2008. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Corneil:2008:UVG

[CK08c] Derek G. Corneil and Richard M.

Krueger. A unified view of graph searching. SIAM Journal on Discrete Mathematics, 22(4):1259–1276, ???? 2008. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cochrane:2011:PGK

[CK11] Todd Cochrane and Sergei Konyagin. Proof of the Goresky Klapper Conjecture on decimations of Lsequences. SIAM Journal on Discrete Mathematics, 25(4): 1812–1831, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v25/i4/p1812_s1.

Chekuri:2014:GRS

[CK14] Chandra Chekuri and Nitish Korula. A graph reduction step preserving elementconnectivity and packing Steiner trees and forests. SIAM Journal on Discrete Mathematics, 28(2):577–597, ???? 2014. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chiu:2018:HDC

[CK18] Man-Kwun Chiu and Matias Korman. High dimensional consistent digital segments. SIAM Journal on Discrete Mathematics, 32(4): 2566–2590, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cunha:2013:ATD

[CKdAHdF13] Luís Felipe I. Cunha, Luis Antonio B. Kowada, Rodrigo de A. Hausen, and Celina M. H. de Figueiredo. Advancing the transposition distance and diameter through lonely permutations. SIAM Journal on Discrete Mathematics, 27(4):1682–1709, ???? 2013. CODEN SJDMEC.

ISSN 0895-4801 (print), 1095-7146 (electronic).

Charikar:2004:MWZ

[CKK^{+}04] Moses Charikar, Jon Kleinberg, Ravi Kumar, Sridhar Rajagopalan, Amit Sahai, and Andrew Tomkins. Minimizing wirelength in zero and bounded skew clock trees. SIAM Journal on Discrete Mathematics, 17 (4):582–595, 2004. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/35262.

Christofides:2010:SDC

[CKKO10] Demetres Christofides, Peter Keevash, Daniela Kühn, and Deryk Osthus. A semiexact degree condition for Hamilton cycles in digraphs. SIAM Journal on Discrete Mathematics, 24(3):709–756, ???? 2010. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Carlson:2021:LBM

[CKL^{+}21] Charles Carlson, Alexandra Kolla, Ray Li, Nitya Mani, Benny Sudakov, and Luca Trevisan. Lower bounds for max-cut in H-free graphs via semidefinite programming. SIAM Journal on Discrete Mathematics, 35(3): 1557–1568, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cardinal:2014:MOC

[CKMU14] Jean Cardinal, Kolja Knauer, Piotr Micek, and Torsten Ueckerdt. Making octants colorful and related covering decomposition problems. SIAM Journal on Discrete Mathematics, 28(4): 1948–1959, ???? 2014. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cygan:2013:SFO

[CKN13] Marek Cygan, Guy Kortsarz, and Zeev Nutov. Steiner forest orientation problems. SIAM Journal on Discrete Mathematics, 27(3):1503– 1513, ???? 2013. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chen:2015:PTD

[CKN^{+}15] Jiehua Chen, Christian Komusiewicz, Rolf Niedermeier, Manuel Sorge, Ondrej Suchý, and Mathias Weller. Polynomial-time data reduction for the subset interconnection design problem. SIAM Journal on Discrete Mathematics, 29(1): 1–25, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cioaba:2016:MOR

[CKNV16] Sebastian M. Cioaba, Jack H.

Koolen, Hiroshi Nozaki, and Jason R. Vermette. Maximizing the order of a regular graph of given valency and second eigenvalue. SIAM Journal on Discrete Mathematics, 30(3):1509–1525, ???? 2016. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chekuri:2005:LPF

[CKNZ05] C. Chekuri, S. Khanna, J. Naor, and L. Zosin. A linear programming formulation and approximation algorithms for the metric labeling problem. SIAM Journal on Discrete Mathematics, 18(3):608–625, 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/39693.

Corneil:2006:LOS

[CKOS06] Derek G. Corneil, Ekkehard Köhler, Stephan Olariu, and Lorna Stewart. Linear orderings of subfamilies of ATfree graphs. SIAM Journal on Discrete Mathematics, 20(1): 105–118, January 2006. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chakrabarty:2013:HLR

[CKP13a] Deeparnab Chakrabarty, Jochen Könemann, and David Pritchard.

Hypergraphic LP relaxations for Steiner trees. SIAM Journal on Discrete Mathematics, 27(1):507–533, ???? 2013. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cheilaris:2013:UMC

[CKP13b] Panagiotis Cheilaris, Balázs Keszegh, and Dömötör Pálvölgyi. Unique-maximum and conflictfree coloring for hypergraphs and tree graphs. SIAM Journal on Discrete Mathematics, 27(4):1775–1787, ???? 2013. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Covert:2016:SPF

[CKP16] David Covert, Doowon Koh, and Youngjin Pi. On the sums of any k points in finite fields. SIAM Journal on Discrete Mathematics, 30(1): 367–382, ???? 2016. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chudnovsky:2021:FLC

[CKP^{+}21] Maria Chudnovsky, Jason King, Michał Pilipczuk, Paweł Rzazewski, and Sophie Spirkl. Finding large Hcolorable subgraphs in hereditary graph classes. SIAM Journal on Discrete Mathematics, 35(4):2357–2386, ???? 2021. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chepoi:2022:ACO

[CKP22] Victor Chepoi, Kolja Knauer, and Manon Philibert. Ample completions of oriented matroids and complexes of uniform oriented matroids. SIAM Journal on Discrete Mathematics, 36(1):509–535, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:/

/epubs.siam.org/doi/10.
1137/20M1355434.

Chudnovsky:2013:LSR

[CKPS13] Maria Chudnovsky, Andrew D. King, Matthieu Plumettaz, and Paul Seymour. A local strengthening of Reed’s ω, Δ, χ conjecture for quasi-line graphs. SIAM Journal on Discrete Mathematics, 27(1):95–108, ???? 2013. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chee:2015:PCA

[CKPS15] Yeow Meng Chee, Han Mao Kiah, Punarbasu Purkayastha, and Patrick Solé. Product construction of affine codes. SIAM Journal on Discrete Mathematics, 29(3): 1540–1552, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chrobak:1991:NRS

[CKPV91] M. Chrobak, H. Karloff, T. Payne, and S. Vishwanathan. New results on server problems. SIAM Journal on Discrete Mathematics, 4(2):172–181, May 1991. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chambers:2009:EPR

[CKPW09] Erin W. Chambers, Bill Kinnersley, Noah Prince, and Douglas B. West. Extremal problems for Roman domination. SIAM Journal on Discrete Mathematics, 23(3): 1575–1586, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cremona:1990:SLP

[CL90] John Cremona and Susan Landau. Shrinking lattice polyhedra. SIAM Journal on Discrete Mathematics, 3(3): 338–348, August 1990. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chan:1991:EHC

[CL91a] Mee Yee Chan and Shiang-Jen Lee. On the existence of Hamiltonian circuits in faulty hypercubes. SIAM Journal on Discrete Mathematics, 4 (4):511–527, November 1991. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chrobak:1991:NAS

[CL91b] Marek Chrobak and Lawrence L.

Larmore. A new approach to the server problem. SIAM Journal on Discrete Mathematics, 4(3):323–328, August 1991. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Coppersmith:2005:CBO

[CL05] Don Coppersmith and Moshe Lewenstein. Constructive bounds on ordered factorizations. SIAM Journal on Discrete Mathematics, 19 (2):301–303, 2005. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/44586.

Chung:2006:VGC

[CL06] Fan Chung and Linyuan Lu. The volume of the giant component of a random graph with given expected degrees. SIAM Journal on Discrete Mathematics, 20(2): 395–411, January 2006. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chee:2007:EGR

[CL07] Yeow Meng Chee and Alan C. H. Ling. On extremal k-graphs without repeated copies of 2-intersecting edges. SIAM Journal on Discrete Mathematics, 21(3):805–821,

????  2007.  CODEN

SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cheng:2011:SRM

[CL11] Christine T. Cheng and Anhua Lin. Stable roommates matchings, mirror posets, median graphs, and the local/global median phenomenon in stable matchings. SIAM Journal on Discrete Mathematics, 25(1): 72–94, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v25/i1/p72_s1.

Cheriyan:2013:AAM

[CL13] J. Cheriyan and B. Laekhanukit. Approximation algorithms for minimum-cost k-(S, T) connected digraphs. SIAM Journal on Discrete Mathematics, 27(3):1450–1481, ???? 2013. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Candrakova:2015:ACF

[CL15a] Barbora Candráková and Robert Lukot’ka. Avoiding 5-circuits in 2-factors of cubic graphs. SIAM Journal on Discrete Mathematics, 29(3): 1387–1405, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Coutinho:2015:NLP

[CL15b] Gabriel Coutinho and Henry Liu. No Laplacian perfect state transfer in trees. SIAM Journal on Discrete Mathematics, 29(4):2179–2188, ???? 2015. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Csikvari:2015:HTP

[CL15c] Péter Csikvári and Zhicong Lin. Homomorphisms of trees into a path. SIAM Journal on Discrete Mathematics, 29(3): 1406–1422, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Candrakova:2016:SCC

[CL16] Barbora Candráková and Robert Lukotka. Short cycle covers on cubic graphs by choosing a 2-factor. SIAM Journal on Discrete Mathematics, 30(4):2086–2106, ???? 2016. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chakraborti:2020:EGL

[CL20a] Debsoumya Chakraborti and Po-Shen Loh. Extremal graphs with local covering conditions. SIAM Journal on Discrete Mathematics, 34(2): 1354–1374, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cranston:2020:CFP

[CL20b] Daniel W. Cranston and Jiaao Li. Circular flows in planar graphs. SIAM Journal on Discrete Mathematics, 34 (1):497–519, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cui:2021:TGC

[CL21] Qing Cui and On-Hei Solomon Lo. Tight gaps in the cycle spectrum of 3-connected planar graphs. SIAM Journal on Discrete Mathematics, 35(3): 2039–2048, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cifre:2022:DLB

[CL22] María A. Hernández Cifre and Eduardo Lucas. On discrete LOG-Brunn–Minkowski type inequalities. SIAM Journal on Discrete Mathematics, 36(3):1748–1760, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1451725.

Chavez-Lomeli:2011:BC

[CLGH11] Laura Chávez-Lomelí, Luis Goddyn, and Winfried Hochstättler. Balancing covectors. SIAM Journal on Discrete Mathematics, 25(3):1315–1318, ???? 2011. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v25/i3/p1315_s1.

Choi:2018:CCO

[CLiO18] Ilkyoo Choi, Chun-Hung Liu, and Sang il Oum. Characterization of cycle obstruction sets for improper coloring planar graphs. SIAM Journal on Discrete Mathematics, 32(2): 1209–1228, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cheng:2018:SGM

[CLLZ18] Jian Cheng, You Lu, Rong Luo, and Cun-Quan Zhang. Signed graphs: From modulo flows to integer-valued flows. SIAM Journal on Discrete Mathematics, 32(2): 956–965, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Canovas:2003:FOP

[CLM03] Lázaro Cánovas, Mercedes Landete, and Alfredo Marín. Facet obtaining procedures for set packing problems. SIAM Journal on Discrete Mathematics, 16(1): 127–155, 2003. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/36678.

Campos:2022:ADG

[CLMS22] Victor Campos, Raul Lopes, Ana Karolinna Maia, and Ignasi Sau. Adapting the directed grid theorem into an FPT algorithm. SIAM Journal on Discrete Mathematics, 36(3):1887–1917, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1452664.

Chan:2009:EFT

[CLS09] W. H. Chan, Peter C. B. Lam, and W. C. Shiu. Edgeface total chromatic number of Halin graphs. SIAM Journal on Discrete Mathematics, 23(3):1646–1654, ???? 2009. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Correa:2015:TTC

[CLS15] José Correa, Omar Larré, and José A. Soto. TSP tours in cubic graphs: Beyond 4/3. SIAM Journal on Discrete Mathematics, 29(2): 915–939, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cusack:2012:CPD

[CLST12] Charles A. Cusack, Timothy Lewis, Daniel Simpson, and Samuel Taggart. The complexity of pebbling in diameter two graphs. SIAM Journal on Discrete Mathematics, 26 (3):919–928, ???? 2012. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cohen:1996:TBS

[CLVZ96] Gerard Cohen, Simon Litsyn, Alexander Vardy, and Gilles Zémor. Tilings of binary spaces. SIAM Journal on Discrete Mathematics, 9(3): 393–412, August 1996. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chen:2009:FSI

[CLW09] William Y. C. Chen, Jiuqiang Liu, and Larry X. W. Wang. Families of sets with intersecting clusters. SIAM Journal on Discrete Mathematics, 23(3):1249–1260, ???? 2009. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cao:2022:NIF

[CLWZ22] Mengyu Cao, Benjian Lv, Kaishun Wang, and Sanming Zhou. Nontrivial tintersecting families for vector spaces. SIAM Journal on Discrete Mathematics, 36 (3):1823–1847, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1357421.

Curran:2005:CDC

[CLY05a] Sean Curran, Orlando Lee, and Xingxing Yu. Chain decompositions of 4-connected graphs. SIAM Journal on Discrete Mathematics, 19 (4):848–880, 2005. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/43459.

Curran:2005:NPC

[CLY05b] Sean Curran, Orlando Lee, and Xingxing Yu. Nonseparating planar chains in 4connected graphs. SIAM Journal on Discrete Mathematics, 19(2):399–419, 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/43458.

Calderbank:1990:QSD

[CM90] A. R. Calderbank and Patrick Morton. Quasi-symmetric 3designs and elliptic curves. SIAM Journal on Discrete Mathematics, 3(2):178–196, May 1990. CODEN SJD-MEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Comellas:2003:SCP

[CM03] Francesc Comellas and Margarida Mitjana. The spectra of cycle prefix digraphs. SIAM Journal on Discrete Mathematics, 16 (3):418–421, 2003. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/38060.

Chaichanavong:2005:SBT

[CM05a] Panu Chaichanavong and Brian H. Marcus. Stabilization of block-type-decodability properties for constrained systems. SIAM Journal on Discrete Mathematics, 19(2):321–344, 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/44367.

Chardon:2005:CGA

[CM05b] Marc Chardon and Aziz Moukrim. The Coffman– Graham algorithm optimally solves UET task systems with overinterval orders. SIAM Journal on Discrete Mathematics, 19(1):109–121, 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/39499.

Cornaz:2007:MIB

[CM07] Denis Cornaz and A. Ridha Mahjoub. The maximum induced bipartite subgraph problem with edge weights. SIAM Journal on Discrete Mathematics, 21(3):662–675,

????  2007.  CODEN

SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Connamacher:2012:STS

[CM12a] Harold Connamacher and Michael Molloy. The satisfiability threshold for a seemingly intractable random constraint satisfaction problem. SIAM Journal on Discrete Mathematics, 26(2): 768–800, ???? 2012. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Coppersmith:2012:BN

[CM12b] Don Coppersmith and Victor S. Miller. Binary nontiles. SIAM Journal on Discrete Mathematics, 26(1): 30–38, ???? 2012. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sidma/resource/
                                        1/sjdmec/v26/i1/p30_s1.

Czygrinow:2014:TCC

[CM14] Andrzej Czygrinow and Theodore Molla. Tight codegree condition for the existence of loose Hamilton cycles in 3graphs. SIAM Journal on Discrete Mathematics, 28(1): 67–76, ???? 2014. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cooper:2016:CSH

[CM16] Jeff Cooper and Dhruv Mubayi. Coloring sparse hypergraphs. SIAM Journal on Discrete Mathematics, 30(2): 1165–1180, ???? 2016. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cleary:2018:ECD

[CM18] Sean Cleary and Roland Maio. Edge conflicts do not determine geodesics in the associahedron. SIAM Journal on Discrete Mathematics, 32 (2):1003–1015, ???? 2018. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chen:2024:PMC

[CM24] Hao Chen and Jie Ma. A property on monochromatic copies of graphs containing a triangle. SIAM Journal on Discrete Mathematics, 38(1): 316–326, January 2024. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Clementi:2010:FTE

[CMM^{+}10] Andrea E. F. Clementi, Claudio Macci, Angelo Monti, Francesco Pasquale, and Riccardo Silvestri. Flooding time of edge-Markovian evolving graphs. SIAM Journal on Discrete Mathematics, 24(4): 1694–1712, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sidma/resource/
                                        1/sjdmec/v24/i4/p1694_s1.

Chandrasekaran:2020:OMC

[CMM20] Karthekeyan Chandrasekaran, Matthias Mnich, and Sahand Mozaffari. Odd multiway cut in directed acyclic graphs. SIAM Journal on Discrete Mathematics, 34(2): 1385–1408, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cruickshank:2024:GRL

[CMM^{+}24] James Cruickshank, Fatemeh Mohammadi, Harshit J. Motwani, Anthony Nixon, and Shin ichi Tanigawa. Global rigidity of line constrained frameworks. SIAM Journal on Discrete Mathematics, 38 (1):743–763, February 2024. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cesmelioglu:2015:BFS

[ÇMP15] Ayça Çesmelioğlu, Wilfried Meidl, and Alexander Pott. Bent functions, spreads, and o-polynomials. SIAM Journal on Discrete Mathematics, 29 (2):854–867, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Caskurlu:2017:PVC

[CMPS17] Bugra Caskurlu, Vahan Mkrtchyan, Ojas Parekh, and K. Subramani. Partial vertex cover and budgeted maximum coverage in bipartite graphs. SIAM Journal on Discrete Mathematics, 31(3): 2172–2184, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cannon:2018:PTR

[CMR18] Sarah Cannon, Sarah Miracle, and Dana Randall. Phase transitions in random dyadic tilings and rectangular dissections. SIAM Journal on Discrete Mathematics, 32(3): 1966–1992, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chapuy:2009:BRM

[CMS09] Guillaume Chapuy, Michel Marcus, and Gilles Schaeffer. A bijection for rooted maps on orientable surfaces. SIAM Journal on Discrete Mathematics, 23(3):1587–1611, ???? 2009. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chappelon:2018:CTT

[CMSM^{+}18] J. Chappelon, L. Martínez-Sandoval, L. Montejano, L. P. Montejano, and J. L. Ramírez Alfonsín. Codimension two and three Kneser transversals. SIAM Journal on Discrete Mathematics, 32(2): 1351–1363, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cheung:2017:PDA

[CMSV17] Maurice Cheung, Julián Mestre, David B. Shmoys, and José Verschae. A primaldual approximation algorithm for min-sum singlemachine scheduling problems. SIAM Journal on Discrete Mathematics, 31(2): 825–838, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chakrabarty:2010:DEO

[CMV10] Deeparnab Chakrabarty, Aranyak Mehta, and Vijay V. Vazirani. Design is as easy as optimization. SIAM Journal on Discrete Mathematics, 24 (1):270–286, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Clark:2016:SFM

[CMvZW16] Ben Clark, Dillon Mayhew, Stefan van Zwam, and Geoff Whittle. The structure of {U_{2,5}, U_{3,5}}-fragile matroids. SIAM Journal on Discrete Mathematics, 30(3): 1480–1508, ???? 2016. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cook:2012:CMG

[CN12] David Cook II and Uwe Nagel. Cohen–Macaulay graphs and face vectors of flag complexes. SIAM Journal on Discrete Mathematics, 26(1):89–101, ???? 2012. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Csaba:2019:EGH

[CNG19] Béla Csaba and Judit Nagy-György. Embedding graphs having Ore-degree at most five. SIAM Journal on Discrete Mathematics, 33(1): 474–508, ???? 2019. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Czerwinski:2021:IBE

[CNP21] Wojciech Czerwiński, Wojciech Nadara, and Marcin Pilipczuk. Improved bounds for the excluded-minor approximation of treedepth. SIAM Journal on Discrete Mathematics, 35(2):934–947, ???? 2021. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Choi:1989:GBM

[CNR89] Hyeong-Ah Choi, Kazuo Nakajima, and Chong S. Rim. Graph bipartization and via minimization. SIAM Journal on Discrete Mathematics, 2(1):38–47, February 1989.

CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chudnovsky:2015:ESA

[CNRS15] Maria Chudnovsky, Sergey Norin, Bruce Reed, and Paul Seymour. Excluding a substar and an antisubstar. SIAM Journal on Discrete Mathematics, 29(1): 297–308, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chudnovsky:2024:CRB

[CNST24] Maria Chudnovsky, Sergey Norin, Paul D. Seymour, and Jérémie Turcotte. Cops and robbers on P_{5}-free graphs. SIAM Journal on Discrete Mathematics, 38(1):845–856, February 2024. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Coja-Oghlan:2010:ESR

[COCF10] Amin Coja-Oghlan, Colin Cooper, and Alan Frieze. An efficient sparse regularity concept. SIAM Journal on Discrete Mathematics, 23(4): 2000–2034, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Coja-Oghlan:2017:WSW

[COHH17] A. Coja-Oghlan, A. Haqshenas, and S. Hetterich. Walksat stalls well below satisfiability. SIAM Journal on Discrete Mathematics, 31(2): 1160–1173, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Coja-Oghlan:2021:CMP

[COHK21] Amin Coja-Oghlan and Max Hahn-Klimroth. The cut metric for probability distributions. SIAM Journal on Discrete Mathematics, 35(2): 1096–1135, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Collins:1998:CS

[Col98] Karen L. Collins. Circulants and sequences. SIAM Journal on Discrete Mathematics, 11 (2):330–339, May 1998. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Coja-Oghlan:2010:FPP

[COL10] Amin Coja-Oghlan and André Lanka. Finding planted partitions in random graphs with general degree distributions. SIAM Journal on Discrete Mathematics, 23(4): 1682–1714, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Coja-Oghlan:2022:IAM

[COLMS22] Amin Coja-Oghlan, Philipp Loick, Balázs F. Mezei, and Gregory B. Sorkin. The Ising antiferromagnet and max cut on random regular graphs. SIAM Journal on Discrete Mathematics, 36 (2):1306–1342, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M137999X.

Constantine:1989:IDR

[Con89] Gregory M. Constantine. On the invertibility of the discrete Radon transform. SIAM Journal on Discrete Mathematics, 2(3):300–306, August 1989. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Constantine:2005:EDI

[Con05] Gregory M. Constantine. Edge-disjoint isomorphic multicolored trees and cycles in complete graphs. SIAM Journal on Discrete Mathematics, 18(3):577–580, 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/39740.

Conlon:2010:LRN

[Con10] David Conlon. On-line Ramsey numbers. SIAM Journal on Discrete Mathematics, 23(4):1954–1963, ???? 2010. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Coja-Oghlan:2012:DPR

[COPP12] Amin Coja-Oghlan and Angelica Y. Pachon-Pinzon. The decimation process in random k-SAT. SIAM Journal on Discrete Mathematics, 26(4): 1471–1509, ???? 2012. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Coregliano:2024:LCP

[Cor24] Leonardo N. Coregliano.

Left-cut-percolation and induced-Sidorenko bigraphs. SIAM Journal on Discrete Mathematics, 38(2):1586–1629, May 2024. CODEN SJD-MEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Corneil:1997:ATF

[COS97] Derek G. Corneil, Stephan Olariu, and Lorna Stewart. Asteroidal triple-free graphs. SIAM Journal on Discrete Mathematics, 10(3): 399–430, August 1997. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Corneil:2010:LSR

[COS10] Derek G. Corneil, Stephan Olariu, and Lorna Stewart. The LBFS structure and recognition of interval graphs. SIAM Journal on Discrete Mathematics, 23(4): 1905–1953, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Coppersmith:1996:QES

[CP96] Don Coppersmith and Steven Phillips. On a question of Erdős on subsequence sums. SIAM Journal on Discrete Mathematics, 9(2): 173–177, May 1996. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cogill:2010:STM

[CP10a] Randy Cogill and Cheng Peng. A spanning tree method for bounding hitting times of random walks on graphs. SIAM Journal on Discrete Mathematics, 24(3): 808–820, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Colomo:2010:LSL

[CP10b] F. Colomo and A. G. Pronko. The limit shape of large alternating sign matrices. SIAM Journal on Discrete Mathematics, 24(4):1558– 1571, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sidma/resource/
                                        1/sjdmec/v24/i4/p1558_s1.

Chung:2016:DRG

[CP16a] Fan Chung and Xing Peng. Decomposition of random graphs into complete bipartite graphs. SIAM Journal on Discrete Mathematics, 30 (1):296–310, ???? 2016. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cifuentes:2016:ECS

[CP16b] Diego Cifuentes and Pablo A.

Parrilo. Exploiting chordal structure in polynomial ideals: a Gröbner bases approach. SIAM Journal on Discrete Mathematics, 30(3): 1534–1570, ???? 2016. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cherkashin:2020:RBM

[CP20a] Danila Cherkashin and Fedor Petrov. Regular behavior of the maximal hypergraph chromatic number. SIAM Journal on Discrete Mathematics, 34(2):1326–1333, ???? 2020. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Coulson:2020:RDT

[CP20b] Matthew Coulson and Guillem Perarnau. A rainbow Dirac’s theorem. SIAM Journal on Discrete Mathematics, 34(3): 1670–1692, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Clement:2022:DDF

[CP22] François Clément and Thang Pham. Distribution of distances in five dimensions and related problems. SIAM Journal on Discrete Mathematics, 36(3):2271–2281, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL

https://epubs.siam.org/
doi/10.1137/21M1447088.

Chan:2023:EPI

[CPP23] Swee Hong Chan, Igor Pak, and Greta Panova. Effective poset inequalities. SIAM Journal on Discrete Mathematics, 37(3):1842–1880, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL

https://epubs.siam.org/
doi/10.1137/22M1532317.

Chudnovsky:2020:MWI

[CPPT20] Maria Chudnovsky, Marcin Pilipczuk, Michał Pilipczuk, and Stéphan Thomassé. On the maximum weight independent set problem in graphs without induced cycles of length at least five. SIAM Journal on Discrete Mathematics, 34(2):1472– 1483, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cygan:2013:SFV

[CPPW13] Marek Cygan, Marcin Pilipczuk, Michal Pilipczuk, and Jakub Onufry Wojtaszczyk. Subset feedback vertex set is fixedparameter tractable. SIAM Journal on Discrete Mathematics, 27(1):290–309, ???? 2013. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Czygrinow:1999:CQR

[CPR99] A. Czygrinow, S. Poljak, and V. Rödl. Constructive quasi-Ramsey numbers and tournament ranking. SIAM Journal on Discrete Mathematics, 12(1):48–63, February 1999. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31830.

Centeno:2013:GNV

[CPRdS13] C. C. Centeno, L. D.

Penso, D. Rautenbach, and V. G. Pereira de Sá. Geodetic number versus hull number in P_{3}-convexity. SIAM Journal on Discrete Mathematics, 27 (2):717–731, ???? 2013. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Corberan:2008:WGR

[CPS08] Angel Corberán, Isaac Plana, and José M. Sanchis. The windy general routing polyhedron: a global view of many known arc routing polyhedra. SIAM Journal on Discrete Mathematics, 22(2): 606–628, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chekuri:2020:LRT

[CQX20] Chandra Chekuri, Kent Quanrud, and Chao Xu. LP relaxation and tree packing for minimum k-cut. SIAM Journal on Discrete Mathematics, 34(2):1334–1353, ???? 2020. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chopra:1996:GAT

[CR96] Sunil Chopra and Giovanni Rinaldi. The graphical asymmetric traveling salesman polyhedron: Symmetric inequalities. SIAM Journal on Discrete Mathematics, 9 (4):602–624, November 1996. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chandran:2004:NMC

[CR04] L. Sunil Chandran and L. Shankar Ram. On the number of minimum cuts in a graph. SIAM Journal on Discrete Mathematics, 18(1):177–194, 2004. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/42713.

Cabello:2010:OCG

[CR10] Sergio Cabello and Günter Rote. Obnoxious centers in graphs. SIAM Journal on Discrete Mathematics, 24(4): 1713–1730, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v24/i4/p1713_s1.

Cranston:2013:CCF

[CR13] Daniel W. Cranston and Landon Rabern. Coloring clawfree graphs with Δ − 1 colors. SIAM Journal on Discrete Mathematics, 27(1): 534–549, ???? 2013. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cranston:2015:GBC

[CR15] Daniel W. Cranston and Landon Rabern. Graphs with χ = Δ have big cliques. SIAM Journal on Discrete Mathematics, 29(4):1792–1814, ???? 2015. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chen:2016:PPA

[CR16] Ricky X. F. Chen and Christian M. Reidys. Plane permutations and applications to a result of Zagier–Stanley and distances of permutations. SIAM Journal on Discrete Mathematics, 30(3): 1660–1684, ???? 2016. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cranston:2017:LCC

[CR17a] Daniel W. Cranston and Landon Rabern. List-coloring claw-free graphs with Δ − 1 colors. SIAM Journal on Discrete Mathematics, 31(2): 726–748, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cranston:2017:SFB

[CR17b] Daniel W. Cranston and Landon Rabern. Short fans and the 5/6 bound for line graphs. SIAM Journal on Discrete Mathematics, 31(3): 2039–2063, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cranston:2019:HZC

[CR19] Daniel W. Cranston and Landon Rabern. The Hilton– Zhao conjecture is true for graphs with maximum degree 4. SIAM Journal on Discrete Mathematics, 33(3): 1228–1241, ???? 2019. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cheraghchi:2023:SCS

[CR23] Mahdi Cheraghchi and João Ribeiro. Simple codes and sparse recovery with fast decoding. SIAM Journal on Discrete Mathematics, 37 (2):612–631, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1465354.

Cranston:2019:AEC

[Cra19] Daniel W. Cranston. Acyclic edge-coloring of planar graphs: Δ colors suffice when Δ is large. SIAM Journal on Discrete Mathematics, 33(2): 614–628, ???? 2019. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Crespi:2004:EFL

[Cre04] Valentino Crespi. Exact formulae for the Lovász theta function of sparse circulant graphs. SIAM Journal on Discrete Mathematics, 17(4):670–674, 2004. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/41852.

Cheng:1989:CSG

[CS89] Ying Cheng and N. J. A.

Sloane. Codes from symmetry groups, and a [32, 17, 8] code. SIAM Journal on Discrete Mathematics, 2(1):28– 37, February 1989. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Coffman:1991:SPU

[CS91] E. G. Coffman, Jr. and P. W.

Shor. A simple proof of the O(\sqrt{n} log^{3/4} n) upright matching bound. SIAM Journal on Discrete Mathematics, 4(1):48–57, February 1991. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Constantine:1994:SPI

[CS94] G. M. Constantine and T. H. Savits. A stochastic process interpretation of partition identities. SIAM Journal on Discrete Mathematics, 7(2):194–202, May 1994. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chor:1998:GAB

[CS98] Benny Chor and Madhu Sudan. A geometric approach to betweenness. SIAM Journal on Discrete Mathematics, 11 (4):511–523, November 1998. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Collins:2002:CCC

[CS02] Karen L. Collins and Benjamin Shemmer. Constructions of 3-colorable cores. SIAM Journal on Discrete Mathematics, 16 (1):74–80, 2002. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/39089.

Chekuri:2009:AID

[CS09] Chandra Chekuri and F. Bruce Shepherd. Approximate integer decompositions for undirected network design problems. SIAM Journal on Discrete Mathematics, 23(1): 163–177, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chudnovsky:2012:GC

[CS12] Maria Chudnovsky and Paul Seymour. Growing without cloning. SIAM Journal on Discrete Mathematics, 26(2): 860–880, ???? 2012. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Candela:2014:CRS

[CS14] Pablo Candela and Olof Sisask. Convergence results for systems of linear forms on cyclic groups and periodic nilsequences. SIAM Journal on Discrete Mathematics, 28 (2):786–810, ???? 2014. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chen:2018:SLM

[CS18a] William Y. C. Chen and Lisa H. Sun. Stanley’s lemma and multiple theta functions. SIAM Journal on Discrete Mathematics, 32(1): 474–504, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chudnovsky:2018:OHB

[CS18b] Maria Chudnovsky and Vaidy Sivaraman. Odd holes in bullfree graphs. SIAM Journal on Discrete Mathematics, 32 (2):951–955, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chudnovsky:2018:CSF

[CS18c] Maria Chudnovsky and Juraj Stacho. 3-colorable subclasses of P_{8}-free graphs. SIAM Journal on Discrete Mathematics, 32(2):1111–1138, ???? 2018. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chen:2019:DCS

[CS19] Guantao Chen and Songling Shan. Dirac’s condition for spanning Halin subgraphs. SIAM Journal on Discrete Mathematics, 33(4):1997– 2022, ???? 2019. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Crew:2021:CMB

[CS21] Logan Crew and Sophie Spirkl. A complete multipartite basis for the chromatic symmetric function. SIAM Journal on Discrete Mathematics, 35(4):2647–2661, ???? 2021. CODEN SJDMEC.

ISSN 0895-4801 (print), 1095-7146 (electronic).

Cheriyan:2001:IAS

[CSS01] J. Cheriyan, A. Sebo, and Z. Szigeti. Improving on the 1.5-approximation of a smallest 2-edge connected spanning subgraph. SIAM Journal on Discrete Mathematics, 14(2):170–180, 2001. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/36207.

Chen:2013:EFG

[CSS13] Guantao Chen, Akira Saito, and Songling Shan. The existence of a 2-factor in a graph satisfying the local Chvátal– Erdős condition. SIAM Journal on Discrete Mathematics, 27(4):1788–1799, ???? 2013. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Czabarka:2017:IBT

[CSW17] Éva Czabarka, László A.

Székely, and Stephan Wagner. Inducibility in binary trees and crossings in random tanglegrams. SIAM Journal on Discrete Mathematics, 31 (3):1732–1750, ???? 2017. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chung:1993:CCQ

[CT93] Fan R. K. Chung and Prasad Tetali. Communication complexity and quasi randomness. SIAM Journal on Discrete Mathematics, 6(1):110– 123, February 1993. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Conlon:2021:RPP

[CT21] David Conlon and Mykhaylo Tyomkyn. Repeated patterns in proper colorings. SIAM Journal on Discrete Mathematics, 35(3):2249–2264, ???? 2021. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chang:2001:MSN

[CtJL01] Gerard J. Chang, Justie Su tzu Juan, and Daphne Der-Fen Liu. Minimum span of no-hole (r + 1)-distant colorings. SIAM Journal on Discrete Mathematics, 14(3):370–380, 2001. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/33945.

Cano:2014:UBC

[CTU14] Javier Cano, Csaba D. Tóth, and Jorge Urrutia. Upper bound constructions for untangling planar geometric graphs. SIAM Journal on Discrete Mathematics, 28(4):

1935–1943, ???? 2014. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Coppersmith:1993:CAR

[CTW93] Don Coppersmith, Prasad Tetali, and Peter Winkler. Collisions among random walks on a graph. SIAM Journal on Discrete Mathematics, 6(3):363–374, August 1993. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cole:2007:GKT

[CuKŠ07] Richard Cole, ukasz Kowalik, and Riste Škrekovski. A generalization of Kotzig’s theorem and its application. SIAM Journal on Discrete Mathematics, 21(1): 93–106, ???? 2007. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cheriyan:2007:AAN

[CV07] Joseph Cheriyan and Adrian Vetta. Approximation algorithms for network design with metric costs. SIAM Journal on Discrete Mathematics, 21(3):612–636, ???? 2007. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Costello:2009:CRD

[CV09] Kevin P. Costello and Van Vu. Concentration of random determinants and permanent estimators. SIAM Journal on Discrete Mathematics, 23(3): 1356–1371, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cambie:2022:MLS

[CvBdJdVK22] Stijn Cambie, Wouter Cames van Batenburg, Rémi de Joannis de Verclos, and Ross J. Kang. Maximizing line subgraphs of diameter at most t. SIAM Journal on Discrete Mathematics, 36 (2):939–950, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1437354.

Cartwright:1992:IGN

[CW92] Donald I. Cartwright and Wolfgang Woess. Infinite graphs with nonconstant Dirichlet finite harmonic functions. SIAM Journal on Discrete Mathematics, 5(3):380–385, August 1992. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Connelly:1996:SOR

[CW96] Robert Connelly and Walter Whiteley. Second-order rigidity and prestress stability for tensegrity frameworks. SIAM Journal on Discrete Mathematics, 9(3):453–491, August 1996. CODEN SJDMEC.

ISSN 0895-4801 (print), 1095-7146 (electronic).

Calkin:1998:NIS

[CW98] Neil J. Calkin and Herbert S.

Wilf. The number of independent sets in a grid graph. SIAM Journal on Discrete Mathematics, 11(1): 54–60, February 1998. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cranston:2009:CRG

[CW09] Daniel W. Cranston and Douglas B. West. Classes of 3regular graphs that are (7, 2)- edge-choosable. SIAM Journal on Discrete Mathematics, 23(2):872–881, ???? 2009. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chen:2014:ICM

[CW14a] Rong Chen and Geoff Whittle. Intertwining connectivity in matroids. SIAM Journal on Discrete Mathematics, 28(3):1402–1404, ???? 2014. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chen:2014:IFQ

[CW14b] Zhixiong Chen and Arne Winterhof. Interpolation of Fermat quotients. SIAM Journal on Discrete Mathematics, 28(1):1–7, ???? 2014. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chien:2016:CPU

[CW16] Tuan-Yow Chien and Shayne Waldron. A characterization of projective unitary equivalence of finite frames and applications. SIAM Journal on Discrete Mathematics, 30(2): 976–994, ???? 2016. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Caro:2000:CDS

[CWY00] Yair Caro, Douglas B. West, and Raphael Yuster. Connected domination and spanning trees with many leaves. SIAM Journal on Discrete Mathematics, 13 (2):202–211, 2000. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/35378.

Cao:2021:BVN

[CWY21] Weiting Cao, Douglas B. West, and Yan Yang. On the bar visibility number of complete bipartite graphs. SIAM Journal on Discrete Mathematics, 35(3):2234–2248, ???? 2021. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chen:2010:MSF

[CWYZ10] William Y. C. Chen, Andrew Y. Z. Wang, Catherine H.

Yan, and Alina F. Y. Zhao. Mixed statistics on 01-fillings of Moon polyominoes. SIAM Journal on Discrete Mathematics, 24(4):1272–1290, ???? 2010. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chen:2008:LPD

[CX08] William Y. C. Chen and Deheng Xu. Labeled partitions and the q-derangement numbers. SIAM Journal on Discrete Mathematics, 22(3): 1099–1104, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Curran:2003:NCC

[CY03] Sean Curran and Xingxing Yu. Nonseparating cycles in 4-connected graphs. SIAM Journal on Discrete Mathematics, 16(4):616–629, 2003. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/39518.

Chang:2008:EBD

[CY08] Mei-Chu Chang and Chui Zhi Yao. An explicit bound on double exponential sums related to Diffie–Hellman distributions. SIAM Journal on Discrete Mathematics, 22(1): 348–359, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chen:2012:PNR

[CY12] Yong-Gao Chen and Quan-Hui Yang. On a problem of Nathanson related to minimal additive complements. SIAM Journal on Discrete Mathematics, 26(4):1532–1536, ???? 2012. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chiba:2018:DFD

[CY18] Shuya Chiba and Tomoki Yamashita. On directed 2-factors in digraphs and 2-factors containing perfect matchings in bipartite graphs. SIAM Journal on Discrete Mathematics, 32(1): 394–409, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cranston:2020:SGN

[CY20] Daniel W. Cranston and Matthew P. Yancey. Sparse graphs are near-bipartite. SIAM Journal on Discrete Mathematics, 34(3):1725– 1768, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cranston:2021:VPI

[CY21] Daniel W. Cranston and Matthew P. Yancey. Vertex partitions into an independent set and a forest with each component small. SIAM Journal on Discrete Mathematics, 35(3):1769–1791, ????

2021. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Charpin:1997:CWD

[CZ97] Pascale Charpin and Victor Zinoviev. On coset weight distributions of the 3-error-correcting BCH-codes. SIAM Journal on Discrete Mathematics, 10(1):128–145, February 1997. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Chen:2005:RVS

[CZLfW05] Robert W. Chen, Alan Zame, Chien-Tai Lin, and Hsiu fen Wu. A random version of Shepp’s urn scheme. SIAM Journal on Discrete Mathematics, 19(1):149–164, 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/41809.

Chen:1998:OAP

[CZOS98] Robert W. Chen, Alan Zame, Andrew M. Odlyzko, and Larry A. Shepp. An optimal acceptance policy for an urn scheme. SIAM Journal on Discrete Mathematics, 11 (2):183–195, May 1998. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dolecek:2010:REC

[DA10] Lara Dolecek and Venkat Anantharam. Repetition error correcting sets: Explicit constructions and prefixing methods. SIAM Journal on Discrete Mathematics, 23(4): 2120–2146, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dahlhaus:1993:FPR

[Dah93] Elias Dahlhaus. Fast parallel recognition of ultrametrics and tree metrics. SIAM Journal on Discrete Mathematics, 6(4):523–532, November 1993. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hausen:2010:UTC

[dAHFdFK10] Rodrigo de A. Hausen, Luerbio Faria, Celina M. H. de Figueiredo, and Luis Antonio B. Kowada. Unitary toric classes, the reality and desire diagram, and sorting by transpositions. SIAM Journal on Discrete Mathematics, 24 (3):792–807, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Daneshgar:2001:FSC

[Dan01]  Amir Daneshgar.  Forc-

ing structures and cliques in uniquely vertex colorable graphs. SIAM Journal on Discrete Mathematics, 14 (4):433–445, 2001. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/30499.

Dang:2009:ASH

[Dan09] Chuangyin Dang. An arbitrary starting homotopy-like simplicial algorithm for computing an integer point in a class of polytopes. SIAM Journal on Discrete Mathematics, 23(2):609–633, ???? 2009. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dasgupta:1999:SMS

[Das99] Samit Dasgupta. On the size of minimum super Arrovian domains. SIAM Journal on Discrete Mathematics, 12 (4):524–534, November 1999. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/33252.

Daubel:2022:IUB

[Däu22] Karl Däubel. An improved upper bound for the ring loading problem. SIAM Journal on Discrete Mathematics, 36(2):867–887, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1319395.

deBerg:2024:SAA

[dBSS24] Mark de Berg, Arpan Sadhukhan, and Frits Spieksma.

Stable approximation algorithms for the dynamic broadcast range-assignment problem. SIAM Journal on Discrete Mathematics, 38(1): 790–827, February 2024. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

deCarvalho:2013:NPM

[dCLM13] Marcelo H. de Carvalho, Cláudio L. Lucchesi, and U. S. R. Murty. On the number of perfect matchings in a bipartite graph. SIAM Journal on Discrete Mathematics, 27(2):940–958, ???? 2013. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Silva:2020:NTD

[dCST20] Marcel K. de Carli Silva and Levent Tunçel. A notion of total dual integrality for convex, semidefinite, and extended formulations. SIAM Journal on Discrete Mathematics, 34(1):470–496, ???? 2020. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ding:2013:CGP

[DD13] Guoli Ding and Stan Dziobiak. On 3-connected graphs of path-width at most three. SIAM Journal on Discrete Mathematics, 27(3):1514– 1526, ???? 2013. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Delucchi:2015:BCL

[DD15] Emanuele Delucchi and Martin Dlugosch. Bergman complexes of lattice path matroids. SIAM Journal on Discrete Mathematics, 29(4): 1916–1930, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dabrowski:2021:TPM

[DDJ^{+}21] Konrad K. Dabrowski, François Dross, Jisu Jeong, Mamadou M. Kanté, O joung Kwon, Sang il Oum, and Daniël Paulusma. Tree pivot-minors and linear rankwidth. SIAM Journal on Discrete Mathematics, 35(4): 2922–2945, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dao:2020:MCM

[DDL20] Hailong Dao, Joseph Doolittle, and Justin Lyle. Minimal Cohen–Macaulay simplicial complexes. SIAM Journal on Discrete Mathematics, 34(3):1602–1608, ???? 2020. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

De:2016:RKI

[DDS16] Anindya De, Ilias Diakonikolas, and Rocco A. Servedio.

A robust Khintchine inequality, and algorithms for computing optimal constants in Fourier analysis and highdimensional geometry. SIAM Journal on Discrete Mathematics, 30(2):1058–1094, ???? 2016. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

DeAmbroggio:2024:SPC

[De 24] Umberto De Ambroggio. A simple path to component sizes in critical random graphs. SIAM Journal on Discrete Mathematics, 38(2): 1492–1525, May 2024. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Donald:1993:SSO

[DE93] John Donald and John Elwin. On the structure of the strong orientations of a graph. SIAM Journal on Discrete Mathematics, 6(1):30–43, February 1993. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dosa:2011:POS

[DE11] György Dósa and Leah Epstein. Preemptive online scheduling with reordering. SIAM Journal on Discrete Mathematics, 25(1): 21–49, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.

siam.org/sidma/resource/
1/sjdmec/v25/i1/p21_s1.

Diestel:2017:DTB

[DEE17] Reinhard Diestel, Philipp Eberenz, and Joshua Erde. Duality theorems for blocks and tangles in graphs. SIAM Journal on Discrete Mathematics, 31(3):1514–1528, ???? 2017. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dudek:2019:RVG

[DEF19] Andrzej Dudek, Sean English, and Alan M. Frieze. A random variant of the game of plates and olives. SIAM Journal on Discrete Mathematics, 33(3):1216–1227, ???? 2019. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

DeVos:2007:CCP

[DEG^{+}07] Matt DeVos, Javad Ebrahimi, Mohammad Ghebleh, Luis Goddyn, Bojan Mohar, and Reza Naserasr. Circular coloring the plane. SIAM Journal on Discrete Mathematics, 21(2):461–465, ???? 2007. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Diao:2020:IRL

[DEH20] Yuanan Diao, Claus Ernst, and Gábor Hetyei. Invariants of rational links represented by reduced alternating diagrams. SIAM Journal on Discrete Mathematics, 34(3): 1944–1968, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Deitmar:2015:IZF

[Dei15] Anton Deitmar. Ihara zeta functions of infinite weighted graphs. SIAM Journal on Discrete Mathematics, 29(4): 2100–2116, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dujmovic:2020:MCG

[DEJ^{+}20] Vida Dujmović, David Eppstein, Gwenaël Joret, Pat Morin, and David R. Wood. Minor-closed graph classes with bounded layered pathwidth. SIAM Journal on Discrete Mathematics, 34(3): 1693–1709, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Do:2022:PGS

[DEK22] Tuan A. Do, Joshua Erde, and Mihyun Kang. Planarity and genus of sparse random bipartite graphs. SIAM Journal on Discrete Mathematics, 36(2):1394–1415, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1341817.

Dogan:2022:MSZ

[DEMT22] M. Levent Dogan, Alperen A.

Ergür, Jake D. Mundo, and Elias Tsigaridas. The multivariate Schwartz–Zippel lemma. SIAM Journal on Discrete Mathematics, 36 (2):888–910, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1333869.

Dereniowski:2012:PCP

[Der12] Dariusz Dereniowski. From pathwidth to connected pathwidth. SIAM Journal on Discrete Mathematics, 26(4): 1709–1732, ???? 2012. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dujmovic:2017:SGL

[DEW17] Vida Dujmović, David Eppstein, and David R. Wood. Structure of graphs with locally restricted crossings. SIAM Journal on Discrete Mathematics, 31(2):805–824, ???? 2017. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dewar:2020:IRN

[Dew20] Sean Dewar. Infinitesimal rigidity in normed planes. SIAM Journal on Discrete Mathematics, 34(2):1205– 1231, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Desmedt:1994:HZK

[DF94] Yvo G. Desmedt and Yair Frankel. Homomorphic zeroknowledge threshold schemes over any finite Abelian group. SIAM Journal on Discrete Mathematics, 7(4):667–679, November 1994. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Doignon:2002:FWO

[DF02] Jean-Paul Doignon and Samuel Fiorini. Facets of the weak order polytope derived from the induced partition projection. SIAM Journal on Discrete Mathematics, 15(1):112–121, 2002. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/36993.

Dougherty:2004:DDP

[DF04] Randall Dougherty and Vance Faber. The degree-diameter problem for several varieties of Cayley graphs I: The Abelian case. SIAM Journal on Discrete Mathematics, 17(3):478–519, 2004. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/37289.

Doerr:2010:HDD

[DF10] Benjamin Doerr and Mahmoud Fouz. Hereditary discrepancies in different numbers of colors II. SIAM Journal on Discrete Mathematics, 24(3):1205–1213, ???? 2010. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Damasdi:2021:CNY

[DFG^{+}21] Gábor Damásdi, Stefan Felsner, António Girão, Balázs Keszegh, David Lewis, Dániel T. Nagy, and Torsten Ueckerdt. On covering numbers, Young diagrams, and the local dimension of posets. SIAM Journal on Discrete Mathematics, 35(2): 915–927, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Demaine:2004:BPL

[DFHT04] Erik D. Demaine, Fedor V.

Fomin, Mohammad Taghi Hajiaghayi, and Dimitrios M. Thilikos. Bidimensional parameters and local treewidth. SIAM Journal on Discrete Mathematics, 18(3): 501–511, 2004. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/43341.

Diemunsch:2015:ETD

[DFJS15] Jennifer Diemunsch, Michael Ferrara, Sogol Jahanbekam, and James M. Shook. Extremal theorems for degree sequence packing and the two-color discrete tomography problem. SIAM Journal on Discrete Mathematics, 29(4):2088–2099, ???? 2015. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Drmota:2011:ASS

[DFK^{+}11] Michael Drmota, Éric Fusy, Mihyun Kang, Veronika Kraus, and Juanjo Rué. Asymptotic study of subcritical graph classes. SIAM Journal on Discrete Mathematics, 25(4):1615–1651, ???? 2011. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v25/i4/p1615_s1.

Dvorak:2021:PAS

[DFK^{+}21] Pavel Dvorák, Andreas E. Feldmann, Dusan Knop, Tomás Masarík, Tomás Toufar, and Pavel Veselý. Parameterized approximation schemes for Steiner trees with small number of Steiner vertices. SIAM Journal on Discrete Mathematics, 35(1): 546–574, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Deschamps:2023:STM

[DFK^{+}23] Quentin Deschamps, Carl Feghali, Frantisek Kardos, Clément Legrand-Duchesne, and Théo Pierron. Strengthening a theorem of Meyniel. SIAM Journal on Discrete Mathematics, 37(2):604–611, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:/

/epubs.siam.org/doi/10.
1137/22M1474394.

deFigueiredo:2004:OBF

[dFM04] Celina M. H. de Figueiredo and Frédéric Maffray. Optimizing bull-free perfect graphs. SIAM Journal on Discrete Mathematics, 18 (2):226–240, 2004. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/33923.

Dudek:2015:RCR

[DFT15] Andrzej Dudek, Alan M.

Frieze, and Charalampos E. Tsourakakis. Rainbow connection of random regular graphs. SIAM Journal on Discrete Mathematics, 29(4): 2255–2266, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dubroff:2021:NED

[DFX21] Quentin Dubroff, Jacob Fox, and Max Wenqiang Xu. A note on the Erdős distinct subset sums problem. SIAM Journal on Discrete Mathematics, 35(1):322–324, ???? 2021. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dvorak:2008:PFH

[DG08] Tomáš Dvořák and Petr Gregor. Partitions of faulty hypercubes into paths with prescribed endvertices. SIAM Journal on Discrete Mathematics, 22(4):1448–1461, ???? 2008. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Didimo:2010:USU

[DGL10] Walter Didimo, Francesco Giordano, and Giuseppe Liotta. Upward spirality and upward planarity testing. SIAM Journal on Discrete Mathematics, 23(4): 1842–1899, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

DiGiacomo:2011:UTB

[DGL11] Emilio Di Giacomo, Francesco Giordano, and Giuseppe Liotta. Upward topological book embeddings of DAGs. SIAM Journal on Discrete Mathematics, 25(2): 479–489, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v25/i2/p479_s1.

Durr:2012:RCG

[DGM12] Christoph Dürr, Flavio Guiñez, and Martin Matamala. Reconstructing 3-colored grids from horizontal and vertical projections is NP-hard: a solution to the 2-atom problem in discrete tomography. SIAM Journal on Discrete Mathematics, 26(1): 330–352, ???? 2012. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Debski:2020:AMR

[DGN^{+}20] Michal Debski, Jarosław Grytczuk, Barbara Nayar, Urszula Pastwa, Joanna Sokół, Michał Tuczyński, Przemysław Wenus, and Krzysztof Wesek. Avoiding multiple repetitions in Euclidean spaces. SIAM Journal on Discrete Mathematics, 34(1):40–52, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

deGevigney:2013:MBP

[dGNS13] Olivier Durand de Gevigney, Viet-Hang Nguyen, and Zoltán Szigeti. Matroidbased packing of arborescences. SIAM Journal on Discrete Mathematics, 27(1):

567–574, ???? 2013. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dulio:2006:DPX

[DGP06] Paolo Dulio, Richard J. Gardner, and Carla Peri. Discrete point X-rays. SIAM Journal on Discrete Mathematics, 20(1):171–188, January 2006. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Doty:1996:CGN

[DGS96] L. L. Doty, R. J. Goldstone, and C. L. Suffel. Cayley graphs with neighbor connectivity one. SIAM Journal on Discrete Mathematics, 9 (4):625–642, November 1996. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

deGraaf:2005:PRS

[dGV05] Mart de Graaf and Paul Valiant. Polynomial representations of symmetric partial Boolean functions. SIAM Journal on Discrete Mathematics, 19(2):481–488, 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/43356.

deCaen:1989:IDC

[dH89] D. de Caen and D. G. Hoffman. Impossibility of decomposing the complete graph on n points into n − 1 isomorphic complete bipartite graphs. SIAM Journal on Discrete Mathematics, 2(1): 48–50, February 1989. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Du:1990:OAS

[DH90] D. Z. Du and F. K. Hwang.

Optimal assembly of an s-stage k-out-of-n system. SIAM Journal on Discrete Mathematics, 3(3):349–354, August 1990. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dress:1991:BSU

[DH91] Andreas W. M. Dress and Timothy F. Havel. Bound smoothing under chirality constraints. SIAM Journal on Discrete Mathematics, 4 (4):535–549, November 1991. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hou:2004:NPN

[dH04] Xiang dong Hou. A note on the proof of Niho’s Conjecture. SIAM Journal on Discrete Mathematics, 18 (2):313–319, 2004. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/43281.

Drmota:2005:BPT

[DH05] Michael Drmota and Hsien-Kuei Hwang. Bimodality and phase transitions in the profile variance of random binary search trees. SIAM Journal on Discrete Mathematics, 19(1):19–45, 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/44013.

Dukes:2020:MDR

[DH20a] Peter J. Dukes and Daniel Horsley. On the minimum degree required for a triangle decomposition. SIAM Journal on Discrete Mathematics, 34(1):597–610, ???? 2020. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dvorak:2020:LCE

[DH20b] Zdenek Dvorák and Xiaolan Hu. (3a: a)-list-colorability of embedded graphs of girth at least five. SIAM Journal on Discrete Mathematics, 34(4): 2137–2165, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dvorak:2020:FCP

[DH20c] Zdenek Dvorák and Xiaolan Hu. Fractional coloring of planar graphs of girth five. SIAM Journal on Discrete Mathematics, 34(1):

538–555, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dujmovic:2013:LTA

[DHJ^{+}13] Vida Dujmović, Daniel J.

Harvey, Gwenaël Joret, Bruce Reed, and David R. Wood. A linear-time algorithm for finding a complete graph minor in a dense graph. SIAM Journal on Discrete Mathematics, 27(4):1770–1774, ???? 2013. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Du:2002:OCC

[DHJN02] Ding-Zhu Du, Frank K.

Hwang, Xiaohua Jia, and Hung Q. Ngo. Optimal consecutive-k-out-of-n: G cycle for n ≤ 2k + 1. SIAM Journal on Discrete Mathematics, 15(3):305–316, 2002. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/37504.

Dress:2011:BCV

[DHKM11] A. W. M. Dress, K. T. Huber, J. Koolen, and V. Moulton. Blocks and cut vertices of the Buneman graph. SIAM Journal on Discrete Mathematics, 25(4):1902– 1919, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.

                                        siam.org/sidma/resource/
                                        1/sjdmec/v25/i4/p1902_s1.

Dolev:2002:NPH

[DHL^{+}02] Danny Dolev, Yuval Harari, Nathan Linial, Noam Nisan, and Michal Parnas. Neighborhood preserving hashing and approximate queries. SIAM Journal on Discrete Mathematics, 15(1):73– 85, 2002. CODEN SJD-MEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/31809.

Dorbec:2013:GPD

[DHL^{+}13] Paul Dorbec, Michael A. Henning, Christian Löwenstein, Mickael Montassier, and André Raspaud. Generalized power domination in regular graphs. SIAM Journal on Discrete Mathematics, 27(3): 1559–1574, ???? 2013. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Devadoss:2014:PCT

[DHS14] Satyan L. Devadoss, Daoji Huang, and Dominic Spadacene. Polyhedral covers of tree space. SIAM Journal on Discrete Mathematics, 28(3): 1508–1514, ???? 2014. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Demaine:2006:BTB

[DHT06] Erik D. Demaine, MohammadTaghi Hajiaghayi, and Dimitrios M. Thilikos. The bidimensional theory of bounded-genus graphs. SIAM Journal on Discrete Mathematics, 20(2):357–371, January 2006. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dor:2001:LBS

[DHUZ01] Dorit Dor, Johan Håstad, Staffan Ulfberg, and Uri Zwick. On lower bounds for selecting the median. SIAM Journal on Discrete Mathematics, 14(3):299–311, 2001. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/30948.

Du:1999:IH

[DHX99] Ding-Zhu Du, Frank K.

Hwang, and Guoliang Xue. Interconnecting highways. SIAM Journal on Discrete Mathematics, 12(2):252–261, May 1999. CODEN SJD-MEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/32653.

Davies:2022:RPT

[DI22] Ewan Davies and Freddie Illingworth. The χ-Ramsey problem for trianglefree graphs. SIAM Journal on Discrete Mathematics, 36 (2):1124–1134, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1437573.

Dieker:2010:IRW

[Die10] A. B. Dieker. Interlacings for random walks on weighted graphs and the interchange process. SIAM Journal on Discrete Mathematics, 24(1): 191–206, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dinitz:2006:FRT

[Din06] Michael Dinitz. Full rank tilings of finite Abelian groups. SIAM Journal on Discrete Mathematics, 20(1): 160–170, January 2006. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ding:2013:CCS

[Din13] Cunsheng Ding. Cyclic codes from some monomials and trinomials. SIAM Journal on Discrete Mathematics, 27(4): 1977–1994, ???? 2013. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dong:2011:ZFI

[DJ11] F. M. Dong and Bill Jackson. A zero-free interval for chromatic polynomials of nearly 3-connected plane graphs. SIAM Journal on Discrete Mathematics, 25(3): 1103–1118, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v25/i3/p1103_s1.

Djidjev:2006:LTA

[Dji06] Hristo N. Djidjev. A lineartime algorithm for finding a maximal planar subgraph. SIAM Journal on Discrete Mathematics, 20(2): 444–462, January 2006. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

DAli:2023:GVS

[DJKKV23] Alessio D’Alì, Martina Juhnke-Kubitzke, Daniel Köhne, and Lorenzo Venturello. On the gamma-vector of symmetric edge polytopes. SIAM Journal on Discrete Mathematics, 37(2):487–515, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1492799.

Diaz:2019:ECR

[DJKO19] Alberto Espuny Díaz, Felix Joos, Daniela Kühn, and Deryk Osthus. Edge correlations in random regular hypergraphs and applications to subgraph testing. SIAM Journal on Discrete Mathematics, 33(4):1837–1863, ???? 2019. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dvorak:2009:MNC

[DJKP09] Zdenĕk Dvořák, Stanislav Jendrol’, Daniel Král’, and Gyula Pap. Matchings and nonrainbow colorings. SIAM Journal on Discrete Mathematics, 23(1):344–348, ???? 2009. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dujmovic:2018:COT

[DJM^{+}18a] Vida Dujmović, Gwenaël Joret, Pat Morin, Sergey Norin, and David R. Wood. Corrigendum: Orthogonal Tree Decompositions of Graphs. SIAM Journal on Discrete Mathematics, 32(4): 3003–3004, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). See [DJM^{+}18b].

Dujmovic:2018:OTD

[DJM^{+}18b] Vida Dujmović, Gwenaël Joret, Pat Morin, Sergey Norin, and David R. Wood. Orthogonal tree decompositions of graphs. SIAM Journal on Discrete Mathematics, 32(2):839–863, ???? 2018. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). See corrigendum [DJM^{+}18a].

deJong:2016:NPT

[dJMS16] J. V. de Jong, J. C. McLeod, and M. Steel. Neighborhoods of phylogenetic trees: Exact and asymptotic counts. SIAM Journal on Discrete Mathematics, 30(4):2265–2287, ???? 2016. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dyer:2021:CWI

[DJMV21] Martin Dyer, Mark Jerrum, Haiko Müller, and Kristina Vusković. Counting weighted independent sets beyond the permanent. SIAM Journal on Discrete Mathematics, 35(2): 1503–1524, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dzindzalieta:2012:OPI

[DJS12] D. Dzindzalieta, T. Juskevicius, and M. Sileikis. Optimal probability inequalities for random walks related to problems in extremal combinatorics. SIAM Journal on Discrete Mathematics, 26(2): 828–837, ???? 2012. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dumitrescu:2015:COI

[DJT15] Adrian Dumitrescu, Minghui Jiang, and Csaba D. Tóth.

Computing opaque interior barriers à la Shermer. SIAM Journal on Discrete Mathematics, 29(3):1372–1386, ???? 2015. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dujmovic:2012:IBF

[DJW12] Vida Dujmović, Gwenaël Joret, and David R. Wood. An improved bound for firstfit on posets without two long incomparable chains. SIAM Journal on Discrete Mathematics, 26(3):1068–1075, ???? 2012. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Delsarte:1989:LRM

[DK89] P. Delsarte and Y. Kamp. Low rank matrices with a given sign pattern. SIAM Journal on Discrete Mathematics, 2(1):51–63, February 1989. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Deogun:2002:DPG

[DK02] Jitender S. Deogun and Dieter Kratsch. Dominating pair graphs. SIAM Journal on Discrete Mathematics, 15(3):353–366, 2002. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/36711.

Dong:2006:GHN

[DK06] F. M. Dong and K. M.

Koh. On graphs having no chromatic zeros in (1, 2). SIAM Journal on Discrete Mathematics, 20(3): 799–810, January 2006. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dong:2010:ZFI

[DK10] F. M. Dong and K. M.

Koh. On zero-free intervals in (1, 2) of chromatic polynomials of some families of graphs. SIAM Journal on Discrete Mathematics, 24(2): 370–378, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dvorak:2014:SIM

[DK14] Zdenek Dvorák and Tereza Klimosová. Strong immersions and maximum degree. SIAM Journal on Discrete Mathematics, 28(1): 177–187, ???? 2014. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Devlin:2016:SEK

[DK16] Pat Devlin and Jeff Kahn. On “stability” in the Erdős–Ko– Rado theorem. SIAM Journal on Discrete Mathematics, 30(2):1283–1289, ???? 2016. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dvorak:2022:PMR

[DK22] Vojtech Dvorak and Ohad Klein. Probability mass of Rademacher sums beyond one standard deviation. SIAM Journal on Discrete Mathematics, 36 (3):2393–2410, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1428212.

Disser:2024:FSM

[DKLW24] Yann Disser, Max Klimm, Annette Lutz, and David Weckbecker. Fractionally subadditive maximization under an incremental knapsack constraint with applications to incremental flows. SIAM Journal on Discrete Mathematics, 38(1):764–789, February 2024. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dellamonica:2012:DAF

[DKM^{+}12] Domingos Dellamonica, Subrahmanyam Kalyanasundaram, Daniel Martin, Vojtech Rödl, and Asaf Shapira. A deterministic algorithm for the Frieze–Kannan regularity lemma. SIAM Journal on Discrete Mathematics, 26 (1):15–29, ???? 2012. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v26/i1/p15_s1.

DeBiasio:2015:AOH

[DKM^{+}15] Louis DeBiasio, Daniela Kühn, Theodore Molla, Deryk Osthus, and Amelia Taylor. Arbitrary orientations of Hamilton cycles in digraphs. SIAM Journal on Discrete Mathematics, 29(3): 1553–1584, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

deKlerk:2006:IBC

[dKMP^{+}06] E. de Klerk, J. Maharry, D. V. Pasechnik, R. B. Richter, and G. Salazar. Improved bounds for the crossing numbers of K_{m,n} and K_{n}. SIAM Journal on Discrete Mathematics, 20(1): 189–202, January 2006. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Disser:2017:PKU

[DKMS17] Yann Disser, Max Klimm, Nicole Megow, and Sebastian Stiller. Packing a knapsack of unknown capacity. SIAM Journal on Discrete Mathematics, 31(3):1477–1497, ???? 2017. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

deKlerk:2013:ILB

[dKPS13] E. de Klerk, D. V. Pasechnik, and G. Salazar. Improved lower bounds on book crossing numbers of complete graphs. SIAM Journal on Discrete Mathematics, 27(2): 619–633, ???? 2013. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dellamonica:2012:URG

[DKRR12] Domingos Dellamonica, Jr., Yoshiharu Kohayakawa, Vojtech Rödl, and Andrzej Ruciński. Universality of random graphs. SIAM Journal on Discrete Mathematics, 26 (1):353–374, ???? 2012. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Drmota:2016:CCC

[DKS16] Michael Drmota, Manuel Kauers, and Lukas Spiegelhofer. On a conjecture of Cusick concerning the sum of digits of n and n + t. SIAM Journal on Discrete Mathematics, 30(2):621–649, ???? 2016. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dowden:2018:ERG

[DKS18] Chris Dowden, Mihyun Kang, and Philipp Sprüssel. The evolution of random graphs on surfaces. SIAM Journal on Discrete Mathematics, 32 (1):695–727, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dumitrescu:2010:VBP

[DKSŻ10] Adrian Dumitrescu, Howi Kok, Ichiro Suzuki, and Pawe Żylinski. Visionbased pursuit-evasion in a grid. SIAM Journal on Discrete Mathematics, 24(3): 1177–1204, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dvorak:2022:TSS

[DKT22] Pavel Dvorák, Dusan Knop, and Tomás Toufar. Target set selection in dense graph classes. SIAM Journal on Discrete Mathematics, 36(1):536–572, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1337624.

DeCarvalho:2020:BNG

[DKWL20] Marcelo H. De Carvalho, Nishad Kothari, Xiumei Wang, and Yixun Lin. Birkhoff–von Neumann graphs that are PM-compact. SIAM Journal on Discrete Mathematics, 34(3):1769–1790, ???? 2020. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Du:1989:CSP

[DL89a] Jianzhong Du and Joseph Y.- T. Leung. Complexity of scheduling parallel task systems. SIAM Journal on Discrete Mathematics, 2(4):473– 487, November 1989. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Du:1989:STS

[DL89b] Jianzhong Du and Joseph Y.- T. Leung. Scheduling treestructured tasks on two processors to minimize schedule length. SIAM Journal on Discrete Mathematics, 2(2): 176–196, May 1989. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ding:2012:CTC

[DL12] Guoli Ding and Cheng Liu. A chain theorem for 3^{+}- connected graphs. SIAM Journal on Discrete Mathematics, 26(1):102–113, ???? 2012. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dvorak:2014:CGS

[DL14] Zdenek Dvorák and Bernard Lidický. 4-critical graphs on surfaces without contractible (≤ 4)-cycles. SIAM Journal on Discrete Mathematics, 28 (1):521–552, ???? 2014. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dvorak:2017:FSC

[DL17] Zdenek Dvorák and Bernard Lidický. Fine structure of 4-critical triangle-free graphs II. Planar triangle-free graphs with two precolored 4-cycles. SIAM Journal on Discrete Mathematics, 31(2):865–874, ???? 2017. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dunik:2018:CTA

[DL18a] Barbora Duník and Robert Lukotka. Cubic TSP: a 1.3approximation. SIAM Journal on Discrete Mathematics, 32(3):2094–2114, ???? 2018. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dvorak:2018:FSCb

[DL18b] Zdenek Dvorák and Bernard Lidický. Fine structure of 4critical triangle-free graphs I. Planar graphs with two triangles and 3-colorability of chains. SIAM Journal on Discrete Mathematics, 32(3): 1775–1805, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dvorak:2018:FSCa

[DL18c] Zdenek Dvorák and Bernard Lidický. Fine structure of 4-critical triangle-free graphs III. General surfaces. SIAM Journal on Discrete Mathematics, 32(1):94–105, ???? 2018. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

DeBiasio:2019:STF

[DL19] Louis DeBiasio and Allan Lo. Spanning trees with few branch vertices. SIAM Journal on Discrete Mathematics, 33(3):1503–1520, ???? 2019. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

deLauney:2009:MNE

[dLL09] Warwick de Launey and David A. Levin. (1, −1)- matrices with near-extremal properties. SIAM Journal on Discrete Mathematics, 23(3): 1422–1440, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Deza:2018:ODS

[DLMO18] Antoine Deza, Asaf Levin, Syed M. Meesum, and Shmuel Onn. Optimization over degree sequences. SIAM Journal on Discrete Mathematics, 32(3):2067–2079, ???? 2018. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

DeBiasio:2021:TTT

[DLMT21] Louis DeBiasio, Allan Lo, Theodore Molla, and Andrew Treglown. Transitive tournament tilings in oriented graphs with large minimum total degree. SIAM Journal on Discrete Mathematics, 35 (1):250–266, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dvorak:2010:CTF

[DLŠ10] Zdeněk Dvořák, Bernard Lidický, and Riste Škrekovski. 3-choosability of trianglefree planar graphs with constraints on 4-cycles. SIAM Journal on Discrete Mathematics, 24(3):934–945, ???? 2010. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dvorak:2011:GTC

[DLS11] Zdenek Dvorák, Bernard Lidický, and Riste Skrekovski. Graphs with two crossings are 5-choosable. SIAM Journal on Discrete Mathematics, 25(4):1746–1753, ???? 2011. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v25/i4/p1746_s1.

Du:1988:MSI

[DM88] Ding Zhu Du and Zevi Miller.

Matroids and subset interconnection design. SIAM Journal on Discrete Mathematics, 1 (4):416–424, November 1988. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dinwoodie:2003:SII

[DM03] I. H. Dinwoodie and Edward Mosteig. Statistical inference for internal network reliability with spatial dependence. SIAM Journal on Discrete Mathematics, 16(4):663–674, 2003. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/39081.

deMier:2007:NFF

[dM07] Anna de Mier. A natural family of flag matroids. SIAM Journal on Discrete Mathematics, 21(1):130–140, ???? 2007. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dragan:2011:NGA

[DM11] Feodor F. Dragan and Martin Matamala. Navigating in a graph by aid of its spanning tree metric. SIAM Journal on Discrete Mathematics, 25 (1):306–332, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v25/i1/p306_s1.

Dick:2013:FCW

[DM13] Josef Dick and Makoto Matsumoto. On the fast computation of the weight enumerator polynomial and the t value of digital nets over finite Abelian groups. SIAM Journal on Discrete Mathematics, 27(3):1335–1359, ???? 2013. CODEN SJDMEC.

ISSN 0895-4801 (print), 1095-7146 (electronic).

Durocher:2015:PTE

[DM15] Stephane Durocher and Debajyoti Mondal. Plane 3trees: Embeddability and approximation. SIAM Journal on Discrete Mathematics, 29 (1):405–420, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dvorak:2017:LIS

[DM17] Zdenek Dvorák and Matthias Mnich. Large independent sets in triangle-free planar graphs. SIAM Journal on Discrete Mathematics, 31(2): 1355–1373, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dau:2018:MSS

[DM18a] Hoang Dau and Olgica Milenkovic. MaxMinSum Steiner systems for access balancing in distributed storage. SIAM Journal on Discrete Mathematics, 32(3): 1644–1671, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Durocher:2018:RGT

[DM18b] Stephane Durocher and Debajyoti Mondal. Relating graph thickness to planar layers and bend complexity. SIAM Journal on Discrete Mathematics, 32(4): 2703–2719, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dyer:2019:CPM

[DM19] Martin Dyer and Haiko Muller. Counting perfect matchings and the switch chain. SIAM Journal on Discrete Mathematics, 33(3): 1146–1174, ???? 2019. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dvorak:2023:DFC

[DM23] Zdenek Dvorák and Bojan Mohar. On density of Z_{3}- flow-critical graphs. SIAM Journal on Discrete Mathematics, 37(2):699–717, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL

                                        https://epubs.siam.org/
                                        doi/10.1137/22M1496529.

Dorbec:2008:PDP

[DMKŠ08] Paul Dorbec, Michel Mollard, Sandi Klavžar, and Simon Špacapan. Power domination in product graphs. SIAM Journal on Discrete Mathematics, 22(2):554–567, ???? 2008. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Draisma:2012:LWD

[DMN12] Jan Draisma, Tyrrell B. McAllister, and Benjamin Nill. Lattice-width directions and Minkowski’s 3^{d}- theorem. SIAM Journal on Discrete Mathematics, 26(3): 1104–1107, ???? 2012. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dvorak:2021:SSI

[DMN21] Zdenek Dvorák, Rose Mc-Carty, and Sergey Norin. Sublinear separators in intersection graphs of convex shapes. SIAM Journal on Discrete Mathematics, 35(2): 1149–1164, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dorn:2013:EAE

[DMNW13] Frederic Dorn, Hannes Moser, Rolf Niedermeier, and Mathias Weller. Efficient algorithms for Eulerian extension and rural postman. SIAM Journal on Discrete Mathematics, 27(1):75–94, ???? 2013. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

deWerra:1993:ECS

[dMP93] D. de Werra, N. V. R. Mahadev, and U. N. Peled. Edge-chromatic scheduling with simultaneity constraints. SIAM Journal on Discrete Mathematics, 6(4):631–641, November 1993. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Diaz:2007:STH

[DMP07] Josep Díaz, Dieter Mitsche, and Xavier Pérez. Sharp threshold for Hamiltonicity of random geometric graphs. SIAM Journal on Discrete Mathematics, 21(1): 57–65, ???? 2007. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Daskalakis:2011:PBB

[DMR11] Constantinos Daskalakis, Elchanan Mossel, and Sebastien Roch. Phylogenies without branch bounds: Contracting the short, pruning the deep. SIAM Journal on Discrete Mathematics, 25(2):872–893, ???? 2011. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sidma/resource/
                                        1/sjdmec/v25/i2/p872_s1.

Dankelmann:2008:ADEb

[DMS08a] Peter Dankelmann, Simon Mukwembi, and Henda C. Swart. Average distance and edge-connectivity I. SIAM Journal on Discrete Mathematics, 22(1):92–101, ???? 2008. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dankelmann:2008:ADEa

[DMS08b] Peter Dankelmann, Simon Mukwembi, and Henda C. Swart. Average distance and edge-connectivity II. SIAM Journal on Discrete Mathematics, 21(4):1035–1052, ???? 2008. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dutta:2012:NLB

[DMS12] Kunal Dutta, Dhruv Mubayi, and C. R. Subramanian. New lower bounds for the independence number of sparse graphs and hypergraphs. SIAM Journal on Discrete Mathematics, 26(3):1134– 1147, ???? 2012. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Deza:2014:CAC

[DMS14] Antoine Deza, Frédéric Meunier, and Pauline Sarrabezolles. A combinatorial approach to colourful simplicial depth. SIAM Journal on Discrete Mathematics, 28(1):306–322, ???? 2014. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dallard:2021:TVC

[DMS21] Clément Dallard, Martin Milanic, and Kenny Storgel. Treewidth versus clique number. I. Graph classes with a forbidden structure. SIAM Journal on Discrete Mathematics, 35(4):2618–2646, ???? 2021. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dvorak:2016:SSS

[DN16] Zdenek Dvorák and Sergey Norin. Strongly sublinear separators and polynomial expansion. SIAM Journal on Discrete Mathematics, 30(2): 1095–1101, ???? 2016. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dragan:1999:CHF

[DNB99] Feodor F. Dragan, Falk Nicolai, and Andreas Brandstädt. Convexity and HHD-free graphs. SIAM Journal on Discrete Mathematics, 12 (1):119–135, February 1999. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/32171.

Dukes:2023:EBN

[DNR23] Peter Dukes, Jonathan Noel, and Abel Romer. Extremal bounds for 3-neighbor bootstrap percolation in dimensions two and three. SIAM Journal on Discrete Mathematics, 37(3):2088–2125, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL

https://epubs.siam.org/
doi/10.1137/22M1534195.

Dankelmann:1994:PTG

[DNS94] Peter Dankelmann, Thomas Niessen, and Ingo Schiermeyer. On path-tough graphs. SIAM Journal on Discrete Mathematics, 7(4): 571–584, November 1994. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dror:2008:COE

[DO08] Moshe Dror and James B. Orlin. Combinatorial optimization with explicit delineation of the ground set by a collection of subsets. SIAM Journal on Discrete Mathematics, 21(4):1019–1034, ???? 2008. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Do:2019:RCS

[Do19] Thao Do. Representation complexities of semialgebraic graphs. SIAM Journal on Discrete Mathematics, 33(4): 1864–1877, ???? 2019. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Bastos:2017:LHC

[dOBMS^{+}17] Josefran de Oliveira Bastos, Guilherme Oliveira Mota, Mathias Schacht, Jakob Schnitzer, and Fabian Schulenburg. Loose Hamiltonian cycles forced by large (k−2)-degree — approximate version. SIAM Journal on Discrete Mathematics, 31(4): 2328–2347, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Contiero:2019:SRT

[dOCHLO19] Lucas de Oliveira Contiero, Carlos Hoppen, Hanno Lefmann, and Knut Odermann. Stability results for two classes of hypergraphs. SIAM Journal on Discrete Mathematics, 33(4):2023–2040, ???? 2019. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Contiero:2021:RER

[dOCHLO21] Lucas de Oliveira Contiero, Carlos Hoppen, Hanno Lefmann, and Knut Odermann. Rainbow Erdős–Rothschild problem for the Fano plane. SIAM Journal on Discrete Mathematics, 35(3):1927– 1951, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dohmen:2003:IIE

[Doh03] Klaus Dohmen. Improved inclusion-exclusion identities and Bonferroni inequalities with reliability applications. SIAM Journal on Discrete Mathematics, 16 (1):156–171, 2003. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/39263.

Day:1994:SDH

[DOS94] D. P. Day, Ortrud R. Oellermann, and Henda C. Swart. Steiner distance-hereditary graphs. SIAM Journal on Discrete Mathematics, 7(3): 437–442, August 1994. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dowd:1988:QRE

[Dow88] Martin Dowd. Questions related to the Erdős–Turán conjecture. SIAM Journal on Discrete Mathematics, 1(1): 142–150, February 1988. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dowd:1991:EIP

[Dow91a]  Martin Dowd.  Erratum:

“An integer program for codes”. SIAM Journal on Discrete Mathematics, 4(4): 589, November 1991. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). See [Dow91b].

Dowd:1991:IPC

[Dow91b] Martin Dowd. An integer program for codes. SIAM Journal on Discrete Mathematics, 4(2):293–303, May 1991. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). See erratum [Dow91a].

Diks:1992:ASG

[DP92] Krzysztof Diks and Andrzej Pelc. Almost safe gossiping in bounded degree networks. SIAM Journal on Discrete Mathematics, 5(3): 338–344, August 1992. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Diks:1996:EGP

[DP96] Krzysztof Diks and Andrzej Pelc. Efficient gossiping by packets in networks with random faults. SIAM Journal on Discrete Mathematics, 9 (1):7–18, February 1996. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dosen:2015:PCG

[DP15] Kosta Dosen and Zoran Petrić. A planarity criterion for graphs. SIAM Journal on Discrete Mathematics, 29(4): 2160–2165, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dudek:2016:ATR

[DP16] Andrzej Dudek and Pawel Pralat. Acquaintance time of random graphs near connectivity threshold. SIAM Journal on Discrete Mathematics, 30(1):555–568, ???? 2016. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dudek:2017:SMR

[DP17] Andrzej Dudek and Pawel Pralat. On some multicolor Ramsey properties of random graphs. SIAM Journal on Discrete Mathematics, 31(3): 2079–2092, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Debsk:2022:FCS

[DPR22] Michał Debsk, Marta Piecyk, and Paweł Rzazewski. Faster 3-coloring of small-diameter graphs. SIAM Journal on Discrete Mathematics, 36 (3):2205–2224, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1447714.

Dourado:2010:HNT

[DPRS10] Mitre C. Dourado, Fábio Protti, Dieter Rautenbach, and Jayme L. Szwarcfiter. On the hull number of trianglefree graphs. SIAM Journal on Discrete Mathematics, 23(4): 2163–2172, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Diaz:2008:WCG

[DPSW08] J. Díaz, X. Pérez, M. J.

Serna, and N. C. Wormald. Walkers on the cycle and the grid. SIAM Journal on Discrete Mathematics, 22(2):

747–775, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ding:2015:PTF

[DQW^{+}15] Cunsheng Ding, Longjiang Qu, Qiang Wang, Jin Yuan, and Pingzhi Yuan. Permutation trinomials over finite fields with even characteristic. SIAM Journal on Discrete Mathematics, 29(1): 79–92, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dey:2004:CCU

[DR04] Bikash Kumar Dey and B. Sundar Rajan. Codes closed under arbitrary Abelian group of permutations. SIAM Journal on Discrete Mathematics, 18(1):1–18, 2004. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/41619.

dAzevedo:2014:BRH

[dR14] Antonio Breda d’Azevedo and Ilda Inácio Rodrigues. Bicontactual regular hypermaps. SIAM Journal on Discrete Mathematics, 28(1): 142–159, ???? 2014. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dellamonica:2019:PPS

[DR19] Domingos Dellamonica, Jr.

and Vojtech Rödl. Packing paths in Steiner triple systems. SIAM Journal on Discrete Mathematics, 33(3): 1669–1690, ???? 2019. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dascalu:2024:TGP

[DR24] Maria Dascălu and Annie Raymond. Tropicalizing the graph profile of some almoststars. SIAM Journal on Discrete Mathematics, 38(2): 1351–1368, April 2024. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dreher:2012:CGC

[Dre12] Deanna Dreher. Cycles in graphs and covers. SIAM Journal on Discrete Mathematics, 26(4):1797–1819, ???? 2012. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dross:2016:FTD

[Dro16] François Dross. Fractional triangle decompositions in graphs with large minimum degree. SIAM Journal on Discrete Mathematics, 30(1): 36–42, ???? 2016. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Deuineko:1998:STE

[DRW98] Vladimir G. Deuineko, Rüdiger Rudolf, and Gerhard J. Woeginger. Sometimes travelling is easy: The Master Tour Problem. SIAM Journal on Discrete Mathematics, 11(1): 81–93, February 1998. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

deWerra:1991:CCC

[dS91] D. de Werra and Ph. Solot. Compact cylindrical chromatic scheduling. SIAM Journal on Discrete Mathematics, 4(4):528–534, November 1991. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Drmota:1997:IPR

[DS97] Michael Drmota and Michèle Soria. Images and preimages in random mappings. SIAM Journal on Discrete Mathematics, 10(2):246–269, May 1997. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Daniel:2005:CGS

[DS05a] Philip Daniel and Charles Semple. A class of general supertree methods for nested taxa. SIAM Journal on Discrete Mathematics, 19(2):463–480, 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/44146.

Dumitriu:2005:TBL

[DS05b] Ioana Dumitriu and Joel Spencer. The two-batch liar game over an arbitrary channel. SIAM Journal on Discrete Mathematics, 19(4): 1056–1064, January 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dvorak:2006:TAC

[DS06] Zdenvek Dvorák and Riste Skrekovski. A theorem about a contractible and light edge. SIAM Journal on Discrete Mathematics, 20(1): 55–61, January 2006. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dvorak:2009:CNG

[DŠ09] Zdenĕk Dvořák and Riste Škrekovski. k-chromatic number of graphs on surfaces. SIAM Journal on Discrete Mathematics, 23(1): 477–486, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dutta:2016:IBI

[DS16] Kunal Dutta and C. R. Subramanian. Improved bounds on induced acyclic subgraphs in random digraphs. SIAM Journal on Discrete Mathematics, 30(3):1848–1865, ????

2016. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Day:2021:CNE

[DS21a] A. Nicholas Day and Amites Sarkar. On a conjecture of Nagy on extremal densities. SIAM Journal on Discrete Mathematics, 35(1): 294–306, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dillon:2021:MDH

[DS21b] Travis Dillon and Pablo Soberón. A Mélange of diameter Helly-type theorems. SIAM Journal on Discrete Mathematics, 35(3): 1615–1627, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dutta:2023:IPH

[DS23] Kunal Dutta and C. R. Subramanian. On induced paths, holes, and trees in random graphs. SIAM Journal on Discrete Mathematics, 37(1):279–303, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1409895.

David:2019:CCP

[DSL19] Roee David, Karthik C. S., and Bundit Laekhanukit. On the complexity of closest pair via polar-pair of pointsets. SIAM Journal on Discrete Mathematics, 33(1): 509–527, ???? 2019. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Debski:2021:SCD

[DŚN21] Michał Debski and Małgorzata Śleszyńska-Nowak. t-strong cliques and the degreediameter problem. SIAM Journal on Discrete Mathematics, 35(4):3017–3029, ???? 2021. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ding:1992:DPP

[DSS92] Guoli Ding, A. Schrijver, and P. D. Seymour. Disjoint paths in a planar graph: a general theorem. SIAM Journal on Discrete Mathematics, 5(1): 112–116, February 1992. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Delon:2012:LMI

[DSS12] Julie Delon, Julien Salomon, and Andrei Sobolevski. Local matching indicators for transport problems with concave costs. SIAM Journal on Discrete Mathematics, 26(2): 801–827, ???? 2012. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dumitrescu:2013:BMM

[DSST13] Adrian Dumitrescu, André Schulz, Adam Sheffer, and Csaba D. Tóth. Bounds on the maximum multiplicity of some common geometric graphs. SIAM Journal on Discrete Mathematics, 27(2): 802–826, ???? 2013. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

DeLoera:2001:EPD

[DST01] Jesús A. De Loera, Francisco Santos, and Fumihiko Takeuchi. Extremal properties for dissections of convex 3-polytopes. SIAM Journal on Discrete Mathematics, 14(2):143–161, 2001. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/36623.

Dvorak:2008:LCS

[DŠT08] Zdenĕk Dvořák, Riste Škrekovski, and Martin Tancer. Listcoloring squares of sparse subcubic graphs. SIAM Journal on Discrete Mathematics, 22(1):139–159, ???? 2008. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dvorak:2008:PGO

[DŠV08] Z. Dvořák, R. Škrekovski, and T. Valla. Planar graphs of odd-girth at least 9 are homomorphic to the Petersen graph. SIAM Journal on Discrete Mathematics, 22(2): 568–591, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Duffus:1990:MCA

[DSW90] D. Duffus, B. Sands, and P. Winkler. Maximal chains and antichains in Boolean lattices. SIAM Journal on Discrete Mathematics, 3(2): 197–205, May 1990. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dietz:2005:TLB

[DSZ05] Paul F. Dietz, Joel I. Seiferas, and Ju Zhang. A tight lower bound for online monotonic list labeling. SIAM Journal on Discrete Mathematics, 18(3):626–637, 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31580.

Demidovich:2023:CSR

[DSZ23] Yury Demidovich, Arkadiy Skorkin, and Maksim Zhukovskii. Cycle saturation in random graphs. SIAM Journal on Discrete Mathematics, 37 (3):1359–1385, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1456479.

Das:2016:RSE

[DT16] Shagnik Das and Tuan Tran. Removal and stability for Erdős–Ko–Rado. SIAM Journal on Discrete Mathematics, 30(2):1102–1114, ???? 2016. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dumitriu:2003:PGT

[DTW03] Ioana Dumitriu, Prasad Tetali, and Peter Winkler. On playing golf with two balls. SIAM Journal on Discrete Mathematics, 16 (4):604–615, 2003. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/40834.

Ducoffe:2021:FAA

[Duc21] Guillaume Ducoffe. Faster approximation algorithms for computing shortest cycles on weighted graphs. SIAM Journal on Discrete Mathematics, 35(2):953–969, ???? 2021. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

DiBattista:1996:APT

[DV96] Giuseppe Di Battista and Luca Vismara. Angles of planar triangular graphs. SIAM Journal on Discrete Mathematics, 9(3):349–359, August 1996. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

DeDeo:2004:RT

[DV04] Michelle R. DeDeo and Elinor Velasquez. The Radon transform on Z_{n}^{k}. SIAM Journal on Discrete Mathematics, 18(3):472–478, 2004. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/43076.

Dvorak:2020:IBT

[DvHT20] Vojtech Dvorak, Peter van Hintum, and Marius Tiba. Improved bound for Tomaszewski’s Problem. SIAM Journal on Discrete Mathematics, 34(4): 2239–2249, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dvorak:2005:HCP

[Dvo05] Tomás Dvorák. Hamiltonian cycles with prescribed edges in hypercubes. SIAM Journal on Discrete Mathematics, 19(1):135–144, 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/43280.

Dahlberg:2018:LLS

[DvW18] Samantha Dahlberg and Stephanie van Willigenburg. Lollipop and lariat symmetric functions. SIAM Journal on Discrete Mathematics, 32(2): 1029–1039, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

DiSumma:2010:LSS

[DW10] Marco Di Summa and Laurence A. Wolsey. Lot-sizing with stock upper bounds and fixed charges. SIAM Journal on Discrete Mathematics, 24 (3):853–875, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). See erratum [?].

DiSumma:2011:ELS

[DW11] Marco Di Summa and Laurence A. Wolsey. Erratum: Lot-sizing with stock upper bounds and fixed charges. SIAM Journal on Discrete Mathematics, 25(2): 731–732, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v25/i2/p731_s1.

See [DW10].

Dong:2022:SGI

[DW22] Zichao Dong and Zhuo Wu. On the stability of the graph independence number. SIAM Journal on Discrete Mathematics, 36(1):229–240, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL

https://epubs.siam.org/
doi/10.1137/21M1405071.

Disser:2024:UGA

[DW24] Yann Disser and David Weckbecker. Unified greedy approximability beyond submodular maximization. SIAM Journal on Discrete Mathematics, 38(1):348–379, January 2024. CODEN SJD-MEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dong:2019:DCP

[DX19] Wei Dong and Baogang Xu.

2-distance coloring of planar graphs without 4-cycles and 5-cycles. SIAM Journal on Discrete Mathematics, 33(3): 1297–1312, ???? 2019. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dong:2024:REC

[DX24] Zichao Dong and Zijian Xu. Rainbow even cycles. SIAM Journal on Discrete Mathematics, 38(2): 1269–1284, April 2024. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dragan:2006:CTS

[DYL06] Feodor F. Dragan, Chenyu Yan, and Irina Lomonosov. Collective tree spanners of graphs. SIAM Journal on Discrete Mathematics, 20(1):

240–260, January 2006. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dor:2001:MSR

[DZ01] Dorit Dor and Uri Zwick. Median selection requires (2 + ϵ)n comparisons. SIAM Journal on Discrete Mathematics, 14(3):312–325, 2001. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/35389.

DelPia:2009:HIV

[DZ09a] Alberto Del Pia and Giacomo Zambelli. Half-integral vertex covers on bipartite bidirected graphs: Total dual integrality and cut-rank. SIAM Journal on Discrete Mathematics, 23(3):1281–1296, ???? 2009. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Dershowitz:2009:MPT

[DZ09b] Nachum Dershowitz and Shmuel Zaks. More patterns in trees: Up and down, young and old, odd and even. SIAM Journal on Discrete Mathematics, 23(1):447–465, ???? 2009. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Eliahou:2011:TGN

[EA11] Shalom Eliahou and Jorge Ramírez Alfonsín. Two-generator numerical semigroups and Fermat and Mersenne numbers. SIAM Journal on Discrete Mathematics, 25(2): 622–630, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v25/i2/p622_s1.

Esperet:2018:ABF

[EdJdVLT18] Louis Esperet, Rémi de Joannis de Verclos, Tien-Nam Le, and Stéphan Thomassé. Additive bases and flows in graphs. SIAM Journal on Discrete Mathematics, 32(1): 534–542, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Engbers:2021:TGC

[EEFH21] John Engbers, Aysel Erey, Jacob Fox, and Xiaoyu He. Tomescu’s graph coloring conjecture for ℓ-connected graphs. SIAM Journal on Discrete Mathematics, 35(2): 1478–1502, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Epstein:2009:OCI

[EEL09] Leah Epstein, Thomas Erlebach, and Asaf Levin. Online capacitated interval coloring. SIAM Journal on Discrete Mathematics, 23(2): 822–841, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Erdos:1991:MDC

[EFF91] Paul Erdős, Peter Fishburn, and Zoltán Füredi. Midpoints of diagonals of convex n-gons. SIAM Journal on Discrete Mathematics, 4(3): 329–341, August 1991. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Erdos:2005:TPS

[EFK05] Péter L. Erdos, Zoltán Füredi, and Gyula O. H. Katona. Two-part and k-Sperner families: New proofs using permutations. SIAM Journal on Discrete Mathematics, 19(2):489–500, 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/44358.

Erdos:2014:SSS

[EFK14] Dóra Erdős, András Frank, and Krisztián Kun. Sinkstable sets of digraphs. SIAM Journal on Discrete Mathematics, 28(4):1651–1674, ???? 2014. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Espig:2018:ECP

[EFK18] Lisa Espig, Alan Frieze, and Michael Krivelevich. Elegantly colored paths and cycles in edge colored random graphs. SIAM Journal on Discrete Mathematics, 32(3): 1585–1618, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Espig:2015:WBG

[EFKP15] Lisa Espig, Alan Frieze, Michael Krivelevich, and Wesley Pegden. Walker– breaker games. SIAM Journal on Discrete Mathematics, 29(3):1476–1485, ???? 2015. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ejov:2008:DLC

[EFMN08] Vladimir Ejov, Jerzy A. Filar, Walter Murray, and Giang T. Nguyen. Determinants and longest cycles of graphs. SIAM Journal on Discrete Mathematics, 22(3): 1215–1225, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Erbes:2018:SPF

[EFMW18] Catherine Erbes, Michael Ferrara, Ryan R. Martin, and Paul Wenger. Stability of the potential function. SIAM Journal on Discrete Mathematics, 32(3):2313–2331, ???? 2018. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Engel:2003:TBC

[EG03] Konrad Engel and Sven Guttmann. Testing bandwidth k for k-connected graphs. SIAM Journal on Discrete Mathematics, 16 (2):301–312, 2003. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/35114.

Egecioglu:2010:BPP

[Ege10]  Ömer Egecioglu.  Bessel

polynomials and the partial sums of the exponential series. SIAM Journal on Discrete Mathematics, 24(4): 1753–1762, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sidma/resource/
                                        1/sjdmec/v24/i4/p1753_s1.

Esperet:2021:IUG

[EGG21] Louis Esperet, Cyril Gavoille, and Carla Groenland. Isometric universal graphs. SIAM Journal on Discrete Mathematics, 35(2):1224– 1237, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ergemlidze:2018:UHL

[EGM18] Beka Ergemlidze, Ervin Györi, and Abhishek Methuku. 3-uniform hypergraphs and linear cycles. SIAM Journal on Discrete Mathematics, 32 (2):933–950, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Engels:2020:EMM

[EGMR20] Christian Engels, Mohit Garg, Kazuhisa Makino, and Anup Rao. On expressing majority as a majority of majorities. SIAM Journal on Discrete Mathematics, 34(1): 730–741, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Eisenstat:2008:CPR

[EGR08] David Eisenstat, Gary Gordon, and Amanda Redlich. Combinatorial properties of a rooted graph polynomial. SIAM Journal on Discrete Mathematics, 22(2): 776–785, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Edelman:2013:SCO

[EGS13] Paul H. Edelman, Tatiana Gvozdeva, and Arkadii Slinko. Simplicial complexes obtained from qualitative probability orders. SIAM Journal on Discrete Mathematics, 27(4):1820–1843, ???? 2013. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Etzion:1991:TLS

[EH91] Tuvi Etzion and Alan Hartman. Towards a large set of Steiner quadruple systems. SIAM Journal on Discrete Mathematics, 4(2): 182–195, May 1991. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Even:2013:LOG

[EH13] Guy Even and Nissim Halabi. Local-optimality guarantees based on paths for optimal decoding. SIAM Journal on Discrete Mathematics, 27(4):2049–2063, ???? 2013. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Enomoto:2001:CCN

[EHJ01] Hikoe Enomoto, Mirko Hornák, and Stanislav Jendrol’. Cyclic chromatic number of 3connected plane graphs. SIAM Journal on Discrete Mathematics, 14(1): 121–137, 2001. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/34615.

Esfandiari:2017:PS

[EHLM17] Hossein Esfandiari, MohammadTaghi Hajiaghayi, Vahid Liaghat, and Morteza Monemizadeh. Prophet secretary. SIAM Journal on Discrete Mathematics, 31(3): 1685–1701, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Erman:2011:CGC

[EHLP11] Rok Erman, Frédéric Havet, Bernard Lidický, and Ondrej Pangrác. 5-coloring graphs with 4 crossings. SIAM Journal on Discrete Mathematics, 25(1):401–422, ???? 2011. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v25/i1/p401_s1.

Ehrenborg:2016:CCP

[Ehr16] Richard Ehrenborg. Cyclically consecutive permutation avoidance. SIAM Journal on Discrete Mathematics, 30(3): 1385–1390, ???? 2016. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ernvall-Hytonen:2018:SGM

[EHV18] Anne-Maria Ernvall-Hytönen and Esa V. Vesalainen. On the secrecy gain of ℓ-modular lattices. SIAM Journal on Discrete Mathematics, 32(2): 1441–1457, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Edelsbrunner:2018:OFL

[EIH18] Herbert Edelsbrunner and Mabel Iglesias-Ham. On the optimality of the FCC lattice for soft sphere packing. SIAM Journal on Discrete Mathematics, 32(1):750–782, ???? 2018. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Egawa:2008:NIC

[EIiK08] Yoshimi Egawa, Katsumi Inoue, and Ken ichi Kawarabayashi. Nonseparating induced cycles consisting of contractible edges in k-connected graphs. SIAM Journal on Discrete Mathematics, 21(4):1061– 1070, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Erickson:2012:NRR

[EIJ^{+}12] Alejandro Erickson, Abraham Isgur, Bradley W. Jackson, Frank Ruskey, and Stephen M. Tanny. Nested recurrence relations with Conolly-like solutions. SIAM Journal on Discrete Mathematics, 26(1):206–238, ???? 2012. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Erlebach:2001:MED

[EJ01] Thomas Erlebach and Klaus Jansen. The maximum edge-disjoint paths problem in bidirected trees. SIAM Journal on Discrete Mathematics, 14(3):326–355, 2001. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/36125.

Erman:2009:ORN

[EJK^{+}09] Rok Erman, Suzana Jurečič, Daniel Král’, Kris Stopar, and Nik Stopar. Optimal real number graph labellings of a subfamily of Kneser graphs. SIAM Journal on Discrete Mathematics, 23(3): 1372–1381, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Elkin:2005:PAI

[EK05] Michael Elkin and Guy Kortsarz. Polylogarithmic additive inapproximability of the radio broadcast problem. SIAM Journal on Discrete Mathematics, 19 (4):881–899, 2005. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/44531.

Elkies:2010:CTI

[EK10] Noam D. Elkies and Scott Duke Kominers. On the classification of type II codes of length 24. SIAM Journal on Discrete Mathematics, 23(4): 2173–2177, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Edwards:2013:BFC

[EK13] Katherine Edwards and Andrew D. King. Bounding the fractional chromatic number of K_{Δ}-free graphs. SIAM Journal on Discrete Mathematics, 27(2):1184–1208, ???? 2013. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Edwards:2015:RHC

[EKK^{+}15] Katherine Edwards, Dong Yeap Kang, Jaehoon Kim, Sang il Oum, and Paul Seymour. A relative of Hadwiger’s conjecture. SIAM Journal on Discrete Mathematics, 29(4): 2385–2388, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Eiben:2019:LKC

[EKM^{+}19] Eduard Eiben, Mithilesh Kumar, Amer E. Mouawad, Fahad Panolan, and Sebastian Siebertz. Lossy kernels for connected dominating set on sparse graphs. SIAM Journal on Discrete Mathematics, 33(3):1743–1771, ???? 2019. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Elbassioni:2009:ADP

[Elb09] Khaled M. Elbassioni. Algorithms for dualization over products of partially ordered sets. SIAM Journal on Discrete Mathematics, 23(1): 487–510, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Elizalde:2009:NPR

[Eli09] Sergi Elizalde. The number of permutations realized by a shift. SIAM Journal on Discrete Mathematics, 23(2): 765–786, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Epstein:2011:IAG

[ELMS11] Leah Epstein, Asaf Levin, Julián Mestre, and Danny Segev. Improved approximation guarantees for weighted matching in the semi-streaming model. SIAM Journal on Discrete Mathematics, 25(3): 1251–1265, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v25/i3/p1251_s1.

Even:1998:MPT

[ELR98] Shimon Even, Ami Litman, and Arnold L. Rosenberg. Monochromatic paths and triangulated graphs. SIAM Journal on Discrete Mathematics, 11(4):546–556, November 1998. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Epstein:2017:PPM

[ELSS17] Leah Epstein, Asaf Levin, Alan J. Soper, and Vitaly A. Strusevich. Power of preemption for minimizing total completion time on uniform parallel machines. SIAM Journal on Discrete Mathematics, 31 (1):101–123, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Epstein:2011:MMO

[ELvS11] Leah Epstein, Asaf Levin, and Rob van Stee. Maxmin online allocations with a reordering buffer. SIAM Journal on Discrete Mathematics, 25(3):1230–1250, ???? 2011. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sidma/resource/
                                        1/sjdmec/v25/i3/p1230_s1.

Enomoto:1999:EGT

[EM99] Hikoe Enomoto and Miki Shimabara Miyauchi. Embedding graphs into a three page book with O(m log n) crossings of edges over the spine. SIAM Journal on Discrete Mathematics, 12(3):337–341, August 1999. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/28031.

Elbassioni:2020:EVC

[EM20] Khaled Elbassioni and Kazuhisa Makino. Enumerating vertices of covering polyhedra with totally unimodular constraint matrices. SIAM Journal on Discrete Mathematics, 34(1):843–864, ???? 2020. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Eirinakis:2014:OSM

[EMM14] Pavlos Eirinakis, Dimitrios Magos, and Ioannis Mourtos. From one stable marriage to the next: How long is the way? SIAM Journal on Discrete Mathematics, 28(4): 1971–1979, ???? 2014. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ellingham:2016:CMF

[EMOT16] M. N. Ellingham, Emily A.

Marshall, Kenta Ozeki, and Shoichi Tsuchiya. A characterization of K_{2,4}-minor-free graphs. SIAM Journal on Discrete Mathematics, 30(2): 955–975, ???? 2016. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Elias:2014:LBG

[EMRPS14] Marek Eliás, Jirí Matousek, Edgardo Roldán-Pensado, and Zuzana Safernová. Lower bounds on geometric Ramsey functions. SIAM Journal on Discrete Mathematics, 28(4): 1960–1970, ???? 2014. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Erdos:2015:DBP

[EMT15] Péter L. Erdős, István Miklós, and Zoltán Toroczkai. A decomposition based proof for fast mixing of a Markov chain over balanced realizations of a joint degree matrix. SIAM Journal on Discrete Mathematics, 29(1): 481–499, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Elkin:2022:LPE

[EN22] Michael Elkin and Ofer Neiman. Lossless prioritized embeddings. SIAM Journal on Discrete Mathematics, 36 (3):1529–1550, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1436221.

Encheva:2005:PCS

[Enc05] Sylvia Encheva. On projective codes satisfying the chain condition. SIAM Journal on Discrete Mathematics, 19(1):145–148, 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31431.

Engebretsen:2004:NNB

[Eng04]  Lars Engebretsen.  The

nonapproximability of non-Boolean predicates. SIAM Journal on Discrete Mathematics, 18(1):114–129, 2004. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/38045.

Elkin:2015:LS

[ENS15] Michael Elkin, Ofer Neiman, and Shay Solomon. Light spanners. SIAM Journal on Discrete Mathematics, 29(3): 1312–1321, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Even:2000:AMS

[ENSZ00] Guy Even, Joseph (Seffi)

Naor, Baruch Schieber, and Leonid Zosin. Approximating minimum subset feedback sets in undirected graphs with applications. SIAM Journal on Discrete Mathematics, 13(2):255–267, 2000. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29187.

Esperet:2016:IGS

[EO16] Louis Esperet and Pascal Ochem. Islands in graphs on surfaces. SIAM Journal on Discrete Mathematics, 30(1):

206–219, ???? 2016. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Emek:2010:TUB

[EP10] Yuval Emek and David Peleg. A tight upper bound on the probabilistic embedding of series-parallel graphs. SIAM Journal on Discrete Mathematics, 23(4):1827–1841, ???? 2010. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Epstein:2006:OPC

[Eps06] Leah Epstein. Online bin packing with cardinality constraints. SIAM Journal on Discrete Mathematics, 20(4): 1015–1030, January 2006. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Erde:2017:RTD

[Erd17] Joshua Erde. Refining a treedecomposition which distinguishes tangles. SIAM Journal on Discrete Mathematics, 31(3):1529–1551, ???? 2017. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Erde:2020:DPD

[Erd20] Joshua Erde. Directed pathdecompositions. SIAM Journal on Discrete Mathematics, 34(1):415–430, ???? 2020. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Eriksson:1996:CFG

[Eri96]  Kimmo Eriksson.  Chip-

firing games on mutating graphs. SIAM Journal on Discrete Mathematics, 9(1): 118–128, February 1996. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Eden:2019:STE

[ERS19] Talya Eden, Dana Ron, and C. Seshadhri. Sublinear time estimation of degree distribution moments: The arboricity connection. SIAM Journal on Discrete Mathematics, 33(4): 2267–2285, ???? 2019. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ezra:2022:CCR

[ERSZ22] Esther Ezra, Orit E. Raz, Micha Sharir, and Joshua Zahl. Counting and cutting rich lenses in arrangements of circles. SIAM Journal on Discrete Mathematics, 36(2):958–974, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1409305.

Espona:1998:CDB

[ES98] M. Espona and O. Serra.

Cayley digraphs based on the de Bruijn networks. SIAM Journal on Discrete Mathematics, 11(2):305–317, May 1998. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Elkin:2011:NSL

[ES11] Michael Elkin and Shay Solomon. Narrow-shallow-low-light trees with and without Steiner points. SIAM Journal on Discrete Mathematics, 25(1):181–210, ???? 2011. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:

//epubs.siam.org/sidma/
resource/1/sjdmec/v25/i1/
p181_s1.

Enright:2014:LCL

[EST14] Jessica Enright, Lorna Stewart, and Gábor Tardos. On list coloring and list homomorphism of permutation and interval graphs. SIAM Journal on Discrete Mathematics, 28(4):1675–1685, ???? 2014. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Etesami:2022:OAM

[Ete22] S. Rasoul Etesami. Online assortment and market segmentation under Bertrand competition with set-dependent revenues. SIAM Journal on Discrete Mathematics, 36 (2):1436–1466, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1400134.

Etingof:2020:ARS

[Eti20]  Pavel Etingof.  Angle-

restricted sets and zerofree regions for the permanent. SIAM Journal on Discrete Mathematics, 34(2): 1248–1260, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Erdos:2013:CDR

[ETT13] Péter L. Erdős, Claude Tardif, and Gábor Tardos. Caterpillar dualities and regular languages. SIAM Journal on Discrete Mathematics, 27(3):1287–1294, ???? 2013. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Etzion:1996:NAP

[Etz96a] Tuvi Etzion. Nonequivalent q-ary perfect codes. SIAM Journal on Discrete Mathematics, 9(3):413–423, August 1996. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Etzion:1996:NPC

[Etz96b] Tuvi Etzion. On the nonexistence of perfect codes in the Johnson scheme. SIAM Journal on Discrete Mathematics, 9(2):201–209, May 1996. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Etzion:1998:PCT

[EV98] Tuvi Etzion and Alexander Vardy. On perfect codes and tilings: Problems and solutions. SIAM Journal on Discrete Mathematics, 11(2): 205–223, May 1998. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Erde:2019:SDS

[EW19] Joshua Erde and Daniel Weißauer. A short derivation of the structure theorem for graphs with excluded topological minors. SIAM Journal on Discrete Mathematics, 33(3):1654–1661, ???? 2019. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Exoo:1989:TCR

[Exo89] Geoffrey Exoo. On two classical Ramsey numbers of the form R(3, n). SIAM Journal on Discrete Mathematics, 2 (4):488–490, November 1989. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fan:1992:CGC

[Fan92] Geng Hua Fan. Covering graphs by cycles. SIAM Journal on Discrete Mathematics, 5(4):491–496, November 1992. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fang:2020:GPD

[Fan20] Jin-Hui Fang. On generalized perfect difference sumsets. SIAM Journal on Discrete Mathematics, 34(4): 2448–2456, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Farzad:2009:PGC

[Far09]  Babak Farzad.  Planar

graphs without 7-cycles are 4choosable. SIAM Journal on Discrete Mathematics, 23(3): 1179–1199, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Feder:2001:CHO

[Fed01]  Tomás Feder.  Classifica-

tion of homomorphisms to oriented cycles and of kpartite satisfiability. SIAM Journal on Discrete Mathematics, 14(4):471–480, 2001. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/38335.

Feder:2006:DTF

[Fed06] Tomás Feder. A dichotomy theorem on fixed points of several nonexpansive mappings. SIAM Journal on Discrete Mathematics, 20(2): 291–301, January 2006. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Feige:2004:AMC

[Fei04] Uriel Feige. Approximating maximum clique by removing subgraphs. SIAM Journal on Discrete Mathematics, 18(2):219–225, 2004. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/40415.

Feige:2010:OSG

[Fei10] Uriel Feige. On optimal strategies for a hat game on graphs. SIAM Journal on Discrete Mathematics, 24(3): 782–791, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fellows:1990:TVP

[Fel90] Michael R. Fellows. Transversals of vertex partitions in graphs. SIAM Journal on Discrete Mathematics, 3(2): 206–215, May 1990. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Felsner:2014:ODP

[Fel14] Stefan Felsner. The order dimension of planar maps revisited. SIAM Journal on Discrete Mathematics, 28(3): 1093–1101, ???? 2014. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Feray:2015:CIE

[Fér15]  Valentin Féray.  Cyclic

inclusion-exclusion. SIAM Journal on Discrete Mathematics, 29(4):2284–2311, ???? 2015. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ferrari:2016:DAI

[Fer16] Luca Ferrari. Dyck algebras, interval temporal logic, and posets of intervals. SIAM Journal on Discrete Mathematics, 30(4):1918–1937, ???? 2016. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fernley:2024:DVM

[Fer24]  John Fernley.  Discursive

voter models on the supercritical scale-free network. SIAM Journal on Discrete Mathematics, 38(2): 1285–1314, April 2024. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Feder:2006:CBB

[FF06] Tomás Feder and Daniel Ford. Classification of bipartite Boolean constraint satisfaction through delta-matroid intersection. SIAM Journal on Discrete Mathematics, 20 (2):372–394, January 2006. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fellows:1994:PNC

[FFHJ94] Michael Fellows, Gerd Fricke, Stephen Hedetniemi, and David Jacobs. The private neighbor cube. SIAM Journal on Discrete Mathematics, 7(1):41–47, February 1994. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fiduccia:1998:GDB

[FFie98] Charles M. Fiduccia, Rodney W. Forcade, and Jennifer S. Zito input epsf. Geometry and diameter bounds of directed Cayley graphs of Abelian groups. SIAM Journal on Discrete Mathematics, 11(1):157–167, February 1998. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fang:2020:CFM

[FFLP20] Xin Fang, Ghislain Fourier, Jan-Philipp Litza, and Christoph Pegel. A continuous family of marked poset polytopes. SIAM Journal on Discrete Mathematics, 34(1): 611–639, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Feldman:1988:WSN

[FFP88] Paul Feldman, Joel Friedman, and Nicholas Pippenger. Wide-sense nonblocking networks. SIAM Journal on Discrete Mathematics, 1 (2):158–173, May 1988. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Feige:2011:DSS

[FFV11] Uriel Feige, Abraham D.

Flaxman, and Dan Vilenchik. On the diameter of the set of satisfying assignments in random satisfiable k-CNF formulas. SIAM Journal on Discrete Mathematics, 25(2): 736–749, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v25/i2/p736_s1.

Fox:1989:RSG

[FG89] Dale R. Fox and S. Kingsley Gnanendran. Recursive solutions for the generalized Erlang queueing systems. SIAM Journal on Discrete Mathematics, 2(2):219–230, May 1989. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fomin:2000:GSI

[FG00] Fedor V. Fomin and Petr A.

Golovach. Graph searching and interval completion. SIAM Journal on Discrete Mathematics, 13 (4):454–464, 2000. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/35047.

Fiol:2001:ACC

[FG01] M. A. Fiol and E. Garriga. An algebraic characterization of completely regular codes in distanceregular graphs. SIAM Journal on Discrete Mathematics, 15(1):1–13, 2001. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/37649.

Fomin:2014:LCL

[FG14] Fedor V. Fomin and Petr A. Golovach. Long circuits and large Euler subgraphs. SIAM Journal on Discrete Mathematics, 28(2):878–892, ???? 2014. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fomin:2021:KWS

[FG21] Fedor V. Fomin and Petr A. Golovach. Kernelization of Whitney switches. SIAM Journal on Discrete Mathematics, 35(2):1298–1336, ???? 2021. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fan:2022:PPO

[FG22] Neil Fan and Peter Guo. Poincaré polynomials of odd diagram classes. SIAM Journal on Discrete Mathematics, 36(3):2225–2237, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1470517.

Furedi:1989:PLG

[FGK89] Z. Furedi, J. R. Griggs, and D. J. Kleitman. Pair labellings with given distance. SIAM Journal on Discrete Mathematics, 2(4):491– 499, November 1989. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Friedrich:2022:SIV

[FGKP22] Tobias Friedrich, Andreas Göbel, Martin S. Krejca, and Marcus Pappik. A spectral independence view on hard spheres via block dynamics. SIAM Journal on Discrete Mathematics, 36 (3):2282–2322, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M143697X.

Fomin:2020:GFD

[FGL^{+}20] Fedor V. Fomin, Petr A.

Golovach, Daniel Lokshtanov, Fahad Panolan, Saket Saurabh, and Meirav Zehavi. Going far from degeneracy. SIAM Journal on Discrete Mathematics, 34(3): 1587–1601, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Florescu:2014:ERR

[FGLP14] Laura Florescu, Shirshendu Ganguly, Lionel Levine, and Yuval Peres. Escape rates for rotor walks in Z^{d}. SIAM Journal on Discrete Mathematics, 28(1):323–334, ???? 2014. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fomin:2018:CVS

[FGLS18] Fedor V. Fomin, Petr A. Golovach, Daniel Lokshtanov, and Saket Saurabh. Covering vectors by spaces: Regular matroids. SIAM Journal on Discrete Mathematics, 32(4):2512–2565, ???? 2018. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Franc:2010:CGI

[FGP10] A. Franc, M. Goulard, and N. Peyrard. Chordal graphs to identify graphical model solutions of maximum of entropy under constraints on marginals. SIAM Journal on Discrete Mathematics, 24(3): 1104–1116, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fomin:2012:CRC

[FGP12] Fedor V. Fomin, Petr A. Golovach, and Pawel Pralat. Cops and robber with constraints. SIAM Journal on Discrete Mathematics, 26(2):

571–590, ???? 2012. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fomin:2019:ECD

[FGPS19] Fedor V. Fomin, Petr Golovach, Fahad Panolan, and Saket Saurabh. Editing to connected F -degree graph. SIAM Journal on Discrete Mathematics, 33(2):795–836, ???? 2019. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fan:2020:LPN

[FGPS20] Neil J. Y. Fan, Peter L. Guo, Simon C. Y. Peng, and Sophie C. C. Sun. Lattice points in the Newton polytopes of key polynomials. SIAM Journal on Discrete Mathematics, 34(2):1281–1289, ???? 2020. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Focke:2021:CHM

[FGRZ21] Jacob Focke, Leslie Ann Goldberg, Marc Roth, and Stanislav Zivný. Counting homomorphisms to K_{4}- minor-free graphs, modulo 2. SIAM Journal on Discrete Mathematics, 35(4): 2749–2814, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fan:2019:PCR

[FGS19] Neil J. Y. Fan, Peter L. Guo, and Sophie C. C. Sun. Proof of a conjecture of Reiner– Tenner–Yong on barely setvalued tableaux. SIAM Journal on Discrete Mathematics, 33(1):189–196, ???? 2019. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fomin:2011:AWP

[FGT11] Fedor V. Fomin, Petr A. Golovach, and Dimitrios M. Thilikos. Approximating width parameters of hypergraphs with excluded minors. SIAM Journal on Discrete Mathematics, 25(3): 1331–1348, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sidma/resource/
                                        1/sjdmec/v25/i3/p1331_s1.

Fomin:2018:SCA

[FGT18] Fedor V. Fomin, Petr A. Golovach, and Dimitrios M. Thilikos. Structured connectivity augmentation. SIAM Journal on Discrete Mathematics, 32(4):2612–2635, ???? 2018. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Focke:2019:CCS

[FGZ19] Jacob Focke, Leslie Ann Goldberg, and Stanislav Zivný. The complexity of counting surjective homomorphisms and compactions. SIAM Journal on Discrete Mathematics, 33(2):1006– 1043, ???? 2019. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Feigenbaum:1989:FES

[FH89] Joan Feigenbaum and Ramsey W. Haddad. On factorable extensions and subgraphs of prime graphs. SIAM Journal on Discrete Mathematics, 2(2):197–218, May 1989. CODEN SJD-MEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fan:1994:SHC

[FH94] Genghua Fan and Roland Häggkvist. The square of a Hamiltonian cycle. SIAM Journal on Discrete Mathematics, 7(2):203–212, May 1994. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fountoulakis:2010:QRS

[FH10] Nikolaos Fountoulakis and Anna Huber. Quasirandom rumor spreading on the complete graph is as fast as randomized rumor spreading. SIAM Journal on Discrete Mathematics, 23(4): 1964–1991, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Feuilloley:2021:GCF

[FH21] Laurent Feuilloley and Michel Habib. Graph classes and forbidden patterns on three vertices. SIAM Journal on Discrete Mathematics, 35(1): 55–90, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Feder:2008:BTT

[FHH08] Tomás Feder, Pavol Hell, and Jing Huang. Brookstype theorems for pair-list colorings and list homomorphisms. SIAM Journal on Discrete Mathematics, 22(1): 1–14, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Feder:2010:RP

[FHJ^{+}10] Tomás Feder, Pavol Hell, Peter Jonsson, Andrei Krokhin, and Gustav Nordh. Retractions to pseudoforests. SIAM Journal on Discrete Mathematics, 24(1):101–112, ???? 2010. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fiorini:2017:EMI

[FHJV17] Samuel Fiorini, Tony Huynh, Gwenaël Joret, and Antonios Varvitsiotis. The excluded minors for isometric realizability in the plane. SIAM Journal on Discrete Mathematics, 31(1):438–453, ????

2017. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Faigle:1996:CNB

[FHK96] Ulrich Faigle, Alan J. Hoffman, and Walter Kern. A characterization of nonnegative box-greedy matrices. SIAM Journal on Discrete Mathematics, 9(1):1– 6, February 1996. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Feder:2003:LP

[FHKM03] Tomas Feder, Pavol Hell, Sulamita Klein, and Rajeev Motwani. List partitions. SIAM Journal on Discrete Mathematics, 16 (3):449–478, 2003. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/38405.

Feder:2013:GAN

[FHL^{+}13a] Tomás Feder, Pavol Hell, Benoît Larose, Cynthia Loten, Mark Siggers, and Claude Tardif. Graphs admitting k-NU operations. Part 1: The reflexive case. SIAM Journal on Discrete Mathematics, 27(4):1940–1963, ???? 2013. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Frieze:2013:GCN

[FHL13b] Alan Frieze, Simcha Haber, and Mikhail Lavrov. On the game chromatic number of sparse random graphs. SIAM Journal on Discrete Mathematics, 27(2):768–790, ???? 2013. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Feder:2014:GAN

[FHL^{+}14] Tomás Feder, Pavol Hell, Benoît Larose, Mark Siggers, and Claude Tardif. Graphs admitting k-NU operations. Part 2: The irreflexive case. SIAM Journal on Discrete Mathematics, 28(2): 817–834, ???? 2014. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fu:2020:ASP

[FHL20] Shishuo Fu, Guo-Niu Han, and Zhicong Lin. karrangements, statistics, and patterns. SIAM Journal on Discrete Mathematics, 34(3): 1830–1853, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Freschi:2021:NCB

[FHLT21] Andrea Freschi, Joseph Hyde, Joanna Lada, and Andrew Treglown. A note on colorbias Hamilton cycles in dense graphs. SIAM Journal on Discrete Mathematics, 35(2):

970–975, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Felsner:1994:IBI

[FHM94] S. Felsner, M. Habib, and R. H. Möhring. On the interplay between interval dimension and dimension. SIAM Journal on Discrete Mathematics, 7(1):32–40, February 1994. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Feder:2003:AHC

[FHM03] Tomás Feder, Pavol Hell, and Bojan Mohar. Acyclic homomorphisms and circular colorings of digraphs. SIAM Journal on Discrete Mathematics, 17(1):161–169, January 2003. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fuji-Hara:2001:OOO

[FHMY01] Ryoh Fuji-Hara, Ying Miao, and Jianxing Yin. Optimal (9v, 4, 1) optical orthogonal codes. SIAM Journal on Discrete Mathematics, 14(2):256–266, 2001. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/37723.

Fluschnik:2018:FKL

[FHNN18] Till Fluschnik, Danny Hermelin, André Nichterlein, and Rolf Niedermeier. Fractals for kernelization lower bounds. SIAM Journal on Discrete Mathematics, 32(1): 656–681, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Francis:2014:BQN

[FHS14] Mathew Francis, Pavol Hell, and Juraj Stacho. Blocking quadruple: a new obstruction to circular-arc graphs. SIAM Journal on Discrete Mathematics, 28(2):631–655, ???? 2014. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fujishige:2005:BFM

[FI05] Satoru Fujishige and Satoru Iwata. Bisubmodular function minimization. SIAM Journal on Discrete Mathematics, 19(4):1065–1073, January 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fill:1989:RT

[Fil89] James Allen Fill. The Radon transform on Z_{n}. SIAM Journal on Discrete Mathematics, 2(2):262–283, May 1989. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Frank:1998:TAD

[FIN98] András Frank, Toshihide Ibaraki, and Hiroshi Nagamochi. Two arc-disjoint paths in Eulerian digraphs. SIAM Journal on Discrete Mathematics, 11(4):557–589, November 1998. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fink:2009:CMG

[Fin09] Jiří Fink. Connectivity of matching graph of hypercube. SIAM Journal on Discrete Mathematics, 23(2): 1100–1109, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fiorini:2006:CLO

[Fio06] Samuel Fiorini. 0, 1/2-cuts and the linear ordering problem: Surfaces that define facets. SIAM Journal on Discrete Mathematics, 20(4): 893–912, January 2006. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fishburn:1990:BPI

[Fis90a] Peter C. Fishburn. Binary probabilities induced by rankings. SIAM Journal on Discrete Mathematics, 3(4):478– 488, November 1990. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fishburn:1990:TOS

[Fis90b] Peter C. Fishburn. Thicknesses of ordered sets. SIAM Journal on Discrete Mathematics, 3(4):489–501, November 1990. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fisher:1994:DFD

[Fis94] David C. Fisher. Domination, fractional domination, 2-packing, and graph products. SIAM Journal on Discrete Mathematics, 7(3):493– 498, August 1994. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fischer:2014:AAQ

[Fis14] Anja Fischer. An analysis of the asymmetric quadratic traveling salesman polytope. SIAM Journal on Discrete Mathematics, 28(1):240–276, ???? 2014. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fujishige:2014:MMT

[FiT14] Satoru Fujishige and Shin ichi Tanigawa. A min-max theorem for transversal submodular functions and its implications. SIAM Journal on Discrete Mathematics, 28(4): 1855–1875, ???? 2014. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fan:2009:RLL

[FJ09] Genghua Fan and Naidan Ji. Relative length of longest paths and cycles in 2-connected graphs. SIAM Journal on Discrete Mathematics, 23(3):1238–1248, ???? 2009. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Frieze:2017:MCM

[FJ17] Alan Frieze and Tony Johansson. Minimum cost matching in a random graph with random costs. SIAM Journal on Discrete Mathematics, 31 (1):489–510, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Francis:2018:URM

[FJJ18] Mathew C. Francis, Dalu Jacob, and Satyabrata Jana. Uniquely restricted matchings in interval graphs. SIAM Journal on Discrete Mathematics, 32(1):148–172, ???? 2018. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Furedi:2019:HCT

[FJK^{+}19] Zoltán Füredi, Tao Jiang, Alexandr Kostochka, Dhruv Mubayi, and Jacques Verstraëte. Hypergraphs not containing a tight tree with a bounded trunk. SIAM Journal on Discrete Mathematics, 33(2):862–873, ???? 2019. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Furedi:2023:EPH

[FJK^{+}23] Zoltán Füredi, Tao Jiang, Alexandr Kostochka, Dhruv Mubayi, and Jacques Verstraëte. Extremal problems for hypergraph blowups of trees. SIAM Journal on Discrete Mathematics, 37 (4):2397–2416, October 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fiala:2003:GSC

[FJLS03] Jirí Fiala, Klaus Jansen, Van Bang Le, and Eike Seidel. Graph subcolorings: Complexity and algorithms. SIAM Journal on Discrete Mathematics, 16(4):635–650, 2003. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/39524.

Fountoulakis:2022:PRG

[FJP22] Nikolaos Fountoulakis, Felix Joos, and Guillem Perarnau. Percolation on random graphs with a fixed degree sequence. SIAM Journal on Discrete Mathematics, 36(1):1–46, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/20M1347607.

Farber:2015:NIE

[FJZ15] Miriam Farber, Charles Johnson, and Leon Zhang. The number of interlacing equalities resulting from removal of a vertex from a tree. SIAM Journal on Discrete Mathematics, 29(3):1245–1258, ???? 2015. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fishburn:1997:SAD

[FK97] Peter C. Fishburn and Jerry S. Kelly. Super Arrovian domains with strict preferences. SIAM Journal on Discrete Mathematics, 10(1): 83–95, February 1997. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Faigle:1998:ACA

[FK98] Ulrich Faigle and Walter Kern. Approximate core allocation for binpacking games. SIAM Journal on Discrete Mathematics, 11(3): 387–399, August 1998. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Filtser:2017:STV

[FK17] Arnold Filtser and Robert Krauthgamer. Sparsification of two-variable valued constraint satisfaction problems. SIAM Journal on Discrete Mathematics, 31(2): 1263–1276, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Friedrich:2018:DHR

[FK18a] Tobias Friedrich and Anton Krohmer. On the diameter of hyperbolic random graphs. SIAM Journal on Discrete Mathematics, 32(2): 1314–1334, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Furedi:2018:KRR

[FK18b] Zoltán Füredi and Ida Kantor. Kneser ranks of random graphs and minimum difference representations. SIAM Journal on Discrete Mathematics, 32(2):1016–1028, ???? 2018. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fulek:2021:SPA

[FK21] Radoslav Fulek and Balázs Keszegh. Saturation problems about forbidden 0–1 submatrices. SIAM Journal on Discrete Mathematics, 35 (3):1964–1977, ???? 2021. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Feng:2005:ETC

[FKK05] Rongquan Feng, Jin Ho Kwak, and Young Soo Kwon. Enumerating typical circulant covering projections onto a circulant graph. SIAM Journal on Discrete Mathematics, 19(1):196–207, 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/43186. See er-

ratum [FKK07].

Feng:2007:EET

[FKK07] Rongquan Feng, Jin Ho Kwak, and Young Soo Kwon. Erratum: Enumerating typical circulant covering projections onto a circulant graph. SIAM Journal on Discrete Mathematics, 21(2): 548–550, ???? 2007. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). See [FKK05].

Friedrich:2018:UDD

[FKK18] Tobias Friedrich, Maximilian Katzmann, and Anton Krohmer. Unbounded discrepancy of deterministic random walks on grids. SIAM Journal on Discrete Mathematics, 32(4):2441–2452, ???? 2018. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Feige:2020:PMC

[FKK20] Uriel Feige, Anne Kenyon, and Shimon Kogan. On the profile of multiplicities of complete subgraphs. SIAM Journal on Discrete Mathematics, 34(1):950–971, ???? 2020. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fekete:2021:MSC

[FKK21] Sándor P. Fekete, Linda Kleist, and Dominik Krupke. Minimum scan cover with angular transition costs. SIAM Journal on Discrete Mathematics, 35(2):1337–1355, ???? 2021. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Feng:1998:ICC

[FKKL98] Rongquan Feng, Jin Ho Kwak, Juyoung Kim, and Jaeun Lee. Isomorphism classes of concrete graph coverings. SIAM Journal on Discrete Mathematics, 11(2): 265–272, May 1998. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Feigenbaum:1993:CRP

[FKL93] Joan Feigenbaum, Jeremy A. Kahn, and Carsten Lund. Complexity results for POM-SET languages. SIAM Journal on Discrete Mathematics, 6(3):432–442, August 1993. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fomin:2019:PSE

[FKL^{+}19] Fedor V. Fomin, Petteri Kaski, Daniel Lokshtanov, Fahad Panolan, and Saket Saurabh. Parameterized single-exponential time polynomial space algorithm for Steiner tree. SIAM Journal on Discrete Mathematics, 33 (1):327–345, ???? 2019. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Floderus:2015:DCS

[FKLL15] Peter Floderus, Mirosław Kowaluk, Andrzej Lingas, and Eva-Marta Lundell. Detecting and counting small pattern graphs. SIAM Journal on Discrete Mathematics, 29(3):1322–1339, ???? 2015. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fishburn:1998:IMC

[FKLW98] P. C. Fishburn, J. H.

Kim, J. C. Lagarias, and P. E. Wright. Interferenceminimizing colorings of regular graphs. SIAM Journal on Discrete Mathematics, 11(1): 15–40, February 1998. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fagin:2006:CPR

[FKM^{+}06] Ronald Fagin, Ravi Kumar, Mohammad Mahdian, D. Sivakumar, and Erik Vee. Comparing partial rankings. SIAM Journal on Discrete Mathematics, 20(3): 628–648, January 2006. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fujita:2013:RMM

[FKM13] Ryo Fujita, Yusuke Kobayashi, and Kazuhisa Makino. Robust matchings and matroid intersections. SIAM Journal on Discrete Mathematics, 27(3):1234–1256, ???? 2013. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fagin:2016:AVV

[FKM^{+}16] Ronald Fagin, Ravi Kumar, Mohammad Mahdian, D. Sivakumar, and Erik Vee. An algorithmic view of voting. SIAM Journal on Discrete Mathematics, 30(4): 1978–1996, ???? 2016. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Furedi:2010:RFC

[FKMS10] Zoltan Füredi, Ida Kantor, Angelo Monti, and Blerina Sinaimeri. On reversefree codes and permutations. SIAM Journal on Discrete Mathematics, 24(3): 964–978, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Felsner:2020:RCF

[FKMS20] Stefan Felsner, Linda Kleist, Torsten Mütze, and Leon Sering. Rainbow cycles in flip graphs. SIAM Journal on Discrete Mathematics, 34(1): 1–39, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Foucaud:2015:LSS

[FKP15] F. Foucaud, M. Krivelevich, and G. Perarnau. Large subgraphs without short cycles. SIAM Journal on Discrete Mathematics, 29(1):65– 78, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fich:2005:GMR

[FKPR05] Faith Ellen Fich, André Kündgen, Michael J. Pelsmajer, and Radhika Ramamurthi. Graph minors and reliable single message transmission. SIAM Journal on Discrete Mathematics, 19(4):815–847, 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/42112.

Frank:1997:IMM

[FKS97] András Frank, Alexander V.

Karzanov, and András Sebö. On integer multiflow maximization. SIAM Journal on Discrete Mathematics, 10(1): 158–170, February 1997. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fagin:2003:CTL

[FKS03] Ronald Fagin, Ravi Kumar, and D. Sivakumar. Comparing top k lists. SIAM Journal on Discrete Mathematics, 17(1):134–160, January 2003. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fiala:2005:BTT

[FKS05a] Jirí Fiala, Daniel Král’, and Riste Skrekovski. A Brookstype theorem for the generalized list T-coloring. SIAM Journal on Discrete Mathematics, 19(3):588–609, 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/43787.

Frieze:2005:SCI

[FKS05b] Alan Frieze, Michael Krivelevich, and Benny Sudakov. The strong chromatic index of random graphs. SIAM Journal on Discrete Mathematics, 19(3):719–727, 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/44575.

Faria:2012:OCT

[FKS12] Luerbio Faria, Sulamita Klein, and Matej Stehlík. Odd cycle transversals and independent sets in fullerene graphs. SIAM Journal on Discrete Mathematics, 26(3): 1458–1469, ???? 2012. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Frieze:1999:PMH

[FKT99] Alan Frieze, Michal Karonski, and Lubos Thoma. On perfect matchings and Hamilton cycles in sums of random trees. SIAM Journal on Discrete Mathematics, 12(2):208–216, May 1999. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31379.

Fekete:2006:HDP

[FKT06] Sándor P. Fekete, Ekkehard Köhler, and Jürgen Teich. Higher-dimensional packing with order constraints. SIAM Journal on Discrete Mathematics, 20(4):1056– 1078, January 2006. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fuhlbruck:2021:IGS

[FKV21] Frank Fuhlbrück, Johannes Köbler, and Oleg Verbitsky. Identifiability of graphs with small color classes by the Weisfeiler–Leman algorithm. SIAM Journal on Discrete Mathematics, 35(3): 1792–1853, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fu:2010:PAM

[FKW10] Bin Fu, Ming-Yang Kao, and Lusheng Wang. Probabilistic analysis of a motif discovery algorithm for multiple sequences. SIAM Journal on Discrete Mathematics, 23(4): 1715–1737, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Friesen:1991:ACP

[FL91] Donald K. Friesen and Michael A. Langston. Analysis of a compound bin packing algorithm. SIAM Journal on Discrete Mathematics, 4(1): 61–79, February 1991. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fellows:1992:WPO

[FL92] Michael R. Fellows and Michael A. Langston. On well-partial-order theory and its application to combinatorial problems of VLSI design. SIAM Journal on Discrete Mathematics, 5(1):117– 126, February 1992. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fraughnaugh:1996:FIS

[FL96] Kathryn Fraughnaugh and Stephen C. Locke. Finding independent sets in trianglefree graphs. SIAM Journal on Discrete Mathematics, 9 (4):674–681, November 1996. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Forcade:2000:LSC

[FL00] Rod Forcade and Jack Lamoreaux. Lattice-simplex coverings and the 84-shape. SIAM Journal on Discrete Mathematics, 13(2): 194–201, 2000. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/34962.

Fujita:2010:BDN

[FL10] Shinya Fujita and Henry Liu.

The balanced decomposition number and vertex connectivity. SIAM Journal on Discrete Mathematics, 24(4): 1597–1616, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v24/i4/p1597_s1.

Fuchs:2015:WIR

[FL15] Michael Fuchs and Chung-Kuei Lee. The Wiener index of random digital trees. SIAM Journal on Discrete Mathematics, 29(1):586–614, ???? 2015. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Foucaud:2022:RIU

[FL22] Florent Foucaud and Tuomo Lehtilä. Revisiting and improving upper bounds for identifying codes. SIAM Journal on Discrete Mathematics, 36(4):2619–2634, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL

                                        https://epubs.siam.org/
                                        doi/10.1137/22M148999X.

Fleischer:2000:AFM

[Fle00]  Lisa K. Fleischer.  Ap-

proximating fractional multicommodity flow independent of the number of commodities. SIAM Journal on Discrete Mathematics, 13 (4):505–520, 2000. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/35575.

Fleiner:2005:DVG

[Fle05] Balázs Fleiner. Detachment of vertices of graphs preserving edge-connectivity. SIAM Journal on Discrete Mathematics, 18(3):581–591, 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/34151.

Fraigniaud:2010:LBN

[FLL10] Pierre Fraigniaud, Emmanuelle Lebhar, and Zvi Lotker. A lower bound for network navigability. SIAM Journal on Discrete Mathematics, 24(1): 72–81, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fraughnaugh:1995:CGS

[FLM^{+}95] Kathryn F. Fraughnaugh, J. Richard Lundgren, Sarah K. Merz, John S. Maybee, and Norman J. Pullman. Competition graphs of strongly connected and Hamiltonian digraphs. SIAM Journal on Discrete Mathematics, 8(2): 179–185, May 1995. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fernandes:2012:AAG

[FLM12] Maria Elisa Fernandes, Dimitri Leemans, and Mark Mixer. All alternating groups A_{n} with n ≥ 12 have polytopes of rank ⌊\frac{n−1}{2}⌋. SIAM Journal on Discrete Mathematics, 26(2):482–498, ???? 2012. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fomin:2016:HFM

[FLM^{+}16] Fedor V. Fomin, Daniel Lokshtanov, Neeldhara Misra, Geevarghese Philip, and Saket Saurabh. Hitting forbidden minors: Approximation and kernelization. SIAM Journal on Discrete Mathematics, 30(1):383–410, ????

2016. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fomin:2018:MRV

[FLM^{+}18] Fedor V. Fomin, Daniel Lokshtanov, S. M. Meesum, Saket Saurabh, and Meirav Zehavi. Matrix rigidity from the viewpoint of parameterized complexity. SIAM Journal on Discrete Mathematics, 32(2):966–985, ???? 2018. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fraiman:2022:PCB

[FLM22] Nicolas Fraiman, Lyuben Lichev, and Dieter Mitsche. On the power of choice for Boolean functions. SIAM Journal on Discrete Mathematics, 36(3):1943–1954, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL

                                        https://epubs.siam.org/
                                        doi/10.1137/21M1449245.

Fouilhoux:2009:GFI

[FLMY09] Pierre Fouilhoux, Martine Labbé, A. Ridha Mahjoub, and Hande Yaman. Generating facets for the independence system polytope. SIAM Journal on Discrete Mathematics, 23(3):1484–1506, ???? 2009. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Forcey:2010:HSM

[FLS10] Stefan Forcey, Aaron Lauve, and Frank Sottile. Hopf structures on the multiplihedra. SIAM Journal on Discrete Mathematics, 24(4): 1250–1271, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fluck:2024:THC

[Flu24] Eva Fluck. Tangles and hierarchical clustering. SIAM Journal on Discrete Mathematics, 38(1):75–92, January 2024. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Felsner:2011:LED

[FM11] Stefan Felsner and Mareike Massow. Linear extension diameter of downset lattices of two-dimensional posets. SIAM Journal on Discrete Mathematics, 25(1): 112–129, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v25/i1/p112_s1.

Fujita:2013:FRS

[FM13] Shinya Fujita and Colton Magnant. Forbidden rainbow subgraphs that force large highly connected monochromatic subgraphs. SIAM Journal on Discrete Mathematics, 27(3):1625–1637, ???? 2013.

CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Filmus:2024:OSQ

[FM24a] Yuval Filmus and Idan Mehalel. Optimal sets of questions for twenty questions. SIAM Journal on Discrete Mathematics, 38(1): 412–452, January 2024. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Frederickson:2024:CDB

[FM24b] Bryce Frederickson and Lukas Michel. Circuit decompositions of binary matroids. SIAM Journal on Discrete Mathematics, 38(2):1193– 1201, April 2024. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Favaron:1993:HPB

[FMMO93] Odile Favaron, Pedro Mago, Consuelo Maulino, and Oscar Ordaz. Hamiltonian properties of bipartite graphs and digraphs with bipartite independence 2. SIAM Journal on Discrete Mathematics, 6 (2):189–196, May 1993. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fife:2020:UFM

[FMOS20] Tara Fife, Dillon Mayhew, James Oxley, and Charles Semple. The unbreakable frame matroids. SIAM Journal on Discrete Mathematics, 34(3):1522–1537, ???? 2020. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Feng:2008:PCS

[FMP08] Qunqiang Feng, Hosam M.

Mahmoud, and Alois Panholzer. Phase changes in subtree varieties in random recursive and binary search trees. SIAM Journal on Discrete Mathematics, 22(1): 160–184, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Felsner:2017:CPO

[FMP17] Stefan Felsner, Irina Mustata, and Martin Pergel. The complexity of the partial order dimension problem: Closing the gap. SIAM Journal on Discrete Mathematics, 31(1): 172–189, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fishburn:1988:UFD

[FMRR88] Peter C. Fishburn, Helen M.

Marcus-Roberts, and Fred S. Roberts. Unique finite difference measurement. SIAM Journal on Discrete Mathematics, 1(3):334–354, August 1988. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fielbaum:2022:WFP

[FMV22] Andrés Fielbaum, Ignacio Morales, and José Verschae. A water-filling primal-dual algorithm for approximating NonLinear covering problems. SIAM Journal on Discrete Mathematics, 36 (4):2889–2915, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1459964.

Flajolet:1990:SAG

[FO90] Philippe Flajolet and Andrew Odlyzko. Singularity analysis of generating functions. SIAM Journal on Discrete Mathematics, 3(2): 216–240, May 1990. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fleischer:2000:ORI

[FO00] Lisa Fleischer and James B. Orlin. Optimal rounding of instantaneous fractional flows over time. SIAM Journal on Discrete Mathematics, 13(2):145–153, 2000. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/34413.

Feige:2008:FMI

[FO08] Uriel Feige and Eran Ofek. Finding a maximum independent set in a sparse random graph. SIAM Journal on Discrete Mathematics, 22(2): 693–718, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fiorini:2010:VPC

[FOST10] Samuel Fiorini, Gianpaolo Oriolo, Laura Sanità, and Dirk Oliver Theis. The VPN problem with concave costs. SIAM Journal on Discrete Mathematics, 24(3): 1080–1090, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fox:2010:CMI

[Fox10] Jacob Fox. Complete minors and independence number. SIAM Journal on Discrete Mathematics, 24(4): 1313–1321, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Flajolet:1999:SNC

[FP99] Philippe Flajolet and Helmut Prodinger. On Stirling numbers for complex arguments and Hankel contours. SIAM Journal on Discrete Mathematics, 12(2): 155–159, May 1999. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/33259.

Flammini:2001:OGL

[FP01] Michele Flammini and Stéphane Pérennès. On the optimality of general lower bounds for broadcasting and gossiping. SIAM Journal on Discrete Mathematics, 14 (2):267–282, 2001. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/36539.

Flammini:2004:LBB

[FP04] Michele Flammini and Stéphane Pérennès. Lower bounds on the broadcasting and gossiping time of restricted protocols. SIAM Journal on Discrete Mathematics, 17(4):521–540, 2004. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/38645.

Fuhr:2013:PPP

[FP13] Hartmut Führ and Isaac Z. Pesenson. Poincaré and Plancherel–Polya inequalities in harmonic analysis on weighted combinatorial graphs. SIAM Journal on Discrete Mathematics, 27(4): 2007–2028, ???? 2013. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fernandes:2022:DRP

[FP22] Maria Elisa Fernandes and Claudio Alexandre Piedade. The degrees of regular polytopes of type [4, 4, 4]. SIAM Journal on Discrete Mathematics, 36(2):1143–1155, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL

https://epubs.siam.org/
doi/10.1137/20M1375012.

Frankl:2022:EPF

[FPP22] Peter Frankl, János Pach, and Dömötör Pálvölgyi. Exchange properties of finite set-systems. SIAM Journal on Discrete Mathematics, 36 (3):2073–2081, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M145149X.

Florescu:2018:OCD

[FPR18] Laura Florescu, Yuval Peres, and Miklós Z. Rácz. Optimal control for diffusions on graphs. SIAM Journal on Discrete Mathematics, 32(4): 2941–2972, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fox:2013:NEQ

[FPS13] Jacob Fox, János Pach, and Andrew Suk. The number of edges in k-quasi-planar graphs. SIAM Journal on Discrete Mathematics, 27(1):

550–561, ???? 2013. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Frieze:2018:DMW

[FPS18] Alan Frieze, Wesley Pegden, and Gregory B. Sorkin. The distribution of minimumweight cliques and other subgraphs in graphs with random edge weights. SIAM Journal on Discrete Mathematics, 32(3):2115–2133, ???? 2018. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fish:2020:LPC

[FPS20] Sara Fish, Cosmin Pohoata, and Adam Sheffer. Local properties via color energy graphs and forbidden configurations. SIAM Journal on Discrete Mathematics, 34(1): 177–187, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fan:2006:NPL

[FPST06] Kangmin Fan, Simon J. Puglisi, W. F. Smyth, and Andrew Turpin. A new periodicity lemma. SIAM Journal on Discrete Mathematics, 20(3):656–668, January 2006. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Frieze:2022:CTE

[FPT22] Alan M. Frieze, Wesley Pegden, and Tomasz Tkocz. On the cover time of the emerging giant. SIAM Journal on Discrete Mathematics, 36 (3):1687–1710, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1441468.

Fishburn:1994:BIG

[FR94] P. C. Fishburn and J. A.

Reeds. Bell inequalities, Grothendieck’s constant, and root two. SIAM Journal on Discrete Mathematics, 7(1): 48–56, February 1994. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Felsner:1999:LED

[FR99] Stefan Felsner and Klaus Reuter. The linear extension diameter of a poset. SIAM Journal on Discrete Mathematics, 12(3):360–373, August 1999. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/32613.

Fishburn:2006:FCT

[FR06] Peter C. Fishburn and Fred S.

Roberts. Full color theorems for L(2, 1)-colorings. SIAM Journal on Discrete Mathematics, 20(2):428–443, January 2006. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Franzblau:1989:PGS

[Fra89] D. S. Franzblau. Performance guarantees on a sweepline heuristic for covering rectilinear polygons with rectangles. SIAM Journal on Discrete Mathematics, 2(3): 307–321, August 1989. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Frankl:1990:CAC

[Fra90] P. Frankl. Canonical antichains on the circle and applications. SIAM Journal on Discrete Mathematics, 3(3): 355–363, August 1990. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Frank:1992:AGM

[Fra92] András Frank. Augmenting graphs to meet edgeconnectivity requirements. SIAM Journal on Discrete Mathematics, 5(1):25–53, February 1992. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Franzblau:1995:CAL

[Fra95]  D. S. Franzblau.  Com-

binatorial algorithm for a lower bound on frame rigidity. SIAM Journal on Discrete Mathematics, 8(3):388– 400, August 1995. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fraenkel:2010:CIF

[Fra10] Aviezri S. Fraenkel. Complementary iterated floor words and the flora game. SIAM Journal on Discrete Mathematics, 24(2):570–588, ???? 2010. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Frieze:2021:EAR

[Fri21] Alan M. Frieze. The effect of adding randomly weighted edges. SIAM Journal on Discrete Mathematics, 35(2): 1182–1200, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Falgas–Ravry:2015:CTG

[FRMPV15] Victor Falgas-Ravry, Edward Marchant, Oleg Pikhurko, and Emil R. Vaughan. The codegree threshold for 3graphs with independent neighborhoods. SIAM Journal on Discrete Mathematics, 29(3):1504–1539, ???? 2015. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fellows:2009:CWN

[FRRS09] Michael R. Fellows, Frances A.

Rosamond, Udi Rotics, and Stefan Szeider. Clique-width is NP-complete. SIAM Journal on Discrete Mathematics, 23(2):909–939, ???? 2009. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Finbow:2012:RSF

[FRW12] Wendy Finbow, Elissa Ross, and Walter Whiteley. The rigidity of spherical frameworks: Swapping blocks and holes. SIAM Journal on Discrete Mathematics, 26(1): 280–304, ???? 2012. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Falgas-Ravry:2016:CTC

[FRZ16] Victor Falgas-Ravry and Yi Zhao. Codegree thresholds for covering 3-uniform hypergraphs. SIAM Journal on Discrete Mathematics, 30(4): 1899–1917, ???? 2016. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Flajolet:1991:CC

[FS91] P. Flajolet and M. Soria. The cycle construction. SIAM Journal on Discrete Mathematics, 4(1):58–60, February 1991. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Friedlander:2001:DDH

[FS01] John B. Friedlander and Igor E. Shparlinski. On the distribution of Diffie– Hellman triples with sparse exponents. SIAM Journal on Discrete Mathematics, 14(2):162–169, 2001. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/36174.

Feder:2005:DTA

[FS05] Tomás Feder and Carlos Subi.

Disks on a tree: Analysis of a combinatorial game. SIAM Journal on Discrete Mathematics, 19(3):543–552, 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/41800.

Ferrara:2009:GLB

[FS09a] Michael J. Ferrara and John Schmitt. A general lower bound for potentially Hgraphic sequences. SIAM Journal on Discrete Mathematics, 23(1):517–526, ???? 2009. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fox:2009:RTP

[FS09b] Jacob Fox and Benny Sudakov. Ramsey-type problem for an almost monochromatic K_{4}. SIAM Journal on Discrete Mathematics, 23(1): 155–162, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Faller:2012:TDE

[FS12a] Beáta Faller and Mike Steel. Trait-dependent extinction leads to greater expected biodiversity loss. SIAM Journal on Discrete Mathematics, 26 (2):472–481, ???? 2012. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Furedi:2012:OMS

[FS12b] Zoltán Füredi and Attila Sali. Optimal multivalued shattering. SIAM Journal on Discrete Mathematics, 26(2): 737–744, ???? 2012. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fomin:2013:PKP

[FSV13] Fedor V. Fomin, Saket Saurabh, and Yngve Villanger. A polynomial kernel for proper interval vertex deletion. SIAM Journal on Discrete Mathematics, 27(4): 1964–1976, ???? 2013. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Finbow-Singh:2013:IBH

[FSW13] Wendy Finbow-Singh and Walter Whiteley. Isostatic block and hole frameworks. SIAM Journal on Discrete Mathematics, 27(2):

991–1020, ???? 2013. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Forst:2005:WCM

[FT05a] Gunnar Forst and Anders Thorup. What costs are minimized by Huffman trees? SIAM Journal on Discrete Mathematics, 19(1):46–68, 2005. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/42172.

Friedman:2005:GAB

[FT05b] Joel Friedman and Jean-Pierre Tillich. Generalized Alon–Boppana theorems and error-correcting codes. SIAM Journal on Discrete Mathematics, 19(3):700–718, 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/40835.

Farzad:2012:RLC

[FT12] Babak Farzad and Dirk Oliver Theis. Random lifts of K_{5}\e are 3-colorable. SIAM Journal on Discrete Mathematics, 26(1):169–176, ???? 2012. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Francke:2017:CPG

[FT17] Andrea Francke and Csaba D. Tóth. A census of plane graphs with polyline edges. SIAM Journal on Discrete Mathematics, 31(2):1174– 1195, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Frieze:2020:RGF

[FT20] Alan M. Frieze and Tomasz Tkocz. Random graphs with a fixed maximum degree. SIAM Journal on Discrete Mathematics, 34(1): 53–61, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Frieze:2024:CMD

[FTS24] Alan M. Frieze, Krzysztof Turowski, and Wojciech Szpankowski. On the concentration of the maximum degree in the duplication–divergence models. SIAM Journal on Discrete Mathematics, 38(1): 988–1006, March 2024. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fuentes:2023:EPT

[Fue23] Melissa M. Fuentes. The extremality of 2-partite Turán graphs with respect to the number of colorings. SIAM Journal on Discrete Mathematics, 37(4):2523–2543, October 2023. CODEN SJD-MEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fujishige:1997:MMT

[Fuj97] Satoru Fujishige. A minmax theorem for bisubmodular polyhedra. SIAM Journal on Discrete Mathematics, 10 (2):294–308, May 1997. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fukunaga:2016:AGT

[Fuk16] Takuro Fukunaga. Approximating the generalized terminal backup problem via half-integral multiflow relaxation. SIAM Journal on Discrete Mathematics, 30(2): 777–800, ???? 2016. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fulek:2014:END

[Ful14] Radoslav Fulek. Estimating the number of disjoint edges in simple topological graphs via cylindrical drawings. SIAM Journal on Discrete Mathematics, 28(1): 116–121, ???? 2014. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Furedi:1991:MIS

[Für91]  Zoltán Füredi.  Maxi-

mal independent subsets in Steiner systems and in planar sets. SIAM Journal on Discrete Mathematics, 4(2): 196–199, May 1991. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fang:1997:TFH

[FV97] Shao C. Fang and Santosh S. Venkatesh. A threshold function for harmonic update. SIAM Journal on Discrete Mathematics, 10(3): 482–498, August 1997. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fischer:2015:CMP

[FvIKS15] Mareike Fischer, Leo van Iersel, Steven Kelk, and Celine Scornavacca. On computing the maximum parsimony score of a phylogenetic network. SIAM Journal on Discrete Mathematics, 29(1): 559–585, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fishburn:2002:IPR

[FW02] Peter C. Fishburn and Paul E. Wright. Interference patterns in regular graphs with bijective colorings. SIAM Journal on Discrete Mathematics, 15 (3):382–402, 2002. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/38234.

Fox:2020:NCG

[FW20] Jacob Fox and Fan Wei.

On the number of cliques in graphs with a forbidden subdivision or immersion. SIAM Journal on Discrete Mathematics, 34(4):2556–2582, ???? 2020. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fan:2014:FSC

[FXYY14] Genghua Fan, Baogang Xu, Tianjun Ye, and Xingxing Yu. Forbidden subgraphs and 3colorings. SIAM Journal on Discrete Mathematics, 28(3): 1226–1256, ???? 2014. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fu:2023:LBL

[FXZ23] Miao Fu, Fei Xue, and Chuanming Zong. Lower bounds on lattice covering densities of simplices. SIAM Journal on Discrete Mathematics, 37(3):1788–1804, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL

https://epubs.siam.org/
doi/10.1137/22M1514155.

Franklin:2004:SHP

[FY04] Matthew Franklin and Moti Yung. Secure hypergraphs: Privacy from partial broadcast. SIAM Journal on Discrete Mathematics, 18 (3):437–450, 2004. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/33521.

Fujita:2000:SCR

[FYK00] Satoshi Fujita, Masafumi Yamashita, and Tiko Kameda. A study on r-configurations — a resource assignment problem on graphs. SIAM Journal on Discrete Mathematics, 13(2):227–254, 2000. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31132.

Foata:1988:LPW

[FZ88] Dominique Foata and Doron Zeilberger. Laguerre polynomials, weighted derangements, and positivity. SIAM Journal on Discrete Mathematics, 1(4):425–433, November 1988. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Fan:2008:OCN

[FZ08] Genghua Fan and Chuixiang Zhou. Ore condition and nowhere-zero 3-flows. SIAM Journal on Discrete Mathematics, 22(1):288–294, ???? 2008. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Furedi:2022:SUH

[FZ22] Zoltán Füredi and Yi Zhao.

Shadows of 3-uniform hypergraphs under a minimum degree condition. SIAM Journal on Discrete Mathematics, 36(4):2523–2533, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1450227.

Gabow:2004:EDA

[Gab04] Harold N. Gabow. An ear decomposition approach to approximating the smallest 3edge connected spanning subgraph of a multigraph. SIAM Journal on Discrete Mathematics, 18(1):41–70, 2004. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/40547.

Gaber:2005:MTL

[Gab05a] I. Gaber. Minimal-time kline broadcasting. SIAM Journal on Discrete Mathematics, 18(4):769–777, 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/38662.

Gabow:2005:IAA

[Gab05b] Harold N. Gabow. An improved analysis for approximating the smallest k-edge connected spanning subgraph of a multigraph. SIAM Journal on Discrete Mathematics, 19(1):1–18, 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/41491.

Gadouleau:2018:FDS

[Gad18] Maximilien Gadouleau. Finite dynamical systems, hat games, and coding theory. SIAM Journal on Discrete Mathematics, 32(3):1922– 1945, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Galtier:2018:CWS

[Gal18] Jerome Galtier. Computing weighted strength and applications to partitioning. SIAM Journal on Discrete Mathematics, 32(4):2747–2782, ???? 2018. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Gao:2013:DSS

[Gao13] Pu Gao. Distributions of sparse spanning subgraphs in random graphs. SIAM Journal on Discrete Mathematics, 27(1):386–401, ???? 2013. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Gao:2015:MPT

[Gao15] Zhihan Gao. On the metric s–t path traveling salesman problem. SIAM Journal on Discrete Mathematics, 29(3): 1133–1149, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Gao:2018:SPC

[Gao18]  Pu Gao.  The stripping

process can be slow: Part II. SIAM Journal on Discrete Mathematics, 32(2): 1159–1188, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Gardy:1992:NLD

[Gar92] Danièle Gardy. Normal limiting distributions for projection and semijoin sizes. SIAM Journal on Discrete Mathematics, 5(2):219–248, May 1992. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Glynn:2012:GOA

[GB12] David G. Glynn and David Byatt. Graphs for orthogonal arrays and projective planes of even order. SIAM Journal on Discrete Mathematics, 26(3):1076–1087, ???? 2012. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Graver:2011:WDC

[GC11] Jack E. Graver and Gerald T. Cargo. When does a curve bound a distorted disk? SIAM Journal on Discrete Mathematics, 25(1): 280–305, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v25/i1/p280_s1.

Gao:2022:SRL

[GCH22] Guorong Gao, An Chang, and Yuan Hou. Spectral radius on linear r-graphs without expanded K_{r+1}. SIAM Journal on Discrete Mathematics, 36(2):1000–1011, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1404740.

Gabrys:2020:SCS

[GDCM20] R. Gabrys, H. S. Dau, C. J. Colbourn, and O. Milenkovic. Set-codes with small intersections and small discrepancies. SIAM Journal on Discrete Mathematics, 34(2): 1148–1171, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Gainer-Dewar:2017:MHS

[GDVL17] Andrew Gainer-Dewar and Paola Vera-Licona. The minimal hitting set generation problem: Algorithms and computation. SIAM Journal on Discrete Mathematics, 31 (1):63–100, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Gerbner:2023:PPN

[Ger23] Dániel Gerbner. The profile polytope of nontrivial intersecting families. SIAM Journal on Discrete Mathematics, 37(4):2265–2275, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1508984.

Garin:2008:AST

[GF08] Federica Garin and Fabio Fagnani. Analysis of serial turbo codes over Abelian groups for symmetric channels. SIAM Journal on Discrete Mathematics, 22(4): 1488–1526, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Gao:2011:CGC

[GG11] Fei Gao and Gennian Ge.

A complete generalization of Clatworthy group divisible designs. SIAM Journal on Discrete Mathematics, 25(4): 1547–1561, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v25/i4/p1547_s1.

Gadouleau:2015:NCB

[GG15] Maximilien Gadouleau and Nicholas Georgiou. New constructions and bounds for Winkler’s hat game. SIAM Journal on Discrete Mathematics, 29(2):823–834, ???? 2015. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Galanis:2022:CAC

[GGHP22] Andreas Galanis, Leslie A. Goldberg, and Andres Herrera-Poyatos. The complexity of approximating the complex-valued Ising model on bounded degree graphs. SIAM Journal on Discrete Mathematics, 36(3):2159– 2204, ???? 2022. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:/

/epubs.siam.org/doi/10.
1137/21M1454043.

Gomez:2007:CDF

[GGI07] Domingo Gómez, Jaime Gutierrez, and Álvar Ibeas. Cayley digraphs of finite Abelian groups and monomial ideals. SIAM Journal on Discrete Mathematics, 21 (3):763–784, ???? 2007. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Gallet:2021:EPM

[GGLS21] Matteo Gallet, Georg Grasegger, Jan Legerský, and Josef Schicho. On the existence of paradoxical motions of generically rigid graphs on the sphere. SIAM Journal on Discrete Mathematics, 35(1): 325–361, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ghosh:2022:PTN

[GGM^{+}22] Debarun Ghosh, Ervin Györi, Ryan R. Martin, Addisu Paulos, and Chuanqi Xiao. Planar Turán number of the 6-cycle. SIAM Journal on Discrete Mathematics, 36 (3):2028–2050, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M140657X.

Garver:2021:CLE

[GGMM21] Alexander Garver, Stefan Grosser, Jacob P. Matherne, and Alejandro Morales. Counting linear extensions of posets with determinants of hook lengths. SIAM Journal on Discrete Mathematics, 35 (1):205–233, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Goucha:2017:RRP

[GGS17] António Pedro Goucha, João Gouveia, and Pedro M. Silva. On ranks of regular polygons. SIAM Journal on Discrete Mathematics, 31(4): 2612–2625, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Geelen:2006:MC

[GGW06] Jim Geelen, Bert Gerards, and Geoff Whittle. Matroid T -connectivity. SIAM Journal on Discrete Mathematics, 20(3):588–596, January 2006. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Goldschmidt:1990:AOL

[GH90a] Olivier Goldschmidt and Dorit S. Hochbaum. Asymptotically optimal linear algorithm for the minimal k-cut in a random graph. SIAM Journal on Discrete Mathematics, 3(1):58–73, February 1990. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Goldschmidt:1990:FPM

[GH90b] Olivier Goldschmidt and Dorit S. Hochbaum. A fast perfect-matching algorithm in random graphs. SIAM Journal on Discrete Mathematics, 3(1):48–57, February 1990. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Geelen:2006:RBC

[GH06] Jim Geelen and Peter J. Humphries. Rota’s basis conjecture for paving matroids. SIAM Journal on Discrete Mathematics, 20(4): 1042–1045, January 2006. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Goldwasser:2013:ETN

[GH13] John Goldwasser and Ryan Hansen. The exact Turán number of F (3, 3) and all extremal configurations. SIAM Journal on Discrete Mathematics, 27(2):910–917, ???? 2013. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Gaspers:2019:FGC

[GH19] Serge Gaspers and Shenwei Huang. (2P_{2}, K_{4})-free graphs are 4-colorable. SIAM Journal on Discrete Mathematics, 33(2):1095–1120, ???? 2019. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Gunasekara:2022:ESR

[GH22] Ajani De Vas Gunasekara and Daniel Horsley. An Evans-style result for block designs. SIAM Journal on Discrete Mathematics, 36(1):47–63, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1373219.

Goldschmidt:1996:AAC

[GHHY96] Oliver Goldschmidt, Dorit S.

Hochbaum, Cor Hurkens, and Gang Yu. Approximation algorithms for the k-clique covering problem. SIAM Journal on Discrete Mathematics, 9(3):492–509, August 1996.

CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ge:2010:MDD

[GHM10] Gennian Ge, Dawei Huang, and Ying Miao. Monotonic directed designs. SIAM Journal on Discrete Mathematics, 24(1):218–238, ???? 2010. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Gould:2014:MCC

[GHM14] Ronald Gould, Paul Horn, and Colton Magnant. Multiply chorded cycles. SIAM Journal on Discrete Mathematics, 28(1):160–172, ???? 2014. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Gao:2021:SOC

[GHM21] Jun Gao, Qingyi Huo, and Jie Ma. A strengthening on odd cycles in graphs of given chromatic number. SIAM Journal on Discrete Mathematics, 35(4):2317–2327, ???? 2021. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Goranci:2020:IGV

[GHP20] Gramoz Goranci, Monika Henzinger, and Pan Peng. Improved guarantees for vertex sparsification in planar graphs. SIAM Journal on Discrete Mathematics, 34(1):

130–162, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Gorgos:2006:NQT

[GHV06] Ion Gorgos, Chính T. Hoàng, and Vitaly Voloshin. A note on quasi-triangulated graphs. SIAM Journal on Discrete Mathematics, 20(3): 597–602, January 2006. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Golovach:2015:HNG

[GHvtHP15] Petr A. Golovach, Pinar Heggernes, Pim van ’t Hof, and Christophe Paul. Hadwiger number of graphs with small chordality. SIAM Journal on Discrete Mathematics, 29(3): 1427–1451, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Geelen:2005:BSM

[GHW05] Jim Geelen, Petr Hlinený, and Geoff Whittle. Bridging separations in matroids. SIAM Journal on Discrete Mathematics, 18 (3):638–646, 2005. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/39638.

Gupta:1997:BSP

[GI97] Arvind Gupta and Russell Impagliazzo. Bounding the size of planar intertwines. SIAM Journal on Discrete Mathematics, 10(3): 337–358, August 1997. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Gijswijt:2005:IDP

[Gij05] Dion Gijswijt. Integer decomposition for polyhedra defined by nearly totally unimodular matrices. SIAM Journal on Discrete Mathematics, 19(3):798–806, 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/43569.

Gittenberger:1999:CRT

[Git99] Bernhard Gittenberger. On the contour of random trees. SIAM Journal on Discrete Mathematics, 12(4):434–458, November 1999. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/28992.

Griggs:2006:RNG

[GJ06] Jerrold R. Griggs and Xiaohua Teresa Jin. Real number graph labellings with distance conditions. SIAM Journal on Discrete Mathematics, 20(2):

302–327, January 2006. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Griggs:2008:RNC

[GJ08] Jerrold R. Griggs and Xiaohua Teresa Jin. Real number channel assignments for lattices. SIAM Journal on Discrete Mathematics, 22(3): 996–1021, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Gutin:2012:NLS

[GJ12] Gregory Gutin and Mark Jones. Note on large subsets of binary vectors with similar distances. SIAM Journal on Discrete Mathematics, 26(3): 1108–1111, ???? 2012. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Gonze:2016:SPF

[GJ16] François Gonze and Raphaël M.

Jungers. On the synchronizing probability function and the triple rendezvous time for synchronizing automata. SIAM Journal on Discrete Mathematics, 30(2): 995–1014, ???? 2016. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ghandehari:2019:OPS

[GJ19] Mahya Ghandehari and Jeannette Janssen. An optimization parameter for seriation of noisy data. SIAM Journal on Discrete Mathematics, 33 (2):712–730, ???? 2019. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Garamvolgyi:2020:GRU

[GJ20] Dániel Garamvölgyi and Tibor Jordán. Global rigidity of unit ball graphs. SIAM Journal on Discrete Mathematics, 34(1):212–229, ???? 2020. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Groenland:2023:DRP

[GJK^{+}23] Carla Groenland, Tom Johnston, Daniel Korandi, Alexander Roberts, Alex Scott, and Jane Tan. Decomposing random permutations into order-isomorphic subpermutations. SIAM Journal on Discrete Mathematics, 37 (2):1252–1261, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M148029X.

Gutin:2016:MCP

[GJW16] Gregory Gutin, Mark Jones, and Magnus Wahlström. The Mixed Chinese Postman Problem parameterized by pathwidth and treedepth. SIAM Journal on Discrete Mathematics, 30(4):2177– 2205, ???? 2016. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Gu:2023:RLV

[GJWZ23] Yangyan Gu, Yiting Jiang, David R. Wood, and Xuding Zhu. Refined list version of Hadwiger’s conjecture. SIAM Journal on Discrete Mathematics, 37 (3):1738–1750, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1522413.

Ge:2020:CIE

[GJXZ20] Gennian Ge, Yifan Jing, Zixiang Xu, and Tao Zhang. Color isomorphic even cycles and a related Ramsey problem. SIAM Journal on Discrete Mathematics, 34(3): 1999–2008, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Guichard:1992:PLG

[GK92] David R. Guichard and John W. Krussel. Pair labelings of graphs. SIAM Journal on Discrete Mathematics, 5(1):144–149, February 1992. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Goldberg:1997:GPU

[GK97] Andrew V. Goldberg and Robert Kennedy. Global price updates help. SIAM Journal on Discrete Mathematics, 10 (4):551–572, November 1997. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Garnier:2002:ELS

[GK02] Josselin Garnier and Leila Kallel. Efficiency of local search with multiple local optima. SIAM Journal on Discrete Mathematics, 15(1):122–141, 2002. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/35522.

Guruswami:2004:HCC

[GK04] Venkatesan Guruswami and Sanjeev Khanna. On the hardness of 4-coloring a 3colorable graph. SIAM Journal on Discrete Mathematics, 18(1):30–40, 2004. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/37679.

Gartner:2007:TNB

[GK07] Bernd Gärtner and Volker Kaibel. Two new bounds for the random-edge simplexalgorithm. SIAM Journal on Discrete Mathematics, 21(1): 178–190, ???? 2007. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Glebov:2013:NHC

[GK13a] Roman Glebov and Michael Krivelevich. On the number of Hamilton cycles in sparse random graphs. SIAM Journal on Discrete Mathematics, 27(1):27–42, ???? 2013. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Gottlieb:2013:PAN

[GK13b] Lee-Ad Gottlieb and Robert Krauthgamer. Proximity algorithms for nearly doubling spaces. SIAM Journal on Discrete Mathematics, 27(4): 1759–1769, ???? 2013. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Guzman:2016:CFG

[GK16] Johnny Guzmán and Caroline Klivans. Chip firing on general invertible matrices. SIAM Journal on Discrete Mathematics, 30(2): 1115–1127, ???? 2016. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Galby:2023:MDP

[GKI^{+}23] Esther Galby, Liana Khazaliya, Fionn Mc Inerney, Roohani Sharma, and Prafullkumar Tale. Metric dimension parameterized by feedback vertex set and other structural parameters. SIAM Journal on Discrete Mathematics, 37(4):2241–2264, ????

2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL

https://epubs.siam.org/
doi/10.1137/22M1510911.

Gyarfas:1999:LCG

[GKL99] András Gyárfás, Zoltán Király, and Jeno Lehel. Online 3-chromatic graphs I. triangle-free graphs. SIAM Journal on Discrete Mathematics, 12(3):385–411, August 1999. CODEN SJD-MEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/31030.

Gimbel:2019:FCM

[GKLT19] John Gimbel, André Kündgen, Binlong Li, and Carsten Thomassen. Fractional coloring methods with applications to degenerate graphs and graphs on surfaces. SIAM Journal on Discrete Mathematics, 33(3):1415–1430, ???? 2019. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Garrod:2012:HCB

[GKM12] Bryn Garrod, Grzegorz Kubicki, and Michal Morayne. How to choose the best twins. SIAM Journal on Discrete Mathematics, 26(1): 384–398, ???? 2012. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Gerbner:2018:LPF

[GKM^{+}18] Dániel Gerbner, Balázs Keszegh, Gábor Mészáros, Balázs Patkós, and Máté Vizer. Line percolation in finite projective planes. SIAM Journal on Discrete Mathematics, 32 (2):864–881, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Goresky:2004:DS

[GKMS04] Mark Goresky, Andrew Klapper, Ram Murty, and Igor Shparlinski. On decimations of ℓ-sequences. SIAM Journal on Discrete Mathematics, 18(1):130–140, 2004. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/40342.

Goke:2022:HWE

[GKMS22] Alexander Göke, Jochen Koenemann, Matthias Mnich, and Hao Sun. Hitting weighted even cycles in planar graphs. SIAM Journal on Discrete Mathematics, 36 (4):2830–2862, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M144894X.

Golovach:2017:PCG

[GKMT17] Petr A. Golovach, Marcin Kamiński, Spyridon Maniatis, and Dimitrios M. Thilikos. The parameterized complexity of graph cyclability. SIAM Journal on Discrete Mathematics, 31(1): 511–541, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Guo:2010:MRM

[GKNU10] Jiong Guo, Christian Komusiewicz, Rolf Niedermeier, and Johannes Uhlmann. A more relaxed model for graph-based data clustering: s-plex cluster editing. SIAM Journal on Discrete Mathematics, 24(4): 1662–1683, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v24/i4/p1662_s1.

Grzesik:2023:QFO

[GkoKK23] Andrzej Grzesik, Daniel Il kovi, Bart omiej Kielak, and Daniel Král. Quasirandomforcing orientations of cycles. SIAM Journal on Discrete Mathematics, 37(4): 2689–2716, November 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Girao:2019:PSC

[GKP19] António Girão, Teeradej Kittipassorn, and Kamil Popielarz. Partite saturation of complete graphs. SIAM Journal on Discrete Mathematics, 33(4):2346–2359, ???? 2019. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Gerbner:2018:NCG

[GKPP18] Dániel Gerbner, Balázs Keszegh, Cory Palmer, and Balázs Patkós. On the number of cycles in a graph with restricted cycle lengths. SIAM Journal on Discrete Mathematics, 32 (1):266–279, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Grohe:2018:CCN

[GKR^{+}18] Martin Grohe, Stephan Kreutzer, Roman Rabinovich, Sebastian Siebertz, and Konstantinos Stavropoulos. Coloring and covering nowhere dense graphs. SIAM Journal on Discrete Mathematics, 32(4):2467–2481, ???? 2018. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Galvin:2015:PCT

[GKRS15] David Galvin, Jeff Kahn, Dana Randall, and Gregory B. Sorkin. Phase coexistence and torpid mixing in the 3-coloring model on Z^{d}. SIAM Journal on Discrete Mathematics, 29(3): 1223–1244, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Grigorescu:2012:SRC

[GKS12] Elena Grigorescu, Tali Kaufman, and Madhu Sudan. Succinct representation of codes with applications to testing. SIAM Journal on Discrete Mathematics, 26(4): 1618–1634, ???? 2012. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ganian:2022:AAT

[GKS22] Robert Ganian, Eun Jung Kim, and Stefan Szeider. Algorithmic applications of tree-cut width. SIAM Journal on Discrete Mathematics, 36(4):2635–2666, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/20M137478X.

Grigorescu:2019:FMF

[GKW19] Elena Grigorescu, Akash Kumar, and Karl Wimmer. Flipping out with many flips: Hardness of testing kmonotonicity. SIAM Journal on Discrete Mathematics, 33(4):2111–2125, ???? 2019. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Gould:2006:MDI

[GKY06] Ronald J. Gould, Alexandr Kostochka, and Gexin Yu. On minimum degree implying that a graph is Hlinked. SIAM Journal on Discrete Mathematics, 20(4): 829–840, January 2006. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Gonzalez:1995:CSS

[GL95] Jaime González and Osvaldo Landaeta. A competitive strong spanning tree algorithm for the maximum bipartite matching problem. SIAM Journal on Discrete Mathematics, 8(2): 186–195, May 1995. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ge:2008:EDA

[GL08] Gennian Ge and Alan C. H.

Ling. On the existence of (K_{5} \ e)-designs with application to optical networks. SIAM Journal on Discrete Mathematics, 21(4): 851–864, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Grynkiewicz:2010:SSC

[GL10] David J. Grynkiewicz and Vsevolod F. Lev. 1-saturating sets, caps, and doublingcritical sets in binary spaces. SIAM Journal on Discrete Mathematics, 24(1):169–190, ???? 2010. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Grimmett:2014:SIC

[GL14] Geoffrey R. Grimmett and Zhongyang Li. Strict inequalities for connective constants of transitive graphs. SIAM Journal on Discrete Mathematics, 28(3):1306–1333, ???? 2014. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Grytczuk:2015:SMN

[GL15] Jaroslaw Grytczuk and Wojciech Lubawski. Splitting multidimensional necklaces and measurable colorings of Euclidean spaces. SIAM Journal on Discrete Mathematics, 29(1):252–258, ???? 2015. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Guruswami:2017:ITP

[GL17] Venkatesan Guruswami and Euiwoong Lee. Inapproximability of H-transversal/ packing. SIAM Journal on Discrete Mathematics, 31(3): 1552–1571, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Guo:2022:SRM

[GLM22] Mingyang Guo, Hongliang Lu, and Dingjia Mao. A stability result on matchings in 3-uniform hypergraphs. SIAM Journal on Discrete Mathematics, 36 (3):2339–2351, ???? 2022.

CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1422720.

Gerbner:2012:AIF

[GLP^{+}12] Dániel Gerbner, Nathan Lemons, Cory Palmer, Balázs Patkós, and Vajk Szécsi. Almost intersecting families of sets. SIAM Journal on Discrete Mathematics, 26(4): 1657–1669, ???? 2012. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Guo:2023:ARN

[GLP23] Mingyang Guo, Hongliang Lu, and Xing Peng. Anti-Ramsey number of matchings in 3-uniform hypergraphs. SIAM Journal on Discrete Mathematics, 37 (3):1970–1987, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1503178.

Glebov:2015:CDN

[GLS15] Roman Glebov, Anita Liebenau, and Tibor Szabó. On the concentration of the domination number of the random graph. SIAM Journal on Discrete Mathematics, 29(3): 1186–1206, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Gu:2020:ARN

[GLS20] Ran Gu, Jiaao Li, and Yongtang Shi. Anti-Ramsey numbers of paths and cycles in hypergraphs. SIAM Journal on Discrete Mathematics, 34 (1):271–307, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Giannopoulou:2016:TDS

[GLSS16] Archontia C. Giannopoulou, Daniel Lokshtanov, Saket Saurabh, and Ondrej Suchý. Tree deletion set has a polynomial kernel but no OPT^{O(∞)} approximation. SIAM Journal on Discrete Mathematics, 30(3):1371– 1384, ???? 2016. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Gritzmann:2011:UDT

[GLW11] Peter Gritzmann, Barbara Langfeld, and Markus Wiegelmann. Uniqueness in discrete tomography: Three remarks and a corollary. SIAM Journal on Discrete Mathematics, 25(4):1589– 1599, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sidma/resource/
                                        1/sjdmec/v25/i4/p1589_s1.

Graham:2007:OTS

[GLY07] Ronald L. Graham, Minming Li, and Frances F. Yao. Optimal tree structures for group key management with batch updates. SIAM Journal on Discrete Mathematics, 21(2): 532–547, ???? 2007. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Glynn:2010:CAT

[Gly10] David G. Glynn. The conjectures of Alon–Tarsi and Rota in dimension prime minus one. SIAM Journal on Discrete Mathematics, 24(2): 394–399, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Glynn:2012:IHP

[Gly12] David G. Glynn. An invariant for hypersurfaces in prime characteristic. SIAM Journal on Discrete Mathematics, 26 (3):881–883, ???? 2012. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Grotschel:1990:IPA

[GM90] Martin Grötschel and Clyde L.

Monma. Integer polyhedra arising from certain network design problems with connectivity constraints. SIAM Journal on Discrete Mathematics, 3(4):502–523, November 1990. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Goles:1991:LFA

[GM91] Eric Goles and Servet Martínez. Lyapunov functionals for automata networks defined by cyclically monotone functions. SIAM Journal on Discrete Mathematics, 4(2): 200–206, May 1991. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). See errata [GM93].

Goles:1993:ELF

[GM93] E. Goles and S. Martínez. Errata: “Lyapunov functionals for automata networks defined by cyclically monotone functions” [SIAM J. Discrete Math. 4 (1991), no. 2, 200– 206; MR 91m:68135]. SIAM Journal on Discrete Mathematics, 6(4):677, November 1993. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). See [GM91].

Grone:1994:LSG

[GM94] Robert Grone and Russell Merris. The Laplacian spectrum of a graph. II. SIAM Journal on Discrete Mathematics, 7(2):221–229, May 1994. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Georges:2003:RGO

[GM03] John P. Georges and David W. Mauro. On regular graphs optimally labeled with a condition at distance two. SIAM Journal on Discrete Mathematics, 17(2):320–331, 2003. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/39124.

Gerke:2004:GIC

[GM04] Stefanie Gerke and Colin McDiarmid. Graph imperfection with a co-site constraint. SIAM Journal on Discrete Mathematics, 17(3):403–425, 2004. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/40251.

Georges:2005:SGN

[GM05] John P. Georges and David W.

Mauro. On the structure of graphs with non-surjective L(2, 1)-labelings. SIAM Journal on Discrete Mathematics, 19(1):208–223, 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/42980.

Greenhill:2013:AES

[GM13] Catherine Greenhill and Brendan D. McKay. Asymptotic enumeration of sparse multigraphs with given degrees. SIAM Journal on Discrete Mathematics, 27(4): 2064–2089, ???? 2013. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Giannopoulou:2016:NGR

[GM16] Archontia C. Giannopoulou and George B. Mertzios. New geometric representations and domination problems on tolerance and multitolerance graphs. SIAM Journal on Discrete Mathematics, 30(3):1685–1725, ???? 2016. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Gao:2020:CVV

[GM20] Jun Gao and Jie Ma. A conjecture of Verstraëte on vertex-disjoint cycles. SIAM Journal on Discrete Mathematics, 34(2):1290–1301, ???? 2020. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Garcia-Marco:2015:MTI

[GMA15] Ignacio García-Marco and Jorge Luis Ramírez Alfonsín. Matroid toric ideals: Complete intersection, minors, and minimal systems of generators. SIAM Journal on Discrete Mathematics, 29(4): 2267–2276, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Gerbner:2020:RPB

[GMOV20] Dániel Gerbner, Abhishek Methuku, Gholamreza Omidi, and Máté Vizer. Ramsey problems for Berge hypergraphs. SIAM Journal on Discrete Mathematics, 34(1): 351–369, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Gutin:2021:PPC

[GMOW21] Gregory Gutin, Diptapriyo Majumdar, Sebastian Ordyniak, and Magnus Wahlström. Parameterized pre-coloring extension and list coloring problems. SIAM Journal on Discrete Mathematics, 35(1): 575–596, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Gao:2022:PMS

[GMP22] Pu Gao, Calum MacRury, and Paweł Prałat. Perfect matchings in the semirandom graph process. SIAM Journal on Discrete Mathematics, 36(2):1274–1290, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1446939.

Goyal:2015:DAM

[GMPZ15] Prachi Goyal, Neeldhara Misra, Fahad Panolan, and Meirav Zehavi. Deterministic algorithms for matching and packing problems based on representative sets. SIAM Journal on Discrete Mathematics, 29(4):1815–1836, ????

2015. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Gu:2021:SCD

[GMR^{+}21] Xiaofeng Gu, Wei Meng, Martin Rolek, Yue Wang, and Gexin Yu. Sufficient conditions for 2-dimensional global rigidity. SIAM Journal on Discrete Mathematics, 35(4): 2520–2534, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Galvin:2011:MHC

[GMRT11] David Galvin, Fabio Martinelli, Kavita Ramanan, and Prasad Tetali. The multistate hard core model on a regular tree. SIAM Journal on Discrete Mathematics, 25(2): 894–915, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sidma/resource/
                                        1/sjdmec/v25/i2/p894_s1.

Georges:2000:LPC

[GMS00] John P. Georges, David W. Mauro, and Melanie I. Stein. Labeling products of complete graphs with a condition at distance two. SIAM Journal on Discrete Mathematics, 14(1):28–35, 2000. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/35185.

Grzesik:2015:DPL

[GMS15] Andrzej Grzesik, Michal Morayne, and Małgorzata Sulkowska. From directed path to linear order — the best choice problem for powers of directed path. SIAM Journal on Discrete Mathematics, 29(1):500–513, ???? 2015. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Gross:2015:LCC

[GMTW15] Jonathan L. Gross, Toufik Mansour, Thomas W. Tucker, and David G. L. Wang. Logconcavity of combinations of sequences and applications to genus distributions. SIAM Journal on Discrete Mathematics, 29(2):1002–1029, ???? 2015. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Gouveia:2019:SRS

[GMTW19] João Gouveia, Antonio Macchia, Rekha R. Thomas, and Amy Wiebe. The slack realization space of a polytope. SIAM Journal on Discrete Mathematics, 33(3): 1637–1653, ???? 2019. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Grotschel:1996:PST

[GMW96] M. Grötschel, A. Martin, and R. Weismantel. Packing Steiner trees: separation algorithms. SIAM Journal on Discrete Mathematics, 9 (2):233–257, May 1996. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ge:2005:CCO

[GMW05] Gennian Ge, Ying Miao, and Lihua Wang. Combinatorial constructions for optimal splitting authentication codes. SIAM Journal on Discrete Mathematics, 18(4):663–678, 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/43546.

Ge:2009:BSS

[GMZ09] Gennian Ge, Ying Miao, and Xiande Zhang. On block sequences of Steiner quadruple systems with error correcting consecutive unions. SIAM Journal on Discrete Mathematics, 23(2):940–958, ???? 2009. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Godsil:2008:COG

[GN08] C. D. Godsil and M. W. Newman. Coloring an orthogonality graph. SIAM Journal on Discrete Mathematics, 22 (2):683–692, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Gimenez:2006:CTP

[GnN06] Omer Giménez, Petr Hlin ný, and Marc Noy. Computing the Tutte polynomial on graphs of bounded cliquewidth. SIAM Journal on Discrete Mathematics, 20(4): 932–946, January 2006. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Gerbner:2020:WBH

[GNPV20] Dániel Gerbner, Dániel T.

Nagy, Balázs Patkós, and Máté Vizer. t-wise Berge and t-heavy hypergraphs. SIAM Journal on Discrete Mathematics, 34(3):1813–1829, ???? 2020. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Guo:2011:PCA

[GNS11] Jiong Guo, Rolf Niedermeier, and Ondrej Suchý. Parameterized complexity of arc-weighted directed Steiner problems. SIAM Journal on Discrete Mathematics, 25(2): 583–599, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v25/i2/p583_s1.

Guneri:2012:BMD

[GÖ12] Cem Günerí and Ferruh Özbudak. A bound on the minimum distance of quasicyclic codes. SIAM Journal on Discrete Mathematics, 26(4):1781–1796, ???? 2012. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Goemans:2001:AES

[Goe01] Michel X. Goemans. Approximate edge splitting. SIAM Journal on Discrete Mathematics, 14(1):138–141, 2001. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/35802.

Golic:1996:CEP

[Gol96] Jovan Dj. Golić. Constrained embedding probability for two binary strings. SIAM Journal on Discrete Mathematics, 9(3):360–364, August 1996. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Goldberg:2006:BPR

[Gol06] Paul W. Goldberg. A bound on the precision required to estimate a Boolean perceptron from its average satisfying assignment. SIAM Journal on Discrete Mathematics, 20(2):328–343, January 2006. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Gordon:1993:DLU

[Gor93] Daniel M. Gordon. Discrete logarithms in GF(p) using the number field sieve. SIAM Journal on Discrete Mathematics, 6(1):124–138, February 1993. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Gelles:2020:EEC

[GOR20] Ran Gelles, Rafail Ostrovsky, and Alan Roytman. Efficient error-correcting codes for sliding windows. SIAM Journal on Discrete Mathematics, 34(1):904–937, ???? 2020. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Gottlieb:2003:HED

[Got03] Eric Gottlieb. On the homology of the h, k-equal Dowling lattice. SIAM Journal on Discrete Mathematics, 17(1):50–71, 2003. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/42248.

Grigni:1991:TBM

[GP91] Michelangelo Grigni and David Peleg. Tight bounds on minimum broadcast networks. SIAM Journal on Discrete Mathematics, 4(2): 207–222, May 1991. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Gavoille:1999:CIR

[GP99] Cyril Gavoille and David Peleg. The compactness of interval routing. SIAM Journal on Discrete Mathematics, 12 (4):459–473, November 1999. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/32863.

Gavoille:2008:ODL

[GP08a] Cyril Gavoille and Christophe Paul. Optimal distance labeling for interval graphs and related graph families. SIAM Journal on Discrete Mathematics, 22(3):1239–1258, ???? 2008. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Gerbner:2008:CPV

[GP08b] Dániel Gerbner and Balázs Patkós. ℓ-chain profile vectors. SIAM Journal on Discrete Mathematics, 22(1): 185–193, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Gilboa:2014:UAP

[GP14] Shoni Gilboa and Rom Pinchasi. On the union of arithmetic progressions. SIAM Journal on Discrete Mathematics, 28(3):1062–1073, ???? 2014. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Gutin:2016:PTS

[GP16] Gregory Gutin and Viresh Patel. Parameterized traveling salesman problem: Beating the average. SIAM Journal on Discrete Mathematics, 30(1):220–238, ???? 2016. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Gerbner:2017:ERB

[GP17] Dániel Gerbner and Cory Palmer. Extremal results for Berge hypergraphs. SIAM Journal on Discrete Mathematics, 31(4):2314–2327, ???? 2017. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Girao:2018:LIS

[GP18] António Girão and Kamil Popielarz. Large induced subgraphs with k vertices of almost maximum degree. SIAM Journal on Discrete Mathematics, 32(1):450–454, ???? 2018. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Glazyrin:2020:RMF

[GP20] Alexey Glazyrin and Josiah Park. Repeated minimizers of p-frame energies. SIAM Journal on Discrete Mathematics, 34(4):2411–2423, ???? 2020. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Gasieniec:2001:WPS

[GPP01] Leszek Gasieniec, Andrzej Pelc, and David Peleg. The wakeup problem in synchronous broadcast systems. SIAM Journal on Discrete Mathematics, 14 (2):207–222, 2001. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/37602.

Galil:2004:TDP

[GPP04] Zvi Galil, Jong Geun Park, and Kunsoo Park. Threedimensional periodicity and its application to pattern matching. SIAM Journal on Discrete Mathematics, 18(2):362–381, 2004. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/39030.

Giannopoulou:2021:LKE

[GPR^{+}21]  Archontia  Giannopoulou,

Michał Pilipczuk, Jean-Florent Raymond, Dimitrios M. Thilikos, and Marcin Wrochna. Linear kernels for edge deletion problems to immersion-closed graph classes. SIAM Journal on Discrete Mathematics, 35(1): 105–151, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Goncalves:2011:DNG

[GPRT11] Daniel Gonçalves, Alexandre Pinlou, Michaël Rao, and Stéphan Thomassé. The domination number of grids. SIAM Journal on Discrete Mathematics, 25(3):1443– 1453, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v25/i3/p1443_s1.

Goldberg:1988:PSB

[GPS88] Andrew V. Goldberg, Serge A.

Plotkin, and Gregory E. Shannon. Parallel symmetrybreaking in sparse graphs. SIAM Journal on Discrete Mathematics, 1(4):434–446, November 1988. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Gravner:2019:MST

[GPS19] Janko Gravner, J. E. Paguyo, and Erik Slivken. Maximal spanning time for neighborhood growth on the Hamming plane. SIAM Journal on Discrete Mathematics, 33 (2):976–993, ???? 2019. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Goldberg:2001:BAG

[GPSS01] Leslie Ann Goldberg, Mike Paterson, Aravind Srinivasan, and Elizabeth Sweedyk. Better approximation guarantees for job-shop scheduling. SIAM Journal on Discrete Mathematics, 14 (1):67–92, 2001. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/32610.

Garnero:2015:ELK

[GPST15] Valentin Garnero, Christophe Paul, Ignasi Sau, and Dimitrios M. Thilikos. Explicit linear kernels via dynamic programming. SIAM Journal on Discrete Mathematics, 29(4):1864–1894, ???? 2015. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Golovach:2015:IDP

[GPvL15] Petr A. Golovach, Daniël Paulusma, and Erik Jan van Leeuwen. Induced disjoint paths in claw-free graphs. SIAM Journal on Discrete Mathematics, 29(1): 348–375, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Graves:2009:GEH

[GPW09] Stephen Graves, Tomaž Pisanski, and Mark E. Watkins. Growth of edgehomogeneous tessellations. SIAM Journal on Discrete Mathematics, 23(1): 1–18, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Guenin:2013:RBP

[GPW13] Bertrand Guenin, Irene Pivotto, and Paul Wollan. Relationships between pairs of representations of signed binary matroids. SIAM Journal on Discrete Mathematics, 27 (1):329–341, ???? 2013. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Goemans:2002:SMS

[GQS^{+}02] Michel X. Goemans, Maurice Queyranne, Andreas S. Schulz, Martin Skutella, and Yaoguang Wang. Single machine scheduling with release dates. SIAM Journal on Discrete Mathematics, 15(2):165–192, 2002. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/36223.

Goldberg:1999:FUU

[GR99] Andrew V. Goldberg and Satish Rao. Flows in undirected unit capacity networks. SIAM Journal on Discrete Mathematics, 12(1): 1–5, February 1999. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/33103.

Giscard:2017:ACT

[GR17] P.-L. Giscard and P. Rochet. Algebraic combinatorics on trace monoids: Extending number theory to walks on graphs. SIAM Journal on Discrete Mathematics, 31(2): 1428–1453, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Graver:1991:RM

[Gra91] Jack E. Graver. Rigidity matroids. SIAM Journal on Discrete Mathematics, 4(3): 355–368, August 1991. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Graver:2004:EFG

[Gra04] Jack E. Graver. Encoding fullerenes and geodesic domes. SIAM Journal on Discrete Mathematics, 17 (4):596–614, 2004. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/39104.

Gray:2007:VMT

[Gra07] Ian D. Gray. Vertex-magic total labelings of regular graphs. SIAM Journal on Discrete Mathematics, 21(1): 170–177, ???? 2007. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Gadouleau:2015:FPB

[GRR15] Maximilien Gadouleau, Adrien Richard, and Søren Riis.

Fixed points of Boolean networks, guessing graphs, and coding theory. SIAM Journal on Discrete Mathematics, 29(4):2312–2335, ???? 2015. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Gonen:2011:CSO

[GRS11] Mira Gonen, Dana Ron, and Yuval Shavitt. Counting stars and other small subgraphs in sublinear-time. SIAM Journal on Discrete Mathematics, 25(3):1365– 1411, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v25/i3/p1365_s1.

Gupta:2012:MRI

[GRS12] Sushmita Gupta, Venkatesh Raman, and Saket Saurabh. Maximum r-regular induced subgraph problem: Fast exponential algorithms and combinatorial bounds. SIAM Journal on Discrete Mathematics, 26(4):1758–1780, ???? 2012. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Gruica:2023:RMC

[GRSZ23] Anina Gruica, Alberto Ravagnani, John Sheekey, and Ferdinando Zullo. Rankmetric codes, semifields, and the average critical problem. SIAM Journal on Discrete Mathematics, 37 (2):1079–1117, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1486893.

Grujic:2017:QFN

[Gru17] Vladimir Grujić. Quasisymmetric functions for nestohedra. SIAM Journal on Discrete Mathematics, 31(4): 2570–2585, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Gutin:2008:MCH

[GRY08] Gregory Gutin, Arash Rafiey, and Anders Yeo. Minimum cost homomorphisms to semicomplete bipartite digraphs. SIAM Journal on Discrete Mathematics, 22(4): 1624–1639, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Goldberg:1989:CMI

[GS89] Mark Goldberg and Thomas Spencer. Constructing a maximal independent set in parallel. SIAM Journal on Discrete Mathematics, 2(3):322– 328, August 1989. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Goldberg:1993:EPA

[GS93a] Mark Goldberg and Thomas Spencer. An efficient parallel algorithm that finds independent sets of guaranteed size. SIAM Journal on Discrete Mathematics, 6(3):443– 459, August 1993. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Gritzmann:1993:MAP

[GS93b] Peter Gritzmann and Bernd Sturmfels. Minkowski addition of polytopes: Computational complexity and applications to Gröbner bases. SIAM Journal on Discrete Mathematics, 6(2):246–269, May 1993. CODEN SJD-MEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Gao:1994:SSE

[GS94a] Jun Gao and J. Michael Steele. Sums of squares of edge lengths and spacefilling curve heuristics for the traveling salesman problem. SIAM Journal on Discrete Mathematics, 7(2):314–324, May 1994. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Goh:1994:RSP

[GS94b] William M. Y. Goh and Eric Schmutz. Random set partitions. SIAM Journal on Discrete Mathematics, 7(3): 419–436, August 1994. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Greenberg:1995:FOO

[GS95] Ronald I. Greenberg and Jau Der Shih. Feasible offset and optimal offset for general single-layer channel routing. SIAM Journal on Discrete Mathematics, 8(4):543– 554, November 1995. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Gerards:1998:GAS

[GS98] A. M. H. Gerards and F. B. Shepherd. The graphs with all subgraphs T -perfect. SIAM Journal on Discrete Mathematics, 11(4):524–545, November 1998. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ginsburg:2000:CDC

[GS00] John Ginsburg and Bill Sands. Completely disconnecting the complete graph. SIAM Journal on Discrete Mathematics, 13 (1):33–47, 2000. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/32655.

Gijswijt:2003:SSP

[GS03] Dion Gijswijt and Alexander Schrijver. On the bstable set polytope of graphs without bad K_{4}. SIAM Journal on Discrete Mathematics, 16(3):511–516, 2003. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/41734.

Goemans:2013:ASS

[GS13] Michel X. Goemans and José A. Soto. Algorithms for symmetric submodular function minimization under hereditary constraints and generalizations. SIAM Journal on Discrete Mathematics, 27(2):1123–1145, ???? 2013. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Grohe:2016:CT

[GS16a] Martin Grohe and Pascal Schweitzer. Computing with tangles. SIAM Journal on Discrete Mathematics, 30(2): 1213–1247, ???? 2016. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Guenin:2016:SCF

[GS16b] Bertrand Guenin and Leanne Stuive. Single commodityflow algorithms for lifts of graphic and CoGraphic matroids. SIAM Journal on Discrete Mathematics, 30(3): 1775–1797, ???? 2016. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Guruswami:2020:RCH

[GS20] Venkatesan Guruswami and Sai Sandeep. Rainbow coloring hardness via low sensitivity polymorphisms. SIAM Journal on Discrete Mathematics, 34(1):520–537, ???? 2020. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Granot:1991:NAR

[GSK91] Daniel Granot and Darko Skorin-Kapov. NC algorithms for recognizing partial 2-trees and 3-trees. SIAM Journal on Discrete Mathematics, 4(3): 342–354, August 1991. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Galeana-Sanchez:1998:SKD

[GSL98] H. Galeana-Sánchez and Xueliang Li. Semikernels and (k, l)-kernels in digraphs. SIAM Journal on Discrete Mathematics, 11(2): 340–346, May 1998. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Galeana-Sanchez:1991:SQK

[GSPRM91] H. Galeana-Sánchez, L. Pastrana Ramírez, and H. A. Rincón-Mejía. Semikernels, quasi kernels, and Grundy functions in the line digraph. SIAM Journal on Discrete Mathematics, 4(1): 80–83, February 1991. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Goldman:2014:GFB

[GSS14] Ron Goldman, Plamen Simeonov, and Yilmaz Simsek. Generating functions for the q-Bernstein bases. SIAM Journal on Discrete Mathematics, 28(3):1009–1025, ???? 2014. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Guruswami:2015:IMV

[GSS15] Venkatesan Guruswami, Sushant Sachdeva, and Rishi Saket. Inapproximability of minimum vertex cover on kuniform k-partite hypergraphs. SIAM Journal on Discrete Mathematics, 29(1): 36–58, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Golovach:2023:CLA

[GST23] Petr A. Golovach, Giannos Stamoulis, and Dimitrios M. Thilikos. Combing a linkage in an annulus. SIAM Journal on Discrete Mathematics, 37(4):2332–2364, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M150914X.

Gregor:2012:QLH

[GSV12] Petr Gregor, Riste Skrekovski, and Vida Vukasinović. Queue layouts of hypercubes. SIAM Journal on Discrete Mathematics, 26(1):77–88, ???? 2012. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Getu:1992:HGG

[GSWW92] Seyoum Getu, Louis W. Shapiro, Wen-Jin Woan, and Leon C. Woodson. How to guess a generating function. SIAM Journal on Discrete Mathematics, 5(4):497– 499, November 1992. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Gupta:2022:TSM

[GSZ22] Sushmita Gupta, Saket Saurabh, and Meirav Zehavi. On treewidth and stable marriage: Parameterized algorithms and hardness results (complete characterization). SIAM Journal on Discrete Mathematics, 36 (1):596–681, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/19M130491X.

Gimbel:1998:WOS

[GT98] John G. Gimbel and Ann N. Trenk. On the weakness of an ordered set. SIAM Journal on Discrete Mathematics, 11 (4):655–663, November 1998. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Goldwasser:2012:VRP

[GT12] John Goldwasser and John Talbot. Vertex Ramsey problems in the hypercube. SIAM Journal on Discrete Mathematics, 26(2):838–853, ???? 2012. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Giannopoulou:2013:OGM

[GT13] Archontia C. Giannopoulou and Dimitrios M. Thilikos. Optimizing the graph minors weak structure theorem. SIAM Journal on Discrete Mathematics, 27(3): 1209–1227, ???? 2013. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Gu:2018:SRS

[Gu18] Xiaofeng Gu. Spanning rigid subgraph packing and sparse subgraph covering. SIAM Journal on Discrete Mathematics, 32(2):1305–1313, ???? 2018. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Gu:2021:PBT

[Gu21] Xiaofeng Gu. A proof of Brouwer’s toughness conjecture. SIAM Journal on Discrete Mathematics, 35(2):

948–952, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Guichard:1998:CGH

[Gui98] David R. Guichard. Competition graphs of Hamiltonian digraphs. SIAM Journal on Discrete Mathematics, 11(1): 128–134, February 1998. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Gunluk:2007:NMC

[Gün07] Oktay Günlük. A new mincut max-flow ratio for multicommodity flows. SIAM Journal on Discrete Mathematics, 21(1):1–15, ???? 2007. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Gutin:1993:FLP

[Gut93] G. Gutin. Finding a longest path in a complete multipartite digraph. SIAM Journal on Discrete Mathematics, 6 (2):270–273, May 1993. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Gargano:1992:MTB

[GV92] Luisa Gargano and Ugo Vaccaro. Minimum time broadcast networks tolerating a logarithmic number of faults. SIAM Journal on Discrete Mathematics, 5(2):

178–198, May 1992. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Gurjar:2021:NCR

[GV21] Rohit Gurjar and Nisheeth K.

Vishnoi. On the number of circuits in regular matroids (with connections to lattices and codes). SIAM Journal on Discrete Mathematics, 35(3): 1688–1705, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Gupta:2008:COC

[GvdHM^{+}08] Arvind Gupta, Jan van den Heuvel, Ján Maňuch, Ladislav Stacho, and Xiaohong Zhao. On the complexity of ordered colorings. SIAM Journal on Discrete Mathematics, 22(2): 832–847, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Gat-Viks:2006:RCF

[GVKSS06] Irit Gat-Viks, Richard M.

Karp, Ron Shamir, and Roded Sharan. Reconstructing chain functions in genetic networks. SIAM Journal on Discrete Mathematics, 20(3): 727–740, January 2006. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Gentile:2006:MCG

[GVW06] C. Gentile, P. Ventura, and R. Weismantel. Mod-2 cuts generation yields the convex hull of bounded integer feasible sets. SIAM Journal on Discrete Mathematics, 20(4): 913–919, January 2006. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Grace:2017:TBM

[GvZ17] Kevin Grace and Stefan H. M. van Zwam. Templates for binary matroids. SIAM Journal on Discrete Mathematics, 31(1):254–282, ???? 2017. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Grace:2019:HCE

[GvZ19] Kevin Grace and Stefan H. M. van Zwam. The highly connected even-cycle and even-cut matroids. SIAM Journal on Discrete Mathematics, 33(1):26–67, ???? 2019. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Goemans:1994:NAA

[GW94] Michel X. Goemans and David P. Williamson. New \frac{3}{4}-approximation algorithms for the maximum satisfiability problem. SIAM Journal on Discrete Mathematics, 7 (4):656–666, November 1994. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Gao:1999:SLC

[GW99] Zhicheng Gao and Nicholas C.

Wormald. The size of the largest components in random planar maps. SIAM Journal on Discrete Mathematics, 12(2):217–228, May 1999. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/29205.

Goemans:2000:TDG

[GW00] Michel X. Goemans and David P. Williamson. Twodimensional Gantt charts and a scheduling algorithm of Lawler. SIAM Journal on Discrete Mathematics, 13(3):281–294, 2000. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/33025.

Geelen:2007:RBC

[GW07] Jim Geelen and Kerri Webb.

On Rota’s basis conjecture. SIAM Journal on Discrete Mathematics, 21(3): 802–804, ???? 2007. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Gutekunst:2019:CIG

[GW19] Samuel C. Gutekunst and David P. Williamson. Characterizing the integrality gap of the subtour LP for the circulant traveling salesman problem. SIAM Journal on Discrete Mathematics, 33(4): 2452–2478, ???? 2019. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Glanzer:2018:NDR

[GWZ18] Christoph Glanzer, Robert Weismantel, and Rico Zenklusen. On the number of distinct rows of a matrix with bounded subdeterminants. SIAM Journal on Discrete Mathematics, 32(3): 1706–1720, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Gao:2022:EIK

[GXY22] Alice L. L. Gao, Matthew H. Y. Xie, and Arthur L. B. Yang. The equivariant inverse Kazhdan–Lusztig polynomials of uniform matroids. SIAM Journal on Discrete Mathematics, 36 (4):2553–2569, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M143995X.

Griggs:1992:LGC

[GY92] Jerrold R. Griggs and Roger K. Yeh. Labelling graphs with a condition at distance 2. SIAM Journal on Discrete Mathematics, 5(4):586–595, November 1992. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Gutin:2023:LBM

[GY23] Gregory Gutin and Anders Yeo. Lower bounds for maximum weighted cut. SIAM Journal on Discrete Mathematics, 37(2):1142–1161, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL

https://epubs.siam.org/
doi/10.1137/21M1411913.

Gyarfas:2019:TNB

[Gyá19] András Gyárfás. The Turán number of Berge K_{4} in triple systems. SIAM Journal on Discrete Mathematics, 33(1): 383–392, ???? 2019. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Guan:1998:MCV

[GZ98] D. J. Guan and Xuding Zhu.

Multiple capacity vehicle routing on paths. SIAM Journal on Discrete Mathematics, 11(4):590–602, November 1998. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Geelen:2006:STI

[GZ06] Jim Geelen and Xiangqian Zhou. A splitter theorem for internally 4-connected binary matroids. SIAM Journal on Discrete Mathematics, 20(3):

578–587, January 2006. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Gorlich:2010:PD

[GŻ10] Agnieszka Görlich and Andrzej Żak. On packable digraphs. SIAM Journal on Discrete Mathematics, 24(2): 552–557, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Gao:2019:PFF

[GZ19] Shiliang Gao and Shira Zerbib. The (2, 2) and (4, 3) properties in families of fat sets in the plane. SIAM Journal on Discrete Mathematics, 33(3):1326–1337, ???? 2019. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Han:1998:NDA

[Han98]  Yunghsiang S. Han.  A

new decoding algorithm for complete decoding of linear block codes. SIAM Journal on Discrete Mathematics, 11 (4):664–671, November 1998. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Han:2009:YAG

[Han09] Guo-Niu Han. Yet another generalization of Postnikov’s hook length formula for binary trees. SIAM Journal on Discrete Mathematics, 23(2):

661–664, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Han:2016:NPM

[Han16a] Jie Han. Near perfect matchings in k-uniform hypergraphs II. SIAM Journal on Discrete Mathematics, 30(3): 1453–1469, ???? 2016. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Han:2016:PMH

[Han16b] Jie Han. Perfect matchings in hypergraphs and the Erdős matching conjecture. SIAM Journal on Discrete Mathematics, 30(3):1351–1357, ???? 2016. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Han:2018:PMT

[Han18] Jie Han. On perfect matchings and tilings in uniform hypergraphs. SIAM Journal on Discrete Mathematics, 32(2): 919–932, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Han:2019:MAD

[Han19a] Guangyue Han. On mergings in acyclic directed graphs. SIAM Journal on Discrete Mathematics, 33(3):1482– 1502, ???? 2019. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hancock:2019:MBR

[Han19b] Robert Hancock. The maker– breaker Rado game on a random set of integers. SIAM Journal on Discrete Mathematics, 33(1):68–94, ???? 2019. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hartvigsen:2001:CRC

[Har01] David Hartvigsen. Compact representations of cuts. SIAM Journal on Discrete Mathematics, 14(1):49–66, 2001. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31233.

Hartvigsen:2010:NSG

[Har10] David Hartvigsen. Neighbor systems and the greedy algorithm. SIAM Journal on Discrete Mathematics, 24(4): 1638–1661, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v24/i4/p1638_s1.

Harangi:2011:ASE

[Har11]  Viktor Harangi.  Acute

sets in Euclidean spaces. SIAM Journal on Discrete Mathematics, 25(3):1212– 1229, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v25/i3/p1212_s1.

Hartvigsen:2018:PMC

[Har18] David Hartvigsen. Packing k-matchings and k-critical graphs. SIAM Journal on Discrete Mathematics, 32(1): 320–351, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Harris:2019:SRC

[Har19]  David G. Harris.  Some

results on chromatic number as a function of triangle count. SIAM Journal on Discrete Mathematics, 33(1): 546–563, ???? 2019. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Haastad:1994:SWT

[Hås94]  Johan Håstad.  On the

size of weights for threshold gates. SIAM Journal on Discrete Mathematics, 7(3): 484–492, August 1994. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Haviv:2019:SFS

[Hav19] Ishay Haviv. Sum-free sets of integers with a forbidden sum. SIAM Journal on Discrete Mathematics, 33(1): 402–424, ???? 2019. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hayamizu:2021:STR

[Hay21] Momoko Hayamizu. A structure theorem for rooted binary phylogenetic networks and its implications for treebased networks. SIAM Journal on Discrete Mathematics, 35(4):2490–2516, ???? 2021. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hertrich:2023:TLB

[HBDS23] Christoph Hertrich, Amitabh Basu, Marco Di Summa, and Martin Skutella. Towards lower bounds on the depth of ReLU neural networks. SIAM Journal on Discrete Mathematics, 37 (2):997–1029, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1489332.

Hwang:1998:NCI

[HC98] Shiow-Fen Hwang and Gerard J. Chang. k-neighborhood-covering and -independence problems for chordal graphs. SIAM Journal on Discrete Mathematics, 11(4):633–643, November 1998. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Heden:2008:PCL

[Hed08] Olof Heden. Perfect codes of length n with kernels of dimension n − log(n + 1) − 3. SIAM Journal on Discrete Mathematics, 22(4): 1338–1350, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hefner:1997:MMT

[Hef97] Andreas Hefner. A min-max theorem for a constrained matching problem. SIAM Journal on Discrete Mathematics, 10(2):180–189, May 1997. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hendrey:2018:SGH

[Hen18]  Kevin Hendrey.  Sparse

graphs of high gonality. SIAM Journal on Discrete Mathematics, 32(2):1400–1407, ???? 2018. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hetyei:2014:HCF

[Het14] Gábor Hetyei. Hurwitzian continued fractions containing a repeated constant and an arithmetic progression. SIAM Journal on Discrete Mathematics, 28(2): 962–985, ???? 2014. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Huang:2020:SNO

[HGY20] Fei Huang, Shanshan Guo, and Jinjiang Yuan. A short note on open-neighborhood conflict-free colorings of graphs. SIAM Journal on Discrete Mathematics, 34(3): 2009–2015, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hassin:1992:MPT

[HH92] Refael Hassin and Moshe Haviv. Mean passage times and nearly uncoupled Markov chains. SIAM Journal on Discrete Mathematics, 5(3): 386–397, August 1992. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Huang:2001:WIT

[HH01] S. H. Huang and F. K. Hwang. When is individual testing optimal for nonadaptive group testing? SIAM Journal on Discrete Mathematics, 14(4):540–548, 2001. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/35924.

Hage:2004:CAS

[HH04] Jurriaan Hage and Tero Harju. A characterization of acyclic switching classes of graphs using forbidden subgraphs. SIAM Journal on Discrete Mathematics, 18(1):159–176, 2004. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/38189.

Hell:2005:CLR

[HH05] Pavol Hell and Jing Huang.

Certifying LexBFS recognition algorithms for proper interval graphs and proper interval bigraphs. SIAM Journal on Discrete Mathematics, 18(3):554–570, 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/43025.

Horan:2013:UCW

[HH13] Victoria Horan and Glenn Hurlbert. Universal cycles for weak orders. SIAM Journal on Discrete Mathematics, 27(3):1360–1371, ???? 2013. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

He:2017:OMD

[HH17] Dayu He and Xin He. Optimal monotone drawings of trees. SIAM Journal on Discrete Mathematics, 31(3): 1867–1877, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hell:2021:SCG

[HHCHL21] Pavol Hell, César Hernández-Cruz, Jing Huang, and Jephian C.-H. Lin. Strong chordality of graphs with possible loops. SIAM Journal on Discrete Mathematics, 35 (1):362–375, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hsieh:2002:CEP

[HHH^{+}02] Sun-Yuan Hsieh, Chin-Wen Ho, Tsan-Sheng Hsu, Ming-Tat Ko, and Gen-Huey Chen. Characterization of efficiently parallel solvable problems on distancehereditary graphs. SIAM Journal on Discrete Mathematics, 15(4):488–518, 2002. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/38988.

Haynes:2002:DGA

[HHHH02] Teresa W. Haynes, Sandra M. Hedetniemi, Stephen T. Hedetniemi, and Michael A. Henning. Domination in graphs applied to electric power networks. SIAM Journal on Discrete Mathematics, 15(4):519–529, 2002. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/37583.

Haslegrave:2022:CLC

[HHK^{+}22] John Haslegrave, Jie Hu, Jaehoon Kim, Hong Liu, Bingyu Luan, and Guanghui Wang. Crux and long cycles in graphs. SIAM Journal on Discrete Mathematics, 36(4):2942–2958, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M143488X.

Hell:2024:SCG

[HHL24] Pavol Hell, Jing Huang, and Jephian C.-H. Lin. Strong cocomparability graphs and slash-free orderings of matrices. SIAM Journal on Discrete Mathematics, 38(1): 828–844, February 2024. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hell:2020:BAC

[HHLM20] Pavol Hell, Jing Huang, Jephian C.-H. Lin, and Ross M. McConnell. Bipartite analogues of comparability and cocomparability graphs. SIAM Journal on Discrete Mathematics, 34(3): 1969–1983, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hamalainen:1995:BBC

[HHLÖ95] Heikki O. Hämäläinen, Iiro S.

Honkala, Simon N. Litsyn, and Patric R. J. Östergård. Bounds for binary codes that are multiple coverings of the farthest-off points. SIAM Journal on Discrete Mathematics, 8(2):196–207, May 1995. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hell:2020:MOD

[HHMR20] Pavol Hell, Jing Huang, Ross M. McConnell, and Arash Rafiey. Min-orderable digraphs. SIAM Journal on Discrete Mathematics, 34(3): 1710–1724, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hirai:2016:SR

[HI16] Hiroshi Hirai and Yuni Iwamasa. On k-submodular relaxation. SIAM Journal on Discrete Mathematics, 30(3): 1726–1736, ???? 2016. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hirai:2023:NCT

[HI23] Hiroshi Hirai and Motoki Ikeda. Node-connectivity terminal backup, separately capacitated multiflow, and discrete convexity. SIAM Journal on Discrete Mathematics, 37(1):351–378, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1361857.

Halldorsson:1999:FSM

[HIKT99] Magnús M. Halldórsson, Kazuo Iwano, Naoki Katoh, and Takeshi Tokuyama. Finding subsets maximizing minimum structures. SIAM Journal on Discrete Mathematics, 12(3):342–359, August 1999. CODEN SJD-MEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/30979.

Hickingbotham:2024:TCG

[HIMW24] Robert Hickingbotham, Freddie Illingworth, Bojan Mohar, and David R. Wood. Treewidth, circle graphs, and circular drawings. SIAM Journal on Discrete Mathematics, 38(1):965–987, March 2024. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hirai:2011:FCM

[Hir11] Hiroshi Hirai. Folder complexes and multiflow combinatorial dualities. SIAM Journal on Discrete Mathematics, 25(3):1119–1143, ???? 2011. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v25/i3/p1119_s1.

Hsu:1994:EPC

[HJ94] D. Frank Hsu and Xing-De Jia. Extremal problems in the construction of distributed loop networks. SIAM Journal on Discrete Mathematics, 7(1):57–71, February 1994. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Han:2018:IBM

[HJ18] Hiˆ e.p Hàn and Andrea Jiménez. Improved bound on the maximum number of clique-free colorings with two and three colors. SIAM Journal on Discrete Mathematics, 32(2):1364–1368, ???? 2018. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Huynh:2024:EPP

[HjK24] Tony Huynh and O joung Kwon. On the Erdős–Pósa property for long holes in C_{4}- free graphs. SIAM Journal on Discrete Mathematics, 38 (1):19–42, January 2024. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Horton:1993:MED

[HK93] J. D. Horton and K. Kilakos. Minimum edge dominating sets. SIAM Journal on Discrete Mathematics, 6(3): 375–387, August 1993. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hoffmann:1996:GCR

[HK96] Frank Hoffmann and Klaus Kriegel. A graph-coloring result and its consequences for polygon-guarding problems. SIAM Journal on Discrete Mathematics, 9(2): 210–224, May 1996. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hager:1999:GPC

[HK99] William W. Hager and Yaroslav Krylyuk. Graph partitioning and continuous quadratic programming. SIAM Journal on Discrete Mathematics, 12(4):500–523, November 1999. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/33582.

Honkala:2002:MBR

[HK02] Iiro Honkala and Andrew Klapper. Multicovering bounds from relative covering radii. SIAM Journal on Discrete Mathematics, 15(2):228–234, 2002. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/37935.

Hsu:2005:OAB

[HK05] Tsan-Sheng Hsu and Ming-Yang Kao. Optimal augmentation for bipartite componentwise biconnectivity in linear time. SIAM Journal on Discrete Mathematics, 19(2):345–362, 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/30321.

Hamel:2011:BPS

[HK11] A. M. Hamel and R. C. King. Bijective proof of a symplectic dual pair identity. SIAM Journal on Discrete Mathematics, 25(2): 539–560, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v25/i2/p539_s1.

Huang:2013:NPM

[HK13] Chien-Chung Huang and Telikepalli Kavitha. Nearpopular matchings in the roommates problem. SIAM Journal on Discrete Mathematics, 27(1):43–62, ???? 2013. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Huber:2014:OTS

[HK14] Anna Huber and Andrei Krokhin. Oracle tractability of skew bisubmodular functions. SIAM Journal on Discrete Mathematics, 28(4): 1828–1837, ???? 2014. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Huber:2015:DMT

[HK15] Katharina T. Huber and George Kettleborough. Distinguished minimal topological lassos. SIAM Journal on Discrete Mathematics, 29(2): 940–961, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hebdige:2016:TCC

[HK16a] Michael Hebdige and Daniel Král’. Third case of the cyclic coloring conjecture. SIAM Journal on Discrete Mathematics, 30(1):525–548, ???? 2016. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Henning:2016:DGP

[HK16b] Michael A. Henning and William B. Kinnersley. Domination game: a proof of the 3/5-conjecture for graphs with minimum degree at least two. SIAM Journal on Discrete Mathematics, 30(1): 20–35, ???? 2016. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hromkovic:2009:SPN

[HKK^{+}09] Juraj Hromkovič, Przemysława Kanarek, Ralf Klasing, Krzysztof Lory, Walter Unger, and Hubert Wagener. On the size of permutation networks and consequences for efficient simulation of hypercube algorithms on bounded-degree networks. SIAM Journal on Discrete Mathematics, 23(3): 1612–1645, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Horsch:2024:RBM

[HKK24] Florian Hörsch, Tomá Kaiser, and Matthias Kriesell. Rainbow bases in matroids. SIAM Journal on Discrete Mathematics, 38(2):1472–1491, May 2024. CODEN SJD-MEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hell:1988:RTF

[HKKK88] Pavol Hell, David Kirkpatrick, Jan Kratochvíl, and Igor Kříž. On restricted two-factors. SIAM Journal on Discrete Mathematics, 1 (4):472–484, November 1988. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

He:1999:LTS

[HKL99] Xin He, Ming-Yang Kao, and Hsueh-I Lu. Lineartime succinct encodings of planar graphs via canonical orderings. SIAM Journal on Discrete Mathematics, 12(3):317–325, August 1999. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/32503.

Harvey:2011:DCB

[HKL11] Nicholas J. A. Harvey, Tamás Király, and Lap Chi Lau. On disjoint common bases in two matroids. SIAM Journal on Discrete Mathematics, 25(4): 1792–1803, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v25/i4/p1792_s1.

Halman:2014:FPT

[HKL^{+}14] Nir Halman, Diego Klabjan, Chung-Lun Li, James Orlin, and David Simchi-Levi. Fully polynomial time approximation schemes for stochastic dynamic programs. SIAM Journal on Discrete Mathematics, 28(4):1725–1796, ???? 2014. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hoppen:2021:QCE

[HKL^{+}21] Carlos Hoppen, Yoshiharu Kohayakawa, Richard Lang, Hanno Lefmann, and Henrique Stagni. On the query complexity of estimating the distance to hereditary graph properties. SIAM Journal on Discrete Mathematics, 35(2): 1238–1251, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hosseini:2020:CRG

[HKM20] Seyyed Aliasghar Hosseini, Fiachra Knox, and Bojan Mohar. Cops and robbers on graphs of bounded diameter. SIAM Journal on Discrete Mathematics, 34(2): 1375–1384, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Huang:2022:AMS

[HKMY22] Chien-Chung Huang, Naonori Kakimura, Simon Mauras, and Yuichi Yoshida. Approximability of monotone submodular function maximization under cardinality and matroid constraints in the streaming model. SIAM Journal on Discrete Mathematics, 36(1):355–382, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL

                                        https://epubs.siam.org/
                                        doi/10.1137/20M1357317.

Hanckowiak:2001:DCC

[HKP01] Michal Hanckowiak, Michal Karonski, and Alessandro Panconesi. On the distributed complexity of computing maximal matchings. SIAM Journal on Discrete Mathematics, 15(1):41–57, 2001. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/37312.

Hladky:2017:ALKa

[HKP^{+}17a] Jan Hladký, János Komlós, Diana Piguet, Miklós Simonovits, Maya Stein, and Endre Szemerédi. The approximate Loebl–Komlós–Sós conjecture I: The sparse decomposition. SIAM Journal on Discrete Mathematics, 31 (2):945–982, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hladky:2017:ALKb

[HKP^{+}17b] Jan Hladký, János Komlós, Diana Piguet, Miklós Simonovits, Maya Stein, and Endre Szemerédi. The approximate Loebl–Komlós–Sós conjecture II: The rough structure of LKS graphs. SIAM Journal on Discrete Mathematics, 31(2): 983–1016, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hladky:2017:ALKc

[HKP^{+}17c] Jan Hladký, János Komlós, Diana Piguet, Miklós Simonovits, Maya Stein, and Endre Szemerédi. The approximate Loebl–Komlós–Sós conjecture III: The finer structure of LKS graphs. SIAM Journal on Discrete Mathematics, 31(2):1017– 1071, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hladky:2017:ALKd

[HKP^{+}17d] Jan Hladký, János Komlós, Diana Piguet, Miklós Simonovits, Maya Stein, and Endre Szemerédi. The approximate Loebl–Komlós–Sós conjecture IV: Embedding techniques and the proof of the main result. SIAM Journal on Discrete Mathematics, 31(2):1072–1148, ???? 2017. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hols:2022:EDB

[HKP22] Eva-Maria C. Hols, Stefan Kratsch, and Astrid Pieterse. Elimination distances, blocking sets, and kernels for vertex cover. SIAM Journal on Discrete Mathematics, 36 (3):1955–1990, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/20M1335285.

Hammer:2000:ESR

[HKR00] Peter L. Hammer, Alexander Kogan, and Uriel G. Rothblum. Evaluation, strength, and relevance of variables of Boolean functions. SIAM Journal on Discrete Mathematics, 13(3):302–312, 2000. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/32787.

Hurkens:2007:VPN

[HKS07] C. A. J. Hurkens, J. C. M.

Keijsper, and L. Stougie. Virtual private network design: a proof of the tree routing conjecture on ring networks. SIAM Journal on Discrete Mathematics, 21(2): 482–503, ???? 2007. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hefetz:2008:PCM

[HKSS08] Dan Hefetz, Michael Krivelevich, Miloš Stojaković, and Tibor Szabó. Planarity, colorability, and minor games. SIAM Journal on Discrete Mathematics, 22 (1):194–212, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hunsaker:2003:OOA

[HKST03] Brady Hunsaker, Anton J.

Kleywegt, Martin W. P. Savelsbergh, and Craig A. Tovey. Optimal online algorithms for minimax resource scheduling. SIAM Journal on Discrete Mathematics, 16(4):555–590, 2003. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/39776.

Haase:2020:LOP

[HKT20] Christian Haase, Florian Kohl, and Akiyoshi Tsuchiya.

Levelness of order polytopes. SIAM Journal on Discrete Mathematics, 34(2): 1261–1280, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Heuberger:2015:CTC

[HKW15] Clemens Heuberger, Daniel Krenn, and Stephan Wagner. Canonical trees, compact prefix-free codes, and sums of unit fractions: a probabilistic analysis. SIAM Journal on Discrete Mathematics, 29(3): 1600–1653, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Henk:2022:LWL

[HKW22] Martin Henk, Stefan Kuhlmann, and Robert Weismantel. On lattice width of lattice-free polyhedra and height of Hilbert bases. SIAM Journal on Discrete Mathematics, 36(3):1918–1942, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M145416X.

Hong:1992:SRL

[HL92] Kwang S. Hong and Joseph Y.-T. Leung. Some results on Liu’s conjecture. SIAM Journal on Discrete Mathematics, 5(4):500–523, November 1992. CODEN SJDMEC.

ISSN 0895-4801 (print), 1095-7146 (electronic).

Hoang:2000:RPS

[HL00] Chính T. Hoàng and Van Bang Le. Recognizing perfect 2split graphs. SIAM Journal on Discrete Mathematics, 13(1):48–55, 2000. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/32908.

Hsieh:2010:PRH

[HL10] Sun-Yuan Hsieh and Chia-Wei Lee. Pancyclicity of restricted hypercube-like networks under the conditional fault model. SIAM Journal on Discrete Mathematics, 23 (4):2100–2119, ???? 2010. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ha:2015:NP

[HL15] Huy Tài Hà and Kuei-Nuan Lin. Normal 0–1 polytopes. SIAM Journal on Discrete Mathematics, 29(1): 210–223, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hartarsky:2024:MRT

[HL24] Ivailo Hartarsky and Lyuben Lichev. The maximal running time of hypergraph bootstrap percolation. SIAM Journal on Discrete Mathematics, 38(2):1462–1471, May 2024. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hlineny:2018:SSR

[Hli18] Petr Hlinený. A simpler self-reduction algorithm for matroid path-width. SIAM Journal on Discrete Mathematics, 32(2):1425–1440, ???? 2018. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Han:2021:IFM

[HLL^{+}21] Miaomiao Han, Jiaao Li, Rong Luo, Yongtang Shi, and Cun-Quan Zhang. Integer flows and modulo orientations of signed graphs. SIAM Journal on Discrete Mathematics, 35(1):376–391, ???? 2021. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Heggernes:2011:CSG

[HLMP11] Pinar Heggernes, Daniel Lokshtanov, Rodica Mihai, and Charis Papadopoulos. Cutwidth of split graphs and threshold graphs. SIAM Journal on Discrete Mathematics, 25(3):1418–1437, ???? 2011. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v25/i3/p1418_s1.

Hoppen:2017:RER

[HLO17] Carlos Hoppen, Hanno Lefmann, and Knut Odermann. A rainbow Erdős–Rothschild problem. SIAM Journal on Discrete Mathematics, 31(4): 2647–2674, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Helmuth:2023:AAR

[HLP^{+}23] Tyler Helmuth, Holden Lee, Will Perkins, Mohan Ravichandran, and Qiang Wu. Approximation algorithms for the random field Ising model. SIAM Journal on Discrete Mathematics, 37(3):1610– 1629, ???? 2023. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:/

/epubs.siam.org/doi/10.
1137/21M1467389.

Heath:1992:CQS

[HLR92] Lenwood S. Heath, Frank Thomson Leighton, and Arnold L. Rosenberg. Comparing queues and stacks as mechanisms for laying out graphs. SIAM Journal on Discrete Mathematics, 5(3):398–412, August 1992. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Horn:2013:JNM

[HLR13] Paul Horn, Steve La Fleur, and Vojtech Rödl. Jumps and nonjumps in multigraphs. SIAM Journal on Discrete Mathematics, 27(2): 1040–1054, ???? 2013. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hajebi:2022:CDL

[HLS22] Sepehr Hajebi, Yanjia Li, and Sophie Spirkl. Complexity dichotomy for list-5-coloring with a forbidden induced subgraph. SIAM Journal on Discrete Mathematics, 36(3):2004–2027, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1443352.

Hajebi:2024:LCO

[HLS24] Sepehr Hajebi, Yanjia Li, and Sophie Spirkl. List-3-coloring ordered graphs with a forbidden induced subgraph. SIAM Journal on Discrete Mathematics, 38(1): 1158–1190, March 2024. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hamidoune:2000:ICV

[HLST00] Y. O. Hamidoune, A. S. Lladó, O. Serra, and R. Tindell. On isoperimetric connectivity in vertex-transitive graphs. SIAM Journal on Discrete Mathematics, 13 (1):139–144, 2000. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/30635.

Hyde:2019:DSK

[HLT19] Joseph Hyde, Hong Liu, and Andrew Treglown. A degree sequence Komlós theorem. SIAM Journal on Discrete Mathematics, 33(4): 2041–2061, ???? 2019. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hu:2023:GTI

[HLWY23] Jie Hu, Hao Li, Yue Wang, and Donglei Yang. Graph tilings in incompatibility systems. SIAM Journal on Discrete Mathematics, 37 (3):1930–1951, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1506353.

Henning:2023:TVT

[HLY23] Michael A. Henning, Christian Löwenstein, and Anders Yeo. The Tuza–Vestergaard theorem. SIAM Journal on Discrete Mathematics, 37 (2):1275–1310, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1475508.

Hu:2013:OBN

[HLZ13] Zhiquan Hu, Ka Ho Law, and Wenan Zang. An optimal binding number condition for bipancyclism. SIAM Journal on Discrete Mathematics, 27 (2):597–618, ???? 2013. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Haimovich:1988:EPH

[HM88] M. Haimovich and T. L. Magnanti. Extremum properties of hexagonal partitioning and the uniform distribution in Euclidean location. SIAM Journal on Discrete Mathematics, 1(1):50–64, February 1988. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hartvigsen:1994:APM

[HM94] David Hartvigsen and Russell Mardon. The all-pairs min cut problem and the minimum cycle basis problem on planar graphs. SIAM Journal on Discrete Mathematics, 7(3):403–418, August 1994. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Harutyunyan:2011:GTL

[HM11] Ararat Harutyunyan and Bojan Mohar. Gallai’s theorem for list coloring of digraphs. SIAM Journal on Discrete Mathematics, 25(1): 170–180, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v25/i1/p170_s1.

Harutyunyan:2012:PGE

[HM12a] Ararat Harutyunyan and Bojan Mohar. Planar graphs have exponentially many 3arboricities. SIAM Journal on Discrete Mathematics, 26 (3):1269–1280, ???? 2012. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Humpert:2012:IHA

[HM12b] Brandon Humpert and Jeremy L.

Martin. The incidence Hopf algebra of graphs. SIAM Journal on Discrete Mathematics, 26(2):555–570, ???? 2012. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Huang:2019:TCH

[HM19] Hao Huang and Jie Ma.

On tight cycles in hypergraphs. SIAM Journal on Discrete Mathematics, 33(1): 230–237, ???? 2019. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hartarsky:2020:CTD

[HM20] Ivailo Hartarsky and Tamás Róbert Mezei. Complexity of twodimensional bootstrap percolation difficulty: Algorithm and NP-Hardness. SIAM Journal on Discrete Mathematics, 34(2):1444–1459, ???? 2020. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hatami:2009:IGS

[HMM09] Hamed Hatami, Avner Magen, and Evangelos Markakis. Integrality gaps of semidefinite programs for vertex cover and relations to ℓ_{1} embeddability of negative type metrics. SIAM Journal on Discrete Mathematics, 23(1): 178–194, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Huang:2021:AAF

[HMM^{+}21] Chien-Chung Huang, Mathieu Mari, Claire Mathieu, Kevin Schewior, and Jens Vygen. An approximation algorithm for fully planar edgedisjoint paths. SIAM Journal on Discrete Mathematics, 35 (2):752–769, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Holzman:1997:LBQ

[HMP97] Ron Holzman, Yosi Marcus, and David Peleg. Load balancing in quorum systems. SIAM Journal on Discrete Mathematics, 10(2): 223–245, May 1997. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Huygens:2004:TED

[HMP04] David Huygens, Ali Ridha Mahjoub, and Pierre Pesneau. Two edge-disjoint hop-constrained paths and polyhedra. SIAM Journal on Discrete Mathematics, 18(2):287–312, 2004. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/41944.

Helman:1993:ECG

[HMS93] Paul Helman, Bernard M. E.

Moret, and Henry D. Shapiro. An exact characterization of greedy structures. SIAM Journal on Discrete Mathematics, 6(2):274–283, May 1993. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Huber:2005:FCS

[HMS05] Katharina T. Huber, Vincent Moulton, and Mike Steel. Four characters suffice to convexly define a phylogenetic tree. SIAM Journal on Discrete Mathematics, 18(4):835–843, 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/41669.

Hommelsheim:2021:HSM

[HMS21] Felix Hommelsheim, Moritz Muehlenthaler, and Oliver Schaudt. How to secure matchings against edge failures. SIAM Journal on Discrete Mathematics, 35(3): 2265–2292, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Huber:2014:RPT

[HMSW14] K. T. Huber, V. Moulton, C. Semple, and T. Wu. Representing partitions on trees. SIAM Journal on Discrete Mathematics, 28(3): 1152–1172, ???? 2014. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Habib:2024:BAB

[HMSZ24] Michel Habib, Lalla Mouatadid, Éric Sopena, and Mengchuan Zou. (α, β)- modules in graphs. SIAM Journal on Discrete Mathematics, 38(1):566–589, January 2024. CODEN SJD-MEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Han:2021:EAS

[HMZ21] Bin Han, Jianxi Mao, and Jiang Zeng. Equidistributions around special kinds of descents and excedances. SIAM Journal on Discrete Mathematics, 35(4):2858–2879, ???? 2021. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hartung:2015:NHF

[HN15] Sepp Hartung and André Nichterlein. NP-hardness and fixed-parameter tractability of realizing degree sequences with directed acyclic graphs. SIAM Journal on Discrete Mathematics, 29(4):1931– 1960, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Huynh:2020:MSP

[HN20] Tony Huynh and Peter Nelson. The matroid secretary problem for minorclosed classes and random matroids. SIAM Journal on Discrete Mathematics, 34(1): 163–176, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ha:2023:EMF

[HN23] Dung M. Ha and Hieu T.

Ngo. Expanders on matrices over a finite chain ring, II. SIAM Journal on Discrete Mathematics, 37 (3):1587–1609, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1514362.

Hng:2022:MDP

[Hng22] Eng Keat Hng. Minimum degrees for powers of paths and cycles. SIAM Journal on Discrete Mathematics, 36 (4):2667–2736, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1359183.

Hoang:2010:CFS

[Hoà10] Chính T. Hoàng. On the complexity of finding a sun in a graph. SIAM Journal on Discrete Mathematics, 23(4): 2156–2162, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Huang:2022:SMP

[HOD22] Yuanqiu Huang, Zhangdong Ouyang, and Fengming Dong. On the size of matchings in 1planar graph with high minimum degree. SIAM Journal on Discrete Mathematics, 36(4):2570–2584, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1459952.

Hofmeister:1995:ECR

[Hof95] M. Hofmeister. Enumeration of concrete regular covering projections. SIAM Journal on Discrete Mathematics, 8(1):51–61, February 1995. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hofmeister:1998:NCC

[Hof98] Michael Hofmeister. A note on counting connected graph covering projections. SIAM Journal on Discrete Mathematics, 11(2):286–292, May 1998. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Horsley:2014:EPS

[Hor14] Daniel Horsley. Embedding partial Steiner triple systems with few triples. SIAM Journal on Discrete Mathematics, 28(3):1199–1213, ???? 2014. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Horn:2019:SGE

[Hor19] Paul Horn. A spacial gradient estimate for solutions to the heat equation on graphs. SIAM Journal on Discrete Mathematics, 33(2): 958–975, ???? 2019. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Heath:1997:SQL

[HP97] Lenwood S. Heath and Sriram V. Pemmaraju. Stack and queue layouts of posets. SIAM Journal on Discrete Mathematics, 10(4):599–625, November 1997. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Huang:2021:LBS

[HP21] Shang-En Huang and Seth Pettie. Lower bounds on sparse spanners, emulators, and diameter-reducing shortcuts. SIAM Journal on Discrete Mathematics, 35(3): 2129–2144, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hu:2023:FCN

[HP23] Xiaolan Hu and Xing Peng. The fractional chromatic number of K_{Δ}-free graphs. SIAM Journal on Discrete Mathematics, 37(4):2486– 2507, October 2023. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Har-Peled:2013:PG

[HPL13] Sariel Har-Peled and Bernard Lidický. Peeling the grid. SIAM Journal on Discrete Mathematics, 27(2):650–655, ???? 2013. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Heydemann:1996:SHD

[HPS96] Marie-Claude Heydemann, Joseph G. Peters, and Dominique Sotteau. Spanners of hypercube-derived networks. SIAM Journal on Discrete Mathematics, 9(1): 37–54, February 1996. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Han:2009:PMU

[HPS09] Hip Hàn, Yury Person, and Mathias Schacht. On perfect matchings in uniform hypergraphs with large minimum vertex degree. SIAM Journal on Discrete Mathematics, 23 (2):732–748, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hoppen:2019:PAR

[HPS19] Carlos Hoppen, Roberto F.

Parente, and Cristiane M. Sato. Packing arborescences in random digraphs. SIAM Journal on Discrete Mathematics, 33(1):438–453, ???? 2019. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hamburger:2009:KRG

[HPW09] Peter Hamburger, Attila Por, and Matt Walsh. Kneser representations of graphs. SIAM Journal on Discrete Mathematics, 23(2):1071–1081, ???? 2009. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hochbaum:2003:MCC

[HQ03] Dorit S. Hochbaum and Maurice Queyranne. Minimizing a convex cost closure set. SIAM Journal on Discrete Mathematics, 16(2):192–207, 2003. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/36958.

Heuer:1988:SGD

[HR88] G. A. Heuer and W. Dow Rieder. Silverman games on disjoint discrete sets. SIAM Journal on Discrete Mathematics, 1(4):485–525, November 1988. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hassin:2002:RM

[HR02] Refael Hassin and Shlomi Rubinstein. Robust matchings. SIAM Journal on Discrete Mathematics, 15 (4):530–537, 2002. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/33215.

Hoang:2005:CSP

[HR05a] Chính T. Hoàng and Bruce Reed. On the co-P_{3}-structure of perfect graphs. SIAM Journal on Discrete Mathematics, 18(3):571–576, 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/37079.

Hwang:2005:POS

[HR05b] Frank K. Hwang and Uriel G. Rothblum. Partition-optimization with Schur convex sum objective functions. SIAM Journal on Discrete Mathematics, 18(3):512–524, 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/34716.

Hell:2012:DMC

[HR12a] Pavol Hell and Arash Rafiey.

The dichotomy of minimum cost homomorphism problems for digraphs. SIAM Journal on Discrete Mathematics, 26(4):1597–1608, ???? 2012. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hell:2012:MPI

[HR12b] Pavol Hell and Arash Rafiey.

Monotone proper interval digraphs and min-max orderings. SIAM Journal on Discrete Mathematics, 26(4): 1576–1596, ???? 2012. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hwang:1993:MOM

[HRS93] Frank K. Hwang, Uriel G.

Rothblum, and Larry Shepp. Monotone optimal multipartitions using Schur convexity with respect to partial orders. SIAM Journal on Discrete Mathematics, 6(4):533– 547, November 1993. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Havet:2012:GYC

[HRS12] Frédéric Havet, Bruce Reed, and Jean-Sébastien Sereni. Griggs and Yeh’s Conjecture and L(p, 1)-labelings. SIAM Journal on Discrete Mathematics, 26(1):145–168, ???? 2012. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Helmberg:2017:CAM

[HRS17] Christoph Helmberg, Israel Rocha, and Uwe Schwerdtfeger. Combinatorial algorithms for minimizing the maximum Laplacian and signless Laplacian eigenvalues of weighted graphs. SIAM Journal on Discrete Mathematics, 31(2):1196–1216, ???? 2017. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Han:2018:VFN

[HRS18] Hiˆ e.p Hàn, Vojtĕch Rödl, and Tibor Szabó. Vertex Folkman numbers and the minimum degree of minimal Ramsey graphs. SIAM Journal on Discrete Mathematics, 32(2):826–838, ???? 2018. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Haddad:1988:RBF

[HS88] Ramsey W. Haddad and Alejandro A. Schäffer. Recognizing Bellman–Ford-orderable graphs. SIAM Journal on Discrete Mathematics, 1(4): 447–471, November 1988. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hakimi:1989:NVA

[HS89a] S. L. Hakimi and E. F.

Schmeichel. A note on the vertex arboricity of a graph. SIAM Journal on Discrete Mathematics, 2(1): 64–67, February 1989. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hurkens:1989:SSS

[HS89b] C. A. J. Hurkens and A. Schrijver. On the size of systems of sets every t of which have an SDR, with an application to the worst-case ratio of heuristics for packing problems. SIAM Journal on Discrete Mathematics, 2(1): 68–72, February 1989. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hajnal:1990:BCP

[HS90] Péter Hajnal and Endre Szemerédi. Brooks coloring in parallel. SIAM Journal on Discrete Mathematics, 3(1): 74–80, February 1990. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hamel:1997:LLC

[HS97] A. M. Hamel and M. A. Steel. The length of a leaf coloration on a random binary tree. SIAM Journal on Discrete Mathematics, 10(3): 359–372, August 1997. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hitczenko:2004:MPR

[HS04] Pawel Hitczenko and Carla D. Savage. On the multiplicity of parts in a random composition of a large integer. SIAM Journal on Discrete Mathematics, 18(2):418–435, 2004. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/36315.

Hoory:2006:NUC

[HS06] Shlomo Hoory and Stefan Szeider. A note on unsatisfiable k-CNF formulas with few occurrences per variable. SIAM Journal on Discrete Mathematics, 20(2): 523–528, January 2006. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hladky:2010:NBG

[HS10] Jan Hladký and Mathias Schacht. Note on bipartite graph tilings. SIAM Journal on Discrete Mathematics, 24 (2):357–362, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Heuer:2022:PPA

[HS22] Dylan Heuer and Jessica Striker. Partial permutation and alternating sign matrix polytopes. SIAM Journal on Discrete Mathematics, 36 (4):2863–2888, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1417958.

Hammer:1988:LSU

[HSLd88] P. L. Hammer, B. Simeone, Th. M. Liebling, and D. de Werra. From linear separability to unimodality: a hierarchy of pseudo-Boolean functions. SIAM Journal on Discrete Mathematics, 1(2): 174–184, May 1988. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Havet:2008:FCP

[HSŠ08] Frédéric Havet, Jean-Sébastien Sereni, and Riste Škrekovski. 3-facial coloring of plane graphs. SIAM Journal on Discrete Mathematics, 22(1): 231–247, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Harks:2019:CUG

[HSS19] Tobias Harks, Anja Schedel, and Manuel Surek. A characterization of undirected graphs admitting optimal cost shares. SIAM Journal on Discrete Mathematics, 33(4): 1932–1996, ???? 2019. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hoefer:2024:SPI

[HSS24] Martin Hoefer, Kevin Schewior, and Daniel Schmand. Stochastic probing with increasing precision. SIAM Journal on Discrete Mathematics, 38(1): 148–169, January 2024. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hancock:2019:ISH

[HST19] Robert Hancock, Katherine Staden, and Andrew Treglown. Independent sets in hypergraphs and Ramsey properties of graphs and the integers. SIAM Journal on Discrete Mathematics, 33(1): 153–188, ???? 2019. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Harrison:2022:LSP

[HST22] Anthony Harrison, Jenya Soprunova, and Patrick Tierney. Lattice size of plane convex bodies. SIAM Journal on Discrete Mathematics, 36(1):92–102, ???? 2022.

CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M137536X.

Harris:2023:LNW

[HSV23] David G. Harris, Hsin-Hao Su, and Hoa T. Vu. On the locality of Nash–Williams forest decomposition and starforest decomposition. SIAM Journal on Discrete Mathematics, 37(2):800–830, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL

https://epubs.siam.org/
doi/10.1137/21M1434441.

Hahnle:2013:SRU

[HSZ13] Nicolai Hähnle, Laura Sanità, and Rico Zenklusen. Stable routing and unique-max coloring on trees. SIAM Journal on Discrete Mathematics, 27 (1):109–125, ???? 2013. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hamidoune:1990:VTS

[HT90] Yahya Ould Hamidoune and Ralph Tindell. Vertex transitivity and super line connectedness. SIAM Journal on Discrete Mathematics, 3 (4):524–530, November 1990. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hujter:1993:NMI

[HT93] Mihály Hujter and Zsolt Tuza. The number of maximal independent sets in triangle-free graphs. SIAM Journal on Discrete Mathematics, 6(2):284–288, May 1993. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hajdu:2013:BAD

[HT13] Lajos Hajdu and Rob Tijdeman. Bounds for approximate discrete tomography solutions. SIAM Journal on Discrete Mathematics, 27(2): 1055–1066, ???? 2013. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hlineny:2018:DPG

[HT18a] Petr Hlinený and Carsten Thomassen. Deciding parity of graph crossing number. SIAM Journal on Discrete Mathematics, 32(3): 1962–1965, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hoyer:2018:FEI

[HT18b] Alexander Hoyer and Robin Thomas. Four edge-independent spanning trees. SIAM Journal on Discrete Mathematics, 32(1):233–248, ???? 2018. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

He:2019:HFB

[HT19] Zhiyang He and Michael Tait.

Hypergraphs with few Berge paths of fixed length between vertices. SIAM Journal on Discrete Mathematics, 33(3): 1472–1481, ???? 2019. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hu:2018:BSC

[HTS18] Sihuang Hu, Itzhak Tamo, and Ofer Shayevitz. A bound on the Shannon capacity via a linear programming variation. SIAM Journal on Discrete Mathematics, 32(3): 2229–2241, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Heggernes:2005:CMT

[HTV05] Pinar Heggernes, Jan Arne Telle, and Yngve Villanger. Computing minimal triangulations in time O(n^{α} log n) = o(n^{2.376}). SIAM Journal on Discrete Mathematics, 19 (4):900–913, 2005. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/44501.

Huang:2014:MID

[Hua14] Hao Huang. On the maximum induced density of directed stars and related problems. SIAM Journal on Discrete Mathematics, 28(1): 92–98, ???? 2014. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Huicochea:2023:KTT

[Hui23] Mario Huicochea. A Knesertype theorem for restricted sumsets. SIAM Journal on Discrete Mathematics, 37(1):83–93, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1476150.

Hurlbert:1994:UCS

[Hur94] Glenn Hurlbert. On universal cycles for k-subsets of an n-set. SIAM Journal on Discrete Mathematics, 7(4):598– 604, November 1994. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hutchinson:1988:SNC

[Hut88] Joan P. Hutchinson. On short noncontractible cycles in embedded graphs. SIAM Journal on Discrete Mathematics, 1(2):185–192, May 1988. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hoogeveen:2000:BPD

[HV00] J. A. Hoogeveen and A. P. A. Vestjens. A best possible deterministic on-line algorithm for minimizing maximum delivery time on a single machine. SIAM Journal on Discrete Mathematics, 13(1):56–63, 2000. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/29682.

Heeger:2017:TCS

[HV17] Klaus Heeger and Jens Vygen. Two-connected spanning subgraphs with at most \frac{10}{7}OP T edges. SIAM Journal on Discrete Mathematics, 31(3):1820–1835, ???? 2017. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hurkens:2007:PRB

[HvIK^{+}07] Cor Hurkens, Leo van Iersel, Judith Keijsper, Steven Kelk, Leen Stougie, and John Tromp. Prefix reversals on binary and ternary strings. SIAM Journal on Discrete Mathematics, 21(3):592–611, ???? 2007. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Heggernes:2013:OBG

[HVLP13] Pinar Heggernes, Pim Van ’T Hof, Daniel Lokshtanov, and Christophe Paul. Obtaining a bipartite graph by contracting few edges. SIAM Journal on Discrete Mathematics, 27(4):2143–2156, ???? 2013. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Heggernes:2012:CCB

[HvtHLN12] Pinar Heggernes, Pim van ’t Hof, Daniel Lokshtanov, and Jesper Nederlof. Computing the cutwidth of bipartite permutation graphs in linear time. SIAM Journal on Discrete Mathematics, 26(3): 1008–1021, ???? 2012. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hartke:2007:FRB

[HVW07] Stephen G. Hartke, Jennifer Vandenbussche, and Paul Wenger. Further results on bar k-visibility graphs. SIAM Journal on Discrete Mathematics, 21(2):523–531, ???? 2007. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Huynh:2014:ICR

[HvZ14] Tony Huynh and Stefan H. M. van Zwam. Intertwining connectivities in representable matroids. SIAM Journal on Discrete Mathematics, 28(1):188–196, ???? 2014. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Haglin:1996:CST

[HW96] David J. Haglin and Marty J. Wolf. On convex subsets in tournaments. SIAM Journal on Discrete Mathematics, 9(1):63–70, February 1996. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Haxell:2010:SPP

[HW10] P. E. Haxell and G. T. Wilfong. On the stable paths problem. SIAM Journal on Discrete Mathematics, 24(3): 1137–1152, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Harvey:2015:CGS

[HW15] Daniel J. Harvey and David R.

Wood. Cycles of given size in a dense graph. SIAM Journal on Discrete Mathematics, 29(4):2336–2349, ???? 2015. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hamann:2016:BCT

[HW16] Matthias Hamann and Daniel Weißauer. Bounding connected tree-width. SIAM Journal on Discrete Mathematics, 30(3):1391–1400, ???? 2016. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hoefer:2017:LSM

[HW17a] Martin Hoefer and Lisa Wagner. Locally stable marriage with strict preferences. SIAM Journal on Discrete Mathematics, 31(1):283–316, ???? 2017. CODEN SJDMEC.

ISSN 0895-4801 (print), 1095-7146 (electronic).

Huang:2017:NNN

[HW17b] Longxiu Huang and Tao Wang. On the number of neighbors in normal tiling. SIAM Journal on Discrete Mathematics, 31(1): 240–253, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Huang:2018:CCP

[HW18] Harry H. Y. Huang and Larry X. W. Wang. The corners of core partitions. SIAM Journal on Discrete Mathematics, 32(3):1887–1902, ???? 2018. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Huang:2023:FAM

[HW23] Chien-Chung Huang and Justin Ward. FPT-algorithms for the ℓ-matchoid problem with a coverage objective. SIAM Journal on Discrete Mathematics, 37 (2):1053–1078, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1442267.

Hickingbotham:2024:SMG

[HW24] Robert Hickingbotham and David R. Wood. Shallow minors, graph products, and beyond-planar graphs.

SIAM Journal on Discrete Mathematics, 38(1):1057– 1089, March 2024. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hu:2018:VDB

[HWS18] Sihuang Hu, Nir Weinberger, and Ofer Shayevitz. On the VC-dimension of binary codes. SIAM Journal on Discrete Mathematics, 32(3): 2161–2171, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hou:2018:BSC

[HWZZ18] Jianfeng Hou, Shufei Wu, Qinghou Zeng, and Wenxing Zhu. The Bollobás–Scott conjecture for 4-uniform hypergraphs. SIAM Journal on Discrete Mathematics, 32 (1):505–521, ???? 2018. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). See erratum [HWZZ20].

Hou:2020:EBS

[HWZZ20] Jianfeng Hou, Shufei Wu, Qinghou Zeng, and Wenxing Zhu. Erratum: The Bollobás–Scott Conjecture for 4-Uniform Hypergraphs. SIAM Journal on Discrete Mathematics, 34(4):2583– 2584, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). See [HWZZ18].

Huang:2018:PCC

[HXZ18] He Huang, Binzhou Xia, and Sanming Zhou. Perfect codes in Cayley graphs. SIAM Journal on Discrete Mathematics, 32(1):548–559, ???? 2018. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hwang:1989:CPM

[HY89] F. K. Hwang and Y.-C. Yao. Cutoff point and monotonicity properties for multinomial group testing. SIAM Journal on Discrete Mathematics, 2 (4):500–507, November 1989. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Henning:2010:STH

[HY10] Michael A. Henning and Anders Yeo. Strong transversals in hypergraphs and double total domination in graphs. SIAM Journal on Discrete Mathematics, 24(4): 1336–1355, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Henning:2012:VDC

[HY12] Michael A. Henning and Anders Yeo. Vertex disjoint cycles of different length in digraphs. SIAM Journal on Discrete Mathematics, 26(2): 687–694, ???? 2012. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hill:2013:RSC

[HY13] Owen Hill and Gexin Yu. A relaxation of Steinberg’s Conjecture. SIAM Journal on Discrete Mathematics, 27(1): 584–596, ???? 2013. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Henning:2015:TTH

[HY15] Michael A. Henning and Anders Yeo. Total transversals in hypergraphs and their applications. SIAM Journal on Discrete Mathematics, 29(1): 309–320, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hyun:2010:SFP

[Hyu10] Jong Yoon Hyun. A subgroup of the full poset-isometry group. SIAM Journal on Discrete Mathematics, 24(2): 589–599, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hell:1995:EHO

[HZ95] Pavol Hell and Xu Ding Zhu. The existence of homomorphisms to oriented cycles. SIAM Journal on Discrete Mathematics, 8(2): 208–222, May 1995. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

He:2008:NOV

[HZ08] Xin He and Huaming Zhang. Nearly optimal visibility representations of plane graphs. SIAM Journal on Discrete Mathematics, 22(4):1364– 1380, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Hatami:2010:FCN

[HZ10] Hamed Hatami and Xuding Zhu. The fractional chromatic number of graphs of maximum degree at most three. SIAM Journal on Discrete Mathematics, 23(4): 1762–1775, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Han:2021:RFA

[HZ21] Dongchun Han and Hanbin Zhang. A reciprocity on finite Abelian groups involving zero-sum sequences. SIAM Journal on Discrete Mathematics, 35(2):1077–1095, ???? 2021. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Iglesias:2014:CVT

[IIL14] Jennifer Iglesias, Nathaniel Ince, and Po-Shen Loh. Computing with voting trees. SIAM Journal on Discrete Mathematics, 28(2):673–684, ???? 2014. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Iosevich:2009:EFD

[IK09] Alex Iosevich and Doowon Koh. The Erdős–Falconer distance problem, exponential sums, and Fourier analytic approach to incidence theorems in vector spaces over finite fields. SIAM Journal on Discrete Mathematics, 23(1): 123–135, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Isler:2006:RPE

[IKK06] Volkan Isler, Sampath Kannan, and Sanjeev Khanna. Randomized pursuit-evasion with local visibility. SIAM Journal on Discrete Mathematics, 20(1):26–41, January 2006. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kawarabayashi:2011:IAH

[iKK11] Ken ichi Kawarabayashi and Yusuke Kobayashi. An improved algorithm for the half-disjoint paths problem. SIAM Journal on Discrete Mathematics, 25(3):1322– 1330, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v25/i3/p1322_s1.

Ito:2022:SRP

[IKK^{+}22] Takehiro Ito, Naonori Kakimura, Naoyuki Kamiyama, Yusuke Kobayashi, and Yoshio Okamoto. Shortest reconfiguration of perfect matchings via alternating cycles. SIAM Journal on Discrete Mathematics, 36(2):1102–1123, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/20M1364370.

Kawarabayashi:2009:CGK

[iKKKL09] Ken ichi Kawarabayashi, Daniel Král’, Jan Kynčl, and Bernard Lidický. 6critical graphs on the Klein bottle. SIAM Journal on Discrete Mathematics, 23(1): 372–383, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Isaak:1992:CG

[IKM^{+}92] Garth Isaak, Suh-Ryung Kim, Terry A. McKee, F. R. McMorris, and Fred S. Roberts. 2-competition graphs. SIAM Journal on Discrete Mathematics, 5(4): 524–538, November 1992. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Imrich:1999:MGT

[IKM99] Wilfried Imrich, Sandi Klavzar, and Henry Martyn Mulder.

Median graphs and trianglefree graphs. SIAM Journal on Discrete Mathematics, 12 (1):111–118, February 1999. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/32349.

Kawarabayashi:2010:SCA

[iKM10] Ken ichi Kawarabayashi and Bojan Mohar. Star coloring and acyclic coloring of locally planar graphs. SIAM Journal on Discrete Mathematics, 24(1):56–71, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kawarabayashi:2009:MTK

[iKMN09] Ken ichi Kawarabayashi, Raiji Mukae, and Atsuhiro Nakamoto. K_{6}-minors in triangulations on the Klein bottle. SIAM Journal on Discrete Mathematics, 23(1): 96–108, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kawarabayashi:2016:CTG

[iKO16] Ken ichi Kawarabayashi and Kenta Ozeki. 5connected toroidal graphs are Hamiltonian-connected. SIAM Journal on Discrete Mathematics, 30(1):112–140, ???? 2016. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kawarabayashi:2004:ONG

[iKSZ04] Ken ichi Kawarabayashi, Chris Stephens, and Xiaoya Zha. Orientable and nonorientable genera for some complete tripartite graphs. SIAM Journal on Discrete Mathematics, 18(3):479–487, 2004. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/42931.

Kawarabayashi:2020:MVV

[iKX20] Ken ichi Kawarabayashi and Chao Xu. Minimum violation vertex maps and their applications to cut problems. SIAM Journal on Discrete Mathematics, 34(4): 2183–2207, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ishikawa:2008:EMS

[IKZ08] Masao Ishikawa, Anisse Kasraoui, and Jiang Zeng. Euler– Mahonian statistics on ordered set partitions. SIAM Journal on Discrete Mathematics, 22(3):1105–1137, ???? 2008. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Isgur:2016:CNF

[ILM^{+}16] A. Isgur, R. Lech, S. Moore, S. Tanny, Y. Verberne, and Y. Zhang. Constructing new families of nested recursions with slow solutions. SIAM Journal on Discrete Mathematics, 30(2):1128–1147, ???? 2016. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Im:2020:BBN

[ILM20] Sungjin Im, Shi Li, and Benjamin Moseley. Breaking 1 − 1/e barrier for nonpreemptive throughput maximization. SIAM Journal on Discrete Mathematics, 34(3): 1649–1669, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Iwata:1996:HPS

[IM96] Satoru Iwata and Kazuo Murota. Horizontal principal structure of layered mixed matrices: Decomposition of discrete systems by designvariable selections. SIAM Journal on Discrete Mathematics, 9(1):71–86, February 1996. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Impagliazzo:2014:EPC

[IMR14] Russell Impagliazzo, Cristopher Moore, and Alexander Russell. An entropic proof of Chang’s inequality. SIAM Journal on Discrete Mathematics, 28(1):173–176, ???? 2014. CODEN SJDMEC.

ISSN 0895-4801 (print), 1095-7146 (electronic).

Iwata:2005:SPC

[IMS05] Satoru Iwata, S. Thomas McCormick, and Maiko Shigeno. A strongly polynomial cut canceling algorithm for minimum cost submodular flow. SIAM Journal on Discrete Mathematics, 19(2):304–320, 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/36153.

Ivanov:2022:QHT

[IN22] Grigory Ivanov and Márton Naszódi. A quantitative Helly-type theorem: Containment in a homothet. SIAM Journal on Discrete Mathematics, 36(2):951–957, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL

                                        https://epubs.siam.org/
                                        doi/10.1137/21M1403308.

Oum:2008:RWW

[iO08] Sang il Oum. Rank-width and well-quasi-ordering. SIAM Journal on Discrete Mathematics, 22(2):666–682, ???? 2008. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ibaraki:1991:WTL

[IP91] T. Ibaraki and S. Poljak. Weak three-linking in Eulerian digraphs. SIAM Journal on Discrete Mathematics, 4(1):84–98, February 1991. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Iriarte:2016:LEL

[Iri16] Benjamin Iriarte. Largest eigenvalue of the Laplacian matrix: Its eigenspace and transitive orientations. SIAM Journal on Discrete Mathematics, 30(4):2146–2161, ???? 2016. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Idury:1993:TTC

[IS93] Ramana M. Idury and Alejandro A. Schäffer. Triangulating three-colored graphs in linear time and linear space. SIAM Journal on Discrete Mathematics, 6(2): 289–293, May 1993. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Iwata:2008:IEF

[IT08] Satoru Iwata and Kenjiro Takazawa. The independent even factor problem. SIAM Journal on Discrete Mathematics, 22(4):1411–1427, ???? 2008. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Iwata:2009:CDA

[IT09] Satoru Iwata and Mizuyo Takamatsu. Computing the degrees of all cofactors in mixed polynomial matrices. SIAM Journal on Discrete Mathematics, 23(2): 647–660, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Tanigawa:2012:GRM

[iT12] Shin ichi Tanigawa. Generic rigidity matroids with Dilworth truncations. SIAM Journal on Discrete Mathematics, 26(3):1412–1439, ???? 2012. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Imase:1991:DST

[IW91] Makoto Imase and Bernard M. Waxman. Dynamic Steiner tree problem. SIAM Journal on Discrete Mathematics, 4(3):369–384, August 1991. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Janssen:2010:RBA

[Ja10] Jeannette Janssen and Pawe Pra at. Rank-based attachment leads to power law graphs. SIAM Journal on Discrete Mathematics, 24(2): 420–440, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Just:2016:LAT

[JA16] Winfried Just and Sungwoo Ahn. Lengths of attractors and transients in neuronal networks with random connectivities. SIAM Journal on Discrete Mathematics, 30(2): 912–933, ???? 2016. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Jacobson:1992:ECC

[Jac92] Michael S. Jacobson. On the p-edge clique cover number of complete bipartite graphs. SIAM Journal on Discrete Mathematics, 5(4): 539–544, November 1992. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Jaeschke:1989:SRC

[Jae89] Gerhard Jaeschke. On solving rational congruences. SIAM Journal on Discrete Mathematics, 2(1):73–90, February 1989. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Janusz:2000:OCP

[Jan00] Gerald J. Janusz. Overlap and covering polynomials with applications to designs and self-dual codes. SIAM Journal on Discrete Mathematics, 13(2):154–178, 2000. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/34176.

Janata:2005:MIP

[Jan05]  Marek Janata.  Matroids

induced by packing subgraphs. SIAM Journal on Discrete Mathematics, 18 (3):525–541, 2005. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/37983.

Jansen:2010:ESJ

[Jan10] Klaus Jansen. An EPTAS for scheduling jobs on uniform processors: Using an MILP relaxation with a constant number of integral variables. SIAM Journal on Discrete Mathematics, 24(2): 457–485, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Janzer:2020:ENS

[Jan20]  Oliver Janzer.  The ex-

tremal number of the subdivisions of the complete bipartite graph. SIAM Journal on Discrete Mathematics, 34(1): 241–250, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Janzer:2022:GRT

[Jan22] Barnabás Janzer. The generalized rainbow Turán problem for cycles. SIAM Journal on Discrete Mathematics, 36(1):436–448, ???? 2022.

CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1351473.

Jiang:2022:EIP

[JB22] Hongyi Jiang and Amitabh Basu. Enumerating integer points in polytopes with bounded subdeterminants. SIAM Journal on Discrete Mathematics, 36 (1):449–460, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M139935X.

Jiang:2006:OIT

[JCB06] Anxiao (Andrew) Jiang, Matthew Cook, and Jehoshua Bruck. Optimal interleaving on tori. SIAM Journal on Discrete Mathematics, 20(4): 841–879, January 2006. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Jordan:2013:GSR

[JDT13] Tibor Jordán, Gábor Domokos, and Krisztina Tóth. Geometric sensitivity of rigid graphs. SIAM Journal on Discrete Mathematics, 27(4): 1710–1726, ???? 2013. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Jedwab:1993:BAE

[Jed93] Jonathan Jedwab. Barker arrays. I. Even number of elements. SIAM Journal on Discrete Mathematics, 6(2): 294–308, May 1993. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Jevtic:1995:FSI

[Jev95] Dušan B. Jevtić. On families of sets of integral vectors whose representatives form sum-distinct sets. SIAM Journal on Discrete Mathematics, 8(4):652–660, November 1995. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Jia:1995:ECD

[Jia95] Xing De Jia. Extremal Cayley digraphs of finite cyclic groups. SIAM Journal on Discrete Mathematics, 8(1): 62–75, February 1995. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Jordan:2022:RRS

[JiT22] Tibor Jordán and Shin ichi Tanigawa. Rigidity of random subgraphs and eigenvalues of stiffness matrices. SIAM Journal on Discrete Mathematics, 36(3):2367–2392, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL

                                        https://epubs.siam.org/
                                        doi/10.1137/20M1349849.

Jackson:2014:CCU

[JJiT14] Bill Jackson, Tibor Jordán, and Shin ichi Tanigawa. Combinatorial conditions for the unique completability of lowrank matrices. SIAM Journal on Discrete Mathematics, 28(4):1797–1819, ???? 2014. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Janzer:2023:PSL

[JJL23] Barnabás Janzer, J. Robert Johnson, and Imre Leader. Partial shuffles by lazy swaps. SIAM Journal on Discrete Mathematics, 37(4):2544– 2557, October 2023. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Janssen:1999:BSS

[JK99] Jeannette Janssen and Kyriakos Kilakos. Bounded stable sets: Polytopes and colorings. SIAM Journal on Discrete Mathematics, 12(2):262–275, May 1999. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/30978.

Jansen:2019:RAO

[JK19] Klaus Jansen and Kim-Manuel Klein. A robust AF-PTAS for online bin packing with polynomial migration. SIAM Journal on Discrete Mathematics, 33(4): 2062–2091, ???? 2019. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Jiradilok:2024:TDB

[JK24] Pakawut Jiradilok and Supanat Kamtue. Transportation distance between probability measures on the infinite regular tree. SIAM Journal on Discrete Mathematics, 38 (1):1113–1157, March 2024. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kwon:2020:SCG

[jKiO20] O joung Kwon and Sang il Oum. Scattered classes of graphs. SIAM Journal on Discrete Mathematics, 34(1): 972–999, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Jeong:2021:FBD

[JKiO21] Jisu Jeong, Eun Jung Kim, and Sang il Oum. Finding branch-decompositions of matroids, hypergraphs, and more. SIAM Journal on Discrete Mathematics, 35(4): 2544–2617, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kwon:2021:PCI

[jKR21] O joung Kwon and Jean-Florent Raymond. Packing and covering induced subdivisions. SIAM Journal on Discrete Mathematics, 35(2): 597–636, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Jin:2017:PGC

[JKSW17] Ligang Jin, Yingli Kang, Michael Schubert, and Yingqian Wang. Plane graphs without 4- and 5-cycles and without ext-triangular 7-cycles are 3colorable. SIAM Journal on Discrete Mathematics, 31(3): 1836–1847, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Jartoux:2022:MRN

[JKSY22] Bruno Jartoux, Chaya Keller, Shakhar Smorodinsky, and Yelena Yuditsky. On multicolor Ramsey numbers and subset coloring of hypergraphs. SIAM Journal on Discrete Mathematics, 36 (3):1848–1860, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1462003.

Joret:2020:PAV

[JL20a] Gwenaël Joret and William Lochet. Progress on the adjacent vertex distinguishing edge coloring conjecture. SIAM Journal on Discrete Mathematics, 34(4):

2221–2238, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Joswig:2020:MTC

[JL20b] Michael Joswig and Georg Loho. Monomial tropical cones for multicriteria optimization. SIAM Journal on Discrete Mathematics, 34(2): 1172–1191, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Jiang:2023:TDG

[JL23] Tao Jiang and Sean Longbrake. Tree-degenerate graphs and nested dependent random choice. SIAM Journal on Discrete Mathematics, 37(3):1805–1817, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1483554.

Jekel:2018:APG

[JLD^{+}18] David Jekel, Avi Levy, Will Dana, Austin Stromme, and Collin Litterell. Algebraic properties of generalized graph Laplacians: Resistor networks, critical groups, and homological algebra. SIAM Journal on Discrete Mathematics, 32(2):1040–1110, ???? 2018. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Jansen:2016:BRT

[JLL16] K. Jansen, F. Land, and K. Land. Bounding the running time of algorithms for scheduling and packing problems. SIAM Journal on Discrete Mathematics, 30(1): 343–366, ???? 2016. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Jedwab:1993:BAI

[JLM93] Jonathan Jedwab, Sheelagh Lloyd, and Miranda Mowbray. Barker arrays. II. Odd number of elements. SIAM Journal on Discrete Mathematics, 6(2):309–328, May 1993. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Jones:2017:PCD

[JLR^{+}17] Mark Jones, Daniel Lokshtanov, M. S. Ramanujan, Saket Saurabh, and Ondrej Suchý. Parameterized complexity of directed Steiner tree on sparse graphs. SIAM Journal on Discrete Mathematics, 31(2):1294–1327, ???? 2017. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Jansen:2020:NLT

[JLR20] Klaus Jansen, Alexandra Lassota, and Lars Rohwedder. Near-linear time algorithm for n-fold ILPs via color coding. SIAM Journal on Discrete Mathematics, 34(4): 2282–2299, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Juvan:1997:OMB

[JM97] Martin Juvan and Bojan Mohar. Obstructions for 2-Möbius band embedding extension problem. SIAM Journal on Discrete Mathematics, 10(1):57–72, February 1997. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Janssen:2017:RSS

[JM17] Jeannette Janssen and Abbas Mehrabian. Rumors spread slowly in a small-world spatial network. SIAM Journal on Discrete Mathematics, 31(4): 2414–2428, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Janzer:2022:TNB

[JMN22] Oliver Janzer, Abhishek Methuku, and Zoltán Lóránt Nagy. On the Turán number of the blow-up of the hexagon. SIAM Journal on Discrete Mathematics, 36 (2):1187–1199, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1428510.

Jacopini:1990:RTM

[JMS90] G. Jacopini, P. Mentrasti, and G. Sontacchi. Reversible Turing machines and polynomial time reversibly computable functions. SIAM Journal on Discrete Mathematics, 3(2):241–254, May 1990. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Jiang:1999:EBG

[JMSW99] Tao Jiang, Dhruv Mubayi, Aditya Shastri, and Douglas B. West. Edgebandwidth of graphs. SIAM Journal on Discrete Mathematics, 12(3):307–316, August 1999. CODEN SJD-MEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/33075. See cor-

rection [JMSW00].

Jiang:2000:CEB

[JMSW00] Tao Jiang, Dhruv Mubayi, Aditya Shastri, and Douglas B. West. Correction to “Edge-Bandwidth of Graphs”. SIAM Journal on Discrete Mathematics, 13(1):1, 2000. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/97411.  See

[JMSW99].

Jungeblut:2023:SBP

[JMU23] Paul Jungeblut, Laura Merker, and Torsten Ueckerdt. A sublinear bound on the page number of upward planar graphs. SIAM Journal on Discrete Mathematics, 37 (4):2312–2331, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1522450.

Joret:2017:PPD

[JMW17] Gwenaël Joret, Piotr Micek, and Veit Wiechert. Planar posets have dimension at most linear in their height. SIAM Journal on Discrete Mathematics, 31(4):2754– 2790, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Joos:2016:IMGa

[JN16] Felix Joos and Viet Hang Nguyen. Induced matchings in graphs of degree at most 4. SIAM Journal on Discrete Mathematics, 30(1): 154–165, ???? 2016. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Jiang:2017:SDS

[JN17] Tao Jiang and Andrew Newman. Small dense subgraphs of a graph. SIAM Journal on Discrete Mathematics, 31 (1):124–142, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Jackson:2021:IBR

[JNiT21] Bill Jackson, Anthony Nixon, and Shin ichi Tanigawa. An improved bound for the rigidity of linearly constrained frameworks. SIAM Journal on Discrete Mathematics, 35 (2):928–933, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Jamison:1995:CHD

[JO95] Beverly Jamison and Stephan Olariu. p-components and the homogeneous decomposition of graphs. SIAM Journal on Discrete Mathematics, 8(3): 448–463, August 1995. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Javadi:2018:QEF

[JO18] Ramin Javadi and Gholamreza Omidi. On a question of Erdős and Faudree on the size Ramsey numbers. SIAM Journal on Discrete Mathematics, 32(3):2217–2228, ???? 2018. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Jonsson:2005:SCG

[Jon05] Jakob Jonsson. Simplicial complexes of graphs and hypergraphs with a bounded covering number. SIAM Journal on Discrete Mathematics, 19(3):633–650, 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/44368.

Jones:2020:WGB

[Jon20] Brant Jones. Weighted games of best choice. SIAM Journal on Discrete Mathematics, 34 (1):399–414, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Joos:2015:RSS

[Joo15] Felix Joos. Random subgraphs in sparse graphs. SIAM Journal on Discrete Mathematics, 29(4):2350– 2360, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Joos:2016:IMGb

[Joo16] Felix Joos. Induced matchings in graphs of bounded maximum degree. SIAM Journal on Discrete Mathematics, 30(3):1876–1882, ???? 2016. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Jordan:2003:CES

[Jor03] Tibor Jordán. Constrained edge-splitting problems. SIAM Journal on Discrete Mathematics, 17(1):88–102, 2003.

CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/36448.

Jansen:2006:PRC

[JP06a] Klaus Jansen and Lorant Porkolab. On preemptive resource constrained scheduling: Polynomial-time approximation schemes. SIAM Journal on Discrete Mathematics, 20(3):545–563, January 2006. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Joswig:2006:COM

[JP06b] Michael Joswig and Marc E.

Pfetsch. Computing optimal Morse matchings. SIAM Journal on Discrete Mathematics, 20(1):11–25, January 2006. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Janson:2012:HTR

[JP12] Svante Janson and Yuval Peres. Hitting times for random walks with restarts. SIAM Journal on Discrete Mathematics, 26(2):537–547, ???? 2012. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Jansen:2018:AKC

[JP18] Bart M. P. Jansen and Marcin Pilipczuk. Approximation and kernelization for chordal vertex deletion. SIAM Journal on Discrete Mathematics, 32(3):2258–2301, ???? 2018. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Joret:2014:HHP

[JPS^{+}14] Gwenaël Joret, Christophe Paul, Ignasi Sau, Saket Saurabh, and Stéphan Thomassé. Hitting and harvesting pumpkins. SIAM Journal on Discrete Mathematics, 28(3): 1363–1390, ???? 2014. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Jiang:2012:DPG

[JPT12] Minghui Jiang, Vincent Pilaud, and Pedro J. Tejada. On a dispersion problem in grid labeling. SIAM Journal on Discrete Mathematics, 26 (1):39–51, ???? 2012. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v26/i1/p39_s1.

Jansen:2021:DPK

[JPvL21] Bart M. P. Jansen, Marcin L. Pilipczuk, and Erik Jan van Leeuwen. A deterministic polynomial kernel for odd cycle transversal and vertex multiway cut in planar graphs. SIAM Journal on Discrete Mathematics, 35(4): 2387–2429, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Jiang:2010:SSS

[JPY10] Tao Jiang, Oleg Pikhurko, and Zelealem Yilma. Set systems without a strong simplex. SIAM Journal on Discrete Mathematics, 24(3): 1038–1045, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Jojic:2021:SNC

[JPZ21] Dusko Jojić, Gaiane Panina, and Rade Zivaljević. Splitting necklaces, with constraints. SIAM Journal on Discrete Mathematics, 35(2): 1268–1286, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Jiang:2020:TNB

[JQ20] Tao Jiang and Yu Qiu. Turán numbers of bipartite subdivisions. SIAM Journal on Discrete Mathematics, 34(1): 556–570, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Joos:2014:IMS

[JRS14] Felix Joos, Dieter Rautenbach, and Thomas Sasse. Induced matchings in subcubic graphs. SIAM Journal on Discrete Mathematics, 28(1): 468–473, ???? 2014. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Jordan:2003:DPL

[JS03] Tibor Jordán and Zoltán Szigeti. Detachments preserving local edge-connectivity of graphs. SIAM Journal on Discrete Mathematics, 17(1):72–87, 2003. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/36393.

Jiang:2012:TNS

[JS12] Tao Jiang and Robert Seiver. Turán numbers of subdivided graphs. SIAM Journal on Discrete Mathematics, 26(3): 1238–1255, ???? 2012. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Jochemko:2014:AMO

[JS14] Katharina Jochemko and Raman Sanyal. Arithmetic of marked order polytopes, monotone triangle reciprocity, and partial colorings. SIAM Journal on Discrete Mathematics, 28(3): 1540–1558, ???? 2014. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Jia:2017:EBF

[JS17] Xingde Jia and Jian Shen. Extremal bases for finite cyclic groups. SIAM Journal on Discrete Mathematics, 31 (2):796–804, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Jansen:2012:PSP

[JSO12] Klaus Jansen and Roberto Solis-Oba. Packing squares with profits. SIAM Journal on Discrete Mathematics, 26 (1):263–279, ???? 2012. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Jansen:2003:MMJ

[JSOS03] Klaus Jansen, Roberto Solis-Oba, and Maxim Sviridenko. Makespan minimization in job shops: a linear time approximation scheme. SIAM Journal on Discrete Mathematics, 16(2):288–300, 2003. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/36390.

Junosza-Szaniawski:2018:OCL

[JSRSW18] Konstanty Junosza-Szaniawski, Paweł Rzazewski, Joanna Sokół, and Krzysztof Wesek. Online coloring and L(2, 1)- labeling of unit disk intersection graphs. SIAM Journal on Discrete Mathematics, 32(2):1335–1350, ???? 2018. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Jiang:2015:MIF

[JSZ15] Tian-Yi Jiang, Ziv Scully, and Yan X. Zhang. Motors and impossible firing patterns in the parallel chipfiring game. SIAM Journal on Discrete Mathematics, 29 (1):615–630, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Jaslar:2011:MWP

[JT11] Steven Jaslar and Sekhar Tatikonda. Maximum weight partial colorings on sparse random graphs. SIAM Journal on Discrete Mathematics, 25(2):934–955, ???? 2011. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:

                                        //epubs.siam.org/sidma/
                                        resource/1/sjdmec/v25/i2/
                                        p934_s1.

Jukna:2016:TCS

[Juk16] Stasys Jukna. Tropical complexity, Sidon sets, and dynamic programming. SIAM Journal on Discrete Mathematics, 30(4):2064–2085, ???? 2016. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Jukna:2021:NHF

[Juk21] Stasys Jukna. Notes on hazard-free circuits. SIAM Journal on Discrete Mathematics, 35(2):770–787, ???? 2021. CODEN SJDMEC.

ISSN 0895-4801 (print), 1095-7146 (electronic).

Jungers:2012:SPF

[Jun12] Raphaël M. Jungers. The synchronizing probability function of an automaton. SIAM Journal on Discrete Mathematics, 26(1): 177–192, ???? 2012. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Juttner:2006:BOP

[Jüt06] Alpár Jüttner. On budgeted optimization problems. SIAM Journal on Discrete Mathematics, 20(4): 880–892, January 2006. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Joret:2018:MFI

[JW18] Gwenaël Joret and David R.

Wood. K_{4}-minor-free induced subgraphs of sparse connected graphs. SIAM Journal on Discrete Mathematics, 32(1):123–147, ???? 2018. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Janssen:2005:LBT

[JWF05] J. C. M. Janssen, T. E.

Wentzell, and S. L. Fitzpatrick. Lower bounds from tile covers for the channel assignment problem. SIAM Journal on Discrete Mathematics, 18(4):679–696, 2005.

CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/38440.

Ji:2023:NCG

[JY23] Naidan Ji and Dong Ye. The number of cliques in graphs covered by long cycles. SIAM Journal on Discrete Mathematics, 37(2):917–924, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL

                                        https://epubs.siam.org/
                                        doi/10.1137/22M1476046.

Ji:2005:RSQ

[JZ05] L. Ji and L. Zhu. Resolvable Steiner quadruple systems for the last 23 orders. SIAM Journal on Discrete Mathematics, 19(2):420–430, 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/44058.

Kahale:1997:IIE

[Kah97] Nabil Kahale. Isoperimetric inequalities and eigenvalues. SIAM Journal on Discrete Mathematics, 10(1):30– 40, February 1997. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kamiyama:2017:PMT

[Kam17] Naoyuki Kamiyama. Popular matchings with ties and matroid constraints. SIAM Journal on Discrete Mathematics, 31(3):1801–1819, ???? 2017. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kamiyama:2019:PSM

[Kam19] Naoyuki Kamiyama. Pareto stable matchings under onesided matroid constraints. SIAM Journal on Discrete Mathematics, 33(3):1431– 1451, ???? 2019. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kamiyama:2022:MGS

[Kam22] Naoyuki Kamiyama. A matroid generalization of the super-stable matching problem. SIAM Journal on Discrete Mathematics, 36 (2):1467–1482, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1437214.

Khosrovshahi:1990:NBT

[KAN90] G. B. Khosrovshahi and S. Ajoodani-Namini. A new basis for trades. SIAM Journal on Discrete Mathematics, 3(3):364–372, August 1990. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kang:2008:LHG

[Kan08] Jeong-Hyun Kang. L(2, 1)- labeling of Hamiltonian graphs with maximum degree 3. SIAM Journal on Discrete Mathematics, 22(1):213–230, ???? 2008. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kao:1996:DSE

[Kao96] Ming-Yang Kao. Data security equals graph connectivity. SIAM Journal on Discrete Mathematics, 9(1):87– 100, February 1996. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kapadia:2014:MMP

[Kap14a] Rohan Kapadia. Matroids with a modular 4-point line. SIAM Journal on Discrete Mathematics, 28(2): 862–877, ???? 2014. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kaplan:2014:MIT

[Kap14b] Nathan Kaplan. MacWilliams identities for m-tuple weight enumerators. SIAM Journal on Discrete Mathematics, 28 (1):428–444, ???? 2014. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Karloff:1989:HLC

[Kar89] Howard J. Karloff. How long can a Euclidean traveling salesman tour be. SIAM Journal on Discrete Mathematics, 2(1):91–99, February 1989. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Karp:1992:TSG

[Kar92] Richard M. Karp. Threestage generalized connectors. SIAM Journal on Discrete Mathematics, 5(2): 259–272, May 1992. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kardos:2020:CAP

[Kar20] Frantisek Kardos. A computerassisted proof of the Barnette– Goodey conjecture: Not only fullerene graphs are Hamiltonian. SIAM Journal on Discrete Mathematics, 34(1): 62–100, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kashyap:2003:MSC

[Kas03] Navin Kashyap. Maximizing the Shannon capacity of constrained systems with two constraints. SIAM Journal on Discrete Mathematics, 17(2):298–319, 2003. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/40275.

Kaski:2005:IFE

[Kas05] Petteri Kaski. Isomorph-free exhaustive generation of designs with prescribed groups of automorphisms. SIAM Journal on Discrete Mathematics, 19(3):664–690, 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/44478.

Kashyap:2008:MPC

[Kas08] Navin Kashyap. Matroid pathwidth and code trellis complexity. SIAM Journal on Discrete Mathematics, 22(1): 256–272, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kavitha:2024:MMP

[Kav24] Telikepalli Kavitha. Maximum matchings and popularity. SIAM Journal on Discrete Mathematics, 38(2): 1202–1221, April 2024. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Khachiyan:2005:CSE

[KBE^{+}05] L. Khachiyan, E. Boros, K. Elbassioni, V. Gurvich, and K. Makino. On the complexity of some enumeration problems for matroids. SIAM Journal on Discrete Mathematics, 19(4): 966–984, January 2005. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kwak:1998:ERG

[KCL98] Jin Ho Kwak, Jang-Ho Chun, and Jaeun Lee. Enumeration of regular graph coverings having finite Abelian covering transformation groups. SIAM Journal on Discrete Mathematics, 11(2):273–285, May 1998. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ku-Cauich:2013:SAC

[KCTR13] J. C. Ku-Cauich and H. Tapia-Recillas. Systematic authentication codes based on a class of bent functions and the Gray map on a Galois ring. SIAM Journal on Discrete Mathematics, 27(2): 1159–1170, ???? 2013. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kao:1994:OPA

[KFHR94] Ming-Yang Kao, Martin Fürer, Xin He, and Balaji Raghavachari. Optimal parallel algorithms for straightline grid embeddings of planar graphs. SIAM Journal on Discrete Mathematics, 7 (4):632–646, November 1994. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kao:1993:EDP

[KG93] Ming-Yang Kao and Dan Gusfield. Efficient detection and protection of information in cross tabulated tables. I. Linear invariant test. SIAM Journal on Discrete Mathematics, 6(3):460–476, August 1993. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kleinberg:1998:LTF

[KG98] Jon Kleinberg and Michel X. Goemans. The lovász theta function and a semidefinite programming relaxation of vertex cover. SIAM Journal on Discrete Mathematics, 11 (2):196–204, May 1998. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Khan:2013:PMU

[Kha13] Imdadullah Khan. Perfect matchings in 3-uniform hypergraphs with large vertex degree. SIAM Journal on Discrete Mathematics, 27(2): 1021–1039, ???? 2013. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kierstead:1988:LFF

[Kie88] H. A. Kierstead. The linearity of first-fit coloring of interval graphs. SIAM Journal on Discrete Mathematics, 1 (4):526–530, November 1988. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kierstead:1997:CGV

[Kie97] H. A. Kierstead. Classes of graphs that are not vertex Ramsey. SIAM Journal on Discrete Mathematics, 10(3): 373–380, August 1997. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kakimura:2012:PDC

[KiK12] Naonori Kakimura and Ken ichi Kawarabayashi. Packing directed circuits through prescribed vertices bounded fractionally. SIAM Journal on Discrete Mathematics, 26(3): 1121–1133, ???? 2012. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kakimura:2017:PED

[KiKK17] Naonori Kakimura, Ken ichi Kawarabayashi, and Yusuke Kobayashi. Packing edgedisjoint odd Eulerian subgraphs through prescribed vertices in 4-edge-connected graphs. SIAM Journal on Discrete Mathematics, 31(2): 766–782, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kim:1991:OPA

[Kim91] Sung Kwon Kim. Optimal parallel algorithms for region labeling and medial axis transform of binary images. SIAM Journal on Discrete Mathematics, 4(3):385– 396, August 1991. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kim:1992:CAB

[Kim92] Dongsu Kim. A combinatorial approach to biorthogonal polynomials. SIAM Journal on Discrete Mathematics, 5(3):413–421, August 1992. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kim:2011:FRS

[Kim11]  Jang Soo Kim.  Front

representation of set partitions. SIAM Journal on Discrete Mathematics, 25(1): 447–461, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sidma/resource/
                                        1/sjdmec/v25/i1/p447_s1.

Kim:2017:USL

[Kim17] Ringi Kim. Unavoidable subtournaments in large tournaments with no homogeneous sets. SIAM Journal on Discrete Mathematics, 31(2): 714–725, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kim:2024:WGS

[KíP^{+}24] Eun Jung Kim, Tomá Masa ík, Marcin Pilipczuk, Roohani Sharma, and Magnus Wahlström. On weighted graph separation problems and flow augmentation. SIAM Journal on Discrete Mathematics, 38(1): 170–189, January 2024. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kiraly:2016:MIA

[Kir16] Csaba Király. On maximal independent arborescence packing. SIAM Journal on Discrete Mathematics, 30(4): 2107–2114, ???? 2016. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kitchens:2002:MCC

[Kit02]  Bruce Kitchens.  Multi-

dimensional convolutional codes. SIAM Journal on Discrete Mathematics, 15 (3):367–381, 2002. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/37849.

Katoh:2013:RTD

[KiT13] Naoki Katoh and Shin ichi Tanigawa. Rooted-tree decompositions with matroid constraints and the infinitesimal rigidity of frameworks with boundaries. SIAM Journal on Discrete Mathematics, 27(1):155–185, ???? 2013. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Krokhin:2004:CSP

[KJJ04] Andrei Krokhin, Peter Jeavons, and Peter Jonsson. Constraint satisfaction problems on intervals and lengths. SIAM Journal on Discrete Mathematics, 17(3): 453–477, 2004. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/41020.

Kohli:1989:APH

[KK89] Rajeev Kohli and Ramesh Krishnamurti. Average performance of heuristics for satisfiability. SIAM Journal on Discrete Mathematics, 2 (4):508–523, November 1989. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Klawe:1990:ALT

[KK90] Maria M. Klawe and Daniel J. Kleitman. An almost linear time algorithm for generalized matrix searching. SIAM Journal on Discrete Mathematics, 3(1):81–97, February 1990. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kortsarz:2001:AAN

[KK01] Guy Kortsarz and Robert Krauthgamer. On approximating the achromatic number. SIAM Journal on Discrete Mathematics, 14 (3):408–422, 2001. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/37957.

Kunszenti-Kovacs:2009:NPA

[KK09]  Dávid  Kunszenti-Kovács.

Network perturbations and asymptotic periodicity of recurrent flows in infinite networks. SIAM Journal on Discrete Mathematics, 23(3): 1561–1574, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Katz:2010:ADE

[KK10] Nets Hawk Katz and Paul Koester. On additive doubling and energy. SIAM Journal on Discrete Mathematics, 24(4):1684–1693, ???? 2010. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v24/i4/p1684_s1.

Kaminski:2014:BVB

[KK14a] Marcin Kamiński and Lukasz Kowalik. Beyond the Vizing’s bound for at most seven colors. SIAM Journal on Discrete Mathematics, 28(3): 1334–1362, ???? 2014. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kocay:2014:RGT

[KK14b] William L. Kocay and Donald L. Kreher. On reconstructing graphs and their complements. SIAM Journal on Discrete Mathematics, 28(3):1026–1034, ???? 2014. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kral:2012:EFP

[KKK^{+}12] Daniel Král’, Matjaz Krnc, Martin Kupec, Borut Luzar, and Jan Volec. Extending fractional precolorings. SIAM Journal on Discrete Mathematics, 26(2):647–660, ???? 2012. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kang:2017:SSC

[KKKS17] Dong Yeap Kang, Jaehoon Kim, Younjin Kim, and Geewon Suh. Sparse spanning kconnected subgraphs in tournaments. SIAM Journal on Discrete Mathematics, 31(3): 2206–2227, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kaiser:2010:SCC

[KKL^{+}10] Tomáš Kaiser, Daniel Král’, Bernard Lidický, Pavel Nejedlý, and Robert Šámal. Short cycle covers of graphs with minimum degree three. SIAM Journal on Discrete Mathematics, 24(1):330–355, ???? 2010. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kim:2019:TCR

[KKL19] Jaehoon Kim, Younjin Kim, and Hong Liu. Two conjectures in Ramsey–Turán theory. SIAM Journal on Discrete Mathematics, 33(1): 564–586, ???? 2019. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kohli:1994:MSP

[KKM94] Rajeev Kohli, Ramesh Krishnamurti, and Prakash Mirchandani. The minimum satisfiability problem. SIAM Journal on Discrete Mathematics, 7(2):275–283, May 1994. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Karchmer:1995:FCC

[KKN95] Mauricio Karchmer, Eyal Kushilevitz, and Noam Nisan. Fractional covers and communication complexity. SIAM Journal on Discrete Mathematics, 8(1):76–92, February 1995. CODEN SJDMEC.

ISSN 0895-4801 (print), 1095-7146 (electronic).

Kim:2016:BHC

[KKO16] Jaehoon Kim, Daniela Kühn, and Deryk Osthus. Bipartitions of highly connected tournaments. SIAM Journal on Discrete Mathematics, 30 (2):895–911, ???? 2016. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kalkowski:2011:NUB

[KKP11] M. Kalkowski, M. Karoński, and F. Pfender. A new upper bound for the irregularity strength of graphs. SIAM Journal on Discrete Mathematics, 25(3):1319– 1321, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sidma/resource/
                                        1/sjdmec/v25/i3/p1319_s1.

Kang:2015:PTM

[KKP15] Mihyun Kang, Christoph Koch, and Angélica Pachón. The phase transition in multitype binomial random graphs. SIAM Journal on Discrete Mathematics, 29(2): 1042–1064, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kardos:2010:LFF

[KKS10] František Kardoš, Dan Král’, and Jean-Sébastien Sereni.

The last fraction of a fractional conjecture. SIAM Journal on Discrete Mathematics, 24(2):699–707, ???? 2010. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Krivelevich:2017:BDS

[KKS17] Michael Krivelevich, Matthew Kwan, and Benny Sudakov. Bounded-degree spanning trees in randomly perturbed graphs. SIAM Journal on Discrete Mathematics, 31(1): 155–171, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kalinowski:2019:ZFN

[KKS19] Thomas Kalinowski, Nina Kamcev, and Benny Sudakov. The zero forcing number of graphs. SIAM Journal on Discrete Mathematics, 33(1): 95–115, ???? 2019. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Koana:2022:ECK

[KKS22] Tomohiro Koana, Christian Komusiewicz, and Frank Sommer. Exploiting c-closure in kernelization algorithms for graph problems. SIAM Journal on Discrete Mathematics, 36(4):2798–2821, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL

                                        https://epubs.siam.org/
                                        doi/10.1137/21M1449476.

Kanj:2020:SPP

[KKSvL20] Iyad Kanj, Christian Komusiewicz, Manuel Sorge, and Erik Jan van Leeuwen. Solving partition problems almost always requires pushing many vertices around. SIAM Journal on Discrete Mathematics, 34(1):640–681, ???? 2020. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kardos:2011:FCC

[KKV11] Frantisek Kardos, Daniel Král’, and Jan Volec. Fractional colorings of cubic graphs with large girth. SIAM Journal on Discrete Mathematics, 25(3):1454– 1476, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sidma/resource/
                                        1/sjdmec/v25/i3/p1454_s1.

Kallal:2017:ILB

[KKW17] Kenz Kallal, Tomoka Kan, and Eric Wang. Improved lower bounds for kissing numbers in dimensions 25 through 31. SIAM Journal on Discrete Mathematics, 31(3): 1895–1908, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Klawe:1992:TLB

[KL92] Maria Klawe and Tom Leighton. A tight lower bound on the size of planar permutation networks. SIAM Journal on Discrete Mathematics, 5(4):558–563, November 1992. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Krokhin:2008:MSF

[KL08] Andrei Krokhin and Benoit Larose. Maximizing supermodular functions on product lattices, with application to maximum constraint satisfaction. SIAM Journal on Discrete Mathematics, 22(1): 312–328, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kim:2014:URG

[KL14] Jeong Han Kim and Sang June Lee. Universality of random graphs for graphs of maximum degree two. SIAM Journal on Discrete Mathematics, 28(3):1467–1478, ???? 2014. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kelk:2019:TKC

[KL19a] Steven Kelk and Simone Linz.

A tight kernel for computing the tree bisection and reconnection distance between two phylogenetic trees. SIAM Journal on Discrete Mathematics, 33(3):1556–1574, ????

2019. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Keller:2019:NLI

[KL19b] Nathan Keller and Noam Lifshitz. A note on large Hintersecting families. SIAM Journal on Discrete Mathematics, 33(1):398–401, ???? 2019. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Klazar:2006:ICN

[Kla06] Martin Klazar. On identities concerning the numbers of crossings and nestings of two edges in matchings. SIAM Journal on Discrete Mathematics, 20(4): 960–976, January 2006. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kleitman:1989:DUC

[Kle89] D. J. Kleitman. Divisors without unit-congruent ratios. SIAM Journal on Discrete Mathematics, 2(3):344– 349, August 1989. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kowaluk:2013:CDS

[KLL13] Miroslaw Kowaluk, Andrzej Lingas, and Eva-Marta Lundell. Counting and detecting small subgraphs via equations. SIAM Journal on Discrete Mathematics, 27(2):

892–909, ???? 2013. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kratsch:2003:ATS

[KLM^{+}03] Dieter Kratsch, Hoàng-Oanh Le, Haiko Müller, Erich Prisner, and Dorothea Wagner. Additive tree spanners. SIAM Journal on Discrete Mathematics, 17(2):332–340, 2003. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29547.

Keevash:2014:SEP

[KLM14] Peter Keevash, John Lenz, and Dhruv Mubayi. Spectral extremal problems for hypergraphs. SIAM Journal on Discrete Mathematics, 28(4): 1838–1854, ???? 2014. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kante:2014:EMD

[KLMN14] Mamadou Moustapha Kanté, Vincent Limouzy, Arnaud Mary, and Lhouari Nourine. On the enumeration of minimal dominating sets and related notions. SIAM Journal on Discrete Mathematics, 28(4):1916–1929, ???? 2014. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kemkes:2003:SNC

[KLMR03] Graeme Kemkes, Chiu Fan Lee, Donatella Merlini, and Bruce Richmond. Stirling numbers for complex arguments: Asymptotics and identities. SIAM Journal on Discrete Mathematics, 16(2):179–191, 2003. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/40111.

Keevash:2013:DGC

[KLMR13] Peter Keevash, Zhentao Li, Bojan Mohar, and Bruce Reed. Digraph girth via chromatic number. SIAM Journal on Discrete Mathematics, 27 (2):693–696, ???? 2013. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kohayakawa:2018:ISS

[KLMR18] Yoshiharu Kohayakawa, Sang June Lee, Carlos Gustavo Moreira, and Vojtech Rödl. Infinite Sidon sets contained in sparse random sets of integers. SIAM Journal on Discrete Mathematics, 32(1): 410–449, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kratochvil:1991:NST

[KLN91] Jan Kratochvíl, Anna Lubiw, and Jaroslav Nešetřil. Noncrossing subgraphs in topological layouts. SIAM Journal on Discrete Mathematics, 4(2):223–244, May 1991. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kortsarz:2010:AMS

[KLN10] Guy Kortsarz, Michael Langberg, and Zeev Nutov. Approximating maximum subgraphs without short cycles. SIAM Journal on Discrete Mathematics, 24(1): 255–269, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

King:2012:FAB

[KLP12] Andrew D. King, Linyuan Lu, and Xing Peng. A fractional analogue of Brooks’ Theorem. SIAM Journal on Discrete Mathematics, 26(2): 452–471, ???? 2012. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kahle:2022:ICA

[KLR22] Thomas Kahle, Dinh Van Le, and Tim Römer. Invariant chains in algebra and discrete geometry. SIAM Journal on Discrete Mathematics, 36(2):975–999, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1385652.

Krivelevich:2010:RPR

[KLS10] Michael Krivelevich, Choongbum Lee, and Benny Sudakov. Resilient pancyclicity of random and pseudorandom graphs. SIAM Journal on Discrete Mathematics, 24 (1):1–16, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kowalik:2018:FGC

[KLS18] Lukasz Kowalik, Juho Lauri, and Arkadiusz Socała. On the fine-grained complexity of rainbow coloring. SIAM Journal on Discrete Mathematics, 32(3):1672–1705, ???? 2018. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kahn:1989:SST

[KLW89] J. Kahn, J. C. Lagarias, and H. S. Witsenhausen. Singlesuit two-person card play. III. The Misère game. SIAM Journal on Discrete Mathematics, 2(3):329–343, August 1989. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kamcev:2021:SRN

[KLWY21] Nina Kamcev, Anita Liebenau, David R. Wood, and Liana Yepremyan. The size Ramsey number of graphs with bounded treewidth. SIAM Journal on Discrete Mathematics, 35(1):281–293,

????  2021.  CODEN

SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Korner:1988:GSE

[KM88] J. Körner and K. Marton. Graphs that split entropies. SIAM Journal on Discrete Mathematics, 1(1): 71–79, February 1988. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Koller:1994:CSS

[KM94] Daphne Koller and Nimrod Megiddo. Constructing small sample spaces satisfying given constraints. SIAM Journal on Discrete Mathematics, 7 (2):260–274, May 1994. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Klawe:1995:ULB

[KM95] Maria Klawe and Brendan Mumey. Upper and lower bounds on constructing alphabetic binary trees. SIAM Journal on Discrete Mathematics, 8(4):638–651, November 1995. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kushilevitz:1997:RPC

[KM97] Eyal Kushilevitz and Yishay Mansour. Randomness in private computations. SIAM Journal on Discrete Mathematics, 10(4):647–661, November 1997. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Korner:2001:CRI

[KM01] János Körner and Angelo Monti. Compact representations of the intersection structure of families of finite sets. SIAM Journal on Discrete Mathematics, 14(2):181–192, 2001. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/34834.

Klavzar:2002:PCC

[KM02] Sandi Klavzar and Henry Martyn Mulder. Partial cubes and crossing graphs. SIAM Journal on Discrete Mathematics, 15(2):235–251, 2002. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/38320.

Kubicki:2005:GTG

[KM05a] Grzegorz Kubicki and Michal Morayne. Graph-theoretic generalization of the secretary problem: The directed path case. SIAM Journal on Discrete Mathematics, 19(3):622–632, 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/44059.

Kushilevitz:2005:CNR

[KM05b] Eyal Kushilevitz and Yishay Mansour. Computation in noisy radio networks. SIAM Journal on Discrete Mathematics, 19(1):96–108, 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/43406.

Korner:2006:PCP

[KM06] János Körner and Claudia Malvenuto. Pairwise colliding permutations and the capacity of infinite graphs. SIAM Journal on Discrete Mathematics, 20(1):203–212, January 2006. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kudekar:2011:DCS

[KM11] Shrinivas Kudekar and Nicolas Macris. Decay of correlations for sparse graph error correcting codes. SIAM Journal on Discrete Mathematics, 25(2):956–988, ???? 2011. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:

//epubs.siam.org/sidma/
resource/1/sjdmec/v25/i2/
p956_s1.

Kakimura:2013:RIS

[KM13a] Naonori Kakimura and Kazuhisa Makino. Robust independence systems. SIAM Journal on Discrete Mathematics, 27(3):1257–1273, ???? 2013. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Korner:2013:DDG

[KM13b] János Körner and Irene Muzi. Degree-doubling graph families. SIAM Journal on Discrete Mathematics, 27(3): 1575–1583, ???? 2013. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kozik:2013:NCN

[KM13c] Jakub Kozik and Piotr Micek. Nonrepetitive choice number of trees. SIAM Journal on Discrete Mathematics, 27(1): 436–446, ???? 2013. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Krakovski:2014:HFW

[KM14] Roi Krakovski and Bojan Mohar. Homological face-width condition forcing K_{6}-minors in graphs on surfaces. SIAM Journal on Discrete Mathematics, 28(3):1257–1275, ???? 2014. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Karthick:2019:SFG

[KM19a] T. Karthick and Frédéric Maffray. Square-free graphs with no six-vertex induced path. SIAM Journal on Discrete Mathematics, 33(2): 874–909, ???? 2019. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kiwi:2019:SLC

[KM19b] Marcos Kiwi and Dieter Mitsche. On the second largest component of random hyperbolic graphs. SIAM Journal on Discrete Mathematics, 33(4):2200–2217, ???? 2019. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Karpenkov:2021:GCR

[KM21a] Oleg Karpenkov and Christian Mueller. Geometric criteria for realizability of tensegrities in higher dimensions. SIAM Journal on Discrete Mathematics, 35(2): 637–660, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kiefer:2021:NIM

[KM21b] Stefan Kiefer and Corto N.

Mascle. On nonnegative integer matrices and short killing words. SIAM Journal on Discrete Mathematics, 35(2): 1252–1267, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kiraly:2022:GRA

[KM22] Csaba Király and András Mihálykó. Globally rigid augmentation of rigid graphs. SIAM Journal on Discrete Mathematics, 36(4):2473– 2496, ???? 2022. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:/

/epubs.siam.org/doi/10.
1137/21M1432417.

Kang:2023:EER

[KM23] Mihyun Kang and Michael Missethan. The early evolution of the random graph process in planar graphs and related classes. SIAM Journal on Discrete Mathematics, 37(1):146–162, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1450616.

Kunisky:2024:SGL

[KM24] Dmitriy Kunisky and Cristopher Moore. The spectrum of the Grigoriev–Laurent pseudomoments. SIAM Journal on Discrete Mathematics, 38 (1):103–131, January 2024. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kwak:2005:EBC

[KML05] Jin Ho Kwak, Alexander Mednykh, and Valery Liskovets. Enumeration of branched coverings of nonorientable surfaces with cyclic branch points. SIAM Journal on Discrete Mathematics, 19(2):388–398, 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/42404.

Kang:2012:IMS

[KMM12] Ross J. Kang, Matthias Mnich, and Tobias Müller. Induced matchings in subcubic planar graphs. SIAM Journal on Discrete Mathematics, 26(3):1383–1411, ???? 2012. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kohayakawa:2021:CEC

[KMMS21] Yoshiharu Kohayakawa, Walner Mendonça, Guilherme Oliveira Mota, and Bjarne Schülke. Covering 3-edge-colored random graphs with monochromatic trees. SIAM Journal on Discrete Mathematics, 35(2): 1447–1459, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kronenberg:2023:GPV

[KMN23] Gal Kronenberg, Adva Mond, and Alon Naor. H-games played on vertex sets of random graphs. SIAM Journal on Discrete Mathematics, 37(2):864–916, ???? 2023.

CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/20M1375991.

Kavut:2018:SSP

[KMÖ18] Selçuk Kavut, Subhamoy Maitra, and Ferruh Özbudak. A super-set of Patterson– Wiedemann functions: Upper bounds and possible nonlinearities. SIAM Journal on Discrete Mathematics, 32(1): 106–122, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kim:2003:PBL

[KMP03] Jon-Lark Kim, Keith E. Mellinger, and Vera Pless. Projections of binary linear codes onto larger fields. SIAM Journal on Discrete Mathematics, 16(4):591–603, 2003. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/40436.

Kung:2014:MSB

[KMPR14] Joseph P. S. Kung, Dillon Mayhew, Irene Pivotto, and Gordon F. Royle. Maximum size binary matroids with no AG(3, 2)-minor are graphic. SIAM Journal on Discrete Mathematics, 28(3): 1559–1577, ???? 2014. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kushilevitz:1995:LUW

[KMR95] Eyal Kushilevitz, Yishay Mansour, and Michael O. Rabin. On lotteries with unique winners. SIAM Journal on Discrete Mathematics, 8(1): 93–98, February 1995. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kaiser:2011:CGF

[KMR11] Tomás Kaiser, Mickaël Montassier, and André Raspaud. Covering a graph by forests and a matching. SIAM Journal on Discrete Mathematics, 25(4): 1804–1811, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v25/i4/p1804_s1.

Kanesh:2023:FEC

[KMR^{+}23] Lawqueen Kanesh, Jayakrishnan Madathil, Sanjukta Roy, Abhishek Sahu, and Saket Saurabh. Further exploiting c-closure for FPT algorithms and kernels for domination problems. SIAM Journal on Discrete Mathematics, 37(4):2626–2669, November 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kneis:2009:BPS

[KMRR09] Joachim Kneis, Daniel Mölle, Stefan Richter, and Peter Rossmanith. A bound on the pathwidth of sparse graphs with applications to exact algorithms. SIAM Journal on Discrete Mathematics, 23(1): 407–427, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Komlos:1998:MNB

[KMS98] János Komlós, Yuan Ma, and Endre Szemerédi. Matching nuts and bolts in O(n log n) time. SIAM Journal on Discrete Mathematics, 11(3): 347–372, August 1998. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Korner:2008:GDP

[KMS08] János Körner, Claudia Malvenuto, and Gábor Simonyi. Graph-different permutations. SIAM Journal on Discrete Mathematics, 22(2): 489–499, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kierstead:2009:TCN

[KMŠ^{+}09] Hal Kierstead, Bojan Mohar, Simon Špacapan, Daqing Yang, and Xuding Zhu. The two-coloring number and degenerate colorings of planar graphs. SIAM Journal on Discrete Mathematics, 23(3): 1548–1560, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Korner:2012:FGD

[KMS12] János Körner, Silvia Messuti, and Gábor Simonyi. Families of graph-different Hamilton paths. SIAM Journal on Discrete Mathematics, 26(1): 321–329, ???? 2012. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Knierim:2023:GNS

[KMS23] Charlotte Knierim, Anders Martinsson, and Raphael Steiner. Hat guessing numbers of strongly degenerate graphs. SIAM Journal on Discrete Mathematics, 37 (2):1331–1347, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1479154.

Kobayashi:2007:OCF

[KMT07] Yusuke Kobayashi, Kazuo Murota, and Ken’ichiro Tanaka. Operations on Mconvex functions on jump systems. SIAM Journal on Discrete Mathematics, 21(1): 107–129, ???? 2007. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kostochka:2015:TPS

[KMV15] Alexandr Kostochka, Dhruv Mubayi, and Jacques Verstraëte. Turán problems and shadows III: Expansions of graphs. SIAM Journal on Discrete Mathematics, 29(2): 868–876, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Keevash:2006:SSN

[KMW06] Peter Keevash, Dhruv Mubayi, and Richard M. Wilson. Set systems with no singleton intersection. SIAM Journal on Discrete Mathematics, 20 (4):1031–1041, January 2006. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kelk:2024:CCA

[KMW24] Steven Kelk, Ruben Meuwese, and Stephan Wagner. Convex characters, algorithms, and matchings. SIAM Journal on Discrete Mathematics, 38(1): 380–411, January 2024. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Krishnamurti:1995:AAP

[KN95] Ramesh Krishnamurti and Bhagirath Narahari. An approximation algorithm for preemptive scheduling on parallel-task systems. SIAM Journal on Discrete Mathematics, 8(4):661–669, November 1995. CODEN SJDMEC.

ISSN 0895-4801 (print), 1095-7146 (electronic).

Klee:2013:FCB

[KN13] Steven Klee and Isabella Novik. From flag complexes to banner complexes. SIAM Journal on Discrete Mathematics, 27(2):1146–1158, ???? 2013. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kante:2016:PTA

[KN16] Mamadou Moustapha Kanté and Lhouari Nourine. Polynomial time algorithms for computing a minimum hull set in distance-hereditary and chordal graphs. SIAM Journal on Discrete Mathematics, 30(1):311–326, ???? 2016. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kiefer:2022:PWL

[KN22] Sandra Kiefer and Daniel Neuen. The power of the Weisfeiler–Leman algorithm to decompose graphs. SIAM Journal on Discrete Mathematics, 36(1):252–298, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL

https://epubs.siam.org/
doi/10.1137/20M1314987.

Khuller:1993:LSF

[KNK93] Samir Khuller, Joseph Naor, and Philip Klein. The lattice structure of flow in planar graphs. SIAM Journal on Discrete Mathematics, 6(3): 477–490, August 1993. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kardos:2023:CCS

[KNNW23] Frantisek Kardos, Jonathan Narboni, Reza Naserasr, and Zhouningxin Wang. Circular (4−ϵ)-coloring of some classes of signed graphs. SIAM Journal on Discrete Mathematics, 37(2):1198–1211, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1436257.

Kostochka:2005:ECD

[KNP05] A. V. Kostochka, K. Nakprasit, and S. V. Pemmaraju. On equitable coloring of ddegenerate graphs. SIAM Journal on Discrete Mathematics, 19(1):83–95, 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/43650.

Kannan:1992:IRG

[KNR92] Sampath Kannan, Moni Naor, and Steven Rudich. Implicit representation of graphs. SIAM Journal on Discrete Mathematics, 5(4): 596–603, November 1992. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Khanna:2005:DND

[KNS05] Sanjeev Khanna, Joseph (Seffi)

Naor, and F. Bruce Shepherd. Directed network design with orientation constraints. SIAM Journal on Discrete Mathematics, 19 (1):245–257, 2005. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/38011.

Knuth:1995:TWR

[Knu95] Donald E. Knuth. Two-way rounding. SIAM Journal on Discrete Mathematics, 8(2): 281–290, May 1995. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Krauthgamer:2014:PTD

[KNZ14] Robert Krauthgamer, Huy L.

Nguy˜ ên, and Tamar Zondiner. Preserving terminal distances using minors. SIAM Journal on Discrete Mathematics, 28 (1):127–141, ???? 2014. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ko:1988:STO

[Ko88] Ker-I Ko. Searching for two objects by underweight feedback. SIAM Journal on Discrete Mathematics, 1(1):65– 70, February 1988. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kuhn:2006:IBT

[KO06a] Daniela Kühn and Deryk Osthus. Improved bounds for topological cliques in graphs of large girth. SIAM Journal on Discrete Mathematics, 20 (1):62–78, January 2006. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kuhn:2006:MHC

[KO06b] Daniela Kühn and Deryk Osthus. Multicolored Hamilton cycles and perfect matchings in pseudorandom graphs. SIAM Journal on Discrete Mathematics, 20(2):273–286, January 2006. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kuhn:2012:PCR

[KO12] Daniela Kühn and Deryk Osthus. On Pósa’s Conjecture for random graphs. SIAM Journal on Discrete Mathematics, 26(3):1440–1457, ???? 2012. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kochol:1998:PIT

[Koc98] Martin Kochol. Partial intersection theorem and flows in abstract networks. SIAM Journal on Discrete Mathematics, 11(3):468–486, August 1998. CODEN SJD-MEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Koppe:2007:PBA

[Köp07] Matthias Köppe. A primal barvinok algorithm based on irrational decompositions. SIAM Journal on Discrete Mathematics, 21(1): 220–236, ???? 2007. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kushilevitz:2021:LUB

[KOP^{+}21] Eyal Kushilevitz, Rafail Ostrovsky, Emmanuel Prouff, Adi Rosén, Adrian Thillard, and Damien Vergnaud. Lower and upper bounds on the randomness complexity of private computations of AND. SIAM Journal on Discrete Mathematics, 35(1):465–484, ???? 2021. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kushilevitz:2003:ARP

[KOR03] Eyal Kushilevitz, Rafail Ostrovsky, and Adi Rosén. Amortizing randomness in private multiparty computations. SIAM Journal on Discrete Mathematics, 16 (4):533–544, 2003. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/35274.

Korandi:2018:RSG

[Kor18] Dániel Korándi. Rainbow saturation and graph capacities. SIAM Journal on Discrete Mathematics, 32(2): 1261–1264, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Khamsemanan:2016:BBU

[KOS16] Nirattaya Khamsemanan, Rafail Ostrovsky, and William E. Skeith. On the black-box use of somewhat homomorphic encryption in noninteractive two-party protocols. SIAM Journal on Discrete Mathematics, 30(1):266–295, ???? 2016. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kuhn:2009:OTT

[KOT09] Daniela Kühn, Deryk Osthus, and Andrew Treglown. An Ore-type theorem for perfect packings in graphs. SIAM Journal on Discrete Mathematics, 23(3):1335–1355, ???? 2009. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kotlar:2013:CSS

[Kot13] Daniel Kotlar. On circuits and serial symmetric basisexchange in matroids. SIAM Journal on Discrete Mathematics, 27(3):1274–1286, ????

2013. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kuhn:2016:RGF

[KOT16] Daniela Kühn, Deryk Osthus, and Amelia Taylor. On the random greedy F -free hypergraph process. SIAM Journal on Discrete Mathematics, 30(3):1343–1350, ???? 2016. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kortsarz:1995:AAM

[KP95] Guy Kortsarz and David Peleg. Approximation algorithms for minimum-time broadcast. SIAM Journal on Discrete Mathematics, 8(3): 401–427, August 1995. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kalyanasundaram:2000:OTP

[KP00] Bala Kalyanasundaram and Kirk R. Pruhs. The online transportation problem. SIAM Journal on Discrete Mathematics, 13 (3):370–383, 2000. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/34231.

Kowalski:2004:FDB

[KP04] Dariusz R. Kowalski and Andrzej Pelc. Faster deterministic broadcasting in ad hoc radio networks. SIAM Journal on Discrete Mathematics, 18(2):332–346, 2004. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/42464.

Kabadi:2006:CMT

[KP06] Santosh N. Kabadi and Abraham P. Punnen. On cost matrices with two and three distinct values of Hamiltonian paths and cycles. SIAM Journal on Discrete Mathematics, 20(4):977–998, January 2006. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Krotov:2009:AQO

[KP09] Denis S. Krotov and Vladimir N. Potapov. n-ary quasigroups of order 4. SIAM Journal on Discrete Mathematics, 23 (2):561–570, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kang:2016:CPG

[KP16] Ross J. Kang and François Pirot. Coloring powers and girth. SIAM Journal on Discrete Mathematics, 30(4): 1938–1949, ???? 2016. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Keszegh:2013:DPG

[KPP13] Balázs Keszegh, János Pach, and Dömötör Pálvölgyi. Drawing planar graphs of bounded degree with few slopes. SIAM Journal on Discrete Mathematics, 27(2): 1171–1183, ???? 2013. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kang:2015:PTQ

[KPPR15] Ross J. Kang, János Pach, Viresh Patel, and Guus Regts. A precise threshold for quasi-Ramsey numbers. SIAM Journal on Discrete Mathematics, 29(3): 1670–1682, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kratsch:2015:FPT

[KPPW15] Stefan Kratsch, Marcin Pilipczuk, MichaL Pilipczuk, and Magnus Wahlström. Fixed-parameter tractability of multicut in directed acyclic graphs. SIAM Journal on Discrete Mathematics, 29(1): 122–144, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kundgen:2010:RSM

[KPR10] André Kündgen, Michael J.

Pelsmajer, and Radhika Ramamurthi. k-robust singlemessage transmission. SIAM Journal on Discrete Mathematics, 24(1):301–321, ???? 2010. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Korandi:2016:RTP

[KPS16] Dániel Korándi, Yuval Peled, and Benny Sudakov. A random triadic process. SIAM Journal on Discrete Mathematics, 30(1):1–19, ???? 2016. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kafer:2019:CDS

[KPS19] Sean Kafer, Kanstantsin Pashkovich, and Laura Sanità. On the circuit diameter of some combinatorial polytopes. SIAM Journal on Discrete Mathematics, 33(1): 1–25, ???? 2019. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Khuller:2020:AON

[KPS20] Samir Khuller, Manish Purohit, and Kanthi K. Sarpatwar. Analyzing the optimal neighborhood: Algorithms for partial and budgeted connected dominating set problems. SIAM Journal on Discrete Mathematics, 34 (1):251–270, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kierstead:1994:LCR

[KPT94] H. A. Kierstead, S. G. Penrice, and W. T. Trotter. Online coloring and recursive graph theory. SIAM Journal on Discrete Mathematics, 7(1):72–89, February 1994. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kierstead:1995:LFF

[KPT95] Henry A. Kierstead, Stephen G.

Penrice, and William T. Trotter. On-line and first-fit coloring of graphs that do not induce P_{5}. SIAM Journal on Discrete Mathematics, 8 (4):485–498, November 1995. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kaibel:2012:SMS

[KPT12] Volker Kaibel, Kanstantsin Pashkovich, and Dirk Oliver Theis. Symmetry matters for sizes of extended formulations. SIAM Journal on Discrete Mathematics, 26(3): 1361–1382, ???? 2012. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Korandi:2020:LHS

[KPT20] Dániel Korándi, János Pach, and István Tomon. Large homogeneous submatrices. SIAM Journal on Discrete Mathematics, 34(4):2532– 2552, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kante:2022:LFP

[KPT22] Mamadou M. Kanté, Christophe Paul, and Dimitrios M. Thilikos. A linear fixed parameter tractable algorithm for connected pathwidth. SIAM Journal on Discrete Mathematics, 36(1):411–435, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL

                                        https://epubs.siam.org/
                                        doi/10.1137/20M1369361.

Knuth:1992:PCR

[KR92] Donald E. Knuth and Arvind Raghunathan. The problem of compatible representatives. SIAM Journal on Discrete Mathematics, 5(3): 422–427, August 1992. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Knopfmacher:1993:RSP

[KR93] A. Knopfmacher and J. N. Ridley. Reciprocal sums over partitions and compositions. SIAM Journal on Discrete Mathematics, 6(3): 388–399, August 1993. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kushilevitz:1998:RRT

[KR98] Eyal Kushilevitz and Adi Rosén. A randomness-rounds tradeoff in private computation. SIAM Journal on Discrete Mathematics, 11(1):61– 80, February 1998. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kindler:2004:DCR

[KR04] Guy Kindler and Dan Romik.

On distributions computable by random walks on graphs. SIAM Journal on Discrete Mathematics, 17(4): 624–633, 2004. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/43106.

Konyagin:2013:NSP

[KR13] Sergei V. Konyagin and Misha Rudnev. On new sum–product-type estimates. SIAM Journal on Discrete Mathematics, 27(2):973–990, ???? 2013. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kurauskas:2015:CIR

[KR15] Valentas Kurauskas and Katarzyna Rybarczyk. On the chromatic index of random uniform hypergraphs. SIAM Journal on Discrete Mathematics, 29(1):541–558, ???? 2015. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kaiser:2016:NZF

[KR16] Tomás Kaiser and Edita Rollová. Nowhere-zero flows in signed series-parallel graphs. SIAM Journal on Discrete Mathematics, 30(2): 1248–1258, ???? 2016. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Krauthgamer:2020:RVS

[KR20] Robert Krauthgamer and Havana Inbal Rika. Refined vertex sparsifiers of planar graphs. SIAM Journal on Discrete Mathematics, 34(1): 101–129, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kirsch:2023:MCB

[KR23] Rachel Kirsch and Jamie Radcliffe. Many cliques in bounded-degree hypergraphs. SIAM Journal on Discrete Mathematics, 37 (3):1436–1456, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1507565.

Kral:2004:CPC

[Krá04] Daniel Král’. Coloring powers of chordal graphs. SIAM Journal on Discrete Mathematics, 18(3):451–461, 2004. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/42407.

Kral:2006:CAP

[Krá06] Daniel Král’. The channel assignment problem with variable weights. SIAM Journal on Discrete Mathematics, 20(3):690–704, January 2006. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kravitz:2007:RSD

[Kra07] David Kravitz. RANDOM 2-SAT does not depend on a giant. SIAM Journal on Discrete Mathematics, 21(2): 408–422, ???? 2007. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kratsch:2018:RPK

[Kra18a] Stefan Kratsch. A randomized polynomial kernelization for vertex cover with a smaller parameter. SIAM Journal on Discrete Mathematics, 32(3): 1806–1839, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Krause:2018:DFM

[Kra18b] Ben Krause. Dimension-free L^{p}-maximal inequalities for spherical means in the hypercube. SIAM Journal on Discrete Mathematics, 32(4): 2493–2511, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Krivelevich:2010:EST

[Kri10] Michael Krivelevich. Embedding spanning trees in random graphs. SIAM Journal on Discrete Mathematics, 24(4):1495–1500, ???? 2010. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Krivelevich:2018:FUE

[Kri18] Michael Krivelevich. Finding and using expanders in locally sparse graphs. SIAM Journal on Discrete Mathematics, 32 (1):611–623, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kociumaka:2016:ERL

[KRR16] Tomasz Kociumaka, Jakub Radoszewski, and Wojciech Rytter. Efficient ranking of Lyndon words and decoding lexicographically minimal de Bruijn sequence. SIAM Journal on Discrete Mathematics, 30(4):2027–2046, ???? 2016. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kaiser:2011:GOC

[KRS11] Tomás Kaiser, Ondrej Rucký, and Riste Skrekovski. Graphs with odd cycle lengths 5 and 7 are 3-colorable. SIAM Journal on Discrete Mathematics, 25(3):1069–1088,

????  2011.  CODEN

SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v25/i3/p1069_s1.

Krivelevich:2015:SAC

[KRS15] Michael Krivelevich, Daniel Reichman, and Wojciech Samotij. Smoothed analysis on connected graphs. SIAM Journal on Discrete Mathematics, 29(3):1654–1669, ???? 2015. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kotlar:2021:SBE

[KRZ21] Dani Kotlar, Elad Roda, and Ran Ziv. On sequential basis exchange in matroids. SIAM Journal on Discrete Mathematics, 35(4):2517–2519, ???? 2021. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Korner:1988:SPS

[KS88] János Körner and Gábor Simonyi. Separating partition systems and locally different sequences. SIAM Journal on Discrete Mathematics, 1(3): 355–359, August 1988. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kalyanasundaram:1992:PCC

[KS92] Bala Kalyanasundaram and Georg Schnitger. The probabilistic communication complexity of set intersection. SIAM Journal on Discrete Mathematics, 5(4):545–557, November 1992. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kratsch:1993:DCG

[KS93] Dieter Kratsch and Lorna Stewart. Domination on cocomparability graphs. SIAM Journal on Discrete Mathematics, 6(3):400–417, August 1993. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Khuller:2000:CCP

[KS00] Samir Khuller and Yoram J. Sussmann. The capacitated K-center problem. SIAM Journal on Discrete Mathematics, 13(3):403–418, 2000. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/32977.

Kratsch:2002:ABM

[KS02] D. Kratsch and L. Stewart. Approximating bandwidth by mixing layouts of interval graphs. SIAM Journal on Discrete Mathematics, 15(4):435–449, 2002. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/35962.

Kashyap:2003:EAC

[KS03a] Navin Kashyap and Paul H.

Siegel. Equalities among capacities of (d, k)-constrained systems. SIAM Journal on Discrete Mathematics, 17 (2):276–297, 2003. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/41371.

Ko:2003:BDS

[KS03b] C. W. Ko and F. B. Shepherd. Bipartite domination and simultaneous matroid covers. SIAM Journal on Discrete Mathematics, 16(4):517–523, 2003. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28371.

Kral:2003:TAC

[KS03c] Daniel Král’ and Riste Skrekovski. A theorem about the channel assignment problem. SIAM Journal on Discrete Mathematics, 16(3):426–437, 2003. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/39944.

Kumar:2003:PFA

[KS03d] Ravi Kumar and D. Sivakumar. On polynomialfactor approximations to the shortest lattice vector length. SIAM Journal on Discrete Mathematics, 16(3):422–425, 2003. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/37998.

Katta:2004:NBT

[KS04] Akshay-Kumar Katta and Jay Sethuraman. A note on bandits with a twist. SIAM Journal on Discrete Mathematics, 18(1):110–113, 2004. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/43354.

Keevash:2005:SSR

[KS05] Peter Keevash and Benny Sudakov. Set systems with restricted cross-intersections and the minimum rank of inclusion matrices. SIAM Journal on Discrete Mathematics, 18(4):713–727, 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/43463.

Kovaleva:2006:AAR

[KS06] Sofia Kovaleva and Frits C. R. Spieksma. Approximation algorithms for rectangle stabbing and interval stabbing problems. SIAM Journal on Discrete Mathematics, 20(3):

748–768, January 2006. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kortsarz:2007:IAA

[KS07] Guy Kortsarz and Sunil Shende. An improved approximation of the achromatic number on bipartite graphs. SIAM Journal on Discrete Mathematics, 21(2): 361–373, ???? 2007. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kaiser:2008:CIE

[KŠ08a] Tomáš Kaiser and Riste Škrekovski. Cycles intersecting edge-cuts of prescribed sizes. SIAM Journal on Discrete Mathematics, 22(3): 861–874, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kral:2008:BRN

[KŠ08b] Daniel Král’ and Petr Škoda.

Bounds for the real number graph labellings and application to labellings of the triangular lattice. SIAM Journal on Discrete Mathematics, 22(4):1559–1569, ???? 2008. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kral:2008:CTF

[KS08c] Daniel Král’ and Matĕj StehlÍk. Coloring of trianglefree graphs on the double torus. SIAM Journal on Discrete Mathematics, 22(2): 541–553, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Krivelevich:2012:OPH

[KS12a] Michael Krivelevich and Wojciech Samotij. Optimal packings of Hamilton cycles in sparse random graphs. SIAM Journal on Discrete Mathematics, 26(3):964–982, ???? 2012. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Krivelevich:2012:CSS

[KS12b] Michael Krivelevich and Reto Spöhel. Creating small subgraphs in Achlioptas processes with growing parameter. SIAM Journal on Discrete Mathematics, 26(2): 670–686, ???? 2012. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kovacs:2019:NSP

[KS19] István Kovács and Daniel Soltész. On k-neighbor separated permutations. SIAM Journal on Discrete Mathematics, 33(3):1691–1711, ???? 2019. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kirova:2023:TCN

[KS23a] Valeriya Kirova and Arsenii Sagdeev. Two-colorings of normed spaces without long monochromatic unit arithmetic progressions. SIAM Journal on Discrete Mathematics, 37(2):718–732, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL

https://epubs.siam.org/
doi/10.1137/22M1483700.

Kummer:2023:MEE

[KS23b] Mario Kummer and Büsra Sert. Matroids on eight elements with the half-plane property and related concepts. SIAM Journal on Discrete Mathematics, 37 (3):2208–2227, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1490648.

Kawase:2019:SMU

[KSF19] Yasushi Kawase, Hanna Sumita, and Takuro Fukunaga. Submodular maximization with uncertain knapsack capacity. SIAM Journal on Discrete Mathematics, 33(3): 1121–1145, ???? 2019. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kowalik:2008:TCP

[KSŠ08] Łukasz Kowalik, Jean-Sébastien Sereni, and Riste Škrekovski.

Total-coloring of plane graphs with maximum degree nine. SIAM Journal on Discrete Mathematics, 22(4):1462– 1479, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kral:2009:NLB

[KSS09] Daniel Král’, Jean-Sébastien Sereni, and Michael Stiebitz. A new lower bound on the number of perfect matchings in cubic graphs. SIAM Journal on Discrete Mathematics, 23(3):1465–1483, ???? 2009. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kaiser:2011:RF

[KSS11a] Tomás Kaiser, Matej Stehlík, and Riste Skrekovski. On the 2-resonance of fullerenes. SIAM Journal on Discrete Mathematics, 25(4):1737– 1745, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v25/i4/p1737_s1.

Kang:2011:EPG

[KSS11b] Ross J. Kang, Jean-Sébastien Sereni, and Matej Stehlík. Every plane graph of maximum degree 8 has an edgeface 9-coloring. SIAM Journal on Discrete Mathematics, 25(2):514–533, ???? 2011. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:

//epubs.siam.org/sidma/
resource/1/sjdmec/v25/i2/
p514_s1.

Kral:2012:MMR

[KSS12] Daniel Král’, Jean-Sébastien Sereni, and Ladislav Stacho. Min-max relations for odd cycles in planar graphs. SIAM Journal on Discrete Mathematics, 26(3):884–895, ???? 2012. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kral:2006:CLG

[KST06] Daniel Král’, Riste Skrekovski, and Martin Tancer. Construction of large graphs with no optimal surjective L(2, 1)- labelings. SIAM Journal on Discrete Mathematics, 20(2): 536–543, January 2006. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kashyap:2005:ART

[KSV05] Navin Kashyap, Paul H.

Siegel, and Alexander Vardy. An application of Ramsey theory to coding for the optical channel. SIAM Journal on Discrete Mathematics, 19(4):921–937, 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/44585.

Kim:2017:WCW

[KSW17] Michael P. Kim, Warut Suksompong, and Virginia Vassilevska Williams. Who can win a single-elimination tournament? SIAM Journal on Discrete Mathematics, 31(3): 1751–1764, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kiss:2018:CEA

[KSY18] Sándor Z. Kiss, Csaba Sándor, and Quan-Hui Yang. On a conjecture of Erdős about sets without k pairwise coprime integers. SIAM Journal on Discrete Mathematics, 32(4):2453–2466, ???? 2018. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kao:1999:LDM

[KT99a] Ming-Yang Kao and Stephen R. Tate. On-line difference maximization. SIAM Journal on Discrete Mathematics, 12 (1):78–90, February 1999. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/30744.

Kratochvil:1999:RDG

[KT99b] Jan Kratochvíl and Zsolt Tuza. Rankings of directed graphs. SIAM Journal on Discrete Mathematics, 12 (3):374–384, August 1999.

CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/33024.

Kakimura:2014:MPD

[KT14] Naonori Kakimura and Mizuyo Takamatsu. Matching problems with delta-matroid constraints. SIAM Journal on Discrete Mathematics, 28(2): 942–961, ???? 2014. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kellner:2016:SSC

[KT16] Kai Kellner and Thorsten Theobald. Sum of squares certificates for containment of H-polytopes in V-polytopes. SIAM Journal on Discrete Mathematics, 30(2):763–776, ???? 2016. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kovacevic:2017:IBS

[KT17] Mladen Kovacević and Vincent Y. F. Tan. Improved bounds on Sidon sets via lattice packings of simplices. SIAM Journal on Discrete Mathematics, 31(3): 2269–2278, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

King:2019:NUB

[KT19] Emily J. King and Xiaoxian Tang. New upper bounds for equiangular lines by pillar decomposition. SIAM Journal on Discrete Mathematics, 33(4):2479–2508, ???? 2019. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Knauer:2024:CDC

[KT24] Kolja Knauer and William T. Trotter. Concepts of dimension for convex geometries. SIAM Journal on Discrete Mathematics, 38(2): 1566–1585, May 2024. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kunisky:2023:DUM

[Kun23] Dmitriy Kunisky. The discrepancy of unsatisfiable matrices and a lower bound for the Komlós conjecture constant. SIAM Journal on Discrete Mathematics, 37 (2):586–603, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M149819X.

Kushilevitz:1992:PCC

[Kus92] Eyal Kushilevitz. Privacy and communication complexity. SIAM Journal on Discrete Mathematics, 5(2): 273–284, May 1992. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kleitman:1990:CBI

[KV90] Daniel J. Kleitman and Rakesh V. Vohra. Computing the bandwidth of interval graphs. SIAM Journal on Discrete Mathematics, 3(3): 373–375, August 1990. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kavitha:2015:SSP

[KV15] Telikepalli Kavitha and Nithin M.

Varma. Small stretch pairwise spanners and approximate D-preservers. SIAM Journal on Discrete Mathematics, 29(4):2239–2254, ???? 2015. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kelk:2012:CKQ

[KvIL^{+}12] Steven Kelk, Leo van Iersel, Nela Lekić, Simone Linz, Celine Scornavacca, and Leen Stougie. Cycle Killer… Qu’estce que c’est? On the comparative approximability of hybridization number and directed feedback vertex set. SIAM Journal on Discrete Mathematics, 26(4): 1635–1656, ???? 2012. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Korman:2023:KGV

[KvRRS23] Matias Korman, André van Renssen, Marcel Roeloffzen, and Frank Staals. Kinetic geodesic Voronoi diagrams in a simple polygon. SIAM Journal on Discrete Mathematics, 37(4):2276–2311, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL

                                        https://epubs.siam.org/
                                        doi/10.1137/20M1384804.

Kanellopoulos:2023:DTH

[KVZ23] Panagiotis Kanellopoulos, Alexandros A. Voudouris, and Rongsen Zhang. On discrete truthful heterogeneous two-facility location. SIAM Journal on Discrete Mathematics, 37(2):779–799, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL

                                        https://epubs.siam.org/
                                        doi/10.1137/22M149908X.

Karchmer:1990:MCC

[KW90] Mauricio Karchmer and Avi Wigderson. Monotone circuits for connectivity require super-logarithmic depth. SIAM Journal on Discrete Mathematics, 3(2): 255–265, May 1990. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kleitman:1991:STM

[KW91] Daniel J. Kleitman and Douglas B. West. Spanning trees with many leaves. SIAM Journal on Discrete Mathematics, 4(1):99–106, February 1991. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kannan:1992:TCG

[KW92] Sampath K. Kannan and Tandy J. Warnow. Triangulating 3-colored graphs. SIAM Journal on Discrete Mathematics, 5(2):249–258, May 1992. CODEN SJD-MEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kratzke:1996:TIN

[KW96] Thomas M. Kratzke and Douglas B. West. The total interval number of a graph. II. trees and complexity. SIAM Journal on Discrete Mathematics, 9(2):339–348, May 1996. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kaul:2008:LLS

[KW08a] Hemanshu Kaul and Douglas B. West. Long local searches for maximal bipartite subgraphs. SIAM Journal on Discrete Mathematics, 22(3):1138–1144, ???? 2008. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kirk:2008:LNS

[KW08b] Russell Kirk and Hua Wang.

Largest number of subtrees of trees with a given maximum degree. SIAM Journal on Discrete Mathematics, 22 (3):985–995, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kemkes:2013:IUB

[KW13] Graeme Kemkes and Nicholas Wormald. An improved upper bound on the length of the longest cycle of a supercritical random graph. SIAM Journal on Discrete Mathematics, 27(1):342–362, ???? 2013. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kolesnik:2014:LBI

[KW14] Brett Kolesnik and Nick Wormald. Lower bounds for the isoperimetric numbers of random regular graphs. SIAM Journal on Discrete Mathematics, 28(1): 553–575, ???? 2014. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kenyon:2017:SCP

[KW17] Richard W. Kenyon and David B. Wilson. The space of circular planar electrical networks. SIAM Journal on Discrete Mathematics, 31 (1):1–28, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kinnersley:2013:EPG

[KWZ13] William B. Kinnersley, Douglas B. West, and Reza Zamani. Extremal problems for game domination number. SIAM Journal on Discrete Mathematics, 27(4): 2090–2107, ???? 2013. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kobayashi:2012:AFM

[KY12] Yusuke Kobayashi and Yuichi Yoshida. Algorithms for finding a maximum non-k-linked graph. SIAM Journal on Discrete Mathematics, 26(2): 591–604, ???? 2012. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Keevash:2018:RMP

[KY18] Peter Keevash and Liana Yepremyan. Rainbow matchings in properly colored multigraphs. SIAM Journal on Discrete Mathematics, 32 (3):1577–1584, ???? 2018. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kim:2021:GSF

[KY21] Jang Soo Kim and Meesue Yoo. Generalized Schur function determinants using the Bazin identity. SIAM Journal on Discrete Mathematics, 35(3):1650–1672, ???? 2021. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kabadi:2009:IEN

[KYDN09] S. N. Kabadi, J. Yan, D. Du, and K. P. K. Nair. Integer exact network synthesis problem. SIAM Journal on Discrete Mathematics, 23(1): 136–154, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kierstead:2004:RTT

[KZ04] H. A. Kierstead and Yingxian Zhu. Radius three trees in graphs with large chromatic number. SIAM Journal on Discrete Mathematics, 17(4):571–581, 2004. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/33986.

Kostochka:2008:ALC

[KZ08] A. V. Kostochka and Xuding Zhu. Adapted list coloring of graphs and hypergraphs. SIAM Journal on Discrete Mathematics, 22(1): 398–408, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Kupavskii:2018:SMS

[KZ18] Andrey Kupavskii and Maksim Zhukovskii. Short monadic second order sentences about sparse random graphs. SIAM Journal on Discrete Mathematics, 32(4): 2916–2940, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Labarre:2013:LBE

[Lab13] Anthony Labarre. Lower bounding edit distances between permutations. SIAM Journal on Discrete Mathematics, 27(3):1410–1428, ???? 2013. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Lagarias:2000:WSL

[Lag00] Jeffrey C. Lagarias. Wellspaced labelings of points in rectangular grids. SIAM Journal on Discrete Mathematics, 13(4):521–534, 2000. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/36224.

Laihonen:2005:OER

[Lai05] Tero Laihonen. On optimal edge-robust and vertexrobust (1, ≤ l)-identifying codes. SIAM Journal on Discrete Mathematics, 18 (4):825–834, 2005. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/44075.

Lai:2008:OD

[Lai08] Hong-Jian Lai. mod(2p + 1)-orientations and K_{1,2p+1}- decompositions. SIAM Journal on Discrete Mathematics, 21(4):844–850, ???? 2008. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Lai:2018:LTH

[Lai18] Tri Lai. Lozenge tilings of a halved hexagon with an array of triangles removed from the boundary. SIAM Journal on Discrete Mathematics, 32 (1):783–814, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Lampis:2020:FTB

[Lam20] Michael Lampis. Finer tight bounds for coloring on cliquewidth. SIAM Journal on Discrete Mathematics, 34(3): 1538–1558, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Lato:2023:SMB

[Lat23] Sabrina Lato. A spectral Moore bound for bipartite semiregular graphs. SIAM Journal on Discrete Mathematics, 37(1):315–331, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL

                                        https://epubs.siam.org/
                                        doi/10.1137/21M1450082.

Lavrov:2016:UBH

[Lav16] Mikhail Lavrov. An upper bound for the Hales–Jewett number HJ(4, 2). SIAM Journal on Discrete Mathematics, 30(2):1333–1342, ???? 2016. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Lal:2009:BSB

[LB09] A. K. Lal and B. Bhattacharjya. Breaking the symmetries of the Book Graph and the Generalized Petersen Graph. SIAM Journal on Discrete Mathematics, 23(3): 1200–1216, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Lau:2004:RPP

[LC04] Lap Chi Lau and Derek G.

Corneil. Recognizing powers of proper interval, split, and chordal graphs. SIAM Journal on Discrete Mathematics, 18(1):83–102, 2004. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/42593.

Liu:2012:RDC

[LC12] Chun-Hung Liu and Gerard J. Chang. Roman domination on 2-connected graphs. SIAM Journal on Discrete Mathematics, 26(1): 193–205, ???? 2012. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Luo:2003:DCG

[LCV03] Yuan Luo, Wende Chen, and A. J. Han Vinck. The determination of the chain good weight hierarchies with high dimension. SIAM Journal on Discrete Mathematics, 17(2):196–209, 2003. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/41668.

Lucchesi:2018:TUP

[LdCKM18] Cláudio L. Lucchesi, Marcelo H. de Carvalho, Nishad Kothari, and U. S. R. Murty. On two unsolved problems concerning matching covered graphs. SIAM Journal on Discrete Mathematics, 32(2): 1478–1504, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Leclerc:1990:MMS

[Lec90]  Bruno Leclerc.  Medians

and majorities in semimodular lattices. SIAM Journal on Discrete Mathematics, 3 (2):266–276, May 1990. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Lee:2017:CLS

[Lee17] James R. Lee. Covering the large spectrum and generalized Riesz products. SIAM Journal on Discrete Mathematics, 31(1):562–572, ???? 2017. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Leighton:1994:LBT

[Lei94] Tom Leighton. A 2d−1 lower bound for two-layer knockknee channel routing. SIAM Journal on Discrete Mathematics, 7(2):230–237, May 1994. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Lenart:1998:GDG

[Len98] Cristian Lenart. A generalized distance in graphs and centered partitions. SIAM Journal on Discrete Mathematics, 11(2):293–304, May 1998. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Letzter:2019:MCG

[Let19] Shoham Letzter. Many Hcopies in graphs with a forbidden tree. SIAM Journal on Discrete Mathematics, 33(4): 2360–2368, ???? 2019. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Levin:2009:AUS

[Lev09] Asaf Levin. Approximating the unweighted k-set cover problem: Greedy meets local search. SIAM Journal on Discrete Mathematics, 23(1): 251–264, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Lev:2015:EIP

[Lev15] Vsevolod F. Lev. Edgeisoperimetric problem for Cayley graphs and generalized Takagi functions. SIAM Journal on Discrete Mathematics, 29(4):2389–2411, ???? 2015. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Lev:2022:SDG

[Lev22] Vsevolod F. Lev. Small doubling in groups with moderate torsion. SIAM Journal on Discrete Mathematics, 36(1):315–335, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1397799.

Lam:2011:GSE

[LGS11] Fumei Lam, Dan Gusfield, and Srinath Sridhar. Generalizing the splits equivalence theorem and four gamete condition: Perfect phylogeny on three-state characters. SIAM Journal on Discrete Mathematics, 25(3): 1144–1175, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v25/i3/p1144_s1.

Lu:1990:DNP

[LHC90] Tung Lin Lu, Pei Hsin Ho, and Gerard J. Chang. The domatic number problem in interval graphs. SIAM Journal on Discrete Mathematics, 3(4):531–536, November 1990. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Li:2018:MOB

[LHHL18] Jiaao Li, Xinmin Hou, Miaomiao Han, and Hong-Jian Lai. Mod (2p + 1)-orientation on bipartite graphs and complementary graphs. SIAM Journal on Discrete Mathematics, 32(1):29–43, ???? 2018. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Liu:2023:CDC

[LHLZ23] Siyan Liu, Rong-Xia Hao, Rong Luo, and Cun-Quan Zhang. 5-cycle double covers, 4-flows, and catlin reduction. SIAM Journal on Discrete Mathematics, 37 (1):253–267, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1472425.

Li:2017:MDS

[Li17] Shuxing Li. The minimum distance of some narrow-sense primitive BCH codes. SIAM Journal on Discrete Mathematics, 31(4):2530–2569, ???? 2017. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Lichtner:1998:IAG

[Lic98]  Jonathan Lichtner.  Iter-

ating an α-ary gray code. SIAM Journal on Discrete Mathematics, 11(3):381–386, August 1998. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Lichiardopol:2014:PCH

[Lic14] Nicolas Lichiardopol. Proof of a conjecture of Henning and Yeo on vertex-disjoint directed cycles. SIAM Journal on Discrete Mathematics, 28(3):1618–1627, ???? 2014. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Lind:1989:PSF

[Lin89] D. A. Lind. Perturbations of shifts of finite type. SIAM Journal on Discrete Mathematics, 2(3):350–365, August 1989. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Linusson:1997:PRB

[Lin97] Svante Linusson. Partitions with restricted block sizes, Möbius functions, and the kof-each problem. SIAM Journal on Discrete Mathematics, 10(1):18–29, February 1997. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Lee:2015:NCG

[LiO15] Choongbum Lee and Sang il Oum. Number of cliques in graphs with a forbidden subdivision. SIAM Journal on Discrete Mathematics, 29(4): 1999–2005, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Lee:2023:ICV

[LiO23] Duksang Lee and Sang il Oum. Intertwining connectivities for vertex-minors and pivot-minors. SIAM Journal on Discrete Mathematics, 37(1):304–314, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1485073.

Liu:2014:UBF

[Liu14] Chun-Hung Liu. An upper bound on the fractional chromatic number of triangle-free subcubic graphs. SIAM Journal on Discrete Mathematics, 28(3):1102–1136, ???? 2014. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Liu:2024:GLS

[Liu24] Lele Liu. Graph limits and spectral extremal problems for graphs. SIAM Journal on Discrete Mathematics, 38(1): 590–608, January 2024. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Laihonen:1999:NBC

[LL99] Tero Laihonen and Simon Litsyn. New bounds on covering radius as a function of dual distance. SIAM Journal on Discrete Mathematics, 12(2):243–251, May 1999. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/33170.

Lee:2014:RPR

[LL14a] Cheng-Wei Lee and Hsueh-I Lu. Replacement paths via row minima of concise matrices. SIAM Journal on Discrete Mathematics, 28(1): 206–225, ???? 2014. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Li:2014:MOG

[LL14b] Ping Li and Hong-Jian Lai. On mod (2s + 1)-orientations of graphs. SIAM Journal on Discrete Mathematics, 28(4): 1820–1827, ???? 2014. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Lin:2015:FLF

[LL15] Qizhong Lin and Yusheng Li. A Folkman linear family. SIAM Journal on Discrete Mathematics, 29(4): 1988–1998, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Liang:2017:MCS

[LL17] Hung-Chun Liang and Hsueh-I Lu. Minimum cuts and shortest cycles in directed planar graphs via noncrossing shortest paths. SIAM Journal on Discrete Mathematics, 31 (1):454–476, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Lin:2021:RNI

[LL21] Qizhong Lin and Xiudi Liu.

Ramsey numbers involving large books. SIAM Journal on Discrete Mathematics, 35 (1):23–34, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Lladser:2006:UFC

[Lla06] Manuel Lladser. Uniform formulae for coefficients of meromorphic functions in two variables. Part I. SIAM Journal on Discrete Mathematics, 20(4):811–828, January 2006. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Li:2017:GCS

[LLL17] Jiaao Li, Hong-Jian Lai, and Rong Luo. Group connectivity, strongly Z_{m}-connectivity, and edge disjoint spanning trees. SIAM Journal on Discrete Mathematics, 31(3): 1909–1922, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Lee:2018:CAS

[LLL18] Kyungyong Lee, Li Li, and Nicholas A. Loehr. A combinatorial approach to the symmetry of q, t-catalan numbers. SIAM Journal on Discrete Mathematics, 32(1): 191–232, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Lavrov:2019:DUG

[LLM19] Mikhail Lavrov, Po-Shen Loh, and Arnau Messegué. Distance-uniform graphs with large diameter. SIAM Journal on Discrete Mathematics, 33(2):994–1005, ???? 2019. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Lin:2004:ICV

[LLS04] Ching-Chi Lin, Hsueh-I Lu, and I-Fan Sun. Improved compact visibility representation of planar graph via Schnyder’s realizer. SIAM Journal on Discrete Mathematics, 18(1):19–29, 2004.

CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/42074.

Lee:2013:SSG

[LLS13] Choongbum Lee, Po-Shen Loh, and Benny Sudakov. Self-similarity of graphs. SIAM Journal on Discrete Mathematics, 27(2):959–972, ???? 2013. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Lu:2020:FSG

[LLS^{+}20] You Lu, Rong Luo, Michael Schubert, Eckhard Steffen, and Cun-Quan Zhang. Flows on signed graphs without long barbells. SIAM Journal on Discrete Mathematics, 34(4): 2166–2182, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Li:2022:FIR

[LLW22] Jiaao Li, Xueliang Li, and Meiling Wang. The flow index of regular class I graphs. SIAM Journal on Discrete Mathematics, 36 (3):1991–2003, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1393169.

Li:2010:RNC

[LLY10] Wei-Tian Li, Linyuan Lu, and Yiting Yang. Routing numbers of cycles, complete bipartite graphs, and hypercubes. SIAM Journal on Discrete Mathematics, 24(4): 1482–1494, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Lih:1999:SEC

[LLZ99] Ko-Wei Lih, Daphne Der-Fen Liu, and Xuding Zhu. Star extremal circulant graphs. SIAM Journal on Discrete Mathematics, 12(4):491–499, November 1999. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/34283.

Lv:2022:PMH

[LLZ22] Zequn Lv, Mei Lu, and Yi Zhang. Perfect matching and Hamilton tight cycle decomposition of complete nbalanced r-partite k-uniform hypergraphs. SIAM Journal on Discrete Mathematics, 36(1):241–251, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/20M1365557.

Liu:1989:LRA

[LM89] Joseph W. H. Liu and Andranik Mirzaian. A linear reordering algorithm for parallel pivoting of chordal graphs. SIAM Journal on Discrete Mathematics, 2(1): 100–107, February 1989. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Leveque:2008:CBF

[LM08] Benjamin Lévêque and Frédéric Maffray. Coloring bullfree perfectly contractile graphs. SIAM Journal on Discrete Mathematics, 21(4): 999–1018, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Lukotka:2010:CGG

[LM10] Robert Lukot’ka and Ján Mazák. Cubic graphs with given circular chromatic index. SIAM Journal on Discrete Mathematics, 24(3): 1091–1103, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Lucier:2011:GDC

[LM11] B. Lucier and M. Molloy.

The Glauber dynamics for colorings of bounded degree trees. SIAM Journal on Discrete Mathematics, 25(2): 827–853, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v25/i2/p827_s1.

Levit:2012:VBA

[LM12] Vadim E. Levit and Eugen Mandrescu. Vertices belonging to all critical sets of a graph. SIAM Journal on Discrete Mathematics, 26(1): 399–403, ???? 2012. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Lo:2014:DTD

[LM14] Allan Lo and Klas Markström. ℓ-degree Turán density. SIAM Journal on Discrete Mathematics, 28(3): 1214–1225, ???? 2014. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Lindzey:2016:LTA

[LM16] Nathan Lindzey and Ross M. McConnell. Linear-time algorithms for finding Tucker submatrices and Lekkerkerker– Boland subgraphs. SIAM Journal on Discrete Mathematics, 30(1):43–69, ???? 2016. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Li:2017:PDD

[LM17] Zhentao Li and Bojan Mohar. Planar digraphs of digirth four are 2-colorable. SIAM Journal on Discrete Mathematics, 31(3):2201–2205, ???? 2017. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Lai:2021:NGC

[LM21] Hong-Jian Lai and Lucian Mazza. A note on group colorings and group structure. SIAM Journal on Discrete Mathematics, 35(4): 2535–2543, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Lason:2022:NSC

[LM22a] Michał Lasoń and Mateusz Michałek. A note on seminormality of cut polytopes. SIAM Journal on Discrete Mathematics, 36 (1):114–117, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M138586X.

Lim:2022:CSO

[LM22b] Tongseok Lim and Robert J.

McCann. On the cardinality of sets in R^{d} obeying a slightly obtuse angle bound. SIAM Journal on Discrete Mathematics, 36 (2):1093–1101, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1403163.

Long:2021:SLP

[LMM21] James A. Long, Jr., Kevin G.

Milans, and Andrea Munaro. Sublinear longest path transversals. SIAM Journal on Discrete Mathematics, 35 (3):1673–1677, ???? 2021. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Lee:2010:MNS

[LMNS10] Jon Lee, Vahab S. Mirrokni, Viswanath Nagarajan, and Maxim Sviridenko. Maximizing nonmonotone submodular functions under matroid or knapsack constraints. SIAM Journal on Discrete Mathematics, 23(4):2053–2078, ???? 2010. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Liu:2019:CHO

[LMP19] Shiping Liu, Florentin Münch, and Norbert Peyerimhoff. Curvature and higher order Buser inequalities for the graph connection Laplacian. SIAM Journal on Discrete Mathematics, 33(1): 257–305, ???? 2019. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Lokshtanov:2017:HSO

[LMRS17] Daniel Lokshtanov, Pranabendu Misra, M. S. Ramanujan, and Saket Saurabh. Hitting selected (odd) cycles. SIAM Journal on Discrete Mathematics, 31(3): 1581–1615, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Liu:2019:GTP

[LMS19] Ricky I. Liu, Karola Mészáros, and Avery St. Dizier. Gelfand– Tsetlin polytopes: a story of flow and order polytopes. SIAM Journal on Discrete Mathematics, 33(4): 2394–2415, ???? 2019. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Lokshtanov:2019:PCF

[LMSZ19] Daniel Lokshtanov, Amer E.

Mouawad, Saket Saurabh, and Meirav Zehavi. Packing cycles faster than Erdős– Posa. SIAM Journal on Discrete Mathematics, 33(3): 1194–1215, ???? 2019. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Lu:2023:ARN

[LMW23] Linyuan Lu, Andrew Meier, and Zhiyu Wang. Anti-Ramsey number of edgedisjoint rainbow spanning trees in all graphs. SIAM Journal on Discrete Mathematics, 37(2):1162–1172, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL

https://epubs.siam.org/
doi/10.1137/21M1428121.

Labbe:2017:BEV

[LN17] Jean-Philippe Labbé and Eran Nevo. Bounds for entries of γ-vectors of flag homology spheres. SIAM Journal on Discrete Mathematics, 31(3):2064–2078, ???? 2017. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Liu:2021:PTR

[LN21] Wenjian Liu and Ning Ning. Phase transition of the reconstructability of a general model with different in-community and outcommunity mutations on an infinite tree. SIAM Journal on Discrete Mathematics, 35 (2):1381–1417, ???? 2021. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Li:2024:GBH

[LN24] Lily Li and Aleksandar Nikolov. On the gap between hereditary discrepancy and the determinant lower bound. SIAM Journal on Discrete Mathematics, 38(2): 1222–1238, April 2024. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Lovasz:1995:SPD

[LNNW95] László Lovász, Moni Naor, Ilan Newman, and Avi Wigderson. Search problems in the decision tree model. SIAM Journal on Discrete Mathematics, 8(1): 119–132, February 1995. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Luby:1996:TBD

[LNO96] Michael G. Luby, Joseph Seffi Naor, and Ariel Orda. Tight bounds for dynamic storage allocation. SIAM Journal on Discrete Mathematics, 9(1): 155–166, February 1996. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ling:2005:IAC

[LÖ05] San Ling and Ferruh Özbudak.

Improved p-ary codes and sequence families from Galois rings of characteristic p^{2}. SIAM Journal on Discrete Mathematics, 19(4): 1011–1028, January 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Lo:2014:ECV

[Lo14] Allan Lo. An edge-colored version of Dirac’s theorem. SIAM Journal on Discrete Mathematics, 28(1): 18–36, ???? 2014. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Loehr:2010:APS

[Loe10] Nicholas A. Loehr. Abacus proofs of Schur function identities. SIAM Journal on Discrete Mathematics, 24(4): 1356–1370, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Long:2021:SIG

[Lon21] Philip M. Long. Superlinear integrality gaps for the minimum majority problem. SIAM Journal on Discrete Mathematics, 35(4): 3004–3016, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Louidor:2010:TFC

[Lou10] Erez Louidor. The tradeoff function for a class of RLL(d, k) constraints. SIAM Journal on Discrete Mathematics, 24(4):1371–1398, ???? 2010. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Lee:2010:ANO

[LOW10] Jon Lee, Shmuel Onn, and Robert Weismantel. Approximate nonlinear optimization over weighted independence systems. SIAM Journal on Discrete Mathematics, 23(4): 1667–1681, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Liestman:1988:BNB

[LP88] Arthur L. Liestman and Joseph G. Peters. Broadcast networks of bounded degree. SIAM Journal on Discrete Mathematics, 1(4):531– 540, November 1988. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Luczak:2019:PHR

[LP19] Tomasz Luczak and Joanna Polcyn. Paths in hypergraphs: a rescaling phenomenon. SIAM Journal on Discrete Mathematics, 33(4): 2251–2266, ???? 2019. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Lidicky:2021:SPR

[LP21a] Bernard Lidický and Florian Pfender. Semidefinite programming and Ramsey numbers. SIAM Journal on Discrete Mathematics, 35(4): 2328–2344, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Lin:2021:LBC

[LP21b] Qizhong Lin and Xing Peng.

Large book-cycle Ramsey numbers. SIAM Journal on Discrete Mathematics, 35(1): 532–545, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Li:2023:RST

[LP23] Yongtao Li and Yuejian Peng. Refinement on spectral Turán’s theorem. SIAM Journal on Discrete Mathematics, 37(4):2462–2485, October 2023. CODEN SJDMEC.

ISSN 0895-4801 (print), 1095-7146 (electronic).

Lichiardopol:2009:STB

[LPS09] Nicolas Lichiardopol, Attila Pór, and Jean-Sébastien Sereni. A step toward the Bermond–Thomassen conjecture about disjoint cycles in digraphs. SIAM Journal on Discrete Mathematics, 23(2): 979–992, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Lokshtanov:2018:BAS

[LPS18] Daniel Lokshtanov, Michał Pilipczuk, and Saket Saurabh. Below all subsets for minimal connected dominating set. SIAM Journal on Discrete Mathematics, 32(3): 2332–2345, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Lotker:2012:RLB

[LPSR12] Zvi Lotker, Boaz Patt-Shamir, and Dror Rawitz. Rent, lease, or buy: Randomized algorithms for multislope ski rental. SIAM Journal on Discrete Mathematics, 26(2): 718–736, ???? 2012. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

LeSaulnier:2013:TAG

[LPW^{+}13] Timothy D. LeSaulnier, Noah Prince, Paul S. Wenger, Douglas B. West, and Pratik Worah. Total acquisition in graphs. SIAM Journal on Discrete Mathematics, 27(4): 1800–1819, ???? 2013. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Lo:2022:HCC

[LQ22] On-Hei Solomon Lo and Jianguo Qian. Hamiltonian cycles in 4-connected planar and projective planar triangulations with few 4separators. SIAM Journal on Discrete Mathematics, 36 (2):1496–1501, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1419556.

Luczak:1991:TMG

[ŁR91] Tomasz Łuczak and Andrzej Ruciński. Tree-matchings in graph processes. SIAM Journal on Discrete Mathematics, 4(1):107–120, February 1991. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Leader:1994:LOI

[LR94] I. Leader and A. J. Radcliffe. Littlewood–Offord inequalities for random variables. SIAM Journal on Discrete Mathematics, 7(1):90– 101, February 1994. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Lozin:2004:BTC

[LR04] V. Lozin and D. Rautenbach. On the band-, tree-, and clique-width of graphs with bounded vertex degree. SIAM Journal on Discrete Mathematics, 18(1):195–206, 2004. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/41975.

Landman:2007:AMS

[LR07] Bruce M. Landman and Aaron Robertson. Avoiding monochromatic sequences with special gaps. SIAM Journal on Discrete Mathematics, 21(3):794–801, ???? 2007. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Li:2011:ISP

[LRN11] Liangpan Li and Oliver Roche-Newton. An improved sum-product estimate for general finite fields. SIAM Journal on Discrete Mathematics, 25(3):1285– 1296, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sidma/resource/
                                        1/sjdmec/v25/i3/p1285_s1.

Levi:2014:TSM

[LRR14] Reut Levi, Dana Ron, and Ronitt Rubinfeld. Testing similar means. SIAM Journal on Discrete Mathematics, 28(4):1699–1724, ???? 2014. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Leader:2021:IPV

[LRR21] Imre Leader, Zarko Randelović, and Eero Räty. Inequalities on projected volumes. SIAM Journal on Discrete Mathematics, 35(3): 1678–1687, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Liu:2021:CKL

[LRS^{+}21] Runrun Liu, Martin Rolek, D. Christopher Stephens, Dong Ye, and Gexin Yu. Connectivity for kite-linked graphs. SIAM Journal on Discrete Mathematics, 35(1): 431–446, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Linusson:2023:GCD

[LRS23] Svante Linusson, Petter Restadh, and Liam Solus. Greedy causal discovery is geometric. SIAM Journal on Discrete Mathematics, 37 (1):233–252, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1457205.

Letchford:2008:OMC

[LRT08] Adam N. Letchford, Gerhard Reinelt, and Dirk Oliver Theis. Odd minimum cut sets and b-matchings revisited. SIAM Journal on Discrete Mathematics, 22(4): 1480–1487, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Lin:2011:SBB

[LRTW11] Henry Lin, Tim Roughgarden, Éva Tardos, and Asher Walkover. Stronger bounds on Braess’s paradox and the maximum latency of selfish routing. SIAM Journal on Discrete Mathematics, 25(4): 1667–1686, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v25/i4/p1667_s1.

Li:2012:PHS

[LRWZ12] Binlong Li, Zdenek Ryjácek, Ying Wang, and Shenggui Zhang. Pairs of heavy subgraphs for Hamiltonicity of 2-connected graphs. SIAM Journal on Discrete Mathematics, 26(3):1088–1103, ???? 2012. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Lawler:1989:DRB

[LS89] Gregory F. Lawler and John Sylvester. Determining resistances from boundary measurements in finite networks. SIAM Journal on Discrete Mathematics, 2(2): 231–239, May 1989. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Liestman:1995:DCN

[LS95a] Arthur L. Liestman and Thomas C. Shermer. Degreeconstrained network spanners with nonconstant delay. SIAM Journal on Discrete Mathematics, 8(2): 291–321, May 1995. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Lin:1995:ASR

[LS95b] Yaw Ling Lin and Steven S.

Skiena. Algorithms for square roots of graphs. SIAM Journal on Discrete Mathematics, 8(1):99–118, February 1995. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Lewenstein:2003:AAM

[LS03a] Moshe Lewenstein and Maxim Sviridenko. A 5/8 approximation algorithm for the maximum asymmetric TSP. SIAM Journal on Discrete Mathematics, 17(2):237–248, 2003. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/40286.

Linusson:2003:CCG

[LS03b] Svante Linusson and John Shareshian. Complexes of t-colorable graphs. SIAM Journal on Discrete Mathematics, 16(3):371–389, 2003. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/36696.

Luz:2005:CQC

[LS05] Carlos J. Luz and Alexander Schrijver. A convex quadratic characterization of the Lovász theta number. SIAM Journal on Discrete Mathematics, 19(2):382–387, 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/42918.

Little:2006:TSC

[LS06] John Little and Hal Schenck. Toric surface codes and Minkowski sums. SIAM Journal on Discrete Mathematics, 20(4):999–1014, January 2006. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Lin:2008:UCA

[LS08] Min Chih Lin and Jayme L. Szwarcfiter. Unit circulararc graph representations and feasible circulations. SIAM Journal on Discrete Mathematics, 22(1):409–423, ???? 2008. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Litsyn:2009:DBF

[LS09] Simon Litsyn and Alexander Shpunt. On the distribution of Boolean function nonlinearity. SIAM Journal on Discrete Mathematics, 23(1): 79–95, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Long:2014:WHH

[LS14] Philip M. Long and Rocco A.

Servedio. On the weight of halfspaces over Hamming balls. SIAM Journal on Discrete Mathematics, 28(3): 1035–1061, ???? 2014. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Lafond:2015:HCG

[LS15] Manuel Lafond and Ben Seamone. Hamiltonian chordal graphs are not cycle extendable. SIAM Journal on Discrete Mathematics, 29(2): 877–887, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Lund:2016:IBB

[LS16] Ben Lund and Shubhangi Saraf. Incidence bounds for block designs. SIAM Journal on Discrete Mathematics, 30(4):1997–2010, ???? 2016. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Laurent:2017:SFS

[LS17] Monique Laurent and Matteo Seminaroti. Similarityfirst search: a new algorithm with application to robinsonian matrix recognition. SIAM Journal on Discrete Mathematics, 31(3): 1765–1800, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Larose:2018:NPR

[LS18] Benoit Larose and Mark Siggers. NU polymorphisms on reflexive digraphs. SIAM Journal on Discrete Mathematics, 32(1):728–749, ???? 2018. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Lavollee:2022:BNE

[LS22] Jérémy Lavollée and Konrad J. Swanepoel. Bounding the number of edges of matchstick graphs. SIAM Journal on Discrete Mathematics, 36(1):777–785, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1441134.

Li:2024:STV

[LS24] Anqi Li and Lisa Sauermann. Sárközy’s theorem in various finite field settings. SIAM Journal on Discrete Mathematics, 38(2): 1409–1416, April 2024. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Lou:1992:OAM

[LSL92] R. D. Lou, M. Sarrafzadeh, and D. T. Lee. An optimal algorithm for the maximum two-chain problem. SIAM Journal on Discrete Mathematics, 5(2):285–304, May 1992. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Liu:2003:LEW

[LSO03] Xiangwei Liu, Gilbert Strang, and Susan Ott. Localized eigenvectors from widely spaced matrix modifications. SIAM Journal on Discrete Mathematics, 16 (3):479–498, 2003. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/40904.

Lang:2017:APE

[LSS17] Richard Lang, Oliver Schaudt, and Maya Stein. Almost partitioning a 3-edge-colored K_{n,n} into five monochromatic cycles. SIAM Journal on Discrete Mathematics, 31(2): 1374–1402, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Los:2024:PFB

[LSS24] Dimitrios Los, Thomas Sauerwald, and John Sylvester. The power of filling in balanced allocations. SIAM Journal on Discrete Mathematics, 38(1):529–565, January 2024. CODEN SJD-MEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Luo:2010:ECS

[LSSY10] Rong Luo, Jean-Sébastien Sereni, D. Christopher Stephens, and Gexin Yu. Equitable coloring of sparse planar graphs. SIAM Journal on Discrete Mathematics, 24(4): 1572–1583, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v24/i4/p1572_s1.

Lokshtanov:2019:BJB

[LSSZ19] Daniel Lokshtanov, Saket Saurabh, Roohani Sharma, and Meirav Zehavi. Balanced judicious bipartition is fixedparameter tractable. SIAM Journal on Discrete Mathematics, 33(4):1878–1911, ????

2019. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Lin:2017:CTS

[LSTY17] Bo Lin, Bernd Sturmfels, Xiaoxian Tang, and Ruriko Yoshida. Convexity in tree spaces. SIAM Journal on Discrete Mathematics, 31(3): 2015–2038, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Lund:2018:FFK

[LSW18] Ben Lund, Shubhangi Saraf, and Charles Wolf. Finite field Kakeya and Nikodym sets in three dimensions. SIAM Journal on Discrete Mathematics, 32(4):2836–2849, ???? 2018. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Li:2014:SCC

[LSX14] Cai Heng Li, Shaohui Sun, and Jing Xu. Self-complementary circulants of prime-power order. SIAM Journal on Discrete Mathematics, 28(1): 8–17, ???? 2014. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Lu:2001:NIA

[LT01] Chi-Jen Lu and Shi-Chun Tsai. A note on iterating an α-ary Gray code. SIAM Journal on Discrete Mathematics, 14(2):237–239, 2001.

CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/36768.

Loten:2011:NUP

[LT11] Cynthia Loten and Claude Tardif. Near-unanimity polymorphisms on structures with finite duality. SIAM Journal on Discrete Mathematics, 25 (2):490–496, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v25/i2/p490_s1.

Laird:2020:RHG

[LTBL20] Lucas Laird, Richard C. Tillquist, Stephen Becker, and Manuel E. Lladser. Resolvability of Hamming graphs. SIAM Journal on Discrete Mathematics, 34(4): 2063–2081, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Lu:2004:DHC

[Lu04] Chi-Jen Lu. Deterministic hypergraph coloring and its applications. SIAM Journal on Discrete Mathematics, 18(2):320–331, 2004. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/36766.

Lu:2008:ECS

[Lu08] Linyuan Lu. Explicit construction of small Folkman graphs. SIAM Journal on Discrete Mathematics, 21(4): 1053–1060, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Lu:2010:ICR

[Lu10] Hsueh-I Lu. Improved compact routing tables for planar networks via orderly spanning trees. SIAM Journal on Discrete Mathematics, 23(4): 2079–2092, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Lubiw:1990:BBP

[Lub90a] Anna Lubiw. The Boolean basis problem and how to cover some polygons by rectangles. SIAM Journal on Discrete Mathematics, 3(1):98– 115, February 1990. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Lubotzky:1990:LSG

[Lub90b] Alexander Lubotzky. Locally symmetric graphs of prescribed girth and Coxeter groups. SIAM Journal on Discrete Mathematics, 3(2): 277–280, May 1990. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Luczak:1998:GAE

[Luc98] Tomasz Luczak. A greedy algorithm estimating the height of random trees. SIAM Journal on Discrete Mathematics, 11(2):318–329, May 1998. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Lucena:2003:NLB

[Luc03]  Brian Lucena.  A new

lower bound for tree-width using maximum cardinality search. SIAM Journal on Discrete Mathematics, 16(3):345–353, 2003. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/39777.

Lukotka:2020:SCC

[Luk20] Robert Lukotka. Short cycle covers of cubic graphs and intersecting 5-circuits. SIAM Journal on Discrete Mathematics, 34(1):188–211, ???? 2020. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Luo:2023:CCM

[Luo23]  Sammy Luo.  On con-

nected components with many edges. SIAM Journal on Discrete Mathematics, 37(2):573–585, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1471365.

Lutz:2004:SEN

[Lut04] Frank H. Lutz. Small examples of nonconstructible simplicial balls and spheres. SIAM Journal on Discrete Mathematics, 18(1): 103–109, 2004. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/43052.

Lefmann:1989:P

[LV89] H. Lefmann and B. Voigt.

Partitions of A^{ω}. SIAM Journal on Discrete Mathematics, 2(4):524–529, November 1989. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Langenhop:1988:CON

[LW88a] Carl E. Langenhop and William E. Wright. Certain occupancy numbers via an algorithm for computing their ratios. SIAM Journal on Discrete Mathematics, 1(3): 360–371, August 1988. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Lempel:1988:SCN

[LW88b] Abraham Lempel and Marcelo J.

Weinberger. Self-complementary normal bases in finite fields. SIAM Journal on Discrete Mathematics, 1(2):193–198, May 1988. CODEN SJD-MEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Li:2003:STC

[LW03] Xueliang Li and Lusheng Wang. Solutions for two conjectures on the inverse problem of the Wiener index of peptoids. SIAM Journal on Discrete Mathematics, 17(2):210–218, 2003. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/38726.

Leizhen:2010:SRG

[LW10] Cai Leizhen and Wang Weifan. The surviving rate of a graph for the firefighter problem. SIAM Journal on Discrete Mathematics, 23(4): 1814–1826, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Li:2017:LTA

[LW17a] Peng Li and Yaokun Wu. A linear time algorithm for the 1-fixed-endpoint path cover problem on interval graphs. SIAM Journal on Discrete Mathematics, 31(1): 210–239, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Lu:2017:TTC

[LW17b] Hongliang Lu and David G. L. Wang. A Tutte-type characterization for graph factors. SIAM Journal on Discrete Mathematics, 31(2): 1149–1159, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Lu:2018:SRN

[LW18] Linyuan Lu and Zhiyu Wang.

On the size-Ramsey number of tight paths. SIAM Journal on Discrete Mathematics, 32(3):2172–2179, ???? 2018. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Lu:2020:ARN

[LW20] Linyuan Lu and Zhiyu Wang.

Anti-Ramsey number of edgedisjoint rainbow spanning trees. SIAM Journal on Discrete Mathematics, 34(4): 2346–2362, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Liu:2010:RGH

[LWW10] Zihui Liu, Jie Wang, and Xin-Wen Wu. On the relative generalized Hamming weights of linear codes and their subcodes. SIAM Journal on Discrete Mathematics, 24(4): 1234–1241, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Lu:2013:EGF

[LWY13] Hongliang Lu, David G. L. Wang, and Qinglin Yu. On the existence of general factors in regular graphs. SIAM Journal on Discrete Mathematics, 27(4):1862–1869, ???? 2013. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Lu:2018:APM

[LWY18] Hongliang Lu, Yan Wang, and Xingxing Yu. Almost perfect matchings in k-partite k-graphs. SIAM Journal on Discrete Mathematics, 32(1): 522–533, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Lu:2022:RPM

[LWY22] Hongliang Lu, Yan Wang, and Xingxing Yu. Rainbow perfect matchings for 4uniform hypergraphs. SIAM Journal on Discrete Mathematics, 36(3):1645–1662, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL

                                        https://epubs.siam.org/
                                        doi/10.1137/21M1442383.

Luo:2008:RDS

[LXZZ08] Rong Luo, Rui Xu, Wenan Zang, and Cun-Quan Zhang. Realizing degree sequences with graphs having nowherezero 3-flows. SIAM Journal on Discrete Mathematics, 22 (2):500–519, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Lu:2010:LBT

[LY10] Linyuan Lu and Yiting Yang.

A lower bound on the transposition diameter. SIAM Journal on Discrete Mathematics, 24(4):1242–1249, ???? 2010. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Lin:2018:TFW

[LY18a] Bo Lin and Ruriko Yoshida.

Tropical Fermat–Weber points. SIAM Journal on Discrete Mathematics, 32(2):1229– 1245, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Lu:2018:RMH

[LY18b] Hongliang Lu and Xingxing Yu. On rainbow matchings for hypergraphs. SIAM Journal on Discrete Mathematics, 32(1):382–393, ???? 2018. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Liu:2021:NHC

[LY21] Xiaonan Liu and Xingxing Yu. Number of Hamiltonian cycles in planar triangulations. SIAM Journal on Discrete Mathematics, 35(2): 1005–1021, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Liu:2023:TNG

[LY23] Xiao-Chuan Liu and Xu Yang. On the Turán number of generalized theta graphs. SIAM Journal on Discrete Mathematics, 37(2):1237–1251, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL

                                        https://epubs.siam.org/
                                        doi/10.1137/21M1408439.

Lynch:1994:RRA

[Lyn94] James F. Lynch. Random resource allocation graphs and the probability of deadlock. SIAM Journal on Discrete Mathematics, 7(3):458– 473, August 1994. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Lu:2021:NPM

[LYY21] Hongliang Lu, Xingxing Yu, and Xiaofan Yuan. Nearly perfect matchings in uniform hypergraphs. SIAM Journal on Discrete Mathematics, 35 (2):1022–1049, ???? 2021. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Larose:2003:CEP

[LZ03] Benoit Larose and László Zádori. The complexity of the extendibility problem for finite posets. SIAM Journal on Discrete Mathematics, 17(1):

114–121, January 2003. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Liu:2005:CDT

[LZ05a] Daphne Der-Fen Liu and Xuding Zhu. Circular distance two labeling and the λ-number for outerplanar graphs. SIAM Journal on Discrete Mathematics, 19 (2):281–293, 2005. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/41429.

Liu:2005:MDL

[LZ05b] Daphne Der-Fen Liu and Xuding Zhu. Multilevel distance labelings for paths and cycles. SIAM Journal on Discrete Mathematics, 19(3):610–621, 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/41776.

Li:2006:DMF

[LZ06] Yusheng Li and Wenan Zang.

Differential methods for finding independent sets in hypergraphs. SIAM Journal on Discrete Mathematics, 20(1): 96–104, January 2006. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Lu:2009:ERH

[LZ09] Linyuan Lu and Yi Zhao. An exact result for hypergraphs and upper bounds for the Turán density of K_{r+1}^{r}. SIAM Journal on Discrete Mathematics, 23(3):1324–1334, ???? 2009. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Lo:2018:CTD

[LZ18a] Allan Lo and Yi Zhao. Codegree Turán density of complete r-uniform hypergraphs. SIAM Journal on Discrete Mathematics, 32(2): 1154–1158, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Lu:2018:SCG

[LZ18b] Changhong Lu and Ping Zhang. K_{5}^{−}-subdivision in 4-connected graphs. SIAM Journal on Discrete Mathematics, 32(4):2900–2915, ???? 2018. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Lafond:2022:CTD

[LZZ22] Manuel Lafond, Binhai Zhu, and Peng Zou. Computing the tandem duplication distance is NP-Hard. SIAM Journal on Discrete Mathematics, 36(1):64–91, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL

https://epubs.siam.org/
doi/10.1137/20M1356257.

MacGillivray:1991:CCV

[Mac91]  Gary MacGillivray.  On

the complexity of colouring by vertex-transitive and arctransitive digraphs. SIAM Journal on Discrete Mathematics, 4(3):397–408, August 1991. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Macajova:2013:BCG

[Mác13] Edita Mácajová. Bridgeless cubic graphs are (7, 2)-edge-choosable. SIAM Journal on Discrete Mathematics, 27(2): 757–767, ???? 2013. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Macourt:2018:IRB

[Mac18] Simon Macourt. Incidence results and bounds of trilinear and quadrilinear exponential sums. SIAM Journal on Discrete Mathematics, 32(2): 815–825, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Makai:2007:MCF

[Mak07] Márton Makai. On maximum cost K_{t,t}-free t-matchings of bipartite graphs. SIAM Journal on Discrete Mathematics, 21(2):349–360, ???? 2007. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Mallows:1989:ISI

[Mal89] C. L. Mallows. Identities satisfied by iterated polynomials and (Q, x)-binomial coefficients. SIAM Journal on Discrete Mathematics, 2(3): 366–370, August 1989. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Malandro:2015:FIF

[Mal15] Martin E. Malandro. Fourier inversion for finite inverse semigroups. SIAM Journal on Discrete Mathematics, 29 (1):269–296, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Marietti:2009:CCC

[Mar09] Mario Marietti. A combinatorial characterization of Coxeter groups. SIAM Journal on Discrete Mathematics, 23 (1):319–332, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Mareche:2020:CGK

[Mar20]  Laure Marêché.  Combi-

natorics for general kinetically constrained spin models. SIAM Journal on Discrete Mathematics, 34(1): 370–384, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Montejano:2022:SDM

[MAR22] Luis Montejano, Jorge L. Ramírez Alfonsín, and Ivan Rasskin. Self-dual maps I: Antipodality. SIAM Journal on Discrete Mathematics, 36 (3):1551–1566, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1367076.

Montejano:2023:SDM

[MAR23] Luis Montejano, Jorge L. Ramírez Alfonsín, and Iván Rasskin. Self-dual maps II: Links and symmetry. SIAM Journal on Discrete Mathematics, 37 (1):191–220, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1426341.

Mattheus:2019:TPF

[Mat19] Sam Mattheus. Trace of products in finite fields from a combinatorial point of view. SIAM Journal on Discrete Mathematics, 33(4): 2126–2139, ???? 2019. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Mayoraz:1996:PDN

[May96] E. N. Mayoraz. On the power of democratic networks. SIAM Journal on Discrete Mathematics, 9(2): 258–268, May 1996. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Mayhew:2008:MCN

[May08] Dillon Mayhew. Matroid complexity and nonsuccinct descriptions. SIAM Journal on Discrete Mathematics, 22 (2):455–466, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Mammoliti:2018:MCC

[MB18] Adam Mammoliti and Thomas Britz. On Mubayi’s conjecture and conditionally intersecting sets. SIAM Journal on Discrete Mathematics, 32 (3):2361–2380, ???? 2018. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

McCormick:1993:WSP

[MC93] S. Thomas McCormick and S. Frank Chang. The weighted sparsity problem: Complexity and algorithms. SIAM Journal on Discrete Mathematics, 6(1):57–69, February 1993. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Milans:2006:CGP

[MC06] Kevin Milans and Bryan Clark. The complexity of graph pebbling. SIAM Journal on Discrete Mathematics, 20(3):769–798, January 2006. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Mertzios:2012:SPA

[MC12] George B. Mertzios and Derek G. Corneil. A simple polynomial algorithm for the longest path problem on cocomparability graphs. SIAM Journal on Discrete Mathematics, 26(3):940–963, ???? 2012. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

McCarty:2021:DIS

[McC21] Rose McCarty. Dense induced subgraphs of dense bipartite graphs. SIAM Journal on Discrete Mathematics, 35(2): 661–667, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

McDonald:2015:LVR

[McD15] Daniel C. McDonald. On-line vertex ranking of trees. SIAM Journal on Discrete Mathematics, 29(1):145–156, ???? 2015. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

McKinley:2019:SCD

[McK19]  Gweneth McKinley.  Su-

perlogarithmic cliques in dense inhomogeneous random graphs. SIAM Journal on Discrete Mathematics, 33(3): 1772–1800, ???? 2019. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

McLeod:2010:AEE

[McL10] Jeanette C. McLeod. Asymptotic enumeration of k-edge-colored k-regular graphs. SIAM Journal on Discrete Mathematics, 23(4):2178– 2197, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

MoranR:2011:MFC

[MD11] Diego A. Morán R. and Santanu S. Dey. On maximal S-free convex sets. SIAM Journal on Discrete Mathematics, 25(1):379–393, ???? 2011. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sidma/resource/
                                        1/sjdmec/v25/i1/p379_s1.

Moran:2016:CIH

[MD16] Diego A. Morán and Santanu S. Dey. Closedness of integer hulls of simple conic sets. SIAM Journal on Discrete Mathematics, 30 (1):70–99, ???? 2016. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Mohammadi:2016:AMT

[MdCW16] Fatemeh Mohammadi, Eduardo Sáenz de Cabezón, and Henry P. Wynn. The algebraic method in tree percolation. SIAM Journal on Discrete Mathematics, 30(2): 1193–1212, ???? 2016. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Mehrabian:2012:DNR

[Meh12] Abbas Mehrabian. On the density of nearly regular graphs with a good edgelabeling. SIAM Journal on Discrete Mathematics, 26(3): 1265–1268, ???? 2012. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Merris:1999:NUB

[Mer99] Russell Merris. Note: An upper bound for the diameter of a graph. SIAM Journal on Discrete Mathematics, 12(3):412, August 1999. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28247.

Mertzios:2015:RST

[Mer15] George B. Mertzios. The recognition of simple-triangle graphs and of linear-interval orders is polynomial. SIAM Journal on Discrete Mathematics, 29(3):1150–1185, ????

2015. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Meszaros:2016:PRT

[Més16a] Karola Mészáros. h-polynomials of reduction trees. SIAM Journal on Discrete Mathematics, 30(2):736–762, ???? 2016. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Meszaros:2016:PDC

[Més16b] Karola Mészáros. Pipe dream complexes and triangulations of root polytopes belong together. SIAM Journal on Discrete Mathematics, 30(1): 100–111, ???? 2016. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ma:2019:OBL

[MG19] Jingxue Ma and Gennian Ge. Optimal binary linear locally repairable codes with disjoint repair groups. SIAM Journal on Discrete Mathematics, 33(4):2509–2529, ???? 2019. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

McKilliam:2014:FCP

[MGC14] Robby G. McKilliam, Alex Grant, and I. Vaughan L. Clarkson. Finding a closest point in a lattice of Voronoi’s first kind. SIAM Journal on Discrete Mathematics, 28(3):

1405–1422, ???? 2014. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

McClosky:2009:CPP

[MH09] Benjamin McClosky and Illya V. Hicks. The co-2-plex polytope and integral systems. SIAM Journal on Discrete Mathematics, 23(3): 1135–1148, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

McKinley:1991:DCR

[MHLHL91] P. K. McKinley, N. Hasan, R. Libeskind-Hadas, and C. L. Liu. Disjoint covers in replicated heterogeneous arrays. SIAM Journal on Discrete Mathematics, 4(2): 281–292, May 1991. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Michini:2021:TCR

[Mic21] Carla Michini. Tight cycle relaxations for the cut polytope. SIAM Journal on Discrete Mathematics, 35(4): 2908–2921, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Mitchell:1997:BSP

[Mit97] Chris J. Mitchell. de Bruijn sequences and perfect factors. SIAM Journal on Discrete Mathematics, 10(2):

270–281, May 1997. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Makino:2001:TRN

[MK01] Kazuhisa Makino and Tiko Kameda. Transformations on regular nondominated coteries and their applications. SIAM Journal on Discrete Mathematics, 14 (3):381–407, 2001. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/37111.

Manada:2009:CSP

[MK09] Akiko Manada and Navin Kashyap. A comparative study of periods in a periodicfinite-type shift. SIAM Journal on Discrete Mathematics, 23(3):1507–1524, ???? 2009. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Martinez:2023:MZS

[MLMR23] F. E. Brochero Martínez, A. Lemos, B. K. Moriya, and S. Ribas. The main zerosum constants over D_{2n} × C_{2}. SIAM Journal on Discrete Mathematics, 37 (3):1496–1508, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1496013.

Munoz:2011:EPR

[MLS11] Xavier Muñoz, Zhentao Li, and Ignasi Sau. Edgepartitioning regular graphs for ring traffic grooming with a priori placement of the ADMs. SIAM Journal on Discrete Mathematics, 25(4): 1490–1505, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v25/i4/p1490_s1.

Machi:1993:GEC

[MM93] Antonio Machì and Filippo Mignosi. Garden of Eden configurations for cellular automata on Cayley graphs of groups. SIAM Journal on Discrete Mathematics, 6(1): 44–56, February 1993. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Manacher:1996:FDP

[MM96] Glenn K. Manacher and Terrance A. Mankus. Finding a domatic partition of an interval graph in time O(n). SIAM Journal on Discrete Mathematics, 9(2):167–172, May 1996. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Magos:2011:FSA

[MM11] D. Magos and I. Mourtos. On the facial structure of the alldifferent system. SIAM Journal on Discrete Mathematics, 25(1): 130–158, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v25/i1/p130_s1.

Maffray:2012:CFG

[MM12] Frédéric Maffray and Grégory Morel. On 3-colorable P_{5}- free graphs. SIAM Journal on Discrete Mathematics, 26(4): 1682–1708, ???? 2012. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Montufar:2015:WDM

[MM15] Guido F. Montúfar and Jason Morton. When does a mixture of products contain a product of mixtures? SIAM Journal on Discrete Mathematics, 29(1):321–347, ???? 2015. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Mutoh:2003:EGB

[MMJF03] Yukiyasu Mutoh, Toshio Morihara, Masakazu Jimbo, and Hung-Lin Fu. The existence of 2 × 4 gridblock designs and their applications. SIAM Journal on Discrete Mathematics, 16(2):173–178, 2003. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/38736.

Matuschke:2020:RFW

[MMO20] Jannik Matuschke, S. Thomas McCormick, and Gianpaolo Oriolo. Rerouting flows when links fail. SIAM Journal on Discrete Mathematics, 34(4): 2082–2107, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Mehrabian:2013:MDG

[MMP13] Abbas Mehrabian, Dieter Mitsche, and Pawel Pralat. On the maximum density of graphs with unique-path labelings. SIAM Journal on Discrete Mathematics, 27(3): 1228–1233, ???? 2013. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

MacRury:2023:PTD

[MMPPG23] Calum MacRury, Tomás Masarík, Leilani Pai, and Xavier Pérez-Giménez. The phase transition of discrepancy in random hypergraphs. SIAM Journal on Discrete Mathematics, 37 (3):1818–1841, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1451427.

Martinez:2010:LEU

[MMPS10] Conrado Martínez, Lucia Moura, Daniel Panario, and Brett Stevens. Locating errors using ELAs, covering arrays, and adaptive testing algorithms. SIAM Journal on Discrete Mathematics, 23(4): 1776–1799, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Martinez-Moro:2006:MCF

[MMR06] E. Martínez-Moro and I. F. Rúa. Multivariable codes over finite chain rings: Serial codes. SIAM Journal on Discrete Mathematics, 20(4): 947–959, January 2006. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Martin:2015:CSC

[MMS15] Barnaby Martin, Florent Madelaine, and Juraj Stacho. Constraint satisfaction with counting quantifiers. SIAM Journal on Discrete Mathematics, 29(2):1065–1113, ???? 2015. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Martin:2017:AMK

[MMS17] Ryan R. Martin, Richard Mycroft, and Jozef Skokan. An asymptotic multipartite Kühn–Osthus theorem. SIAM Journal on Discrete Mathematics, 31(3):1498– 1513, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Momihara:2008:CWC

[MMSJ08] Koji Momihara, Meinard Müller, Junya Satoh, and Masakazu Jimbo. Constant weight conflict-avoiding codes. SIAM Journal on Discrete Mathematics, 21(4): 959–979, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ma:2023:PDP

[MMSW23] Yulai Ma, Davide Mattiolo, Eckhard Steffen, and Isaak H. Wolf. Pairwise disjoint perfect matchings in r-edge-connected r-regular graphs. SIAM Journal on Discrete Mathematics, 37 (3):1548–1565, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1500654.

McConnell:2015:LTR

[MN15] Ross M. McConnell and Yahav Nussbaum. Linear-time recognition of probe interval graphs. SIAM Journal on Discrete Mathematics, 29(4): 2006–2046, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ma:2018:CGT

[MN18] Jie Ma and Bo Ning. Coloring graphs with two odd cycle lengths. SIAM Journal on Discrete Mathematics, 32 (1):296–319, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Mertzios:2018:LTA

[MNN18] George B. Mertzios, André Nichterlein, and Rolf Niedermeier. A linear-time algorithm for maximumcardinality matching on cocomparability graphs. SIAM Journal on Discrete Mathematics, 32(4):2820–2835, ???? 2018. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Motwani:2008:LBL

[MNP08] Rajeev Motwani, Assaf Naor, and Rina Panigrahy. Lower bounds on locality sensitive hashing. SIAM Journal on Discrete Mathematics, 21(4): 930–935, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Mouawad:2017:SRP

[MNPR17] Amer E. Mouawad, Naomi Nishimura, Vinayak Pathak, and Venkatesh Raman. Shortest reconfiguration paths in the solution space of Boolean formulas. SIAM Journal on Discrete Mathematics, 31(3): 2185–2200, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Mousset:2014:NGL

[MNS14] Frank Mousset, Rajko Nenadov, and Angelika Steger. On the number of graphs without large cliques. SIAM Journal on Discrete Mathematics, 28(4):1980–1986, ???? 2014. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Matoya:2022:PPP

[MO22] Kazuki Matoya and Taihei Oki. Pfaffian pairs and parities: Counting on linear matroid intersection and parity problems. SIAM Journal on Discrete Mathematics, 36(3):2121–2158, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1421751.

Mohar:1999:LTA

[Moh99]  Bojan Mohar.  A linear

time algorithm for embedding graphs in an arbitrary surface. SIAM Journal on Discrete Mathematics, 12(1):6–26, February 1999. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29248.

Moisio:2008:DBH

[Moi08] Marko Moisio. On the duals of binary hyper-Kloosterman codes. SIAM Journal on Discrete Mathematics, 22(1):

273–287, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Mollard:2011:VPH

[Mol11] Michel Mollard. On vertex partitions of hypercubes by isometric trees. SIAM Journal on Discrete Mathematics, 25(2):534–538, ???? 2011. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sidma/resource/
                                        1/sjdmec/v25/i2/p534_s1.

Momihara:2013:SHD

[Mom13]  Koji Momihara.  Skew

Hadamard difference sets from cyclotomic strongly regular graphs. SIAM Journal on Discrete Mathematics, 27 (2):1112–1122, ???? 2013. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Montgomery:2015:AAF

[Mon15] Richard Montgomery. Almost all friendly matrices have many obstructions. SIAM Journal on Discrete Mathematics, 29(1):421–430, ???? 2015. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Morgenstern:1994:RD

[Mor94] Moshe Morgenstern. Ramanujan diagrams. SIAM Journal on Discrete Mathematics, 7(4):560–570, November 1994. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Moreira:2021:RGC

[Mor21] Luiz Moreira. Ramsey goodness of clique versus paths in random graphs. SIAM Journal on Discrete Mathematics, 35(3):2210–2222, ???? 2021. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Mota:2019:TCB

[Mot19] G. O. Mota. Three-color bipartite Ramsey number for graphs with small bandwidth. SIAM Journal on Discrete Mathematics, 33(1): 197–208, ???? 2019. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Mouaha:1994:SCO

[Mou94] C. Mouaha. On systematic codes obtained from linear codes over GF(q^{2}). SIAM Journal on Discrete Mathematics, 7(3):337–343, August 1994. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Malitz:1994:ARP

[MP94] Seth Malitz and Achilleas Papakostas. On the angular resolution of planar graphs. SIAM Journal on Discrete Mathematics, 7(2):

172–183, May 1994. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

McMorris:1995:MPF

[MP95] F. R. McMorris and R. C. Powers. The median procedure in a formal theory of consensus. SIAM Journal on Discrete Mathematics, 8 (4):507–516, November 1995. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Mitchell:1998:PFC

[MP98] Chris J. Mitchell and Kenneth G. Paterson. Perfect factors from cyclic codes and interleaving. SIAM Journal on Discrete Mathematics, 11 (2):241–264, May 1998. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Mustafa:2004:LYN

[MP04] Nabil H. Mustafa and Aleksandar Pekec. Listen to your neighbors: How (not) to reach a consensus. SIAM Journal on Discrete Mathematics, 17(4):634–660, 2004. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/40821.

Matousek:2008:LMC

[MP08] JiřÍ Matoušek and Aleš PřÍvĕtivý. Large monochromatic components in twocolored grids. SIAM Journal on Discrete Mathematics, 22 (1):295–311, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Marcus:2013:CBE

[MP13] Brian Marcus and Ronnie Pavlov. Computing bounds for entropy of stationary Z^{d} Markov random fields. SIAM Journal on Discrete Mathematics, 27(3):1544–1558, ???? 2013. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Majerski:2014:ISD

[MP14] Piotr Majerski and Jakub Przybylo. On the irregularity strength of dense graphs. SIAM Journal on Discrete Mathematics, 28(1): 197–205, ???? 2014. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Meier:2017:PFS

[MP17a] Florian Meier and Ueli Peter.

Push is fast on sparse random graphs. SIAM Journal on Discrete Mathematics, 31 (1):29–49, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Mitsche:2017:TRP

[MP17b] Dieter Mitsche and Guillem Perarnau. On treewidth and related parameters of random geometric graphs. SIAM Journal on Discrete Mathematics, 31(2):1328–1354, ???? 2017. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Mixon:2021:GOS

[MP21] Dustin G. Mixon and Hans Parshall. Globally optimizing small codes in real projective spaces. SIAM Journal on Discrete Mathematics, 35 (1):234–249, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Martinez-Penas:2022:GFM

[MP22] Umberto Martínez-Peñas. A general family of MSRD codes and PMDS codes with smaller field sizes from extended Moore matrices. SIAM Journal on Discrete Mathematics, 36(3):1868– 1886, ???? 2022. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:/

/epubs.siam.org/doi/10.
1137/20M1386001.

Morales:2017:HFS

[MPP17] Alejandro H. Morales, Igor Pak, and Greta Panova. Hook formulas for skew shapes II. Combinatorial proofs and enumerative applications. SIAM Journal on Discrete Mathematics, 31(3): 1953–1989, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Melczer:2020:CPI

[MPP20] Stephen Melczer, Greta Panova, and Robin Pemantle. Counting partitions inside a rectangle. SIAM Journal on Discrete Mathematics, 34(4): 2388–2410, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Masarik:2022:CCB

[MPRS22] Tomás Masarík, Marcin Pilipczuk, Paweł Rzazewski, and Manuel Sorge. Constant congestion brambles in directed graphs. SIAM Journal on Discrete Mathematics, 36(2):922–938, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1417661.

Ma:2006:CD

[MPS06] Jun Ma, Liqun Pu, and Hao Shen. Cycle decompositions of K_{n,n} −I. SIAM Journal on Discrete Mathematics, 20(3): 603–609, January 2006. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Matousek:2008:HMP

[MPŠ08] Jiří Matoušek, Aleš Přívĕtivý, and Petr Škovroň. How many points can be reconstructed from k projections? SIAM Journal on Discrete Mathematics, 22(4):1605–1623, ???? 2008. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Morton:2009:CRT

[MPS^{+}09] Jason Morton, Lior Pachter, Anne Shiu, Bernd Sturmfels, and Oliver Wienand. Convex rank tests and semigraphoids. SIAM Journal on Discrete Mathematics, 23(3): 1117–1134, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Macourt:2020:BTT

[MPSS20] Simon Macourt, Giorgis Petridis, Ilya D. Shkredov, and Igor E. Shparlinski. Bounds of trilinear and trinomial exponential sums. SIAM Journal on Discrete Mathematics, 34(4):2124–2136, ???? 2020. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

McCormick:2021:PTA

[MPSV21] S. Thomas McCormick, Britta Peis, Robert Scheidweiler, and Frank Vallentin. A polynomial time algorithm for solving the closest vector problem in zonotopal lattices. SIAM Journal on Discrete Mathematics, 35(4): 2345–2356, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Meesum:2019:RVC

[MPSZ19] Syed M. Meesum, Fahad Panolan, Saket Saurabh, and Meirav Zehavi. Rank vertex cover as a natural problem for algebraic compression. SIAM Journal on Discrete Mathematics, 33(3):1277–1296, ???? 2019. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Miers:2004:CSGa

[MR04a] C. R. Miers and F. Ruskey.

Counting strings with given elementary symmetric function evaluations I: Strings over z_{p} with p prime. SIAM Journal on Discrete Mathematics, 17(4):675–685, 2004. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/41538.

Miers:2004:CSGb

[MR04b] C. R. Miers and F. Ruskey.

Counting strings with given elementary symmetric function evaluations II: Circular strings. SIAM Journal on Discrete Mathematics, 18(1):71–82, 2004. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/42805.

Mayhew:2012:ICB

[MR12] Dillon Mayhew and Gordon Royle. The internally 4-connected binary matroids with no M(K_{5}\e)- minor. SIAM Journal on Discrete Mathematics, 26(2): 755–767, ???? 2012. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Moore:2015:ARA

[MR15a] Cristopher Moore and Alexander Russell. Approximate representations, approximate homomorphisms, and lowdimensional embeddings of groups. SIAM Journal on Discrete Mathematics, 29(1): 182–197, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Moore:2015:OBS

[MR15b] Cristopher Moore and Alexander Russell. Optimal ε-biased sets with just a little randomness. SIAM Journal on Discrete Mathematics, 29(3): 1303–1311, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

McDevitt:2021:AWQ

[MR21] M. McDevitt and N. Ruskuc. Atomicity and well quasiorder for consecutive orderings on words and permutations. SIAM Journal on Discrete Mathematics, 35(1): 495–520, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Mrazovic:2017:OPC

[Mra17] Rudi Mrazović. One-point concentration of the clique and chromatic numbers of the random Cayley graph on F_{2}^{n}. SIAM Journal on Discrete Mathematics, 31(1): 143–154, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Medina:2019:NUD

[MRAS19] Carolina Medina, Jorge Ramírez-Alfonsín, and Gelasio Salazar. On the number of unknot diagrams. SIAM Journal on Discrete Mathematics, 33(1):306–326, ???? 2019. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Murphy:2015:VSP

[MRNS15] Brendan Murphy, Oliver Roche-Newton, and Ilya Shkredov. Variations on the sum-product problem. SIAM Journal on Discrete Mathematics, 29(1):514–540, ???? 2015. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Murphy:2017:VSP

[MRNS17] Brendan Murphy, Oliver Roche-Newton, and Ilya D.

Shkredov. Variations on the sum–product problem II. SIAM Journal on Discrete Mathematics, 31(3): 1878–1894, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Mansour:2020:FAP

[MRR20] Toufik Mansour, Reza Rastegar, and Alexander Roitershtein. Finite automata, probabilistic method, and occurrence enumeration of a pattern in words and permutations. SIAM Journal on Discrete Mathematics, 34(2): 1011–1038, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Motwani:1989:PGO

[MRS89] Rajeev Motwani, Arvind Raghunathan, and Huzur Saran. Perfect graphs and orthogonally convex covers. SIAM Journal on Discrete Mathematics, 2(3):371– 392, August 1989. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Macajova:2019:CCS

[MRS19] Edita Mácajová, Edita Rollová, and Martin Skoviera. Circuit covers of signed Eulerian graphs. SIAM Journal on Discrete Mathematics, 33 (2):933–957, ???? 2019. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Miracle:2016:SCO

[MRST16] Sarah Miracle, Dana Randall, Amanda Pascoe Streib, and Prasad Tetali. Sampling and counting 3-orientations of planar triangulations. SIAM Journal on Discrete Mathematics, 30(2):801–831, ???? 2016. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Montanari:2011:RCR

[MRT11] Andrea Montanari, Ricardo Restrepo, and Prasad Tetali. Reconstruction and clustering in random constraint satisfaction problems. SIAM Journal on Discrete Mathematics, 25(2):771–808, ???? 2011. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:

//epubs.siam.org/sidma/
resource/1/sjdmec/v25/i2/
p771_s1.

Mauhar:2017:RKP

[MRV17] Marie Mauhar, Joseph Rusinko, and Zoe Vernon. Hrepresentation of the Kimura-3 polytope for the m-claw tree. SIAM Journal on Discrete Mathematics, 31(2): 783–795, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Muir:2005:ADS

[MS05] James A. Muir and Douglas R. Stinson. Alternative digit sets for nonadjacent representations. SIAM Journal on Discrete Mathematics, 19(1):165–191, 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/43765.

Meagher:2014:EKR

[MS14] Karen Meagher and Pablo Spiga. An Erdős–Ko–Rado theorem for the derangement graph of PGL_{3}(q) acting on the projective plane. SIAM Journal on Discrete Mathematics, 28(2):918–941, ???? 2014. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Mani:2016:NLC

[MS16] Arun P. Mani and Rebecca J. Stones. The number of labeled connected graphs modulo prime powers. SIAM Journal on Discrete Mathematics, 30(2):1046–1057, ???? 2016. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Macajova:2017:NZF

[MS17a] Edita Mácajová and Martin Skoviera. Nowhere-zero flows on signed Eulerian graphs. SIAM Journal on Discrete Mathematics, 31(3):

1937–1952, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Macajova:2017:ODE

[MS17b] Edita Mácajová and Martin Skoviera. Odd decompositions of Eulerian graphs. SIAM Journal on Discrete Mathematics, 31(3):1923– 1936, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Misra:2019:BES

[MS19] Pratik Misra and Seth Sullivant. Bounds on the expected size of the maximum agreement subtree for a given tree shape. SIAM Journal on Discrete Mathematics, 33(4): 2316–2325, ???? 2019. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Manurangsi:2020:WDE

[MS20] Pasin Manurangsi and Warut Suksompong. When do envyfree allocations exist? SIAM Journal on Discrete Mathematics, 34(3):1505–1521, ???? 2020. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Macajova:2021:CGN

[MS21a] Edita Mácajová and Martin Skoviera. Cubic graphs with no short cycle covers. SIAM Journal on Discrete Mathematics, 35(3):2223–2233, ???? 2021. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Macajova:2021:PMI

[MS21b] Edita Mácajová and Martin Skoviera. Perfect matching index versus circular flow number of a cubic graph. SIAM Journal on Discrete Mathematics, 35(2): 1287–1297, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Manurangsi:2021:CGA

[MS21c] Pasin Manurangsi and Warut Suksompong. Closing gaps in asymptotic fair division. SIAM Journal on Discrete Mathematics, 35(2): 668–706, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Miracle:2024:RMB

[MS24] Sarah Miracle and Amanda Pascoe Streib. Rapid mixing of k-class biased permutations. SIAM Journal on Discrete Mathematics, 38(1): 702–725, February 2024. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Miller:2019:KNT

[MSD19] Stephen D. Miller and Noah Stephens-Davidowitz. Kissing numbers and transference theorems from generalized tail bounds. SIAM Journal on Discrete Mathematics, 33(3):1313–1325, ???? 2019. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Morris:1993:EFM

[MSK93] Macdonald Morris, Gabriel Schachtel, and Samuel Karlin. Exact formulas for multitype run statistics in a random ordering. SIAM Journal on Discrete Mathematics, 6(1):70–86, February 1993. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

McKay:2014:CNQ

[MSS14] Brendan D. McKay, Pascal Schweitzer, and Patrick Schweitzer. Competition numbers, quasi-line graphs, and holes. SIAM Journal on Discrete Mathematics, 28 (1):77–91, ???? 2014. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Miskuf:2009:BCG

[MŠT09] Jozef Miškuf, Riste Škrekovski, and Martin Tancer. Backbone colorings and generalized Mycielski graphs. SIAM Journal on Discrete Mathematics, 23(2):1063–1070, ???? 2009. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Magnard:2021:SSI

[MST21] Thomas Magnard, Michael Skotnica, and Martin Tancer. Shellings and sheddings induced by collapses. SIAM Journal on Discrete Mathematics, 35(3):1978–2002, ???? 2021. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Mertzios:2010:NIM

[MSZ10] George B. Mertzios, Ignasi Sau, and Shmuel Zaks. A new intersection model and improved algorithms for tolerance graphs. SIAM Journal on Discrete Mathematics, 23(4):1800–1813, ???? 2010. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Manoussakis:1990:PAF

[MT90] Y. Manoussakis and Z. Tuza. Polynomial algorithms for finding cycles and paths in bipartite tournaments. SIAM Journal on Discrete Mathematics, 3(4):537–543, November 1990. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Metelsky:2003:LGH

[MT03] Yury Metelsky and Regina Tyshkevich. Line graphs of Helly hypergraphs. SIAM Journal on Discrete Mathematics, 16(3):438–448, 2003. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/36521.

Maffray:2005:APC

[MT05] Frédéric Maffray and Nicolas Trotignon. Algorithms for perfectly contractile graphs. SIAM Journal on Discrete Mathematics, 19(3): 553–574, 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/44252.

Mezard:2011:SSC

[MT11] Marc Mézard and Prasad Tetali. Special section on constraint satisfaction problems and message passing algorithms. SIAM Journal on Discrete Mathematics, 25(2): 733–735, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v25/i2/p733_s1.

Mubayi:2020:ETL

[MT20] Dhruv Mubayi and Caroline Terry. Extremal theory of locally sparse multigraphs. SIAM Journal on Discrete Mathematics, 34(3): 1922–1943, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Mudgal:2005:BTC

[MTGK05] Apurva Mudgal, Craig Tovey, Sam Greenberg, and Sven Koenig. Bounds on the travel cost of a Mars Rover prototype search heuristic. SIAM Journal on Discrete Mathematics, 19(2):431–447, 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/44256.

Mohammadian:2014:FDF

[MTR14] A. Mohammadian and B. Tayfeh-Rezaie. On a family of diamond-free strongly regular graphs. SIAM Journal on Discrete Mathematics, 28(4): 1906–1915, ???? 2014. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Maffray:2008:ASF

[MTV08] Frédéric Maffray, Nicolas Trotignon, and Kristina Vušković. Algorithms for square-3PC(·, ·)-free Berge graphs. SIAM Journal on Discrete Mathematics, 22(1): 51–71, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Munoz:2005:IMC

[Muñ05]  S. Muñoz.  On identify-

ing maximal covers. SIAM Journal on Discrete Mathematics, 18(4):749–768, 2005.

CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/40958.

Munarini:2006:CIC

[Mun06] Emanuele Munarini. A combinatorial interpretation of the Chebyshev polynomials. SIAM Journal on Discrete Mathematics, 20(3): 649–655, January 2006. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Murota:1996:VMIa

[Mur96a] Kazuo Murota. Valuated matroid intersection. I. Optimality criteria. SIAM Journal on Discrete Mathematics, 9 (4):545–561, November 1996. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Murota:1996:VMIb

[Mur96b] Kazuo Murota. Valuated matroid intersection. II. Algorithms. SIAM Journal on Discrete Mathematics, 9 (4):562–576, November 1996. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Murota:2006:CFJ

[Mur06] Kazuo Murota. M-convex functions on jump systems: a general framework for minsquare graph factor problem. SIAM Journal on Discrete Mathematics, 20(1): 213–226, January 2006. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Murai:2010:FVB

[Mur10] Satoshi Murai. On face vectors of barycentric subdivisions of manifolds. SIAM Journal on Discrete Mathematics, 24(3):1019–1037, ???? 2010. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Musin:2021:MME

[Mus21] Oleg R. Musin. Majorization and minimal energy on spheres. SIAM Journal on Discrete Mathematics, 35(3): 1578–1591, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Mohammadi:2018:GPP

[MUWY18] Fatemeh Mohammadi, Caroline Uhler, Charles Wang, and Josephine Yu. Generalized permutohedra from probabilistic graphical models. SIAM Journal on Discrete Mathematics, 32(1):64– 93, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Mahajan:1999:DOA

[MV99] Meena Mahajan and V. Vinay. Determinant: Old algorithms, new insights. SIAM Journal on Discrete Mathematics, 12(4):474–490, November 1999. CODEN SJD-MEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/33882.

Megow:2018:DTS

[MV18] Nicole Megow and José Verschae. Dual techniques for scheduling on a machine with varying speed. SIAM Journal on Discrete Mathematics, 32(3):1541–1571, ???? 2018. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Muller:2013:IRC

[MvLvL13] Tobias Müller, Erik Jan van Leeuwen, and Jan van Leeuwen. Integer representations of convex polygon intersection graphs. SIAM Journal on Discrete Mathematics, 27 (1):205–231, ???? 2013. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Moran:1990:OPB

[MW90] Shlomo Moran and Yaron Wolfstahl. One-page book embedding under vertexneighborhood constraints. SIAM Journal on Discrete Mathematics, 3(3):376–390, August 1990. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Myers:2003:SNA

[MW03] Amy N. Myers and Herbert S. Wilf. Some new aspects of the coupon Collector’s problem. SIAM Journal on Discrete Mathematics, 17(1):1–17, 2003. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/40307.

McKay:2008:CSL

[MW08] Brendan D. McKay and Ian M. Wanless. A census of small Latin hypercubes. SIAM Journal on Discrete Mathematics, 22(2): 719–736, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Marx:2014:IHE

[MW14] Dániel Marx and Paul Wollan. Immersions in highly edge connected graphs. SIAM Journal on Discrete Mathematics, 28(1):503–520, ???? 2014. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Melczer:2019:HDL

[MW19a] Stephen Melczer and Mark C. Wilson. Higher dimensional lattice walks: Connecting combinatorial and analytic behavior. SIAM Journal on Discrete Mathematics, 33(4): 2140–2174, ???? 2019. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Micheli:2019:RUM

[MW19b] Giacomo Micheli and Violetta Weger. On rectangular unimodular matrices over the algebraic integers. SIAM Journal on Discrete Mathematics, 33(1):425–437, ???? 2019. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Milans:2020:DTF

[MW20a] Kevin G. Milans and Michael C.

Wigal. A dichotomy theorem for first-fit chain partitions. SIAM Journal on Discrete Mathematics, 34(1): 586–596, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Murphy:2020:GSF

[MW20b] Brendan Murphy and James Wheeler. Growth in some finite three-dimensional matrix groups. SIAM Journal on Discrete Mathematics, 34(3): 1984–1998, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Moulton:2022:PSN

[MW22] Vincent Moulton and Taoyang Wu. Posets and spaces of k-noncrossing RNA structures. SIAM Journal on Discrete Mathematics, 36 (3):1586–1611, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1413316.

Mayhew:2011:ODT

[MWvZ11] Dillon Mayhew, Geoff Whittle, and Stefan H. M. van Zwam. An obstacle to a decomposition theorem for near-regular matroids. SIAM Journal on Discrete Mathematics, 25(1): 271–279, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v25/i1/p271_s1.

McMorris:1994:TVC

[MWW94a] F. R. McMorris, Tandy J. Warnow, and Thomas Wimer. Triangulating vertex-colored graphs. SIAM Journal on Discrete Mathematics, 7(2): 296–306, May 1994. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Mitchell:1994:CPA

[MWW94b] Chris Mitchell, Michael Walker, and Peter Wild. The combinatorics of perfect authentication schemes. SIAM Journal on Discrete Mathematics, 7(1):102–107, February 1994. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Makhul:2020:SKP

[MWW20] Mehdi Makhul, Audie Warren, and Arne Winterhof. The spherical Kakeya problem in finite fields. SIAM Journal on Discrete Mathematics, 34(4): 2502–2509, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Madiman:2021:EIS

[MWW21] Mokshay Madiman, Liyao Wang, and Jae Oh Woo. Entropy inequalities for sums in prime cyclic groups. SIAM Journal on Discrete Mathematics, 35(3):1628–1649, ???? 2021. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ma:2020:SPC

[MWWZ20] Xuanlong Ma, Gary L. Walls, Kaishun Wang, and Sanming Zhou. Subgroup perfect codes in Cayley graphs. SIAM Journal on Discrete Mathematics, 34(3):1909–1921, ???? 2020. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Myung:2001:EAR

[Myu01] Young-Soo Myung. An efficient algorithm for the ring loading problem with integer demand splitting. SIAM Journal on Discrete Mathematics, 14(3):291–298, 2001. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/35870.

McGinnis:2022:LTF

[MZ22] Daniel McGinnis and Shira Zerbib. Line transversals in families of connected sets in the plane. SIAM Journal on Discrete Mathematics, 36 (4):2916–2919, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1408920.

Naatz:2000:GLE

[Naa00] Michael Naatz. The graph of linear extensions revisited. SIAM Journal on Discrete Mathematics, 13 (3):354–369, 2000. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/35260.

Naatz:2001:NQC

[Naa01] Michael Naatz. A note on a question of C. D. Savage. SIAM Journal on Discrete Mathematics, 14 (1):116–120, February 2001. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/36919.  See

[Sav97].

Nagle:2010:CFI

[Nag10] Brendan Nagle. On computing the frequencies of induced subhypergraphs. SIAM Journal on Discrete Mathematics, 24(1):322–329, ???? 2010. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Nagle:2017:CPT

[Nag17] Brendan Nagle. Constructive packings of triple systems. SIAM Journal on Discrete Mathematics, 31(4): 2479–2516, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Naor:1991:LBP

[Nao91] Moni Naor. A lower bound on probabilistic algorithms for distributive ring coloring. SIAM Journal on Discrete Mathematics, 4(3):409– 412, August 1991. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Nasre:2014:PMS

[Nas14] Meghana Nasre. Popular matchings: Structure and strategic issues. SIAM Journal on Discrete Mathematics, 28(3):1423–1448, ???? 2014. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Neelamani:2007:NOL

[NDB07] Ramesh Neelamani, Sanjeeb Dash, and Richard G. Baraniuk. On nearly orthogonal lattice bases and random lattices. SIAM Journal on Discrete Mathematics, 21(1): 199–219, ???? 2007. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Nelson:2015:MDT

[Nel15]  Peter Nelson.  Matroids

denser than a projective geometry. SIAM Journal on Discrete Mathematics, 29(2): 730–735, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Neuen:2024:ITP

[Neu24]  Daniel Neuen.  Isomor-

phism testing parameterized by genus and beyond. SIAM Journal on Discrete Mathematics, 38(1):453–484, January 2024. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Newman:2020:IAM

[New20]  Alantha Newman.  An

improved analysis of the Mömke–Svensson algorithm for graph-TSP on subquartic graphs. SIAM Journal on Discrete Mathematics, 34(1): 865–884, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Nguyen:2010:ARM

[Ngu10]  Viet-Hang Nguyen.  On

abstract rigidity matroids. SIAM Journal on Discrete Mathematics, 24(2):363–369, ???? 2010. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Nguyen:2013:SRC

[Ngu13] Hoi H. Nguyen. On the singularity of random combinatorial matrices. SIAM Journal on Discrete Mathematics, 27 (1):447–458, ???? 2013. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Nguyen:2024:HCS

[Ngu24] Tung H. Nguyen. Highly connected subgraphs with large chromatic number. SIAM Journal on Discrete Mathematics, 38(1):243–260, January 2024. CODEN SJD-MEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ng:1991:TDS

[NH91] Cheng Ng and Daniel S.

Hirschberg. Three-dimensional stable matching problems. SIAM Journal on Discrete Mathematics, 4(2):245–252, May 1991. CODEN SJD-MEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Nagamochi:1992:CEC

[NI92] Hiroshi Nagamochi and Toshihide Ibaraki. Computing edge-connectivity in multigraphs and capacitated graphs. SIAM Journal on Discrete Mathematics, 5(1): 54–66, February 1992. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Nierhoff:2000:TBI

[Nie00] Till Nierhoff. A tight bound on the irregularity strength of graphs. SIAM Journal on Discrete Mathematics, 13(3):313–323, 2000. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/31429.

Nishizeki:1990:ECM

[NK90] Takao Nishizeki and Kenichi Kashiwagi. On the 1.1 edge-coloring of multigraphs. SIAM Journal on Discrete Mathematics, 3(3):391–410, August 1990. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Noguchi:2021:BCM

[NLJM21] Satoshi Noguchi, Xiao-Nan Lu, Masakazu Jimbo, and Ying Miao. BCH codes with minimum distance proportional to code length. SIAM Journal on Discrete Mathematics, 35(1):179–193, ???? 2021. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Niazadeh:2024:BFF

[NLS24] Rad Niazadeh, Renato Paes Leme, and Jon Schneider. Bernoulli factories for flowbased polytopes. SIAM Journal on Discrete Mathematics, 38(1):726–742, February 2024. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Nelson:2022:SFT

[NN22] Peter Nelson and Kazuhiro Nomoto. The structure of I_{4}-free and triangle-free binary matroids. SIAM Journal on Discrete Mathematics, 36(3):1711–1729, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1343099.

Nagamochi:1997:CAS

[NNI97] Hiroshi Nagamochi, Kazuhiro Nishimura, and Toshihide Ibaraki. Computing all small cuts in an undirected network. SIAM Journal on Discrete Mathematics, 10(3): 469–481, August 1997. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Nakamoto:2015:EET

[NNO15] Atsuhiro Nakamoto, Kenta Noguchi, and Kenta Ozeki. Extension to even triangulations. SIAM Journal on Discrete Mathematics, 29(4): 2075–2087, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Nakamoto:2019:ECT

[NNO19] Atsuhiro Nakamoto, Kenta Noguchi, and Kenta Ozeki. Extension to 3-colorable triangulations. SIAM Journal on Discrete Mathematics, 33 (3):1390–1414, ???? 2019. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Niederreiter:2008:IAB

[NÖ08] Harald Niederreiter and Ferruh Özbudak. Improved asymptotic bounds for codes using distinguished divisors of global function fields. SIAM Journal on Discrete Mathematics, 21(4):865–899, ???? 2008. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Nielsen:2009:STC

[NO09] Morten H. Nielsen and Ortrud R. Oellermann. Steiner trees and convex geometries. SIAM Journal on Discrete Mathematics, 23(2): 680–693, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Noguchi:2024:SBQ

[Nog24] Kenta Noguchi. Spanning bipartite quadrangulations of triangulations of the projective plane. SIAM Journal on Discrete Mathematics, 38(2): 1250–1268, April 2024. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Nakamoto:2012:BET

[NOO12] Atsuhiro Nakamoto, Katsuhiro Ota, and Kenta Ozeki. Book embedding of toroidal bipartite graphs. SIAM Journal on Discrete Mathematics, 26(2):661–669, ???? 2012. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Nixon:2012:RFS

[NOP12] A. Nixon, J. C. Owen, and S. C. Power. Rigidity of frameworks supported on surfaces. SIAM Journal on Discrete Mathematics, 26(4): 1733–1757, ???? 2012. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Nixon:2014:CGR

[NOP14] A. Nixon, J. C. Owen, and S. C. Power. A characterization of generically rigid frameworks on surfaces of revolution. SIAM Journal on Discrete Mathematics, 28(4):

2008–2028, ???? 2014. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Norine:2011:TGN

[Nor11] Serguei Norine. Turán graphs and the number of colorings. SIAM Journal on Discrete Mathematics, 25(1): 260–266, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sidma/resource/
                                        1/sjdmec/v25/i1/p260_s1.

Novik:2018:ASC

[Nov18] Isabella Novik. From acute sets to centrally symmetric 2neighborly polytopes. SIAM Journal on Discrete Mathematics, 32(3):1572–1576, ???? 2018. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Nguyen:2018:EPI

[NP18] Danny Nguyen and Igor Pak. Enumerating projections of integer points in unbounded polyhedra. SIAM Journal on Discrete Mathematics, 32(2): 986–1002, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Nelson:2020:RTB

[NP20a] Peter Nelson and Sophia Park. A Ramsey theorem for biased graphs. SIAM Journal on Discrete Mathematics, 34(4):2270–2281, ???? 2020. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Nenadov:2020:RTV

[NP20b] Rajko Nenadov and Yanitsa Pehova. On a Ramsey– Turán variant of the Hajnal– Szemerédi theorem. SIAM Journal on Discrete Mathematics, 34(2):1001–1010, ???? 2020. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Nederlof:2023:HCP

[NPSW23] Jesper Nederlof, Michał Pilipczuk, Celine M. F. Swennenhuis, and Karol Wegrzycki. Hamiltonian cycle parameterized by treedepth in single exponential time and polynomial space. SIAM Journal on Discrete Mathematics, 37(3):1566–1586, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL

https://epubs.siam.org/
doi/10.1137/22M1518943.

Nisan:1989:PAZ

[NS89] Noam Nisan and Danny Soroker. Parallel algorithms for zero-one supply-demand problems. SIAM Journal on Discrete Mathematics, 2(1): 108–125, February 1989. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Nobili:1993:PPC

[NS93] P. Nobili and A. Sassano. Polyhedral properties of clutter amalgam. SIAM Journal on Discrete Mathematics, 6(1):139–151, February 1993. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Nesetril:2007:SDD

[NŠ07] Jaroslav Nešetřil and Ida Švejdarová. Small diameters of duals. SIAM Journal on Discrete Mathematics, 21(2): 374–384, ???? 2007. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Nozaki:2011:BDS

[NS11] Hiroshi Nozaki and Sho Suda. Bounds on s-distance sets with strength t. SIAM Journal on Discrete Mathematics, 25(4):1699–1713, ???? 2011. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sidma/resource/
                                        1/sjdmec/v25/i4/p1699_s1.

Nguyen:2016:CCC

[NS16] Phong Q. Nguyen and Igor E. Shparlinski. Counting cocyclic lattices. SIAM Journal on Discrete Mathematics, 30(3):1358–1370, ???? 2016. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Naves:2022:WDG

[NS22] Guyslain Naves and Bruce Shepherd. When do Gomory– Hu subtrees exist? SIAM Journal on Discrete Mathematics, 36(3):1567–1585, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL

https://epubs.siam.org/
doi/10.1137/20M1356968.

Nixon:2018:RFE

[NSiTW18] Anthony Nixon, Bernd Schulze, Shin ichi Tanigawa, and Walter Whiteley. Rigidity of frameworks on expanding spheres. SIAM Journal on Discrete Mathematics, 32(4): 2591–2611, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Nesetril:2005:SAE

[NT05] Jaroslav Nesetril and Claude Tardif. Short answers to exponentially long questions: Extremal aspects of homomorphism duality. SIAM Journal on Discrete Mathematics, 19(4):914–920, 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/44563.

Nakamigawa:2012:CLP

[NT12] Tomoki Nakamigawa and Norihide Tokushige. Counting lattice paths via a new cycle lemma. SIAM Journal on Discrete Mathematics, 26 (2):745–754, ???? 2012. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Nenadov:2021:SFR

[NT21] Rajko Nenadov and Milos Trujić. Sprinkling a few random edges doubles the power. SIAM Journal on Discrete Mathematics, 35(2): 988–1004, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Nagle:2024:SCT

[NT24] Brendan Nagle and John Theado. Some cubic time regularity algorithms for triple systems. SIAM Journal on Discrete Mathematics, 38(1): 668–701, February 2024. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Nie:2022:RNN

[NV22] Jiaxi Nie and Jacques Verstraëte. Ramsey numbers for nontrivial Berge cycles. SIAM Journal on Discrete Mathematics, 36(1):103–113, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL

                                        https://epubs.siam.org/
                                        doi/10.1137/21M1396770.

Nelson:2018:DEM

[NvdP18] Peter Nelson and Jorn van der Pol. Doubly exponentially many Ingleton matroids. SIAM Journal on Discrete Mathematics, 32(2): 1145–1153, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Nelson:2019:NBG

[NvdP19] Peter Nelson and Jorn van der Pol. On the number of biased graphs. SIAM Journal on Discrete Mathematics, 33(1):373–382, ???? 2019. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Nomoto:2024:TCB

[NvdP24] Kazuhiro Nomoto and Jorn van der Pol. Tuza’s conjecture for binary geometries. SIAM Journal on Discrete Mathematics, 38(2): 1676–1685, May 2024. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Nelson:2015:MRF

[NvZ15] Peter Nelson and Stefan H. M. van Zwam. Matroids representable over fields with a common subfield. SIAM Journal on Discrete Mathematics, 29(2):796–810, ???? 2015. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Newman:1995:LBF

[NW95] Ilan Newman and Avi Wigderson. Lower bounds on formula size of Boolean functions using hypergraph entropy. SIAM Journal on Discrete Mathematics, 8(4):536– 542, November 1995. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Nelson:2022:EFE

[NW22] Peter Nelson and Zach Walsh. The extremal function for excluding geometry minors over prime fields. SIAM Journal on Discrete Mathematics, 36 (3):2100–2120, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1410750.

Nassar:2021:AFA

[NY21] Safwat Nassar and Raphael Yuster. All feedback arc sets of a random Turán tournament have ⌊n/k⌋ − k + 1 disjoint k-cliques (and this is tight). SIAM Journal on Discrete Mathematics, 35(2): 1460–1477, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Naserasr:2023:PSS

[NY23] Reza Naserasr and Weiqiang Yu. Packing signatures in signed graphs. SIAM Journal on Discrete Mathematics, 37(4):2365–2381, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1411883.

Nakashima:2020:FSH

[NYKY20] Junpei Nakashima, Yukiko Yamauchi, Shuji Kijima, and Masafumi Yamashita. Finding submodularity hidden in symmetric difference. SIAM Journal on Discrete Mathematics, 34(1):571–585, ???? 2020. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

o:2009:LBI

[o09] Jakub Przyby o. Linear bound on the irregularity strength and the total vertex irregularity strength of graphs. SIAM Journal on Discrete Mathematics, 23(1): 511–516, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Obrenic:1993:EBS

[Obr93] Bojana Obrenić. Embedding de Bruijn and shuffleexchange graphs in five pages. SIAM Journal on Discrete Mathematics, 6(4):642–654, November 1993. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

O:2010:ECE

[OC10] Suil O and Sebastian M.

Cioabă. Edge-connectivity, eigenvalues, and matchings in regular graphs. SIAM Journal on Discrete Mathematics, 24(4):1470–1481, ???? 2010. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Oursler:2019:TMD

[OC19] Roy Oursler and Andrzej Czygrinow. Tight minimum degree condition for the existence of loose cycle tilings in 3-graphs. SIAM Journal on Discrete Mathematics, 33(4): 1912–1931, ???? 2019. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Offner:2008:PCS

[Off08] David Offner. Polychromatic colorings of subcubes of the hypercube. SIAM Journal on Discrete Mathematics, 22(2): 450–454, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ogata:2006:OSS

[OKS06] Wakaha Ogata, Kaoru Kurosawa, and Douglas R. Stinson. Optimum secret sharing scheme secure against cheating. SIAM Journal on Discrete Mathematics, 20(1):79– 95, January 2006. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Onn:1991:GCC

[Onn91] Shmuel Onn. On the geometry and computational complexity of Radon partitions in the integer lattice. SIAM Journal on Discrete Mathematics, 4(3):436–446, August 1991. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Onn:2003:CMO

[Onn03] Shmuel Onn. Convex matroid optimization. SIAM Journal on Discrete Mathematics, 17(2):249–253, 2003. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/40855.

Ozeki:2019:BEG

[ONN19] Kenta Ozeki, Atsuhiro Nakamoto, and Takayuki Nozawa. Book embedding of graphs on the projective plane. SIAM Journal on Discrete Mathematics, 33(4):1801–1836, ???? 2019. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ordentlich:2012:AEB

[OPR12] Erik Ordentlich, Farzad Parvaresh, and Ron M. Roth. Asymptotic enumeration of binary matrices with bounded row and column sums. SIAM Journal on Discrete Mathematics, 26(4): 1550–1575, ???? 2012. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Oh:2021:FAM

[OPS21] Hoon Oh, Ariel D. Procaccia, and Warut Suksompong. Fairly allocating many goods with few queries. SIAM Journal on Discrete Mathematics, 35(2):788–813, ???? 2021. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

OCarroll:2014:DAP

[OPVV14] Liam O’Carroll, Francesc Planas-Vilanova, and Rafael H. Villarreal. Degree and algebraic properties of lattice and matrix ideals. SIAM Journal on Discrete Mathematics, 28 (1):394–427, ???? 2014. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ordentlich:2004:ISR

[OR04] Erik Ordentlich and Ron M. Roth. Independent sets in regular hypergraphs and multidimensional runlengthlimited constraints. SIAM Journal on Discrete Mathematics, 17(4):615–623, 2004. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/41976.

Orlitsky:1993:ICB

[Orl93] Alon Orlitsky. Interactive communication of balanced distributions and of correlated files. SIAM Journal on Discrete Mathematics, 6 (4):548–564, November 1993. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Orrick:2008:SOH

[Orr08] William P. Orrick. Switching operations for Hadamard matrices. SIAM Journal on Discrete Mathematics, 22(1): 31–50, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Olesky:1992:FLR

[OS92] D. D. Olesky and T. A.

Slater. A forward lower restricted ordering algorithm for digraphs. SIAM Journal on Discrete Mathematics, 5(1):67–73, February 1992. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ostafe:2011:PDF

[OS11] Alina Ostafe and Igor E. Shparlinski. Pseudorandomness and dynamics of Fermat quotients. SIAM Journal on Discrete Mathematics, 25(1):50–71, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.

                                        siam.org/sidma/resource/
                                        1/sjdmec/v25/i1/p50_s1.

ONeal:2013:UVM

[OS13a] Allen O’Neal and Peter J. Slater. Uniqueness of vertex magic constants. SIAM Journal on Discrete Mathematics, 27(2):708–716, ???? 2013. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Osthus:2013:AHD

[OS13b] Deryk Osthus and Katherine Staden. Approximate Hamilton decompositions of robustly expanding regular digraphs. SIAM Journal on Discrete Mathematics, 27(3): 1372–1409, ???? 2013. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Onn:2015:HUF

[OS15] Shmuel Onn and Pauline Sarrabezolles. Huge unimodular n-fold programs. SIAM Journal on Discrete Mathematics, 29(4):2277–2283, ???? 2015. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ordentlich:2016:UBS

[OS16] Or Ordentlich and Ofer Shayevitz. An upper bound on the sizes of multiset-union-free families. SIAM Journal on Discrete Mathematics, 30(2):1032–1045, ???? 2016. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Omidi:2017:DRN

[OS17] G. R. Omidi and M. Shahsiah. Diagonal Ramsey numbers of loose cycles in uniform hypergraphs. SIAM Journal on Discrete Mathematics, 31 (3):1634–1669, ???? 2017. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Oxley:2022:SCB

[OS22] James Oxley and Jagdeep Singh. The smallest classes of binary and ternary matroids closed under direct sums and complements. SIAM Journal on Discrete Mathematics, 36(3):2051–2072, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1453852.

Oshima:2021:IRA

[Osh21] Hiroki Oshima. Improved randomized algorithm for ksubmodular function maximization. SIAM Journal on Discrete Mathematics, 35(1): 1–22, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Osting:2023:EGR

[Ost23] Braxton Osting. Extremal graph realizations and graph Laplacian eigenvalues. SIAM Journal on Discrete Mathematics, 37(3):1630–1644, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL

https://epubs.siam.org/
doi/10.1137/22M1504421.

Oxley:2016:WWT

[OSW16] James Oxley, Charles Semple, and Geoff Whittle. A wheelsand-whirls theorem for 3connected 2-polymatroids. SIAM Journal on Discrete Mathematics, 30(1):493–524, ???? 2016. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ostergaard:2004:REF

[ÖV04] Patric R. J. Östergård and Alexander Vardy. Resolving the existence of fullrank tilings of binary Hamming spaces. SIAM Journal on Discrete Mathematics, 18(2):382–387, 2004. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/43580.

Ollivier:2012:CBM

[OV12] Y. Ollivier and C. Villani. A curved Brunn–Minkowski inequality on the discrete hypercube, or: What is the Ricci curvature of the discrete hypercube? SIAM Journal on Discrete Mathematics, 26 (3):983–996, ???? 2012. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Oxley:2022:MM

[OW22a] James Oxley and Zach Walsh.

2-modular matrices. SIAM Journal on Discrete Mathematics, 36(2):1231–1248, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL

https://epubs.siam.org/
doi/10.1137/21M1419131.

Oxley:2022:SCM

[OW22b] James Oxley and Zach Walsh.

Small cocircuits in minimally vertically 4-connected matroids. SIAM Journal on Discrete Mathematics, 36 (4):2822–2829, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1454869.

Owen:2011:CGD

[Owe11] Megan Owen. Computing geodesic distances in tree space. SIAM Journal on Discrete Mathematics, 25(4): 1506–1529, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v25/i4/p1506_s1.

Oxley:2019:MER

[Oxl19] James Oxley. A matroid extension result. SIAM Journal on Discrete Mathematics, 33 (1):138–152, ???? 2019. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Oh:2013:RGS

[OYY13] Suho Oh, Hwanchul Yoo, and Taedong Yun. Rainbow graphs and switching classes. SIAM Journal on Discrete Mathematics, 27(2): 1106–1111, ???? 2013. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ozeki:2018:ECG

[OZ18] Kenta Ozeki and Carol T. Zamfirescu. Every 4connected graph with crossing number 2 is Hamiltonian. SIAM Journal on Discrete Mathematics, 32(4): 2783–2794, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ozeki:2013:STB

[Oze13]  Kenta Ozeki.  Spanning

trees with bounded maximum degrees of graphs on surfaces. SIAM Journal on Discrete Mathematics, 27(1): 422–435, ???? 2013. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Padrol:2016:ECP

[Pad16] Arnau Padrol. Extension complexity of polytopes with few vertices or facets. SIAM Journal on Discrete Mathematics, 30(4):2162–2176, ???? 2016. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Pal:2012:BAC

[Pal12]  Soumik Pal.  Brownian

approximation to counting graphs. SIAM Journal on Discrete Mathematics, 26(3): 1181–1188, ???? 2012. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Paterson:1988:UCW

[Pat88] M. S. Paterson. Universal chains and wiring layouts. SIAM Journal on Discrete Mathematics, 1(1):80– 85, February 1988. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Payne:2017:BLP

[Pay17] Michael S. Payne. Bichromatic lines in the plane. SIAM Journal on Discrete Mathematics, 31(2):857–864, ???? 2017. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Provan:1997:CPA

[PC97] J. Scott Provan and Manoj K.

Chari. Counting problems associated with Steiner trees in graphs. SIAM Journal on Discrete Mathematics, 10(3):

436–446, August 1997. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Proenca:2021:DHB

[PdCSC21] Nathan Benedetto Proença, Marcel K. de Carli Silva, and Gabriel Coutinho. Dual Hoffman bounds for the stability and chromatic numbers based on semidefinite programming. SIAM Journal on Discrete Mathematics, 35(4): 2880–2907, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Pegden:2014:EMT

[Peg14] Wesley Pegden. An extension of the Moser–Tardos algorithmic local lemma. SIAM Journal on Discrete Mathematics, 28(2):911–917, ???? 2014. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Penev:2012:CBF

[Pen12] Irena Penev. Coloring bullfree perfect graphs. SIAM Journal on Discrete Mathematics, 26(3):1281–1309, ???? 2012. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Perrett:2016:CRM

[Per16] Thomas Perrett. Chromatic roots and minor-closed families of graphs. SIAM Journal on Discrete Mathematics, 30(3):1883–1897, ???? 2016. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Pettie:2011:OND

[Pet11] Seth Pettie. Origins of nonlinearity in Davenport–Schinzel sequences. SIAM Journal on Discrete Mathematics, 25(1): 211–233, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v25/i1/p211_s1.

Petersen:2013:SIO

[Pet13] T. Kyle Petersen. On the shard intersection order of a Coxeter group. SIAM Journal on Discrete Mathematics, 27(4):1880–1912, ???? 2013. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Pettie:2015:TGD

[Pet15] Seth Pettie. Three generalizations of Davenport–Schinzel sequences. SIAM Journal on Discrete Mathematics, 29(4): 2189–2238, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Pfender:2015:CDC

[Pfe15]  Florian Pfender.  Com-

bined degree and connectivity conditions for H-linked graphs. SIAM Journal on Discrete Mathematics, 29(2):

716–729, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Padro:2006:RSI

[PG06] Carles Padró and Ignacio Gracia. Representing small identically self-dual matroids by self-dual codes. SIAM Journal on Discrete Mathematics, 20(4):1046–1055, January 2006. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Picollelli:2014:FSF

[Pic14] Michael E. Picollelli. The final size of the C_{ℓ}-free process. SIAM Journal on Discrete Mathematics, 28(3): 1276–1305, ???? 2014. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Pikhurko:2003:ASR

[Pik03]  Oleg Pikhurko.  Asymp-

totic size Ramsey results for bipartite graphs. SIAM Journal on Discrete Mathematics, 16(1):99–113, 2003. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/38408.

Pinchasi:2008:MND

[Pin08] Rom Pinchasi. The minimum number of distinct areas of triangles determined by a set of n points in the plane. SIAM Journal on Discrete Mathematics, 22(2): 828–831, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Pinchasi:2014:FFP

[Pin14] Rom Pinchasi. A finite family of pseudodiscs must include a “Small” pseudodisc. SIAM Journal on Discrete Mathematics, 28(4):1930–1934, ???? 2014. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Pippenger:1989:RSA

[Pip89] Nicholas Pippenger. Random sequential adsorption on graphs. SIAM Journal on Discrete Mathematics, 2(3): 393–401, August 1989. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Pippenger:1991:ECC

[Pip91] Nicholas Pippenger. The expected capacity of concentrators. SIAM Journal on Discrete Mathematics, 4(1):121– 129, February 1991. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Pippenger:1995:ARA

[Pip95] Nicholas Pippenger. Analysis of a recurrence arising from a construction for nonblocking networks. SIAM Journal on Discrete Mathematics, 8(2):322–345, May 1995. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Pippenger:2001:EET

[Pip01] Nicholas Pippenger. Enumeration of equicolorable trees. SIAM Journal on Discrete Mathematics, 14 (1):93–115, 2001. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/36846.

Pippenger:2002:EMI

[Pip02] Nicholas Pippenger. Enumeration of matchings in the incidence graphs of complete and complete bipartite graphs. SIAM Journal on Discrete Mathematics, 16(1):47–64, 2002. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/39569.

Pippenger:2006:LPD

[Pip06] Nicholas Pippenger. The linking probability of deep spiderweb networks. SIAM Journal on Discrete Mathematics, 20(1):143–159, January 2006. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Pittel:1989:ANS

[Pit89]  Boris Pittel.  The aver-

age number of stable matchings. SIAM Journal on Discrete Mathematics, 2(4):530– 549, November 1989. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Pittel:2023:ENI

[Pit23]  Boris Pittel.  Expected

number of induced subtrees shared by two independent copies of a random tree. SIAM Journal on Discrete Mathematics, 37(1):1–16, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1416771.

Pippenger:1994:FTC

[PL94] Nicholas Pippenger and Geng Lin. Fault-tolerant circuitswitching networks. SIAM Journal on Discrete Mathematics, 7(1):108–118, February 1994. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Phelps:1999:NPC

[PL99] Kevin T. Phelps and Mike LeVan. Nonsystematic perfect codes. SIAM Journal on Discrete Mathematics, 12(1):27–34, February 1999. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/31220.

Padro:1998:FTF

[PMM98] C. Padró, P. Morillo, and X. Muñoz. Fault-tolerant fixed routings in some families of digraphs. SIAM Journal on Discrete Mathematics, 11(3):501–509, August 1998. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Polyanskiy:2019:HSA

[Pol19] Yury Polyanskiy. Hypercontractivity of spherical averages in Hamming space. SIAM Journal on Discrete Mathematics, 33(2):731–754, ???? 2019. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Pongracz:2020:BLC

[Pon20] András Pongrácz. Binary linear codes with near-extremal maximum distance. SIAM Journal on Discrete Mathematics, 34(4):2300–2317, ???? 2020. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Przytycka:1990:LBS

[PP90] T. Przytycka and J. H. Przytycki. On a lower bound for short noncontractible cycles in embedded graphs. SIAM Journal on Discrete Mathematics, 3(2):281–293, May 1990. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Paletta:2007:NAA

[PP07a] Giuseppe Paletta and Paolamaria Pietramala. A new approximation algorithm for the nonpreemptive scheduling of independent jobs on identical parallel processors. SIAM Journal on Discrete Mathematics, 21(2):313–328, ???? 2007. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Perles:2007:FSP

[PP07b] Micha A. Perles and Rom Pinchasi. Forbidden k-sets in the plane. SIAM Journal on Discrete Mathematics, 21(2): 385–395, ???? 2007. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Pinchasi:2010:DSU

[PP10] Rom Pinchasi and Allan Pinkus. Dominating subsets under projections. SIAM Journal on Discrete Mathematics, 24(3):910–920, ???? 2010. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Peserico:2012:HCC

[PP12] Enoch Peserico and Luca Pretto. HITS can converge slowly, but not too slowly, in score and rank. SIAM Journal on Discrete Mathematics, 26(3):1189–1209, ???? 2012. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Papalamprou:2013:DBS

[PP13] Konstantinos Papalamprou and Leonidas Pitsoulis. Decomposition of binary signedgraphic matroids. SIAM Journal on Discrete Mathematics, 27(2):669–692, ???? 2013. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Pach:2022:SAS

[PP22] Péter Pál Pach and Richárd Palincza. Sets avoiding six-term arithmetic progressions in Z_{6}^{n} are exponentially small. SIAM Journal on Discrete Mathematics, 36 (2):1135–1142, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1413766.

Paesani:2022:FVS

[PPR22] Giacomo Paesani, Daniël Paulusma, and Paweł Rzazwski. Feedback vertex set and even cycle transversal for H-free graphs: Finding large block graphs. SIAM Journal on Discrete Mathematics, 36(4):2453–2472, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1468864.

Pournaki:2022:SCS

[PPTY22] M. R. Pournaki, M. Poursoltani, N. Terai, and S. Yassemi. Simplicial complexes satisfying Serre’s condition versus the ones which are Cohen– Macaulay in a fixed codimension. SIAM Journal on Discrete Mathematics, 36 (4):2506–2522, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1439687.

Pemantle:1992:TPI

[PPU92] Robin Pemantle, James Propp, and Daniel Ullman. On tensor powers of integer programs. SIAM Journal on Discrete Mathematics, 5(1): 127–143, February 1992. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Pruesse:1991:GLE

[PR91] Gara Pruesse and Frank Ruskey. Generating the linear extensions of certain posets by transpositions. SIAM Journal on Discrete Mathematics, 4(3):413–422, August 1991. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Poonen:1998:NIP

[PR98] Bjorn Poonen and Michael Rubinstein. The number of intersection points made by the diagonals of a regular polygon. SIAM Journal on Discrete Mathematics, 11(1): 135–156, February 1998. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Plaut:2020:AEF

[PR20] Benjamin Plaut and Tim Roughgarden. Almost envyfreeness with general valuations. SIAM Journal on Discrete Mathematics, 34(2): 1039–1068, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Pralat:2013:SGF

[Pra13] Pawel Pralat. Sparse graphs are not flammable. SIAM Journal on Discrete Mathematics, 27(4):2157–2166, ???? 2013. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Palmer:2002:CCF

[PRR02] Edgar M. Palmer, Ronald C. Read, and Robert W. Robinson. Counting claw-free cubic graphs. SIAM Journal on Discrete Mathematics, 16(1):65–73, 2002. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/27477.

Poljak:1988:TRE

[PRS88] Svatopluk Poljak, Vojtěch Rödl, and Joel Spencer. Tournament ranking with expected profit in polynomial time. SIAM Journal on Discrete Mathematics, 1(3): 372–376, August 1988. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Proskurowski:1998:AAL

[PRS98] Andrzej Proskurowski, Frank Ruskey, and Malcolm Smith. Analysis of algorithms for listing equivalence classes of k-ary strings. SIAM Journal on Discrete Mathematics, 11(1):94–109, February 1998. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Parnas:2002:TBB

[PRS02] Michal Parnas, Dana Ron, and Alex Samorodnitsky. Testing basic Boolean formulae. SIAM Journal on Discrete Mathematics, 16 (1):20–46, 2002. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/40744.

Pebody:2003:FSP

[PRS03] L. Pebody, A. J. Radcliffe, and A. D. Scott. Finite subsets of the plane are 18-reconstructible. SIAM Journal on Discrete Mathematics, 16(2):262–275, 2003. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/39164.

Philip:2018:GPD

[PRS18] Geevarghese Philip, Ashutosh Rai, and Saket Saurabh. Generalized pseudoforest deletion: Algorithms and uniform kernel. SIAM Journal on Discrete Mathematics, 32 (2):882–901, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Peyerimhoff:2023:PCC

[PRS^{+}23] Norbert Peyerimhoff, Marc Roth, Johannes Schmitt, Jakob Stix, Alina Vdovina, and Philip Wellnitz. Parameterized counting and Cayley graph expanders. SIAM Journal on Discrete Mathematics, 37(2):405–486, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1479804.

Prymak:2023:NBH

[Pry23] Andriy Prymak. A new bound for Hadwiger’s covering problem in E^{3}. SIAM Journal on Discrete Mathematics, 37(1):17–24, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL

https://epubs.siam.org/
doi/10.1137/22M1490314.

Przybylo:2013:NDE

[Prz13] Jakub Przybylo. Neighbor distinguishing edge colorings via the combinatorial Nullstellensatz. SIAM Journal on Discrete Mathematics, 27(3): 1313–1322, ???? 2013. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Peer:1997:RIG

[PS97] Itsik Pe’er and Ron Shamir.

Realizing interval graphs with size and distance constraints. SIAM Journal on Discrete Mathematics, 10(4):662–687, November 1997. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Park:2017:CCM

[PS17] Sejun Park and Jinwoo Shin.

Convergence and correctness of max-product belief propagation for linear programming. SIAM Journal on Discrete Mathematics, 31(3): 2228–2246, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Pokrovskiy:2020:RGC

[PS20] Alexey Pokrovskiy and Benny Sudakov. Ramsey goodness of cycles. SIAM Journal on Discrete Mathematics, 34(3):

1884–1908, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Pirot:2021:FCN

[PS21] François Pirot and Jean-Sébastien Sereni. Fractional chromatic number, maximum degree, and girth. SIAM Journal on Discrete Mathematics, 35(4):2815–2843, ???? 2021. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Patel:2022:PTA

[PS22] Viresh Patel and Fabian Stroh. A polynomial-time algorithm to determine (almost) Hamiltonicity of dense regular graphs. SIAM Journal on Discrete Mathematics, 36(2):1363–1393, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1356191.

Pu:2008:CSC

[PSML08] Liqun Pu, Hao Shen, Jun Ma, and San Ling. Cycle systems in the complete bipartite graph plus a onefactor. SIAM Journal on Discrete Mathematics, 21(4): 1083–1092, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Pelsmajer:2009:SHT

[PSS09] Michael J. Pelsmajer, Marcus Schaefer, and Despina Stasi. Strong Hanani–Tutte on the projective plane. SIAM Journal on Discrete Mathematics, 23(3):1317–1323, ???? 2009. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Pelsmajer:2010:RIE

[PSŠ10] Michael J. Pelsmajer, Marcus Schaefer, and Daniel Štefankovič. Removing independently even crossings. SIAM Journal on Discrete Mathematics, 24(2):379–393, ???? 2010. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Patt-Shamir:2000:EAE

[PSTF00] Boaz Patt-Shamir, Yiannis Tsiounis, and Yair Frankel. Exact analysis of exact change: The k-payment problem. SIAM Journal on Discrete Mathematics, 13 (4):436–453, 2000. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/31811.

Pittel:2008:NFP

[PSV08] B. Pittel, L. Shepp, and E. Veklerov. On the number of fixed pairs in a random instance of the stable marriage problem. SIAM Journal on Discrete Mathematics, 21(4): 947–958, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Proskurowski:1996:PET

[PSW96] Andrzej Proskurowski, Maciej M. Sysło, and Paweł Winter. Plane embeddings of 2-trees and biconnected partial 2-trees. SIAM Journal on Discrete Mathematics, 9 (4):577–596, November 1996. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Phillips:1997:TSN

[PSW97] Cynthia Phillips, Clifford Stein, and Joel Wein. Task scheduling in networks. SIAM Journal on Discrete Mathematics, 10(4):573–598, November 1997. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Paat:2024:CNF

[PSWX24] Joseph Paat, Ingo Stallknecht, Zach Walsh, and Luze Xu. On the column number and forbidden submatrices for Δ-modular matrices. SIAM Journal on Discrete Mathematics, 38(1): 1–18, January 2024. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Petit:1990:TPR

[PT90] Jean-Luc Petit and Eric Térouanne. A theory of proportional representation. SIAM Journal on Discrete Mathematics, 3(1):116–139, February 1990. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Poljak:1994:BST

[PT94] Svatopluk Poljak and Zsolt Tuza. Bipartite subgraphs of triangle-free graphs. SIAM Journal on Discrete Mathematics, 7(2):307–313, May 1994. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Peng:2014:ITP

[PT14] Xing Peng and Craig Timmons. Infinite Turán problems for bipartite graphs. SIAM Journal on Discrete Mathematics, 28(2):702–710, ???? 2014. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Penev:2016:IHC

[PTT16] Irena Penev, Stéphan Thomassé, and Nicolas Trotignon. Isolating highly connected induced subgraphs. SIAM Journal on Discrete Mathematics, 30(1): 592–619, ???? 2016. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Pitassi:1995:CHC

[PU95] Toniann Pitassi and Alasdair Urquhart. The complexity of the Hajós calculus. SIAM Journal on Discrete Mathematics, 8(3):464–483, August 1995. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Pak:2010:RYT

[PV10a] Igor Pak and Ernesto Vallejo. Reductions of Young tableau bijections. SIAM Journal on Discrete Mathematics, 24(1): 113–145, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Paletta:2010:SNA

[PV10b] Giuseppe Paletta and Francesca Vocaturo. A short note on an advance in estimating the worst-case performance ratio of the MPS algorithm. SIAM Journal on Discrete Mathematics, 23(4):2198–2203, ???? 2010. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Panizzut:2022:TLC

[PV22] Marta Panizzut and Magnus Dehli Vigeland. Tropical lines on cubic surfaces. SIAM Journal on Discrete Mathematics, 36(1):383–410, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL

https://epubs.siam.org/
doi/10.1137/20M136520X.

Pineda-Villavicencio:2022:RMP

[PVS22] Guillermo Pineda-Villavicencio and Benjamin Schröter. Reconstructibility of matroid polytopes. SIAM Journal on Discrete Mathematics, 36 (1):490–508, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1401176.

Pineda-Villavicencio:2018:EDP

[PVUY18] Guillermo Pineda-Villavicencio, Julien Ugon, and David Yost. The excess degree of a polytope. SIAM Journal on Discrete Mathematics, 32(3): 2011–2046, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Pineda-Villavicencio:2022:LBT

[PVY22] Guillermo Pineda-Villavicencio and David Yost. The lower bound theorem for dpolytopes with 2d + 1 vertices. SIAM Journal on Discrete Mathematics, 36 (4):2920–2941, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M144832X.

Peleg:2002:HES

[PW02] David Peleg and Avishai Wool. How to be an efficient snoop, or the probe complexity of quorum systems. SIAM Journal on Discrete Mathematics, 15(3):416–433, 2002. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/34381.

Payne:2013:GPS

[PW13] Michael S. Payne and David R. Wood. On the general position subset selection problem. SIAM Journal on Discrete Mathematics, 27(4): 1727–1733, ???? 2013. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Pellicer:2018:PRT

[PW18] Daniel Pellicer and Gordon Ian Williams. Pyramids over regular 3-tori. SIAM Journal on Discrete Mathematics, 32(1):249–265, ???? 2018. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Pretzel:1990:CLG

[PY90] Oliver Pretzel and Dale Youngs. Cycle lengths and graph orientations. SIAM Journal on Discrete Mathematics, 3(4):544–553, November 1990. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Poznanovic:2009:CNT

[PY09] Svetlana Poznanović and Catherine Yan. Crossings and nestings of two edges in set partitions. SIAM Journal on Discrete Mathematics, 23(2): 787–804, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Paturi:1998:DPB

[PZ98] Ramamohan Paturi and Francis Zane. Dimension of projections in Boolean functions. SIAM Journal on Discrete Mathematics, 11 (4):624–632, November 1998. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Pike:2005:DCP

[PZ05] David A. Pike and Yubo Zou.

Decycling Cartesian products of two cycles. SIAM Journal on Discrete Mathematics, 19(3):651–663, 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/44016.

Pan:2010:MCC

[PZ10] Zhishi Pan and Xuding Zhu.

Multiple coloring of cone graphs. SIAM Journal on Discrete Mathematics, 24(4): 1515–1526, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v24/i4/p1515_s1.

Qu:2014:DPS

[QD14] Longjiang Qu and Cunsheng Ding. Dickson polynomials of the second kind that permute Z_{m}. SIAM Journal on Discrete Mathematics, 28(2): 722–735, ???? 2014. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Qureshi:2015:RAF

[QP15] Claudio Qureshi and Daniel Panario. Rédei actions on finite fields and multiplication map in cyclic group. SIAM Journal on Discrete Mathematics, 29(3):1486–1503, ???? 2015. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Quilodran:2010:JP

[Qui10] René Quilodrán. The joints problem in R^{n}. SIAM Journal on Discrete Mathematics, 23(4):2211–2213, ???? 2010. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Rabern:2006:GA

[Rab06] Landon Rabern. On graph associations. SIAM Journal on Discrete Mathematics, 20(2): 529–535, January 2006. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Rabern:2008:NRC

[Rab08] Landon Rabern. A note on Reed’s Conjecture. SIAM Journal on Discrete Mathematics, 22(2):820–827, ???? 2008. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Rahman:2016:FIP

[Rah16]  Mustazee Rahman.  Fac-

tor of IID percolation on trees. SIAM Journal on Discrete Mathematics, 30(4): 2217–2242, ???? 2016. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ramachandran:1990:MAT

[Ram90] Vijaya Ramachandran. A minimax arc theorem for reducible flow graphs. SIAM Journal on Discrete Mathematics, 3(4):554–560, November 1990. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ramachandramurthi:1997:SNO

[Ram97a] Siddharthan Ramachandramurthi. The structure and number of obstructions to treewidth. SIAM Journal on Discrete Mathematics, 10(1): 146–157, February 1997. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ramras:1997:CFO

[Ram97b] Mark Ramras. Congestionfree optimal routings of hypercube automorphisms. SIAM Journal on Discrete Mathematics, 10(2):201–208, May 1997. CODEN SJD-MEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ramras:1998:CFR

[Ram98] Mark Ramras. Congestionfree routings of linear complement permutations. SIAM Journal on Discrete Mathematics, 11(3):487–500, August 1998. CODEN SJD-MEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ramnath:2004:DDC

[Ram04]  Sarnath Ramnath.  Dy-

namic digraph connectivity hastens minimum sum-of-diameters clustering. SIAM Journal on Discrete Mathematics, 18(2):272–286, 2004. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/39609.

Ranto:2002:IFD

[Ran02] Kalle Ranto. Infinite families of 3-designs from Z_{4}-Goethals codes with block size 8. SIAM Journal on Discrete Mathematics, 15(3):289–304, 2002. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/38588.

Raychaudhuri:1994:FRC

[Ray94] Arundhati Raychaudhuri.

Further results on T -coloring and frequency assignment problems. SIAM Journal on Discrete Mathematics, 7 (4):605–613, November 1994. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Razborov:2010:HFV

[Raz10] Alexander A. Razborov. On 3-hypergraphs with forbidden 4-vertex configurations. SIAM Journal on Discrete Mathematics, 24(3):946–963, ???? 2010. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Raz:2020:UTB

[Raz20] Abigail Raz. Upper tail bounds for cycles. SIAM Journal on Discrete Mathematics, 34(3):1559–1586, ???? 2020. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Rispoli:1998:BDS

[RC98] Fred J. Rispoli and Steven Cosares. A bound of 4 for the diameter of the symmetric traveling salesman polytope. SIAM Journal on Discrete Mathematics, 11(3): 373–380, August 1998. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Rada:2018:NAE

[RC18] Miroslav Rada and Michal Cerný. A new algorithm for enumeration of cells of hyperplane arrangements and a comparison with Avis and Fukuda’s reverse search. SIAM Journal on Discrete Mathematics, 32(1):455–473, ???? 2018. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ray-Chaudhuri:1988:EDL

[RCS88] D. K. Ray-Chaudhuri and N. M. Singhi. On existence of t-designs with large v and λ. SIAM Journal on Discrete Mathematics, 1(1):98– 104, February 1988. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ramras:2011:AGJ

[RD11] Mark Ramras and Elizabeth Donovan. The automorphism group of a Johnson graph. SIAM Journal on Discrete Mathematics, 25(1): 267–270, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sidma/resource/
                                        1/sjdmec/v25/i1/p267_s1.

Reading:2008:CNP

[Rea08] Nathan Reading. Chains in the noncrossing partition lattice. SIAM Journal on Discrete Mathematics, 22(3): 875–886, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Reading:2015:NAD

[Rea15]  Nathan Reading.  Non-

crossing arc diagrams and canonical join representations. SIAM Journal on Discrete Mathematics, 29(2): 736–750, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Redlich:2017:PTC

[Red17] Amanda Redlich. A power-of-two-choices unbalanced allocation process. SIAM Journal on Discrete Mathematics, 31 (1):477–488, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Remila:2002:SSS

[Rém02]  Eric Rémila.  On the

structure of some spaces of tilings. SIAM Journal on Discrete Mathematics, 16(1):1–19, 2002. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/37329.

Ravazzi:2012:GRI

[RF12] Chiara Ravazzi and Fabio Fagnani. On the growth rate of the input-output weight distribution of convolutional encoders. SIAM Journal on Discrete Mathematics, 26(3): 1310–1345, ???? 2012. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Rhoades:2015:ADR

[Rho15] Brendon Rhoades. Alexander duality and rational associahedra. SIAM Journal on Discrete Mathematics, 29(1): 431–460, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Richter:2014:DRR

[Ric14] Christian Richter. A density result for random packings of unit disks. SIAM Journal on Discrete Mathematics, 28(1): 122–126, ???? 2014. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Rifa:1999:WOS

[Rif99] Josep Rifà. Well-ordered Steiner triple systems and 1-perfect partitions of the n-cube. SIAM Journal on Discrete Mathematics, 12 (1):35–47, February 1999. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/33072.

Riordan:1998:OED

[Rio98]  Oliver Riordan.  An or-

dering on the even discrete torus. SIAM Journal on Discrete Mathematics, 11(1): 110–127, February 1998. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Rizzi:2002:FFB

[Riz02] Romeo Rizzi. Finding 1factors in bipartite regular graphs and edge-coloring bipartite graphs. SIAM Journal on Discrete Mathematics, 15(3):283–288, 2002. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/35113.

Rais:1993:LDD

[RJS93] Bonita Rais, Philippe Jacquet, and Wojciech Szpankowski. Limiting distribution for the depth in Patricia tries. SIAM Journal on Discrete Mathematics, 6(2):197–213, May 1993. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Rokicki:2013:DRC

[RKDD13] Tomas Rokicki, Herbert Kociemba, Morley Davidson, and John Dethridge. The diameter of the Rubik’s cube group is twenty. SIAM Journal on Discrete Mathematics, 27(2):1082–1105, ???? 2013.

CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Rong:2021:CEH

[RLWY21] Guozhen Rong, Wenjun Li, Jianxin Wang, and Yongjie Yang. Cycle extendability of Hamiltonian strongly chordal graphs. SIAM Journal on Discrete Mathematics, 35(3): 2115–2128, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Richmond:1997:SNC

[RM97] Bruce Richmond and Donatella Merlini. Stirling numbers for complex arguments. SIAM Journal on Discrete Mathematics, 10(1): 73–82, February 1997. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ruskey:2001:NIP

[RMS01] F. Ruskey, C. R. Miers, and J. Sawada. The number of irreducible polynomials and Lyndon words with given trace. SIAM Journal on Discrete Mathematics, 14(2):240–245, 2001. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/36805.

Roche-Newton:2021:SPD

[RN21] Oliver Roche-Newton. Sums, products, and dilates on sparse graphs. SIAM Journal on Discrete Mathematics, 35 (1):194–204, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ramachandra:2023:TPB

[RN23] Arjun Kodagehalli Ramachandra and Karthik Natarajan. Tight probability bounds with pairwise independence. SIAM Journal on Discrete Mathematics, 37(2):516–555, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1408294.

Romik:2006:SPT

[Rom06] Dan Romik. Shortest paths in the Tower of Hanoi graph and finite automata. SIAM Journal on Discrete Mathematics, 20(3):610–622, January 2006. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ronyai:1992:GGF

[Rón92] Lajos Rónyai. Galois groups and factoring polynomials over finite fields. SIAM Journal on Discrete Mathematics, 5(3):345–365, August 1992. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Rosendahl:2009:CMV

[Ros09] Petri Rosendahl. On Cusick’s method and value sets of certain polynomials over finite fields. SIAM Journal on Discrete Mathematics, 23(1): 333–343, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Rozhon:2019:LAE

[Roz19] Václav Rozhon. A local approach to the Erdős–Sós conjecture. SIAM Journal on Discrete Mathematics, 33(2): 643–664, ???? 2019. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ruckenstein:2003:BLD

[RR03] Gitit Ruckenstein and Ron M. Roth. Bounds on the listdecoding radius of Reed– Solomon codes. SIAM Journal on Discrete Mathematics, 17(2):171–195, 2003. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/39550.

Rao:2018:ATC

[RR18] Michaël Rao and Matthieu Rosenfeld. Avoiding two consecutive blocks of same size and same sum over Z^{2}. SIAM Journal on Discrete Mathematics, 32(4):2381–2397, ???? 2018. CODEN SJDMEC.

ISSN 0895-4801 (print), 1095-7146 (electronic).

Rodl:2007:RPR

[RRS07] Vojtĕch Rödl, Andrzej Ruciński, and Mathias Schacht. Ramsey properties of random k-partite, k-uniform hypergraphs. SIAM Journal on Discrete Mathematics, 21(2): 442–460, ???? 2007. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Redman:2022:SMS

[RRW22] Maximus Redman, Lauren Rose, and Raphael Walker. A small maximal Sidon set in Z_{2}^{n}. SIAM Journal on Discrete Mathematics, 36 (3):1861–1867, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1454663.

Ruskey:1993:HCE

[RS93] Frank Ruskey and Carla Savage. Hamilton cycles that extend transposition matchings in Cayley graphs of S_{n}. SIAM Journal on Discrete Mathematics, 6(1):152–166, February 1993. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Rodl:2008:RMG

[RS08] V. Rödl and M. Siggers.

On Ramsey minimal graphs. SIAM Journal on Discrete Mathematics, 22(2):467–488, ???? 2008. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ravi:2010:AAM

[RS10] R. Ravi and Amitabh Sinha. Approximation algorithms for multicommodity facility location problems. SIAM Journal on Discrete Mathematics, 24(2):538–551, ???? 2010. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Rubinfeld:2011:STA

[RS11] Ronitt Rubinfeld and Asaf Shapira. Sublinear time algorithms. SIAM Journal on Discrete Mathematics, 25(4): 1562–1588, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sidma/resource/
                                        1/sjdmec/v25/i4/p1562_s1.

Rautenbach:2014:TLP

[RS14] Dieter Rautenbach and Jean-Sébastien Sereni. Transversals of longest paths and cycles. SIAM Journal on Discrete Mathematics, 28(1): 335–341, ???? 2014. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Royle:2015:LBT

[RS15] Gordon F. Royle and Alan D. Sokal. Linear bound in terms of maxmaxflow for the chromatic roots of series-parallel graphs. SIAM Journal on Discrete Mathematics, 29(4): 2117–2159, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Rudnev:2016:UKQ

[RS16] Misha Rudnev and J. M.

Selig. On the use of the Klein quadric for geometric incidence problems in two dimensions. SIAM Journal on Discrete Mathematics, 30(2): 934–954, ???? 2016. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Rhodes:2022:UBM

[RS22] John Rhodes and Anne Schilling. Upper bounds on mixing time of finite Markov chains. SIAM Journal on Discrete Mathematics, 36 (4):3031–3057, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1421994.

Regev:2017:IGL

[RSD17] Oded Regev and Noah Stephens-Davidowitz. An inequality for Gaussians on lattices. SIAM Journal on Discrete Mathematics, 31(2): 749–757, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ravi:1996:STS

[RSM^{+}96] R. Ravi, R. Sundaram, M. V. Marathe, D. J. Rosenkrantz, and S. S. Ravi. Spanning trees — short or small. SIAM Journal on Discrete Mathematics, 9(2):178–200, May 1996. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ragde:1988:PCE

[RSSW88] Prabhakar Ragde, William Steiger, Endre Szemerédi, and Avi Wigderson. The parallel complexity of element distinctness is Ω\sqrt{log n}. SIAM Journal on Discrete Mathematics, 1(3):399–410, August 1988. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Restrepo:2014:PTG

[RSV^{+}14] Ricardo Restrepo, Daniel Stefankovic, Juan C. Vera, Eric Vigoda, and Linji Yang. Phase transition for Glauber dynamics for independent sets on regular trees. SIAM Journal on Discrete Mathematics, 28(2):835–861, ???? 2014. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Radcliffe:2005:RTF

[RSW05] A. J. Radcliffe, A. D. Scott, and E. L. Wilmer. Reversals and transpositions over finite alphabets. SIAM Journal on Discrete Mathematics, 19(1):224–244, 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/43355.

Ruskey:2012:BSF

[RSW12] Frank Ruskey, Joe Sawada, and Aaron Williams. De Bruijn sequences for fixedweight binary strings. SIAM Journal on Discrete Mathematics, 26(2):605–617, ???? 2012. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Rubinov:1998:SNO

[RT98] Anatoly R. Rubinov and Vadim G. Timkovsky. String noninclusion optimization problems. SIAM Journal on Discrete Mathematics, 11(3): 456–467, August 1998. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Richerby:2009:GSC

[RT09] David Richerby and Dimitrios M. Thilikos. Graph searching in a crime wave. SIAM Journal on Discrete Mathematics, 23(1):349–368, ???? 2009. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Richerby:2011:SVL

[RT11] David Richerby and Dimitrios M. Thilikos. Searching for a visible, lazy fugitive. SIAM Journal on Discrete Mathematics, 25(2): 497–513, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sidma/resource/
                                        1/sjdmec/v25/i2/p497_s1.

Riquelme:2018:STP

[RT18] Edgardo Riquelme and Nicolas Thériault. Symbolic trisection polynomials for genus 2 curves in odd characteristic. SIAM Journal on Discrete Mathematics, 32(4): 2421–2440, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Radhakrishnan:2000:BDE

[RTS00] Jaikumar Radhakrishnan and Amnon Ta-Shma. Bounds for dispersers, extractors, and depth-two superconcentrators. SIAM Journal on Discrete Mathematics, 13(1):2–24, 2000. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/32950.

Rubinstein:1997:STT

[RTW97] J. H. Rubinstein, D. A. Thomas, and N. C. Wormald.

Steiner trees for terminals constrained to curves. SIAM Journal on Discrete Mathematics, 10(1):1–17, February 1997. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Raghavachari:1999:AAM

[RV99] Balaji Raghavachari and Jeyakesavan Veerasamy. A 3/2-approximation algorithm for the mixed postman problem. SIAM Journal on Discrete Mathematics, 12(4): 425–433, November 1999. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/33145.

Rodl:1989:RTT

[RW89] Vojtěch Rödl and Peter Winkler. A Ramsey-type theorem for orderings of a graph. SIAM Journal on Discrete Mathematics, 2(3): 402–406, August 1989. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Rok:2019:OGB

[RW19] Alexandre Rok and Bartosz Walczak. Outerstring graphs are χ-bounded. SIAM Journal on Discrete Mathematics, 33(4):2181–2199, ???? 2019. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Rawlins:1988:HPR

[RWW88] Gregory J. E. Rawlins, Peter Widmayer, and Derick Wood. Hole problems for rectangles in the plane. SIAM Journal on Discrete Mathematics, 1(1):86–97, February 1988. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Roberts:1988:OSC

[RX88] Fred S. Roberts and Yonghua Xu. On the optimal strongly connected orientations of city street graphs. I. Large grids. SIAM Journal on Discrete Mathematics, 1(2): 199–222, May 1988. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Remmel:1991:SRH

[RY91] J. B. Remmel and M. Yang. Special rim hook tabloids and some new multiplicity-free Sseries. SIAM Journal on Discrete Mathematics, 4(2): 253–274, May 1991. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Rao:2022:AEG

[RY22] Anup Rao and Amir Yehudayoff. Anticoncentration and the exact gap-Hamming problem. SIAM Journal on Discrete Mathematics, 36 (2):1071–1092, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1435288.

Ren:2023:GBR

[RY23] Xin Ren and Kohji Yanagawa. Gröbner bases of radical Li–Li type ideals associated with partitions. SIAM Journal on Discrete Mathematics, 37(4):2382–2396, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL

https://epubs.siam.org/
doi/10.1137/23M1547627.

Ryan:1996:IH

[Rya96] Jennifer Ryan. IIS-hypergraphs.

SIAM Journal on Discrete Mathematics, 9(4):643–653, November 1996. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ryan:2007:TSP

[Rya07] Jeffrey A. Ryan. A twoset problem on coloring the integers. SIAM Journal on Discrete Mathematics, 21(3): 731–736, ???? 2007. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Robins:2005:TBG

[RZ05] Gabriel Robins and Alexander Zelikovsky. Tighter bounds for graph Steiner tree approximation. SIAM Journal on Discrete Mathematics, 19(1):122–134, 2005.

CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/39315.

Sherali:1990:HRB

[SA90] Hanif D. Sherali and Warren P. Adams. A hierarchy of relaxations between the continuous and convex hull representations for zero-one programming problems. SIAM Journal on Discrete Mathematics, 3(3):411–430, August 1990. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Saad:1993:CFI

[Saa93] Rachid Saad. Complexity of the forwarding index problem. SIAM Journal on Discrete Mathematics, 6(3):418– 427, August 1993. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Safe:2021:CCO

[Saf21]  Martín D. Safe.  Cir-

cularly compatible ones, D-circularity, and proper circular-arc bigraphs. SIAM Journal on Discrete Mathematics, 35(2):707–751, ???? 2021. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Sagan:1988:NIS

[Sag88] Bruce E. Sagan. A note on independent sets in trees. SIAM Journal on Discrete Mathematics, 1(1):105–108, February 1988. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Saks:1989:RNP

[Sak89] Michael Saks. A robust noncryptographic protocol for collective coin flipping. SIAM Journal on Discrete Mathematics, 2(2):240–244, May 1989. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Sakai:1994:LCG

[Sak94]  Denise Sakai.  Labelling

chordal graphs: Distance two condition. SIAM Journal on Discrete Mathematics, 7(1): 133–140, February 1994. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Sanders:1996:LRT

[San96] Daniel P. Sanders. On linear recognition of tree-width at most four. SIAM Journal on Discrete Mathematics, 9(1): 101–117, February 1996. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Sarangarajan:1997:LBA

[Sar97] A. Sarangarajan. A lower bound for adjacencies on the traveling salesman polytope. SIAM Journal on Discrete Mathematics, 10(3):

431–435, August 1997. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Savage:1990:GPD

[Sav90] Carla D. Savage. Generating permutations with kdifferences. SIAM Journal on Discrete Mathematics, 3 (4):561–573, November 1990. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Savage:1997:SCG

[Sav97]  Carla Savage.  A sur-

vey of combinatorial Gray codes. SIAM Review, 39 (4):605–629, December 1997. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29527.  See

note [Naa01].

Savitsky:2014:EPS

[Sav14] Thomas J. Savitsky. Enumeration of 2-polymatroids on up to seven elements. SIAM Journal on Discrete Mathematics, 28(4):1641–1650, ???? 2014. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Sawada:2002:FAG

[Saw02] Joe Sawada. A fast algorithm for generating nonisomorphic chord diagrams. SIAM Journal on Discrete Mathematics, 15(4):546–561, 2002.

CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/37797.

Sawada:2007:SGC

[Saw07] J. Sawada. A simple Gray code to list all minimal signed binary representations. SIAM Journal on Discrete Mathematics, 21(1):16–25, ???? 2007. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Siu:1991:PTC

[SB91] Kai-Yeung Siu and Jehoshua Bruck. On the power of threshold circuits with small weights. SIAM Journal on Discrete Mathematics, 4(3): 423–435, August 1991. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Szwarcfiter:1994:CGC

[SB94] Jayme L. Szwarcfiter and Claudson F. Bornstein. Clique graphs of chordal and path graphs. SIAM Journal on Discrete Mathematics, 7(2): 331–336, May 1994. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Stolk:2010:AFD

[SB10] Arjen Stolk and K. Joost Batenburg. An algebraic framework for discrete tomography: Revealing the structure of dependencies. SIAM Journal on Discrete Mathematics, 24(3):1056– 1079, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Sidiropoulos:2019:AAL

[SBD^{+}19] Anastasios Sidiropoulos, Mihai Badoiu, Kedar Dhamdhere, Anupam Gupta, Piotr Indyk, Yuri Rabinovich, Harald Racke, and R. Ravi. Approximation algorithms for low-distortion embeddings into low-dimensional spaces. SIAM Journal on Discrete Mathematics, 33(1): 454–473, ???? 2019. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Sarkar:2017:UBS

[SC17] Kaushik Sarkar and Charles J. Colbourn. Upper bounds on the size of covering arrays. SIAM Journal on Discrete Mathematics, 31(2): 1277–1293, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Scarabotti:2003:FAC

[Sca03] Fabio Scarabotti. Fourier analysis of a class of finite Radon transforms. SIAM Journal on Discrete Mathematics, 16(4):545–554, 2003. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/41639.

Scarabotti:2005:DST

[Sca05] Fabio Scarabotti. The discrete sine transform and the spectrum of the finite qary tree. SIAM Journal on Discrete Mathematics, 19 (4):1004–1010, January 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Scheinerman:1991:NPG

[Sch91a] Edward R. Scheinerman. A note on planar graphs and circle orders. SIAM Journal on Discrete Mathematics, 4(3): 447–450, August 1991. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Schrijver:1991:EDH

[Sch91b] A. Schrijver. Edge-disjoint homotopic paths in straightline planar graphs. SIAM Journal on Discrete Mathematics, 4(1):130–138, February 1991. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Scheinerman:1992:GCN

[Sch92] Edward R. Scheinerman.

Generalized chromatic numbers of random graphs. SIAM Journal on Discrete Mathematics, 5(1):74–80, February 1992. CODEN SJDMEC.

ISSN 0895-4801 (print), 1095-7146 (electronic).

Schirokauer:2002:SFF

[Sch02a] Oliver Schirokauer. The special function field sieve. SIAM Journal on Discrete Mathematics, 16(1):81–98, 2002. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/37266.

Schrijver:2002:SPB

[Sch02b] Alexander Schrijver. Strong T-perfection of bad-K_{4}-free graphs. SIAM Journal on Discrete Mathematics, 15 (3):403–415, 2002. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/40110.

Schachtel:2004:EWR

[Sch04] Gabriel A. Schachtel. Exact word-run statistics in random orderings. SIAM Journal on Discrete Mathematics, 18(1):141–158, 2004. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/39057.

Schweig:2009:CED

[Sch09] Jay Schweig. A convexear decomposition for rankselected subposets of supersolvable lattices. SIAM Journal on Discrete Mathematics, 23(2):1009–1022, ???? 2009. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Schulze:2010:SSC

[Sch10] Bernd Schulze. Symmetry as a sufficient condition for a finite flex. SIAM Journal on Discrete Mathematics, 24(4): 1291–1312, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Schaefer:2022:HTH

[Sch22] Marcus Schaefer. Hanani– Tutte and hierarchical partial planarity. SIAM Journal on Discrete Mathematics, 36(4):2959–2974, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1464749.

Solomon:2014:BDD

[SE14] Shay Solomon and Michael Elkin. Balancing degree, diameter, and weight in Euclidean spanners. SIAM Journal on Discrete Mathematics, 28(3):1173–1198, ???? 2014. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Seiden:2001:OOA

[Sei01] Steven S. Seiden. An optimal online algorithm for bounded space variable-sized bin packing. SIAM Journal on Discrete Mathematics, 14(4):458–470, 2001. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/36994.

Sendrier:1997:DH

[Sen97] Nicolas Sendrier. On the dimension of the hull. SIAM Journal on Discrete Mathematics, 10(2):282–293, May 1997. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Seress:1988:QGC

[Ser88] Ákos Seress. Quick gossiping by conference calls. SIAM Journal on Discrete Mathematics, 1(1):109–120, February 1988. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Servatius:1989:BP

[Ser89] Brigitte Servatius. Birigidity in the plane. SIAM Journal on Discrete Mathematics, 2 (4):582–589, November 1989. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Setiabrata:2021:FRP

[Set21] Linus Setiabrata. Faces of root polytopes. SIAM Journal on Discrete Mathematics, 35(3):2093–2114, ???? 2021. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Sommer:2009:FCL

[SFS09] Naftali Sommer, Meir Feder, and Ofir Shalvi. Finding the closest lattice point by iterative slicing. SIAM Journal on Discrete Mathematics, 23 (2):715–731, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Shangguan:2016:SHF

[SG16] Chong Shangguan and Gennian Ge. Separating hash families: a Johnson-type bound and new constructions. SIAM Journal on Discrete Mathematics, 30(4): 2243–2264, ???? 2016. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

ski:2024:GRP

[sG24] Sebastian Babi ski and Andrzej Grzesik. Graphs without a rainbow path of length 3. SIAM Journal on Discrete Mathematics, 38(1): 629–644, February 2024. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Sikiric:2020:VCC

[SGM20] Mathieu Dutour Sikirić, Alexey Garber, and Alexander Magazinov. On the Voronoi conjecture for combinatorially Voronoi parallelohedra in dimension 5. SIAM Journal on Discrete Mathematics, 34(4):2481–2501, ???? 2020. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Shaw:2013:TIP

[Sha13] Kristin M. Shaw. A tropical intersection product in matroidal fans. SIAM Journal on Discrete Mathematics, 27 (1):459–491, ???? 2013. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Shakan:2020:LGD

[Sha20] George Shakan. A large gap in a dilate of a set. SIAM Journal on Discrete Mathematics, 34(4):2553–2555, ???? 2020. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Shangguan:2023:DTD

[Sha23]  Chong Shangguan.  De-

generate Turán densities of sparse hypergraphs II: A solution to the Brown–Erdős– Sós problem for every uniformity. SIAM Journal on Discrete Mathematics, 37 (3):1920–1929, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1541290.

Shen:2018:PIL

[She18] Bobby Shen. Parametrizing an integer linear program by an integer. SIAM Journal on Discrete Mathematics, 32(1): 173–190, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Shioura:2012:NSJ

[Shi12] Akiyoshi Shioura. Neighbor systems, jump systems, and bisubmodular polyhedra. SIAM Journal on Discrete Mathematics, 26(1): 114–144, ???? 2012. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Shparlinski:2010:DOK

[Shp10] Igor E. Shparlinski. On the distribution of orbits of PGL_{2}(q) in F_{q^{n}} and the Klapper Conjecture. SIAM Journal on Discrete Mathematics, 23(4):2093–2099, ???? 2010. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Shparlinski:2013:ADS

[Shp13] Igor E. Shparlinski. Additive decompositions of subgroups of finite fields. SIAM Journal on Discrete Mathematics, 27(4):1870–1879, ???? 2013. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Shparlinski:2015:CGG

[Shp15] Igor E. Shparlinski. Cayley graphs generated by small degree polynomials over finite fields. SIAM Journal on Discrete Mathematics, 29(1): 376–381, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Shparlinski:2022:MPH

[Shp22] Igor E. Shparlinski. Multiplicative properties of Hilbert cubes. SIAM Journal on Discrete Mathematics, 36 (2):1064–1070, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1470396.

Schaller:2023:OFG

[SHS23] David Schaller, Marc Hellmuth, and Peter F. Stadler. Orientation of Fitch graphs and reconciliation-free inference of horizontal gene transfer in gene trees. SIAM Journal on Discrete Mathematics, 37(3):2172–2207, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M150736X.

Shukla:2023:VRC

[Shu23] Samir Shukla. On Vietoris– Rips complexes (with scale 3) of hypercube graphs. SIAM Journal on Discrete Mathematics, 37(3):1472–1495, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL

https://epubs.siam.org/
doi/10.1137/22M1481440.

Sidorenko:2018:EPH

[Sid18] Alexander Sidorenko. Extremal problems on the hypercube and the codegree Turán density of complete rgraphs. SIAM Journal on Discrete Mathematics, 32(4): 2667–2674, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Siggers:2010:NPC

[Sig10] Mark H. Siggers. A new proof of the H-coloring dichotomy. SIAM Journal on Discrete Mathematics, 23(4): 2204–2210, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Simon:1990:ASC

[Sim90] Hans Ulrich Simon. On approximate solutions for combinatorial optimization problems. SIAM Journal on Discrete Mathematics, 3(2): 294–310, May 1990. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Simson:2013:CGC

[Sim13] Daniel Simson. A Coxeter– Gram classification of positive simply laced edgebipartite graphs. SIAM Journal on Discrete Mathematics, 27(2):827–854, ???? 2013. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Sissokho:2021:GMS

[Sis21] Papa A. Sissokho. Geometry of the minimal solutions of a linear Diophantine equation. SIAM Journal on Discrete Mathematics, 35(3): 1952–1963, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Schulze:2015:IRS

[SiT15] Bernd Schulze and Shin ichi Tanigawa. Infinitesimal rigidity of symmetric bar-joint frameworks. SIAM Journal on Discrete Mathematics, 29 (3):1259–1286, ???? 2015. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Sivashankar:2023:EUR

[Siv23]  Varun Sivashankar.  Ex-

tremal uniquely resolvable multisets. SIAM Journal on Discrete Mathematics, 37 (1):268–278, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1481129.

ski:2024:PVC

[sJS24] Marcin Bria ski, Gwenaël Joret, and Micha T. Seweryn. Pathwidth vs cocircumference. SIAM Journal on Discrete Mathematics, 38(1): 857–866, February 2024. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Skutella:2016:NRL

[Sku16] Martin Skutella. A note on the ring loading problem. SIAM Journal on Discrete Mathematics, 30(1): 327–342, ???? 2016. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Sherali:1995:SSL

[SL95] Hanif D. Sherali and Youngho Lee. Sequential and simultaneous liftings of minimal cover inequalities for generalized upper bound constrained knapsack polytopes. SIAM Journal on Discrete Mathematics, 8(1):133–153, February 1995. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Shen:1996:GAC

[SL96] Yinan N. Shen and Fabrizio Lombardi. Graph algorithms for conformance testing using the rural Chinese postman tour. SIAM Journal on Discrete Mathematics, 9(4):511– 528, November 1996. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Slivkins:2010:PTE

[Sli10] Aleksandrs Slivkins. Parameterized tractability of edgedisjoint paths on directed acyclic graphs. SIAM Journal on Discrete Mathematics, 24 (1):146–157, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Smith:2001:PSP

[Smi01] Gregory L. Smith. Partitions and (m and n) sums of products — two cell partition. SIAM Journal on Discrete Mathematics, 14 (3):356–369, 2001. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/34901.

Sakashita:2009:MTP

[SMNF09] Mariko Sakashita, Kazuhisa Makino, Hiroshi Nagamochi, and Satoru Fujishige. Minimum transversals in posimodular systems. SIAM Journal on Discrete Mathematics, 23(2):858–871, ???? 2009. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Smorodinsky:2007:CNG

[Smo07] Shakhar Smorodinsky. On the chromatic number of geometric hypergraphs. SIAM Journal on Discrete Mathematics, 21(3):676–687, ???? 2007. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Sales:2008:PQP

[SMR08] Cláudia Linhares Sales, Frédéric Maffray, and Bruce Reed. On planar quasi-parity graphs. SIAM Journal on Discrete Mathematics, 22(1):329–347, ???? 2008. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Sniady:2014:RSK

[Śni14] Piotr Śniady. Robinson– Schensted–Knuth algorithm, Jeu de Taquin, and Kerov– Vershik measures on infinite tableaux. SIAM Journal on Discrete Mathematics, 28(2): 598–630, ???? 2014. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Snook:2013:CDP

[Sno13] Michael Snook. The complexity of deletion problems for matroids. SIAM Journal on Discrete Mathematics, 27(1): 402–421, ???? 2013. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Solymosi:2006:DAL

[Sol06] József Solymosi. Dense arrangements are locally very dense. I. SIAM Journal on Discrete Mathematics, 20(3): 623–627, January 2006. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Solomon:2012:MSD

[Sol12] Shay Solomon. The MST of symmetric disk graphs (in arbitrary metric spaces) is light. SIAM Journal on Discrete Mathematics, 26(1): 250–262, ???? 2012. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Soma:2014:FDA

[Som14] Tasuku Soma. Fast deterministic algorithms for matrix completion problems. SIAM Journal on Discrete Mathematics, 28(1):490–502, ???? 2014. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Soto:2015:IAM

[Sot15] José A. Soto. Improved analysis of a max-cut algorithm based on spectral partitioning. SIAM Journal on Discrete Mathematics, 29(1): 259–268, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Simoes-Pereira:1988:OCG

[SP88] J. M. S. Simões-Pereira. An optimality criterion for graph embeddings of metrics. SIAM Journal on Discrete Mathematics, 1(2):223–229, May 1988. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Spacapan:2007:OLT

[Špa07] Simon Špacapan. Optimal Lee-type local structures in Cartesian products of cycles and paths. SIAM Journal on Discrete Mathematics, 21(3): 750–762, ???? 2007. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Speyer:2008:TLS

[Spe08] David E. Speyer. Tropical linear spaces. SIAM Journal on Discrete Mathematics, 22(4): 1527–1558, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Spinrad:1989:PTS

[Spi89] Jeremy Spinrad. Prime testing for the split decomposition of a graph. SIAM Journal on Discrete Mathematics, 2(4):590–599, November 1989. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Spinrad:1995:NFM

[Spi95] Jeremy P. Spinrad. Nonredundant 1’s in Γ-free matrices. SIAM Journal on Discrete Mathematics, 8(2): 251–257, May 1995. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Spink:2019:OSC

[Spi19] Hunter Spink. Orthogonal symmetric chain decompositions of hypercubes. SIAM Journal on Discrete Mathematics, 33(2):910–932, ???? 2019. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Spiro:2023:IHM

[Spi23] Sam Spiro. On t-intersecting hypergraphs with minimum positive codegrees. SIAM Journal on Discrete Mathematics, 37(2):675–683, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL

https://epubs.siam.org/
doi/10.1137/21M1465998.

Sprague:1994:ABI

[Spr94]  Alan P. Sprague.  An

O(n log n) algorithm for bandwidth of interval graphs. SIAM Journal on Discrete Mathematics, 7(2):213–220, May 1994. CODEN SJD-MEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Siu:1994:ODT

[SR94] Kai-Yeung Siu and Vwani P.

Roychowdhury. On optimal depth threshold circuits for multiplication and related problems. SIAM Journal on Discrete Mathematics, 7(2): 284–292, May 1994. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Sugihara:1989:OAP

[SS89] Kazuo Sugihara and Ichiro Suzuki. Optimal algorithms for a pursuit-evasion problem in grids. SIAM Journal on Discrete Mathematics, 2(1): 126–143, February 1989. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Shapiro:1991:BPS

[SS91] Louis Shapiro and A. B.

Stephens. Bootstrap percolation, the Schroder numbers, and the N-kings problem. SIAM Journal on Discrete Mathematics, 4(2): 275–280, May 1991. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Sen:1994:IDG

[SS94] M. Sen and B. K. Sanyal. Indifference digraphs: a generalization of indifference graphs and semiorders. SIAM Journal on Discrete Mathematics, 7(2):157–165, May 1994. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Snyder:1995:EAD

[SS95] Timothy Law Snyder and J. Michael Steele. Equidistribution in all dimensions of worst-case point sets for the traveling salesman problem. SIAM Journal on Discrete Mathematics, 8(4):678– 683, November 1995. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Steiner:2000:SRI

[SS00] George Steiner and Paul Stephenson. Subset-restricted interchange for dynamic min-max scheduling problems. SIAM Journal on Discrete Mathematics, 13 (4):419–435, 2000. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/34341.

Schulz:2002:SUM

[SS02a] Andreas S. Schulz and Martin Skutella. Scheduling unrelated machines by randomized rounding. SIAM Journal on Discrete Mathematics, 15(4):450–469, 2002. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/35707.

Steel:2002:IRF

[SS02b] Michael A. Steel and László A.

Székely. Inverting random functions II: Explicit bounds for discrete maximum likelihood estimation, with applications. SIAM Journal on Discrete Mathematics, 15(4):562–575, 2002. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/38790.

Shachnai:2004:FLI

[SS04] Hadas Shachnai and Aravind Srinivasan. Finding large independent sets in graphs and hypergraphs. SIAM Journal on Discrete Mathematics, 18(3):488–500, 2004. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/41973.

Schaefer:2005:SGI

[SS05] Marcus Schaefer and Daniel Stefankovic. Solvability of graph inequalities. SIAM Journal on Discrete Mathematics, 19(3):728–743, 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/36065.

Solymosi:2008:EIT

[SS08] József Solymosi and Konrad J. Swanepoel. Elementary incidence theorems for complex numbers and quaternions. SIAM Journal on Discrete Mathematics, 22(3): 1145–1148, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Soprunov:2009:TSC

[SS09] Ivan Soprunov and Jenya Soprunova. Toric surface codes and Minkowski length of polygons. SIAM Journal on Discrete Mathematics, 23 (1):384–400, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Soprunov:2010:BTC

[SS10] Ivan Soprunov and Jenya Soprunova. Bringing toric codes to the next dimension. SIAM Journal on Discrete Mathematics, 24(2):655–665, ???? 2010. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Safari:2011:CFA

[SS11a] MohammadAli Safari and Mohammad R. Salavatipour. A constant factor approximation for minimum λedge-connected k-subgraph with metric costs. SIAM Journal on Discrete Mathematics, 25(3):1089–1102, ???? 2011. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.

siam.org/sidma/resource/
1/sjdmec/v25/i3/p1089_s1.

Scott:2011:BCR

[SS11b] Alex Scott and Benny Sudakov. A bound for the Cops and Robbers Problem. SIAM Journal on Discrete Mathematics, 25(3): 1438–1442, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v25/i3/p1438_s1.

Sawhney:2018:SCN

[SS18a] Mehtaab Sawhney and David Stoner. On symmetric but not cyclotomic numerical semigroups. SIAM Journal on Discrete Mathematics, 32 (2):1296–1304, ???? 2018. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Seacrest:2018:LBT

[SS18b] Tyler Seacrest and Francis Edward Su. A lower bound technique for triangulations of simplotopes. SIAM Journal on Discrete Mathematics, 32(1):1–28, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Schoen:2019:UBW

[SS19] Tomasz Schoen and Ilya D.

Shkredov. An upper bound for weak B_{k}-sets. SIAM Journal on Discrete Mathematics, 33(2):837–844, ???? 2019. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Shkredov:2021:UCA

[SS21] Ilya Shkredov and Jozsef Solymosi. The uniformity conjecture in additive combinatorics. SIAM Journal on Discrete Mathematics, 35(1): 307–321, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Sundaram:1994:TCA

[SSR94] Ravi Sundaram, Karan Sher Singh, and C. Pandu Rangan. Treewidth of circulararc graphs. SIAM Journal on Discrete Mathematics, 7 (4):647–655, November 1994. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Schmidt:1995:CHB

[SSS95] Jeanette P. Schmidt, Alan Siegel, and Aravind Srinivasan. Chernoff–Hoeffding bounds for applications with limited independence. SIAM Journal on Discrete Mathematics, 8(2):223–250, May 1995. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Shioura:2013:SOA

[SSS13] Akiyoshi Shioura, Natalia V. Shakhlevich, and Vitaly A.

Strusevich. A submodular optimization approach to bicriteria scheduling problems with controllable processing times on parallel machines. SIAM Journal on Discrete Mathematics, 27(1): 186–204, ???? 2013. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Scott:2022:PPV

[SSS22] Alex Scott, Paul Seymour, and Sophie Spirkl. Pure pairs VI: Excluding an ordered tree. SIAM Journal on Discrete Mathematics, 36(1):170–187, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1368331.

Scott:2024:PPI

[SSS24] Alex Scott, Paul Seymour, and Sophie T. Spirkl. Pure pairs. IX. Transversal trees. SIAM Journal on Discrete Mathematics, 38(1):645–667, February 2024. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Santoro:1989:GCP

[SSSU89] Nicola Santoro, Jeffrey B.

Sidney, Stuart J. Sidney, and Jorge Urrutia. Geometric containment and partial orders. SIAM Journal on Discrete Mathematics, 2(2):

245–254, May 1989. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Sloan:2008:TDL

[SST08] Robert H. Sloan, Balázs Szörényi, and György Turán. On k-term DNF with the largest number of prime implicants. SIAM Journal on Discrete Mathematics, 21(4): 987–998, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Schulze:2022:FCE

[SST22] Bernd Schulze, Hattie Serocold, and Louis Theran. Frameworks with coordinated edge motions. SIAM Journal on Discrete Mathematics, 36(4):2602–2618, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/20M1377539.

Schrijver:1998:RLP

[SSW98] Alexander Schrijver, Paul Seymour, and Peter Winkler. The ring loading problem. SIAM Journal on Discrete Mathematics, 11(1):1– 14, February 1998. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Shioura:2007:PTA

[ST07] Akiyoshi Shioura and Ken’ichiro Tanaka. Polynomial-time algorithms for linear and convex optimization on jump systems. SIAM Journal on Discrete Mathematics, 21(2): 504–522, ???? 2007. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Steffens:2010:CGT

[ST10] Reinhard Steffens and Thorsten Theobald. Combinatorics and genus of tropical intersections and Ehrhart Theory. SIAM Journal on Discrete Mathematics, 24(1):17–32, ???? 2010. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Scheidweiler:2013:LBC

[ST13] Robert Scheidweiler and Eberhard Triesch. A lower bound for the complexity of monotone graph properties. SIAM Journal on Discrete Mathematics, 27(1): 257–265, ???? 2013. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Schillewaert:2014:CCA

[ST14] Jeroen Schillewaert and Koen Thas. Construction and comparison of authentication codes. SIAM Journal on Discrete Mathematics, 28(1): 474–489, ???? 2014. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Song:2017:HCG

[ST17a] Zi-Xia Song and Brian Thomas. Hadwiger’s conjecture for graphs with forbidden holes. SIAM Journal on Discrete Mathematics, 31(3): 1572–1580, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Staden:2017:DSF

[ST17b] Katherine Staden and Andrew Treglown. On degree sequences forcing the square of a Hamilton cycle. SIAM Journal on Discrete Mathematics, 31(1):383–437, ???? 2017. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Shangguan:2020:NTE

[ST20a] Chong Shangguan and Itzhak Tamo. New Turán exponents for two extremal hypergraph problems. SIAM Journal on Discrete Mathematics, 34(4): 2338–2345, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Shangguan:2020:SHA

[ST20b] Chong Shangguan and Itzhak Tamo. Sparse hypergraphs with applications to coding theory. SIAM Journal on Discrete Mathematics, 34(3): 1493–1504, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Skotnica:2023:NHC

[ST23] Michael Skotnica and Martin Tancer. NP-hardness of computing PL geometric category in dimension 2. SIAM Journal on Discrete Mathematics, 37(3):2016–2029, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1523960.

Stachowiak:1992:HPG

[Sta92] Grzegorz Stachowiak. Hamilton paths in graphs of linear extensions for unions of posets. SIAM Journal on Discrete Mathematics, 5(2): 199–206, May 1992. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Stanton:2011:IBI

[Sta11] Brendon Stanton. Improved bounds for r-identifying codes of the hex grid. SIAM Journal on Discrete Mathematics, 25 (1):159–169, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v25/i1/p159_s1.

Steel:1988:DSD

[Ste88]  M. A. Steel.  Distribu-

tion of the symmetric difference metric on phylogenetic trees. SIAM Journal on Discrete Mathematics, 1(4):541– 551, November 1988. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Steel:2000:SCT

[Ste00] Mike Steel. Sufficient conditions for two tree reconstruction techniques to succeed on sufficiently long sequences. SIAM Journal on Discrete Mathematics, 14(1):36–48, 2000. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/34357.

Stephen:2007:MBE

[Ste07] Tamon Stephen. A majorization bound for the eigenvalues of some graph Laplacians. SIAM Journal on Discrete Mathematics, 21(2): 303–312, ???? 2007. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Stein:2010:EVS

[Ste10] Maya Stein. Ends and vertices of small degree in infinite minimally k-(Edge)- Connected graphs. SIAM Journal on Discrete Mathematics, 24(4):1584–1596, ???? 2010. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v24/i4/p1584_s1.

Steiner:2022:DCL

[Ste22] Raphael M. Steiner. Disjoint cycles with length constraints in digraphs of large connectivity or large minimum degree. SIAM Journal on Discrete Mathematics, 36(2):1343–1362, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1382398.

Stones:2012:FAT

[Sto12] Douglas S. Stones. Formulae for the Alon–Tarsi conjecture. SIAM Journal on Discrete Mathematics, 26(1): 65–70, ???? 2012. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v26/i1/p65_s1.

Strang:2020:SMV

[Str20]  Alexander Strang.  Solu-

tions to the minimum variance problem using Delaunay triangulation. SIAM Journal on Discrete Mathematics, 34(4):2510–2531, ???? 2020. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Sleator:1992:SEE

[STT92] Daniel D. Sleator, Robert E.

Tarjan, and William P. Thurston. Short encodings of evolving structures. SIAM Journal on Discrete Mathematics, 5(3):428–450, August 1992. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Sturmfels:1988:SAA

[Stu88] Bernd Sturmfels. Some applications of affine Gale diagrams to polytopes with few vertices. SIAM Journal on Discrete Mathematics, 1(1): 121–133, February 1988. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Stucky:2021:PUC

[Stu21]  Eric N. Stucky.  Parity-

unimodality and a cyclic sieving phenomenon for necklaces. SIAM Journal on Discrete Mathematics, 35(3): 2049–2069, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Srinivasan:2021:PDS

[STV21] Srikanth Srinivasan, Utkarsh Tripathi, and S. Venkitesh. On the probabilistic degrees of symmetric Boolean functions. SIAM Journal on Discrete Mathematics, 35(3): 2070–2092, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Serafini:1989:MMP

[SU89] Paolo Serafini and Walter Ukovich. A mathematical model for periodic scheduling problems. SIAM Journal on Discrete Mathematics, 2 (4):550–581, November 1989. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Sudakov:2008:RNS

[Sud08] Benny Sudakov. Ramsey numbers and the size of graphs. SIAM Journal on Discrete Mathematics, 21(4): 980–986, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Suk:2013:DTI

[Suk13] Andrew Suk. Density theorems for intersection graphs of t-monotone curves. SIAM Journal on Discrete Mathematics, 27(3):1323–1334, ???? 2013. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Sullivant:2005:SCT

[Sul05] Seth Sullivant. Small contingency tables with large gaps. SIAM Journal on Discrete Mathematics, 18 (4):787–793, 2005. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/44409.

Sullivant:2012:DNP

[Sul12] Seth Sullivant. The disentangling number for phylogenetic mixtures. SIAM Journal on Discrete Mathematics, 26(2): 856–859, ???? 2012. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Suzuki:2010:REM

[Suz10] Yusuke Suzuki. Re-embeddings of maximum 1-planar graphs. SIAM Journal on Discrete Mathematics, 24(4):1527– 1540, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sidma/resource/
                                        1/sjdmec/v24/i4/p1527_s1.

Stinson:1988:CAT

[SV88] D. R. Stinson and S. A. Vanstone. A combinatorial approach to threshold schemes. SIAM Journal on Discrete Mathematics, 1(2): 230–236, May 1988. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Sudeep:2008:MFD

[SV08] K. S. Sudeep and Sundar Vishwanathan. Matchedfactor d-domatic coloring of graphs. SIAM Journal on Discrete Mathematics, 21(4): 1071–1082, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Stefankovic:2011:FCM

[SV11] Daniel Stefankovic and Eric Vigoda. Fast convergence of Markov chain Monte Carlo algorithms for phylogenetic reconstruction with homogeneous data on closely related species. SIAM Journal on Discrete Mathematics, 25(3): 1194–1211, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v25/i3/p1194_s1.

Sudakov:2018:TRE

[SV18] Benny Sudakov and Pedro Vieira. Two remarks on eventown and oddtown problems. SIAM Journal on Discrete Mathematics, 32(1): 280–295, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Silvanus:2020:FSP

[SV20] Jannik Silvanus and Jens Vygen. Few sequence pairs suffice: Representing all rectangle placements. SIAM Journal on Discrete Mathematics, 34(4):2017–2032, ???? 2020. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Spiro:2021:RTP

[SV21] Sam Spiro and Jacques Verstraëte. Relative Turán problems for uniform hypergraphs. SIAM Journal on Discrete Mathematics, 35(3): 2170–2191, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Sviridenko:2003:MMN

[Svi03] M. Sviridenko. Makespan minimization in no-wait flow shops: a polynomial time approximation scheme. SIAM Journal on Discrete Mathematics, 16(2):313–322, 2003. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/37080.

Siersma:2008:NMG

[SvM08] D. Siersma and M. van Manen. The nine Morse generic tetrahedra. SIAM Journal on Discrete Mathematics, 22(2): 737–746, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Stinson:1998:CPC

[SW98] D. R. Stinson and R. Wei. Combinatorial properties and constructions of traceability schemes and frameproof codes. SIAM Journal on Discrete Mathematics, 11(1): 41–53, February 1998. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Servatius:1999:CPC

[SW99] Brigitte Servatius and Walter Whiteley. Constraining plane configurations in computer-aided design: Combinatorics of directions and lengths. SIAM Journal on Discrete Mathematics, 12 (1):136–153, February 1999. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/30734.

Shi:2001:SDW

[SW01] Weiping Shi and Douglas B.

West. Structural diagnosis of wiring networks: Finding connected components of unknown subgraphs. SIAM Journal on Discrete Mathematics, 14(4):510–523, 2001. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/37128.

Saliola:2004:CPC

[SW04] Franco Saliola and Walter Whiteley. Constraining plane configurations in CAD: Circles, lines, and angles in the plane. SIAM Journal on Discrete Mathematics, 18(2):246–271, 2004. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/37413.

Suen:2010:SRA

[SW10] Stephen Suen and Kevin P. Wagner. Some results for the acceptance urn model. SIAM Journal on Discrete Mathematics, 24(3):876–891, ???? 2010. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Scott:2014:HBD

[SW14] Alex Scott and Elizabeth Wilmer. Hypergraphs of bounded disjointness. SIAM Journal on Discrete Mathematics, 28(1):372–384, ???? 2014. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Shparlinski:2021:ESS

[SW21] Igor E. Shparlinski and Qiang Wang. Exponential sums with sparse polynomials over finite fields. SIAM Journal on Discrete Mathematics, 35 (2):976–987, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Sheu:2023:PCF

[SW23a] Wen-Horng Sheu and Biing-Feng Wang. Parameterized complexity for finding a perfect phylogeny from mixed tumor samples. SIAM Journal on Discrete Mathematics, 37(3):2049–2071, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1449269.

Shokrieh:2023:TSS

[SW23b] Farbod Shokrieh and Cameron Wright. On torsor structures on spanning trees. SIAM Journal on Discrete Mathematics, 37(3):2126–2147, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL

https://epubs.siam.org/
doi/10.1137/22M149154X.

Swartz:2005:LBV

[Swa05] E. Swartz. Lower bounds for h-vectors of k-CM, independence, and broken circuit complexes. SIAM Journal on Discrete Mathematics, 18(3):647–661, 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/41135.

Sun:2010:DCS

[SWKP10] Hung-Min Sun, Huaxiong Wang, Bying-He Ku, and Josef Pieprzyk. Decomposition construction for secret sharing schemes with graph access structures in polynomial time. SIAM Journal on Discrete Mathematics, 24(2): 617–638, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Shetler:2012:SMR

[SWR12] Daniel S. Shetler, Michael A. Wurtz, and Stanislaw P. Radziszowski. On some multicolor Ramsey numbers involving K_{3} + e and K_{4} − e. SIAM Journal on Discrete Mathematics, 26(3): 1256–1264, ???? 2012. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Shen:2013:AAF

[SX13] Hong Shen and Shihong Xu. Approximation algorithms for fault tolerant facility allocation. SIAM Journal on Discrete Mathematics, 27(3): 1584–1609, ???? 2013. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Snir:2011:LTA

[SY11] Sagi Snir and Raphael Yuster. A linear time approximation scheme for maximum quartet consistency on sparse sampled inputs. SIAM Journal on Discrete Mathematics, 25(4): 1722–1736, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sidma/resource/
                                        1/sjdmec/v25/i4/p1722_s1.

Shiraga:2018:DRW

[SYKY18] Takeharu Shiraga, Yukiko Yamauchi, Shuji Kijima, and Masafumi Yamashita. Deterministic random walks for rapidly mixing chains. SIAM Journal on Discrete Mathematics, 32(3):2180–2193, ???? 2018. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Sole:1994:CAS

[SZ94] Patrick Solé and Thomas Zaslavsky. A coding approach to signed graphs. SIAM Journal on Discrete Mathematics, 7 (4):544–553, November 1994. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Stein:2013:FLC

[SZ13] Maya Stein and José Zamora.

Forcing large complete (topological) minors in infinite graphs. SIAM Journal on Discrete Mathematics, 27(2): 697–707, ???? 2013. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Stanley:2015:CCA

[SZ15] Richard P. Stanley and Fabrizio Zanello. The Catalan case of Armstrong’s Conjecture on simultaneous core partitions. SIAM Journal on Discrete Mathematics, 29(1): 658–666, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Szabo:2006:FFA

[Sza06]  Sándor Szabó.  Factor-

ing finite Abelian groups by subsets with maximal span. SIAM Journal on Discrete Mathematics, 20(4): 920–931, January 2006. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Szabo:2008:MPJ

[Sza08] Jácint Szabó. Matroid parity and jump systems: a solution to a conjecture of Recski. SIAM Journal on Discrete Mathematics, 22(3): 854–860, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Tang:1993:RBC

[TA93] K. Wendy Tang and Bruce W. Arden. Representations of Borel Cayley graphs. SIAM Journal on Discrete Mathematics, 6(4):655–676, November 1993. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Takacs:1990:NDF

[Tak90] Lajos Takács. On the number of distinct forests. SIAM Journal on Discrete Mathematics, 3(4):574–581, November 1990. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Takata:2008:WCA

[Tak08]  Ken Takata.  A worst-

case analysis of the sequential method to list the minimal hitting sets of a hypergraph. SIAM Journal on Discrete Mathematics, 21(4): 936–946, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Takazawa:2014:OMF

[Tak14] Kenjiro Takazawa. Optimal matching forests and valuated delta-matroids. SIAM Journal on Discrete Mathematics, 28(1):445–467, ???? 2014. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Takazawa:2022:EFB

[Tak22]  Kenjiro Takazawa.  Ex-

cluded t-factors in bipartite graphs: Unified framework for nonbipartite matchings, restricted 2-matchings, and matroids. SIAM Journal on Discrete Mathematics, 36(1):702–727, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/18M1176737.

Tamir:1988:ICB

[Tam88] Arie Tamir. Improved complexity bounds for center location problems on networks by using dynamic data structures. SIAM Journal on Discrete Mathematics, 1(3): 377–396, August 1988. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Tamir:1991:OFL

[Tam91] Arie Tamir. Obnoxious facility location on graphs. SIAM Journal on Discrete Mathematics, 4(4):550–567, November 1991. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Tan:2010:ADD

[Tan10]  Ngo Dac Tan.  3-arc-

dominated digraphs. SIAM Journal on Discrete Mathematics, 24(3):1153–1161, ???? 2010. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Tan:2021:TBT

[Tan21] Ngo Dac Tan. Tournaments and bipartite tournaments without vertex disjoint cycles of different lengths. SIAM Journal on Discrete Mathematics, 35(1):485–494, ???? 2021. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Tardos:1988:PBC

[Tar88] Gábor Tardos. Polynomial bound for a chip firing game on graphs. SIAM Journal on Discrete Mathematics, 1(3): 397–398, August 1988. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Tassiulas:1997:WCL

[Tas97] L. Tassiulas. Worst case length of nearest neighbor tours for the Euclidean traveling salesman problem. SIAM Journal on Discrete Mathematics, 10(2):171–179, May 1997. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Tenner:2009:OLE

[Ten09] Bridget Eileen Tenner. Optimizing linear extensions. SIAM Journal on Discrete Mathematics, 23(3):1450– 1454, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Tuan:1990:RRS

[TH90] Tai-Ching Tuan and S. Louis Hakimi. River routing with a small number of jogs. SIAM Journal on Discrete Mathematics, 3(4):585–597, November 1990. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Taylor:2011:LFN

[TH11] Joshua A. Taylor and Franz S.

Hover. Laplacians for flow networks. SIAM Journal on Discrete Mathematics, 25(3): 1349–1364, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v25/i3/p1349_s1.

Thatte:2008:CPC

[Tha08] Bhalchandra D. Thatte.

Combinatorics of pedigrees I: Counterexamples to a reconstruction question. SIAM Journal on Discrete Mathematics, 22(3):961–970, ???? 2008. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Timar:2008:SH

[Tim08] Ádám Timár. Split hypergraphs. SIAM Journal on Discrete Mathematics, 22(3): 1155–1163, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Terada:2018:SLR

[TKA18] I. Terada, R. C. King, and O. Azenhas. The symmetry of Littlewood–Richardson coefficients: a new hive model involutory bijection. SIAM Journal on Discrete Mathematics, 32(4):2850–2899, ???? 2018. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Tang:2019:MMM

[TKMM19] Deng Tang, Selçuk Kavut, Bimal Mandal, and Subhamoy Maitra. Modifying Maiorana– McFarland type bent functions for good cryptographic properties and efficient implementation. SIAM Journal on Discrete Mathematics, 33(1): 238–256, ???? 2019. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Todd:1989:TGC

[Tod89] Philip Todd. A k-tree generalization that characterizes consistency of dimensioned engineering drawings. SIAM Journal on Discrete Mathematics, 2(2):255–261, May 1989. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Todd:2014:IKK

[Tod14]  Michael J. Todd.  An

improved Kalai–Kleitman bound for the diameter of a polyhedron. SIAM Journal on Discrete Mathematics, 28(4):1944–1947, ???? 2014. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Torocsik:1993:SB

[Tör93] Jenö Töröcsik. n+1 segments beat n. SIAM Journal on Discrete Mathematics, 6(3): 491–500, August 1993. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Toth:2008:AAS

[Tót08] Csaba D. Tóth. Axis-aligned subdivisions with low stabbing numbers. SIAM Journal on Discrete Mathematics, 22(3):1187–1204, ???? 2008. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Toth:2010:UCI

[Tót10] Ágnes Tóth. The ultimate categorical independence ratio of complete multipartite graphs. SIAM Journal on Discrete Mathematics, 23(4): 1900–1904, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Tovey:1990:SAR

[Tov90] Craig A. Tovey. A simplified anomaly and reduction for precedence constrained multiprocessor scheduling. SIAM Journal on Discrete Mathematics, 3(4):582–584, November 1990. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Telle:1997:AVP

[TP97] Jan Arne Telle and Andrzej Proskurowski. Algorithms for vertex partitioning problems on partial ktrees. SIAM Journal on Discrete Mathematics, 10(4): 529–550, November 1997. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Tian:2009:ADD

[TQ09] Tian Tian and Wen-Feng Qi. Autocorrelation and distinctness of decimations of ℓsequences. SIAM Journal on Discrete Mathematics, 23(2): 805–821, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Tan:2013:FSN

[TQLT13] Yin Tan, Longjiang Qu, San Ling, and Chik How Tan. On the Fourier spectra of new APN functions. SIAM Journal on Discrete Mathematics, 27(2):791–801, ???? 2013. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Trevisan:2004:LVG

[Tre04] Luca Trevisan. On local versus global satisfiability. SIAM Journal on Discrete Mathematics, 17(4):541–547, 2004. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/32652.

Tropp:2015:IFP

[Tro15] Joel A. Tropp. Integer factorization of a positive-definite matrix. SIAM Journal on Discrete Mathematics, 29(4): 1783–1791, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Tapia-Recillas:2003:LQC

[TRV03] H. Tapia-Recillas and G. Vega.

On Z_{2} k-linear and quaternary codes. SIAM Journal on Discrete Mathematics, 17(1): 103–113, January 2003. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Tsai:1996:LBR

[Tsa96] Shi-Chun Tsai. Lower bounds on representing Boolean functions as polynomials in Z_{m}. SIAM Journal on Discrete Mathematics, 9(1):55– 62, February 1996. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

To:2004:MCL

[TSN04] Vu Dong Tô and Reihaneh Safavi-Naini. On the maximal codes of length 3 with the 2-identifiable parent property. SIAM Journal on Discrete Mathematics, 17(4):548–570, 2004. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/40042.

Tsuchiya:2023:CSM

[Tsu23] Akiyoshi Tsuchiya. Cayley sums and Minkowski sums of lattice polytopes. SIAM Journal on Discrete Mathematics, 37(2):1348–1357, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL

https://epubs.siam.org/
doi/10.1137/22M1507991.

Tipnis:1989:NPP

[TT89] S. K. Tipnis and L. E. Trotter, Jr. Node-packing problems with integer rounding properties. SIAM Journal on Discrete Mathematics, 2(3): 407–412, August 1989. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Tamassia:1991:RGC

[TT91] Roberto Tamassia and Ioannis G. Tollis. Representations of graphs on a cylinder. SIAM Journal on Discrete Mathematics, 4(1):139–149, February 1991. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Todd:1993:NTS

[TT93] Michael J. Todd and Levent Tunçel. A new triangularization for simplicial algorithms. SIAM Journal on Discrete Mathematics, 6(1): 167–180, February 1993. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Tardos:2007:CST

[TT07] Gábor Tardos and Géza Tóth. Crossing stars in topological graphs. SIAM Journal on Discrete Mathematics, 21 (3):737–749, ???? 2007. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Tait:2016:ISP

[TT16] Michael Tait and Craig Timmons. Independent sets in polarity graphs. SIAM Journal on Discrete Mathematics, 30(4):2115–2129, ???? 2016. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Tuza:2008:HDG

[Tuz08] Zsolt Tuza. Hereditary domination in graphs: Characterization with forbidden induced subgraphs. SIAM Journal on Discrete Mathematics, 22(3): 849–853, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Trachtenberg:2003:FRT

[TV03] Ari Trachtenberg and Alexander Vardy. Full-rank tilings of F_{!2}^{8} do not exist. SIAM Journal on Discrete Mathematics, 16(3):390–392, 2003. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/41021.

Terlep:2012:GGK

[TW12] T. Arthur Terlep and Jason Williford. Graphs from generalized Kac–Moody algebras. SIAM Journal on Discrete Mathematics, 26(3): 1112–1120, ???? 2012. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Tardif:2019:HCS

[TW19] Claude Tardif and Marcin Wrochna. Hedetniemi’s conjecture and strongly multiplicative graphs. SIAM Journal on Discrete Mathematics, 33(4):2218–2250, ???? 2019. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Tait:2021:IBS

[TW21] Michael Tait and Robert Won. Improved bounds on sizes of generalized caps in AG(n, q). SIAM Journal on Discrete Mathematics, 35(1): 521–531, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Tillich:1997:OCC

[TZ97] Jean-Pierre Tillich and Gilles Zémor. Optimal cycle codes constructed from Ramanujan graphs. SIAM Journal on Discrete Mathematics, 10(3): 447–459, August 1997. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Thapper:2015:NCT

[TZ15] Johan Thapper and Stanislav Zivný. Necessary conditions for tractability of valued CSPs. SIAM Journal on Discrete Mathematics, 29(4): 2361–2384, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Tan:2019:IPV

[TZ19] Zihan Tan and Liwei Zeng. On the inequalities of projected volumes and the constructible region. SIAM Journal on Discrete Mathematics, 33(2):694–711, ???? 2019. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). See erratum [TZ20].

Tan:2020:EIP

[TZ20] Zihan Tan and Liwei Zeng. Erratum for “On the Inequalities of Projected Volumes and the Constructible Region”. SIAM Journal on Discrete Mathematics, 34(1): 1000, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). See [TZ19].

Tzanaki:2008:FGC

[Tza08] Eleni Tzanaki. Faces of generalized cluster complexes and noncrossing partitions. SIAM Journal on Discrete Mathematics, 22(1):15–30, ???? 2008. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Ueno:1993:MFT

[UBHS93] S. Ueno, A. Bagchi, S. L. Hakimi, and E. F. Schmeichel. On minimum faulttolerant networks. SIAM Journal on Discrete Mathematics, 6(4):565–574, November 1993. CODEN SJDMEC.

ISSN 0895-4801 (print), 1095-7146 (electronic).

Ullrich:2014:SWF

[Ull14] Mario Ullrich. Swendsen– Wang is faster than singlebond dynamics. SIAM Journal on Discrete Mathematics, 28(1):37–48, ???? 2014. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Umanna:2015:DTC

[UV15] Valérie Gauthier Umanña and Mauricio Velasco. Dual toric codes and polytopes of degree one. SIAM Journal on Discrete Mathematics, 29 (1):683–692, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Urbaniak:1997:VBA

[UWZ97] Regina Urbaniak, Robert Weismantel, and Günter M. Ziegler. A variant of the Buchberger algorithm for integer programming. SIAM Journal on Discrete Mathematics, 10(1):96–108, February 1997. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

vanderVeen:1994:NCP

[van94] Jack A. A. van der Veen.

A new class of pyramidally solvable symmetric traveling salesman problems. SIAM Journal on Discrete Mathematics, 7(4):585–592, November 1994. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Vavasis:1989:GEP

[Vav89] Stephen A. Vavasis. Gaussian elimination with pivoting is P-complete. SIAM Journal on Discrete Mathematics, 2(3):413–423, August 1989. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Vazirani:2012:RCP

[Vaz12] Vijay V. Vazirani. Rational convex programs and efficient algorithms for 2-player Nash and nonsymmetric bargaining games. SIAM Journal on Discrete Mathematics, 26 (3):896–918, ???? 2012. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Vazirani:2013:NCU

[Vaz13] Vijay V. Vazirani. Nonseparable, concave utilities are easy — in a perfect price discrimination market model. SIAM Journal on Discrete Mathematics, 27(1): 266–273, ???? 2013. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

vanBevern:2015:NBV

[vBBC^{+}15] René van Bevern, Robert Bredereck, Jiehua Chen, Vincent Froese, Rolf Niedermeier, and Gerhard J. Woeginger. Network-based vertex dissolution. SIAM Journal on Discrete Mathematics, 29(2): 888–914, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

vanBatenburg:2017:CJR

[vBEM17] Wouter Cames van Batenburg, Louis Esperet, and Tobias Müller. Coloring Jordan regions and curves. SIAM Journal on Discrete Mathematics, 31(3):1670–1684, ???? 2017. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

vanBatenburg:2020:EPB

[vBJU20] Wouter Cames van Batenburg, Gwenaël Joret, and Arthur Ulmer. Erdős–Pósa from ball packing. SIAM Journal on Discrete Mathematics, 34(3):1609–1619, ???? 2020. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

vanDalen:2011:BLR

[vD11] Birgit van Dalen. Boundary length of reconstructions in discrete tomography. SIAM Journal on Discrete Mathematics, 25(2): 645–659, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v25/i2/p645_s1.

Vegh:2011:AUN

[Vég11] László A. Végh. Augmenting undirected node-connectivity by one. SIAM Journal on Discrete Mathematics, 25(2): 695–718, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sidma/resource/
                                        1/sjdmec/v25/i2/p695_s1.

vanIersel:2016:HNT

[vIKL^{+}16] Leo van Iersel, Steven Kelk, Nela Lekić, Chris Whidden, and Norbert Zeh. Hybridization number on three rooted binary trees is EPT. SIAM Journal on Discrete Mathematics, 30(3):1607–1631, ???? 2016. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

vanIersel:2014:AAN

[vIKLS14] Leo van Iersel, Steven Kelk, Nela Lekić, and Leen Stougie. Approximation algorithms for nonbinary agreement forests. SIAM Journal on Discrete Mathematics, 28(1): 49–66, ???? 2014. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

vanIersel:2018:LRP

[vIM18] Leo van Iersel and Vincent Moulton. Leaf-reconstructibility of phylogenetic networks. SIAM Journal on Discrete Mathematics, 32(3):2047– 2066, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Vince:1993:RT

[Vin93] Andrew Vince. Replicating tessellations. SIAM Journal on Discrete Mathematics, 6(3):501–521, August 1993. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Vince:2007:DLW

[Vin07] A. Vince. Discrete lines and wandering paths. SIAM Journal on Discrete Mathematics, 21(3):647–661, ???? 2007. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Vinh:2011:EFD

[Vin11] Le Anh Vinh. The Erdős– Falconer distance problem on the unit sphere in vector spaces over finite fields. SIAM Journal on Discrete Mathematics, 25(2):681–684, ???? 2011. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v25/i2/p681_s1.

Vinh:2012:PMR

[Vin12] Le Anh Vinh. On the permanents of matrices with restricted entries over finite fields. SIAM Journal on Discrete Mathematics, 26(3):

997–1007, ???? 2012. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Vinh:2013:FVE

[Vin13] Le Anh Vinh. On fourvariable expanders in finite fields. SIAM Journal on Discrete Mathematics, 27(4): 2038–2048, ???? 2013. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Vodicka:2021:NKP

[Vod21] Martin Vodicka. Normality of the Kimura 3-parameter model. SIAM Journal on Discrete Mathematics, 35(3): 1536–1547, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Voigt:2007:PEC

[Voi07]  Margit Voigt.  Precolor-

ing extension for 2-connected graphs. SIAM Journal on Discrete Mathematics, 21(1): 258–263, ???? 2007. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Vsemirnov:2005:APS

[Vse05] Maxim Vsemirnov. Automorphisms of projective spaces and min-wise independent sets of permutations. SIAM Journal on Discrete Mathematics, 18(3):592–607, 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/41818.

VanderGeer:1994:GHW

[VV94] G. Van der Geer and M. Van der Vlugt. Generalized Hamming weights of Melas codes and dual Melas codes. SIAM Journal on Discrete Mathematics, 7(4):554–559, November 1994. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Vera:2015:IBP

[VVY15] Juan C. Vera, Eric Vigoda, and Linji Yang. Improved bounds on the phase transition for the hard-core model in 2 dimensions. SIAM Journal on Discrete Mathematics, 29(4):1895–1915, ???? 2015. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Vandenbussche:2009:MEH

[VW09] Jennifer Vandenbussche and Douglas B. West. Matching extendability in hypercubes. SIAM Journal on Discrete Mathematics, 23(3): 1539–1547, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

vanHoesel:1994:PCE

[vWW94] C. P. M. van Hoesel, A. P. M.

Wagelmans, and L. A.

Wolsey. Polyhedral characterization of the economic lotsizing problem with start-up costs. SIAM Journal on Discrete Mathematics, 7(1):141– 151, February 1994. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Vandenbussche:2009:PT

[VWY09] Jennifer Vandenbussche, Douglas B. West, and Gexin Yu. On the pagenumber of k-trees. SIAM Journal on Discrete Mathematics, 23(3): 1455–1464, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Vygen:2016:RTT

[Vyg16] Jens Vygen. Reassembling trees for the traveling salesman. SIAM Journal on Discrete Mathematics, 30(2): 875–894, ???? 2016. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Vollmar:1993:EST

[VZ93] Kenneth R. Vollmar and Yanjun Zhang. On the enumeration of Steiner-tree topologies for the points on a circle. SIAM Journal on Discrete Mathematics, 6(2): 329–333, May 1993. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

vonzurGathen:2013:CRP

[vzGVZ13] Joachim von zur Gathen, Alfredo Viola, and Konstantin Ziegler. Counting reducible, powerful, and relatively irreducible multivariate polynomials over finite fields. SIAM Journal on Discrete Mathematics, 27(2):855–891, ???? 2013. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Wagner:1996:MS

[Wag96] David G. Wagner. Multipartition series. SIAM Journal on Discrete Mathematics, 9 (4):529–544, November 1996. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Wagner:2007:CGT

[Wag07] Stephan G. Wagner. Correlation of graph-theoretical indices. SIAM Journal on Discrete Mathematics, 21(1): 33–46, ???? 2007. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Wagner:2018:CGM

[Wag18] Donald K. Wagner. A characterization of graphic matroids based on circuit orderings. SIAM Journal on Discrete Mathematics, 32(2): 1139–1144, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Wagner:2021:NCB

[Wag21] Donald K. Wagner. Nonseparating cocircuits in binary matroids. SIAM Journal on Discrete Mathematics, 35(4): 3030–3049, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Walter:2019:CDM

[Wal19]  Matthias Walter.  Com-

plete description of matching polytopes with one linearized quadratic term for bipartite graphs. SIAM Journal on Discrete Mathematics, 33(2): 1061–1094, ???? 2019. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Wang:2002:FCT

[Wan02a]  Biing-Feng Wang.  Find-

ing a 2-core of a tree in linear time. SIAM Journal on Discrete Mathematics, 15(2):193–210, 2002. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/37424.

Wang:2002:CBG

[Wan02b]  Hong Wang.  On cover-

ing a bipartite graph with cycles. SIAM Journal on Discrete Mathematics, 15 (1):86–96, 2002. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/31048.

Wang:2008:SCG

[Wan08] Susan S. Wang. Structure and coloring of graphs with only small odd cycles. SIAM Journal on Discrete Mathematics, 22(3):1040–1072, ???? 2008. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Wang:2023:BSL

[Wan23] Yan Wang. Balanced subdivisions of a large clique in graphs with high average degree. SIAM Journal on Discrete Mathematics, 37 (2):1262–1274, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1511266.

Wayne:2001:NPF

[Way01] Kevin D. Wayne. A new property and a faster algorithm for baseball elimination. SIAM Journal on Discrete Mathematics, 14(2):223–229, 2001. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/34884.

Wang:1990:BSM

[WB90] Muzhong Wang and Ian F.

Blake. Bit serial multiplication in finite fields. SIAM Journal on Discrete Mathematics, 3(1):140–148, February 1990. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Wu:2012:CCC

[WC12] Hsin-Lung Wu and Jen-Chun Chang. Constructing constant composition codes via distance-increasing mappings. SIAM Journal on Discrete Mathematics, 26(1): 375–383, ???? 2012. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Wang:2015:VFI

[WCLZ15] Yi Wang, Jian Cheng, Rong Luo, and Cun-Quan Zhang. Vector flows and integer flows. SIAM Journal on Discrete Mathematics, 29(4): 2166–2178, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Weller-Davies:2021:CPB

[WDSH21] Oliver Weller-Davies, Mike Steel, and Jotun Hein. Combinatorics of polymer-based models of early metabolism. SIAM Journal on Discrete Mathematics, 35(4):2304– 2316, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Webb:2008:VSG

[Web08] K. P. Webb. The vector space over GF (2) generated by the common bases of two binary matroids. SIAM Journal on Discrete Mathematics, 22(3): 1205–1214, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Weissauer:2019:BNG

[Wei19] Daniel Weißauer. On the block number of graphs. SIAM Journal on Discrete Mathematics, 33(1):346–357, ???? 2019. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Weng:1997:LST

[Wen97] J. F. Weng. Linear Steiner trees for infinite spirals. SIAM Journal on Discrete Mathematics, 10(3):388–398, August 1997. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Whittlesey:1995:NRG

[WGM95] Marshall A. Whittlesey, John P. Georges, and David W. Mauro. On the λ-number of Q_{n} and related graphs. SIAM Journal on Discrete Mathematics, 8(4):499–506, November 1995. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Will:2004:PM

[WH04] Todd G. Will and Heather Hulett. Parsimonious multigraphs. SIAM Journal on Discrete Mathematics, 18 (2):241–245, 2004. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/43247.

Wang:2015:CAV

[WH15] Weifan Wang and Danjun Huang. A characterization on the adjacent vertex distinguishing index of planar graphs with large maximum degree. SIAM Journal on Discrete Mathematics, 29(4): 2412–2431, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Whiteley:1988:UMR

[Whi88] Walter Whiteley. The union of matroids and the rigidity of frameworks. SIAM Journal on Discrete Mathematics, 1(2):237–255, May 1988. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Wang:2014:PGE

[WHW14] Yiqiao Wang, Xiaoxue Hu, and Weifan Wang. Planar graphs with Δ ≥ 9 are entirely (Δ+2)-colorable. SIAM Journal on Discrete Mathematics, 28(4):1892–1905, ????

2014. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Will:1999:SDB

[Wil99] Todd G. Will. Switching distance between graphs with the same degrees. SIAM Journal on Discrete Mathematics, 12(3):298–306, August 1999. CODEN SJD-MEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/33115.

Wild:2005:ANB

[Wil05] Marcel Wild. The asymptotic number of binary codes and binary matroids. SIAM Journal on Discrete Mathematics, 19(3):691–699, 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/44553.

Wildon:2016:CPP

[Wil16] Mark Wildon. A combinatorial proof of a plethystic Murnaghan–Nakayama rule. SIAM Journal on Discrete Mathematics, 30(3): 1526–1533, ???? 2016. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Winkler:1988:CMR

[Win88] Peter Winkler. The complexity of metric realization. SIAM Journal on Discrete Mathematics, 1(4):552– 559, November 1988. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Windisch:2016:RMM

[Win16] Tobias Windisch. Rapid mixing and Markov bases. SIAM Journal on Discrete Mathematics, 30(4):2130–2145, ???? 2016. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Watanabe:2012:AAP

[WIS12] Saori Watanabe, Kota Ishii, and Masanori Sawa. A qanalogue of the addressing problem of graphs by Graham and Pollak. SIAM Journal on Discrete Mathematics, 26 (2):527–536, ???? 2012. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Wang:2002:CEC

[WL02] Wei-Fan Wang and Ko-Wei Lih. Choosability and edge choosability of planar graphs without intersecting triangles. SIAM Journal on Discrete Mathematics, 15(4):538–545, 2002. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/37625.

Wang:2003:LPG

[WL03] Wei-Fan Wang and Ko-Wei Lih. Labeling planar graphs with conditions on girth and distance two. SIAM Journal on Discrete Mathematics, 17(2):264–275, 2003. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/39044.

Wood:2010:TCA

[WL10] David R. Wood and Svante Linusson. Thomassen’s choosability argument revisited. SIAM Journal on Discrete Mathematics, 24(4): 1632–1637, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v24/i4/p1632_s1.

Wolfe:2009:ERL

[WLD09] A. J. Wolfe, A. C. H. Ling, and J. H. Dinitz. The existence of N_{2} resolvable Latin squares. SIAM Journal on Discrete Mathematics, 23(3): 1217–1237, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Woeginger:1993:ISB

[Woe93] Gerhard Woeginger. Improved space for boundedspace, on-line bin-packing. SIAM Journal on Discrete Mathematics, 6(4):575–581, November 1993. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Wollan:2010:BHC

[Wol10] Paul Wollan. Bridges in highly connected graphs. SIAM Journal on Discrete Mathematics, 24(4):1731– 1741, ???? 2010. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sidma/resource/
                                        1/sjdmec/v24/i4/p1731_s1.

Worley:1988:VR

[Wor88] R. T. Worley. The Voronoĭ region of E_{7}^{∗}. SIAM Journal on Discrete Mathematics, 1(1): 134–141, February 1988. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Wrochna:2020:HRH

[Wro20] Marcin Wrochna. Homomorphism reconfiguration via homotopy. SIAM Journal on Discrete Mathematics, 34(1): 328–350, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Wang:1996:GCN

[WS96] Terry Min Yih Wang and Carla D. Savage. A Gray code for necklaces of fixed density. SIAM Journal on Discrete Mathematics, 9(4):654– 673, November 1996. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Weinard:2006:GSC

[WS06] Maik Weinard and Georg Schnitger. On the greedy superstring conjecture. SIAM Journal on Discrete Mathematics, 20(2):502–522, January 2006. CODEN SJD-MEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Wendemuth:2012:MDO

[WS12] Andreas Wendemuth and Italo Simonelli. Majority decisions in overlapping committees and asymptotic size of dichotomies. SIAM Journal on Discrete Mathematics, 26(4):1537–1549, ???? 2012. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Wang:2017:GIR

[WS17a] Lele Wang and Ofer Shayevitz.

Graph information ratio. SIAM Journal on Discrete Mathematics, 31(4):2703– 2734, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Wang:2017:SNF

[WS17b] Yinghui Wang and Richard P.

Stanley. The Smith normal form distribution of a random integer matrix. SIAM Journal on Discrete Mathematics, 31(3):2247–2268, ???? 2017. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Weinberger:2019:SPB

[WS19] Nir Weinberger and Ofer Shayevitz. Self-predicting Boolean functions. SIAM Journal on Discrete Mathematics, 33(2):665–693, ???? 2019. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Wu:2009:LAG

[Wu09] Jian-Liang Wu. The linear arboricity of graphs on surfaces of negative Euler characteristic. SIAM Journal on Discrete Mathematics, 23(1): 54–58, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Wang:1991:LDR

[WW91] Stanley Xi Wang and Edward C. Waymire. A large deviation rate and central limit theorem for Horton ratios. SIAM Journal on Discrete Mathematics, 4(4):575– 588, November 1991. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Wilson:2018:AFC

[WW18] Tim E. Wilson and David R. Wood. Anagram-free colorings of graph subdivisions. SIAM Journal on Discrete Mathematics, 32(3): 2346–2360, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Wanless:2022:GFH

[WW22] Ian M. Wanless and David R.

Wood. A general framework for hypergraph coloring. SIAM Journal on Discrete Mathematics, 36 (3):1663–1677, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1421015.

Wang:2011:EBN

[WWKY11] Shaofan Wang, Renhong Wang, Dehui Kong, and Baocai Yin. Estimate of the Bézout number for linear piecewise algebraic curves over arbitrary triangulations. SIAM Journal on Discrete Mathematics, 25(2): 561–582, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v25/i2/p561_s1.

Wang:2023:DCD

[WWY23] Hong Wang, Yun Wang, and Jin Yan. Disjoint cycles in a digraph with partial degree. SIAM Journal on Discrete Mathematics, 37(1):221–232, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1460594.

Wang:2013:ICP

[WX13] Yingqian Wang and Lingji Xu. Improper choosability of planar graphs without 4cycles. SIAM Journal on Discrete Mathematics, 27(4): 2029–2037, ???? 2013. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Weimann:2010:CGP

[WY10] Oren Weimann and Raphael Yuster. Computing the girth of a planar graph in O(n log n) time. SIAM Journal on Discrete Mathematics, 24(2):609–616, ???? 2010. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Wu:2020:MJS

[WY20] Yezhou Wu and Dong Ye. Minimum T -joins and signedcircuit covering. SIAM Journal on Discrete Mathematics, 34(2):1192–1204, ???? 2020. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Wang:2023:EPM

[WY23] Jian Wang and Jie You. Extremal problem for matchings and rainbow matchings on direct products. SIAM Journal on Discrete Mathematics, 37(3):2030–2048, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1457990.

Wu:2014:NZF

[WYZZ14] Yezhou Wu, Dong Ye, Wenan Zang, and Cun-Quan Zhang. Nowhere-zero 3-flows in signed graphs. SIAM Journal on Discrete Mathematics, 28(3):1628–1637, ???? 2014. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Wang:2008:NMP

[WZ08] Jun Wang and Huajun Zhang. Normalized matching property of restricted subspace lattices. SIAM Journal on Discrete Mathematics, 22 (1):248–255, ???? 2008. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Wang:2018:IFS

[WZ18] Jun Wang and Huajun Zhang. Intersecting families in symmetric unions of direct products of set families. SIAM Journal on Discrete Mathematics, 32(1): 372–381, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Wu:2009:CAC

[WZZ09] Yezhou Wu, Wenan Zang, and Cun-Quan Zhang. A characterization of almost CIS graphs. SIAM Journal on Discrete Mathematics, 23 (2):749–753, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Wang:2018:SNC

[WZZ18] Yanping Wang, Weiguo Zhang, and Zhengbang Zha. Six new classes of permutation trinomials over F_{2^{n}}. SIAM Journal on Discrete Mathematics, 32(3):1946– 1961, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Xu:2020:EFD

[XG20] Zixiang Xu and Gennian Ge. Erdős–Falconer distance problem under Hamming metric in vector spaces over finite fields. SIAM Journal on Discrete Mathematics, 34(4):2208–2220, ???? 2020. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Xu:2015:MAD

[XGG15] Shou-Jun Xu, Rob Gysel, and Dan Gusfield. Minimum average distance clique trees. SIAM Journal on Discrete Mathematics, 29(3): 1706–1734, ???? 2015. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Xu:2006:AMP

[XQ06] Hong Xu and Wen-Feng Qi. Autocorrelations of maximum period FCSR sequences. SIAM Journal on Discrete Mathematics, 20(3): 568–577, January 2006. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Xu:2011:MCL

[XSR11] Xiaodong Xu, Zehui Shao, and Stanisław P. Radziszowski. More constructive lower bounds on classical Ramsey numbers. SIAM Journal on Discrete Mathematics, 25(1):394–400, ???? 2011. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:

//epubs.siam.org/sidma/
resource/1/sjdmec/v25/i1/
p394_s1.

Xu:2009:CPG

[Xu09] Baogang Xu. On (3, 1)^{∗}- coloring of plane graphs. SIAM Journal on Discrete Mathematics, 23(1):205–220, ???? 2009. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Xi:2015:CIP

[XY15] Jing Xi and Ruriko Yoshida.

The characteristic imset polytope of Bayesian networks with ordered nodes. SIAM Journal on Discrete Mathematics, 29(2):697–715, ???? 2015. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Xu:2022:SFC

[XZ22] Rongxing Xu and Xuding Zhu. The strong fractional choice number and the strong fractional paint number of graphs. SIAM Journal on Discrete Mathematics, 36 (4):2585–2601, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1434556.

Yaman:2005:PAU

[Yam05] Hande Yaman. Polyhedral analysis for the uncapacitated hub location problem with modular arc capacities. SIAM Journal on Discrete Mathematics, 19(2):501–522, 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/43915.

Yaman:2007:IKC

[Yam07] Hande Yaman. The integer knapsack cover polyhedron. SIAM Journal on Discrete Mathematics, 21(3): 551–572, ???? 2007. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Yamaguchi:2016:PPG

[Yam16] Yutaro Yamaguchi. Packing A-paths in group-labelled graphs via linear matroid parity. SIAM Journal on Discrete Mathematics, 30(1): 474–492, ???? 2016. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Yamamoto:2020:LDT

[Yam20] Ken Yamamoto. Large deviation theorem for branches of the random binary tree in the Horton–Strahler analysis. SIAM Journal on Discrete Mathematics, 34(1): 938–949, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Yang:2016:LBP

[YAT16] Jun Yang, Fady Alajaji, and Glen Takahara. Lower bounds on the probability of a finite union of events. SIAM Journal on Discrete Mathematics, 30(3):1437–1452, ???? 2016. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Yu:1997:HMF

[YCH97] B. Yu, J. Cheriyan, and P. E.

Haxell. Hypercubes and multicommodity flows. SIAM Journal on Discrete Mathematics, 10(2):190–200, May 1997. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Yen:1994:NMP

[Yen94] Lily Yen. A note on multiset permutations. SIAM Journal on Discrete Mathematics, 7(1):152–155, February 1994. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Yen:1997:CPS

[Yen97] Lily Yen. A combinatorial proof for Stockhausen’s problem. SIAM Journal on Discrete Mathematics, 10(3): 499–504, August 1997. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Yao:1988:FMG

[YH88] Y.-C. Yao and F. K. Hwang. A fundamental monotonicity in group testing. SIAM Journal on Discrete Mathematics, 1(2):256–259, May 1988. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Yokoi:2019:MCF

[Yok19] Yu Yokoi. Matroidal choice functions. SIAM Journal on Discrete Mathematics, 33(3): 1712–1724, ???? 2019. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Yoshida:2019:MMS

[Yos19] Yuichi Yoshida. Maximizing a monotone submodular function with a bounded curvature under a knapsack constraint. SIAM Journal on Discrete Mathematics, 33(3): 1452–1471, ???? 2019. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Yan:2022:ILD

[YP22] Zilong Yan and Yuejian Peng. An irrational Lagrangian density of a single hypergraph. SIAM Journal on Discrete Mathematics, 36(1):786–822, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1410798.

Ye:2009:RFG

[YQZ09] Dong Ye, Zhongbin Qi, and Heping Zhang. On k-resonant Fullerene graphs. SIAM Journal on Discrete Mathematics, 23(2):1023–1044, ???? 2009. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Yeap:1995:SFG

[YS95] Gary K. H. Yeap and Majid Sarrafzadeh. Sliceable floorplanning by graph dualization. SIAM Journal on Discrete Mathematics, 8(2): 258–280, May 1995. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Yu:2017:NBE

[Yu17] Wei-Hsuan Yu. New bounds for equiangular lines and spherical two-distance sets. SIAM Journal on Discrete Mathematics, 31(2):908–917, ???? 2017. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Yuster:2003:ECU

[Yus03] Raphael Yuster. Equitable coloring of k-uniform hypergraphs. SIAM Journal on Discrete Mathematics, 16(4):524–532, 2003. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/40276.

Yuster:2009:DCA

[Yus09]  Raphael Yuster.  Dis-

joint color-avoiding triangles. SIAM Journal on Discrete Mathematics, 23(1): 195–204, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Yuster:2014:EDC

[Yus14]  Raphael Yuster.  Edge-

disjoint cliques in graphs with high minimum degree. SIAM Journal on Discrete Mathematics, 28(2):893–910, ????

2014. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Yuster:2024:PCG

[Yus24] Raphael Yuster. Packing and covering a given directed graph in a directed graph. SIAM Journal on Discrete Mathematics, 38(1): 43–54, January 2024. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Yang:2021:IPG

[YWWL21] Wanshun Yang, Yiqiao Wang, Weifan Wang, and Ko-Wei Lih. IC-planar graphs are 6choosable. SIAM Journal on Discrete Mathematics, 35(3): 1729–1745, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Yuster:1997:FEC

[YZ97] Raphael Yuster and Uri Zwick. Finding even cycles even faster. SIAM Journal on Discrete Mathematics, 10 (2):209–222, May 1997. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Yang:2017:BOP

[YZ17] Arthur L. B. Yang and Philip B. Zhang. Brenti’s open problem on the realrootedness of q-Eulerian polynomials of type D. SIAM Journal on Discrete Mathematics, 31(2):918–926, ???? 2017. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Zak:2014:PTG

[Zak14]  Andrzej Zak.  On pack-

ing two graphs with bounded sum of sizes and maximum degree. SIAM Journal on Discrete Mathematics, 28(4): 1686–1698, ???? 2014. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Zamfirescu:2021:HG

[Zam21] Carol T. Zamfirescu. K_{2}- hamiltonian graphs: I. SIAM Journal on Discrete Mathematics, 35(3):1706–1728, ???? 2021. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Zamfirescu:2022:RGF

[Zam22] Carol T. Zamfirescu. Regular graphs with few longest cycles. SIAM Journal on Discrete Mathematics, 36 (1):755–776, ???? 2022. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1414048.

Zehavi:2017:MMV

[Zeh17] Meirav Zehavi. Maximum minimal vertex cover parameterized by vertex cover. SIAM Journal on Discrete Mathematics, 31(4):2440–2456, ???? 2017. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Zeng:1990:CSP

[Zen90] Jiang Zeng. Calcul saalschützien des partitions et des dérangements colorés. (French) [Saalschutz calculation of colored partitions and derangements]. SIAM Journal on Discrete Mathematics, 3(1):149–156, February 1990. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Zerbib:2011:ZCV

[Zer11]  Shira Zerbib.  On the

zone complexity of a vertex. SIAM Journal on Discrete Mathematics, 25(2): 719–730, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v25/i2/p719_s1.

Zhang:2021:CD

[ZG21] Tao Zhang and Gennian Ge.

On the codegree density of P G_{m}(q). SIAM Journal on Discrete Mathematics, 35(3): 1548–1556, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Zhu:2023:ECM

[ZGH^{+}23] Xiutao Zhu, Ervin Györi, Zhen He, Zequn Lv, Nika Salia, Casey Tompkins, and Kitti Varga. Edges not covered by monochromatic bipartite graph. SIAM Journal on Discrete Mathematics, 37 (4):2508–2522, October 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Zhang:2009:EGB

[ZGL^{+}09] Rucong Zhang, Gennian Ge, Alan C. H. Ling, Hung-Lin Fu, and Yukiyasu Mutoh. The existence of r × 4 grid-block designs with r = 3, 4. SIAM Journal on Discrete Mathematics, 23(2): 1045–1062, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Zhang:1990:FCI

[Zha90] Cun Quan Zhang. Finding critical independent sets and critical vertex subsets are polynomial problems. SIAM Journal on Discrete Mathematics, 3(3):431–438, August 1990. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Zhang:1993:PSS

[Zha93] Cun Quan Zhang. Parity subgraph, shortest cycle cover, and postman tour. SIAM Journal on Discrete Mathematics, 6(3):428–431, August 1993. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Zhang:1994:SNB

[Zha94] Guo-Hui Zhang. Some new bounds on single-change covering designs. SIAM Journal on Discrete Mathematics, 7(2):166–171, May 1994. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Zhang:1999:OBM

[Zha99]  Louxin Zhang.  Optimal

bounds for matching routing on trees. SIAM Journal on Discrete Mathematics, 12(1):64–77, February 1999. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/32315.

Zhao:2009:BGT

[Zha09] Yi Zhao. Bipartite graph tiling. SIAM Journal on Discrete Mathematics, 23(2): 888–900, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Zhao:2011:BST

[Zha11] Yufei Zhao. The bipartite swapping trick on graph homomorphisms. SIAM Journal on Discrete Mathematics, 25 (2):660–680, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sidma/resource/
                                        1/sjdmec/v25/i2/p660_s1.

Zhang:2023:FCH

[Zha23] Tao Zhang. Fuglede’s conjecture holds in Z_{p} ×Z_{p^{n}}. SIAM Journal on Discrete Mathematics, 37(2):1180–1197, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL

                                        https://epubs.siam.org/
                                        doi/10.1137/22M1493598.

Zheng:2016:MBT

[Zhe16] Hailun Zheng. Minimal balanced triangulations of sphere bundles over the circle. SIAM Journal on Discrete Mathematics, 30(2):1259–1268, ???? 2016. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Zhou:1988:AER

[Zho88]  Bing Zhou.  Almost ev-

ery randomly near-traceable graph has diameter at most two. SIAM Journal on Discrete Mathematics, 1(2): 260–268, May 1988. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Zhou:1992:NOC

[Zho92] Hui Shan Zhou. On the nonmultiplicativity of oriented cycles. SIAM Journal on Discrete Mathematics, 5(2): 207–218, May 1992. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Zhou:1993:CHP

[Zho93] Hui Shan Zhou. Characterization of the homomorphic preimages of certain oriented cycles. SIAM Journal on Discrete Mathematics, 6(1): 87–99, February 1993. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Zhou:2005:LCG

[Zho05] Sanming Zhou. Labelling Cayley graphs on Abelian groups. SIAM Journal on Discrete Mathematics, 19(4): 985–1003, January 2005. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Zhou:2009:CAT

[Zho09] Sanming Zhou. A class of arc-transitive Cayley graphs as models for interconnection networks. SIAM Journal on Discrete Mathematics, 23(2): 694–714, ???? 2009. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Zhu:2018:PIS

[Zhu18] Bao-Xuan Zhu. Positivity of iterated sequences of polynomials. SIAM Journal on Discrete Mathematics, 32(3): 1993–2010, ???? 2018. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Zhu:2021:TPE

[Zhu21] Bao-Xuan Zhu. Total positivity from the exponential Riordan arrays. SIAM Journal on Discrete Mathematics, 35(4): 2971–3003, ???? 2021. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Zitnik:1994:DGS

[Žit94]  Arjana Žitnik.  Drawing

graphs on surfaces. SIAM Journal on Discrete Mathematics, 7(4):593–597, November 1994. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Zdeborova:2011:QPL

[ZK11] Lenka Zdeborová and Florent Krzakala. Quiet planting in the locked constraint satisfaction problems. SIAM Journal on Discrete Mathematics, 25 (2):750–770, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v25/i2/p750_s1.

Zhu:2020:TNB

[ZKNS20] Hui Zhu, Liying Kang, Zhenyu Ni, and Erfang Shan. The Turán number of Berge– K_{4} in 3-uniform hypergraphs. SIAM Journal on Discrete Mathematics, 34(3):1485– 1492, ???? 2020. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Zhang:2011:ESC

[ZL11] Zhao Zhang and Hao Li. Eulerian subgraphs containing given vertices. SIAM Journal on Discrete Mathematics, 25 (2):611–621, ???? 2011. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v25/i2/p611_s1.

Zhang:2008:RGB

[ZLS08] Heping Zhang, Peter Che Bor Lam, and Wai Chee Shiu. Resonance graphs and a binary coding for the 1-factors of benzenoid systems. SIAM Journal on Discrete Mathematics, 22(3):971–984, ???? 2008. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Zhu:2012:LPG

[ZLWC12] Hai-Yang Zhu, Xin-Zhong Lu, Cui-Qi Wang, and Ming Chen. Labeling planar graphs without 4,5-cycles with a condition on distance two. SIAM Journal on Discrete Mathematics, 26 (1):52–64, ???? 2012. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v26/i1/p52_s1.

Zelikovsky:2002:PAA

[ZM02] Alexander Z. Zelikovsky and Ion I. Mandoiu. Practical approximation algorithms for zero- and boundedskew trees. SIAM Journal on Discrete Mathematics, 15(1):97–111, 2002. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/37836.

Zhou:2008:ODS

[ZN08] Xiao Zhou and Takao Nishizeki. Orthogonal drawings of seriesparallel graphs with minimum bends. SIAM Journal on Discrete Mathematics, 22 (4):1570–1604, ???? 2008. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Zhang:2011:SCA

[ZSW11] Yijin Zhang, Kenneth W. Shum, and Wing Shing Wong. Strongly conflictavoiding codes. SIAM Journal on Discrete Mathematics, 25(3):1035–1053, ???? 2011. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sidma/resource/
                                        1/sjdmec/v25/i3/p1035_s1.

Zuckerman:1992:TLB

[Zuc92] David Zuckerman. A technique for lower bounding the cover time. SIAM Journal on Discrete Mathematics, 5(1): 81–87, February 1992. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Zunic:2011:NNT

[Zun11] Jovisa Zunić. Note on the number of two-dimensional threshold functions. SIAM Journal on Discrete Mathematics, 25(3):1266–1268, ???? 2011. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (elec-

tronic). URL http://epubs.
siam.org/sidma/resource/
1/sjdmec/v25/i3/p1266_s1.

Zhao:2023:BIL

[ZWW23] Qinghong Zhao, Bing Wei, and Haidong Wu. Bonds intersecting long paths in kconnected graphs. SIAM Journal on Discrete Mathematics, 37(2):733–747, ???? 2023. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL

https://epubs.siam.org/
doi/10.1137/22M1481105.

Zamfirescu:1992:HPG

[ZZ92] Christina Zamfirescu and Tudor Zamfirescu. Hamiltonian properties of grid graphs. SIAM Journal on Discrete Mathematics, 5(4): 564–570, November 1992. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Zhang:2017:EDC

[ZZBL17] Zan-Bo Zhang, Xiaoyan Zhang, Hajo Broersma, and Dingjun Lou. Extremal and degree conditions for path extendability in digraphs. SIAM Journal on Discrete Mathematics, 31(3): 1990–2014, ???? 2017. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Zhang:2013:DHC

[ZZW13] Zan-Bo Zhang, Xiaoyan Zhang, and Xuelian Wen. Directed Hamilton cycles in digraphs and matching alternating Hamilton cycles in bipartite graphs. SIAM Journal on Discrete Mathematics, 27 (1):274–289, ???? 2013. CO-DEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).